A Complete Bibliography of Publications in Computing

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

12 April 2025 Version 2.56

Title word cross-reference

#13753 [1136]. #14962 [1172]. #15949 [1129].

(2b + 4) [2745]. (6, 3) [2073]. (A − λB)x = 0 [1423]. (nα) [2557]. + [2938]. 0 [1915, 1320, 1092, 2559]. 1 [1778, 2723, 1915, 2206, 1320, 2019, 1790, 2441, 1092, 8, 2559, 7, 711]. 1/||x − y|| [2551]. 10 [371, 601]. 11 [400]. 11 (12) [1492]. 12 [414]. 13 [454]. 14 [528]. 15 [1856, 573]. 16 [574]. 19 [790]. 2 [1907, 2235, 1717, 2434, 4317, 2019, 1652, 1878, 2065, 2123, 2032, 2454, 2394, 2783, 2466, 1872, 2172, 58, 1933, 2183, 2229, 25, 711, 2356, 1104, 2197, 2878, 1942]. 2 − 1/k [2692]. 20 [791]. 24 [942, 874]. 25 [889]. 26 [897]. 27 [1382, 930]. 28 [931]. 29 [932]. 2^{k} [2612]. 2^{p} [2040]. 2k + e [1893]. 2k − 1 [1763].

2n + 1 [1102]. 3 ; [2785, 2838, 1434, 2696, 2851, 4280, 2019, ; 2660, 1972, 4323, 2722, 1865, 2135, 3038, ; 4364, 2884, 82, 2741, 2742, 1104, 2878]. 36 ; [1086]. 37 [1092]. 38 [1093]. 39 [1144, 1257]. ; 3^{rd} [2139]. 4 [4321, 2677, 133]. 40 [1171]. 42 ; [1202]. 45 [1288]. 47 [1320]. 4k + 2 [2423]. 5 ; [1910, 167, 1077, 166, 168]. 51 [1407]. 6 [244]. ; 68 [1247]. 7 [461, 278]. 8 [298, 345]. 9 [346]. ; < [2325]. < χ > [2346]. > [2325]. ^{m} [2440]. _{3} ; [4001]. A [923, 309, 1332]. A(α) [1084, 1077]. ; A + B^{∗}Φ(CX) [843]. α [1865]. α^{1/p} [1479]. ; AX + Y B = C [1424]. AX − XB = C [1590]. ; Ax = b [937]. Ax = λBx [1311]. AXB = C ; [2747]. B [1532, 1533, 1640, 1597, 1726, 1710, ; 1813, 1677]. c [3701]. C^{1} [2117, 2196, 4206, ; 2451, 1636, 1921, 1953, 4219, 2719, 1873]. C^{2} ; [1956, 2608, 2713, 1806, 1976]. C^{k}[a, b] [968]. ; C^{r} [4216]. C_{4} [4319]. H [2740, 2245, 2418]. ; H^{2} [2390]. D [504, 1430, 2375]. ; d^{2}x/dt^{2} = f(t, x, dx/dt) [888]. Δu = bu^{2}

[1661]. E [1299]. E_{s}^{α} [2063]. ℓ [2289, 2411, 2385]. ℓ_{1} [2471]. ϵ [518, 1782]. exp(A) [1821]. F [1352]. F (T) = 0 [162, 162]. f(x) = x [1299]. F_{2} [2562]. G [298]. G^{1} [2718, 2719]. G^{2} [4309]. G_{1} [4321]. H [2421, 2472, 2669, 562, 2580, 2356, 725, 2500, 2503, 2641, 2618, 2499]. H^{−1} [4233]. H^{−1/2} [2230]. H^{1} [2563, 2620]. H^{2} [2552, 2502]. H^{2}(E_{ρ}) [878]. H_{2}^{R} [1485]. xBR^{+} [3331]. hp [2850, 2853, 2861, 2860, 2865, 2908, 2870, 2873, 2859, 2887, 2129]. I(R) [711]. ∫_{0}^{1} exp(As) ds [1821]. k [1893, 3489, 3558, 2008, 2921, 988, 3390, 3659, 3068, 1971, 3735, 1137, 1443, 1581, 3342, 502, 1847, 791, 930, 4285]. k(t, t) ≡ 0 [1282]. Kr [2334]. L [908]. L^{1} [504, 2507]. L^{2}(E_{ρ}) [878]. L_{2}^{R} [1485]. L_{1} [1287, 846, 1060, 951]. L_{2} [1762, 2557, 762]. L_{P} [804, 1669, 457, 1012, 1424]. L_{s} [1252]. λ [921]. LI_{p} [457]. LU [3174, 1516]. M [872, 936, 1196, 777, 1404, 2471, 1461, 988, 1800]. M/G/1 [1132, 1356]. IR [4112]. μ [3949, 2605]. N [2801, 2171, 680, 78, 2765, 2145, 3228, 461, 896, 278, 1687, 920, 2745, 1538, 1845, 18, 877, 1125, 459, 2178, 1714]. n + 1 [459]. N + m [1800]. O(log ∗n) [2600]. O(n) [1557]. O(n^{2}) [2243]. O(n log n) [1745]. O(|V | · |E|) [775]. p [2504, 2628, 2812, 4342, 2580, 2356]. P1 [1987]. p = 2k − m + 1 [988]. p = 3k − m + 2 [988]. p^{m} − 12 + τ(Z) [199]. P_{3}^{mod}/P_{2}^{disc} [2534]. P_{4} [2067]. PRESC^{2} [2967]. ψ [1911]. Q [2144]. Q_{k} − P_{k−1}^{disc} [2395]. R [1334, 1367, 1542, 1582, 4103, 1949, 1261, 1217, 1326]. R^{3} [1045, 2209]. R^{d} [1313, 2183]. S [2733, 1873, 1921, 1953, 4343]. ^{n} \sqrt{2} [2455]. st [1357]. t [2647, 1229]. T^{ω} [471]. u_{xy} = f(x, y, u, u_{x}, u_{y}) [1000]. V [2036]. KDVEM [3068]. W [2222]. ℘(z) [4232]. X [2747, 1220]. y′′ = f(x, y) [1492]. y^{(m)} = f(x, y) [988, 988]. y^{IV} + f · y = g [1194]. y′′ = f(x, y) [1253]. Z [961, 227]. z^{y} [1399]. z_{st} = f(s, t, z) [1175].

-adaptive [2850, 2861, 2860, 2865, 2908, 2873]. -algorithm [518, 1745]. -anonymous [3659]. -Approximation [804, 762, 457]. -Approximations [1252]. -arithmetic [2641]. -ary [2745]. -Band-Simulation [502]. -based [3294, 3481]. -bases [2605]. -Baum-Technik [961]. -Bifurcation [1872]. -boxes [2812]. -center [2723]. -clubs [2921]. -completeness [1461]. -convergence [1532, 1533, 1640, 1597, 1710, 1813, 1677]. -convergent [1726]. -Convex [1921, 1953, 1873]. -convexe [1953]. -Cycle [2036]. -cyclic [4342]. -D [2441, 4364, 1942, 2356]. -D-Thermoelasticity [1790]. -decompositions [1516]. -degree [3068]. -DG [2870]. -Dimensional [2375, 1538, 1581, 2019, 1845, 18, 1847, 1714]. -dimensionaler [18]. -Discrepancy [2557]. -error [1762]. -Estimator [2471]. -FEM [2853, 2504, 2628, 2859, 2887]. -finites [1956]. -Fit [1971]. -Fitting [2507]. -fold [1687]. -Formulae [4285]. -Galerkin [2563]. -gram [2765]. -grams [3228]. -head [502]. -Height-Balanced [1137]. -Histopolation [1953, 1953]. -Interpolation [1921, 1921, 2620]. -intertwined [2801]. -konvexe [1921]. -Kopf-Turing-Maschinen [502]. -kürzesten [791, 930]. -Lattice [2346]. -matchings [1652]. -Matrices [2502, 2500, 2245, 921, 2421, 2472, 2390, 2740, 1196, 2418]. -Matrix [2503, 2499, 2552, 777, 1404, 2618]. -Matrixintervalls [777]. -means [3342]. -Median [2206]. -medoids [3390]. -member [3558]. -methods [1299, 2222]. -Nearest [3735]. -Norm [457]. -Optimality [1782]. -Order [2144, 1949, 1334, 1367, 1542, 1582, 1261]. -ordering [1357]. -Ordnung [1949]. -Packing [2334]. -Partitioning [1972]. -Point [1856, 2745]. -predictions [3489].

-Processor-sharing-System [1356]. -regular [1865]. -Schrittverfahren [988]. -Sequences [2557]. -Shape-Preserving [2117]. -Skeletons [2325]. -solution [1424]. -spline [1220, 1636, 1806]. -Splines [872, 908, 872, 936, 908, 1873]. -Split [4321]. -stabile [1084, 1332]. -Stability [309, 923, 1430]. -stable [1084, 1077, 1332]. -stage [2785]. -Subdivision [4309]. -tape [502]. -ter [78, 988, 877, 1125]. -term [2183]. -th [2145, 988, 877, 791, 930]. -th-order [1125]. -Track [2008]. -Transform [227, 1911]. -tree [961]. -trees [3331]. -Triangulationen [1933]. -Tuple [896]. -tuply [1217, 1326]. -type [2733]. -Version [2129, 2669]. -vertex [1865]. -Vollständigkeit [1461]. -wave [3701]. -Well [2434].

/Algorithmus [1288].

1 [185, 215, 915, 3843, 758, 730]. 1-D [3843]. 110 [2744]. 1130 [603]. 12 [1129]. 1216/1216a [46]. 1216a [46]. 135 [389]. 14 [1172]. 1401 [305]. 17 [534]. 170 [1592]. 170/720 [1592]. 18 [1136]. 19 [3730, 3647, 3734, 3733, 3735, 3736, 3732, 4082, 3737, 3638]. 1959 [68]. 1972 [830]. 1975 [856]. 1994 [4378]. 1996 [4380]. 1997 [4379].

2 [28, 302, 1265, 916, 203, 233, 4240]. 2. [168, 202]. 2.0 [2668]. 2012 [2958]. 2013 [3059]. 2015 [3315]. 2016 [3224]. 2020 [3866]. 22 [1257]. 2D [3808, 2301]. 2D-Stokes [2301]. 2D/3D [3808]. 2nd [297].

36 [1714]. 360 [492, 351]. 3D [3808, 3986]. 3rd [4380]. 3Vs [3419].

4-round [2690]. 4.-6 [830]. 41 [1845, 1813]. 42 [2176]. 44 [427].

50th [3114]. 52 [42, 1129]. 53 [1172]. 56 [1136]. 5G [3514, 3782, 3485]. 5G-based [3485].

6 [104]. 60 [465, 609, 460, 441]. 60th

[439, 106]. 65 [9]. 65th [605]. 6LoWPAN [4052, 3527]. 6th [387].

709- [120]. 7090- [120]. 7094 [120]. 7094-Computers [120]. 720 [1592]. 7th [4378].

802.11s [3055]. 802.15.4 [3469, 3221]. 802.15.4e [3623]. 802.15.4k [3664]. 87j [1714].

90 [1913, 4239]. 90b [1845, 1813]. 90j [2176]. ’97 [4379]. 9th [3138].

= [4369].

A* [3502]. A-DCNN [4016]. a-posteriori [2865]. A. [335, 375, 539, 490, 826, 173, 603, 580, 732, 305, 114, 214, 659, 100, 336, 144, 552, 153, 1845, 30, 38, 263, 116, 203, 254, 204, 233, 334, 62, 282, 348, 541, 123]. A.-R [211]. Abadie [644]. Abbildung [1062, 112, 23]. Abbildungen [766]. Abbrechkriterium [1076]. Abbruch [1142]. Abbruchkriterium [1075]. ABC [2125]. ABE [3308]. abeitungsfreies [414]. Abel [2664, 2670, 1081, 2220]. abelian [2683]. Ableitung [20, 555]. Ableitungen [77, 1438, 567]. Ableitungsfreie [780, 296, 164]. abnormal [2892]. abortability [3678]. Abramowitz [134, 69]. ABS [1968, 1819, 1849, 1968, 1737]. Abschätzung [1021, 1683]. Abschätzungen [780]. Abschnitten [1032]. abscissas [1101, 1335]. absolut [468]. Absolute [1844, 1593, 310, 115]. Absolutely [468]. Abständen [710]. Abstandsbegriff [440]. Abstract [907, 1901, 746, 318, 2653, 54, 773, 2949, 1024]. abstraction [2753, 3567]. abstractly [1114]. abstrakt [1114]. Abstrakte [673, 907, 54, 773, 1024]. academia [2799]. Accelerated [1949, 2982, 3238, 1196, 3521, 2890]. Accelerating [3187, 2512, 864, 1528, 910]. Acceleration [2528, 688, 1093, 1090, 4226, 2990, 3540]. acceleration-based [2990]. Acceptable [1689]. acceptors [1329]. Access [1634, 4186, 2152, 2027, 2963, 3400, 3977, 3156, 3690, 3277, 3768, 4002, 3295, 3487, 3955, 3839]. According [2439, 346, 3622]. Account [2416]. Accounting [45]. Accumulated [1057, 1654]. accumulation [2774]. accumulator [2813]. Accuracy [1794, 2341, 2043, 4183, 1468, 1835, 1189, 2373, 1342, 739, 183, 1615, 1482, 3012, 2199, 4182, 2703, 4184, 1399, 3637]. Accurate [1512, 2619, 2338, 4305, 2478, 1895, 1259, 3966, 2645, 3348, 3808, 4239, 4240, 3279, 4098, 1958]. achieve [4059]. achievements [2917]. Achieving [3675, 1748, 3520]. achsensymmetrische [1941]. acknowledges [3674]. Ackoff [205]. ACO [3983]. acoustic [3353, 2860, 1630]. acquire [3497]. Acquisition [2261]. ACRITH [1634, 4186]. Across [2216, 3785, 3096]. ACS [3027]. Action [681, 3944]. activation [3953, 3296]. Active [2257, 3744, 2947]. activities [3582]. activity [3526, 2892, 2990, 3997, 3241, 3346, 3327, 3212]. actor [2997, 3650]. actor-learner [3650]. actuator [2872, 2874]. ad [2986, 2952, 3853, 3884, 2988, 4057, 3458, 2831]. ad-hoc [2952, 3884, 4057, 2831]. ADAGE [2831]. Adam [114, 123]. Adams [377, 2180]. adapt [2913]. adaptability [3815]. adaptable [2994]. Adaptation [3050, 3533, 2763, 3787, 2912, 2911, 2830, 976, 3747]. Adapted [1524, 2340, 2263]. Adapting [1675]. Adaption [338]. Adaptive [2388, 2146, 2235, 2408, 2094, 2, 2502, 2134, 2430, 2986, 2253, 2434, 2578, 3861, 3863, 3814, 2545, 2095, 1779, 1876, 2429, 2472, 2500, 2226, 2415, 2390, 1881, 4322, 2330, 2708, 2259, 2931, 2097, 3858, 2194, 3151, 3078, 3855, 4329, 3516, 2937, 2323, 2271, 2632, 733, 1733, 3286, 3291, 1525, 1552, 3781, 3780, 3272, 3070, 1242, 2653, 2654, 2664, 2670, 2727, 2850, 2919, 2254, 3335, 3152, 3234, 2834, 2753, 4000, 2861, 2860, 2865, 2908, 2368, 3636, 2870, 2873, 3719, 3444, 3203, 3496, 3764, 2615, 4162, 3867, 3803, 2995, 3843, 3149, 3100, 3942, 2859, 2887, 1339, 1526, 2742, 2178]. adaptive [3480, 3277, 3776, 3725, 1678, 4077, 2956, 2923, 3856, 1358]. Adaptive-Multilevel [2937]. Adaptively [2371]. adaptiven [1339]. adaptiver [1526]. Adaptivity [3694, 2584, 2875]. adäquaten [1349]. ADBench [3607]. Addendum [1553]. adder [2813]. Addition [2582, 1658, 2767, 2770, 2766]. Additional [1969, 3]. Additive [2215, 2180, 2246, 2627, 2075, 2356, 2069, 661]. address [3982, 3023]. Addressing [1587, 3177]. adequate [1349]. Adini [2373]. Adjacent [2294, 4220]. adjustable [3048]. adjusted [3531]. Adjustment [807, 1503]. Admissibility [2483]. admission [3027]. ADMs [2725]. Advance [153]. Advanced [268, 3736, 3325, 3138, 3209]. Advances [157, 3729, 3224, 2974, 3616, 3740, 3739, 3605]. Advantages [1913]. Advection [2288, 2480]. Advection-Diffusion [2288, 2480]. Adversarial [3699, 4056, 3937]. advertising [2977, 3237]. advising [3672]. aerial [3387, 3976, 3925, 4047]. aeroelastic [2857]. AES [2690]. AES-like [2690]. affect [3081, 3364]. affect-sensitive [3081]. affecting [3248]. Affine [4326, 4212, 2812, 2687]. after [3387]. Against [2082, 3629, 3240, 3801, 3485, 3456, 3574, 3905]. agenda [3583]. agent [3552, 2763, 3716, 4058, 3119, 3964, 3225, 3761, 3011, 4040, 3491]. agent-based [4058]. agents [3413]. aggregate [3281]. Aggregation [4342, 2111, 3054, 3326, 3570, 3136, 2627]. aggregation-based [2627]. Aggregation/Disaggregation [4342]. Agile [2991]. Agreement [3113, 3805]. agriculture [3811, 3489, 3971]. agro [3226]. agro-climatic [3226]. Agros2D [2868]. ahead [3880, 3827]. AHP [3633]. AHP-TOPSIS [3633]. Ahrens [1129, 2393]. AI [3629, 3969, 3993, 3584, 4083, 3926, 3336, 3732]. AI-enabled [3732]. AICA [387]. AICA-IFIP [387]. Aid [55, 112]. Aided [539, 478, 1025, 3081, 3994, 2798, 3295]. Air [3468, 4057]. Aircraft [4296]. Airey [893]. Aireyschen [893]. Airtime [3055]. Aiserman [115]. Aizenshtadt [212]. Akilov [72]. al [732]. ALBLA [4000]. Albrecht [1508]. ALCOR [121]. ALCOR-Gruppe [121]. Alefeld [1891]. alert [3313]. Algebra [1145, 4153, 625, 653, 4152, 4370, 4371, 2695, 4238, 27, 485]. Algebraic [4096, 1934, 143, 2366, 2099, 4140, 1998, 2448, 4151, 2413, 1871, 1877, 957, 2065, 2515, 2554, 4150, 1775, 2252, 1501, 2910, 3002, 2172, 4164, 1848, 4136, 155, 1072, 1504, 2571, 2149, 247, 130, 161, 2461, 2111, 2311, 620, 4370, 4371, 1540, 1744, 507, 584, 2717, 2655, 2618, 1642, 1487, 4102, 436]. algebraische [1998, 4102, 436]. algebraischen [130, 161]. algebraischer [4096]. Algebras [747, 2683]. Algebren [747]. algébriques [176]. Algebro [1934]. Algebro-Differentialgleichungen [1934]. Algebrodifferentialgleichungen [1945]. ALGOL [346, 741, 609, 512, 357, 405, 376, 346, 1247, 741, 609, 514, 441, 141, 93, 177, 81, 82, 465, 99, 460, 273]. ALGOL- [357]. ALGOL-60 [441, 460, 441, 465, 460]. ALGOL-Procedures [346, 741]. ALGOL-Programme [81, 82]. ALGOL-Programming [177]. ALGOL-Prozeduren [346, 741]. Algorithm [2431, 1625, 2445, 906, 2286, 2101, 2366, 4336, 1122, 1925, 1397, 2206, 1612, 1281, 1390, 2683, 926, 2243, 2085, 1827, 2137, 2120, 1592, 2158, 4279, 1819, 2590, 1304, 897, 942, 1400, 2535, 1320, 21, 167, 1923, 1643, 2011, 2148, 2057, 1086, 1171, 1788, 1093, 752, 2191, 2393, 2484, 874, 931, 2370, 775, 2207, 2378, 2385, 1853, 2546, 345, 2024, 2127, 1401, 1999, 4323, 2124, 2265, 2422, 803, 4360, 2077, 1092, 1787, 806, 790, 2325, 8, 58, 168, 927, 2072, 1816, 841, 4309, 2121, 1142, 932, 1288, 1920, 2562, 3639, 1931, 2304, 574, 1144, 1257, 840, 2323, 2271, 573]. Algorithm [4264, 791, 930, 1407, 2013, 2109, 1525, 3588, 1552, 3887, 1694, 3305, 3925, 4004, 3188, 3199, 2003, 1287, 1158, 692, 1262, 710, 740, 3611, 1745, 2903, 4025, 1245, 3230, 1313, 1494, 2850, 2696, 1688, 2680, 1974, 2217, 3915, 2839, 3234, 3627, 3876, 3861, 3600, 1557, 1459, 3223, 1247, 1482, 1463, 1202, 3908, 3446, 1377, 3040, 1632, 3319, 3751, 3830, 1237, 2640, 2006, 1481, 3587, 1374, 3441, 1361, 3281, 1652, 3853, 1039, 1118, 3502, 3187, 3953, 3854, 1732, 964, 3529, 2182, 3119, 3808, 3687, 1620, 1428, 3906, 3939, 3500, 1579, 3822, 2781, 2894, 1449, 1375, 3364, 3334, 3601, 3354, 2920]. algorithm [3085, 3435, 3669, 3667, 3082, 3998, 4033, 2616, 3444, 3659, 960, 3350, 3738, 1293, 3525, 1561, 3068, 3522, 4074, 3881, 1022, 1613, 1103, 1095, 1651, 3764, 518, 1503, 2844, 1681, 3836, 996, 3888, 2926, 2622, 1015, 570, 2625, 3484, 3470, 3078, 3632, 1136, 3975, 1054, 2995, 3843, 3845, 1566, 1100, 1317, 1487, 3766, 3705, 1562, 614, 1383, 2041, 1056, 4236, 1206, 3760, 3633, 3389, 3519, 3660, 714, 3942, 3553, 2596, 3790, 2895, 3873, 1341, 1737, 973, 3028, 3769, 3135, 2662, 2658, 3692, 720, 1439, 3640, 2178, 1248, 1272, 1156, 1992, 1322, 495, 2607, 992, 967, 1975, 4008, 2932, 3531, 1076]. algorithm [4077, 3289, 3558, 1356, 3869, 3046, 3498, 3613, 3047, 2736, 3279, 1611, 978, 346, 454, 3159, 1202, 298, 278, 400, 1092, 166, 7, 25, 414, 1113, 371, 601, 82, 133, 528, 791, 930, 244]. Algorithm/Algorithmus [1202].

Algorithmen [346, 465, 1961, 915, 916, 1950, 594, 612, 953, 833, 740, 1993, 741, 379, 742]. Algorithmenklasse [2000]. Algorithmic [1546, 4352, 1050, 294, 306, 1513, 2220, 2808]. Algorithmical [2523]. algorithmische [141]. algorithmisches [294, 306]. Algorithms [1441, 1177, 1360, 465, 4097, 1484, 2270, 2409, 2335, 2586, 1903, 1993, 1985, 4287, 2490, 2061, 954, 1437, 1386, 754, 2470, 2540, 2145, 1134, 1004, 870, 2105, 4197, 1112, 1961, 1121, 1802, 4191, 1798, 1860, 2351, 1309, 1970, 2000, 1197, 399, 1548, 1306, 2556, 1273, 1331, 4190, 1057, 1166, 915, 916, 1120, 4117, 4118, 4180, 1950, 751, 835, 1809, 2200, 2284, 1538, 2042, 932, 229, 2126, 4273, 1504, 3710, 1319, 953, 449, 4185, 2169, 4122, 892, 1205, 2175, 165, 1563, 1505, 3565, 1198, 2930, 346, 3966, 1129, 2752, 2020, 3537, 3253, 3310, 4125, 2018, 3353, 3714, 1231]. algorithms [1264, 737, 2600, 4044, 1570, 4157, 2019, 2660, 3479, 3541, 3931, 2676, 1403, 1344, 1719, 1743, 2176, 1108, 3701, 379, 2818, 1017, 3284, 2972, 3193, 3215, 944, 3229, 2721, 1845, 594, 612, 3556, 3331, 3563, 3282, 3596, 3794, 3362, 1362, 1528, 1607, 3644, 4036, 2904, 3259, 3916, 3703, 3283, 1692, 3063, 1714, 981]. Algorithmus [4097, 692, 740, 1632, 167, 1923, 2011, 461, 964, 345, 1999, 960, 8, 58, 168, 570, 1920, 1931, 601, 840, 720, 495, 791, 930, 2013, 346, 454, 710, 889, 1202, 897, 1320, 21, 1361, 298, 278, 400, 166, 1142, 1288, 7, 25, 414, 371, 82, 133, 528, 791, 930, 1076, 244]. alias [1734]. Alibaba [4030]. Aligned [2348]. Aligning [2797]. alignment [3050, 3796]. All-dummy [3659]. All-Integer [752]. ALLC [4062]. allen [156, 807]. aller [692, 298]. Allgemeine [185, 708, 303, 4104]. allgemeinen [307, 904, 1016, 867]. allgemeiner [767]. allgemeinere [6, 751]. allgemeinerem [440]. Allocation [2187, 865, 3482, 3580, 1147, 3269, 3615, 3940, 4022, 3319, 3901, 3351, 3104, 3765, 3832, 4072, 3756, 3193, 4066, 3881, 3106, 4082, 4021, 4064, 3602]. allocations [3684]. allreduce [3750]. alpha [4316]. alpha-Sectors [4316]. Alphabets [304]. alphanumerischer [234]. Alternate [1450]. Alternating [4312, 687, 1297, 1416, 2474, 3576, 3820]. Alternative [985, 2535, 1113, 3129]. Alternativroutensuche [1113]. Always [887, 823]. AMAE [3699]. Amazon [3546, 4064]. AmazonAICloud [4064]. ambiguity [557]. Amendment [668]. Amer [211]. America [104]. AMG [2505]. Ammon [136]. among [3133, 2770, 3323]. amplifying [2756]. Analog [493, 645, 6, 51]. Analogie [112]. Analogiques [491]. Analogrechenschaltungen [456]. Analogrechentechnik [136]. Analogrechnen [795]. Analogrechner [6]. Analogrechners [51]. analogs [407]. Analogue [89, 322, 802, 456]. analogy [112]. Analyse [405, 1942]. analysers [100]. Analysing [3547, 4020]. Analysis [542, 83, 2335, 72, 1723, 1122, 4244, 4297, 1448, 1803, 113, 1830, 2187, 3954, 1450, 1851, 1386, 4290, 249, 1722, 2107, 4250, 4197, 2011, 2148, 2351, 1849, 2038, 1784, 4112, 2422, 94, 1197, 2595, 1854, 2373, 2556, 886, 709, 781, 1379, 1498, 2327, 2087, 29, 3185, 266, 1319, 4036, 4122, 2074, 2109, 2529, 1942, 3447, 3306, 1644, 4368, 4382, 3664, 3299, 3180, 3802, 1497, 3148, 3615, 2891, 2917, 1211, 2918, 1459, 3670, 1822, 3788, 3276, 3394, 2676, 1403, 460, 2976, 2677, 2892, 3274, 3772, 3218, 3401, 441, 3098, 3969, 3485, 2674, 2735, 3349, 3645, 3358, 3699, 3499, 2848]. analysis [3842, 4010, 1220, 4055, 3266, 3128, 1986, 3080, 2831, 2823, 2821, 2882, 3965, 2761, 951, 1221, 694, 3142, 1526, 3222, 1130, 1766, 3280, 4054, 3418, 3383, 139, 174, 509]. analysis-based [3080]. Analytic [1378, 1591, 375, 109, 2216, 1624, 1549, 993, 2265, 2247, 850, 2365, 1192, 1010, 983]. Analytical [3218, 301, 3626, 3443, 3877]. analytics [3397, 3403, 3029, 3416]. Analytischen [850, 993]. analytischer [983]. analyze [4072]. analyzer [3127]. Analyzing [3750, 506, 3546]. anderen [807]. Änderung [244]. Andree [213]. Android [3662]. Anfänger [832]. Anfangs [1045]. Anfangs-Randwertproblems [1045]. Anfangsrandwertaufgaben [1254, 445]. Anfangswertaufgaben [988, 912, 1125, 1011]. Anfangswertprobleme [18, 1958, 1339]. Anfangswertproblems [599]. Anfangwertaufgaben [877]. angepassten [854]. angepaßtes [4127]. Angermann [214]. Angewandte [818]. Angewandten [403, 16, 357, 10, 809, 404]. Angle [2266]. Anisotropic [2648, 1886, 2204, 2159, 4277, 2626]. anisotropically [2865]. Anisotropy [2348, 2433]. Anke [702]. Ankündigungen [734]. Ankündigungen/Announcements [734]. Anlagen [749, 958]. Anleitung [512, 270]. ANN [3665]. ANN-based [3665]. Annealing [2577, 2326, 1657, 3236]. anniversary [3114]. Annotated [4279]. annotation [2965, 3519]. Annotations [1319]. Announcements [734]. Annual [419]. Annuities [4351]. Anomalies [3323, 3534]. Anomaly [3726, 3893, 3495, 3895, 3798]. anonymity [3068]. anonymization [3558]. anonymous [3057, 3716, 3471, 3579, 3753, 3659, 3438]. Anregungen [904]. Ansatzes [1107]. Ansatzfunktionen [760]. answers [3949]. Ant [3334, 3230, 3914, 4005]. Anti [2423]. Anti-Symmetric [2423]. antifragility [3590]. antisymmetric [600, 720]. antisymmetrische [720]. antisymmetrischen [600]. Antwortzeitgesteuerte [1042]. Anwendung [611, 467, 51, 4128, 770, 868, 482, 982, 314, 751]. Anwendungen [455, 185, 200, 238, 54, 1939, 990, 1955, 77, 181, 336, 436]. Anwendungsgebiete [389, 481]. Anwendungsmöglichkeiten [4103]. any [1097, 1062, 1380]. AOR [1721, 1196]. apnea [3251]. apparent [3828]. applicable [2926]. Application [2431, 419, 2352, 455, 1804, 2934, 2262, 77, 2578, 2591, 149, 1623, 2057, 2287, 2329, 2330, 770, 2675, 378, 3411, 2466, 900, 4131, 751, 1782, 2195, 2960, 227, 1594, 868, 3063, 2878, 1044, 440, 1222, 3152, 3785, 3786, 1247, 982, 611, 1364, 1210, 3787, 2677, 3504, 4129, 3354, 3951, 571, 1422, 4167, 3173, 3362, 980, 3548, 3644, 694, 1911, 1780, 3640, 2932, 3829, 2989, 3934, 1418, 314]. Application-oriented [2960]. Applications [4332, 4293, 1251, 1167, 1886, 1769, 1955, 4153, 54, 388, 4178, 894, 2384, 71, 2028, 1400, 2531, 1939, 1878, 2551, 1881, 4191, 4356, 4383, 1108, 4114, 1658, 2510, 2499, 4347, 4119, 4275, 2963, 3911, 4376, 2913, 3162, 2891, 3398, 3004, 4372, 185, 3621, 238, 3410, 3062, 3814, 3061, 3686, 4027, 3056, 2640, 3333, 3749, 2908, 2720, 4087, 1542, 1582, 3572, 3943, 3500, 2983, 1560, 3964, 4028, 3603, 3740, 3739, 2832, 3341, 3960, 3460, 2830, 3325, 2794, 1702, 3833, 3071, 229, 3936, 4053, 3209, 2900, 1708, 4065, 2751, 2964, 2975, 3482, 3385, 268, 289, 265, 359]. Applied [372, 2509, 178, 35, 36, 784, 476, 1927, 174, 42, 2349, 237, 61, 40, 2128, 1161, 1761, 1227, 1707, 3204, 1080, 2863, 1349, 1661, 3193, 452, 3224, 2881]. Appraisal [1592]. Approach [4332, 109, 1915, 771, 4156, 4202, 2565, 1893, 2025, 1877, 4254, 2303, 1867, 1197, 399, 4271, 1391, 1842, 4273, 2056, 3744, 3369, 3213, 3133, 3374, 3387, 3552, 3355, 3013, 2752, 4222, 3070, 3614, 3618, 3398, 3536, 3301, 2639, 2918, 3081, 3174, 1479, 3785, 3755, 3754, 4039, 2834, 1657, 1744, 1991, 2965, 3403, 950, 4058, 3928, 4000, 3089, 3982, 2991, 3196, 3953, 4087, 2938, 2912, 3672, 1669, 3504, 3390, 3555, 3069, 3994, 3122, 3345, 3595, 3693, 3969, 3391, 3812, 3903, 3597, 3617, 3841, 3896, 2612, 3208, 2888, 3311, 3967, 3599, 3584, 1485, 2694, 2833, 3226, 3022, 3735, 3205]. approach [3051, 2966, 3920, 2618, 1266, 2822, 3975, 3731, 3433, 2837, 3622, 3965, 3366, 3489, 3100, 3191, 3494, 1597, 3357, 4082, 3593, 3760, 3759, 3077, 3332, 1221, 3190, 3692, 3535, 3222, 3424, 3449, 4018, 1345, 3860, 3457, 2942, 4013, 3434, 3722, 4061, 3707, 3762, 2736, 4089]. Approaches [516, 192, 3559, 2851, 3961, 3924, 3828, 3630, 3242, 4024, 3203, 3589, 2758, 2004, 3112]. Appropriate [1397, 598]. Approximants [1786, 1675, 1636]. Approximate [1378, 2930, 1128, 4204, 824, 2082, 1878, 1877, 2383, 2314, 2404, 1527, 782, 464, 4216, 394, 865, 3614, 1667, 2703, 3069, 1050, 1742, 1762, 445, 1577, 1457, 3157, 849]. Approximated [705]. Approximating [1783, 1134, 1940, 2344, 4070, 1920, 1772]. Approximation [1591, 1360, 1801, 4241, 2408, 2, 1409, 2467, 2502, 2134, 52, 1621, 16, 285, 2279, 1825, 578, 4300, 618, 687, 660, 846, 1304, 211, 1940, 780, 2496, 2362, 2449, 2390, 2551, 2542, 2543, 4209, 4101, 2336, 761, 1885, 878, 457, 1381, 4, 66, 804, 768, 849, 2189, 2533, 1000, 2410, 805, 4346, 2213, 1267, 2416, 1094, 1920, 1848, 1824, 500, 318, 956, 4367, 390, 4275, 762, 368, 1063, 2482, 1608, 2305, 209, 2197, 4263, 110, 1037, 1287, 1149, 1418, 1010, 113, 1089, 2145, 843, 1019, 2660, 339, 504, 382, 693, 1561, 292, 1001, 1069, 2637]. approximation [2886, 1322, 992, 1026, 1055, 967, 1012, 1193, 159, 1376, 1683, 1409, 206, 949, 585, 504, 693, 707, 152, 110]. Approximationen [52, 1013, 370, 19]. Approximations [4110, 4111, 4246, 4203, 2292, 2166, 2464, 370, 1252, 1303, 2358, 1029, 1958, 537, 1466, 2633, 2861, 1060, 1173, 2750, 19, 1305, 1244, 2041, 329]. Approximationsmaße [368]. Approximationsprobleme [1089, 585]. Approximationsproblemen [339]. Approximationstheorie [66, 324, 209, 113]. approximieren [445]. April [4379, 830]. äquidistanten [1102, 785].

Äquivalenzkalkül [223]. Arbeitstagung [245]. Arbitrarily [1062]. Arbitrarily-Prescribed [1062]. Arbitrary [1864, 2227, 2476, 2395, 1233, 2248, 684, 3701, 2815, 3, 3997, 4236]. arc [2712]. arc-length [2712]. arced [3029]. Architecting [3780, 3779, 3116]. Architectural [1899, 3139, 2918, 3783, 3124]. Architecture [2919, 2917, 3057, 4019, 3533, 3594, 3923, 4017, 3060, 3163, 3832, 3177, 3435, 3371, 2827, 3467, 3775, 3521, 3075, 3564, 3778, 3760, 3826, 4065, 3809, 3851]. Architecture-based [2919, 2917]. Architectures [2028, 3781, 3037, 3748, 3880, 3740, 3739, 3072, 3842, 3650, 3201, 3784]. archive [3873]. Arcs [4275]. Area [3275, 2560, 4365, 3206, 3924]. Areas [855]. Arguments [4183, 1436, 4184, 2661, 269]. Arising [2010, 2355, 2144, 2242, 517, 2700, 1572]. Arithmetic [4109, 159, 128, 1512, 1496, 2585, 1187, 1183, 4151, 1680, 1189, 2245, 2287, 1879, 458, 764, 883, 4366, 4104, 4113, 2466, 1914, 4114, 4115, 1904, 1919, 1181, 1268, 1550, 2473, 2587, 343, 501, 1371, 2193, 868, 4108, 2492, 861, 4124, 1035, 1075, 3626, 1497, 2812, 1765, 1176, 4182, 1152, 1545, 1186, 1308, 2768, 2641, 4074, 2814, 2816, 4234, 1443, 408, 1571, 852]. Arithmetic-Average [2473]. arithmetic-based [3626]. Arithmetical [4180]. Arithmetically [1513]. Arithmetics [2552, 2421, 2503, 885, 2819]. Arithmetik [1545, 883, 4109, 764, 4366, 1914]. Arithmetiken [852]. armed [4056]. Arnold [575]. AronszajnBazley [682, 682]. AronszajnBazley-Fox [682]. AronszajnBazley-Fox-Verfahrens [682]. Array [4126, 3400, 3748, 4129, 1642, 1487]. Arrays [1587, 1385, 2001, 1056]. Arrib [170]. Arscott [11]. Art [619, 314, 3026, 2777, 1095, 1248, 1469, 1957].

ART-based [2777]. artefacts [3466]. Arterial [2441]. arterio [2754]. artery [3290]. artery-specific [3290]. article [1129, 3033, 2895, 3189]. artifact [3087, 3391, 3424, 3434]. artifact-centric [3087, 3434]. artifact-driven [3424]. Artificial [3159, 2798, 2060, 3792, 3726, 3032, 3349, 3525, 3718, 3717, 3888, 3125, 3707]. ary [2745]. Asaithambi [2622]. ASH [2843]. Asian [2473]. ASN [592, 592]. ASN-functions [592]. ASN-Funktionen [592]. Aspects [1330, 485, 4359, 1816, 620, 943, 2690, 1707, 695, 663, 2886]. Aspekte [695]. ASRNN [3958, 4091]. ASRNN-bi-directional [3958, 4091]. Assembler [290, 829]. assembly [3254, 2798]. assembly-free [3254]. assess [2810, 2960]. Assessing [3341]. Assessment [3731, 3249, 3626, 4015, 3265, 3083, 3039]. Assets [2568]. Assigned [1866]. assigning [3388, 3433]. Assignment [2008, 1281, 1390, 1047, 1450, 1639, 2282, 2283, 2186, 2169, 1509, 1556, 1620, 3633, 593, 3084, 3768]. assimilation [2677]. assistance [3179]. assistant [3248]. Assisted [1932, 1841, 3688, 3374, 3529, 3846]. Associated [1869, 5, 1457]. Association [4096, 2280, 4102, 3920, 2792]. associative [1123]. associativity [3041]. associées [258]. Assuming [2366]. Assumption [2491]. Assumptions [2065, 2644]. assurance [3061]. astronomical [3843]. Asymptotic [1769, 1421, 641, 1854, 1887, 163, 572, 2192, 999, 1458, 307, 1021, 1447, 4, 992, 1408, 1260]. Asymptotically [52, 2093]. asymptotisch [52, 1260]. asymptotische [307, 163]. asymptotischer [4, 572]. asymtotischen [1021]. Asynchronous [1484, 479, 4339, 2841, 2797, 3579, 3257, 3843, 4034, 3576]. Athans [169]. atomic [3753]. atoms [2796]. ATT&CK [3970]. attack [2689, 3136, 3415, 3527]. Attacks [3514, 3863, 3801, 4052, 3456]. Attainable [2295, 2516]. attention [3885, 3771, 3378, 3883, 3941, 3934, 3796, 3989]. attentional [3468]. attentions [3905]. Attentive [4265]. attracted [3279]. attribute [3977, 3895, 2794, 3535]. attribute-based [3977]. Attributed [4283, 3557, 3532, 3175, 3267]. attributes [3371, 3703]. attribution [3228, 3943]. audio [3386, 4071]. auditing [3931, 3572, 4075]. aufbau [570]. Aufbaünd [1170]. Aufgabe [1884]. Aufgaben [186, 286, 262, 576, 972]. Aufgabensammlung [65]. Aufl. [44, 817, 512, 626, 39]. Auflösung [465, 241, 413]. Aufstellung [224]. auftretende [679]. Auftretenden [4105]. Aufzählbarkeit [626]. Aufzeichnung [240]. augmentation [2723, 3982]. Augmented [2123, 3560]. Augmenting [1047, 2169, 1117, 1620]. Ausbauchens [1943]. Ausbildung [643]. Ausdrücke [478]. Ausführung [234]. Ausgabe [1376, 952]. ausgedehnter [855]. Ausgleich [1547, 209]. Ausgleichsaufgaben [973]. Ausgleichsformeln [1128]. Ausgleichsprobleme [2011]. Ausgleichsproblemen [1349]. Ausgleichsrechnung [195, 431]. Ausnahmen [1269]. Außenraumproblems [385]. Aussteuerung [267]. Auswahl [1318]. Auswahldisziplin [1356]. Auswertung [114, 460, 1944, 441]. Auswertungsalgorithmen [981]. Auswirkung [248]. authenticated [3805]. authentication [3573, 3478, 3912, 3643, 3816, 2985, 2988, 3973, 3340, 3635, 3102]. Author [4155]. authoring [3809]. authorship [3228]. auto [2839, 3699, 3149]. auto-encoder [3699]. auto-scaling [2839, 3149]. autoconfiguration [3023]. autoencoders [3516]. Automat [79]. Automata [438, 1146, 49, 299, 361, 131, 344, 911, 590, 479, 507, 2797, 4000, 2992, 4369, 2844, 715, 834, 3309, 4137, 422, 659, 746]. automata-based [2844]. automate [3792]. Automated [3276, 4215, 4258, 3619, 3686, 3670, 3589, 3875, 3638, 4001]. Automaten [650, 412, 673, 590, 438, 49, 131, 4369, 671, 743, 198, 344, 175]. automatentheoretishcer [342]. Automatentheorie [698, 686, 102, 816]. Automatic [3369, 353, 354, 419, 73, 2909, 1803, 639, 979, 2965, 240, 76, 2086, 3958, 4169, 709, 162, 841, 1716, 1288, 394, 29, 389, 481, 3865, 2323, 183, 259, 2883, 3164, 950, 3172, 1039, 1547, 4373, 3363, 4119, 3171, 2754, 4091, 1062, 81]. automatical [1483]. Automatically [1480, 1553]. Automating [3099, 3665]. Automation [311, 210, 138, 486, 197, 3109]. Automatische [240, 201, 224, 1062, 81]. automatischen [338]. automatischer [1547]. automatisierte [282]. automatisierten [643]. Automatisierung [252]. Automaton [239, 1897, 589, 342, 1031, 2744, 412]. automaton-theoretical [342]. Automorphism [208, 79, 239]. Automorphismen [903]. Automorphismengruppe [208, 239]. Automorphisms [589, 911, 903, 1064]. autonomen [971]. autonomer [182]. Autonomic [3110, 3447]. Autonomous [1168, 4062, 2785, 3976, 3223, 971, 3832, 3122, 3605, 3813, 3607, 3302]. autoscaling [4072]. AUV [3914]. Auxiliary [604, 2113]. Availability [3180, 565, 3786, 3066, 3594, 3347, 3351, 2969, 2996, 3304, 3775, 3096, 2929, 3564, 4023, 3222]. availability-aware [3786]. Average [749, 1231, 2473, 2224, 1327, 3646, 1217, 3844]. Averages [2344]. AVL [1148]. avoid [1705]. avoidance [3569]. Avoiding [2235, 2025]. awakens [3258]. Aware [3204, 2306, 3625, 3043, 3272, 3402, 3220, 3533, 3461, 3197, 3657, 2854, 3223, 3786, 3159, 3009, 3545, 3923, 3103, 3538, 3853, 3691, 3255, 3854, 3787, 3141, 3897, 3687, 3390, 3822, 3181, 4028, 3354, 3669, 3774, 3896, 3719, 3109, 3468, 3584, 3427, 3960, 4010, 3650, 3217, 3932, 3278, 3864, 3149, 2846, 3942, 3962, 3448, 3367, 3465, 3084, 3886, 4018, 3869, 3797, 3715, 4089, 3496, 3653]. awareness [3233]. axes [3029]. Axiomatisierung [1852]. axiomatization [1852]. Axis [2177]. Axisymmetric [2134, 1941, 2533].

B [543, 48, 492, 289, 418, 427, 147, 1700, 101, 576, 75, 744, 63, 125, 247, 351, 481, 4312, 1025, 1003, 4246, 1596, 4249, 1395, 4306, 4211, 2791, 1442, 1354, 1548, 2880, 4311]. B-Rep [4312]. B-spine [1003]. B-Spline [4306, 1548, 4311, 1025, 4246, 4249, 1395, 2791, 1354]. B-Spline-Kurven [1003, 1395]. B-Splines [1596, 4211]. B-Trees [1442]. B. [420, 542, 178, 509, 70, 485, 249, 180, 701, 42, 117, 260, 802, 548]. B.V.P.s [1878]. B2B [3012]. Babuska [135, 2546]. Babuska-Summation [2546]. back [3400, 3925]. Back-and-Forth [3925]. back-end [3400]. backed [3606]. backpropagation [3685]. Backward [2087, 1880, 1983, 1221]. Bad [2526, 2989]. BaF [3925]. bag [4027, 3367]. bag-of-tasks [4027]. Bagging [3350, 3738]. Bailey [71, 2903, 4384]. Balance [2581, 3950]. Balanced [2096, 2384, 2472, 3090, 1137, 2069, 1104, 4262, 3745, 3200, 3242]. Balancing [2273, 3530, 2914, 3441, 4000, 4060, 3595, 2837, 3760, 3826, 2840, 3268]. Balas [880, 880]. Balintfy [228]. Ball [4242, 4347, 1929, 2608, 1754, 1190]. Banach [2269, 985, 526, 4274, 934]. Banach-Räumen [526]. Band [330, 28, 238, 418, 579, 646, 674, 46, 10, 576, 66, 7, 389, 481, 124, 203, 758, 502, 511, 658, 1066, 4028, 1311, 560]. Banded [1772, 1699]. bandits [4056]. Bandmatrix [7]. Bandwidth [735, 2251, 939, 1393, 3580, 3221, 717, 3184, 3494]. Bank [119, 63, 3124].

Bankd [488]. banking [3160]. Barnett [246]. Barrier [2386, 3499, 1236]. barrier-synchronized [3499]. barriers [1710, 1813]. Barron [373]. Bartsch [171]. Bary [172]. Barycentric [1786, 2556, 1842]. base [3560, 3551, 1959]. Based [2146, 1170, 1723, 2585, 1993, 2275, 1771, 1962, 2376, 4290, 2236, 4318, 2412, 2226, 2245, 2483, 2086, 4254, 2342, 2282, 2505, 2554, 2441, 2374, 4255, 2455, 4325, 2072, 2240, 4214, 2060, 1770, 4281, 2460, 2104, 1772, 2462, 4285, 2379, 1250, 4056, 3043, 3821, 3559, 3365, 3655, 3744, 2953, 3641, 3430, 3474, 3781, 3956, 3440, 3488, 3400, 3857, 2811, 3013, 3680, 3734, 3537, 3992, 3199, 3626, 3665, 3070, 3682, 3634, 2849, 2501, 4384, 3834, 2917, 1064, 3301, 2696, 3244, 2488, 4019, 3915, 2919, 3164, 3643, 2788, 3234, 3353, 3627, 3914, 3876, 3940, 2729, 3755, 3754, 4039, 3657, 3223, 2777]. based [3945, 3034, 3066, 3009, 3594, 3711, 3523, 4027, 2756, 3712, 3581, 3893, 3526, 2965, 3575, 3751, 3830, 3453, 4085, 3239, 3783, 4078, 4058, 3570, 3587, 1374, 3859, 3441, 3961, 3610, 3351, 3394, 3982, 3502, 2991, 3033, 3789, 3980, 3933, 2992, 3399, 3854, 2114, 3141, 3706, 3927, 3572, 3529, 3727, 3765, 2912, 3409, 3079, 3832, 3930, 3897, 3943, 2931, 4048, 2969, 3375, 3894, 3990, 3995, 3687, 4379, 3672, 2701, 3504, 3695, 3555, 2892, 2985, 3069, 3300, 3585, 3344, 3772, 3401, 2990, 3658, 3811, 3356, 3181, 3334, 3977, 4024, 4072, 3485, 3812, 3287, 3354, 2704, 2988, 3179, 3349, 3435, 3774, 3896, 3944, 3996, 3895, 3998, 3109]. based [3439, 3444, 3468, 3659, 3628, 3704, 3522, 4074, 3881, 3511, 3761, 4038, 3307, 1516, 2888, 3458, 3599, 3251, 3554, 3513, 4063, 2694, 2833, 3120, 2844, 3226, 3030, 3902, 1752, 3836, 3265, 3735, 4086, 3460, 2926, 3793, 3650, 3926, 4051, 3121, 3521, 2830, 3567, 3266, 2987, 3713, 3044, 3096, 3484, 3309, 2793, 3470, 3997, 3382, 3136, 2995, 3459, 3080, 3999, 3019, 3991, 4052, 4042, 3146, 2824, 2821, 4158, 3314, 3023, 2738, 991, 3517, 3563, 3282, 3838, 3456, 3864, 3705, 2021, 3489, 3100, 3312, 2771, 3330, 2627, 3357, 3250, 3393, 4053, 3826, 3352, 3477, 3633, 3759, 2802, 4034, 3011, 3389]. based [3519, 3660, 3635, 4040, 3147, 3245, 3553, 4020, 4065, 2792, 3377, 3873, 3962, 2665, 3028, 3476, 3769, 2482, 3815, 3791, 3620, 3703, 3795, 3283, 3694, 3871, 3640, 3386, 3424, 3604, 4071, 3527, 3277, 4008, 2975, 3380, 3776, 3798, 3725, 3805, 3768, 3839, 3819, 4064, 2477, 1563, 3236, 3978, 4067, 4077, 3294, 3376, 3685, 3979, 3378, 3851, 4002, 3383, 3539, 3829, 3892, 3860, 3457, 3483, 3157, 3799, 3233, 2994, 3206, 3883, 3046, 3231, 3498, 3613, 3609, 3707, 3797, 2989, 3288, 3709, 3777, 3934, 4011, 3985, 3420, 3796, 2736, 2981, 3237, 3481, 3296, 3437, 3443]. Bases [4242, 2502, 2252, 1029, 4331, 1502, 1522, 2605]. BASIC [832, 4151, 4355, 2309, 4190, 333, 681, 1283, 1196, 2867, 2768]. Basics [238]. basing [1678]. Basis [2134, 4251, 2451, 764, 883, 2214, 2547, 2060, 4264, 4331, 2641, 2140, 1022, 2715]. basis-inverse [1022]. Basiserweiterung [1959]. Basiswahl [764, 883]. BAT [3790]. Batch [3346, 588, 1321, 3340]. Batchelor [68]. Bates [279]. BATMAN [2936]. Bauer [46, 10, 534, 389, 481]. Bauknecht [672]. Baum [961]. Baumann [121]. Baumgartner [44]. Bauwesen [200]. Bayer [510]. Bayes [3143, 3825]. BBR [3460]. BBR-based [3460]. Bd [357, 362, 390, 404, 403]. BDDC [2644, 2666, 2937]. BDPS [3645]. be [2818, 2175, 2114, 2741, 2742]. beacon [4077]. beam [1708]. beamforming [3671]. Bearing [3587]. Bechenbach [61]. Bedeutung [92, 32, 282, 110]. bedingte [456]. Bedingung [500]. Bedingungen [1922]. Bee [3159, 3667, 3525, 3707]. Begriff [4107]. Begriffe [54]. Begriffes [78]. Behandlung [1943, 440, 16, 1089, 559, 972, 1046, 986, 1043, 1126, 994, 973, 789, 1226, 1339].

Behavior [1879, 2516, 130, 161, 1602, 3407, 2804, 1231, 3161, 3343, 3894, 2899, 3586, 1107, 3791, 148, 192]. behavioral [3164, 2954, 3434]. Behaviour [4364, 435, 524, 1861, 950, 480, 1577, 3313, 1408]. Behnke [238, 418]. Beim [1058, 556, 752, 4105, 2009, 1142, 862, 209, 1947]. Beispeil [840]. Beitrag [597, 112, 980]. Beitrage [29, 30, 37, 31, 33, 32, 38, 43, 158]. Bekey [535]. beliebig [1062]. beliebigen [684, 3]. belief [3628]. BELL [81, 82, 4331, 624, 81, 82]. BELL-Programmen [81, 82]. Bell-Shaped [4331]. Bellman [319, 280, 300]. belonging [1558]. below [3674]. BEM [2501, 2500, 2010, 2702, 2790, 2889]. Bemerkung [438, 572, 24, 1256]. Bemerkungen [879, 607, 693, 381]. Benchmark [2434, 2907, 3064, 3691, 3548]. benchmarking [2960, 3607]. benchmarks [3749]. Bending [4177, 1708]. Beneficial [163]. benefit [3433]. benefits [3833]. Benutzung [603]. Beobachtungsergebnissen [114]. Berechenbarkeit [626]. Berechneten [705]. Berechnung [823, 1957, 692, 432, 632, 1540, 466, 586, 638, 21, 240, 1999, 1754, 1402, 162, 1236, 534, 1070, 1082, 908, 444, 863, 592]. Berechnungen [747]. Berezin [12]. Berge [347]. Berichte [200]. Berichtigung [883, 601, 668]. Bernoulli [1596, 2649]. Bernstein [2412, 2080, 2556, 2822, 1100, 1317]. BERT [3979]. Bertin [320]. Bertram [238]. Berücksichtigung [185, 85, 384, 389, 481]. beschleunigter [1949]. besondere [113]. besonderer [185, 85, 384, 389, 481]. Bessel [1957, 20]. Besselfunktion [1957]. Besselfunktionen [20]. Best [4203, 1924, 660, 777, 370, 1971, 1115, 791, 930, 3767, 457, 3246]. Best- [1971]. bester [370]. Bestimmung [710, 598, 859, 914, 1547, 370, 848, 298, 379, 960, 163, 1931, 665, 903, 182, 791, 930, 293, 208, 684, 780, 763]. Beta [788, 1129, 1952, 2325]. beta-verteilten [1952]. BETI [2651, 2696]. Betrachtungen [252]. betriebliche [248]. Betriebsmittelsteuerung [1170]. Betriebssysteme [105]. Betriebssystemen [808]. Betriebswirtschaft [123]. Betriebswirtschaftslehre [250]. Betriebswissenschaftlichen [579]. Better [1529, 2796]. Between [2269, 79, 278, 4293, 1472, 2727, 2939, 2819, 3751, 4268, 1659, 2454, 1182, 1660, 3099, 2947, 4059, 3566, 1148, 3136, 2266, 1439, 2128, 4220, 4285, 2971, 3637]. beun [397]. Beweis [694]. Beweisalgorithmen [165]. Bewertete [840]. bewerteten [293, 741, 737]. Bewley [217]. Beyond [4020]. Bézier [1395, 4245, 4301, 4208, 1395, 2453, 2454, 4210, 4256, 4219, 4220, 2719]. Bézier-Kurven [1395]. BFGS [2789]. BGP [2947]. Bi [3199, 4268, 3989, 3958, 4091, 2108]. Bi- [4268]. Bi-CG [2108]. Bi-directional [3989]. Bi-objective [3199]. BIAD [3273]. BIAS [2044]. biased [3510]. Bibliography [67, 1319]. BiCGStab [2411]. biconnected [1357]. Bicriteria [1782]. Bicubic [959, 4308]. bicycle [3134]. bidirectional [3919]. Bifurcation [1872, 1523, 1863, 2184, 2165, 1907, 1717, 1753, 1422, 1686, 1226, 1696]. Big [3127, 3419, 3422, 2944, 3417, 3196, 3566, 3412, 3071, 3195, 3416, 3455, 3381, 3211, 3383, 3410, 3548]. BigData [3029]. biharmonic [974, 1156]. Bijleveld [73]. bike [3408]. bikubischer [959]. bilinear [2650, 2726]. Billeter [757, 417]. billets [2877]. Billing [175]. BiLSTM [3379]. BIM [3198]. BIM-enabled [3198]. Bin [1643, 1947, 1971, 2109, 2210, 2679, 1557, 2019, 1344, 3708, 1884]. Bin-Packing [1643, 1947, 1557, 1884]. Binäre [549]. binären [880, 501]. binärer [55]. Binary [55, 501, 2068, 2150, 3310, 3723, 316, 1691]. binary-wise [3723]. Binärziffern [316]. binding [3292]. Bingham [1941, 1941]. Bingham-Typus [1941]. Binomial [788, 2236, 1216, 1129, 4227, 592, 1780]. Binomialverteilung [592]. Bio [3213, 3335, 3544, 3305, 2833, 3582, 2796]. bio-geography [3582]. Bio-inspired [3213, 3335, 2833, 2796]. bio-inspried [3305]. Bio-signal [3544]. Bioelectromagnetic [2852]. Bioinspired [2829]. biology [2796]. Biorthogonal [2447]. Biostatistics [640]. bipartite [2678]. bipolar [4071]. biquadratic [1873]. biquadratischen [863]. birds [3888]. Birkhoff [1555]. Birthday [439, 106, 605]. Bisection [1569, 1311]. Bisectors [4316]. Bit [1658, 1342, 1371]. bits [2796]. BitTorrent [3097]. Bivariate [2556, 1933, 1139]. bivariaten [1933]. Black [2275, 685, 3190]. Black-Box [2275, 3190]. Blackman [2352]. Blackman-Tukey [2352]. Blaschke [1786]. Bleitungsfreie [413]. Blending [4200, 2071, 1922, 4216, 2492, 2442, 1922]. Blending-Funktionen [1922]. Block [2773, 2366, 2289, 784, 2512, 5, 2203, 2030, 2159, 1772, 2734, 3145, 3959, 1359, 4236, 2744, 2681, 2739]. block-by-block [1359]. Block-Factorization [2030, 2159]. Block-Pentadiagonal [2366]. Block-Stochastic [5]. Blockchain [3572, 4002, 3483, 3956, 4016, 3648, 3931, 3789, 4038, 3599, 3833, 3456, 3681, 3653, 3851, 3539, 3679]. Blockchain-based [3572, 4002, 3483, 3789, 3539]. blockchain-enabled [3653]. blockchains [3954]. Blocking [1032, 1751]. blocks [3540]. Blockunderrelaxation [522, 523]. Blockwise [573]. Blood [2441]. Blossoming [4320, 2607]. Blow [2388, 2549]. Blow-up [2388, 2549]. Bluetooth [3727, 3396]. Blur [4267].

Bodies [855]. Body [3206, 3924]. Bóhling [743]. Böhme [216]. Boltjanskij [271, 187]. Boltzmann [1767, 2397, 1693, 2970, 1603]. Bonus [3592]. Book [89, 304, 246, 335, 372, 353, 354, 374, 375, 391, 392, 420, 421, 419, 422, 542, 543, 544, 539, 545, 673, 814, 898, 44, 72, 45, 178, 143, 184, 171, 220, 232, 134, 234, 268, 602, 509, 492, 490, 792, 817, 826, 173, 14, 533, 531, 330, 483, 73, 416, 331, 185, 603, 105, 70, 87, 26, 67, 68, 221, 225, 27, 16, 35, 28, 102, 41, 69, 186, 179, 200, 238, 169, 285, 286, 289, 271, 319, 325, 429, 418, 426, 427, 551, 577, 579, 580, 646, 639, 12, 156, 388, 485, 484]. Book [578, 675, 674, 36, 652, 811, 85, 15, 74, 46, 71, 147, 126, 201, 148, 137, 181, 149, 249, 261, 262, 536, 424, 377, 60, 65, 138, 337, 226, 118, 145, 283, 727, 732, 305, 99, 215, 235, 302, 347, 349, 114, 214, 180, 218, 140, 211, 355, 350, 360, 358, 514, 430, 512, 515, 756, 757, 698, 157, 122, 654, 264, 299, 328, 361, 575, 659, 799, 195, 222, 431, 486, 90, 100, 336, 144, 139, 120, 401, 119, 641, 650, 796, 193, 224, 230, 151, 323, 281, 291, 357, 648, 649, 797, 406]. Book [815, 10, 121, 329, 552, 301, 236, 13, 97, 101, 103, 88, 94, 92, 64, 11, 194, 153, 174, 290, 279, 303, 327, 287, 362, 378, 488, 623, 627, 576, 549, 625, 547, 701, 700, 656, 146, 626, 581, 670, 671, 729, 743, 746, 42, 190, 191, 197, 288, 812, 75, 66, 104, 425, 744, 794, 333, 321, 253, 532, 198, 188, 628, 513, 98, 540, 34, 39, 237, 702, 685, 487, 530, 534, 582, 657, 651, 644, 653, 816, 338, 29, 30, 37, 95, 31, 33, 32, 38, 43, 61, 229, 332, 263, 356, 363, 677, 116, 642, 647]. Book [829, 832, 745, 192, 96, 141, 154, 389, 481, 510, 172, 251, 390, 63, 47, 86, 40, 117, 216, 227, 155, 219, 212, 152, 257, 265, 699, 809, 818, 124, 176, 203, 170, 187, 205, 142, 217, 196, 260, 266, 267, 254, 91, 93, 125, 135, 177, 127, 150, 199, 204, 213, 233, 228, 324, 247, 269, 270, 320, 258, 280, 300, 334, 351, 373, 402, 352, 404, 405, 376, 604, 537, 546, 643, 759, 758, 813, 827, 800, 801, 793, 62, 282, 248, 252, 250, 256, 676, 182, 115, 183, 231, 136, 284, 322, 403, 482]. Book [428, 462, 489, 491, 493, 535, 645, 802, 810, 831, 795, 538, 833, 359, 423, 417, 798, 828, 202, 326, 348, 548, 624, 550, 541, 672, 733, 629, 686, 728, 730, 830, 511, 640, 658, 655, 123, 175, 158, 223, 731, 884, 259]. Bool [255]. Boolean [327, 1494, 1139, 2687, 928, 1562, 839]. Boolesche [1139]. Boolescher [839]. Booth [299]. Bootstrap [2236]. bootstrapping [3969]. Bopp [260]. Border [1434, 3253]. bordered [1672]. Bork [603]. Borsky [322]. Borwein [2903, 4384]. both [3623]. botnets [3660]. Bottleneck [1281, 1047, 1450, 2186, 2121, 1117]. bottom [2780, 3321]. bottom-up [2780, 3321]. Bound [2590, 4344, 162, 591, 2072, 2559, 1816, 3188, 3417, 1742, 1375, 1317, 1672]. Boundaries [4110, 2070, 1942]. Boundary [2055, 2173, 2233, 1484, 1840, 1934, 1300, 1397, 4337, 2567, 2253, 1808, 2294, 2376, 1254, 1588, 1834, 2264, 4205, 760, 1045, 2362, 1788, 619, 2566, 2103, 2368, 2438, 2129, 1573, 2370, 2394, 722, 947, 1999, 1465, 4359, 2435, 1315, 2291, 1290, 1174, 750, 849, 994, 2522, 1841, 1926, 2588, 1267, 2416, 2339, 2553, 4170, 295, 2465, 2356, 2197, 1536, 2367, 4171, 275, 753, 1502, 297, 1161, 1242, 1436, 1615, 1324, 2141, 2254, 1227, 2669, 1364, 995, 3859, 1224, 2649, 1013, 2663, 1609, 1742, 1762, 2791, 2732, 2735, 384, 1725, 2682, 2809, 1162, 2655, 410, 1572, 2787]. boundary [1107, 3476, 1755, 1156, 1741, 689, 2171, 1483, 1193, 1611]. Boundary-element [2566]. Bounded [2112, 2590, 1947, 2538, 900, 920, 1288, 2195, 1898, 681, 1216, 2657, 2823]. Bounded-Time [1898]. Bounding [4332, 1251, 1798, 2454, 2479, 1479, 1695, 1690, 1704, 2652, 470]. Bounds [2592, 2357, 1868, 2110, 4351, 1610, 466, 2405, 760, 1680, 4100, 1860, 4356, 1573, 1867, 2282, 1185, 1853, 442, 1174, 4154, 2022, 242, 464, 1926, 1009, 4120, 4164, 860, 839, 1151, 164, 2482, 2316, 275, 2171, 273, 721, 1558, 2671, 2509, 634, 2679, 704, 2606, 1924, 2624, 1735, 2019, 369, 1762, 4232, 379, 470, 1740, 2714, 1162, 1100, 1978, 1571, 1483, 1936, 773]. Box [4336, 2275, 2372, 2591, 2049, 2225, 1715, 3190, 1996]. Box-Constrained [4336]. Box-Methoden [1996]. Box-Scheme [2372, 2591]. Box-Splitting [2049]. boxes [2812, 1276]. BP [3234]. BPMN [3713]. BPX [1987]. BPX-preconditioner [1987]. Bracken [388]. Bracketed [955, 1352]. Bracketing [599, 366, 1059, 458]. Bradley [402]. BRADO [3836]. Brains [170]. Brammen [1943]. Branch [2072, 1816, 1375, 591, 591]. branch-and-bound [591]. Branch-and-Bound-Verfahren-Generator [591]. Branching [4206]. Brauer [549]. Brayton [2888]. Brazilian [2939]. breakdown [2828]. Breaking [1142]. Breakthrough [1265]. Bregman [2636]. Briatenberg [793]. Bridge [2243, 2936]. Bridge-Connected [2243]. Brief [4243, 3115]. Brillhart [1362, 1362]. Bringing [3357]. broadcast [3677, 4085, 3753, 3438]. Broadcasting [2177]. broker [2854, 3589, 3278]. brokers [3124]. Brosowski [701]. brother [1148]. Brouwer [868]. Brouwerschen [868]. Brown [826, 1699]. Brownian [1449]. Brownlee [137]. browser [3841, 2884]. browsers [3265]. Broyden [2642, 1892, 1230]. Brusberg [85]. BSP [3877]. Buchbesprechungen [884]. Buchta [65]. Budak [42]. budget [3872, 3612]. budget-deadline [3612]. budgeted [3614]. buffer [3968, 412]. buffer-tape [412]. Bühler [321]. build [3746, 3589]. Building [4280, 3062]. Built [1057, 1654]. Built-In [1057, 1654]. Bukovics [856]. Bulging [1943]. Bulk [3331]. bulk-insertion [3331]. Bulk-loading [3331]. bump [2603]. bump-surfaces [2603]. Bundesrepublik [85]. Buneman [4236]. Burdick [228].

Bürger [551]. Burgers [1431]. burning [4050]. Büro [252]. bursts [2756]. bus [2001]. Busacker [323]. Buses [2137]. Business [3955, 238, 3464, 3161, 3087, 3088, 3089, 3684, 3520, 3391, 3109, 3241, 3346, 3270, 3434, 3012]. business- [3161]. Business-to-Business [3012]. Busy [2105]. Butcher [819]. Butzer [66, 390]. BVPs [1694]. Byzantine [3263, 3882].

C [372, 374, 391, 419, 220, 792, 731, 674, 337, 350, 575, 230, 357, 88, 532, 251, 177, 247, 269, 604, 831, 1969, 3495, 2815, 2814, 1640, 4262, 4014]. C-section [4014]. C-support [3495]. C-Tree [4262]. C-XSC [2815]. C. [89, 268, 347, 816, 745, 699, 351, 640]. C4.5 [4033]. Cache [2028, 2306, 3430, 2967, 2737, 3277]. Cache-Aware [2306]. cache-based [3430]. cache-enabled [3277]. cache-oblivious [2737]. caching [2624, 3994, 2967, 3277]. Cactus [2206, 2629]. CAD [4132]. CAGD [2456, 2050]. calcium [3290]. Calculate [2227, 1733]. Calculateurs [320]. Calculating [2766, 2624, 922, 1003, 1479]. Calculation [1522, 1868, 1569, 1521, 4279, 466, 1365, 724, 1844, 1519, 1350, 162, 666, 522, 523, 1297, 1179, 969, 1566, 340, 2855, 1754]. Calculus [225, 1140, 231, 2068, 2150, 3870, 827]. calibrated [4028]. Calibration [4269, 3348, 2872]. Calligaert [86]. camera [3943, 3348]. Can [3838, 2175, 2818, 1992]. Cancellation [2347]. Canonical [2012, 3357, 1969]. Cantor [2191]. Cantor-Zassenhaus [2191]. Capable [723]. capacitance [1067]. capacities [3359]. Capacity [1105, 1979, 3690, 3028]. capillary [1001]. CAPRA [3051]. captioning [3958, 4091]. capture [3619]. Carbon [4048, 3536, 2848]. cardiac [2883, 3290]. cardinality [2665]. care [3730, 2934, 2892, 3829]. Carlo [1915, 1915, 1693, 2079, 2237]. Carmichael [1143]. carry [2813]. carry-save [2813]. Cascadic [2399, 2304]. Case [933, 2559, 2224, 1929, 2109, 2170, 1458, 3626, 2729, 1924, 1019, 4093, 1325, 3541, 4230, 1476, 3503, 2016, 3548, 1002, 1736, 3232]. Cases [542, 2066, 3239]. cash [3548]. Caslin [220, 88, 269, 604]. Cast [1943]. Cat [3953]. Catastrophe [1505]. Catastrophic [2347]. Categorical [3875, 3551]. categories [3967]. categorization [3661, 3875]. caterpillars [2623]. Catfish [3121]. Catfish-PSO [3121]. Cauchy [1296, 907, 517, 1963, 918, 1150, 1662, 1771, 1271, 1238, 1085, 1790, 2093, 1351, 1392, 2486]. Cauchy-Problem [907, 907]. Cauchy-type [1238]. causal [3676, 3112]. CDF [2404]. Cebysev [910]. Cebysev-Entwicklungen [910]. Celestial [301]. CELF [3892]. cell [2883, 2827]. cells [1313, 2862]. Cellular [1897, 2744, 4137, 4138]. center [2723, 3103, 3888, 4036]. Centered [1514, 1489, 1581, 1455, 3014, 1308, 1190, 1292, 1298, 1372, 1490]. Centers [1773, 3070, 3536, 2939, 3197, 3441, 3928, 3165, 3085, 3719, 4074, 3459, 3125, 3518, 3191, 3429]. central [2750]. centrality [3758, 3889, 3921]. centric [3086, 3087, 3571, 3088, 3089, 2976, 3007, 3345, 3617, 3124, 3864, 3434, 3609, 3139]. centroid [3532]. Centrosymmetric [1912]. CEP [3849]. Certain [4332, 1963, 954, 476, 2355, 2543, 2359, 1527, 2223, 782, 1909, 2527, 517, 2752, 1366, 1222, 2700, 396, 1171, 584, 497, 4176, 1750, 2005, 1978]. certification [3013, 167, 345, 8, 58, 168]. CFR [3624]. CFS [3231]. CG [1707, 2108]. CG-method [1707]. Ch. [372]. chain [3915, 4075, 3539]. chained [3183, 3475]. Chains [2582, 4342, 2770, 2766, 962]. challenges [4006, 2917, 3506, 3733, 3833, 3732, 3936]. championship [3660]. Chance [197]. change [3327]. changes [3755, 3754, 2899, 3226, 2945, 4137, 807]. changing [3343, 2870]. channel [3664, 3309, 2879, 3850]. channels [3438]. chaos [3911]. Chaotic [1874, 3486]. Character [1586, 4141]. Characteristic [1506, 2336, 1638, 1510, 1175, 1494, 2733, 2795, 2616]. Characteristics [1254, 2319, 1061, 2645, 3143, 2014, 3470]. characteristics-based [3470]. Characterization [768, 4212, 1726, 3055, 1907, 339, 1402, 711]. Characterized [1146]. Characterizing [3407, 1195]. Charakterisierung [339, 1402, 711]. Charakteristikenverfahren [1254]. Charge [1825, 3592]. charging [3773, 3846, 3535, 3807, 3475]. Charkewitsch [348]. Chase [3260]. chatbot [3488, 3583]. Chawla [2633]. Chebotarev [301]. Chebyshev [370, 52, 1836, 618, 843, 1304, 504, 382, 693, 910, 768, 1427, 905, 991, 110]. Chebyshev-approximation [504]. Check [1586, 3585]. check-in [3585]. Checking [3676, 3466, 3711, 3841, 3090, 3465]. checkpoint [2916]. Checkpoints [2479]. cheetah [4039]. Chemical [283, 150, 587]. chemischen [587]. cherry [3250]. chest [3737]. Chevyshev [766]. Chiang [640]. Chiffrierverfahren [1347]. Chinal [828]. Chinese [3378]. Chintschin [173]. chip [3034, 3037, 3810, 3039]. Choice [1595, 475, 1397, 683, 764, 883, 821, 1665, 1058]. Cholbd [8]. Cholesky [1212, 7]. Chomsky [342]. Chomsky-Sprachen [342]. choosing [1564]. Chorafas [349]. Chord [1058]. Chordal [2073]. choreographies [3091]. choreography [3783, 3713]. choreography-based [3783]. Christoffel [265]. Chromatic [874]. chromosome [3479, 3632]. Chronos [4064]. Chu [228]. churn [3597, 3235]. CIEDE2000 [3249]. Cimmino [1472]. Cimmino-method [1472]. ciphers [2690, 2694]. Circle [1531, 2560, 1396, 943, 1372, 2697]. Circles [852, 2133, 1094]. Circuit [278, 513, 3021].

Circuits [422, 539, 928, 133, 406].

Circulant [1255, 2030, 2159]. Circular [4293, 1680, 1268, 4108, 1155, 1443, 1702, 1298]. Citation [3906]. Cities [3137, 3447, 3473, 3974, 3636, 3424, 3396, 3472, 3501, 3517]. citizen [3585]. citizens [3424]. Clamped [2227]. Clarence [492]. clarity [3932]. Clason [639]. Class [2270, 2561, 2366, 4356, 2345, 2249, 2000, 1965, 567, 1166, 927, 863, 2121, 2052, 1164, 2160, 1504, 1413, 2474, 2034, 2320, 1727, 517, 4222, 2731, 1745, 3891, 1090, 3918, 3982, 4045, 1344, 1579, 1576, 2818, 1208, 473, 594, 612, 3185, 1737, 1703]. Classes [4332, 208, 1588, 751, 2396, 2965, 1318, 2063, 873]. Classical [3763, 2071, 678, 1763, 384, 3683, 364, 467, 888]. classical-quantum [3683]. Classification [711, 3618, 3405, 3804, 3591, 3931, 3927, 3495, 3948, 3143, 3284, 4043, 3554, 3793, 3568, 3975, 3843, 3282, 3639, 3759, 3865, 2904, 3885, 3624, 3386, 3235, 3819, 3979, 3378, 3231, 3934, 3762]. classifier [3230, 2777, 3836, 3735, 3292]. classify [3225]. Claus [698, 671]. clause [525]. Clenshaw [395, 1471, 667, 395]. client [3342, 3882]. client-side [3342]. clients [3099, 3637]. climate [3226, 3604]. climatic [3226]. clinical [4082, 4014]. Clique [1112, 2003, 3998]. Cliques [1864]. Closed [4204, 4144, 4206, 2319, 1885, 1644, 3355, 1682]. closeness [3758]. Closure [897, 892, 3852, 1383]. Cloud [3070, 3481, 3244, 3093, 3788, 3419, 3948, 3109, 3761, 3888, 3182, 3864, 3245, 3176, 3625, 2759, 3606, 3074, 3911, 3641, 3430, 3887, 3026, 3966, 3355, 3530, 3199, 3272, 3615, 3432, 2891, 3621, 3024, 3422, 3657, 3945, 3066, 4027, 4022, 3712, 2757, 3575, 3923, 3441, 3928, 4000, 3610, 3726, 3204, 3351, 3113, 3394, 3104, 3192, 3980, 3933, 3255, 3399, 3697, 3118, 3727, 3907, 3029, 3119, 3550, 3832, 3995, 3687, 3504, 3390, 3822, 3122, 4060, 3964, 3334, 3818, 3969, 3977, 4024, 4069, 4072, 3085, 3156, 3592, 3667, 3774, 3756, 3719, 3998, 4033, 3193, 3439, 4059, 4074, 3708, 3229, 3589, 3120, 3862, 3718, 3717]. cloud [3950, 3735, 4010, 3368, 3121, 3124, 3459, 3080, 3433, 3999, 3125, 3991, 3126, 3071, 3855, 3856, 3518, 3075, 3794, 3362, 3149, 3342, 3668, 3191, 3760, 3633, 3759, 3123, 3635, 3147, 3942, 3553, 3790, 3947, 2901, 3077, 3562, 3367, 3612, 3190, 3986, 3222, 3640, 2760, 3140, 2975, 3561, 3886, 4064, 3095, 3073, 3076, 3268, 3423, 3722, 3869, 3707, 3602, 3308, 3139, 4029, 2913, 4087, 3787, 2911, 4089, 3729, 3826]. cloud-aware [3255]. Cloud-backed [3606]. cloud-based [3657]. cloud-edge [3561]. cloud-fog [3964, 3942]. cloud-hosted [3342]. cloud-P2P [3075]. cloud/fog/IoT [3826]. cloudlet [3354]. cloudlets [3155]. Clouds [4323, 3094, 3132, 4012, 2854, 3066, 3000, 3645, 3096, 4023, 3027]. clubs [2921]. Cluley [350]. Cluster [3265, 4230, 4057, 4030, 951, 3962, 2971]. Cluster-based [3265]. clustered [3347, 2969, 3274, 3671]. Clustering [2472, 4322, 2296, 3214, 3365, 3324, 3661, 3788, 3928, 3549, 3854, 2922, 3307, 3427, 4086, 3175, 3794, 3692, 3551, 2754, 3776, 3819, 3482, 3558, 742]. clustering-based [3307]. clusters [3741, 3110, 3804]. Clusterwise [1593, 1257, 1341, 1144]. CM [4240]. CM-2 [4240]. CMAC [3217]. CMCSF [3560]. CNN [3526]. CNN-GRU [3526]. CNNs [3843]. co [3918, 2926, 2932]. co-design [2926]. co-occurring [3918]. co-scheduling [2932]. Coalescence [2581]. Coarse [2042]. Coarse-Grain [2042]. Coarsening [2566, 2627]. COBOL [425]. COBOL-Praxis [425]. code [3207, 3918, 2868, 2964, 3620, 1678, 2611]. Coded [2061, 1974, 3479]. Codes [304, 1529, 1460, 3086, 1093, 3184, 1382, 996, 3017, 1078, 363, 543]. Coding [940, 120, 2219, 1973, 3130, 3184, 3344, 1691]. Coefficients [2426, 4097, 443, 2405, 1491, 366, 922, 838, 908, 782, 4327, 2553, 1606, 1018, 1350, 2627, 1990, 973]. coercive [2696]. coevolving [3332]. cognition [3606]. cognitive [3688, 3374, 3626, 3665, 3383]. cognitive-assisted [3374]. coincidentally [3239]. cold [3961, 4007, 3412]. collaborating [3391]. collaboration [3410, 2757, 532, 4067, 3602]. Collaborative [3915, 3514, 3108, 2959, 3689, 3897, 3560, 3862, 3917, 3849, 3329, 3864, 3624, 3491, 4011, 3074]. Collaterals [4132]. Collatz [439, 418, 324, 610]. Collection [42, 402, 3353, 3471, 3129, 2920, 3216]. collective [4058]. Collectively [1167]. Collinear [2270, 2409, 2020]. Collins [215]. Collision [2463, 3150]. Collocation [1934, 2408, 2545, 2544, 2012, 902, 2010, 2295, 1314, 1465, 2406, 2485, 753, 1149, 1436, 1157, 1099, 1684, 1209, 2791, 1332]. Collocation-Type [753]. Colloquium [202]. Colony [3159, 3334, 3914, 3667, 3525, 4005, 3707]. Colorations [940]. Coloring [1653, 2410, 4196, 1007]. colors [2623]. Column [244, 1470]. Columns [1595]. Combination [1230, 669, 2673]. Combinations [2309, 2326, 1660]. Combinatorial [738, 261, 61, 1319, 1594]. combinatorics [3929]. Combined [3530, 1310, 3484, 2987, 3135, 3235]. Combining [3762, 3266, 3722]. Command [881]. Comment [438]. Comments [2380, 1548, 1136, 1884]. commercial [3026]. Common [2308, 4279, 942, 507]. Communication [4194, 2551, 3324, 3643, 2959, 3662, 3702, 3806, 2933, 3924, 3218, 3242, 3038, 3136, 2960, 3476, 3620, 3480, 4071, 2955, 3102, 3609, 3952, 2150, 2324, 2325]. communication-based [3609]. Communications [830, 3387, 3665, 416]. communities [3032, 2998, 3632]. Community [3600, 3714, 3959, 3852, 3895, 3168, 3588, 3598, 3446, 3033, 3409, 3631, 3837, 3831, 3705, 3330, 3389, 3097, 3757]. Compact [1167, 2633, 1561]. compactly [1990]. Company [204]. comparability [1007]. Comparative [420, 3358, 3148, 2939, 3880, 1382, 3847]. compare [2810]. Comparing [2883, 2996, 2980]. Comparison [2173, 2215, 2257, 2488, 2469, 2597, 3443, 1764, 1929, 1394, 3310, 3066, 2819, 3788, 1743, 2176, 1474, 2652, 1343, 2879, 1362, 861]. Comparisons [1309, 437, 1705, 1977]. compensation [2802]. competing [3313]. competition [3237]. Competitive [3324, 3322, 3932]. Compilation [2403]. compiler [3006, 3177]. Complementarity [2417, 2671, 1689, 2598]. Complementary [1868]. Complements [2314, 2505]. Complete [2541, 2518, 604, 692, 298]. Completed [1188]. Completely [162, 409, 433]. Completeness [939, 1461]. completing [608]. completion [3850]. Complex [1378, 1591, 4244, 4183, 2081, 505, 618, 2261, 366, 1680, 565, 852, 4340, 2383, 1186, 1519, 2365, 1268, 1550, 1809, 435, 4328, 739, 524, 269, 604, 862, 4108, 1283, 2731, 2734, 3042, 2893, 3015, 3901, 3103, 480, 1155, 3852, 3823, 3032, 4184, 4043, 473, 3189, 3847, 272, 1350, 1181, 1702, 3632, 1127, 1298, 1372, 1739, 3766, 3705, 3312, 3899, 3800, 3871, 1963]. complexes [78]. COMPLEXIS [3410]. Complexity [1177, 2552, 2278, 2138, 4194, 2391, 781, 1581, 839, 747, 2224, 1766, 1168, 1283, 3188, 3410, 3859, 1007, 3316, 3844, 1412, 2835, 1490, 1048, 962, 564, 668, 3637, 1079]. compliant [3697, 3262]. component [3682, 3004, 3662, 2899, 3567, 3096, 3080, 3047, 2981]. component-based [3682, 3096, 2981]. Components [1723, 2243, 469, 524, 4010]. componentwise [1558, 1978, 1049, 1703]. Composing [3528]. Composite [1389, 2233, 2118, 2553, 2037, 2197, 3988, 4162, 1077, 3462, 2697, 2719]. Composition [4210, 1888, 3272, 3009, 3910, 3399, 3907, 3687, 1560, 3669, 4063, 3050, 3971, 3436]. Compositional [2576]. compositions [2912, 3713]. Compound [2023, 2174, 2093, 2217, 1991, 1204]. compound-Poisson [1204]. Comprehensive [67, 4090, 4037, 3165, 3051, 4052]. compressible [2866, 2870]. Compression [2101, 2259, 2451, 4274, 3680, 3763, 506, 3196, 3801, 3386]. compression-based [3680]. Computability [729]. Computable [2405, 2361, 1978]. Computation [823, 1648, 1957, 616, 2561, 1028, 1634, 4186, 1258, 2570, 1814, 2137, 1433, 380, 21, 1440, 4353, 1928, 240, 2537, 2464, 2355, 1673, 387, 4102, 4103, 2090, 4232, 2441, 2369, 1527, 1914, 436, 806, 990, 2473, 1427, 2471, 1288, 481, 2165, 1862, 196, 125, 322, 802, 678, 2486, 1216, 4368, 1907, 3614, 1821, 1717, 598, 4370, 4371, 979, 632, 2199, 2819, 1210, 2797, 2643, 3930, 4229, 3500, 1214, 4373, 1354, 3844, 960, 2040, 3049, 1402, 2841, 1236, 3258, 4119, 1317, 1070, 389, 1737, 1002, 2754, 1741, 3874, 1483, 842, 1563, 1369, 908, 444, 863, 86, 535]. Computational [354, 375, 943, 1287, 1783, 2690, 1450, 1773, 1946, 516, 2554, 75, 804, 1420, 1816, 2186, 1607, 150, 759, 892, 4374, 1008, 1168, 620, 3043, 3042, 1564, 1363, 3238, 3354, 3224, 2886, 2759, 1692, 3212, 3574, 651, 747]. Computations [4265, 4105, 4187, 2302, 862, 4123, 1123, 2815, 2742, 2632, 1630, 157]. Compute [890, 710, 2683, 1712, 3359, 3969, 1687]. computed [2221]. Computer [788, 1233, 246, 420, 419, 544, 539, 4109, 1129, 483, 1187, 4370, 4371, 289, 4153, 2576, 1183, 377, 1189, 569, 401, 193, 329, 4366, 4104, 532, 1841, 2225, 4272, 745, 3250, 1637, 91, 204, 213, 228, 247, 352, 259, 478, 830, 4152, 4368, 456, 1025, 3755, 3754, 1249, 966, 1093, 51, 397, 2983, 4378, 243, 3742, 463, 229, 2798, 4238, 4138, 4109, 230, 4366, 1932, 980, 548]. Computer-Aided [539, 478].

Computer-Arithmetik [4109, 4366]. Computer-Assisted [1841, 1932]. Computer-Communications [830]. Computer-Geometrie [980]. Computer-gesteuerter [569]. computer-nets [4138]. computer-oriented [4368, 463]. computer-programmed [569]. Computer-Technik [548]. Computergestutzte [1932]. Computerized [723]. Computers [372, 1512, 45, 891, 551, 639, 749, 311, 2177, 4178, 2028, 360, 397, 103, 623, 4105, 437, 233, 2580, 681, 620, 6, 949, 272, 615, 89, 180, 355, 350, 120, 513, 685, 96, 322, 493, 84]. Computersystems [603]. Computing [1441, 2592, 372, 1129, 1316, 4350, 2734, 2828, 1846, 2243, 2578, 1172, 12, 4144, 1468, 4315, 2196, 4290, 167, 1188, 2188, 716, 4355, 1419, 2176, 1867, 1870, 281, 1999, 2505, 2265, 2422, 4148, 4150, 920, 1906, 2325, 4141, 4149, 1845, 1243, 1523, 1626, 1863, 2184, 1136, 1781, 3779, 2266, 2885, 1813, 1257, 2665, 367, 561, 718, 2316, 1510, 871, 3235, 1399, 3580, 1714, 4056, 3573, 3688, 3655, 3641, 2826, 3392, 3857, 3162, 3987, 3199, 3763, 2653, 2654, 3615, 3866, 3170, 3398, 3220, 1494, 4019, 3915, 3506, 3643, 3024, 3940, 3945, 4022, 4094, 1540, 586]. computing [638, 3891, 3901, 1237, 1481, 3961, 3113, 3104, 3980, 3933, 3953, 3983, 3118, 3706, 2968, 3727, 4007, 3029, 3765, 3119, 3550, 1753, 3832, 3930, 2911, 1669, 3994, 2675, 2983, 3993, 3085, 3156, 3560, 3592, 3756, 3858, 4033, 3542, 3193, 4066, 3522, 4041, 3467, 3229, 2832, 3307, 3303, 1422, 3566, 2694, 3718, 3717, 3950, 2966, 2622, 4055, 3917, 3849, 3696, 3003, 3182, 1054, 3321, 3138, 3314, 3794, 592, 1636, 3633, 2846, 3123, 3690, 3635, 3553, 3790, 3947, 2901, 3415, 3879, 1439, 3640, 2759, 2796, 4239, 2975, 3561, 3805, 3768, 3839, 3820, 4064, 4018, 1082, 4067, 2950, 2971, 3869, 4061, 3707, 3295, 3487]. computing [3308, 2829, 3261, 3501, 3152, 763, 2911, 530, 437, 645, 3361, 3674, 3114]. computing-assisted [3688]. Concave [1444, 2412, 1463, 1633, 2007, 1743, 2176]. Concentrated [2394]. concentration [3234]. Concept [1203, 4253, 343, 589, 4107, 78, 4221, 3245, 3436, 2603]. Concepts [1795, 2097, 876, 2867, 3936, 372]. conceptual [3492]. Concerning [4351, 621, 607, 1805, 662]. Concurrency [1442]. concurrent [3743, 3499, 3259]. Condition [438, 4340, 722, 2534, 2314, 2395, 2193, 822, 1455, 3743, 1981, 585, 1706, 500, 1672]. Conditional [378, 456, 3340]. Conditionally [1647]. Conditioned [1029]. Conditioning [1488, 1035]. Conditions [1484, 1826, 891, 2565, 2294, 619, 2438, 1573, 878, 2242, 2132, 2416, 2263, 2197, 1147, 3664, 3134, 1224, 1762, 2732, 1445]. Conductivity [2522]. Cone [4293, 1444, 2638]. Conference [387, 318, 59]. conferencing [3079]. confidence [3744, 3417, 4234, 3476]. configuration [3814, 2984]. Conflict [2027, 3988]. Conformal [2663, 1396, 112, 23]. conformance [3466, 3090, 3465]. Conforming [2349, 2594]. confused [3322]. congested [3082]. Congestion [3569, 1032, 4046, 3974, 3460, 2955]. Congress [83, 455, 9, 387, 237]. Congruence [210]. Congruences [4301]. Congruential [2174, 1991]. Conic [2059]. Conjugate [1051, 780, 1121, 1823, 1120, 864, 1984]. Connected [1723, 2243, 2177, 210, 740, 3657, 2652, 617, 3512, 3652, 1103, 3513, 112, 4365]. Connection [1772, 1196]. Connections [1472]. connectivity [3529, 3845]. Conquer [1903]. Conrol [259]. Consensus [3579, 3753, 3323]. Conservation [4102, 3125]. Conservative [1224]. Conserving [2756, 3335]. Consideration [1055, 3370]. considering [3523, 2936, 3304, 3433, 3176, 3957]. considers [2826]. Consistency [2321, 1388, 1084, 3520, 3676]. Consistent [907, 1540, 3434]. consolidation [3923, 3697, 3818, 3085, 3888, 3991, 3100]. consortium [4075]. Constant [1856, 4200, 838, 1419, 908, 2418, 2001, 4344, 1824]. Constant-Radius [4200]. Constant-time [2001]. Constants [851]. constellations [4044]. constituents [43]. Constrained [4336, 1910, 2590, 4250, 2011, 2148, 2297, 2033, 1381, 4347, 4259, 1902, 1218, 2058, 2773, 3600, 4087, 1743, 2176, 3307, 1729, 2846, 3283, 2786, 2589]. Constraint [3711, 4325, 3448, 1969, 3580, 3037, 3872, 2820, 2747, 3444]. Constraint-Based [4325, 3711]. Constraints [1500, 2369, 1922, 1097, 1202, 3110, 1123, 1650, 558, 3761, 3, 3368, 594, 612, 2637, 3028, 3612, 4088]. constriction [3236]. Constructing [2765, 2729, 2378, 2298, 2672, 1847]. Construction [4110, 2445, 4313, 872, 2421, 4356, 400, 2641, 1950, 1888, 4365, 636, 4108, 1817, 2311, 3384, 2248, 2599, 2640, 1676, 3723, 1900, 397, 2966, 1162, 1686, 3076, 3198, 984]. constructions [434]. Constructive [2375, 1985, 485]. Constructor [4274, 4062]. consumer [3399, 3124]. consumer-centric [3124]. consumers [3322]. consumption [3641, 3101, 3145, 2933, 3347, 3163, 3719, 3304, 3341, 3248, 3191, 2901, 3604, 3396]. Contact [1997, 855, 4359, 2696]. contactless [3731]. container [4090, 4037, 4041]. containerized [4053]. Containing [2138]. Containment [2587]. Contemporary [1258, 3630]. content [3571, 3930, 4075]. content-centric [3571]. contention [3562]. contents [3277]. Context [1133, 3545, 941, 726, 3011, 1119, 3501, 3043, 3272, 3533, 3461, 3643, 4028, 3895, 312, 3548, 3720, 3233]. Context-aware [3545, 3043, 3272, 3533, 3461, 4028]. Context-based [3011]. Context-Free [941, 1119, 726]. Context-Sensitive [1133, 312]. Contextual [3973, 3296]. Conti [35]. continous [3959]. Continuation [2475, 2573, 2487, 933, 2190, 1817, 2668, 1752]. Continued [1360, 1135, 1338]. Continuity [4220, 1445]. Continuous [2062, 1386, 4207, 3787, 1815, 3521, 1785, 4260, 696, 2930, 4025, 2850, 409, 433, 2908, 4133, 2823, 474, 3784, 1279, 3694, 1967]. continuous-time [2823]. Continuously [1943]. continuum [3605, 3855, 3856]. Contour [4304, 272, 1056]. Contraction [4289, 295]. Contractive [1929]. Contractivity [1413]. contracts [3931]. contrast [3290]. contrastive [4029]. Contrasts [2591]. contribution [3517, 1026, 112]. Contributions [622, 1183, 211, 3042, 43]. Control [465, 2094, 2, 483, 1145, 225, 289, 427, 639, 2576, 149, 337, 467, 888, 1491, 2166, 2164, 2449, 2454, 1442, 900, 4257, 866, 651, 1791, 4106, 553, 266, 2060, 183, 952, 2589, 2963, 3306, 3780, 2664, 3335, 3571, 364, 2861, 1349, 2712, 2674, 3156, 3460, 497, 2824, 2823, 2646, 3778, 4040, 3809, 1280, 3604, 2650, 2726, 2955, 3027, 4002, 4001, 3609, 3955, 169, 377, 138, 378, 802]. control-volume [2646]. Controllability [1783]. Controllable [2380, 2443]. Controlled [2261, 1105, 3490, 1033, 2698, 3493]. controllers [3442]. controlling [2710, 3007, 4136]. Convection [2341, 2096, 2146, 2235, 2337, 2372, 2338, 2319, 2340, 2029, 2291, 2418, 2349, 2339, 2474, 2733, 2595, 2682, 2775, 2787, 2646, 2764]. Convection-Diffusion [2341, 2146, 2235, 2337, 2372, 2319, 2340, 2029, 2349, 2339, 2474, 2418, 2733, 2595, 2775, 2787, 2646, 2764]. Convection-Dominated [2096, 2291, 2682]. Convergence [736, 1282, 108, 2270, 1795, 1477, 1150, 1930, 1789, 4299, 1819, 536, 1254, 779, 2491, 1874, 1353, 1090, 4226, 4354, 2144, 924, 2493, 1348, 2340, 4358, 447, 2208, 682, 2332, 1948, 910, 1368, 915, 916, 2274, 2318, 1986, 2132, 163, 1537, 1164, 2263, 1920, 2646, 2579, 606, 1977, 1270, 2528, 1466, 783, 407, 688, 2350, 4348, 1380, 1388, 307, 1110, 1655, 1727, 2752, 1532, 1662, 1533, 607, 1325, 1093, 1403, 1704, 1085, 1308, 1405, 1095, 1115, 445, 2016, 1685, 1845, 2183, 1728, 296, 1640, 1597, 1710, 1813, 1677, 1248, 1392, 2614, 1219, 1601, 1235, 766, 563, 1438]. Convergence [384, 1011, 1053, 1107]. Convergent [823, 4338, 887, 1874, 1429, 1174, 2528, 1245, 1374, 1422, 2879, 1726]. convergents [1545]. Converging [893, 1275]. conversation [3634]. converse [1493]. Conversion [81, 82]. Conversions [4295, 1182]. Converting [4312]. Convex [1441, 766, 906, 2270, 52, 889, 1518, 2412, 2033, 4360, 879, 1192, 1920, 1921, 1953, 2458, 636, 1969, 783, 1264, 1060, 1650, 2143, 2882, 1636, 1873, 1439, 1235]. Convex-Concave [2412]. Convex-Decomposable [4360, 2143]. convexe [1953]. Convexification [2538]. Convexity [674, 1559, 1202]. conveyor [2876]. Convolution [1993, 1959, 2617, 1687, 3997]. Convolutional [3634, 3741, 3750, 3958, 4091, 3865, 3850, 3285, 3231, 3385, 3796]. convolutions [2612]. Cooley [538]. Cooling [1604]. Cooperative [3277, 2953, 2931, 4069, 4003, 4057, 3236]. Coordinate [760, 1325]. coordinated [2946]. Coordinates [2537, 4214, 3029, 1180]. copy [3980, 3933]. copycat [3527]. coral [3530]. Corank [1872, 1907, 1717]. Corank- [1872]. core [2914, 4076, 3007, 3001, 3258, 3003, 3815, 4054]. core/periphery [3815]. coreference [3148, 3720]. coreness [3823]. Corner [2382]. Corners [498, 1301, 1126]. corona [2909, 3300]. coronary [3290]. coronavirus [3629]. Corporate [3840]. Correct [1870, 3239]. corrected [2850]. Correcting [543]. Correction [317, 2146, 4156, 958, 3755, 4090, 2107, 1421, 1451, 3980, 3271, 4165, 4091, 883, 3740, 3718, 3856, 4367, 1257, 4372, 1073, 1364, 4157, 4161, 4163, 2910, 3002, 3842, 1244, 4158, 1160, 3722, 601]. Corrections [2244, 2513, 4166, 1615, 4167]. correctness [3680, 3743]. Corrector [310, 2906]. correlate [3299]. correlated [3528, 3106]. Correlation [3803, 3944, 3215, 2947, 583, 3126, 3330, 3420]. correlations [3145, 3248]. Correlators [2352]. correspondence [1148]. Corresponding [908]. Corrigendum [1813]. cosine [3876, 3998]. Cost [1178, 1355, 754, 4010, 3368, 1782, 4282, 1008, 1552, 3066, 3788, 3370, 3111, 4059, 3888, 3519, 3942, 3886]. Cost-driven [3368]. cost-informed [3111]. Cost-Optimal [1178, 1552]. costs [3887]. cosymmetrical [4353]. Cotes [1101, 2527, 2697]. counselor [3011]. count [3417, 3829]. count-min [3417]. counting [1980, 3036, 3318]. Coupled [784, 619, 2144, 2301, 2162, 2478, 3598, 2778, 1334, 1367, 2784, 2114, 3808, 2868, 2595, 1261, 2889, 3091]. coupled-seeds [3598]. Coupling [3629, 2254, 2253, 1832, 2141, 2856, 4081, 3185, 2741, 2742]. Courant [219]. Courier [3990]. Course [40, 3215]. CovaDel [3653]. coverage [3275, 3525, 2844]. Covering [835, 2593, 2632]. COVID [3730, 3647, 3734, 3733, 3735, 3736, 3732, 4082, 3737, 3653, 3638]. COVID-19 [3730, 3647, 3734, 3733, 3735, 3736, 3732, 4082, 3737, 3638]. COVID-like [3653]. COVID-shield [3638]. Covolume [2437]. Cowan [196]. CP [3308]. CPS [2931, 3155]. CPU [4079, 3006, 3443, 2951, 3722]. CPU-intensive [2951]. CPU/GPU [3006]. CPUs [3463]. Crank [2474]. crash [3579, 3753, 3257]. crash-prone [3257]. crash-recovery [3579]. Crawler [3244]. CRCW [2600]. create [3225]. Creating [386, 3749]. creation [3954, 2798]. Credible [1898]. credit [2833, 3624]. Crete [3138]. crisis [3699]. crisis-related [3699].

Criteria [1595, 2439, 2409, 2448, 3655, 3502, 2998, 3232, 982]. Criterion [2577, 2483, 294, 306, 309, 1075, 3253, 2717, 554, 1806, 1076]. Critical [4177, 3296, 1032, 3134, 3673]. criticality [3611]. crop [3934]. crops [3226]. Cross [4205, 2305, 3148, 3339, 3547, 3228, 2717, 3812, 3169, 3309, 3796]. cross-device [3339]. cross-document [3148]. cross-domain [3547]. cross-language [3169]. cross-layer [3812, 3309]. cross-linguistic [3796]. cross-section [2717]. cross-topic [3228]. Crossing [2541, 1439]. Crouzeix [2722]. Crow [3639, 3887]. crowdfunded [3546]. Crowdsourcing [3269, 3932]. Crowns [1599]. cryptography [2691, 3977, 3999]. cryptosystems [1347]. crystal [2908]. CSAMOMC [3887]. CT [3737, 3290]. Cubature [1668, 1195, 2547, 1817, 1763, 2248, 1676, 1402, 2135, 1686]. Cubatures [2010]. Cube [721, 3418]. Cubes [2381]. Cubic [959, 1025, 2544, 1771, 1618, 2174, 1866, 1353, 2451, 722, 2406, 2104, 1792, 1836, 2605, 1202, 946, 2021, 1806, 1682, 2614]. cuckoo [3950]. CUDA [3392, 3173, 3877]. cultural [2904]. Cumulative [1178, 4227]. curation [3455]. curb [3733]. currency [3926]. current [2943, 3506, 3828, 2701, 2741, 2742, 2759]. Curtis [1471, 667, 395]. Curvature [4336, 1518, 2450, 4209, 1419, 4260, 4365, 2710]. Curvature-Continuous [4260]. Curve [2449, 2712, 946, 2058]. Curved [1944]. Curves [268, 958, 4245, 4298, 4314, 2448, 4203, 4300, 4249, 1940, 240, 2368, 2454, 4256, 4257, 2189, 2457, 4261, 4275, 2572, 1916, 4311, 1025, 1003, 2606, 1395, 1711, 1243, 1682, 2719]. Curvilinear [2112, 1787, 1918]. custom [3339]. Customer [3597, 3210, 3051, 3975, 3235, 3762]. customers [2739]. Customized [3759, 3123]. Cut [2009]. Cuts [2089]. Cutting [904, 1652]. cyber [3654, 3775, 2929, 3155, 3116, 2928, 1592]. cyber-physical [3654, 3775, 3116]. Cybernetic [328]. Cybernetics [375]. Cycle [4083, 790, 2036, 3976, 1447, 1412, 1272]. Cycles [869, 741, 737]. Cyclic [911, 1044, 4342, 1077]. Cyclide [4293]. Cycling [1992, 3408]. Cylindrical [4324, 2871, 2877].

D

[184, 87, 215, 235, 302, 641, 13, 3038, 196, 228, 643, 759, 758, 359, 1969, 1778, 2838, 1434, 2235, 2696, 2851, 4280, 2660, 1790, 2677, 2454, 2394, 4323, 2441, 2783, 2466, 2722, 4364, 3843, 2229, 2884, 2741, 2742, 2878, 1942, 2356]. D-Discontinuity [2229]. D-grid [1778]. D. [353, 354, 48, 87, 27, 147, 126, 349, 514, 13, 97, 101, 290, 63, 373, 813, 827, 645, 359, 624]. D2HCP [2984]. Daas [4063]. DAEs [1945]. DAG [3940]. Dagstuhl [4380, 2601, 2706, 4378, 4381, 4377]. Dahl [792]. Dahlquist [970]. DAIP [3786]. Dale [302]. damage [3291]. damit [4369]. Damped [1965, 2036]. Dämpfung [56, 1036]. Damping [1036, 56]. Daniel [3451]. Darstellung [786, 2012, 893, 43, 998, 958, 1395, 236]. Data [2417, 1464, 1512, 3127, 1723, 2101, 551, 639, 2490, 3326, 3353, 1437, 2569, 3880, 1265, 2364, 2256, 2390, 575, 4172, 3401, 552, 2374, 2097, 3372, 3512, 3652, 3960, 2035, 2053, 4213, 3917, 2326, 3404, 4217, 2229, 4218, 116, 4329, 2547, 742, 4185, 1608, 1131, 3440, 3400, 3528, 3763, 3070, 3682, 3264, 3616, 3618, 3536, 598, 2943, 2939, 3397, 1675, 3024, 3422, 3913, 3197, 1557, 3657, 2842, 3547, 3471, 3087, 1981, 3088, 3909, 3403, 3103, 3067, 3343, 2673, 3570, 3859, 3441, 3928, 2944, 3165, 3417, 3089, 3982, 3196, 3691, 3130, 3529, 3577, 3723, 3129, 2677, 2969, 4070, 3296, 3695, 3143]. data [3337, 2990, 3345, 3811, 3360, 3595, 3242, 2920, 3048, 2978, 3085, 3156, 3284, 3841, 3719, 2972, 3215, 3704, 4003, 3846, 4074, 3708, 2922, 1516, 3497, 2694, 3534, 3226, 3888, 4086, 3793, 2715, 4055, 3325, 3997, 3382, 3663, 3363, 3136, 3459, 3999, 2831, 3125, 3019, 3845, 3412, 3071, 3195, 3416, 3622, 3518, 3443, 3824, 43, 2902, 3825, 3489, 3191, 3455, 4082, 3544, 3209, 3245, 3332, 4036, 3142, 3879, 3624, 3450, 3386, 3424, 2954, 3025, 3214, 3297, 3429, 3381, 3798, 3819, 4075, 3091, 1407, 3280, 3978, 3851, 3073, 3211, 3418, 3383, 3829, 3807, 2942, 3076, 3426, 3385, 3922, 3430, 3410, 3419, 94, 3548, 402, 537]. data-centric [3088, 3089]. Data-Dependent [4329]. Data-driven [3404, 3091, 3426]. data-intensive [3325]. Data-sparse [2390]. Database [2484]. databases [3422, 3748, 1079, 3357, 3676]. datacenter [3575]. datacenters [3448]. Dataflow [2212]. dataflows [3746]. dataset [3691, 4035]. datasets [3745, 2949, 3982, 3695, 3258]. date [550]. Datenbank [814]. Datenbanken [1079]. Datenerfassung [699]. Datenfernverarbeitung [796]. Datenmodell [1265]. Datenmodelle [1265]. Datenstrukturen [598]. Datenübertragung [424]. Datenverabeitung [201, 545, 14]. Datenverarbeitung [114, 282, 248, 255, 643, 624]. Datenverarbeitungsanlagen [105, 102, 235]. Datenverarbeitungssystemen [417]. Datenverarbeitungstechniken [248]. Davidon [783, 2020, 783]. Davidon-Fletcher-Powell-Method [783]. Davidon-Fletcher-Powell-Verfahrens [783]. Davidson [372]. day [3880]. day-ahead [3880]. DBSO [4025]. DC [3021, 3350, 3738]. DC-Bagging [3350, 3738]. DCNN [4016]. DCT [3414]. DDPG [4008]. dead [1973]. deadline [4027, 3822, 3368, 3612]. deadline-aware [3822]. deadline-sensitive [4027].

Deadlines [1881, 1042, 3708]. Deadlock [3499, 3027]. deadlocks [3593]. Dean [70]. debugging [3711, 826]. Decaying [2154]. Decentralized [3035, 2922, 3809, 3780, 2839, 3571, 4043, 3826, 3882]. deception [3831]. deceptive [3379]. Decidability [2529, 312, 729]. deciding [943]. Decimal [1512]. Decision [2966, 3580, 3992, 3199, 3230, 3502, 4060, 3371, 2794, 3146, 3759, 3535, 967, 4014, 3139, 4089]. decision-centric [3139]. decisions [3243, 2833, 3415]. declarative [2834]. Decomposable [4360, 2143, 1246]. Decomposing [2042]. Decomposition [2501, 2504, 2231, 2376, 2535, 1429, 1878, 1877, 1337, 2495, 1879, 2288, 1823, 2241, 2036, 2214, 2165, 2075, 2740, 4161, 2875, 2888, 1725, 2625, 2793, 1206, 2786]. Decomposition-based [2501]. Decompositions [2289, 1516, 984, 365]. dedicated [3623]. deduce [694]. deduction [3154]. Deep [4019, 3591, 3650, 3937, 3288, 3934, 3369, 3966, 3834, 4025, 3893, 3399, 3828, 3832, 3375, 3958, 4091, 4024, 3411, 3774, 3944, 3996, 3628, 3970, 3770, 3975, 3737, 3377, 4008, 3380, 3509, 3373, 3978, 3376, 3383, 3860, 3799, 3231, 4014, 3762]. DeepWalk [3557]. Defect [2146, 4156, 4166, 4372, 1599, 2107, 1421, 4157, 1451, 1615, 1073, 1364, 3178, 4161, 4163, 4158, 3692, 1160]. Defect-Correction [2146]. Deferred [4167]. deficient [3802, 1476]. Definable [876]. defined [2648, 3478, 3495, 3569, 1711, 3967, 1243, 3078, 3268]. definierbare [876]. Definite [2578, 7, 2323, 867, 3152, 2806, 2861, 1672]. definiten [867]. definiteness [1471, 1627]. definiter [7]. Definition [665, 834, 3436, 273]. definitions [3492]. Deformation [4218]. Degenerate [1830, 1684]. Degradation [1941, 1941]. Degraded [4267]. Degree [917, 4298, 872, 4319, 857, 522, 4311, 2853, 1763, 2248, 2729, 2643, 1060, 3601, 3068, 2135, 2718]. Dehib [485]. Dekomposition [744]. Delaunay [2640, 1870, 3874]. Delay [1967, 2091, 2353, 3614, 2436, 3786, 3356, 2947, 1359, 3460, 3563, 2661, 1678, 2748]. delay-efficient [3786]. delay-integro-differential [2748]. delay-sensitive [3460]. Delayed [1898]. Delays [2295, 3440, 3929]. delivering [3071]. delivery [4063, 3653]. demand [3773, 3281, 3967, 3433, 3535, 3140, 3829]. demands [3275]. Demidowitsch [509]. Demo [2949]. Dempster [2821]. dengue [3724]. deniable [2988]. Denmead [402]. Denoising [2538, 2507, 2195, 3722]. Dense [2042, 2126, 1699, 1620, 3083]. Densities [1417]. Density [2317, 1116, 3559, 1463, 1071, 3705, 3819]. density-based [3559]. DEPAS [2839]. Dependability [3013, 3201]. Dependable [3031, 3779, 2939, 3261]. dependence [3394]. Dependencies [1858, 3834, 2727, 3025]. dependency [4010, 2971]. Dependent [1936, 2376, 1775, 2407, 1966, 4329, 2151, 2611, 2746, 3842, 2682, 2741, 2742, 2764]. Depending [1626, 1243]. Deployment [3387, 2893, 4019, 2688, 3970, 3951, 3793, 4018]. deployments [3362]. depth [1327, 1293]. deque [3260]. deren [1]. Derivation [1612, 3172, 20, 1096, 1822, 413]. Derivation-free [413]. Derivative [894, 2781, 2894, 2265, 2422, 1775, 2514, 2168, 2032, 2895, 780, 164]. Derivative-Free [2265, 2422, 780, 164]. Derivatives [1355, 2196, 4205, 567, 555, 2195, 1240, 2486, 2633, 1097, 1438, 1271, 1746, 296, 1528]. Derived [477, 2532]. Deriving [874, 1772]. Descent [2383, 3843, 1002]. Describe [4343, 4138]. describing [1318]. Description [1723, 3252, 3073]. descriptive [980]. Design [89, 539, 2352, 4247, 4250, 355, 4301, 2449, 4208, 4302, 4127, 2976, 4305, 406, 4255, 513, 3683, 251, 4273, 266, 3036, 3815, 828, 133, 3026, 3680, 1025, 3326, 2915, 2952, 3037, 3654, 4093, 2611, 3896, 3371, 3072, 2694, 2926, 3793, 3182, 3321, 3393, 2798, 2736, 421, 103]. designed [3178]. Designing [3548, 3042, 4168]. designs [3453, 697]. Desirable [718]. dessen [694]. details [3177, 243]. Detect [3299, 4281, 3338, 737]. Detecting [2899, 1908, 2989, 741, 2997, 3032, 2876, 3632, 3593, 3660]. Detection [1112, 3828, 3349, 3534, 2229, 4282, 2463, 3821, 3559, 3588, 3369, 3598, 3634, 3608, 3618, 4016, 3861, 3600, 3714, 3581, 3893, 3446, 3984, 3239, 3343, 3570, 3726, 3931, 3959, 3691, 3852, 4045, 3150, 3630, 3409, 3495, 3938, 3631, 3988, 3895, 3379, 2998, 3251, 3735, 3926, 3314, 3023, 3705, 3330, 3389, 3516, 3414, 3415, 3791, 3694, 3798, 3725, 4009, 3882, 4047, 3168, 3420, 3989, 3919, 3617]. Detector [3567, 2014]. Detector-based [3567]. detectors [3817]. Determinants [2355, 1366]. Determination [1434, 1444, 1812, 780, 914, 76, 705, 1889, 370, 1857, 1549, 848, 298, 163, 707, 903, 1259, 1191, 1296, 1547, 470, 1091, 1716, 4237, 293, 208, 616, 684, 859, 53]. Determining [1411, 1425, 1480, 1931, 1313, 396, 1343, 1422, 1412, 1553, 1911, 1235]. Deterministic [2857, 1767, 1979, 3022, 3005]. DetLock [3005]. Deutsche [33]. Deutschland [85]. developers [3099]. Developing [4231, 3976, 2832]. Development [2446, 2923, 2991, 4084, 3266, 3011, 2980, 70, 893]. Deviations [1593]. Device [2100, 2934, 3339, 3893, 2985, 3251]. Devices [435, 3846, 3973, 3513, 2987, 3128, 3868, 2951, 1322, 3396]. Dew [3688]. DG [2865, 2870]. diabetics [3230]. diagnosing [3230, 590]. Diagnosis [1898, 3817, 3587, 3349, 3404, 3737, 3561]. Diagnostische [590]. diagonal [703]. diagonaldominanten [108]. Diagonalization [2203]. Diagonally [108, 2324, 1325]. Diagonally-Dominant [108]. diagonals [1272]. diagrams [2838].

Diagraphen [741]. Diameter [1781]. diamond [2848]. diarrhea [3688]. Dictionary [551, 639, 2082]. did [3465]. dielectrics [969]. Diemer [282]. Dienstplanerstellung [400]. dietary [3902]. Dieter [1129]. Difference [1464, 1856, 2146, 2092, 2575, 1506, 784, 1834, 2118, 748, 4320, 619, 1624, 761, 2340, 2321, 912, 2332, 1887, 1000, 1406, 1234, 231, 1135, 4171, 1445, 2633, 989, 1822, 1224, 506, 1080, 2750, 2595, 1200, 571, 445, 4176, 1244, 4168, 1279, 1563, 984, 3777]. difference-schemes [445]. difference/finite [2750]. Differences [2226, 231, 4331, 2619, 975]. differencing [3214]. Different [2851, 2155, 1832, 2819, 2879, 3176]. Differentiability [1826, 770]. differentiable [4340]. Differential [2426, 4110, 2532, 1167, 1389, 1934, 2092, 1141, 2561, 2, 4201, 774, 35, 2102, 2536, 1768, 1051, 4178, 1830, 2180, 2440, 2112, 2155, 2091, 779, 888, 129, 207, 218, 1451, 748, 1043, 619, 2427, 2481, 2353, 755, 838, 100, 2065, 641, 2336, 1624, 2295, 1776, 2310, 1927, 11, 1775, 1501, 854, 1887, 1932, 750, 849, 2172, 444, 2022, 1120, 851, 1391, 1000, 39, 707, 1534, 1480, 2239, 2315, 2396, 135, 449, 759, 877, 767, 798, 1466, 789, 1168, 3486, 2442, 1589, 297, 2785, 2782, 78, 1761, 2555, 1175, 1149, 1436, 2634, 1667, 2436, 2690, 1720, 949]. differential [988, 1492, 1073, 4157, 975, 986, 585, 1263, 1108, 3119, 1085, 1700, 1762, 24, 1398, 2750, 1709, 3358, 1359, 1983, 450, 571, 1225, 4176, 1266, 410, 1716, 1194, 614, 527, 1553, 2139, 2661, 976, 1125, 1339, 1358, 1279, 1160, 1083, 1226, 1678, 446, 984, 3799, 2748]. Differential- [218]. Differential-Algebraic [1934, 1775, 1501, 2172]. Differential-Difference [1624]. Differential-Functional [2092]. Differential-operators [444]. Differentialgleichung [78, 446]. Differentialgleichungen [297, 1993, 949, 888, 988, 1492, 975, 986, 619, 838, 119, 1932, 1225, 1000, 614, 758, 877, 1125, 1339, 984, 1934, 429, 1994, 779, 129, 207, 1043, 24, 450, 571, 849, 410, 527, 707, 767, 182, 1226]. Differentialgleichungssysteme [976, 1358]. Differentialgleichungssystemen [585, 854, 851, 789]. Differentially [3913]. differentiated [3509, 3952]. Differentiation [1464, 1546, 4352, 1256, 611, 2808, 1711, 1983, 4119, 611]. Differenzengleichungen [482]. Differenzenschemata [619]. Differenzenverfahren [445, 1000, 984, 912]. Differenzierbarkeitsbegriffe [770]. Difficult [1106]. Diffusion [2341, 2146, 2235, 2433, 2337, 4338, 2372, 1834, 2525, 2319, 761, 2340, 2029, 2288, 2349, 2371, 2318, 2339, 2474, 4277, 3802, 2639, 969, 2733, 1224, 2746, 2418, 2478, 2595, 2775, 2480, 2626, 2787, 2646, 3477, 2890, 2764]. diffusion-tensor [2890]. diffusive [2837]. Digital [4178, 949, 355, 240, 2303, 230, 1638, 188, 513, 116, 437, 125, 233, 828, 3014, 2640, 4045, 1095, 2979, 3605, 3936, 4075, 541, 80]. Digital- [949]. Digitale [799, 672, 756, 240, 80]. digitalen [234]. digitaler [74]. Digitalrechner [356, 546, 334]. Digitalrechnern [437]. Digits [997, 73]. Digraph [897, 737, 1103]. Digraphen [737]. digraphs [741]. Dijkstra [792]. dilation [1990]. DIMA [4053]. Dimension [674, 2429, 711, 2683, 3639, 4236]. Dimensional [4110, 2375, 1122, 1825, 2372, 1771, 1722, 1400, 2292, 1834, 1643, 2496, 2287, 1666, 761, 2314, 2291, 1538, 1581, 2339, 1820, 217, 742, 4278, 1278, 2170, 634, 2743, 2504, 2909, 2853, 2591, 632, 4224, 2019, 1045, 2542, 2543, 3029, 3723, 965, 1665, 613, 2616, 1351, 1845, 18, 1056, 414, 3865, 371, 1847, 3978, 2603, 1714, 574]. dimensionaler [18, 414]. dimensionality [3763]. Dimensionen [680]. Dimensions [2366, 2228, 411, 2549, 2306, 2395, 1855, 3026, 680, 2649, 2063, 2178]. Diner [576]. Dinkelbach [511]. Diophantine [962]. Direct [2173, 926, 411, 239, 2297, 2381, 516, 2394, 2240, 561, 493, 1270, 2461, 1907, 2806, 3479, 1693, 4158, 432]. direct-coded [3479]. directed [293, 710, 1206, 3771]. Direction [1923, 3833, 3576]. Directional [894, 1933, 3958, 4091, 2844, 3937, 3989]. directionality [2829]. Directions [4336, 864, 3506, 3583, 3828, 3589]. directives [3006]. Directory [544]. Direkte [432]. direkten [239]. Dirichlet [4224, 411, 1878, 1877, 1661, 2663, 1762, 1302, 2522, 4214, 2416, 767]. Dirichlet-boundary [1762]. Dirichlet-Problem [767]. Dirichletproblems [767]. Disaggregation [4342]. disambiguation [3395]. disaster [3387, 3180, 3923]. disaster-recovery-as-a-service [3180]. discipline [1132, 3891]. Disciplines [2136]. disclosure [3840]. discomfort [3408]. disconnected [2652]. Discontinuity [2229]. Discontinuous [443, 1927, 2553, 789, 2860, 1738, 43]. discount [3889]. discovered [3955]. Discovering [3215, 3219, 3226]. discovery [3878, 3910, 3473, 2992, 2947, 3030, 3568, 3382, 3510, 3065, 3762, 3955, 3244]. Discrepancy [4315, 2185, 2557, 2316, 1629, 1982]. Discrete [1794, 2062, 1300, 4287, 1227, 1312, 2427, 2549, 1277, 1309, 2685, 700, 682, 804, 2240, 4135, 813, 4331, 2170, 1138, 1287, 2825, 1089, 1633, 3830, 3359, 2710, 2007, 2701, 382, 2641, 1095, 2017, 2738, 3766, 414, 3865, 3874, 1012, 407, 2736]. discrete-time [3874]. discretisation [450]. Discretisations [2262, 2387, 2482]. Discretization [2112, 2118, 207, 1887, 2312, 798, 1502, 1694, 1778, 1044, 2753, 2700, 2630, 999]. Discretizations [2096, 1826, 2504, 2301, 2314, 2571, 767, 2436, 2667, 2910, 3002]. Discretized [1524, 1348, 1720]. discriminant [2892]. discriminative [2990]. Discussion [2216, 84]. disease [3250]. diseases [3934]. Disjoint [3758, 3656]. Disk [1087, 2028, 1088, 2756, 2804, 3145, 2803]. diskontinuierlichen [43]. diskrete [4135]. diskretem [682]. diskreten [382, 628]. diskreter [1089]. Diskretisierung [975, 571]. Diskretisierungen [767]. Diskretisierungsfehler [207]. Diskretisierungsverfahren [450]. Disks [57]. Diskussionsitzung [84]. Dispatch [1939, 3135]. Dispatching [808]. dispersed [3765]. dispersion [3401]. Dispersive [1406]. disruption [3571]. Dissipative [1994, 2022, 1726]. dissipativer [1994]. Distance [2707, 4266, 2454, 440, 3968, 2819, 2657, 3370, 3189, 3249, 3245, 3694, 1439, 3273]. distance-based [3694]. distance-hereditary [2657]. Distances [1811, 710, 4228]. distant [963]. distinguishing [3210]. Distributed [738, 2256, 2836, 3849, 3065, 2273, 1637, 4056, 4062, 3133, 3413, 3054, 3057, 3108, 3750, 3197, 1557, 2893, 3471, 3817, 3319, 3804, 3783, 4017, 3130, 2992, 664, 3344, 3337, 1693, 3896, 3082, 4033, 3542, 3740, 3739, 3862, 2835, 3556, 3362, 2902, 3593, 3438, 3155, 2984, 3576, 3486, 3076, 2994, 3676, 2981, 3302, 3115]. distributing [3993]. Distribution [2167, 2236, 821, 1387, 2583, 1159, 789, 1734, 3915, 2853, 966, 2647, 316, 2876, 583, 2848, 3442, 3404, 637, 4136, 3879, 3047, 556]. Distributionen [789]. Distributions [788, 2062, 1496, 1619, 1216, 1233, 1644, 1129, 1633, 2007, 2972, 592]. Distributive [2477, 1231, 1176, 1408]. Distributivgesetze [1176]. Distributivität [408]. Distributivity [408]. disturbances [2823]. Dittrich [743]. div [2358]. div-stable [2358]. divergence [1861, 2641, 2667, 384]. divergence-free [2641, 2667]. Divergenz [384]. Divergenzverhalten [1861]. diverse [3266].

Diversity [3479, 4039, 3364, 3312, 3757]. Divide [1903]. Divide-and-Conquer [1903]. Divided [4320, 2619]. Dividend [2386]. division [3365, 1313, 2041, 714]. DNA [3999]. DNN [3685]. do [3161]. Docker [3951, 3793]. Docker-based [3793]. Document [3228, 3148, 2956]. documents [3123, 3158]. does [3364]. Domain [2504, 1925, 1901, 2376, 2261, 1878, 1877, 2495, 1126, 2875, 2520, 2288, 1823, 1725, 2241, 2214, 3547, 4161, 2676, 3439, 3099, 2625, 2632, 1718, 3952]. Domains [498, 1301, 4151, 2112, 2359, 2533, 1151, 2580, 2070, 453, 2680]. Dominant [108, 2324, 1325]. Dominated [2096, 2291, 2682]. domination [2657]. donkey [3845]. Doppelschichtpotential [1107]. Doppelschichtpotential-Ansatzes [1107]. Dorn [281, 13]. Dörr [686]. Dot [2024, 2768]. Dot-Product [2024]. Dotzauer [809]. Double [1591, 1558, 2743, 2672, 1760, 1108, 1107]. double-ended [2672]. double-periodic [1760]. Doubling [1486]. Doubly [932]. Douglas [279]. Down [1142]. downtime [3590]. DP [2644]. dqds [2286]. Dr [605]. Draft [4104]. dragonfly [3135]. Draper [174]. drawing [1016]. dreidimensionalen [965]. Dreieckskonstruktionen [434]. Dreiecksund [222]. Drenick [325]. drift [969]. drifted [3997]. Driven [2569, 1042, 3785, 4045, 3636, 3360, 3951, 3368, 2830, 2831, 3416, 3404, 3868, 3690, 3638, 3424, 3091, 3426, 3472, 2878, 3493]. drivers [3626]. Drives [3331, 3145, 2803]. driving [3607]. DRL [3893]. drone [3653]. drones [3629]. drop [2909]. dropping [3136]. drug [4068]. Drums [1153]. DSOL [2834]. DTN [3673]. Dual [1751, 2453, 2282, 4289, 2539, 3627, 3145, 1636]. dual-block [3145]. Dualität [879]. Duality [1289, 879, 2223]. Due [1290]. Duffin [391]. dummy [3659]. dumping [615]. Dumps [615]. dune [2720, 2654, 2753]. dune-subgrid [2720]. Dunngitterprobleme [1961]. duopoly [3237]. Dupin [4293]. duplicate [3023]. Durand [176, 1671]. Duration [1940]. durch [159, 52, 855, 1940, 4101, 1545, 606, 935, 1234, 953, 1916]. Durchbruch [1265]. Durchführbarkeit [4097]. Durchschnittsproblem [1923]. DV [749]. DV-Anlagen [749]. DVFS [3172, 3538, 3007, 3181]. Dworatschek [756]. DX9 [2489]. Dynamic [3615, 107, 4183, 2569, 3538, 1881, 2378, 3884, 700, 3603, 3313, 4193, 2284, 1205, 3839, 2389, 3425, 3941, 3426, 4089, 3070, 3627, 3223, 3523, 3804, 3923, 4085, 3752, 3697, 4087, 3924, 3296, 2925, 4184, 3964, 4069, 4072, 4086, 3173, 4053, 3155, 3116, 3252, 2984, 3613, 3905, 2981, 3237]. Dynamical [1146, 2825, 2221, 1716]. Dynamically [4262, 2870, 3433, 2923]. dynamically-changing [2870]. dynamics [3658, 3847, 4030, 3574]. Dynamik [60]. Dynamische [319, 794, 232, 426, 577]. dynamischen [657]. Dynamischer [831]. dynamisehen [194]. dynamizing [1246].

e-banking [3160]. E-D2HCP [2984].

e-KYC [3789]. e-Learning

[2757, 3081, 3285]. E. [483, 551, 181, 536, 226, 302, 757, 222, 581, 1136, 29, 43, 61, 25, 227, 809, 176, 402, 643, 417, 658]. EAC [3529]. Each [1823]. earlier [3439]. Early [4273, 3382]. easing [3004]. eavesdropping [3485]. Ebene [1062]. EC [2440]. ECC [3643, 3635, 3805]. ECC-based [3643, 3635, 3805]. ECG [3349, 3865]. echo [4071, 3142]. Echtzeit [1938]. Echtzeit-Optimierung [1938]. Ecken [1126]. ECMA [622]. eco [3248]. eco-assistant [3248]. Econometric [706, 690]. Economic [148, 185]. economic- [185]. economical [1274, 3119]. Economics [426, 4383, 581, 540, 530, 534]. ecosystem [3572, 3011]. ecosystems [3014]. ECP [3307]. ED [3880]. Eddison [48]. eddy [2701, 2970, 2741, 2742]. Edge [3481, 3506, 3716, 2290, 803, 3951, 3696, 2009, 3314, 4282, 3879, 2462, 4067, 3580, 3573, 2723, 3478, 3763, 3940, 3686, 3891, 3782, 4000, 3983, 3930, 4070, 3994, 3388, 4024, 3993, 3560, 3858, 3068, 4041, 3534, 3862, 3917, 3663, 3855, 3856, 3312, 3760, 3690, 3561, 3805, 3768, 3839, 3820, 4018, 4061, 3295, 3487, 3602]. Edge-based [3314]. edge-cloud [4000]. Edge-driven [3951]. edge-fog-cloud [3760]. Edge-Partitioning-Oriented [803]. Edges [2533, 2621, 3652]. Editing [2457]. editor [2895, 2896]. Editorial [2509, 2962, 3543, 3092, 3315, 2425, 2285, 2897, 2730, 3397, 3010, 2724, 2755, 3114, 2444, 2601, 2706, 2684, 2983, 3508, 2805, 3381, 3674, 3024, 3087, 3209, 2993]. Editors [48, 238]. Education [1145, 3503, 3833, 3383]. educational [3431, 3284, 3496]. EDV [958, 642]. EDV-Anlagen [958]. EDV-Organisation [642]. EEG [3804]. EELB [4060]. EELRP [3578]. Effect [2076, 997, 3055, 2221, 4234, 3916, 3874, 3237]. Effective [3365, 542, 4351, 4228, 1749, 3282, 2343, 3239, 3904, 2920, 3996, 3586, 2932, 3027, 3232, 1345, 2728, 3715]. effectively [3440]. effectiveness [2855]. effects [3262]. Effertz [60]. efficacy [3238]. Efficiency [1849, 3000, 3928, 3924, 3930, 3964, 3756, 4059, 1362, 3171, 2975]. Efficient [4004, 2403, 2642, 4384, 1456, 2914, 2180, 2440, 621, 870, 1588, 1797, 2698, 3473, 2355, 2581, 2103, 1798, 1860, 2032, 3150, 3370, 3872, 3300, 2394, 3344, 3938, 4024, 1635, 3428, 3005, 4213, 4187, 932, 2884, 4284, 3986, 2463, 3561, 4264, 2477, 132, 3685, 3655, 2953, 1233, 3641, 3552, 3355, 4046, 2903, 3536, 1688, 2883, 2952, 3174, 3223, 3786, 1556, 3773, 3804, 3403, 3923, 3359, 3610, 3931, 3578, 3104, 3706, 2645, 3823, 2663, 1579, 3822, 2985, 1560, 1375, 4060, 3595, 3818, 2978, 3085, 3156, 3603, 3669, 2616, 3215, 3642, 3257, 3340, 4051, 2655, 2005, 3442, 2995, 3459, 2967, 2951, 3100]. efficient [2596, 4065, 3800, 3671, 3972, 3769, 3612, 3604, 3235, 3318, 1670, 3102, 3429, 3820, 4075, 4047, 3807, 4061, 3308, 1086, 1027]. efficiently [1698]. Effiziente [1027]. Effizienzvergleich [1362]. effort [3898]. Ehlers [745]. Eichmeier [330]. Eigenfunction [2521]. eigenfunctions [704]. Eigenpairs [2383]. Eigenproblem [1517, 4159, 1739]. eigenproblems [2863]. Eigenschaften [4103]. Eigensolver [2274]. Eigenvalue [2047, 2083, 1546, 1903, 2267, 505, 2294, 468, 1828, 4340, 2383, 2242, 2466, 2268, 1391, 553, 2197, 1416, 661, 1558, 4175, 1760, 2861, 2865, 369, 1709, 470, 1244, 1067, 381, 1426, 2859, 679, 1630, 1457, 1984]. Eigenvalues [1868, 1936, 1569, 914, 4103, 442, 682, 1912, 1894, 1895, 1297, 2170, 4175, 704, 409, 433, 2643, 393, 2014, 1740, 4176, 1507, 340, 1584, 679, 2171]. eigenvector [2715, 1457]. Eigenvectors [469, 914, 442, 1059, 632, 393]. Eigenvektoren [632, 393, 442]. Eigenvektors [469, 1059]. Eigenverktoren [914]. Eigenwert [468]. Eigenwertaufgaben [369]. Eigenwerte [1936, 409, 433, 393, 442, 682, 470, 679]. Eigenwerten [914]. Eigenwertproblemen [553]. Eigenwertprolemen [381]. Eignung [614]. Eilon [26]. ein- [779]. Ein/Ausgabe [952]. Einbettung [695, 567]. Einbettungsalgorithmen [933]. Eindeutigkeitsaussagen [78]. Eindeutigkeitsfragen [504]. einem [692, 466, 607, 239, 1126, 499, 895, 791, 930, 1356]. einfach [112]. einfache [159, 609]. einfachen [162]. Einführende [252, 236]. Einführung [545, 232, 602, 14, 221, 579, 60, 515, 1545, 303, 581, 321, 657, 653, 95, 263, 829, 141, 510, 548, 730, 201, 4128]. Einführungskursus [426]. Eingabelung [470]. Eingrenzung [503, 606]. Einheiten [863]. Einhullende [1916]. einige [964, 436, 343, 54, 1939, 1269, 879, 1884]. einiger [1347, 21, 1949, 495]. eins [456]. Einsartz [397]. Einsatz [417, 702]. einschließ [1955]. einschließlich [162]. Einschliessung [1037, 599, 366, 935, 1059, 409, 433, 469, 777, 1515, 1234, 458]. Einschließungassätze [446]. Einschließungssatz [468, 662]. Einschliessungsverfahren [160, 1038]. Einschrittverfahren [1125, 1452]. einseitige [987]. einseitigen [693]. Einsicht [186, 286]. einspringenden [1126]. Einzeldarstellungen [362, 488, 389, 481]. Einzelschrittverfahren [1018, 1075, 1098]. Ekman [177]. Elastic [855, 2531, 3680, 2866, 2967]. Elasticity [2401, 2300, 3166, 3096]. elastischer [855]. elastoplastic [2631, 3772]. elastostatics [2696, 2783]. elderly [2892]. electric [3247, 3404]. Electrical [211, 217, 2888, 112, 4036]. Electricity [3886]. Electricity-cost-aware [3886]. Electromagnetic [2881, 142, 2771]. Electromagnetism [2567]. Electronic [89, 1825, 350, 188, 322, 3213, 2769, 243, 272, 4001]. Electrostatic [1825]. elektrischen [112]. Elektrodynamik [260]. Elektronenrechners [33]. Elektronik [330]. Elektronische [102, 98, 256, 14, 349, 282]. elektronischen [615]. elektronischer [200, 97, 272]. Element [2055, 2401, 1886, 2439, 443, 498, 2094, 1301, 2313, 4246, 1803, 2434, 1825, 2576, 2294, 2292, 1956, 2563, 2415, 2205, 2506, 1013, 2129, 2336, 1573, 1604, 1666, 2301, 2387, 2370, 901, 2534, 2124, 1465, 2435, 2097, 1432, 2064, 2373, 2395, 2424, 4344, 2358, 2579, 2571, 2302, 2580, 2356, 1502, 2594, 2849, 2669, 2660, 2860, 2566, 2863, 1661, 2795, 2701, 3695, 3437, 3772, 2874, 1660, 2656, 2704, 2750, 2732, 2140, 2722, 2910, 3002, 2533, 2655, 2646, 3352, 2632, 2764]. element-based [3695]. Element-Boundary [2370]. elemental [3437, 3254]. Elementare [185, 470, 998]. elementarer [21]. Elementary [2335, 21, 99, 4118, 2744, 4122, 185, 1585, 1687, 2040, 470, 998]. Elemente [194, 935, 956, 1038, 4367]. Elementen [994, 1063]. Elements [2204, 2586, 2399, 2253, 71, 2103, 139, 2290, 1277, 1972, 2468, 2349, 75, 994, 2553, 935, 956, 4367, 2428, 1063, 2907, 1037, 2594, 2850, 2141, 2254, 2674, 2615, 1162, 1987, 1986, 2787, 1038, 3352, 335]. elevating [3770]. eleventh [237]. Elimination [1434, 683, 4144, 621, 2297, 1154, 1495, 2324, 2562, 1880, 1497, 1022, 1571, 630]. elite [3627, 3876]. ellipsoid [4234]. Ellipsoids [2463]. Elliptic [2426, 109, 2096, 2233, 2497, 443, 220, 2504, 1301, 2433, 2099, 1051, 2591, 2112, 2118, 2496, 2226, 1878, 2382, 901, 2554, 88, 2291, 2309, 2030, 2159, 1872, 1174, 849, 2268, 2510, 1120, 1926, 2437, 2553, 2428, 2304, 269, 604, 2580, 767, 2197, 2111, 1502, 2773, 1242, 2642, 1019, 1364, 1760, 975, 986, 2660, 2865, 2621, 1013, 1661, 947, 1762, 4232, 2674, 1200, 450, 571, 1725, 4233, 1572, 2627, 2859, 1755, 2650, 2726, 1741]. Ellipticity [2591]. elliptische [1013]. elliptischen [986, 450]. elliptischer [975, 571, 849, 767]. Elsevier [639]. embedded [2918, 3018, 3035, 4005, 3084]. Embedding [1187, 567, 2400, 4199, 1474, 3996, 3532, 3336, 3477, 3462]. embeddings [3228, 3380]. emergency [3974]. emergent [3447]. Emerging [3410, 3329]. emissions [3536]. emotion [3647, 3383]. emotion-sensitive [3383]. emotional [3860]. emphasis [4382, 185]. Empirical [3097, 3911, 3755, 3754, 4081, 3266, 2737]. Employing [3145, 3890, 3540]. empowered [4068]. Emulating [3264]. enable [3637]. enabled [3730, 3181, 3926, 3732, 3653, 3671, 3277, 3957, 3198, 3912, 3737]. Enabling [3413, 3091, 3074, 3623]. enactment [3346]. Enclosing [1797, 1516, 4173, 2053, 2051, 935, 1414, 4171, 1222, 4227, 409, 433]. Enclosure [2519, 2052, 1741, 2819, 2598, 606]. Enclosures [2173, 2578, 3152]. encoder [3699]. encoding [3466, 2995]. Encountered [4105]. encrypted [3422, 3927, 3342]. End [2210, 3400, 2934, 987, 3929, 3436, 3290]. end-device [2934]. end-to-end [3929, 3290]. ende [1955]. ended [2672]. endlich [414]. endlich-dimensionaler [414]. endlichdimensionalen [632]. Endliche [652, 323, 743]. endlichen [293, 710, 740, 617, 570, 747, 665, 903]. endlicher [49, 763, 1114]. endorsement [3154]. endowed [3212]. endpoint [3703]. Endres [812]. Energie [987]. Energy [2953, 3536, 4249, 3923, 3103, 3172, 3706, 3550, 3007, 3356, 3242, 2402, 3085, 3774, 2252, 3770, 3038, 3125, 3173, 2951, 3671, 3962, 3769, 3612, 3827, 2975, 3084, 3580, 3641, 3552, 3221, 3101, 3132, 3170, 3915, 2939, 3008, 2952, 3335, 3174, 3223, 2756, 3145, 3880, 3804, 3347, 3359, 3165, 3578, 3104, 3854, 3727, 3163, 3822, 4060, 3595, 987, 3181, 3603, 3896, 3304, 2888, 3584, 3341, 3427, 4010, 3248, 3205, 3442, 3794, 3100, 3191, 3171, 3942, 2901, 3972, 3604, 4008, 3429, 3776, 3820, 3807, 3423, 3396, 3869]. energy-adaptive [3776]. Energy-aware [3923, 3103, 3774, 3962, 3084, 3854, 3896, 3427, 3942, 3869]. Energy-centric [3007]. energy-consumption [2901]. Energy-efficient [3536, 3706, 3085, 2951, 3769, 3612, 2952, 3174, 3223, 3804, 3359, 3822, 4060, 3603, 3442, 3429, 3820]. Energy-makespan [3550]. energy-proportional [3132]. energy-saving [4008]. enforcement [3426]. enforces [4125]. Enforcing [2963]. Engine [4296]. Engineering [335, 542, 455, 1634, 4186, 90, 4383, 2213, 251, 359, 2461, 3911, 4046, 3616, 3397, 2604, 3967, 3325, 4168, 2759, 3211, 137, 283, 3493, 217, 259]. Engineering/Scientific [4186]. Engineers [540, 360, 150]. Englischen [33]. English [3377]. Enhanced [3390, 3524, 3707, 2826, 3912, 3689, 3994, 3977, 3970, 3735, 4031, 3633, 3962, 2984, 3378, 3073, 4047, 3609, 3159]. enhancement [3107, 3848, 3532, 3358, 2929]. Enhancing [4025, 3876, 3891, 3928, 3930, 4041, 3462, 4013, 3046, 4014, 3976, 4063, 3346]. enriching [3577]. ensemble [3369, 3488, 3992, 4012, 4045, 3967, 3282, 3737, 3235, 3292]. ensembles [3395, 3405, 3516]. ensembling [3827]. ensure [3654]. Entdeckung [286]. Entdocken [186]. Entering [2482]. enterprise [3270]. entitled [3674]. entity [3744, 3395, 4002, 3796]. Entrant [1126]. entropy [3672, 2728]. entropy-based [3672]. Entscheidbarkeit [626]. Entscheidbarkeitsprobleme [312]. Entwicklung [1021, 32]. Entwicklungen [4, 910, 163]. Entwicklungsstand [85]. Entwurf [4127, 146, 133]. Enumerating [906]. Enumeration [869]. Envelopes [1916]. Environment [2045, 2256, 3625, 1023, 3580, 3688, 2913, 3398, 3861, 4027, 3712, 3103, 3610, 3726, 3697, 3504, 3822, 4060, 2936, 3938, 3818, 4072, 3522, 3708, 3761, 3083, 4031, 3759, 3942, 3077, 3367, 3851, 3073]. environment-and-blockchain-based [3851]. environmental [3547, 3815]. Environments [2569, 2838, 3530, 3834, 3940, 3786, 2756, 3752, 4076, 4024, 4033, 3203, 3512, 3652, 3030, 3650, 4051, 3991, 3149, 3826, 3660, 3252, 3285, 4088]. EOR [1473]. EPiC [3318]. epidemic [2801]. epidemics [3734]. episodes [3355]. equal [2704, 3352]. equal-order [2704, 3352]. Equality [2590, 2011, 2148, 2297, 525, 2369]. Equality-Constrained [2148]. Equation [1794, 465, 1484, 2366, 1631, 2348, 2570, 2372, 2095, 1046, 1834, 1045, 1767, 4100, 2563, 1499, 1777, 2301, 761, 1126, 2306, 2466, 1030, 1565, 2511, 990, 1774, 4364, 1406, 7, 1848, 1036, 1820, 2056, 871, 2128, 1168, 1198, 517, 78, 1175, 2670, 1081, 2851, 1684, 969, 1431, 2613, 988, 2611, 1061, 2397, 1585, 1661, 1900, 2875, 385, 1208, 1693, 2735, 2143, 1590, 2014, 1603, 1194, 1328, 614, 2220, 2015, 446, 2004, 1424, 1713]. Equations [2426, 4110, 2341, 4097, 4174, 109, 317, 277, 2532, 1167, 1389, 1826, 1769, 443, 1967, 1934, 2092, 2555, 778, 1141, 2585, 2561, 276, 4201, 1301, 1995, 313, 875, 824, 774, 35, 2102, 2567, 1808, 1930, 2161, 2536, 1893, 1488, 1768, 1051, 4178, 1830, 2591, 2115, 2180, 2440, 2112, 784, 870, 2155, 2091, 779, 2264, 888, 902, 129, 207, 1451, 4353, 1797, 748, 1045, 1043, 619, 2413, 2481, 2353, 2355, 755, 2156, 838, 4355, 957, 787, 2368, 2438, 1365, 2336, 1624, 1968, 1857, 2295, 1776, 1666, 1917, 1927, 1185, 1679, 765, 241, 919, 2208, 1315, 1635, 2309]. Equations [2162, 2478, 1567, 1605, 1501, 4173, 1818, 854, 1887, 933, 1932, 750, 849, 1373, 1906, 1855, 1425, 2189, 2022, 2371, 2533, 1120, 851, 4361, 4345, 2406, 1000, 2240, 39, 2485, 1523, 1626, 2480, 925, 2051, 394, 4120, 1195, 606, 1414, 707, 1534, 2160, 1606, 1480, 2239, 2315, 2396, 2272, 1504, 449, 877, 2465, 825, 767, 1270, 1135, 1466, 789, 2474, 1259, 2477, 2442, 2589, 1589, 1138, 453, 1525, 297, 2785, 1018, 1110, 2782, 1282, 1472, 1761, 1149, 1436, 1066, 2664, 2727, 2778, 1323, 1157, 1099, 1667, 2436, 704, 1721, 1482, 989, 4224, 949, 396, 1492, 1073, 4157]. equations [1699, 2806, 975, 986, 585, 2784, 1263, 584, 1080, 1074, 1659, 1698, 1805, 1209, 1108, 1238, 1085, 2795, 3808, 2703, 1299, 4229, 1749, 2781, 2894, 1700, 1050, 1762, 24, 1214, 2870, 1398, 1208, 2656, 2704, 2750, 2732, 2595, 1200, 1359, 1983, 384, 470, 473, 1115, 450, 571, 1343, 1738, 112, 554, 1225, 948, 452, 4176, 1266, 410, 1716, 1243, 1307, 1572, 1759, 1701, 527, 413, 1990, 1454, 1528, 1553, 2139, 398, 414, 2661, 1764, 1708, 2895, 4237, 962, 976, 1125, 1339, 1358, 1279, 459, 1096, 1757, 1160, 1083, 1226, 1678, 1082, 1563, 984, 1800, 2789, 2748, 1201]. Equations [641, 11, 117, 135, 759, 231, 798, 176]. Equidistant [785, 522, 523, 1102, 2617, 1295]. Equidistribution [2332]. Equilibria [4353]. equilibrium [3459]. Equivalence [2216, 2396, 1270, 985, 2687]. Equivalent [1386, 941, 1103]. era [3616, 3211]. erasure [3130, 3344]. erasure-coding-based [3344]. Erens [390]. EREW [2772]. Erfahrungen [880, 965, 397]. erfüllter [500]. Ergänzungsverfahren [608]. Ergebnisse [16, 338]. Ergodic [840]. Erhaltung [4102]. Erich [856]. Erkennen [22]. Erlang [1644]. Ermittlung [707]. Erratum [2558, 2419, 3152, 1172, 2176, 2894, 3002, 2662]. Error [465, 2671, 2096, 2509, 2204, 1769, 1804, 1496, 2094, 2313, 1448, 2636, 1720, 2405, 1491, 1421, 1680, 2164, 2148, 2861, 913, 1285, 1511, 2205, 752, 1860, 1867, 1303, 2093, 2422, 1197, 2556, 1887, 1057, 782, 1844, 2714, 1379, 1220, 4344, 1498, 162, 242, 2327, 2087, 2416, 1654, 905, 860, 1534, 1486, 125, 4122, 2482, 2106, 2377, 2293, 1458, 2648, 1558, 1761, 1497, 2664, 1338, 2141, 2729, 2606, 1822, 2669, 2865, 2908, 1749, 1050, 1742, 1762, 379, 4158, 398, 1978, 1280, 1452, 1571, 1049, 2741, 2742, 2614, 1483, 1733, 3722, 543, 760, 2011, 4100, 983, 393]. Error [765, 241, 395, 464, 4106, 553]. Error-Bounds [1680, 1860, 2509]. Errorbounds [4174, 394, 921]. Errors [556, 207, 2024, 627, 1290, 4121, 155, 3306, 3240, 2659, 1010, 3574]. Erstellung [386]. ersten [20]. erster [1957, 469, 984]. Erwe [218]. erweiterten [159, 1176, 4103]. erweitertes [1153]. Erweiterung [970]. Erweiterungen [4102]. Erzeugende [785]. erzeugter [1946]. Erzeugung [1916, 1952]. Erzeugungen [637]. ES2DE [3246]. escaping [1031]. eScience [3402]. ESOD [4047]. ESOD-YOLO [4047]. especially [113]. Essen [105]. Essential [2927]. established [559]. Estimate [468, 1338, 2141, 2865, 3417, 2016, 1107, 3289]. Estimates [2096, 2204, 2313, 2253, 1598, 2093, 241, 1290, 1057, 4344, 2241, 1654, 2377, 2648, 2254, 1720, 1749, 947, 393, 395, 2809, 4158, 398, 1049, 2741, 2742]. Estimating [2939, 1010, 3333, 2972]. Estimation [2352, 4245, 2205, 765, 4271, 2416, 2195, 1534, 1486, 209, 1602, 3665, 1021, 2636, 1695, 1050, 2848, 3921, 3879, 3874, 3860, 1452]. Estimations [782, 3665]. Estimationsproblem [209]. Estimator [2471, 1004, 2908]. estimators [2828, 2669]. ETH [579]. Ethereum [3931, 3791]. ETPC [2933]. ETYS [3674]. Eucalyptus [3066]. Eucalyptus-based [3066]. Euclidean [1847]. Euler [1761, 1494, 1481, 2549, 2870]. Euler-characteristic [1494]. EuroMPI [2958]. evaluable [473]. evaluate [3253, 3449]. evaluated [293]. Evaluating [1355, 2804, 4052, 4264, 3160, 1097, 460, 441]. Evaluation [1378, 2446, 1456, 2392, 2228, 1771, 4182, 3238, 2202, 2581, 1811, 1944, 4181, 2556, 2424, 3813, 981, 1371, 3644, 2527, 2106, 2907, 3655, 3956, 3537, 432, 2940, 2619, 2919, 2659, 3062, 1662, 4094, 2933, 4090, 4037, 950, 3421, 1382, 1271, 4228, 3021, 3167, 4129, 379, 3969, 3458, 1351, 1750, 3564, 3201, 4023, 3784, 2901, 3815, 3273, 1670, 3297, 3515]. evaluations [1570]. evaporated [969]. even [1412, 1865, 725]. Event [3126, 2893, 3015, 2997, 3510, 2942, 3426, 3955]. Events [2068, 2150, 3299, 3547, 2946, 2945]. every [1031]. everyone [2852]. Evidence [2257]. evolution [2946, 3631, 3542, 3713, 2798, 3327, 2981]. Evolutionary [4288, 3966, 3479, 3529, 3632, 4036]. evolutionary-EAC [3529]. Evolutions [2452]. evolved [4025]. evolving [2800]. exa [3170]. exa-scale [3170]. Exact [1360, 1464, 2536, 2251, 1440, 1365, 2119, 1306, 3835, 1387, 4364, 1198, 1684, 2617, 2768, 2774, 1070]. exactness [2729]. Example [2347, 1447, 3900]. Examples [621]. excluded [2790]. exclusion [3853, 3678, 3601]. executable [3039]. execute [4072]. executing [3141, 3684].

Execution [2403, 1898, 3781, 3255, 3334, 3519, 3800, 3851]. exemption [3600]. Existence [4334, 4337, 2423, 2469, 1788, 2464, 912, 2208, 1841, 1926, 1190, 1151, 297, 78, 2597, 597, 4237]. Existenz [78, 912]. Existenz- [78]. Existenzaussagen [297]. Existenzkriterien [597]. exit [507]. exit-automata [507]. exp [1005]. exp-trig [1005]. Expanded [1153]. Expansion [2412, 1673, 2521, 1887, 3598, 1021, 1338, 2675, 725]. Expansions [1769, 1421, 2281, 2351, 1427, 2192, 1458, 2783]. Expectations [2188]. Expected [1851, 2690, 1264]. Experience [880, 2149, 1287, 3333, 3698]. Experiences [804, 892, 965, 4168, 397]. Experimental [2202, 2872, 94, 2936, 1006, 1774, 251, 2876, 697]. Experiments [246, 2434, 1981, 1451, 590, 3341, 1607, 3091]. expert [4056, 2991, 4231]. expert-based [4056]. expertise [3012]. explainable [3949, 3285, 4014, 3617]. explained [2744]. explaining [2800]. explanation [3828]. Explicit [1044, 1910, 786, 1506, 1991, 2549, 1573, 761, 2064, 2172, 2432, 1926, 2557, 666, 1505, 2785, 3161, 1431, 1490, 1976, 1678, 129]. explizite [786, 129]. exploitation [320]. Exploiting [4327, 3935, 3721, 3130, 3177, 3437]. exploration [4004, 3317, 3716, 3867, 2973]. exploratory [3911, 3698, 3404]. explorer [2935]. Exploring [3911, 3453, 4093, 4081, 2765]. Exponential [206, 2066, 1998, 1827, 1559, 4211, 893, 1194, 383, 574, 1021, 1756, 2619, 1547, 1236, 976, 1671]. Exponential-fitting [1194]. Exponentialanpassung [976]. Exponentialfunktion [1236]. Exponentialintegral [893]. Exponentialintegrals [1021]. Exponentially [854]. Exponentials [2551]. Exponentialsplines [1547]. exponentialsummen [206]. exponentiell [854]. exponentielle [1998]. expression [2814, 3383, 3860]. Expressions [1801, 1871, 4101, 1879, 1371, 478, 4182, 3321]. extendable [1200]. Extended [159, 2788, 855, 4103, 4112, 3756, 1432, 1983, 4114, 1818, 1538, 1845, 1523, 3142, 1714, 2633, 1324, 1433, 1176, 2933, 3246, 2748]. Extendible [1587, 1385]. Extending [3163]. Extensibility [1900]. extensible [2599]. Extension [2366, 1634, 4186, 1959, 2811, 3959, 2948, 970, 2956]. Extensions [2412, 4102, 2131, 4150, 4149, 1785, 1184, 1190]. Extensive [858]. Exterior [2376, 2114, 385, 2889]. External [1781, 3716]. extracting [3355, 2949, 2954]. Extraction [2515, 3704]. Extrapolation [1804, 611, 2155, 448, 1529, 1452, 1536, 1474, 1560, 2775, 1280]. extremal [1754, 1566]. extremaler [1754]. Extreme [4312, 2835, 3529]. extremum [53]. Extremwertermittlung [53].

F [169, 271, 60, 515, 329, 532, 61, 389, 481, 251, 187, 127, 402, 115, 493, 658, 421, 3984]. F-yolov7 [3984]. F. [374, 143, 184, 429, 60, 727, 218, 430, 195, 193, 10, 11, 279, 198, 389, 481, 352]. FAACP [3023]. FAARAF [3626]. FaaS [3507]. FAC [4162]. face [3731, 3147, 3286]. faceted [3405]. Fachberichte [124]. Fachtagung [728]. Fachwörterbuch [423]. Facial [3975, 3383, 3860]. facilitate [3182]. facilitating [3951]. facility [3269]. facing [3713]. Factor [239, 3635, 3236]. factorable [1333]. Factored [718]. Factoring [2481, 1362]. Factorization [2518, 2335, 1109, 4146, 1432, 2030, 2159, 1605, 55, 2705, 2734, 1212, 1467, 3174, 2703, 2641, 1200, 1737]. Factorizations [1255, 1772]. Factorized [2585]. Factors [893, 163, 2482, 3556]. factory [3468]. factory-aware [3468]. Faddejew [27]. Faddejewa [27]. fading [3664]. Fahrpläne [648]. Failure [2940, 846, 3439, 3817, 3039]. Failures [4306, 3514, 3579, 3753, 1269]. Fair [3106, 3193, 3438]. Fairing [4249]. Fairness [2448, 2207, 3756, 3932]. faithful [1236]. fake [3982, 3926, 3617]. Faktor [239]. Faktorisierungsverfahren [1362]. Falb [169]. fall [3314]. false [526]. falsi [1447, 595, 526, 503, 1234, 595, 503]. Families [52, 1253, 687, 2411, 4191, 1380]. Family [4353, 2594, 2659, 3431, 2692, 1293, 1686]. Fano [326]. far [3507]. Farkas [1759]. Farkas-type [1759]. farming [3976]. Fasol [462]. Fassung [461]. Fast [3384, 1133, 2082, 2228, 2061, 1346, 940, 1712, 494, 1121, 2617, 2610, 2181, 2574, 2024, 2124, 2998, 4162, 4361, 4345, 805, 2121, 998, 2465, 2013, 2003, 3912, 2997, 3984, 3403, 2010, 3529, 3808, 2044, 314, 2774, 1017, 3082, 1613, 2655, 3921, 981, 1052, 1341, 1755, 1156, 1462, 3571, 566]. Fast-MICDTN [3571]. fast-optimaler [981]. faster [4084]. fastest [1198]. Fastkörpers [4103]. Fat [4313]. Fault [1897, 3611, 3040, 3263, 3570, 3587, 4045, 3354, 3146, 2821, 3404, 3232, 3561]. Fault-Tolerant [1897, 3611, 3040, 3263]. faults [3262]. faulty [3040]. FC4CD [3244]. FCA [3217]. FE [2262, 2114, 2482, 2741, 2742]. FE-Discretisations [2262, 2482]. FE/BE [2114, 2741, 2742]. Feasibility [4097]. Feast [2277]. Feature [2258, 4318, 3421, 4215, 4258, 3330, 3819, 3276, 3858, 3379, 3704, 4043, 4035, 3967, 3867, 4005, 3975, 3582, 2802, 3624, 3524, 3725, 3531, 3378, 3418, 3883]. Feature-Based [4318]. feature-enhanced [3378]. feature-model-guided [3867]. Featuremodelling [4202]. Features [4259, 4281, 3730, 2762, 3878, 3248, 3377, 3235, 4009, 4138]. FedBPM [4043]. Federated [3936, 3746, 3913, 3863, 3782, 3587, 4026, 3994, 3858, 4043, 4068, 3793, 4031, 3868, 4034, 3561, 3882]. FedFlow [3746]. Feedback [2153, 266, 655, 2824, 4013]. Feedback-Shift-Register [655]. feedforward [2824]. feeding [3888]. Fehlerabschátzuang [395]. Fehlerabschätzung [765, 241, 162, 393]. Fehlerabschätzungen [860, 398, 760, 782]. Fehleranalyse [2011]. Fehlererfassung [465, 4106, 553]. Fehlernormen [983]. Fehlerschätzung [1452]. Fehlerschranken [4100, 379, 242, 464]. Feichtinger [650, 315]. Felder [1131]. Feldverarbeitung [4126]. FEM [2853, 2504, 2628, 2394, 3254, 2859, 2887, 2128, 2377, 2753]. Fenyö [646]. Fernbach [623]. Fernwirkteehnik [124]. Fertigungssteuerung [256]. Feshchenko [641]. Festkommamaschine [556]. Feststellung [740, 617]. FETI [2644]. FETI-DP [2644]. Fettis [220, 88, 269, 604]. Few [1917, 3728, 2067, 4029]. Few-shot [3728, 4029]. fewer [3194, 2623]. FFDMLC [3975]. FFH [2109]. FFTs [2612]. Fictitious [2676]. Fiducial [1321]. Field [2375, 402, 3264, 2114, 2877, 1001, 2480, 3214, 3091]. Fields [4144, 4103, 863, 217, 2639, 2771, 2638]. Fifth [455, 1253, 364, 2791]. fifth- [364]. fight [3629]. File [2152, 3994, 3344, 4023]. filesystem [2803]. filing [1064]. Filippi [578]. Fill [635, 4307]. Fill-in [635, 4307]. filleting [2608]. Film [2848, 2909]. Filter [880, 3023, 4024]. Filter-based [3023]. Filtering [2163, 3108, 2639, 2982, 3689, 3897, 3420]. Filtermethode [880]. Filters [211]. FIML [1602]. FIML-Estimation [1602]. financial [3548]. find [740]. Finding [1864, 2921, 2251, 897, 1724, 1624, 955, 1875, 1368, 3248, 3586, 1949, 2132, 4363, 4282, 2190, 2218, 1264, 1384, 1103, 1352, 1272, 1345]. fine [3982, 2961, 3883]. fine-grain [2961]. fine-grained [3982, 3883]. fingerprint [2985]. fingerprints [3384]. Finish [2278]. Finish-Start [2278]. Finite [2055, 4110, 2401, 374, 1464, 2096, 2233, 1886, 2204, 2439, 2146, 443, 498, 2092, 2375, 2586, 2094, 1301, 2313, 4246, 1803, 2253, 2434, 1825, 2576, 674, 2294, 210, 1822, 2292, 1834, 2118, 1956, 2226, 2563, 2415, 2205, 49, 2506, 1277, 2336, 2568, 1573, 1604, 1666, 2301, 2387, 2370, 763, 1944, 901, 2321, 2124, 912, 5, 4143, 2468, 2029, 2097, 2750, 2332, 1432, 2064, 2349, 4115, 2424, 994, 2053, 2533, 4142, 2358, 715, 2471, 2787, 2553, 2579, 2428, 2571, 2302, 747, 665, 903, 231, 1502, 293, 2594, 2849, 710, 740, 1064, 2633, 1313, 2850, 2141, 2254, 979, 632, 975]. finite [2660, 2860, 506, 2863, 1013, 1661, 4228, 2795, 1114, 2701, 1665, 3437, 3772, 2685, 617, 3658, 2874, 2674, 1660, 2656, 2704, 2732, 2140, 2595, 2722, 2615, 571, 2910, 3002, 1162, 570, 1244, 1986, 2063, 2646, 414, 3352, 1392, 2632, 4137, 2697, 2764]. Finite-Difference [912, 2633, 2595]. Finite-Dimensional [4110, 1665, 414]. Finite-Element [1502, 2533]. finite-Element- [1013]. Finite-Part [1944, 1392, 2697]. Finite-State [374, 5]. finite-time [3658]. Finite-Volume [2096]. finiten [994]. finites [1956]. firefly [3046, 3498, 3279]. First [2092, 1957, 771, 824, 1488, 469, 88, 854, 2160, 604, 1282, 2633, 2664, 2670, 1157, 1099, 1715, 20, 1209, 1238, 1060, 1293, 2014, 1115, 976, 2796, 1156, 984]. First-Kind [824]. First-Order [771]. Fit [1643, 1971, 1387]. fitness [3622]. fits [2871]. Fitted [2568, 854]. Fitting [2673, 4302, 4321, 2507, 2326, 2133, 1202, 946, 1194, 1682]. Five [1839]. Fixed [556, 2048, 1185, 562, 868, 1544, 2634, 451, 1111, 1453, 1299, 3035, 1214, 470, 1663, 3366, 2002, 3155, 2932]. Fixed- [562]. Fixed-Point [556, 451]. Fixpunkte [1453]. Fixpunkteigenschaft [451]. Fixpunktsatzes [868]. Flachen [1944, 958, 833]. flächenabgleichssplines [1053]. Flächenstücken [1016]. flame [3421]. flash [2803]. fleets [3976]. Fletcher [783]. Flexible [2431, 1682, 3348, 2800]. flight [3440, 4063]. flit [3810]. Floating [1496, 4098, 4099, 997, 1285, 1511, 1970, 764, 883, 885, 938, 1306, 1914, 1619, 1497, 1571].

Floating-Point [4098, 4099, 1511, 764, 883, 885, 938, 1914, 1571]. floodplain [3504]. Florian [3451]. Flow [2943, 2576, 2414, 754, 1773, 2342, 4305, 2441, 2443, 696, 2302, 2489, 2125, 559, 2698, 2857, 2864, 1118, 2866, 964, 2869, 2972, 80, 3778, 474, 2889, 3985]. Flow-inspector [2943]. Flowchart [882, 1172]. Flows [2384, 2292, 1941, 1337, 2879, 1782, 2858, 2880, 3352, 593, 600]. Fluid [2344, 4305, 2866, 2880, 3574]. fluid-structure [2866]. fluids [2869]. Flüsse [600]. Flußprobleme [964]. flux [2850, 947, 2910, 3002]. flux-corrected [2850]. fly [2898]. flying [3853, 4057]. Fock [4166]. focus [3210, 1979]. Focused [3244]. focusing [3885]. Fog [3481, 3594, 3761, 4065, 3972, 3482, 4089, 3655, 3641, 3857, 3398, 3507, 4019, 3643, 3814, 3786, 4027, 3901, 3953, 4076, 3706, 3550, 3816, 3822, 3964, 3522, 3881, 3467, 3229, 4051, 3760, 3826, 3942, 3644, 3186, 3864]. Fog-based [3594, 4065]. fog-cloud [4027, 3550]. Fog-Cloud-IoT [3864]. fold [1687]. Folding [4188]. Folgen [664]. Fóllinger [810]. Follow [2403]. Follow-up [2403]. Following [1872]. Fonctions [258]. food [3902]. footprint [3794]. foraging [3908, 3155]. force [3258]. Forcing [2269, 1578]. forecast [3020, 1733]. Forecasting [375, 3880, 2867, 4048, 3411, 3967, 3770, 3827, 3457, 4014]. foreign [3327]. forensic [3128, 3126, 3668]. forensics [2948]. forest [3968, 3403]. forestry [3385]. Foreword [3118, 3543, 2796]. Form [2446, 456, 4298, 2227, 988, 1492, 4321, 2080, 2336, 2508, 712, 713, 1308, 2556, 4325, 2365, 1000, 4259, 4218, 1606, 718, 1505, 1455, 1494, 1022, 1292, 1298, 1372, 1490]. Formal [3662, 929, 478, 4087, 3045, 3515]. Formale [698, 686]. formalen [478]. formaler [17]. formalism [3171, 3795, 4001]. Formalization [885, 938, 3362]. Formel [1070]. Formeln [1128, 364, 467, 888, 171]. Formelpaar [1492]. Formelsammlung [190]. Formen [55]. Forms [1514, 941, 2427, 1489, 55, 1581, 718, 985, 1211, 1034, 1329, 1068, 1646]. Formula [2548, 77, 1247, 888, 2557, 1388, 1492, 2182, 1238, 1070]. formula-pair [1492]. Formulae [1786, 1668, 2093, 309, 562, 1654, 905, 1195, 199, 2527, 1817, 4285, 1471, 2659, 1763, 563, 1983, 1485, 2135, 1686]. Formulas [1128, 69, 918, 467, 1491, 619, 993, 1950, 2104, 17, 2248, 364, 1210, 1676, 3227, 3271, 1738, 1402, 2021, 1764, 1335, 1201, 144]. formulating [569]. Formulation [2012, 2480, 998, 2849]. formulations [2676, 2114]. Formulierung [569]. Formulierungsprozeß [30]. Forschung [647]. Forsythe [247]. Fortbildung [643]. Forth [3925]. FORTRAN [602, 817, 87, 4127, 281, 357, 64, 253, 359, 430, 13, 1634, 4186, 515, 4127, 4239, 1258, 942, 2385, 1913]. FORTRAN-Programmen [357]. FORTRAN-SC [1634, 4186]. FORTRAN-XSC [4239]. Forward [1668]. forwarding [4052]. Foster [401]. Foundations [4268, 700, 4115, 4277, 1347, 4378, 4109, 4096, 128, 4366]. Four [551, 1839, 3909, 1933, 634, 1732, 312]. four-dimensional [634]. Four-Directional [1933]. Four-language [551]. Fourier [578, 2743, 1675, 1346, 2611, 716, 2621, 4229, 314, 1350, 2533, 1427, 991, 998, 1052, 2074]. Fourier-representations [4229]. Fourier-Tschebyscheff-Approximation [578]. Fouriertransformation [314, 998]. Fourth [2267, 467, 2111, 2442, 1589, 975, 986, 1194, 2171]. fourth-order [1194, 2171]. Fox [157, 682, 830, 682]. FPGA [3453, 3521, 3540]. FPGA-based [3521]. FPGAs [2926]. Fractal [2550, 3225]. Fractal-like [2550]. fraction [1338]. Fractional [2101, 1551, 2440, 3119, 3358]. Fractions [1360, 1135, 292]. fragile [3414]. fragments [3713]. frame [3704]. Framework [4287, 3103, 2282, 3655, 3133, 3488, 3492, 3626, 3665, 2653, 4025, 3514, 3230, 2943, 3093, 4016, 3339, 3686, 3666, 4027, 3445, 3788, 3749, 3113, 3399, 3577, 3684, 3167, 3772, 4072, 3993, 3560, 3944, 3496, 3764, 3793, 3493, 2929, 3540, 3442, 3999, 2831, 3932, 3456, 3864, 3681, 3690, 3653, 3800, 3798, 4047, 3212, 2994, 2956, 4014, 3290]. frameworks [3169, 3936]. France [4379]. Frank [221, 28, 1530]. Franks [283]. fraud [3782, 3791]. Fredholm [276, 1749, 2160, 1764]. Free [2446, 4298, 780, 941, 4321, 2265, 2422, 1411, 4325, 2027, 2514, 1267, 4218, 164, 1119, 899, 3680, 456, 3207, 2567, 1227, 2649, 726, 3890, 1609, 2781, 2894, 3437, 2641, 3592, 2667, 3254, 3921, 3843, 2895, 2718, 413]. Free-Form [2446, 4321, 4325]. free-of-charge [3592]. freedom [3675]. Freeform [2458]. FreeRec [3057]. freien [899]. Frequency [2036, 2075, 2163, 583]. frequent [2944, 3082, 2902, 3065]. Frequenzkennlinien [462]. freshness [3891]. Freud [607, 607]. Friedrichs [197]. friendly [3540]. Frisch [795]. Frobenius [859]. Frobeniuswurzel [859]. Fröberg [177]. Fromme [223]. Frommer [2597]. Front [1312]. Front-Tracking [1312]. Frontier [4291]. Frontier-Region [4291]. frontiers [3542]. fronts [2887]. frugal [3921]. Fruitful [4293]. Full [1961, 1232, 2632, 2779]. full-rank [2779]. Fully [1794, 2838, 2701, 3319, 3753, 3939]. Fully-Discrete [1794]. Function [1648, 52, 882, 1172, 618, 1121, 922, 1889, 2249, 1291, 920, 1922, 990, 2060, 496, 1578, 6, 1338, 440, 2776, 1019, 20, 1269, 4232, 1404, 1541, 667, 3505, 1236, 53, 1333, 3725, 842, 1191, 1718, 4061, 370]. function-theoretic [6]. Functional [109, 2092, 72, 313, 2536, 2303, 2295, 2310, 1149, 1436, 2436, 4229, 1709, 3920, 694]. Functional-Differential [2295, 1709]. Functionality [4201]. Functionals [2090, 878, 4261, 2539, 1044, 1763].

Functions [1378, 1591, 890, 2270, 1957, 143, 2134, 4183, 69, 285, 1468, 1596, 954, 1346, 2147, 21, 760, 4252, 2542, 705, 4101, 2538, 845, 1549, 2131, 2574, 1314, 722, 993, 2127, 2265, 2422, 850, 1116, 785, 1166, 4327, 2365, 1124, 4118, 841, 555, 1920, 839, 2547, 199, 269, 537, 4185, 4331, 1027, 2361, 909, 2671, 2752, 2628, 1010, 2199, 149, 1237, 2687, 460, 1453, 1060, 1650, 983, 4184, 441, 1687, 2140, 1071, 1140, 470, 2040, 272, 1225, 2715, 1702, 592, 53, 471, 1256, 1240, 1218, 134, 212, 220]. Functoids [4165]. Fundamental [2341, 485, 142, 3492, 2735, 2783, 1754, 3542, 1566]. Fundamentals [4368, 138, 1188, 533]. Fundamentalsysteme [1754]. funds [2769]. fünfter [364]. Funktion [990]. Funktionalanalysis [324]. funktionalanalytischer [694]. Funktionale [878]. Funktionen [16, 21, 4101, 785, 272, 592, 53, 890, 705, 983, 993, 850, 457, 1922, 555, 1920, 839, 1256]. Funktionenfamilien [52]. funktionentheoretischer [6]. Funktionentheorie [179, 119]. Funktionentheorien [6]. Funktionsebenen [30]. Funktionstafeln [191]. Funktionswerten [53]. furr [614]. Further [2107]. fused [3249]. Fusing [3415]. Fusion [2257, 3730, 3547, 3343, 3570, 3555, 4043, 3975, 3378, 3883]. FusionCL [3555]. Futh [642]. Future [3583, 2917, 3506, 3661, 3828, 3875, 3494]. futures [3979]. fuzzing [4040]. Fuzzy [2777, 3523, 3812, 3226, 3136, 2060, 4285, 2782, 3298, 3626, 3230, 2931, 3939, 3208, 3761, 3458, 3100, 4021, 2728, 3158]. Fuzzy-based [3136]. Fuzzy-IoT [3298]. FVM [2857, 2128].

G

[353, 354, 375, 421, 72, 416, 238, 271, 429, 388, 377, 607, 515, 323, 357, 103, 576, 743, 34, 141, 47, 187, 91, 324, 255, 320, 537, 970, 686, 1700]. G-B-S-method [1700]. G.

[602, 315, 16, 35, 186, 285, 286, 577, 607, 283, 698, 650, 648, 301, 197, 657, 829, 510, 216, 257, 266, 247, 970, 182, 535]. GA [4045, 3181]. Gaede [15]. Gaines [377]. gains [964]. Galerkin [836, 760, 2115, 760, 995, 1199, 2860, 2908, 2563, 1209, 1465, 1660, 2312, 2356]. Galerkin-Verfahren [760]. Gallager [416]. Galler [180]. Game [3494, 192, 3821, 3133, 3768, 3839]. Games [148]. Gamkrelidze [271, 187]. GAMM [2509]. GAMM-Workshop [2509]. Gamma [788, 1159, 1129, 1370, 1229]. GAN [3982]. Ganin [576]. Gantmacher [115]. Ganzhorn [812]. Ganzzahligen [889]. Garner [337]. gat2vec [3267]. Gate [4188]. gateway [2976, 3309]. gathering [3242, 4086, 3019, 3845, 3807]. Gauging [2290]. Gauss [1963, 1961, 1961, 1142, 2707, 2594, 1963, 2158, 708, 1403, 1659, 1698, 2324, 2049, 1142, 905, 2302, 1880]. Gauss-Turan [905]. Gaussian [4097, 1497, 1338, 683, 1433, 1533, 621, 2181, 2562, 1571, 630]. Gaussians [2326]. Gauss’sche [708]. Gaussschen [4097]. GbR [4379]. Gebiet [1126]. Gebiete [112]. Gebieten [156]. Gebildeten [477]. Gebrauch [512, 270]. Gebrucg [485]. Geburtstag [439, 106, 605]. gedampfter [1965]. Gehirngespinste [793]. Gehring [756]. gekoppelte [619]. gekrummten [1944]. gemacht [288]. gemessenen [316]. Gemischt [889]. Gemischt-Ganzzahligen [889]. gemischten [619]. Gemischter [849]. Genaue [4098, 1958]. General [1723, 904, 2025, 819, 4356, 4340, 2383, 4104, 2333, 1495, 886, 751, 767, 867, 1388, 307, 1161, 4222, 4175, 1245, 1667, 1090, 2014, 1983, 1246, 834, 1052, 185]. Generalisation [2191, 1268, 1757]. Generalization [4279, 2261, 1143, 456, 2812, 1891, 2888, 519, 78, 6]. Generalizations [2150]. Generalized [4242, 1051, 2167, 2236, 2484, 1185, 1314, 2546, 912, 1432, 4147, 1674, 4193, 2407, 2995, 1234, 589, 1159, 367, 496, 561, 825, 2056, 1135, 678, 2681, 3829, 1539, 440, 1707, 2006, 1481, 1574, 2869, 314, 1305, 2880, 1739, 3593, 1219, 1736, 1630, 1984, 3046, 1128, 1954, 988, 240, 770, 3334]. Generalizing [3878]. generate [1293]. Generated [1346, 1946]. Generating [111, 931, 4118, 4198, 2442, 2767, 1262, 1463, 3520, 1229, 785, 1916, 1952]. Generation [2582, 3207, 4244, 1803, 926, 1592, 1417, 1568, 738, 1871, 2086, 2181, 726, 1169, 821, 790, 4328, 1637, 2192, 3135, 1778, 2903, 2883, 3949, 1274, 1249, 2862, 3770, 570, 3247, 3875, 3494, 2964, 2749, 637]. Generative [1336, 1105, 2990]. generative/discriminative [2990]. Generator [2154, 1851, 494, 4384, 1274, 1633, 2007, 2647, 591, 591]. Generators [1843, 2106, 1370]. generic [2653, 2654, 2639, 2753]. genetic [4004, 3199, 3230, 3751, 2867, 3667, 2926, 2771, 3633, 3519]. genetic-based [3519]. Genetics [815]. genome [4055]. Genrich [224]. geo [3197]. geo-distributed [3197]. Geodätische [222]. Geodesic [1781]. GEODIS [3197]. geographic [2952]. geography [3582]. geolocated [2942]. Geometric [391, 4125, 4247, 4381, 4314, 1400, 4375, 4380, 4301, 4377, 2444, 2601, 1867, 4255, 4213, 4180, 4220, 2785, 3925, 1025, 2706]. Geometrically [2472, 4326]. Geometrie [238, 980]. Geometry [268, 4301, 238, 1363, 2640, 3029, 980, 1692]. George [352]. geospatial [3363, 2964]. geotechnical [3198]. gerichteten [293, 710]. Germain [492, 351]. German [297, 1, 293, 346, 454, 907, 4109, 4096, 307, 465, 477, 766, 823, 4097, 1037, 1018, 1110, 1655, 1861, 439, 455, 680, 634, 78, 6, 159, 160, 128, 108, 890, 1376, 1683, 917, 1075, 880, 599, 1967, 1934, 1957, 1170, 208, 945, 616, 692, 778, 1955, 17, 710, 740, 432, 1936, 456, 1997, 2039, 1021, 1915, 1943, 1102, 1993, 440, 521, 4098, 4099, 958, 959, 1003, 590, 4126, 1995, 1153, 1409, 185, 56, 206, 598, 904, 1543, 438, 836, 1128, 52, 451, 2008, 238, 106, 605, 77, 749, 889, 1963, 1998, 1954, 1994, 365, 54, 113]. German [245, 786, 1089, 1962, 872, 1139, 1176, 468, 409, 433, 469, 632, 741, 737, 684, 1540, 466, 559, 556, 1347, 772, 972, 1254, 434, 1632, 1946, 2012, 982, 586, 638, 611, 708, 837, 607, 563, 779, 887, 1046, 949, 1941, 364, 467, 888, 239, 396, 695, 902, 609, 21, 129, 207, 988, 1492, 167, 366, 386, 773, 855, 1923, 690, 706, 1265, 859, 760, 1852, 1940, 240, 780, 1947, 2011, 1956, 914, 663, 1960, 975, 986, 1045, 585, 1043, 1361, 619, 1961, 1084, 4100, 1939, 9, 2010, 57, 59, 369, 752, 49, 1016, 852, 461]. German [1349, 1438, 777, 776, 1547, 838, 1098, 51, 20, 4127, 1042, 339, 504, 1395, 460, 705, 787, 4101, 1515, 4128, 568, 588, 569, 1020, 664, 770, 595, 526, 924, 971, 1013, 370, 964, 131, 342, 1968, 1453, 1970, 848, 397, 965, 763, 1114, 1545, 1917, 1269, 820, 1126, 4102, 4103, 1972, 1938, 1944, 722, 294, 306, 24, 712, 713, 878, 983, 993, 298, 278, 316, 345, 617, 314, 608, 1062, 1999, 2000, 908, 382, 558, 693, 987, 724, 393, 379, 458, 441, 442, 765, 1308, 764, 883, 499, 850, 912, 4105, 4366, 4104, 1965, 241, 385]. German [457, 567, 400, 613, 1032, 1754, 682, 1948, 1154, 395, 910, 960, 884, 1914, 662, 879, 384, 470, 436, 473, 4, 19, 610, 450, 571, 3, 854, 933, 1260, 292, 785, 1503, 1932, 1030, 849, 994, 1079, 782, 445, 1922, 112, 272, 554, 1225, 1332, 1919, 1402, 8, 58, 444, 4132, 893, 497, 866, 166, 162, 168, 242, 464, 915, 916, 990, 570, 23, 452, 591, 1236, 55, 851, 1000, 876, 1950, 807, 805, 751, 835, 1949, 410, 519, 597, 594, 612, 163, 572, 863, 2009, 1959, 968, 1884, 1011, 343, 501, 555, 853, 4106, 1142, 981, 80]. German [1966, 1318, 1933, 1951, 553, 43, 1920, 7, 25, 381, 1362, 18, 860, 500, 1070, 274, 592, 697, 1036, 1058, 637, 614, 527, 22, 437, 53, 296, 413, 503, 474, 471, 606, 935, 956, 1038, 1053, 4367, 1921, 1996, 1953, 344, 4135, 707, 1234, 312, 839, 520, 50, 566, 711, 895, 1256, 669, 998, 1931, 398, 414, 961, 980, 1958, 1945, 1113, 408, 371, 601, 973, 953, 1059, 694, 840, 368, 1063, 164, 1107, 502, 560, 747, 877, 976, 1125, 1339, 1358, 1452, 1526, 970, 81, 82, 868, 600, 593, 665, 720, 903, 1964]. German [767, 130, 161, 1461, 862, 4107, 4108, 615, 209, 1969, 165, 495, 478, 789, 412, 689, 1008, 1023, 1131, 133, 1119, 1916, 1027, 1226, 528, 564, 668, 791, 930, 629, 881, 587, 110, 446, 679, 783, 1462, 275, 808, 952, 1076, 1082, 273, 861, 1935, 984, 1918, 2013, 1356, 1952, 84, 867, 1147, 244, 822, 899, 1942, 1024]. Germany [4380, 4378, 4377]. Gerschgorinkreise [57]. Gershgorin [57]. GERT [1165]. gerundeter [4096]. Gerüste [848]. Gesamt [1075]. Gesamt- [1075]. geschlossener [592]. Gessner [794]. Gestalt [473]. gestellter [972]. gesteuerter [569]. gestörter [1226]. gestortes [1966]. Gewichtsoptimale [993]. Gewinnung [297]. gewisse [1963, 782]. gewisser [396, 497]. gewohnliche [1993, 614, 1358]. gewöhnlichen [297, 78, 779, 585, 1226]. gewohnlicher [1994, 470, 1225, 877, 1125, 1339, 758]. Ghouila [347]. Ghouila-Houri [347]. Gigas [222]. Gilbert [89]. Ginzburg [138, 925]. gitterpositiver [1683]. Given [931, 4117, 1712, 970, 593, 1718]. Givens [2705, 847, 1613]. Glaserfeld [29]. glasses [3620]. glatte [371]. glatter [833]. gleichartige [1972]. Gleichung [1126]. Gleichungen [4097, 1018, 606, 787, 2000, 933, 413]. Gleichungs [450, 2011, 452]. Gleichungs-Nebenbedingungen [2011]. Gleichungs-systeme [450]. Gleichungs-systemen [452]. Gleichungssystem [7]. Gleichungssysteme [465, 1540, 1917, 4100, 1968, 990, 414]. Gleichungssystemen [384, 1110, 521, 765, 554, 1036]. Gleichungssystems [1030]. gleichverteilten [664]. Gleitkomma [883, 1914, 764]. Gleitkomma- [1914]. Gleitkomma-Arithmetik [883, 764]. Gleitkommazahlen [4098, 4099]. Gleitpunktalgorithmen [1970]. Global [2270, 2142, 1477, 2322, 1830, 2131, 1348, 447, 2309, 4341, 1887, 1919, 2022, 1649, 2822, 1538, 1774, 1845, 2049, 1534, 2175, 1458, 3908, 4058, 1325, 1152, 4230, 1616, 1758, 3425, 2923, 3420, 1714]. globale [1919]. Globalisierungen [1008]. Globalization [1008, 1574, 1575]. Globally [4338, 1429, 1245, 1374, 1743, 2176]. glycation [3293]. Goldberg [482]. Goldman [62]. Göldner [489]. Gomory [752]. Good [494, 2237, 2599]. Google [3394]. Görtler [237]. GoSCAN [2922]. Göschen [46]. Gould [811]. Gov [3681]. governance [3681]. GPGPUs’ [3049]. GPR [3384]. GPS [4355, 3577, 2954]. GPSS [4127, 4127]. GPSS-FORTRAN [4127, 4127]. GPU [4079, 3310, 2982, 3009, 3238, 3006, 3437, 3772, 3016, 3658, 3969, 3072, 2966, 3254, 3443, 4030, 2890, 3144]. GPU-accelerated [3238, 2890]. GPU-based [3437, 3009, 3772]. GPUs [3741, 2856, 3808, 3463, 2970]. Grades [872]. Gradient [1595, 1051, 1823, 866, 1120, 2339, 2489, 1712, 2673, 1574, 382, 3843, 2726, 1984]. Gradientenverfahren [382]. Gradientenverfahrens [866]. Gradients [4336]. Gradual [1284]. Grain [2042, 2961]. grained [3982, 3883]. gram [2765, 1719]. Grammar [941, 1034, 1329]. Grammars [1336, 1410, 1105, 1119, 1033, 1065]. Grammatiken [1119]. Grammatische [38]. grams [3228]. granular [3245]. Graph [4288, 1336, 4290, 1410, 4197, 2202, 2153, 1653, 4286, 874, 4379, 1062, 4289, 803, 1411, 4190, 4188, 4189, 4291, 2410, 4282, 4284, 903, 1555, 4281, 791, 4196, 3716, 3848, 4050, 1007, 1020, 964, 3689, 963, 3069, 3994, 298, 3345, 1065, 1293, 1103, 4131, 3935, 3778, 1206, 3885, 3850, 3791, 4374, 3462, 3328, 3797, 3905, 3796, 3715, 3989, 751, 835]. Graph-Based [4290, 4281]. Graph-Coloring [2410]. graph-theoretical [964]. Graphen [598, 1020, 323, 712, 713, 298, 1062, 293, 208, 710, 740, 684, 848, 617, 1318, 665, 903, 791, 930]. Graphenalgorithmen [751]. Graphenalogrithmen [835]. graphentheoretischer [964, 495]. Graphentheorie [649]. Graphic [727]. graphical [3834]. Graphics [2489, 2828, 2903, 3187]. Graphs [2541, 1864, 2073, 208, 69, 2308, 2363, 4283, 2201, 4280, 4192, 1724, 2026, 1112, 2531, 4191, 2281, 931, 2219, 848, 775, 726, 712, 713, 4292, 929, 4187, 4195, 665, 4198, 4285, 1905, 293, 710, 740, 598, 2921, 2692, 3900, 684, 1357, 2657, 2629, 2067, 1669, 617, 3902, 1865, 3175, 1318, 3267, 3389, 1847, 3097, 495, 2739]. GRASP [3521]. grasshopper [3627]. gratefully [3674]. Gratis [1622]. Gravitational [3444, 1001]. gray [3641, 996]. greater [1272]. Greatest [4279]. Greece [3138]. Greedy [2101, 2490, 1972, 3925, 3598, 4050, 3983, 3364]. Greedy-Heuristic [1972]. Greedy-Heuristik [1972]. Green [4068, 3915, 2830, 3396, 3113]. Greenberg [281]. greener [4080]. greenhouse [3604]. Greenspan [813]. Grey [4286]. Grid [2426, 2154, 2215, 1997, 2098, 1615, 1568, 1961, 2226, 2086, 2329, 2330, 2332, 3030, 3182, 2826, 3392, 1778, 1683, 3188, 2653, 2654, 2634, 1621, 2567, 2673, 2630, 2397, 2667, 2916, 4162, 3136, 3476, 3916, 3805, 2992]. grid-positive [1683]. Grids [2233, 2262, 2430, 2112, 2118, 2364, 2234, 2427, 2495, 2321, 2314, 1674, 2194, 2360, 2263, 1921, 1953, 2113, 1380, 2584, 2633, 1822, 2699, 2611, 2617, 2615, 1215]. Griffin [624]. Grigorieff [758]. Grinten [284].

GRL [3994]. GRL-aided [3994]. Grochla [643]. Gronwal [456]. Gronwall [456]. Gross [670, 3808]. Große [1931, 316]. grosser [55, 50, 566, 267]. grössten [468]. ground [3651, 4057]. ground-air [4057]. Group [208, 178, 184, 36, 239, 819, 1337, 567, 2239, 2315, 1431, 3872, 2693, 3693, 570, 2794, 2139, 3297, 567]. Group-Theoretic [178]. grouping [2811]. Groups [79, 806, 4141, 390, 1931, 1064, 1171, 2685, 3693, 2759]. growth [2799]. GRU [4025, 3526, 3285]. Gruenberger [91]. Grundbegriffe [343]. Grundbegriffen [876]. Grundlage [1993, 224]. Grundlagen [173, 330, 1347, 424, 757, 581, 489, 4109, 4096, 128, 4366]. Grundlehren [674, 362, 488, 389, 481, 390]. Grundschaltungen [756]. Grundzüge [238, 418, 74, 46]. Gruppe [121, 570]. Gruppentheorie [44]. Gruyter [510]. Gschwind [355]. GSLA [3113]. GSPN [3795]. GSPN-based [3795]. Guarantee [1499]. Guaranteed [2509, 1868, 4344, 4164, 1739, 4182, 3874]. Guard [997]. Guest [3024, 3087, 3508, 3209, 2993]. GUHA [76]. Guide [87, 4144]. Guided [3672, 3532, 3867, 3728]. guidelines [3240, 3326]. guider [3941]. Guillemin [226]. günstiger [163]. Gunston [625]. Gunzenhäuser [392]. Gute [654]. Guttman [90]. GWO [3962, 3776]. Gymnasiem [238]. Gymnasien [418]. gymnasium [238].

H [372, 542, 673, 814, 1793, 1129, 817, 731, 68, 418, 484, 811, 377, 60, 65, 211, 756, 575, 281, 357, 174, 290, 743, 746, 794, 702, 651, 647, 389, 481, 390, 155, 219, 150, 233, 228, 324, 334, 352, 404, 546, 827, 462, 831, 730, 658, 3723, 2466, 2287]. H-Matrix [2466, 2287]. H. [304, 543, 171, 220, 14, 28, 102, 179, 200, 238, 418, 426, 579, 675, 85, 46, 424, 65, 118, 235, 355, 799, 431, 189, 224, 357, 797, 88, 92, 627, 626, 288, 1937, 321, 237, 534, 582, 37, 95, 356, 677, 642, 150, 269, 270, 404, 604, 800, 801, 252, 403, 428, 831, 833, 629, 728, 730, 175, 158]. H.-J. [1937]. Haar [2310, 500]. Haarscher [500]. Hackl [816]. Hadamard [2697]. HADAS [3248]. Hadley [577]. Hagelschurer [744]. Halbgruppen [763, 1114]. halblinearer [1999]. halblogarithmischer [1852]. Halbordnung [54]. Half [1910, 2172, 1396]. Half-Explicit [1910, 2172]. Halley [1792, 2244, 1601]. Halley-Like [2244]. Halpern [419]. Hamiltonicity [2067]. Hammer [327]. Hamming [3245]. Hand [1858, 789, 1279]. Handbook [134, 69, 147, 101, 389, 481]. Handbuch [492, 331]. Händler [531]. handling [3141]. Handover [3573]. hanging [2620]. Hankel [2734, 1627]. Hanoi [1570, 1732]. Hansen [2597]. Hansjörg [1859]. Harat [329]. Hard [1881, 2051, 2803, 1865, 3540]. Hardware [4129, 2489, 952, 3177, 2926, 2816, 3556]. hardware-based [2926]. Hardware-Software-System [952]. Harmonic [2354, 2794, 3921]. Harmonics [1440]. Harnessing [3733]. Harris [249, 3830, 3854, 3524]. Hartmann [32]. Hartree [4166]. harvesting [3671]. HaSGP [3715]. hash [3680, 2985, 2039]. Hash-Speicherung [2039]. hashing [3549, 3938, 2039]. hate [3369]. HATMOG [3633]. Hauptwertintegrale [1963]. Haus [105]. Hausdorff [1811, 4228]. Having [2247]. Hawgood [93]. Hawkins [513]. Hawks [3830, 3854, 3524]. Hayes [537]. Hays [353, 354]. HCI [3045]. head [502, 3962]. header [3927]. header-based [3927]. healing [3621, 3110, 2758]. Health [536, 3455, 3730, 3746, 3422, 2892, 3999, 3622, 4031, 4020, 3142, 3829, 3291]. health-cloud [3999]. healthcare [4006, 3912, 3471, 3403, 3654, 3729, 2925, 3735, 3936, 3724, 3962]. Heaps [4193, 2751].

Heapsort [1924]. heart [1418]. heartbeat [3327]. Heat [467, 1045, 2466]. heating [2877]. Heaviside [827]. hedonic [3120]. Heft [29, 30, 37, 31, 33, 32, 38, 43]. Height [1137, 1104, 1217]. Heinhold [15, 10]. Heinz [529]. Helical [2213]. Hellfors [248]. Helmholtz [2129, 2579, 1713]. Helmholtzsche [385]. Helmut [1543]. Help [458, 3716]. Helps [1947]. Hennie [374]. Henrici [139]. Herausgeber [70, 28, 69, 200, 238, 249, 302, 157, 122, 121, 197, 75, 66, 237, 61, 192, 124, 205, 267, 254, 204, 233, 324, 269, 258, 259, 123, 175, 158, 156, 153, 287, 203]. hereditary [2657]. Herleitung [694]. Hermes [626]. Hermite [2564, 1028, 2117, 1744, 902, 1842]. Hermite-Collocation [902]. Hermitesche [902, 1000]. Hermiteschen [914]. Hermitian [2046, 2591, 1000]. Hern [605]. Herschel [512, 376]. Hersonissos [3138]. Herstellung [272]. Hertz [48]. Herzberger [1891]. Hessian [2538, 1333]. Hestenes [225]. heteroclinic [4237]. Heterogeneity [3149, 4038]. Heterogeneity-aware [3149]. Heterogeneous [3622, 3365, 3976, 3188, 4079, 3174, 3666, 3848, 3263, 3319, 3702, 3804, 3752, 3591, 3463, 3275, 3904, 3896, 3193, 4043, 3584, 3902, 3842, 3151, 3518, 2846, 3800, 3448, 3429, 3886, 3715]. heterogeneous-aware [3715]. Heuristic [1985, 2138, 3939, 1972, 1401, 399, 803, 1273, 2072, 1935, 3565, 4050, 3334, 3896, 3511, 3921, 3596, 3149, 3957]. Heuristic-Based [2072]. Heuristics [2187, 2026, 2202, 4268, 2283, 4196, 3303, 3584]. Heuristik [1972]. hexa [2827]. hexa-cell [2827]. Hexaedral [2439]. hexahedral [2620, 2615]. Hg [353, 354, 543, 490, 418, 485, 362, 488, 623, 701, 656, 644, 481, 624, 541, 672, 686]. HHSE [3848]. Hiagh [183]. Hiagh-Accuracy [183]. Hidden [1400]. Hidden-Line [1400]. Hiding [1731, 4071]. Hierarchial [2234]. Hierarchical [4046, 2585, 2501, 2494, 2134, 2496, 2413, 2483, 2451, 1635, 1163, 2194, 2625, 1502, 2642, 3723, 3958, 4091, 2892, 2675, 3030, 1206, 3222, 3515]. Hierarchically [2314]. Hierarchies [2081]. High [2586, 2267, 1831, 2591, 2166, 2496, 2860, 1835, 4207, 2338, 2746, 2373, 3254, 1945, 522, 1964, 2212, 753, 3952, 3384, 4079, 2828, 1615, 2680, 2939, 2883, 3174, 3594, 2199, 4182, 2857, 3351, 1084, 2643, 2863, 2968, 3029, 3723, 2969, 2996, 583, 3775, 3096, 3185, 3273, 3978]. high-availability [3775]. High-Dimensional [2496, 3723]. High-Index [1945]. High-Level [2212]. High-Order [2166, 2338, 1964, 2586, 2860, 2680, 2857]. High-order/ [2860]. High-performance [3952, 4079]. high-speed [3273]. Higher [2426, 823, 1514, 317, 277, 1187, 1468, 872, 2415, 1277, 2874, 1425, 1888, 2195, 606, 40, 2428, 2571, 2539, 2367, 1727, 2594, 3848, 986, 1690, 3835, 292, 3503]. Higher-Order [2539, 2367, 2874, 2571, 3848, 1690, 3835]. Highlights [3246]. Highly [2424, 841, 2579, 2680, 3005, 2627, 3546]. highway [80]. Higman [420, 178]. Hilbert [409, 433, 836, 1128, 468, 409, 433, 1629, 783]. Hilbert-Räumen [409, 433, 468]. Hilbertraum [836, 783]. Hilberträumen [1128]. Hilbertsche [118]. Hilfe [2012, 21, 1944, 458, 457, 112, 866, 55, 503, 478, 789, 984]. Hilfsmittel [287, 362, 488, 656]. Hill [421, 575]. Hinderer [700]. Hinode [3345]. histo [1053]. Histograms [1902]. Histopolation [1799, 1953]. historical [3469, 3345]. History [533, 3111]. HITS [3906]. HKT [2543]. Hoare [792, 731, 1136, 1015]. hoc [2986, 2952, 3853, 3884, 2988, 4057, 3458, 2831]. Hochrechnungsproblems [1503]. Hochschulen [576]. Hochschulstundenplänen [386]. Hochschultaschenbücher [63]. Hoffmann [796]. Höhenkarten [272]. Höhenmessung [222]. höher [1084, 1964]. höherem [1945]. höheren [872]. höherer [823, 986, 292, 296, 606]. hole [2709]. holistic [3461]. Hollberg [745]. Holomorphic [2422]. HOM4PS [2668]. HOM4PS-2.0 [2668]. homes [3212]. homoclinic [4237]. homogeneous [3220, 704, 989, 1199, 3518]. Homogenization [2331]. homology [1171]. Homomorphic [4148, 3184]. Homomorphically [3342]. Homotopiemethode [1349]. Homotopy [2475, 2573, 2487, 1349, 2668]. honesty [873]. hop [3242]. Hopf [1863, 2087]. Hoppe [30]. hopscotch [1708]. hormone [3292]. hormone-binding [3292]. Horner [2392, 1379, 2327]. hospital [3564]. host [2984]. hosted [3342]. Hotes [356]. Hotz [698, 686]. Hough [2001]. Houri [347]. Householder [389, 481]. households [3133]. Howard [694]. hp} [2368, 2336]. hp}-adaptive [2368]. hp}-Finite [2336]. HPC [2891, 3004, 3969]. HPMaX [3463]. HRCT [4082]. Hrsg [731, 727, 812, 818, 801, 728, 830, 745]. HSS [2731]. HTTP [3277]. HTTP-based [3277]. Hu [422]. Huang [2182, 1737]. Huard [2158]. Huber [2664, 2670, 2727, 1004, 2471]. Huffman [1973, 133]. Huffman-Mealy [133]. huge [3226]. Hull [1857, 2458, 636]. Hulls [1441, 1264]. Human [2441, 3137, 3861, 3526, 3894, 2892, 3997, 3582, 3617]. huMAN-centric [3617]. hurdle [3829]. Hurwitz [1366]. HW [2926]. HW/SW [2926]. Hybrid [3966, 1789, 1643, 387, 4045, 588, 2309, 4035, 2771, 4260, 2108, 1063, 535, 645, 3576, 3608, 4025, 3301, 3081, 4012, 4039, 2854, 3712, 3891, 949, 2803, 3928, 3192, 3184, 568, 2938, 3687, 3823, 3032, 3390, 2990, 4069, 3667, 3967, 2682, 3736, 3673, 3075, 3633, 3790, 3916, 3692, 3827, 3386, 3084, 2786, 3869, 3707, 3302]. Hybrid-System [588]. hybride [568]. hybriden [1063]. Hybrides [491]. Hybridrechnern [949]. Hydraulie [257].

Hydrodynamic [2312]. Hydroelectric [1938]. hydrological [2867]. HYGOL [568]. Hyndman [645]. hyper [2827]. Hyperbeln [766]. Hyperbolas [766]. Hyperbolic [1506, 1610, 1254, 129, 2310, 4103, 1315, 2404, 1151, 2343, 1175, 995, 1199, 1061, 1208, 18, 984]. hyperbolische [129, 207]. Hyperbolischen [4103, 1254, 18, 984]. Hyperbus [2177]. Hypergeometric [2586]. Hypergraph [4279]. Hypergraphs [4279, 2944, 3835]. Hypernormbälle [895, 773]. Hypernormballs [895, 773]. hyperplane [2786]. Hyperpower [1860]. Hyperrectangles [2089, 4228]. Hypersingular [2527]. hypervisor [3938]. Hypotheses [76]. hypothetical [3949]. hyrax [3553]. Hysteresis [1626]. Hyvärinen [540].

I.

[12, 167, 90, 712, 103, 576, 651, 141, 135, 250]. IaaS [3093, 4012, 2757, 3120, 3121, 2760]. IBM [490, 603, 305, 120, 288, 204, 351, 492]. IBM-System [490]. IBM-System/360 [490]. IBM/360 [288, 492]. ichi [815]. ICIRS [3767]. ICN [3930]. ICN-based [3930]. ICO [3766]. Iconic [1730]. ICSOC [3866]. ICUs [4082]. ideation [3618]. Identical [2250, 2420]. Identification [3767, 3589, 3292, 1, 4026, 2987, 3250, 1755, 2942, 3883]. Identifizierung [1]. identify [3178, 3693]. Identifying [3056, 3996, 2902, 3312, 3871, 4054, 3918, 3824]. Identities [1871]. Identity [2988]. Identity-based [2988]. idle [2804]. IDS [3610]. IEEE [3055, 3469, 3664, 3221, 3623]. if [943]. IFIP [9, 387, 9, 423]. IFIP-Fachwörterbuch [423]. IFIP-Kongreß [9]. ihre [234, 440, 16, 77, 181, 982, 467, 4128, 770, 1308, 614, 282, 482]. II [433, 119, 306, 488, 63, 2497, 2654, 1334, 1836, 2254, 433, 1019, 638, 218, 706, 1451, 1878, 2287, 2543, 1582, 1656, 664, 306, 713, 2327, 612, 546, 800, 2742, 161, 2106, 2104, 681].

IIA [2353]. IID [3913]. III [200, 362, 1640, 63, 604, 1367, 1899]. IIoT [4045]. Ikm [455]. Ill [2084, 1029, 1035, 2738, 2771, 1982]. Ill-Conditioned [1029]. Ill-Conditioning [1035]. Ill-Posed [2084, 2738, 2771, 1982]. Illetschko [250]. Illumination [3725, 3286]. Illumination-robust [3725]. Illustrative [2711]. im [814, 836, 200, 1045, 840, 1131, 783, 867]. Image [1723, 4288, 1730, 4290, 2256, 2303, 2342, 2259, 3811, 2521, 4292, 2027, 2200, 4274, 2489, 4277, 2379, 2636, 2965, 4043, 3358, 3072, 2637, 3800]. Image-Based [2379]. Images [2260, 4267, 2507, 2122, 2374, 4148, 2255, 2379, 1691, 2979, 3249, 4047, 3288]. imbalance [3982]. imbalanced [3618, 3284, 3484, 3825]. Imbedding [1397, 1312, 447, 1324, 1227, 695, 1374, 340]. IMDCS [3757]. Immersing [3424]. immune [3032]. Impact [45, 4080, 3703, 3911, 2909, 2939, 3750, 3008, 3547, 4052, 3794, 3510, 3527, 3465]. Impact-aware [3465]. impacts [3192]. impaired [3137, 3517]. impairments [3620]. Imperfect [435]. imperfections [2987]. implementable [1698]. Implementation [2401, 3310, 1569, 1612, 1808, 1833, 1869, 1189, 2573, 885, 938, 2674, 2489, 1160, 4240, 3423, 3202, 2654, 2903, 4384, 3015, 3817, 2863, 2702, 2032, 1743, 2176, 2976, 3072, 2948, 3083, 2005, 4134, 3036, 2937, 3815, 3262]. implementation-level [3262]. Implementations [2592, 1892]. implemented [2851, 3021]. Implementing [4290, 1699, 2353, 3345, 3548]. Implications [1040]. Implicit [2043, 2179, 2102, 1768, 1833, 1254, 1834, 691, 1927, 1654, 1413, 4330, 1458, 1761, 1988, 2819, 2549, 1398, 2715, 1011, 1452, 1677]. Implicitization [2605, 2572]. implicitly [1711, 1243]. impliziten [1452]. impliziter [1011]. implizites [1254]. Implussysteme [264]. Improper [850, 1288, 1369]. improperly [972]. Improve [1488, 1879, 3430, 3099, 3517, 3235]. Improved [3559, 2286, 2582, 2679, 3945, 3908, 3446, 2703, 3001, 1465, 3284, 1513, 2126, 310, 3297, 3474, 3627, 3587, 4048, 3990, 1700, 3818, 3525, 3522, 3888, 3254, 3149, 1562, 1383, 4082, 3873, 3916, 3871, 3498, 1098]. Improvement [1903, 4303, 1558, 3111, 1545, 442, 1131]. Improvements [1883, 2754]. Improving [3588, 2959, 3012, 3471, 2383, 1384, 3243, 4038, 3053, 1750, 2284, 3240, 3530, 2862, 4028, 4035, 3624, 1457]. Impulses [1624]. Impulsive [2092]. imputation [4078]. Inaccurate [2205, 1516]. incentive [3356]. Including [1955, 1415, 2268, 162, 1517, 871, 1838]. Inclusion [1141, 4175, 468, 469, 2412, 777, 1860, 4383, 2080, 4361, 1550, 2244, 4362, 2514, 2513, 1234, 4222, 1175, 1515, 1542, 1582, 1308, 662, 2014, 4176, 1190, 503]. inclusion-methods [1542, 1582]. Inclusions [4177, 4337, 4355, 4185, 2634, 1720, 1739]. Inclusive [4360, 2143]. Incomplete [438, 2289, 2258, 344, 2305, 1772, 2705, 1516]. Incompressible [2300, 3352]. inconsistent [459]. Incorporating [2867, 3143, 3379, 3905, 3878, 3595, 4013]. increase [3523, 3555]. increasing [3041]. incremental [3751, 3960, 3593]. indecomposability [2737]. Indefinite [1649, 2579, 2304]. Independence [1826, 1161]. Independent [863, 1262, 1293]. Index [4155, 156, 2065, 2172, 1945, 4262, 1765, 3120, 1079]. Indexing [940, 4284, 3328, 3922, 2979]. Indexselektion [1079]. indicate [2799]. indicator [3873]. indicator-based [3873]. indices [2842, 2277]. Indirect [1863, 1270]. individual [4020]. individuals [3137]. Indoor [4008, 3384, 4004, 3727, 3898].

Induced [1941]. induction [2871, 4028, 2877, 2889]. Industrial [483, 1881, 4016, 4040, 3623, 3561]. Industrie [238, 418, 26, 587]. Industriebetrieben [92]. Industry [4217, 238, 4068, 587, 2799, 3539, 3762]. induzierter [1941]. Inequalities [1997, 2313, 772, 1777, 4347, 2070, 661, 2651, 1363, 1143]. inequality [456]. inertial [3554]. Inexact [2269, 2270, 4283, 1823, 2276]. Inf [2395, 2534]. Inf-Sup [2395]. infarction [3289]. infection [3724]. Inference [2060, 3987, 2931, 3801, 4070]. Infinite [1940, 4341, 2692, 979, 499, 3795, 999, 1563, 1193]. infinite-server [3795]. infinitely [558, 613]. infiniten [1940]. Influence [3452, 3830, 3892, 3709, 3565, 3614, 3601, 3996, 3313, 3366, 3766, 3757]. influencers [3767]. influential [3847, 3312]. Influenza [3020]. Informatik [809, 818]. Information [304, 544, 416, 4279, 2026, 103, 540, 4217, 801, 4062, 3299, 3042, 3616, 2692, 3059, 3891, 3375, 3990, 3200, 3098, 3904, 3837, 3439, 1503, 3153, 2935, 3975, 3075, 229, 3477, 3703, 3383, 3860, 2974, 3613, 2989, 3198]. informationization [3280]. Informations [581, 252]. Informations-theorie [581]. Informations-Verarbeitung [252]. Informationssystem [814]. Informationstheorie [303]. Informationsübertragung [326]. Informationsverabeitung [31, 33, 647, 256]. Informationsverarbeitung [29, 30, 37, 32, 38, 43, 392, 423, 158]. Informationsverlustes [1503]. informed [3111]. infrastructure [3781, 3814, 3204, 4010, 2830, 3182, 4053, 3176, 3805, 3423]. Ingenieur [15, 547]. Ingenieur-Mathenatuj [547]. Ingenieur-Statistik [15]. Ingenieure [154]. Ingenieurmathematik [119, 63]. Ingenieurs [362, 488, 287, 656, 226]. ingenieurschulen [216]. Ingenieurwissenschaften [455, 200]. inherent [557]. Inheritance [2081]. Inhomogeneous [2370, 470]. inhomogener [470]. Initial [599, 1141, 1254, 1314, 912, 1315, 1501, 2276, 2132, 2052, 1958, 1151, 877, 1578, 2785, 1175, 1667, 2218, 1748, 988, 995, 445, 2682, 2809, 1989, 1011, 18, 1726, 2041, 1977, 1125, 1339, 1275, 1045]. initial-boundary [2809]. Initial-Boundary-Value [1254, 1315, 2682]. Initial-Value [1141, 1314, 912, 1667, 445, 1011, 18]. Initialisation [2670]. Initializing [1129]. initiatives [3833]. inkblot [950]. Inklusionsisotonie [1308]. inkorrekt [972]. Inner [2337, 1498, 2643]. inneren [1922]. Input [4110, 4172, 151, 952, 3526, 1204]. Input-Output-Modelle [151]. Input/Output [952]. ins [113, 33]. insensitive [1678]. Insertion [1330, 2086, 3331]. insight [3668]. insights [4030, 3426]. Insignificant [567]. Inspection [2446]. inspector [2943]. inspired [3688, 3213, 3335, 4022, 2833, 3845, 3596, 2796, 2799, 2800]. inspried [3305]. instabiler [1999]. instable [1066]. Instance [3306, 3529, 3978]. instance-learning-based [3529]. instances [3178, 3969, 4059, 3241, 3173]. Institutes [579]. instructions [569]. InTec [3993]. Integer [2772, 889, 1551, 752, 644, 3751, 1375, 3205, 2822, 2995, 1362, 3219, 1722]. Integral [1967, 456, 276, 2467, 2502, 824, 2567, 1488, 1830, 870, 1046, 2264, 902, 2537, 1045, 2362, 957, 2368, 1777, 241, 1565, 2511, 893, 2406, 2240, 2485, 2473, 2160, 604, 2465, 1270, 517, 1021, 2664, 2670, 2727, 1323, 1157, 1099, 2680, 704, 1684, 969, 4224, 2613, 1210, 1171, 1209, 1238, 1749, 3358, 470, 1738, 112, 1572, 471, 1764, 1082, 2697, 1201]. Integral-Ungleichun [456]. Integrale [850, 368]. Integralen [1944]. Integralgleichung [112, 1046].

Integralgleichungen [1967, 241, 470, 902, 145, 1082]. Integralgleichungsmethode [1045]. Integralrechnung [218]. Integrals [1378, 1591, 2494, 1612, 1963, 918, 2578, 1827, 1771, 2023, 2071, 2010, 2581, 2038, 1944, 2093, 88, 850, 1527, 1427, 1288, 2037, 368, 1909, 2323, 2527, 2486, 1296, 2727, 1150, 3152, 1662, 979, 1210, 716, 1271, 1351, 2135, 1750, 1392, 1686, 992, 1563, 1369, 143]. integrand [979]. Integrands [2185]. Integrated [3093, 3103, 406, 1442, 952, 3651, 2628, 4016, 2961, 3993]. Integrating [4250, 4009]. Integration [1389, 680, 634, 1993, 1963, 2565, 1962, 2381, 769, 2185, 309, 1875, 854, 1057, 4142, 851, 841, 1820, 164, 2078, 1909, 2323, 2125, 1770, 2104, 1622, 1578, 4025, 1086, 1039, 1687, 1398, 470, 3363, 1194, 3622, 2063, 2021, 3450, 164]. Integrations [2495]. Integrationsformeln [1963]. integrator [1263]. integriertes [952]. Integro [2555, 774, 1830, 1667, 2436, 2750, 1359, 1983, 2661, 2748]. Integro-Differential [774, 1830, 2555, 1667, 2436, 2750, 1359, 1983, 2661]. Intelligence [3688, 215, 302, 3094, 3472, 3043, 3413, 3042, 3464, 3081, 4058, 3726, 3541, 3812, 3718, 3717, 3224, 3972, 3792]. intelligence-aided [3081]. Intelligence-driven [3472]. Intelligence-inspired [3688]. Intelligent [3697, 3637, 3857, 3481, 3107, 3915, 3581, 3909, 3150, 2931, 3636, 3733, 3337, 3225, 3458, 2979, 3724, 2904, 3482, 3539]. Intended [399]. Intensity [2056, 3860]. intensive [2918, 3024, 3325, 2951, 3025]. Inter [4044, 2914, 3662, 3520, 3391]. inter-component [3662]. inter-core [2914]. Inter-mesh [4044]. inter-organizational [3520, 3391]. interaction [2866, 3860]. interactions [3099, 3902]. Interactive [3016, 3067]. Interation [1008]. Intercalation [688]. Interchanges [244]. Interconnecting [2435]. Interconnection [4199]. Interdata [9]. interdisciplinary [3583, 3595]. Interface [2433, 2359, 901, 2468, 1824, 2653, 2654, 2849, 2753, 3067, 1224]. interface-preserving [2849]. Interference [3625]. Interior [1922]. interlinking [3695]. Internal [1434]. International [419, 544, 83, 455, 622, 35, 646, 36, 3782, 42, 66, 40, 237]. Internationalen [200]. Internationaler [455]. INTERNET [83, 3294, 3976, 4085, 3698, 4026, 4042, 3821, 3641, 4006, 3413, 3298, 3481, 3410, 3901, 3117, 2941, 3630, 3884, 3275, 3419, 3656, 2946, 3812, 3372, 3229, 3511, 3311, 2948, 4086, 2945, 3137, 3527, 4077, 3851, 4054, 3295, 3487, 3475]. Internetware [3047]. interoperability [3956, 2898, 2925, 3455]. Interpolant [2450, 2648]. Interpolants [1518, 1839, 1628, 2456, 1770, 1806]. Interpolarity [837]. Interpolating [1028, 1814, 1962, 2196, 2091, 666, 1438, 2708, 944]. Interpolation [1886, 2204, 1102, 2467, 2548, 1448, 2117, 1475, 1771, 1618, 1304, 780, 1956, 1866, 1353, 1559, 2342, 845, 4211, 2505, 1854, 785, 1674, 2035, 4218, 4260, 1614, 1921, 2168, 1842, 2059, 739, 371, 383, 2572, 1964, 1278, 2058, 945, 2619, 2606, 1139, 1744, 2620, 1238, 2017, 2714, 1873, 2596, 1295, 1976, 1736, 1975, 1678, 1215, 1407, 945, 1139, 1956, 499, 1933, 371]. Interpolations [2328, 1053]. interpolations- [1053]. Interpolationspolynomen [1438]. Interpolationsquadraturen [837]. Interpolatory [4319, 1101, 1485, 1053]. interpolierenden [1962]. interpretable [3829]. Interpretation [3363, 1250]. interpretations [1034]. interpretative [3871]. interrogating [2711]. Interrogative [4294]. Intersection [4299, 1923, 2050, 1544, 2530, 2018]. intertwined [2801]. Interval [4097, 1801, 2417, 159, 160, 128, 917, 616, 778, 1028, 4183, 1783, 1998, 2322, 2392, 2578, 1869, 466, 1599, 2105, 1188, 1068, 777, 4101, 1600, 1641, 1857, 1784, 1811, 4172, 1858, 4102, 4103, 4112, 1879, 2131, 724, 458, 1184, 1404, 1415, 1489, 1580, 1647, 4113, 1411, 4114, 436, 4173, 1904, 4343, 4116, 1919, 2814, 990, 1192, 1268, 4363, 2587, 1276, 343, 501, 555, 853, 2049, 1142, 2042, 1239, 1894, 2051, 666, 857, 1009, 1646, 1890, 1606, 2272, 1758, 2054, 2193, 868, 2492, 861, 721, 4124, 1035, 1655, 1727, 1861, 1075, 2811, 692, 1366, 1222, 1211, 2788, 3152, 1695]. interval [1721, 1482, 1176, 4223, 4225, 2806, 1155, 776, 1547, 460, 1656, 1690, 1704, 1152, 1453, 2652, 2044, 4184, 4232, 2815, 441, 1308, 1405, 1576, 499, 2818, 2768, 944, 1140, 473, 2817, 2816, 1180, 1307, 1759, 1701, 1069, 1100, 1127, 1292, 1317, 1256, 1454, 1528, 4236, 2015, 1616, 1664, 1333, 408, 973, 3476, 1757, 999, 692]. Interval-Arithmetic [1904]. Interval-majorant [1758]. interval-Newton-method [1655]. interval-Newton-operators [1576]. Interval-Valued [555, 1256]. Intervall [777, 1655, 1861, 466, 499, 1142]. Intervall-Einschließung [777]. Intervall-Gauß-Algorithmus [1142]. Intervall-Newton-Verfahrens [1655, 1861]. Intervallarithmetik [159, 1308, 868, 1176, 458, 1919, 501, 408]. intervallarithmetische [1075, 343]. intervallarithmetischer [861]. Intervallausdrücke [4101]. Intervallen [4097, 1018, 862]. Intervallfunktionen [1453]. Intervallgleichungssysteme [473]. Intervallgleichungssystemen [778]. Intervallkoeffizienten [973]. Intervallmäßerung [441]. Intervallmäßige [460]. Intervallmatrizen [776]. Intervallmethoden [1998]. Intervallpolynomen [917, 853]. Intervallrechnung [4102, 4103, 724, 436, 990, 616]. Intervallteilungspunkte [1547]. intervallwertigen [555, 1256]. Intervals [2154, 4352, 1188, 862, 1018, 1253, 1321, 1193]. intervention [2709]. intrinsic [2640]. Introducing [3015, 3360]. Introduction [372, 354, 421, 4243, 3018, 126, 337, 360, 2245, 361, 2968, 64, 4139, 253, 685, 3003, 177, 352, 231, 640, 1545, 3115, 538]. Introductory [90, 193]. intrusion [3608, 4016, 3630, 3122, 3516, 3919]. intrusive [3619]. intuitionistic [2728]. invading [1609]. Invariance [19]. Invariant [1598, 1312, 4266, 4326, 806, 4212, 340, 4261, 1558, 1324, 1227, 776]. invariante [776]. Invarianten [1970]. Invariants [1970, 1111, 472]. Invarianz [19]. invasive [3281]. Inventories [333]. Inverse [2173, 2431, 2047, 2723, 107, 505, 1134, 380, 2496, 777, 3227, 1798, 1860, 2519, 2249, 4184, 2520, 2521, 2053, 2524, 956, 4367, 2272, 1862, 718, 678, 661, 1037, 2636, 1515, 1690, 3894, 1022, 1305, 2809, 444, 2771, 381, 1049, 2779, 3271]. Inverse-Monotone [2173]. Inversen [777, 444, 935, 1515, 381, 1037]. inverser [956, 4367]. Inverses [844, 935, 367, 496, 561, 825, 1772, 1539, 1704]. Inversion [2404, 719, 1038, 280, 132, 608, 1487, 2665, 573, 867]. inversive [1991]. Invertibility [2510]. Invertierung [608]. Inverting [1333, 630]. Investigating [4059]. Investigation [476, 1024, 583, 3097, 207, 130, 161]. Investigations [2102, 2186, 3916, 1843, 1118, 953]. investment [3237]. Involving [1786, 2366, 2528, 296, 1275]. Inzinger [106, 106, 605, 1223]. IoMT [3912]. IoMT-enabled [3912]. ion [969]. IoT [3298, 3481, 3294, 4062, 3730, 3781, 3440, 3734, 3398, 3301, 4019, 3643, 3940, 4016, 3861, 3863, 3814, 3785, 3786, 4027, 3347, 3752, 3854, 4087, 3787, 4076, 3729, 3296, 3890, 3964, 3485, 3996, 4003, 3846, 3513, 3534, 3649, 3735, 3650, 3926, 3568, 3997, 3663, 3336, 4052, 3855, 3856, 3868, 3517, 3456, 3864, 3681, 3971, 3660, 3737, 3516, 4065, 3620, 3638, 3450, 3604, 3297, 3798, 3212, 4089, 3919, 3826].

IoT-based [3296, 4027, 3926, 3517, 4065]. IoT-cloud [3729, 3735]. IoT-edge-cloud [3855, 3856]. IoT-enabled [3737, 3730]. IoT-endowed [3212]. IoT-Gov [3681]. IOTA [3413]. IoTBDS [3410]. IoTs [3449]. IoV [3884, 3642]. IPv6 [4085]. IPv6-based [4085]. I’rozesse [262]. Irrational [719]. Irreducible [1059, 586, 638]. irreduziblen [1059]. irreduzibler [586, 638]. Irregular [769, 3914]. irrigation [4065]. ISO [3006]. ISO/TTI [3006]. Isodynamical [2384]. Isolated [2366]. isolation [1198]. isomorpher [1318]. isomorphic [1318]. Isomorphie [740, 617, 1020]. Isomorphism [2073, 803, 4285, 740, 1020, 1428, 617, 1555]. Isoparametric [2204]. Isotone [1647]. isotonic [1190]. Isotonicity [2412, 1308]. ISP [2749]. Issue [3256, 2509, 3162, 3674, 3481, 3042, 3616, 3866, 3170, 3320, 3315, 3464, 3490, 3397, 3410, 3024, 3018, 3034, 3087, 3052, 2601, 2706, 3118, 2968, 2911, 2983, 3881, 3229, 3718, 3717, 3153, 3105, 3605, 3003, 3779, 3209, 3899, 2796, 3381, 3373, 2928, 3294, 3261]. Issues [1808, 1869, 3026, 2891, 3661, 3713]. Item [3689]. items [2902]. Iterated [1073, 1451, 3154, 2104, 1364, 2021]. Iterates [2269, 2575, 1421, 4354, 2778]. Iteration [1856, 521, 882, 1185, 1879, 447, 384, 4361, 519, 2126, 861, 1694, 1778, 1110, 2731, 4159, 1172, 4163, 1515, 2143, 1470, 473, 292, 4233, 3937, 410, 694, 689, 1601, 1696, 1703]. Iterations [1874, 4339, 2103, 1573, 2383, 1415, 4114, 1268, 1517, 1929, 2070, 2636, 1690, 4084, 1184]. Iterationsalgorithmen [614]. iterationsfähiger [473]. Iterationsverfahren [772, 777, 787, 1515, 292, 410, 689, 1008, 24, 477, 23, 861]. Iterationsverfahrens [1110, 519]. Iterative [477, 1558, 2417, 1141, 313, 2275, 427, 2536, 1134, 772, 2086, 777, 787, 1798, 2144, 1719, 4340, 919, 2064, 4173, 1305, 2358, 1848, 1606, 1059, 2354, 802, 4123, 1505, 1175, 521, 1570, 1423, 4225, 1704, 1732, 24, 4160, 1516, 1446, 2841, 2005, 3254, 1454, 1616, 2859, 1221, 1755, 1975, 3798, 297, 1948, 450, 23, 614]. iteratively [2707]. Iterativen [1037, 1059]. itinerary [3552]. IV [238, 510, 602, 87, 238]. Iversen [244, 244]. Iverson [140]. IVP [2810]. IVPs [1908]. IWOBI [3246]. IWT [3414].

J [391, 421, 419, 1793, 1129, 171, 220, 792, 173, 330, 73, 71, 305, 235, 514, 361, 796, 224, 281, 10, 88, 576, 66, 532, 1136, 116, 389, 481, 40, 227, 196, 260, 266, 228, 269, 320, 280, 604, 322, 535, 798, 624, 1714]. J. [232, 533, 185, 68, 289, 427, 484, 674, 15, 148, 350, 264, 328, 401, 329, 94, 194, 303, 729, 1937, 513, 487, 644, 338, 832, 155, 124, 267, 91, 93, 320, 537, 759, 798, 828, 686, 830]. Jacobi [2628, 1340, 380, 57, 2512, 1101, 1335]. Jacobian [220, 1699, 1325, 2424]. Jacobians [1476]. Jacobirotation [57]. Jaffray [91]. Jäger [580]. Java [3022, 3022]. Job [2357, 1294, 1837, 2198, 2826, 1688, 3197, 1979, 3774, 2951, 3553]. Job-Shop [1837, 1688, 1979]. Jobs [1853, 2391, 2476, 1688, 4019, 2002]. John [98]. Johnson [251, 47]. joint [2694, 3290]. Jointly [3580]. Jordan [2324, 4275, 1880]. Journal [3946, 3963, 3981, 4032, 4049, 4073, 4092, 4095, 3361, 104, 3406, 3674]. journey [3407]. July [3138]. June [245, 3138]. Juni [105, 245]. Jurksen [235]. Jury [227]. just [3974]. just-in-time [3974]. JXTA [2934].

K-anonymization [3558]. K. [817, 156, 137, 337, 99, 140, 230, 236, 92, 549, 700, 743, 190, 812, 188, 702, 653, 677, 256, 462, 489, 672]. K.-H. [92, 677]. Kaczmarz [1472]. Kaczmarz-method [1472]. Kaden [424]. Kahan [2546]. Kalman [4024]. Kaltenecker [702]. Kamp [730]. kann [1947]. Kanonische [1969]. kanonischer [2012]. Kanten [2009].

Kanten-Relaxationen [2009]. Kantengewinnen [964]. Kantorovich [72, 4334, 1286]. Kaps [92, 1430]. Karhunen [2675]. Karmazina [199]. Karplus [535]. Karzanov [1118]. Kategorie [38]. Kaufmann [252]. KdV [2784]. Keller [2372]. Kemp [1326]. Ken [815]. Ken-ichi [815]. Kernel [1488, 1730, 1684, 3555, 2892, 3435, 991, 3995]. Kernels [2138, 1830, 2264, 241, 665, 3006, 1738, 3877]. Kernen [665, 241]. Kerner [141, 1671]. Kernfunktionen [118]. Kernreaktortechnik [795]. key [3912, 3643, 2688, 2985, 3357, 3805, 3922]. key-based [3357]. key-hash [2985]. key-value [3922]. keys [3327]. keyword [2978, 3208]. Khattri [2895]. killer [3369]. Kilter [720]. Kimaber [486]. Kind [1957, 824, 1488, 870, 1174, 2160, 604, 2465, 4171, 1282, 2664, 2670, 2727, 1081, 1157, 1099, 1684, 1209, 1238, 314, 1660, 470, 1115, 1201, 88]. Kindler [65]. Kinetic [1855, 2004]. kinetics [2848]. King [142, 1352]. Kirchhoff [2227]. Kirchhoff-Plates [2227]. kite [3247]. Klaba [280]. Klasse [1965, 473, 594, 612]. Klassen [208, 763, 567, 1318]. Klassenzahlen [863]. Klassifizierung [711]. Klassische [364, 467, 888]. klassischen [384, 679]. Klein [143]. Kleines [255]. kleinsten [195, 431, 797]. Klimesch [814]. Klir [328]. Knapsack [1320, 1273, 2284, 1205, 1657, 662, 1092, 662]. knapsack-problem [662, 662]. Kneser [179]. KNN [3804, 3258]. Knodel [1884, 1554, 1884]. Knoll [330]. Knot [1062, 4329]. Knoten [1102, 785, 1332, 807, 2009]. Knoten- [2009]. Knotenbilder [1062]. Knotenmatrizen [1318]. Knotenmengen [1020]. Knots [4246, 1668, 4252, 2449, 785, 522, 523, 1102]. Knotting [4096]. Knowledge [2261, 4255, 3616, 3900, 3033, 2688, 3099, 3226, 2924, 3850, 3462]. Knowledge-Based [4255]. Known [4350, 2593]. Koeffizienten [4097, 1018, 366, 838, 908]. Koeffizientenmatrix [1540]. Koenig [372]. Kogbetliantz [4354]. Kojima [815]. Kolberg [60]. Kollaterale [4132]. Kollokation [1934, 2012, 902, 2010]. Kollokationsverfahren [1332]. Kolloquiums [200]. Komarow [576]. Kommandosprachen [881]. Kommentare [1884]. komplexe [852, 1963]. Komplexen [78, 366, 272, 862]. komplexer [473, 4108]. Komplexität [839, 564, 668]. Komplexitätserfordernisse [1079]. Komponenten [469]. Konditionszahl [822]. Konfidenzbereichs [1918]. Konfidenzbereichs-Verfahren [1918]. konformen [112, 23]. Kong [228]. Kongreß [455, 9]. Konjugierten [780]. Konjugierten-Bestimmung [780]. konsistente [907]. konsistenter [1540]. Konsistenzordnung [1084]. konstanten [838, 908]. Konstituenten [43]. Konstruction [833]. Konstruktion [1003, 872, 1950, 4108, 984]. konstruktiven [980]. Kontaktfläche [855]. Kontaktproblemen [1997]. Kontaktproblems [855]. kontextfreier [1119]. kontextsensitive [312]. kontinuierliches [474]. Kontrolle [467, 888, 364]. Kontrollprobleme [866, 497]. Konvergente [823, 887]. Konvergentes [1038]. Konvergenz [766, 1110, 108, 1438, 924, 1308, 682, 915, 916, 783, 1254, 563, 1948, 1011, 1053]. Konvergenzbeschleunigte [384]. Konvergenzbeschleunigung [910]. Konvergenzfaktoren [893, 163]. Konvergenzgeschwindigheit [296]. Konvergenzgeschwindigkeit [307, 606]. Konvergenzordnung [1655, 445]. Konvergenzordnungen [779]. Konvergenzrate [1920]. Konvergenzverhalten [1107]. konvexe [52, 1921, 1969, 783]. konvexen [766, 889, 879]. konvexer [1920]. Konvolutionssumme [1959]. konvolutiven [1993]. Konzept [4104]. Kopf [502]. Korevaar [66]. Korfhage [229]. Korobov [2599]. Körper [855]. Korrectur [4367]. Korrektur [958, 807]. korrigierbare [363]. Koschmieder [31]. Kowalik [487]. Kraftfahrzeugsströmungen [474]. Krawczyk [1528, 1703]. Krawczyk-like [1528]. Krawczyk-Moore [1703]. Kreis [425]. Kreisarithmetiken [4108]. Kreise [852]. Kreiss [1615]. Kreiss-method [1615]. Krelle [579, 582]. Kreyszig [181]. Kriterien [982]. kritischen [1032]. kritischer [37]. Kronecker [2488, 2464, 2542, 2543, 2614]. Kronecker-based [2488]. Kronecker-product [2542, 2543]. Kronecker-Rank [2464]. Kronecker-sequences [2614]. Kronrod [1963, 144, 1963]. krummlinigem [1918]. Krylov [212, 2222]. Krylov- [2222]. Kshell [3871]. Kubaturformeln [1402]. Kubisch [1038]. kubische [722]. kubischer [959]. Kuipers [331]. Kunsemüller [799]. Kunz [647]. Künzi [14, 426, 579, 357, 582]. Kuo [96]. Kurven [958, 1395, 833, 1916, 1003, 1395]. Kürzeste [461, 278]. kürzesten [791, 930]. kürzester [598, 807]. Kusko [539]. KUTTA [364, 207, 1967, 467, 888, 129, 1492, 1628, 1617, 1358, 1964, 2785, 1967, 1910, 2179, 1532, 1839, 1831, 1954, 1768, 1833, 1988, 1830, 1533, 2384, 1588, 2091, 364, 467, 888, 1491, 2166, 129, 207, 923, 995, 2065, 631, 691, 1747, 1057, 2172, 1989, 1654, 1726, 1339, 1358, 1526, 1677, 1964, 1770, 1083, 725, 1678, 2293, 2748]. Kutta-Formeln [467]. Kutta-Methoden [1967, 1358, 1339, 1526]. Kutta-methods [725]. Kutta-Nystrom [888]. Kutta-Verfahren [129, 1964, 1954]. Kybernetik [203, 254]. Kybernetische [332]. kybernetischer [28]. KYC [3789].

L [391, 1896, 418, 652, 215, 299, 641, 323, 10, 279, 327, 378, 576, 66, 610, 39, 657, 389, 481, 251, 390, 205, 199, 228, 250, 256, 493, 640, 655, 1145]. L-A-S [1145]. L. [44, 72, 331, 271, 261, 262, 157, 122, 610, 540, 33, 47, 187, 217, 125, 127, 324, 258]. label [3588, 3600, 3900, 3446, 3918, 3168]. labeling [1556]. Labelled [929, 1065]. Laboratory [2885]. labyrinth [1031]. Labyrinths [1883]. LAD [2385]. Lag [1820]. lagerhaltung [92]. Lagrange [1139, 2369, 2572, 1407]. Lagrange-Interpolation [1139]. Lagrangian [716, 2123, 666]. Lags [2278]. Laguerre [212, 4363, 4346]. Laguerre-like [4346]. Lamb [3291]. Lambda [1517]. Lambda-Matrix [1517]. Lambert [759]. Lamé [2199, 2533]. Lamprecht [602]. lamps [3972]. LAN [2136]. Lanczos [2738]. Landau [925]. Landis [658]. Landsat [3358]. landslide [3382]. landstraße [80]. lane [3581, 3343, 2607]. Lang [2597]. Lang-Schnurr [2597]. Länge [860]. Langmaack [728]. Language [354, 1145, 180, 3369, 3374, 353, 551, 2765, 568, 569, 3169, 3965, 3373, 140, 290, 279]. Language/One [279]. Languages [420, 533, 1187, 858, 896, 726, 5, 929, 881, 17, 342, 3021, 557, 312, 639, 727]. Lapa [375]. Laplace [280, 2431, 1108, 2735, 719, 1067, 992, 1563, 132]. Laplace-type [1563]. Laplacian [3995, 2466]. Large [2046, 1724, 2256, 2413, 2464, 2033, 2064, 2371, 4213, 55, 3965, 50, 3183, 3780, 3127, 2921, 3547, 1340, 3453, 2949, 2931, 2946, 3218, 3337, 1017, 2970, 3669, 3835, 2835, 3649, 4055, 3258, 2843, 4235, 3705, 566, 3389, 2964, 2890, 2994]. Large-Scale [2033, 3780, 3547, 2946, 3218, 3669, 4055, 2843, 3705, 3389, 2994]. largely [1233]. Largest [468]. Lastdauerlinien [1940]. Lastverteilung [1939]. Latency [3625, 3573, 3961]. Lateness [1294, 1882]. latent [3098]. Lattice [2526, 806, 2346, 2583, 2237, 1622, 1522, 2691, 2729, 2970, 3340, 3245]. Lattices [763, 1875]. Laugwitz [63]. launch [3306]. launch-time [3306]. Lavi [153]. law [2765, 1176, 2869, 3426]. law-enforcement [3426]. Lawes [514]. Lax [985]. Layer [2340, 750, 2263, 3723, 3812, 2209, 3309, 1107]. Layer-Adapted [2340, 2263]. layered [3578]. layers [3730, 2733, 2787]. Laying [2026]. Layout [4188]. Lazy [2039, 3390]. lead [3949]. leading [3668]. leaf [3652, 3934]. leaf-edges [3652]. league [3660]. learned [3228]. learner [3650]. Learning [2757, 723, 3285, 3728, 3744, 3369, 3213, 3966, 3387, 3976, 3857, 3734, 3530, 3987, 3992, 3608, 3741, 3834, 4025, 3230, 4019, 3081, 3627, 3914, 3876, 3861, 3913, 3863, 4039, 3893, 3319, 3751, 3782, 2699, 4078, 3587, 3441, 3928, 4000, 3538, 3591, 3931, 3918, 3980, 3933, 2992, 3399, 4045, 3729, 3529, 3727, 4007, 3828, 3630, 4026, 3375, 3894, 3687, 3500, 3390, 3555, 3636, 3994, 3904, 4024, 3597, 3993, 3284, 3411, 3435, 3667, 3774, 3858, 3944, 3996, 3970, 3973, 4074, 4068, 3742, 3229, 3967, 3303, 3584, 3554, 3867, 2844, 3770, 3793, 3650, 3875, 3309, 3975, 3855, 3856, 3868, 3556, 4134, 3282, 3936]. learning [3864, 4082, 4053, 3760, 3267, 4034, 4040, 3377, 3720, 3640, 4008, 3380, 3561, 3509, 3373, 3978, 3486, 3289, 3383, 3829, 3860, 3882, 3799, 3063, 3046, 4014, 3905, 3288, 3290, 3777, 3934, 3762, 4029]. learning-based [3734, 4019, 3980, 3933, 3894, 3996, 3864, 3383, 3799, 3934]. Least [844, 1128, 4252, 2011, 2297, 1835, 2249, 2369, 1787, 1029, 1342, 1371, 1593, 2133, 2302, 209, 2389, 2170, 3744, 2707, 3802, 1707, 1547, 1651, 1476, 2715, 1986, 1069, 1737, 973, 2764, 1036]. least-mean [3802]. Least-Square [1128]. Least-Squares [1787, 2133, 2302, 2170, 2707, 1476, 1986, 2764, 1036]. Least-Squares-Lösungen [1036]. leaves [3250]. Lecture [426, 650, 648, 649, 627, 700, 581, 744, 628, 530, 534, 582, 651, 672]. Lectures [540]. LEDA [4289]. ledger [4062]. Lee [377]. Leg [2350]. Legendre [77, 2594, 77, 258]. Legendre-Polynome [77]. LEGOL [1265]. LEGOL-2 [1265]. Lehrbuch [510]. Lehren [186, 286]. Lehrer [238, 418]. Lehrmaschinen [28]. Lehrprogramme [569]. leicht [288]. leichte [456]. Leipholz [403]. Leistung [749]. leistungsfähiger [791]. Leistungsgrössen [1409]. Leitfaden [404, 403, 357, 10]. Leitungstätigkeit [248]. Lekhtman [138]. lemma [2088]. lemniscatic [1019]. lending [2833]. Length [2061, 860, 2457, 2058, 2723, 3802, 2770, 1974, 3646, 3479, 2712, 2972, 2612, 1412, 1272]. Lengths [807]. Lens [2713]. Lens-shaped [2713]. Lentin [670]. LEO [4044, 3768, 3839]. LEO-multi-access [3839, 3768]. Leone [251]. Lernen [186, 286, 338]. lernender [198]. Lernprozesse [650]. Lernschritte [654]. Lesion [3728]. LESPP [3102]. less [1370]. Letter [2896, 2895]. Lev [3260]. Level [2275, 3113, 2368, 2387, 1267, 2075, 2356, 2212, 1380, 2849, 3004, 1765, 1695, 3056, 3817, 2506, 2635, 4160, 4038, 3427, 2209, 1011, 2637, 2751, 3262, 2749, 2973, 2878, 2940]. level-index [1765]. level-set [2637, 2878]. Levenberg [2491, 1403, 4348]. Levenberg-Morrison-Marquardt [1403]. leveraging [2804, 3757]. Levering [2924]. Lévy [2442]. lexicographically [1293]. Lexikon [255]. LGAT [3885]. Liapunov [1995, 1995, 982]. Liapunovstabilität [982]. Libraries [486]. library [2044, 4239]. license [3697]. Lidstone [522]. lie [943, 2683]. Liebling [649]. liefert [777]. Lienard [446]. Liénardschen [446]. Life [4083]. lifecycle [3391]. Light [3040, 3022, 3258, 3885, 2980]. Light-weight [3040, 3022]. lighting [4015]. Lightweight [3478, 4062, 3643, 3060, 3102, 3805, 4009].

Like [1809, 2244, 4187, 2690, 3161, 2028, 1698, 4358, 1818, 2848, 3736, 4346, 4362, 2513, 1645, 2550, 1528, 2813, 3653, 2149, 1445]. Likelihood [4278]. Limit [1432, 2516, 2486]. limitation [2765]. limitations [4094]. limited [3591]. limits [3748, 470, 446]. Line [2270, 2409, 2260, 1400, 4301, 1045, 2250, 2307, 685, 4365, 2210, 2530, 2540, 1947, 2019, 2476, 2840]. Linear [844, 2341, 465, 2417, 1934, 2555, 2366, 2552, 1397, 1477, 875, 1207, 1289, 289, 2206, 1047, 36, 772, 2166, 706, 4317, 4267, 1639, 4339, 2411, 2011, 2057, 748, 1835, 2464, 2481, 4100, 1499, 838, 2181, 1777, 641, 1857, 4172, 1867, 1858, 820, 878, 2242, 2000, 724, 627, 241, 919, 2369, 2203, 1635, 309, 1290, 4173, 4343, 1174, 1932, 1030, 4344, 2510, 990, 562, 2318, 448, 2695, 2516, 2326, 2186, 2471, 1307, 1239, 2051, 4120, 4164, 1890, 7, 2562, 1195, 839, 117, 895, 1606, 2396, 2272, 4275, 1593, 247, 825, 1460, 2175, 4281, 209, 1135, 4263, 789, 1219]. Linear [4171, 2389, 2350, 2070, 1388, 1250, 2907, 1018, 1110, 2782, 517, 1282, 2731, 1745, 1066, 2727, 2696, 1097, 1324, 2217, 1089, 1721, 2606, 1482, 1117, 989, 1264, 632, 2095, 1540, 1340, 3751, 396, 690, 2806, 584, 1689, 2114, 1108, 1620, 2652, 1650, 382, 2874, 1017, 2783, 1200, 384, 1516, 2721, 1115, 1872, 450, 3, 2809, 452, 1572, 1759, 1706, 1701, 4235, 500, 413, 1977, 1144, 1257, 1341, 1737, 973, 1221, 962, 1452, 459, 2650, 1322, 1240, 2598, 550, 1424, 2807, 3219, 1145, 1547, 613, 398]. Lineare [580, 1934, 772, 760, 4100, 820, 878, 2000, 990, 653, 500, 839, 613, 398]. Linearen [1018, 1932, 1110, 880, 27, 2011, 264, 382, 724, 384, 3, 744, 1030, 452, 628, 895, 209, 789]. linearer [465, 325, 1089, 632, 1540, 396, 690, 706, 838, 241, 413, 1547]. Lineares [291, 7]. linearizability [3316]. Linearization [2249, 2282]. Linearized [1767, 2156]. Linearly [2179, 2033, 1381, 1729, 1743, 2176, 1677]. Linearly-Implicit [2179]. Lines [1804, 1466, 3431, 2881]. linguistic [2794, 3796]. Linguistics [354, 43]. Linguistik [670]. Link [3903, 3902, 3557, 3107, 3894, 3904, 3631, 4028, 3837, 3019, 3899, 3476, 3703, 3462, 2956]. LinkD [3695]. linked [3430, 3695, 3622]. LinkedIn [3161]. linking [3169, 3106]. LINKORD [3019]. links [3387, 3040, 3623]. Liouville [1740, 2053, 1244, 340, 1895]. Lippmann [33]. Lipschitz [2680, 1404, 1541]. Lipschitzian [4338]. liquid [2909]. List [786, 401, 1032, 3522]. list-based [3522]. listed [3674]. Listenstrukturen [786]. Literature [4144, 3789, 3897, 3428, 3303, 3663, 3792]. live [1459, 3445, 3575, 3980, 3933, 2836, 3423]. Liveness [3677]. living [3529]. Ljapunow [203, 254]. LLNs [3456]. Load [1940, 1881, 2837, 3217, 2273, 3530, 3398, 3441, 4000, 3200, 4060, 3442, 3342, 3760, 3826, 3827, 2840, 3268, 3722]. Load-aware [3217]. load-balancing [2840]. loading [3331]. loads [3911]. Lobatto [1934, 1934, 1640]. Lobatto-Kollokation [1934]. Local [1618, 4249, 2107, 2733, 2202, 4161, 2344, 2029, 1948, 1829, 4214, 1164, 3825, 1486, 2718, 3338, 2664, 1418, 1822, 3853, 1403, 2656, 2998, 1280, 1452, 3703, 3420]. Locality [3220, 1131, 3197, 3708]. Locality-aware [3220, 3197]. localization [3384, 4004, 3253, 3727, 4028, 1616, 3232, 3291]. Localized [4251, 2522]. Localizing [4321]. Locally [2173, 2134, 2057, 3053, 2787, 3825, 275]. Locating [2077, 1908, 1507, 1584]. Location [240, 2124, 2723, 3269, 3200, 3585, 3659, 3053, 2995, 3278, 3028, 2975, 3799, 3487]. location-sensitive [3278]. locations [2845, 3058]. Lochkartenverfahren [189]. lock [3680, 3041, 3843]. lock-free [3680, 3843]. Locking [2152, 2760]. Loeb [846]. Loeckx [729]. Loève [2675]. Log [926, 4265, 1463, 1633, 2007, 3955]. log-concave [1463, 1633, 2007]. Log-Normal [926]. Log-Polar [4265]. Logarithmic [2577, 1496, 1488, 2167, 2486, 1497, 1852, 703]. logarithmischer [1852]. logarithms [1705, 2685]. logging [3299]. Logic [771, 406, 229, 2761, 2060, 165, 629, 2068, 3523, 2797, 3146, 4021, 3262]. logic-based [3523]. Logical [374, 421]. Logics [882, 1172]. Logik [263]. logische [146]. Logs [3270, 3127, 3510]. Lohnes [538]. lokal [1948, 275]. Lokale [1452]. Lokalität [1131]. Long [2563, 3394, 3420, 3927, 2768, 2813, 3273]. long-distance [3273]. Long-term [3420]. look [2917]. Loop [2447, 886, 962, 1970]. loop-chains [962]. Loop-Subdivision [2447]. Lorentz [36]. Lorenz [4237]. Lornende [175]. Lösen [186, 286, 410]. Loskett [280]. loss [1503]. Lossless [2259, 3069]. Lössungen [912]. lossy [4085, 3438]. Lösung [454, 599, 778, 1915, 904, 1994, 1254, 1632, 396, 988, 855, 1045, 838, 1968, 965, 1917, 382, 385, 450, 933, 1503, 1030, 112, 1225, 497, 866, 1000, 1011, 18, 860, 414, 1107, 877, 1125, 495, 689, 984, 2013]. Lösungen [1540, 1013, 1999, 849, 503, 606, 707, 182, 446, 1082, 262, 1036, 1958]. Lösungsfunktion [470]. Lösungsmethode [500]. Lothar [439, 439]. lots [3513]. loudspeaker [2873]. Low [2408, 2467, 4315, 1617, 2464, 2542, 2543, 2185, 1581, 3588, 3004, 4085, 2613, 3727, 3898, 1490, 3342]. Low-Discrepancy [4315, 2185]. low-effort [3898]. low-power [4085]. Low-Rank [2408, 2467, 2542, 2543, 2613]. low-significance [3588]. Lower [3188, 467, 2019, 1573, 2282, 2559, 839, 2679, 3536, 1924, 2624, 1627, 1740]. Lower-bound [3188]. Lowest [2349, 2704, 3352, 2171]. LP [3751]. LP-relaxation [3751]. LPs [1579]. LPT [1853]. LPT-Algorithm [1853]. LSTM [3840, 4025, 3880, 3927, 3958, 4091, 3818, 3468, 4063, 3476, 4013, 3919].

LSTM-based [4063]. LU [2501]. Lukacs [536]. LUMF [4009]. LUMF-YOLO [4009]. Lumping [2377]. Lutz [814]. Lyapunov [2585]. Lyon [4379]. lysine [3293].

M [391, 603, 377, 328, 120, 401, 649, 103, 11, 279, 576, 549, 42, 487, 266, 135, 284, 326, 728]. M. [246, 419, 134, 330, 225, 69, 169, 36, 652, 360, 796, 625, 670, 532, 192, 265, 170, 199, 115, 183]. MAC [2953, 3664, 3924]. MacCormack [2119]. Machine [159, 128, 3734, 3987, 1294, 2278, 77, 2470, 556, 215, 302, 2307, 3630, 3636, 2198, 2593, 3720, 3374, 3387, 3976, 3608, 3741, 3230, 2217, 4019, 2414, 3861, 2756, 3263, 3575, 609, 3441, 3928, 3538, 3931, 3980, 3933, 3729, 3727, 4007, 3555, 3597, 3993, 3973, 3742, 3229, 3303, 3584, 3554, 3770, 3888, 3151, 3855, 3856, 3556, 3864, 1070, 4136, 4082, 4053, 3769, 3640, 3424, 564, 668, 3376, 3289, 3829, 3423, 3869]. machine-based [3424]. machine-learning [3555, 3289]. Machine-Technical [77]. machinelle [189]. Machines [2357, 1837, 2232, 2540, 299, 2250, 2420, 170, 2593, 3164, 507, 3085, 3459, 3125, 3100, 502, 560, 3190, 681, 374, 551, 655]. Maechanik [357]. Maess [182]. Magnetic [539, 2114, 2877, 2889]. Magnetohydrodynamic [516]. magnets [2877]. main [3661, 3067, 3084]. main-memory [3067]. maintainability [62]. Maintaining [3682, 3345]. maintenance [4015]. Major [622]. majorant [1758]. majority [3992]. Makespan [2198, 3887, 3550]. making [3502, 2794, 3535]. Malcolme [514]. Malcolme-Lawes [514]. malicious [3863, 3938]. Malov [138]. Mammitzsch [797]. manage [3089]. Management [1170, 126, 248, 3641, 3651, 3298, 3857, 4046, 3912, 3915, 3093, 3643, 3339, 3110, 3165, 3953, 3200, 3419, 3296, 3636, 3048, 3512, 3652, 3166, 3888, 3917, 3078, 3195, 3864, 3548, 4053, 3077, 3142, 3972, 3551, 3429, 3095, 3073, 3539, 2994, 3198]. Management-Datenverarbeitung-Operations [248]. Managing [3103, 3096, 2845, 3058]. Mandel [94]. Mandelbrot [2765]. MANET [3335, 2936, 3673]. MANETs [3107, 3023]. MANIFESTO [3617]. manifold [2699]. Manifolds [2531, 1885]. Manipulation [1944, 1247, 614]. Mann [543]. Manteuff [653]. Manual [121]. Manufacture [4244]. manufacturing [3907, 3687, 3667, 4068, 272, 3539]. Many [2110, 1386, 3765, 805, 558, 3001, 3003, 3873, 3877]. Many-BSP [3877]. many-core [3001, 3003]. many-objective [3873]. Many-to-many [3765]. Map [2522, 3858, 2995]. MAPE [4072]. Mapped [2395]. Mapping [3974, 1062, 4083, 1396, 1929, 1418, 3408, 3670, 4094, 3204, 2663, 3031, 3304, 3203, 2615, 3049, 112, 3663, 4021, 3501]. Mappings [766, 2247, 295]. MapReduce [3065, 3028, 3318, 3074]. Maps [4287, 1560, 272, 23, 1275]. March [4378]. Marchl [158]. marginal [3614]. Margulis [117]. Mark [259]. market [3322, 3237]. marketplace [3483]. Markhoff [768]. Markle [654]. Markoff [243]. Markoff-type [243]. MARKOV [378, 4060, 4342, 840, 3562, 967, 978]. Markovian [2488, 3428]. Markovprozesse [840]. MARLFuzz [4040]. Maron [509]. Marquardt [2491, 1403, 4348]. Martensen [701]. Martin [655]. Maschinelle [37]. maschinellen [1070]. Maschinen [652, 332, 502]. Maschinenintervallarithmetik [159, 128]. maschinentechnische [77]. mask [2745, 3728]. Mason [1143]. Mass [209, 3234]. Massive [2709, 3048]. Massively [2100, 2169, 3694]. Maßtheorie [46]. Matazke [200]. Matching [2262, 2082, 2061, 4283, 1551, 684, 4318, 4127, 2303, 775, 1832, 4284, 1974, 2678, 3765, 3287, 3768].

Matchings [1551, 1652]. Matchmaker [723]. material [2876]. materials [2874]. math [488]. Mathematica [2863]. Mathematical [335, 354, 422, 134, 69, 426, 536, 283, 1939, 650, 648, 649, 815, 627, 700, 581, 42, 243, 628, 540, 34, 530, 534, 582, 651, 4272, 233, 862, 478, 672, 1347, 2876, 744, 39, 2879]. Mathematician [811]. mathematicians [238]. Mathematics [375, 455, 646, 675, 36, 1811, 281, 40, 4221, 35, 238, 54, 42, 2509, 75, 61, 170, 213]. Mathematik [455, 16, 200, 238, 418, 54, 675, 357, 10, 576, 66, 154, 216, 219, 300, 403, 831, 404, 331, 95, 324]. Mathematiker [238, 418]. Mathematische [171, 173, 271, 226, 287, 362, 488, 656, 670, 190, 187, 334, 546, 676, 489, 579, 646, 576, 263, 191]. mathematischen [674, 1347, 1939, 357, 362, 701, 581, 389, 481, 390, 862]. mathematischer [186, 286, 156, 478]. Mathenatuj [547]. Mathews [532]. Matlab [2401]. Matrices [108, 1856, 2408, 1936, 1255, 1569, 2366, 2502, 1390, 2046, 735, 505, 1109, 859, 914, 2421, 2413, 2472, 2500, 2245, 2390, 2483, 922, 2383, 2508, 1858, 2127, 2242, 2298, 5, 2324, 2424, 1912, 2042, 921, 2550, 718, 1772, 1297, 822, 899, 1539, 1035, 307, 1683, 2642, 4175, 2734, 2740, 1222, 717, 847, 586, 1237, 472, 1196, 1530, 776, 2675, 3437, 1627, 608, 393, 2418, 2721, 463, 1236, 1572, 1318, 1706, 1672, 630, 1049, 1311]. Matrix [2047, 178, 2585, 2043, 2552, 2567, 1521, 2384, 1134, 4339, 2413, 2245, 2287, 57, 777, 787, 1798, 1860, 4340, 442, 625, 2503, 938, 2203, 2466, 1775, 928, 4188, 2371, 4361, 4345, 2499, 1966, 1894, 553, 1517, 7, 22, 1862, 2054, 953, 1059, 868, 495, 573, 867, 244, 661, 4079, 1066, 1756, 2639, 1540, 638, 2686, 1515, 1690, 1704, 3463, 3437, 1404, 2747, 1590, 2612, 1305, 2618, 2625, 1067, 1487, 1562, 2638, 4240, 2779, 1424, 1515, 953, 1059, 867]. Matrix-Algorithmus [495]. Matrix-Dependent [1966]. matrix-fields [2639]. Matrix-free [2567, 3437]. Matrix-Multiplikation [953]. matrixabhangigen [1966]. Matrixdarstellungen [787]. Matrixintervalls [777]. Matrixnormen [57]. Matrizen [1683, 638, 404, 307, 108, 1936, 586, 859, 914, 608, 393, 442, 1236, 553, 868]. Matroids [1337]. Matyäs [322, 3893]. Matyas-Meyer [3893]. Max [2443, 2009, 679]. Max-Cut [2009]. Maximal [2239, 2315, 2492, 822, 1262, 2683, 684, 1118, 298, 3998, 1293, 2139, 1078, 1453]. Maximale [1453]. maximalen [684, 298]. Maximaler [822]. maximally [963]. maximisation [3601]. Maximization [2279, 3565, 3614, 3830, 3996, 3366, 3766, 3028, 3448, 3757, 3892, 3709]. maximize [3586]. Maximizing [2105]. Maximum [1864, 2043, 1294, 1802, 1013, 1277, 775, 767, 2003, 1089, 2678, 1640, 1616, 694, 1399]. maximum-minimum-principle [1013]. Maximum-Minimum-Prinzip [1013]. Maximum-Principle [1277]. Maximumprinzips [694]. Maximumsnorm [1089]. Maxwell [2348, 2438]. MB* [4213]. MC [3021]. MC/DC [3021]. McCalla [253]. McCormick [388]. McCracken [87, 13, 359]. Mconvkgc [3850]. McRae [45]. MCRDR [2991]. MDER [3601]. MDPCluster [3389]. Mealy [133]. Mean [1644, 2363, 1211, 660, 1419, 2365, 2056, 1455, 3802, 2794]. Meaning [110]. Means [21, 726, 2190, 789, 716, 1574, 1060, 3342, 53]. Measure [1409, 4300, 368, 1035, 3488, 3338, 3889]. measured [316]. measurement [3253, 3988]. measurements [4129, 2947, 337]. measures [3178, 3312, 2728]. measuring [3719, 3158]. Mechanical [1910, 4328]. Mechanics [237, 301, 237]. Mechanik [10, 404, 173]. Mechanisierung [434]. mechanism [3680, 3571, 4066, 4068, 3049, 3378, 3883, 3602, 3934, 4011, 3074]. Mechanisms [3216, 3816, 3169, 3106, 3176]. Mechanization [434]. mechanized [3260]. Meddal [3708]. Media [2292, 3360, 3647, 3618, 3767, 506, 3078, 3071, 4013]. media-optimized [3078, 3071]. Medial [2177]. Median [2629, 2739, 2206]. mediators [2915]. medical [2934]. medium [1609]. Medizin [745]. medoids [3390]. Meerov [183]. Meet [4310]. meeting [3708]. Mehrdeutigkeit [457]. Mehrdimensionale [593]. mehreren [53]. mehrfachen [1332]. Mehrfachregelungen [428]. Mehrgittermethode [1997]. Mehrgitterverfahren [1955, 1966]. Mehrprozessorsysteme [1409]. Mehrschrittverfahren [907, 779, 1011, 851, 976]. Mehrstufige [1356]. Mei [1805]. Meinardus [16, 285, 324]. Meixner [260]. Melentjew [358]. Mellin [1462]. Mellin-Transformation [1462]. member [3558]. Membership [3801]. membrane [1426]. meme [3313]. memoriam [1937]. memorizing [3632]. Memory [2503, 2027, 3041, 3675, 3056, 3748, 2803, 3067, 2938, 3927, 3049, 3084]. Menes [332]. Mengen [4106]. Merbach [627]. merge [3391]. Merges [1087]. Merging [2842, 2772]. Mertens [818]. Mertin [699]. Meschkowski [118, 95]. Mesh [1826, 1803, 2253, 2177, 2137, 2124, 1933, 3055, 1161, 2709, 2254, 4044, 2999, 3309, 3217, 2178]. Mesh-Connected [2177]. Mesh-Independence [1826]. Mesh-Refinement [2253, 2254]. Meshes [2439, 4318, 2340, 2424, 1996, 2459, 2218, 2733, 2865, 2620, 2870, 2616, 2787, 2632]. meshless [2784, 3808]. Meß [284]. Meß- [284]. message [3257, 3885]. message-passing [3257]. messages [3299]. Messen [123]. Meßverfahren [222]. Messwerte [209]. meta [3565, 2826, 3334, 3896, 4043, 3303, 2814, 3596]. meta-heuristic [3565, 3334, 3896]. meta-heuristics [3303]. meta-method [4043]. meta-scheduling [2826]. metacomputing [3592]. metaheuristic [3909]. Metel’skii [212]. Method [1595, 2055, 2062, 2401, 766, 1886, 160, 890, 2518, 880, 1804, 599, 2146, 443, 498, 2092, 2375, 2211, 1997, 4166, 1995, 2399, 2134, 1910, 904, 2262, 2275, 875, 2179, 1783, 2227, 2161, 4338, 1051, 2115, 1047, 738, 1901, 2376, 846, 772, 1254, 2216, 2147, 887, 2491, 2107, 2118, 129, 366, 1940, 2297, 1045, 1767, 1524, 1828, 2475, 2563, 1878, 1877, 4321, 76, 4354, 2512, 777, 922, 4355, 1600, 1641, 2129, 370, 1624, 2359, 2568, 2519, 1604, 4172, 1666, 2301, 4306, 1917, 901, 1348, 2247, 241, 2374, 919, 4358, 447, 1465, 1315, 2404, 1850, 2520, 2288, 1310, 1116, 1567, 2349]. Method [1904, 1818, 1823, 4325, 1030, 1368, 1922, 1906, 1855, 2268, 2510, 1425, 866, 2371, 242, 2432, 55, 2318, 1391, 1000, 2241, 2358, 1550, 2244, 163, 1863, 1537, 2524, 2588, 4347, 2049, 2036, 2437, 3362, 1848, 4260, 1058, 22, 606, 707, 1435, 4170, 2571, 1059, 2075, 2465, 2580, 2356, 2305, 2197, 871, 1466, 1838, 133, 4171, 783, 2389, 688, 2113, 4348, 2589, 297, 1216, 1734, 2785, 3430, 1, 307, 1018, 1110, 1655, 1861, 2782, 1161, 3857, 1472, 6, 3598, 1907, 3992, 1242, 1149, 1436, 3557, 2633, 2664, 2670, 2727, 1756, 1615, 2651, 1418, 1097, 1212, 704, 1792, 1836, 1467]. method [2788, 1684, 1695, 1721, 2982, 969, 1274, 3523, 1533, 1117, 1447, 4223, 1340, 1431, 1707, 396, 2669, 3239, 1712, 1699, 1735, 975, 986, 2784, 1892, 3204, 584, 506, 1061, 1080, 1074, 4161, 2643, 1196, 3549, 2863, 1068, 2649, 1349, 3983, 3854, 1515, 1656, 1704, 2032, 595, 3927, 1698, 1209, 3907, 1152, 3723, 1585, 1661, 2795, 3375, 3823, 1700, 3007, 947, 1384, 1560, 1650, 1891, 24, 617, 1214, 3658, 608, 2870, 2746, 385, 3287, 2791, 1398, 2668, 2656, 2732, 2735, 2970, 3293, 2140, 3719, 1709, 4043, 4057, 395, 1071, 1359, 384, 1230, 1470, 1422, 1663, 1286, 3050, 3427, 1725, 4086]. method [4005, 1015, 497, 23, 2655, 1266, 3484, 3309, 410, 519, 3673, 1352, 1645, 1136, 1067, 3080, 1572, 2626, 3824, 1701, 1728, 2646, 2771, 2761, 18, 500, 527, 1038, 520, 1454, 414, 3201, 3352, 1616, 1758, 3660, 1708, 934, 2859, 3415, 1755, 630, 1757, 2726, 1976, 1670, 1219, 1697, 689, 1601, 3576, 3776, 1671, 1483, 132, 3882, 2786, 1457, 3799, 3883, 2789, 3385, 2764, 3475, 3796, 1696, 3715, 1713, 3515, 2878, 1408, 889, 914, 2573, 805, 227]. Methode [890, 366, 195, 431, 797, 1917, 1030, 163, 133, 396, 986]. Methoden [1967, 509, 27, 238, 646, 181, 226, 357, 701, 300, 334, 546, 1358, 676, 854, 1996, 1339, 1526]. Methodenlehre [185]. Methodenunter [185]. Methodik [4128]. Methodology [137, 3503, 2915, 3654, 2710, 3177, 4065, 3158]. Methods [2426, 736, 4110, 788, 375, 907, 477, 823, 2417, 1464, 1389, 108, 1840, 4350, 1856, 4335, 1967, 1955, 2094, 1301, 836, 4337, 774, 2488, 2102, 1998, 2434, 2322, 1831, 1954, 2565, 1930, 2576, 1768, 2100, 1833, 1830, 4299, 2294, 2384, 1598, 2180, 1568, 2440, 1869, 261, 2216, 2405, 2155, 1588, 2091, 779, 207, 1628, 1617, 4251, 1797, 1928, 1429, 2071, 2226, 2549, 2353, 755, 819, 2506, 4322, 787, 1860, 4383, 2368, 2065, 2538, 924, 641, 2495, 1849, 1968, 2515, 4340, 2295, 1776, 2310, 691, 1185, 1679, 1314, 2487, 955, 2090, 301, 2306, 13, 327, 912, 2435, 2029]. Methods [2097, 1832, 682, 2064, 1169, 2255, 1775, 1381, 4342, 1605, 854, 933, 750, 1373, 4116, 2276, 2172, 253, 1124, 851, 1785, 2240, 34, 2084, 1949, 2132, 4363, 4346, 4362, 2514, 2513, 2186, 487, 530, 651, 2523, 2087, 2052, 864, 1164, 4218, 96, 935, 956, 4367, 1414, 1234, 1606, 1820, 2108, 2149, 2302, 93, 233, 759, 2078, 1413, 561, 1964, 802, 1270, 2237, 4123, 828, 2074, 2462, 1008, 1416, 2474, 1602, 1168, 2343, 2300, 861, 753, 1918, 2293, 2350, 1278, 1380, 2034, 1388, 1589, 1250, 1138, 1233, 2707, 1044, 1727, 1282, 1129, 1075, 4222, 2705, 2731, 432]. methods [3148, 2634, 4372, 185, 56, 1157, 1099, 238, 2636, 1532, 1792, 1836, 1253, 1988, 741, 2672, 1423, 988, 1364, 4225, 923, 2806, 995, 2860, 2865, 2908, 1084, 4226, 1098, 1542, 1582, 971, 1574, 1085, 2635, 2701, 1474, 1299, 2652, 1229, 2873, 1660, 2704, 1747, 3835, 3371, 4160, 473, 1516, 4162, 1115, 571, 1343, 1260, 292, 1077, 1446, 1577, 1332, 2209, 1246, 1603, 1011, 1194, 1328, 3732, 413, 1726, 1710, 1813, 4168, 1575, 3393, 1977, 1002, 1221, 694, 976, 1125, 1339, 1358, 1452, 1526, 1677, 1279, 1275, 1096, 1160, 2650, 1083, 725, 1012, 1678, 2222, 984, 1235, 1984, 1445]. methods [2748, 1703, 1430, 178, 285, 12, 586, 638, 64, 296]. Metric [736, 1381, 436, 1164, 440, 451, 1669, 1561, 3189]. metrics [3492, 3714, 4039, 3479]. metrische [440, 436]. metrischen [451]. Mexico [3503]. Meyer [3893]. MFS [2663]. MGOPT [2726]. MHD [2889]. MICAI [3224]. MICDTN [3571]. Mice [1883]. Michel [146]. Michie [215, 302]. Micro [3406, 4134, 3508, 3916]. micro-grid [3916]. Micro-journal [3406]. Micro-modules [4134]. micro-services [3508]. microdata [1503]. microgrid [3595]. microgrids [3133, 3809]. Micromagnetics [2499]. microprocessors [3008]. Microservice [3713, 3783, 3964, 4081, 4041, 3775, 3567, 3952]. microservice-based [3567]. microservice-oriented [3783, 4041]. Microservices [4083, 3949, 3923, 3354]. Microstructures [2434]. middleware [3447, 3093, 2996, 3140, 2981]. Midpoint [1927, 1458]. migration [3445, 3575, 4090, 4037, 3980, 3933, 3085, 4010, 3125, 3640, 3423, 3139]. Mikhlin [34]. Mild [1826]. mildew [3250]. Miller [532, 231]. millimeter [3926]. MIMD [2200]. mimic [3692]. min [3417, 679]. Min-max-principles [679]. Min-Max-Prinzipien [679]. miner [3230]. Minimal [2582, 2110, 2260, 1518, 1668, 1763, 754, 1123, 2281, 2351, 848, 2024, 2000, 1485, 2135, 572, 860, 1195, 1820, 4107, 1008, 2770, 1102, 989, 339, 987, 1103, 1295, 1026, 1055, 2479]. Minimal-Length [860]. Minimaleigenschaft [1102]. minimalen [1503, 4107]. minimaler [293, 848, 987, 860]. minimales [1008]. Minimalglied [572]. Minimalisierungsverfahren [56]. minimalist [2644]. Minimallösungen [339]. Minimally [1523]. Minimax [1773, 1309, 1381, 459, 1561]. minimierende [2000]. Minimierung [1008, 1082]. Minimierungsaufgaben [783]. Minimization [1122, 2251, 2725, 2033, 1310, 2443, 939, 1649, 2198, 703, 496, 678, 783, 2849, 56, 1097, 1467, 2631, 1743, 2176, 1984, 1008]. minimization-based [2849]. Minimizer [1829]. Minimizing [1294, 735, 2147, 2303, 1166, 2709, 717, 3646, 1650, 3125, 1082]. Minimum [3101, 2308, 1013, 1782, 4282, 1552, 293, 1652, 3184, 1503, 3415, 2718, 1439, 1191]. minimum-risk [3415]. Mining [3210, 2364, 3585, 3406, 3471, 2944, 3529, 3082, 3215, 3439, 3835, 4063, 3920, 2761, 3209, 2792]. mining-oriented [3920]. Minkowski [1522]. Minkowski-reduced [1522]. Minmax [2223]. Minsky [652]. Minty [346, 346]. Miranda [4334, 2469]. Miranker [1896]. mirrorsymmetric [2747]. Miscenko [271, 187]. Missing [3837, 3817, 4078, 3332]. Mitarbeit [756]. mitigate [3136]. Mitigating [4007]. mitigation [3974, 3495, 3122, 3970, 2940]. Mitteilung [83, 455, 113, 386, 9, 59, 669, 84]. Mitteilung/Note [83, 113, 386, 9, 59, 84]. Mitteilungen [245]. mittels [912, 1032, 272, 1919, 614, 1107, 1125, 1452, 1952]. mittlerer [749]. Mix [454]. Mixed [4313, 889, 619, 2563, 4163, 2123, 2387, 1465, 849, 2437, 4029, 3611, 4070, 3242, 2735, 2667, 1725, 2822, 1986, 2646, 2878]. mixed-criticality [3611]. mixed-hop [3242]. mixed-integer [2822]. mixed-mode [2878]. mixed-precision [4070]. mixes [2939]. Mixing [454]. MKP [2927]. ML [3686, 4080, 3960, 3443]. ML-based [3443]. MMODPAD [3893]. MMODPAD-DRL [3893]. Mmsc [1592]. mmWave [3926]. MOACO [3983]. Mobile [3059, 3147, 3580, 3573, 2963, 2930, 3413, 3552, 3272, 3514, 2977, 2986, 3061, 3716, 3891, 3473, 3983, 3706, 3500, 2985, 3994, 3354, 2978, 3560, 3858, 4041, 3497, 3341, 3508, 2987, 3128, 3813, 3470, 3849, 3336, 2951, 2964, 3606, 3252, 2975, 3776, 3820, 4067, 3892, 2950, 3396, 2974, 2989, 3472, 2978]. Mobility [3651, 3573, 2843, 3517, 3295]. MobiWIS [3059]. modal [4070, 3793]. Mode [2074, 3283, 2878]. Model [4312, 3537, 1153, 4290, 2261, 1265, 4017, 2329, 2330, 516, 2249, 2402, 3325, 4217, 43, 696, 4135, 2059, 2069, 1024, 3160, 3781, 3013, 4004, 3665, 3608, 3741, 1088, 3004, 3915, 3785, 3755, 3754, 3015, 3711, 3767, 3712, 3014, 3817, 559, 3662, 3575, 2933, 3654, 2698, 4058, 3172, 3421, 3959, 3502, 3980, 3933, 2976, 2931, 3388, 2990, 2872, 4060, 3904, 4069, 2875, 2876, 3628, 4035, 3742, 3458, 3599, 3867, 1079, 3045, 3735, 2817, 3039, 3568, 3567, 3044, 3877, 3282, 474, 3593, 2802, 4034, 3690, 4020, 2901, 3850, 3562, 3871, 3638, 3827, 2801, 4064, 1733, 3376, 3483, 3231, 3679, 3985, 3919, 3493]. Model-based [3537, 3781, 3013, 3755, 3754, 3599, 3044, 2802]. model-driven [3785, 3690, 3493]. model-oriented [3654]. modeled [3743]. Modeling [4314, 283, 2601, 2873, 4259, 3929, 2749, 3180, 2769, 3066, 4015, 2610, 2706, 3684, 3296, 2874, 3346, 3623]. Modell [1153, 288, 474, 4135]. Modelle [690, 706, 151, 198, 587, 1024]. Modellen [1079]. Modelling [4204, 2100, 2444, 2441, 4215, 3491, 4381, 2776, 2871, 3347, 4375, 4380, 4377, 3218, 3345, 3099, 328]. Models [374, 1386, 706, 1265, 1285, 2581, 2342, 4271, 2136, 4023, 4276, 2461, 4138, 3987, 2898, 3405, 2883, 2765, 3088, 3751, 3648, 4015, 690, 3276, 3351, 4080, 2862, 3982, 3409, 3801, 2996, 3016, 3428, 3341, 2899, 2879, 2881, 3173, 3404, 3443, 3965, 3362, 2845, 3058, 2884, 4020, 3262, 587, 3829, 3063, 3434, 3955, 813]. Modern [375, 575, 3008, 3062]. Moderne [646, 10, 95]. moderner [248]. Modification [1743, 2176, 2247, 4223, 3890, 1214, 3068, 1613, 1470]. Modifications [1488, 887, 1904, 1601, 1361]. Modified [1934, 2705, 2731, 2023, 2093, 1368, 346, 3557, 3421, 2182, 3906, 608, 3601, 1503, 520, 3660, 1737, 1452]. Modifikation [887]. Modifikationen [1361]. modifizierte [1934, 608, 520]. modifizierten [346]. modifizierter [1452]. modifiziertes [1503]. Modifying [4327, 3930]. module [2753, 2720, 4035, 4239]. modules [4134]. Modulus [4339, 604]. MOHBA [3790]. molecular [2702, 2790]. Molecules [2554]. Moler [247]. Moment [1363, 753]. Moment-And [753]. Moments [2428, 2648]. Monegato [1662]. money [3804]. money-saving [3804]. Monge [2201]. MongoDB [2969]. MongoDB-based [2969]. monitor [4072]. monitor-analyze-plan-execute [4072]. Monitoring [3311, 2900, 3474, 3026, 3533, 3461, 3575, 3935, 3732, 3423, 3291]. Monograph [420, 539, 427]. Monographs [35, 36, 536, 401, 42, 40]. monoton [477, 1955]. Monotone [2173, 1694, 1955, 2575, 2778, 1417, 1866, 2080, 1573, 928, 1174, 4345, 1537, 1838, 1778, 1222, 1645, 1535, 1038, 1921, 1953, 619, 935, 1234]. monotoner [619]. monotonic [3101, 2143]. Monotonically [477]. Monotonicity [1960, 1600, 1303, 448, 1656, 1690, 595, 767]. Monotonie [767]. Monotonie- [767]. Monotonienebenbedingungen [1960]. Monotoniesätze [595]. monotonous [2017, 935]. Monroe [116]. Monte [1915, 1915, 1693, 2079, 2237]. Monte-Carlo [1915, 1915]. mood [3406].

Moore [532, 346, 4334, 2469, 2162, 2622, 1664, 827, 2779, 1703]. Morgenstern [148]. Morozov [1965, 1982]. Morozov’scher [1965]. Morphology [4272]. Morrison [1403, 1362, 1362]. Morrison-Brillhart [1362, 1362]. Morse [3620]. Mortar [2399, 2468, 2722]. Mortaring [2382, 2525, 2621]. Mosaic [2305, 3385]. Mosaic-Skeleton [2305]. Moser [2888]. moth [3421]. moth-flame [3421]. motifs [3835]. Motion [4265, 4267, 4269, 4009, 2603]. motion-planning [2603]. Motivated [4326]. motivations [3894]. motive [3161]. motive-structure [3161]. mounted [3820]. Moving [1834, 2463, 1218]. Moyle [360]. MPI [2940, 3750, 2959, 3204, 2938, 2961, 3576]. MPI-Performance-Aware-Reallocation [3204]. MPI/OpenMP [3576]. MPSoCs [3205]. MR [1172, 2176, 1845, 1136, 1813, 1714, 1129]. Mráz [1992]. mReno [2955]. MRI [3289]. Mufti [651]. Müller [817, 426, 97, 627, 658]. Müller-Merbach [627]. Multi [1203, 2215, 1997, 1837, 2440, 3900, 3526, 2287, 2542, 2543, 3918, 819, 1849, 2387, 3761, 3950, 3094, 2929, 3518, 3779, 1072, 3971, 2075, 3283, 3280, 1380, 3613, 3295, 4056, 3730, 3887, 3552, 1075, 3834, 2763, 3675, 3891, 3878, 3000, 3441, 3928, 3281, 3931, 3502, 3983, 4045, 4076, 3119, 3200, 4070, 3390, 3007, 3964, 3001, 3988, 2667, 3193, 4043, 4057, 3225, 4074, 3881, 3090, 2998, 3764, 4038, 3793, 4051, 3175, 3309, 3003, 1011, 3217, 3760, 3633, 3690, 3635, 4040, 3809, 4036, 2904, 3850, 3986, 3535, 3640, 3480, 3491, 3232, 3768, 3839, 3886, 3820, 3091, 3425, 3483, 3957, 3707, 3385, 3487, 3777]. multi [3420, 3449, 3453, 3790]. multi- [1075]. Multi-access [3295, 3690, 3487]. Multi-agent [3761, 2763, 3119, 3964, 3225, 4040, 3491]. multi-algorithm [3764]. multi-armed [4056]. multi-attribute [3535]. multi-authoring [3809]. multi-channel [3309, 3850]. multi-class [3891, 4045]. multi-classification [3931]. multi-cloud [3390]. multi-core [4076, 3007, 3001, 3003]. multi-criteria [3502, 3232]. Multi-cyber [2929]. Multi-Dimensional [2287, 2542, 2543]. Multi-example [3900]. multi-factor [3635]. multi-field [3091]. multi-FPGA [3453]. multi-graphs [3175]. Multi-Grid [2215, 1997, 2667]. Multi-input [3526]. Multi-IoTs [3449]. multi-issue [3881]. Multi-label [3918]. Multi-Level [2387, 2075, 1380, 4038, 1011]. multi-mobile [3552]. multi-modal [4070, 3793]. multi-mode [3283]. Multi-objective [3950, 3094, 3518, 3971, 3283, 3280, 3613, 3887, 3441, 3928, 3281, 3983, 4051, 3760, 3633, 4036, 2904, 3707, 3777, 3790]. multi-objectives [4074]. Multi-paradigm [3779]. multi-path [4057]. multi-perspective [3090]. Multi-Processor [1203, 1072]. Multi-Purpose [1837]. multi-radio [3309, 3217]. multi-resource [3193, 3483]. multi-robot [3480]. multi-scale [4043, 2998, 3986, 3091, 3420]. multi-sink [3200]. multi-stage [3820]. multi-stakeholder [3988]. Multi-step [1849]. multi-swarm [3425]. multi-temporal [3385]. multi-tenancy [3957]. Multi-Term [2440]. multi-threaded [3000]. multi-tier [3640]. multi-truth [3878]. multi-type [3764]. Multi-Value [819]. Multi-variate [2542]. multi-version [3675]. multi-workflow [3886]. Multicast [2999, 3217, 3112]. multicoloring [2623]. Multicomputers [2187, 2200, 2042]. multiconstraint [1657]. Multicore [3744, 2914, 3538, 3035, 3181]. Multicriterion [1784, 1228]. Multiderivative [755]. Multidimensional [1869, 2283, 3293, 1622, 3338, 2673, 1091]. MULTIFIT [2232]. multifunction [3629].

Multigrid [1955, 2094, 2348, 2399, 2099, 2262, 2427, 2303, 2301, 2306, 2554, 2252, 2331, 2318, 2116, 2036, 2304, 2571, 2650, 2074, 2111, 1838, 2477, 2300, 2113, 1525, 1080, 2674, 2655, 2626, 1966]. Multigrid-Method [1838]. Multilayer [1825]. Multilevel [2488, 2337, 2660, 2362, 2226, 2123, 2290, 2505, 2230, 2937, 2302, 2311, 1242, 2666, 2615, 1961, 1356]. Multimedia [3656, 3513, 3080, 3075, 3077]. multimodal [3547, 3982, 3699, 3045]. Multipatch [4219]. multipath [3758, 3107, 3656, 4042, 2955]. Multiple [1378, 2211, 4246, 2515, 631, 2815, 2247, 1273, 900, 1809, 2184, 925, 2293, 3992, 3594, 1004, 923, 2970, 3444, 1663, 3050, 1332, 3096, 2794, 2063, 3377, 3450, 4088, 3934]. Multiple/arbitrary [2815]. multiplex [3959, 3313]. Multiplication [1106, 2043, 997, 1285, 1511, 4117, 953, 3463, 1562, 566, 4240, 50]. Multiplicative [2215, 2075, 1969]. Multiplicators [475]. Multiplicity [1889, 871]. Multipliers [2495, 2369, 3576]. Multiplikation [566, 50, 953]. multiplikativen [1969]. Multiply [932, 2350]. Multiplying [1775]. multipoint [1611]. Multipole [2465, 2655]. Multiprocessor [1409, 1897, 4127, 4128, 1853, 1006, 3183, 3307, 3842]. Multiprogramming [749]. Multiprozessor [4128]. Multiprozessor-Rechenanlagen [4128]. Multiprozessorsystem [4127]. Multiquadric [4252]. Multirate [2155]. Multiresolution [2564, 4297, 2415, 2455, 2457]. multiresource [3756]. multiscale [2942]. Multisection [2322]. multisensor [3343, 3511]. multiset [3870]. Multispace [2666]. Multisplitting [1928, 1735]. Multistage [755]. Multistep [907, 779, 755, 309, 1638, 562, 1413, 2350, 1388, 1282, 988, 1199, 1085, 1115, 1077, 1275, 1201, 851, 976]. Multistep-Multistage-Multiderivative [755]. multithreading [3005]. multivalued [1275]. Multivariable [845]. Multivariate [2430, 1475, 2385, 3274, 1674, 2327, 4072, 1215, 587]. MUMIPLOP [4076]. Murnaghan [184]. Murphy [98]. Muschelischwili [145]. MUSCL [1789]. MUSCL-Type [1789]. music [3225, 532]. musical [3050]. Mutability [3995]. mutation [3627, 4074]. mutual [3853, 3678, 3287]. MVs [3807]. myocardial [3289]. myocarditis [3349].

N [27, 12, 349, 641, 42, 745, 199, 643]. N. [71, 145, 215, 174, 333, 172, 251, 152, 491]. Nacelle [4296]. nach [346, 239, 1045, 195, 431, 298, 1503, 7, 25, 133]. Nachrichtentechnische [124]. Nachrichtenübertragung [326]. Nachriehtentechnik [236]. Nachweis [1020, 912]. Naherungen [1958]. Näherungsbrüche [292]. Näherungskoeffizienten [782]. Näherungsmethoden [358]. Näherungsverfahren [707]. näherungsweise [705]. Näherungswerten [464]. Naimark [36]. naive [3143, 3825]. Nake [727]. named [3744, 3395]. Nanoinformatics [2832]. nanomedicine [2832]. Narrow [4351]. Nash [3459]. native [3911]. Natural [16, 2376, 71, 3369, 3374, 17, 2765, 1238, 1350, 3373]. Nature [3845, 3596, 3523, 4022, 2799, 2800]. Nature-inspired [3845, 4022, 2799, 2800]. natürlichen [679]. natürlicher [17]. Naturwissenschaften [156]. Navier [2506, 2795, 2732, 2589]. navigation [2930, 3946, 3963, 3981, 4032, 4049, 4073, 4092, 4095]. Naylor [228]. NB [4003]. NCLab [2885]. nd [1878, 2123, 2032, 2197]. nd-order [1878]. Near [4103, 4357, 981, 716, 1911]. Near-Fields [4103]. Near-optimal [981, 716, 1911]. Nearby [2165]. Nearest [3735]. Nearly [2010, 2037, 2300, 1558]. Nebenbedingung [585, 1969]. Nebenbedingungen [878, 3, 594, 612, 2011, 558]. Nebennäherungsbrüchen [1545]. Necessarily [682]. necessary [1455]. Needed [2479]. Nef [672, 4243, 4295]. Neg-Weighted [2206, 2739]. Negative [4336, 2236, 307, 741, 737, 638]. negativen [108]. negativer [741]. negotiable [2762]. negotiation [3881, 3761]. neighbor [3679]. Neighborhood [2066, 4292, 3312, 1847]. neighboring [3370, 3872]. neighbors [3462]. neighbour [3366, 3735]. neighbourhood [3823, 1669]. Nekrasov [2511]. Nelson [361]. Nernhauser [657]. Nested [4297, 2387, 2194, 2667]. nestings [1154]. Net [2378, 1394, 3634]. nets [4138, 2829, 3219]. Network [83, 2490, 3646, 2441, 4199, 3365, 3306, 3840, 3537, 3514, 3557, 2943, 3461, 2952, 3234, 3863, 3034, 2842, 3037, 3581, 3893, 3161, 3848, 3545, 2698, 2944, 3591, 3924, 3927, 3689, 3630, 3832, 2931, 3032, 3810, 3958, 4091, 3569, 2983, 3274, 3818, 4069, 3631, 2920, 3944, 3532, 3031, 3304, 3468, 3628, 960, 3525, 4038, 2888, 2948, 3505, 1228, 3875, 3484, 3470, 3997, 3136, 2945, 3125, 3845, 3217, 3404, 3278, 3838, 2761, 669, 3065, 3477, 3865, 3516, 4020, 3885, 3879, 3815, 593, 3436, 2801, 3480, 2749, 3509, 4009, 3771, 3376, 3378, 2942, 3231, 3679, 3941, 3797, 2779, 3487, 3796, 4089, 3637, 3919]. network-aware [3278]. network-based [3034, 2931]. network-data [2842]. network-on-chip [3037, 3810]. network-oriented [3815]. networked [3268]. networking [3495]. Networks [2043, 2120, 3649, 927, 807, 1113, 2060, 830, 3055, 3559, 3365, 3588, 3565, 2953, 1644, 1751, 3651, 3478, 3552, 3452, 3598, 3221, 3802, 3626, 3741, 3338, 2986, 3750, 3353, 3677, 2997, 3646, 3773, 3319, 3661, 3571, 4044, 2725, 3830, 4085, 1165, 3570, 4000, 3473, 3853, 2678, 3959, 3578, 3502, 3852, 3924, 964, 3872, 3129, 3409, 2999, 2688, 3823, 3701, 3032, 3275, 3300, 3890, 3585, 3218, 3401, 3356, 3242, 3098, 3601, 3904, 3354, 2988, 3179, 3349, 3996, 3082, 3895, 4003, 4057, 3642, 3458, 3189, 3497, 3862, 2844, 3503, 3847, 3427, 3313, 3831, 2841, 1473, 3586, 3875, 3309, 3673, 3921]. networks [3078, 3632, 2843, 3019, 3845, 3935, 3217, 3563, 3366, 3766, 3705, 2845, 3058, 3312, 3494, 3330, 3270, 3899, 3737, 3929, 3036, 3791, 600, 720, 3535, 3871, 3186, 3273, 3527, 2955, 3277, 3285, 3776, 3820, 1345, 3892, 3807, 3063, 3957, 3168, 3385, 2989, 3709, 3472, 3989, 3200, 3206]. NETYS [3315, 3052]. NETYS’14 [3105]. NETYS’17 [3320]. NETYS’2016 [3256]. Netzen [960]. Netzplantechnik [669, 676]. Netzwerk [807]. Netzwerk-Veränderungen [807]. Netzwerke [323, 720]. Netzwerken [964, 600]. Netzwerkflüsse [593]. Neue [222, 77, 953]. neuer [840, 930]. Neuere [582]. Neuhold [1136]. Neumann [148, 534, 4224, 385, 2522]. Neumannschen [385]. Neural [3634, 2060, 3840, 3741, 3750, 3234, 3848, 2862, 3591, 3832, 3958, 4091, 4069, 3349, 3468, 3997, 3125, 3935, 3217, 2761, 3737, 3865, 3791, 3285, 3376, 3378, 3231, 3385, 3797, 2779, 3796]. neuro [2771]. neuro-genetic [2771]. neurons [3238]. Neutral [2295, 2748]. Newell [732]. news [3982, 3617]. Newton [1655, 1861, 1260, 1503, 2707, 1655, 1861, 1161, 2269, 1097, 1212, 4159, 1467, 1930, 2085, 2788, 4223, 1869, 1828, 2381, 1068, 2649, 1403, 595, 1659, 1698, 2247, 379, 1576, 1470, 1605, 1904, 1818, 1260, 1503, 1286, 1844, 1446, 1577, 1752, 519, 1537, 1645, 4347, 527, 2041, 3393, 2149, 2302, 2527, 1096, 1757, 1215, 1445, 2697, 2589]. Newton-Algorithmen [379]. Newton-Iteration [519]. Newton-like [1698, 1818, 1645, 2149, 1445]. Newton-SOR-methods [1260]. Newton-SOR-Verfahren [1260]. Newton-type [2707, 3393]. Newton-Verfahren [1503, 595]. Newtonian [2869, 2880]. Newtonschen [519]. Newtonsches [527]. Next [1643]. Next-Fit [1643]. NFC [3060]. NFIS [3916]. NFV [3957]. NFV-enabled [3957]. Nichols [575]. Nicht [392, 108, 760, 914, 682, 500, 867]. nicht-Hermiteschen [914]. nicht-negativen [108]. Nicht-numerische [392]. nichtdeterministischen [344]. nichtglatten [1942]. nichtlinear [679]. Nichtlineare [577, 582, 810, 1955, 369, 1526]. nichtlinearen [521, 1349, 339, 787, 971, 765, 1260, 554, 1036]. nichtlinearer [440, 695, 504, 1968, 1917, 450, 571, 933, 994, 497, 866, 410, 606, 707, 414, 767, 182, 689, 1082]. Nichtlinearitäten [130, 161]. nichtmonotones [1918]. nichtnegativen [307, 1059]. nichtnegativer [586, 638, 859, 1236]. Nickel [167]. Nicolson [2474]. Nicot [99]. Niederreiter [2562]. niedrigerer [467]. Niehtlineare [558]. Niemeyer [829]. Nieveelt [426]. Nievergelt [658]. Nikolenko [641]. Nine [1856]. Nine- [1856]. Nitsche [2382, 2525, 2621]. Nitsche-mortaring [2621]. NN [3336]. No [539, 67, 427, 401, 1129, 1172, 2176, 1845, 1136, 1813, 1257, 1714]. NoC [3540]. NoCs [3038]. Node [4056, 3821, 2853, 3823, 3996, 2843, 3885, 3679, 1020, 2009]. Node.js [3062]. Nodes [837, 144, 631, 3588, 3600, 1433, 923, 3959, 2620, 3996, 4038, 3847, 1332, 3586, 3800, 1295, 2971]. Noise [2490, 2180, 196]. Noisy [2564, 2147, 3284, 2715]. Nomograph [865]. Non [2564, 1464, 1512, 4297, 313, 2262, 875, 2046, 4338, 2095, 2540, 1109, 4353, 3541, 157, 2438, 4340, 2387, 2131, 2391, 700, 2667, 2288, 2241, 2882, 1195, 4365, 911, 2193, 2580, 550, 1942, 307, 517, 2594, 3054, 2727, 2909, 704, 3913, 3748, 638, 2687, 2617, 2795, 3243, 2870, 2874, 3969, 3287, 457, 1200, 450, 3619, 1056, 2859, 3290, 873]. non-conforming [2594]. non-contrast [3290]. Non-convex [2882].

Non-cosymmetrical [4353]. Non-Cyclic [911]. Non-Decimal [1512]. Non-differentiable [4340]. non-equidistant [2617]. Non-Exact [1464]. Non-Hermitian [2046]. non-homogeneous [704]. non-IID [3913]. non-intrusive [3619]. Non-Iterative [313]. Non-Linear [1195, 2095, 550, 517, 2727, 2874, 1200, 450]. Non-Lipschitzian [4338]. Non-Matching [2262]. non-negative [307, 638]. Non-Nested [4297, 2387, 2667]. non-normal [2909]. Non-Numerical [157]. non-observable [3243]. Non-Overestimation [2193]. Non-Overlapping [2288, 2241]. non-parametric [3969, 3287]. Non-Periodic [1109]. non-permanent [3054]. Non-preemptive [2540, 2391]. Non-Reflecting [2438]. non-renameability [873]. Non-singly-connected [4365]. Non-Smooth [2131, 1942, 2564, 2580]. non-smoothing [1056]. Non-stationary [700, 2795, 2870]. Non-swarm [3541]. non-symmetric [2859]. non-uniform [3748]. non-uniqueness [457]. Nonconforming [2055, 2262, 2253, 2301, 2373, 2349, 2254, 2656, 2910, 3002]. Nonconservative [2400]. Nonconvex [1429, 1829, 1002, 1992]. Nondeterministic [344]. Nondifferentiable [1930, 1166]. Nondirected [1112]. nonfixed [2002]. NonHermitian [914]. nonholonomic [2603]. Nonisolated [1425, 1259]. noniterative [1585]. Nonlinear [1794, 2386, 4174, 109, 317, 277, 2083, 2509, 1484, 1546, 1826, 1955, 2561, 2575, 4337, 1477, 1478, 2570, 35, 1808, 1930, 1893, 388, 1051, 1819, 2376, 760, 4353, 1797, 748, 2303, 787, 4383, 1624, 2359, 1849, 1968, 1573, 1604, 1666, 1790, 2249, 1917, 1879, 2131, 2090, 765, 1381, 1567, 1605, 1787, 1818, 933, 994, 1373, 4116, 1906, 2510, 1425, 2189, 866, 1120, 1774, 644, 1926, 1523, 1626, 1342, 4164, 2126, 1036, 606, 1414, 2328, 707, 2160, 295, 1504, 266, 2149, 762, 1413, 767, 871, 1259, 753, 2070, 4382, 2671, 2642, 1245, 2778, 440, 521, 1081, 1564, 695, 1699]. nonlinear [1735, 2698, 369, 1349, 2114, 339, 504, 1382, 971, 1659, 1698, 1805, 1585, 1661, 2781, 2894, 1050, 1560, 1214, 2732, 1561, 1651, 571, 1343, 1476, 1260, 554, 4233, 497, 1266, 2822, 410, 1243, 1276, 1982, 1454, 1528, 414, 2661, 1664, 2895, 1221, 1096, 1757, 1741, 689, 3874, 679, 1082, 1483, 1718, 1800, 2789, 2748, 558, 1526]. Nonlinearities [130, 161]. Nonlocal [2292, 2542, 2543]. nonmagnetic [2877]. Nonmatching [2495]. Nonmonotone [2034, 1918]. Nonnegative [108, 859, 2336, 1059, 586, 1236]. Nonpreemptive [1853]. Nonsingularity [2491]. Nonsmooth [2161]. nonsparse [1096]. Nonstandard [2197]. Nonsymmetric [1502]. Nonuniform [2544, 2333, 1380, 2633]. nonvalued [710]. Norkin [75]. Norm [380, 2402, 457, 4326, 444, 1089]. Normal [836, 926, 468, 469, 2508, 712, 713, 1505, 1734, 4384, 2909, 1274]. normale [836]. normaler [468, 469]. Normalformen [712, 713]. Normalität [110]. Normality [110]. Normalization [997]. Normalized [4340, 2715]. Normally [738]. normals [2609]. Normed [72, 1218]. Norms [57, 983, 4005, 2230, 703]. nose [3213]. NoSQL [3422, 3357]. Notation [4121]. Note [2409, 1795, 2232, 2102, 717, 380, 2393, 2402, 4358, 2317, 3738, 2324, 562, 3058, 2539, 1807, 978, 2962, 1129, 1370, 1157, 1264, 1327, 2088, 1753, 1241, 1747, 518, 1446, 1845, 1352, 1326, 3071, 3195, 1100, 1372, 1575, 1664, 1469, 1584, 1714, 83, 455, 113, 386, 9, 59, 669, 84]. Notes [622, 426, 650, 648, 649, 627, 700, 581, 2443, 744, 628, 540, 530, 534, 582, 651, 672, 605, 245, 387, 381]. notifications [3339]. Notion [933]. Notions [3467, 54]. notwendig [682]. Novel [3529, 3531, 3821, 3992, 3618, 3230, 4039, 3657, 2982, 3086, 3773, 3854, 3727, 2969, 3810, 3287, 3669, 3841, 3350, 3738, 3307, 3836, 3926, 3673, 3632, 3766, 3456, 3635, 3414, 2901, 3332, 3850, 2840, 3551, 4071, 3292, 3289, 3860, 3457, 3233, 3722, 4011, 3919]. nozzles [2864]. NP [1865, 939, 2051]. NP-Completeness [939]. NP-Hard [2051, 1865]. Nr [63]. NS [4022]. NS-OWACC [4022]. NTG [84]. NTG-Diskussionsitzung [84]. Null [2011]. Nullraum [2011]. Nullraum-Algorithmus [2011]. Nullsstellen [166]. Nullstelle [162]. Nullstellen [692, 432, 366, 20, 379, 458, 823, 160]. Nullstellenbestimmung [890, 705, 616]. Number [2541, 1668, 494, 2181, 2281, 2351, 1365, 1889, 1411, 1022, 104, 1959, 911, 1843, 2106, 2479, 273, 1380, 822, 620, 3745, 2903, 4384, 1274, 773, 1852, 2862, 1005, 960, 637, 1672, 4137]. Numbers [475, 2039, 4098, 4099, 926, 738, 556, 2174, 2314, 1306, 821, 1638, 55, 863, 1619, 50, 862, 4384, 1991, 1981, 316, 1101, 1706, 566, 3323, 2662, 2658]. Numeration [2825]. Numeric [1624, 915, 916]. Numeric-Analytic [1624]. Numerical [1296, 4333, 2341, 2047, 1648, 109, 317, 277, 1464, 2083, 2497, 2555, 276, 2634, 1943, 1258, 1995, 2548, 2433, 4337, 875, 2179, 774, 1289, 285, 646, 2227, 2434, 1893, 1488, 1051, 2871, 2216, 1707, 1451, 2858, 1043, 157, 2010, 2864, 1118, 2866, 838, 139, 957, 2144, 2631, 1309, 2495, 2038, 2397, 769, 2310, 1790, 2868, 1126, 1314, 2869, 2487, 2185, 301, 13, 64, 4105, 241, 1208, 2747, 1850, 1709, 1711, 1501, 900, 854, 750, 2276, 1855, 2511, 2268, 2522, 2510, 253, 1629, 851, 1391, 2473, 1926, 2880, 2087, 1267, 981, 925, 2063, 2021, 1824, 96, 1195]. Numerical [1435, 4168, 1820, 1958, 2547, 522, 523, 2149, 93, 164, 135, 127, 280, 537, 1909, 2323, 4122, 1908, 2527, 1270, 2056, 862, 2104, 2343, 2004, 132, 2786, 1578, 1525, 4368, 4382, 517, 2752, 1149, 2664, 1756, 1081, 2851, 1010, 1662, 54, 969, 972, 982, 611, 1981, 988, 1210, 1086, 1760, 975, 986, 2784, 2808, 1074, 1108, 1271, 1585, 1749, 314, 2872, 2746, 385, 1687, 1359, 1983, 1351, 2682, 1213, 1750, 1266, 1091, 2879, 1067, 1011, 4119, 1194, 1328, 2882, 2661, 1426, 1764, 1107, 1697, 1026, 1055, 1226, 1076, 1369, 1713, 880, 904, 113, 1962, 559, 1046]. Numerical [695, 663, 965, 1944, 994, 1225, 18, 877, 1125, 1942]. Numerically [1887, 1002, 1407]. numerics [4376, 4231]. Numerik [663, 404, 758]. numérique [176, 405]. Numériques [258]. Numerische [509, 27, 113, 357, 324, 392, 16, 982, 611, 838, 1126, 1950, 1958, 164, 880, 1962, 559, 1046, 695, 2010, 965, 1754, 994, 1225, 18, 877, 1125, 1942]. Numerischen [66, 1943, 54, 972, 988, 986, 1043, 1944, 4105, 854, 851, 1011, 1107, 862, 1226, 1076]. numerischer [915, 916, 981]. numerisches [1995, 241, 385]. Numerov [2633, 1820]. Numerov-Type [1820]. NURBS [1916, 2710, 1916, 2719]. NURBS-Kurven [1916]. NUTS [3191]. Nystrom [888, 1358, 888, 1492, 1358, 1526, 1492, 1253, 1628, 1617, 1526]. Nyström-Formeln [888]. Nyström-Formelpaar [1492]. Nyström-Runge [1526, 1526].

O [426, 485, 148, 141, 699, 402, 658]. O. [792, 810]. OaaS [3186]. Ober [680, 634, 664]. Oberblick [9]. Oberflachenintegrale [2010]. Oberwolfach [245]. Object [2403, 2257, 4202, 3872, 3948, 3185, 2924, 3720, 4009, 4047, 3420]. Object-Oriented [2403, 4202, 3185, 2924]. objective [3887, 3199, 3441, 3928, 3281, 3983, 3950, 4051, 3094, 3518, 3760, 3633, 3971, 3873, 4036, 2904, 3283, 3280, 3613, 3707, 3777, 3790]. Objectives [1072, 3172, 4074]. Objects [2081, 1722, 4213, 4278, 2819]. OBL [4039]. oblivious [3317, 2737]. Obreshkov [2906]. observable [3243]. Obstacle [4333, 1524, 2521, 1435]. obstacles [3270, 1911, 1055]. Obtained [4201]. obtaining [297, 704, 2614, 1766]. OC [592, 592]. OC- [592, 592]. occupancy [3254]. occur [1992]. occurrence [2811]. Occurring [862, 3918, 2771]. Octtrees [4310]. Odd [1599]. ODE [1994, 1977, 1280]. ODEs [1840, 1993, 2032, 1474, 1607, 1908, 2700]. Oetker [702]. Oettli [1070]. Off [913, 627, 1398, 709, 3637]. off-step [1398]. offer [3761]. offerings [2980]. Office [551]. offline [3806]. offload [3186]. Offloading [3580, 3940, 3500, 3354, 3428, 4065, 3839, 4067, 4089]. offs [4059]. Offset [4241, 3659]. Offsets [2130]. Ohne [413]. Ökonometrie [511, 658]. ökonometrischer [690, 706]. Old [1491]. Oldroyd [2880]. Oliver [481]. Olver [389]. omission [3579]. on-chip [3034]. On-demand [3140, 3535]. On-Line [2250, 2307, 685, 2210, 2540, 1947, 2476, 2019, 2840]. on-the-fly [2898]. One [2062, 1122, 1825, 2372, 2591, 1771, 894, 779, 2292, 1834, 4339, 1776, 1666, 761, 2249, 2400, 2432, 1137, 2072, 1816, 2339, 1824, 2165, 1820, 742, 1205, 1168, 2350, 1278, 2170, 1138, 2853, 1684, 1657, 1377, 3177, 693, 987, 1470, 1351, 4199, 1031, 3865, 1125, 1452, 3827, 279]. One-Dimensional [1122, 2372, 1771, 2292, 1834, 1666, 761, 2339, 1820, 742, 2170, 2591, 2853, 1351, 3865]. One-Leg [2350]. One-Parameter [2400]. One-Phase [1824]. One-Sided [894, 1137, 1377, 693]. One-Step [1776, 2432, 1168, 1138, 1684, 1125, 1452, 3827]. one-symbol [1031]. One-Table [2062]. Online [2470, 2105, 3898, 3747, 2109, 2679, 3210, 3806, 3196, 1344, 3500, 3098, 4066, 3867, 3094, 2593, 3602, 3237, 3989]. online/offline [3806]. only [1893, 1279, 1800]. onto [3183, 1396]. ontologies [3329]. Ontology [3382, 3900, 3628]. Ontology-based [3382, 3628]. ontology-label [3900]. Open [4350, 1985, 2210, 3733, 4081, 3169, 3622, 2716, 3430]. Open-End [2210]. open-source [4081]. OpenCL [3555]. OpenMP [3576, 3745]. OpenStack [3306, 3423]. Operating [731, 808, 3143, 3265, 4138]. Operational [1391, 827, 1266]. Operationen [1170]. Operations [48, 1170, 70, 67, 68, 426, 858, 650, 151, 648, 649, 327, 627, 700, 581, 938, 4113, 1495, 104, 744, 628, 540, 530, 534, 582, 651, 205, 248, 672, 1283, 51, 1022, 1180, 1170, 104]. Operator [1826, 1995, 4338, 1185, 4360, 1414, 1234, 453, 396, 1541, 2143, 444, 1067, 398]. Operator-Equations [1185]. Operatoren [477, 836, 468, 409, 433, 469, 632, 908, 398]. Operatorgleichungen [1995, 396, 398]. Operators [477, 1251, 1167, 2467, 2502, 836, 1444, 1930, 468, 469, 2537, 2496, 2362, 2542, 2543, 908, 1489, 1580, 1647, 1495, 2510, 1966, 390, 2311, 409, 433, 632, 2643, 1404, 1576, 3358, 2809, 444, 2715, 2794, 968, 398, 1240, 1235]. OPHAencoder [3693]. opinion [3647]. Opler [490, 204]. OppNets [3202]. opportunistic [3884, 3356, 3454]. opportunity [3410]. opposite [1074]. opposition [3914, 3876, 4039, 3046]. opposition-based [3914, 3876, 4039, 3046]. Optical [2342, 552]. Optimal [1178, 1441, 454, 1786, 2018, 1128, 4298, 225, 169, 427, 1222, 683, 2544, 1551, 3317, 2590, 1599, 2449, 1189, 1939, 1173, 1252, 1303, 993, 378, 1875, 1567, 900, 3497, 1844, 1638, 4344, 3247, 651, 633, 1317, 2041, 1882, 865, 275, 2113, 2089, 2589, 1552, 1044, 1376, 1683, 4019, 2853, 2766, 3056, 3908, 716, 4076, 3939, 2674, 4059, 1260, 1079, 1629, 3845, 981, 1640, 1597, 4034, 1571, 1911, 1439, 2650, 2726, 1345, 778, 878, 764, 883, 593]. Optimale [648, 1683, 1128, 777, 1260, 275, 778, 764, 883]. optimalen [1376, 1939, 1079]. optimaler [271, 534, 981, 187]. Optimality [4258, 1782]. Optimieren [291]. Optimierung [325, 579, 1939, 357, 558, 613, 1919, 880, 232, 889, 1938, 549, 879, 794, 594, 612, 653, 1951, 697, 1918]. Optimierungsaufgabe [1969]. Optimierungsaufgaben [245, 1027]. Optimierungsproblemen [628, 1915]. Optimierungsrechnung [221]. Optimierungstechniken [1969]. Optimierungstheorie [770, 597]. Optimierungsverfahren [151]. optimisation [4058]. Optimization [2142, 4336, 1207, 1289, 2322, 2085, 245, 674, 2196, 1429, 770, 1784, 2033, 2131, 1938, 153, 2927, 3435, 3896, 4341, 3970, 2877, 879, 2252, 3584, 1919, 1124, 1120, 927, 2084, 1538, 487, 530, 2049, 1951, 2059, 1594, 2175, 1322, 1027, 1918, 1278, 3392, 2020, 3530, 3221, 2773, 3915, 3353, 3627, 3914, 3197, 3945, 3908, 3646, 2965, 3909, 3830, 1374, 3961, 3281, 4080, 3421, 3854, 1152, 3550, 1574, 3687, 4230, 3948, 3890, 3218, 3964, 2874, 613, 3841, 3193, 3525, 4041, 3311, 3303, 3847, 3205, 4131, 4005, 4051, 1729, 1228, 3038, 3813, 2822, 2836, 1845, 597, 3518, 3596, 697, 1575, 3582, 3971, 1758]. optimization [3553, 3873, 2798, 2904, 3962, 3692, 3524, 3604, 2650, 4008, 3232, 3798, 3820, 3531, 1218, 2799, 2800, 3280, 3236, 3425, 3486, 3613, 3957, 3777, 2736, 1714, 3953, 4292, 594, 612]. optimization-based [3945, 3961]. optimizations [3710]. optimize [3204, 3907, 3191]. Optimized [3466, 2057, 4302, 4003, 3374, 3107, 3445, 3078, 3071, 3582]. optimizer [4039, 4074, 3950, 3776, 3334]. Optimizing [4079, 3417, 3933, 1500, 3205, 4055, 3519, 4021, 3580, 3887, 3369, 4022, 3453, 3399, 3980]. Optimum [454, 476, 1583, 1791]. Optimum-Mix [454]. option [3062]. Options [997, 2568, 2473]. optoelectronic [2827]. oracle [3956]. oracle-based [3956]. Orbital [2031]. orbits [2221, 4237]. orchestration [3432, 3507, 3862]. orchestrations [2834]. Order [1794, 2426, 823, 1514, 317, 277, 2092, 771, 2267, 1910, 2099, 1796, 2289, 1839, 2337, 1831, 2565, 1827, 1468, 1506, 2423, 467, 2166, 1617, 1639, 4207, 2415, 838, 2338, 2144, 2123, 1277, 2295, 4102, 908, 2349, 854, 933, 1932, 1368, 1425, 1550, 1949, 1888, 2296, 2195, 1710, 1813, 2037, 2428, 2239, 2315, 2539, 877, 2465, 1964, 2237, 2197, 2111, 2367, 2222, 753, 2442, 1589, 1578, 297, 1655, 1727, 1161, 78, 2594, 2586, 2680, 1334, 1367, 1253, 1533, 3848, 364, 988, 1492, 2857, 975, 986, 1084, 1878, 1061, 1715, 2863, 1542, 1582, 1690, 1980, 2874, 2746, 2791, 2704]. order [3835, 2014, 292, 1077, 444, 1194, 1535, 1038, 1261, 1640, 4168, 1052, 2139, 2396, 3352, 2571, 2638, 1078, 976, 1125, 1358, 3562, 2171, 1563, 984, 1235, 1780]. order-convex [1235]. order/ [2860]. Ordered [4111, 2219, 2314, 742]. Ordering [2291, 847, 1357, 3019]. ordering-based [3019]. Orders [1599, 4192, 1874, 4193, 985, 779]. Ordinal [3860, 3383]. Ordinary [1389, 1141, 1768, 708, 779, 755, 2239, 449, 759, 877, 798, 4171, 297, 2785, 78, 6, 4157, 585, 1085, 1700, 1709, 470, 1225, 1716, 614, 2139, 1125, 1339, 1358, 1279, 1083, 1226]. Ordnung [823, 78, 364, 467, 988, 1492, 838, 908, 854, 1949, 1125, 984, 297, 975, 986, 292, 1932, 444, 877, 976, 1358, 1964]. ordnungs [4102]. ordnungs- [4102]. Ordnungsbegriff [933]. Oregonator [4179]. Organisation [642]. Organisational [2763]. organization [3672]. organizational [3520, 3391, 4081]. organizations [3824]. organizing [3016]. Orientation [3691]. Orientations [2333, 1544, 3661, 1074]. Oriented [2403, 4202, 2081, 803, 4368, 3866, 3654, 3783, 2911, 3364, 4041, 3920, 463, 3096, 3138, 3185, 2960, 3947, 2924, 3815]. origin [4237]. Origins [1863]. ornach [58]. Ornstein [2613]. Orten [461, 278]. Örtern [1062].

Orthoconvex [2279]. Orthogonal [4312, 2705, 3227, 3271, 2641, 1642, 1737, 407, 820]. Orthogonale [820]. orthogonalization [1719]. Orthogons [2279]. Orthonormalization [25, 308]. Orthonormalized [635]. Orthonormierung [25]. Osborne [487]. Oscillating [2528]. oscillations [971, 1560]. Oscillatory [841, 1750]. Oseas [3893]. Oseen [2358]. Osprey [3666]. Other [1400]. OTIS [2827]. out-of-core [3258]. out-of-kilter-algorithm [720]. Out-Of-Kilter-Algorithmus [720]. outages [2945]. outcome [3099]. outdoor [2936]. outlier [3559, 3694]. outlook [2759]. Output [151, 2325, 952, 1376, 3045]. Output-Sensitive [2325]. outsourced [3297]. Over- [452]. Over-Set [2329]. overcoming [3270]. Overestimation [990, 2193]. overflow [456]. overflow-free [456]. Overhauser [4248]. overhead [3240]. overheads [2960]. overlap [2959]. overlapped [3178]. Overlapping [2288, 2241, 3598, 3446, 3409, 2998, 3705]. Overlay [2934, 2842]. overlays [2836]. Overrelaxation [1477, 1478, 1604, 1196, 1661]. oversampling [3484]. Overview [9, 1899, 3026, 3663, 3563, 3140]. OWACC [4022]. Owen [147, 101]. ownership [4081]. oximetry [3251].

P [89, 246, 72, 509, 14, 426, 579, 12, 388, 360, 757, 357, 153, 576, 540, 582, 142, 266, 417, 2440]. P. [673, 74, 358, 139, 151, 327, 746, 66, 425, 744, 794, 390, 86, 818, 284, 538, 550]. P.-R. [538]. P2P [2934, 2836, 3075]. PaaS [3166, 3096, 2980]. package [2668]. Packaging [1401]. Packet [3927, 3136]. Packing [1643, 1947, 2334, 1971, 2560, 2109, 2210, 2679, 1557, 2019, 1344, 3708, 1884]. PACKTER [2948]. pädagogisscher [28]. Padé [292, 1848]. Padésche [292]. Paging [1153, 1023, 962, 1153]. paging-complexity [962]. Paging-Trommeln [1153]. Pair [2363, 2294, 1492, 2665]. Pairs [1839, 2906, 2704, 1989]. palindromes [2218]. pan [3215]. pan-correlation [3215]. pandemic [3629, 3733, 3736]. pandemics [3653]. Panel [2296]. Pape [234]. Paper [315, 317, 1805, 1136]. Papers [2958, 541, 3320, 3052, 3105, 3246]. Parabolic [2388, 317, 277, 1925, 4353, 619, 2338, 1887, 2377, 1589, 1138, 2778, 1822, 2630, 1074, 2703, 2750, 445, 2775, 2809]. parabolische [445]. parabolischer [619]. paradigm [3336, 3779, 3211]. Parallel [1441, 1203, 2352, 1840, 1569, 2179, 4314, 1336, 311, 1994, 2085, 2120, 2218, 2100, 2187, 2470, 4290, 2028, 2152, 1744, 1928, 2057, 2256, 1892, 2573, 858, 2114, 4191, 2338, 1679, 4369, 2127, 2503, 2441, 1306, 3740, 3739, 1914, 3072, 1331, 2721, 4190, 1823, 1232, 1691, 2157, 2027, 2200, 974, 274, 341, 1058, 1898, 2273, 1040, 524, 2169, 2580, 2074, 1297, 1416, 2479, 2163, 2764, 3144, 1552, 3305, 2653, 2654, 2903, 4125, 3220, 3207, 1988, 3034, 2753, 3751, 1735, 3549, 3196, 2938, 3029, 3463, 3344, 3243, 2970, 3435, 3944, 4033, 3303, 3566, 1681, 2966, 4133, 4055, 3352, 2937, 3694, 2632]. parallel [3747, 1601, 3576, 1993, 4127, 1964]. Parallel-Chord-Method [1058]. parallel-network [3944]. Parallel-Processor [311]. Parallel-Verarbeitung [1914]. Parallelarbeit [614]. parallele [1993, 1994, 4369, 1964]. parallelen [274]. Parallelenverfahren [1058]. Parallelism [3658, 4125, 3056, 1123, 2973]. parallelization [3008, 4131, 3685]. Parallelizations [2042]. Parallelstruktur [4127]. Parameter [1936, 4302, 700, 2400, 2682, 1863, 2195, 1370, 1564, 1695, 1665, 2746, 1243, 2823, 1058, 1755]. Parameter-Dependent [1936]. parameter-robust [2746]. Parameter-uniform [2682]. parameterabhangiger [1936].

Parameterdarstellung [959]. Parameterisation [2460]. Parameterization [2130, 2459]. Parameterized [4280, 2210, 3227, 3271]. Parametern [899]. Parameters [1478, 1776, 820, 1626, 1058, 4135, 2367, 899, 1260, 2848, 2798]. Parametric [2445, 4200, 4241, 2137, 4203, 4250, 2450, 2454, 4271, 4328, 2328, 2050, 1945, 2572, 2602, 1374, 3969, 3287, 2807, 959]. Parametrisation [4296]. parametrische [511]. parametrization [2605, 2702, 1990]. parametrized [2609]. parcels [4136]. Pareto [3445, 3283]. Pareto-based [3283]. parity [2657]. parking [3513]. Parodi [1143]. PARSEC [3064]. parser [2780]. Parsing [726]. Part [2515, 1944, 2022, 915, 916, 481, 4276, 2170, 1392, 2697, 2653, 2654, 185, 409, 433, 1451, 1878, 1877, 2245, 2287, 2542, 2543, 294, 306, 2741, 2742]. Partagée [320]. Partial [2426, 2092, 4201, 1051, 2158, 129, 207, 748, 2481, 2188, 2336, 2249, 1032, 4193, 1120, 39, 4106, 1480, 747, 1466, 2442, 1589, 1108, 1762, 1709, 3053, 1553, 984]. Partially [837, 3132, 1279, 3279]. Participant [84]. participants [3054]. Particle [1767, 2090, 2927, 3777, 2982, 3847, 3205, 3531, 3236, 3425, 3486]. Particle-Method [1767]. Particular [368]. partielle [129, 207]. Partiellen [4106, 747]. partieller [984]. Partition [1583, 931, 1774, 3844, 4034, 2665, 3715]. partition-based [2665]. Partitioned [919, 2172, 2293, 3845]. Partitionierung [1972]. Partitioning [2138, 2202, 1401, 803, 4079, 3453, 1547, 1020, 3354, 1972]. Partitions [2544, 932, 1345]. Parton [230]. Parts [4328]. ParVec [3064]. Parvicursor [3182]. Paschkis [493]. passende [598]. passing [3257, 3885]. Passive [3700]. PASTA [2846]. Pastel [266]. Pasting [4324]. Patch [2445]. Patches [4205, 4251, 4207, 4308, 4220, 4330]. Patchwork [1952, 1952].

Patchwork-Verwerfung [1952]. Path [1047, 2487, 1331, 1872, 2623, 4195, 4282, 2169, 1918, 3305, 3925, 4046, 598, 3914, 1117, 3646, 3479, 1620, 3837, 4057, 1681, 1642, 1487, 3710, 2751, 1911]. Paths [1810, 2201, 2281, 848, 807, 293, 960, 1487, 3679]. patients [3349]. Pattern [122, 2830, 2944, 4379, 4134, 3978]. Pattern-driven [2830]. pattern-recognition [4134]. Patterns [4280, 3573, 3400, 3324, 4093, 3585, 3274, 3082, 3215, 4063, 3393, 3747, 2954]. Paul [728]. Paulin [255]. payment [2976]. payments [3060]. PC [1078]. PC-codes [1078]. PCA [3350, 3738]. PCS [2845, 3058]. PCSA [3639]. PDE [2773, 4244, 4246, 2680, 2698, 2635, 4328, 4168, 2460]. PDE-constrained [2773]. PDE-controlled [2698]. PDEs [2226, 2377]. Pebbling [1393]. Pedell [256]. peer [2838, 2763, 2841, 2837, 2840]. peer-to-peer [2838, 2763, 2841, 2837, 2840]. pegs [1732]. penalties [2673]. Penalty [1524, 1435, 1218, 986]. penalty-functions [1218]. Penalty-method [986]. Penalty-Methode [986]. Pencils [2203]. Penrose [2779]. Pentadiagonal [2366]. people [1731]. PEPPHER [3009]. Perfect [4191, 1652]. perform [2933]. Performability [2769]. Performance [3447, 3956, 3664, 2891, 749, 3008, 1592, 1570, 337, 3006, 3204, 3192, 2393, 2704, 1513, 3083, 3564, 3623, 3273, 2277, 3430, 3240, 3537, 4079, 1088, 3004, 1323, 2883, 3062, 3174, 3009, 3086, 3748, 3347, 4090, 4037, 1382, 2968, 2976, 3555, 4129, 3964, 3645, 4059, 4035, 3742, 3584, 3341, 3248, 3770, 4055, 3254, 3003, 3877, 3100, 3510, 3222, 3624, 3524, 3285, 2786, 3952, 3515, 1409]. performance-aware [3009]. performance-cost [4059]. Periodic [4179, 2518, 4178, 2294, 936, 1596, 1109, 1121, 1624, 11, 707, 1578, 1149, 1010, 1760, 1989, 1764, 1682, 2932, 446, 3807].

Periodical [1958]. Periodically [2400]. periodicity [1253]. periodische [446]. periodischen [1958]. periodischer [707, 182]. periods [2218]. periphery [3815]. permanence [3831]. permanent [3054, 2877]. PERMDEC [3831]. permissioned [3956]. Permutation [1169]. Permutations [869]. Perron [456, 456]. Perrot [731]. persistent [4130]. Person [3883]. personal [3422, 3314, 3824]. Personalised [2964, 3902]. personality [3828]. personalization [3057]. Personalized [3343, 3272, 3212, 3498]. perspective [3961, 3163, 3090, 272, 3124, 3126, 2792, 3509]. perspective-relief [272]. Perspektiven [338]. perspektivischen [272]. Perturbation [2405, 2329, 2330, 1850, 1966, 2293, 2350, 4163, 1697]. Perturbations [1290, 1945, 1, 1981, 3905]. Perturbed [2055, 2525, 2338, 2402, 2478, 2400, 2339, 1414, 1536, 453, 947, 2746, 2595, 2775, 1226]. pervasive [3534, 3960, 3917]. Peters [303]. Peterson [391, 421, 363]. Petri [1394]. Pfad [1918]. Pfade [598]. Pfanzagi [185, 185]. pharmaceutical [4068]. Phase [1824, 1534, 1820, 3907, 2014, 1526]. Phase-Lag [1820]. Phasenuntersuchung [1526]. Phasic [3944]. Philip [120]. Philosophy [16, 4304]. PHoM [2475]. PHoMpara [2573]. photonic [2908]. photovoltaic [3880]. Physical [623, 251, 3654, 3742, 3775, 2929, 3116, 2928]. Physically [4305]. Physics [4383, 42, 39, 2796, 34]. Physik [260, 300, 701]. Physikalisch [222]. Physikalisch-Geodätische [222]. Physiker [154]. physiological [3731]. pictures [1095, 3128, 1248]. PID [3916]. piece [3817]. piecemeal [637]. pieces [1016]. Piecewise [1207, 1289, 2117, 1618, 4317, 1866, 1314, 2485, 1824, 4275, 4263, 2606, 946]. Piehler [232]. Pierce [532]. piles [3475]. pilot [3314]. ping [3055, 2960]. ping-pong [3055, 2960]. Pintér [1714]. pipe [2889]. pipeline [3841]. pipelines [3993]. Pitaevskii [3808]. Pitchfork [2184]. Pivot [1595, 847, 2806]. Pivoting [2158, 776]. Pivots [683]. Pivotsuche [776]. Pixel [4292]. PL [279, 270, 730]. PL/1 [279, 730]. PL/I [270]. PLA [4188]. Place [220]. Placement [1531, 4329, 3164, 3786, 3748, 3441, 3928, 3983, 4076, 3991, 3518, 3644, 3769, 1322, 3869, 3957, 4061]. placements [3787]. Plan [1791, 4072]. Planar [2446, 2153, 712, 713, 1062, 4189, 1288, 4275, 3269, 2606, 1736]. planaren [1062]. planarer [712, 713]. Planaritätskriterium [294, 306]. Planarity [294, 306]. Plane [268, 904, 4265, 1062, 1465, 4180, 1396, 4329, 3306, 1313, 1494, 1652, 1217, 1326]. Planes [2463]. Planning [542, 83, 891, 3552, 3925, 3407, 3914, 3281, 3599, 3584, 3690, 3176, 2603, 4088]. Plant [427, 1938]. plantation [3226]. planted [1217, 1326]. Plates [2227]. platform [3387, 3746, 3433, 3455, 2981]. platforms [3160, 3785, 3453, 3007, 4055, 2967]. player [3106]. Plette [305]. PLL [2014]. plotter [980]. plugin [3172]. Plunungs [649]. Plunungs- [649]. plus [2938]. PM2.5 [3234]. PMINA [3881]. PML [2208]. PNS [2345]. POF [3445]. POF-SVLM [3445]. POI [4013]. Point [2564, 1856, 1496, 4098, 4099, 4183, 1808, 2258, 556, 784, 1588, 1788, 997, 1285, 1511, 565, 2086, 1970, 1185, 4323, 2124, 764, 883, 885, 938, 1306, 2520, 1914, 4257, 1841, 2588, 1619, 2165, 1908, 868, 1929, 2320, 1497, 451, 1299, 1742, 4184, 1398, 3998, 1663, 2745, 1162, 2793, 4170, 1571, 3986, 2681]. Point-Location [2124]. Point-Wise [565]. Points [2235, 2375, 2526, 2048, 845, 278, 1850, 1872, 1906, 1523, 1626, 2184, 2087, 2583, 2237, 1622, 1907, 1621, 1111, 461, 1453, 1753, 1214, 3497, 1343, 1422, 1243, 3136, 1697, 1696]. Pointwise [2340, 2332]. Poisson [1129, 1126, 788, 1216, 1484, 2851, 1249, 1851, 2095, 966, 1126, 2306, 1204, 1387, 974, 2128]. Poisson-Gleichung [1126]. Polar [4265]. policies [3670, 3321]. Policy [3321, 3093, 3804, 3218, 3179, 3944, 3965, 694, 3155, 3640, 3396]. policy-based [3179]. Politikiteration [694]. Pollard [1361]. Pollard-Algorithmus [1361]. polling [2959]. pollution [3468]. Polroot [167]. Polya [186, 286, 4264]. Polychotomy [1807]. Polycrystalline [1594]. Polygon [2454, 1396, 1781, 2737]. Polygonal [1810, 1722, 4207, 1519, 4365]. polygonaler [637]. Polygons [2452, 1544, 1731, 1439]. Polyhedra [4312, 4243, 4295, 4308, 1494]. Polyhedral [2475, 2487, 2668, 1911, 2573]. Polyhedron [906]. Polymorphic [4280]. Polynome [77, 1998, 2012, 1940, 981]. Polynomen [366, 760, 432]. Polynomial [1360, 1251, 2532, 4350, 2561, 2548, 1812, 1618, 618, 466, 366, 4319, 2475, 2573, 2383, 874, 1252, 2077, 4147, 1368, 166, 2410, 2485, 1268, 1550, 1809, 1949, 2244, 4346, 4362, 2514, 2513, 1029, 666, 857, 1009, 1069, 2190, 1510, 1198, 943, 4222, 2619, 2659, 2853, 2700, 1155, 946, 2668, 1754, 944, 2714, 2822, 1566, 1127, 1317, 637, 2596, 1736, 968]. Polynomially [2066]. Polynomials [4332, 1177, 2592, 917, 1028, 77, 1814, 1998, 4145, 1355, 2012, 760, 1680, 1940, 2412, 4356, 2515, 4146, 4181, 2556, 1674, 1638, 464, 4216, 853, 633, 212, 4264, 1493, 432, 2628, 1366, 1005, 1438, 1900, 1213, 991, 1100, 1292, 1670, 1766, 1671, 407, 981]. Polynomoperatoren [968]. Polynomräume [1754]. Polynoms [466, 166]. Polynomwur [1949]. Polynomwur-zelbestimmung [1949]. Polynomwurzeln [464]. Polytope [1946]. Polytopes [1946]. pong [3055, 2960]. Pontrjagin [271, 187, 694, 694]. POOSC [3806]. Popularity [3546, 3477, 2800]. Population [2581, 815, 3751]. poro [506]. poro- [506]. Poroelasticity [2477]. Porous [2292]. portability [3004, 3785, 3096, 3003].

Portable [4124, 3251, 3005, 4239]. Portrait [2046]. Pos [2206, 2739]. Pos/Neg [2206, 2739]. Pos/Neg-Weighted [2206, 2739]. Posed [2084, 972, 2738, 2771, 1982]. posets [1262]. Position [2450, 3289]. positioning [3898]. Positions [1062]. positiv [7, 867]. positiv-definiter [7]. Positive [4332, 2592, 2278, 1444, 4356, 1291, 948, 7, 935, 1614, 1059, 867, 1037, 1683, 409, 433, 2806, 1038, 398, 1672, 1240, 837, 1921, 1953]. Positiven [1037, 1059]. Positiver [1038, 409, 433, 935]. positivity [1101, 1806, 1335]. Possibilistic [2257]. Possible [777, 2106]. Post [2691, 3686]. post-processing [3686]. Post-quantum [2691]. Posteriori [2253, 2164, 2205, 1867, 2241, 2416, 860, 2096, 2141, 2254, 2669, 2865, 2908, 1665, 2809, 4158, 1571, 2742]. Posteriori-Fehlerabschätzungen [860]. Potential [1825, 2381, 572, 2973, 1525, 1418, 2701, 2209, 1107, 3214]. potential-based [2701]. Potentials [2486]. Potenzreihen [572]. Pouzet [2748]. powdery [3250]. Powell [783, 2035, 783]. Power [1336, 1521, 1939, 1938, 3133, 3537, 2914, 3335, 3319, 2933, 3110, 4085, 3172, 3591, 1090, 2869, 1687, 3719, 3846, 3742, 3247, 2846, 714, 3769, 3429, 4077, 3423]. power-aware [2846]. power-law [2869]. powered [3132]. Powerful [791]. powers [725]. PP [3943]. PP-PRNU [3943]. Prabhu [333]. Practical [1903, 3702, 1617, 2531, 625, 1089, 2990, 2955, 238]. Practice [1928, 3170, 3371, 3261]. practices [3061]. Prädikate [876]. Prädikatenlogik [165]. Prager [64, 135, 1070]. Prager-Oettli-Schranke [1070]. Praktikum [809]. Praktische [238, 675, 33, 154, 658, 417]. PraktischeMathematik [800]. praktischen [1089, 512, 336, 270]. PRAM [2772, 2600]. Prandtl [1565]. präsentierter [1114]. Praxis [657, 425]. pre [3686, 4080, 3980, 3933, 4007, 1900, 3047]. pre-copy [3980, 3933]. pre-distribution [3047]. pre-processing [3686]. pre-trained [4080]. pre-warming [4007]. pre-wavelets [1900]. Preassigned [837]. precedence [3307, 2846]. precedence-constrained [2846]. Precise [4174, 4181, 2788, 2662, 2658]. precisely [3417]. Precision [2267, 4115, 4117, 1371, 4070, 2815]. precomputed [3053]. Preconditioned [2123, 2064, 1823, 2274, 1340, 1423]. Preconditioner [2504, 2567, 2655, 1987, 2738, 2681]. Preconditioners [2501, 2264, 2030, 2356, 1772, 2773, 2618, 2793, 2005]. Preconditioning [4335, 2234, 2156, 2246, 2159, 2113, 1572]. Predicate [165]. Predicates [876]. predict [3968]. Predicting [3745, 3134, 3742, 3440, 3443, 3546, 4020, 3724]. Prediction [3234, 3504, 3631, 127, 3562, 3688, 3966, 3840, 3355, 1376, 1683, 3834, 3514, 3557, 3712, 3012, 3575, 3403, 3782, 3178, 3980, 3933, 3894, 3995, 3390, 3904, 4024, 4069, 3903, 3597, 3293, 3837, 3439, 3468, 3350, 3738, 4035, 3511, 4063, 3902, 3241, 3935, 3146, 3173, 3404, 3826, 3477, 3899, 3332, 3476, 3692, 3703, 3524, 3235, 3285, 3462, 3722, 3985, 4029]. prediction-based [3575, 3511, 3826]. predictions [4025, 3489]. predictive [4007, 2780, 3935, 3979, 3268]. Predictor [2906, 2259, 310]. Predictor-Corrector [310, 2906]. predicts [3161]. predistribution [2688]. preemptions [3194]. Preemptive [2357, 2540, 2198, 1935, 2391, 1204]. Preemptives [1935]. Preface [2238, 2299, 2498, 3361, 2517, 2847, 2905]. Preference [3881, 3613]. prepaging [1130]. preplanned [1088]. Preplanning [1087]. preprocessing [3595]. Prescence [196]. Prescribed [1062, 2298]. prescription [2762]. Presence [1498, 1911]. presented [1114]. Preserving [2260, 2117, 4319, 1559, 448, 4260, 2457, 1799, 2849, 3054, 2977, 3108, 3471, 3943, 3340, 2841, 3635, 3318, 3102, 3799]. preset [590]. prespecified [1482]. Pressure [2732, 2856]. preventing [3724]. Prewavelets [2407]. price [3120, 3322]. prices [4048]. Pricing [3120, 3155, 3839]. Prigogine [559]. Prime [2335, 55]. Primfaktorzerlegung [55]. Primitive [909]. Principal [1963, 918, 1771, 2699, 2531, 2093, 1296, 1150, 1662, 1271, 1351, 3080, 4158, 1392]. principal-value [1392]. Principle [1826, 1277, 767, 1161, 1013, 1629, 1982, 694]. Principles [542, 1868, 1965, 86, 2753, 679]. Printing [1883, 1031]. Prinzip [1013, 1503]. Prinzipien [1965, 679]. Prior [3728]. Prior-mask-guided [3728]. Priori [709, 1338, 1162, 2741]. Prioritäten [1356]. priorities [1041, 1356]. prioritization [3706]. prioritizing [3920]. Priority [3940, 565, 2826, 2672, 480, 3853, 3035, 2751, 2932]. Priority-based [3940]. Prism [4313]. Privacy [2977, 3108, 3402, 3054, 3912, 4016, 3670, 3471, 3943, 3659, 3340, 2841, 4031, 3965, 3635, 3318, 3102, 3486, 3799, 3295, 3487, 4011, 3955]. privacy-aware [3402]. privacy-enhanced [3912]. Privacy-preserving [3108, 3943, 3340, 2841, 3635, 3318]. Private [2757, 3913, 3066, 4023, 3222]. PRNU [3943]. PRNU-based [3943]. Proactive [4078, 3095, 3337, 3818, 4064]. Probabilistic [1497, 2257, 4197, 1319, 1250, 2839, 3711, 4085, 3150, 1980, 3388, 2845, 3058]. probabilities [4227, 715, 1746, 3415]. Probability [1116, 2069, 2690, 1071]. Probe [2523]. Problem [1178, 2388, 2055, 2173, 907, 2047, 107, 2433, 1985, 904, 4337, 2008, 2066, 2206, 1281, 1390, 1531, 2251, 1551, 4351, 411, 1320, 855, 1923, 211, 1643, 1639, 1045, 1043, 1835, 2549, 1828, 2334, 2153, 1802, 461, 2329, 2506, 2382, 2330, 1309, 2359, 2519, 1790, 2344, 2249, 2282, 2283, 278, 1401, 2402, 1465, 399, 2466, 2288, 1273, 1331, 1872, 610, 1829, 2053, 2358, 939, 751, 2186, 2407, 2009, 4364, 2524, 1267, 2416, 1966, 553, 1824, 22, 2214, 1435, 2272, 2169, 865, 252, 767, 1205, 209, 1536, 930, 661, 1942, 3565, 2003, 3253, 3870, 1717, 1418, 1688, 1459, 3159]. problem [4224, 1657, 1632, 1570, 2657, 3479, 2649, 4163, 2631, 4007, 1661, 1085, 1428, 2701, 3939, 4230, 3177, 1609, 1375, 382, 385, 662, 1092, 1740, 3, 2844, 3050, 4130, 1162, 1228, 1884, 1067, 1642, 1487, 2771, 500, 3971, 1426, 973, 1048, 1107, 2751, 1755, 3283, 2741, 2742, 1156, 1686, 679, 1630, 3882, 2739, 2736, 2878, 907, 662, 2009, 1966]. Probleme [338, 33, 203, 254, 440, 695, 495]. Problemklassen [751]. Problems [844, 4333, 2154, 454, 2417, 2083, 2096, 2509, 2233, 1546, 4350, 2497, 1795, 599, 2146, 2235, 1934, 1141, 1997, 1915, 2504, 1300, 2267, 1925, 2099, 1397, 2278, 2031, 2337, 1808, 4338, 245, 2591, 505, 2294, 1047, 1450, 1610, 754, 2470, 2590, 2376, 1254, 1400, 1588, 467, 2118, 760, 2011, 2148, 2297, 2496, 1429, 1788, 476, 1961, 1524, 2287, 2103, 2525, 2338, 4383, 2065, 2123, 2129, 2345, 2420, 1604, 4340, 2383, 4172, 2310, 2319, 2387, 2370, 2033, 1314, 901, 2340, 1999, 2554, 912, 42, 2468, 4359, 2369, 1315, 2029, 2291, 2418, 1832, 1850, 2030, 2159, 2560]. Problems [1787, 1501, 1290, 2349, 900, 2223, 2312, 1174, 849, 994, 4188, 2276, 2268, 2510, 2157, 866, 4189, 2432, 1120, 1192, 2318, 1391, 4180, 1649, 2084, 835, 530, 1841, 1926, 1863, 2121, 2052, 1239, 2339, 4195, 2437, 4164, 2553, 2579, 2050, 4170, 2428, 2304, 295, 1958, 762, 1151, 91, 877, 2151, 2580, 2356, 2197, 2111, 2367, 1027, 4171, 783, 1544, 753, 2389, 2293, 2350, 1502, 2320, 1578, 2907, 297, 2785, 4382, 2671, 2723, 3269, 517, 1161, 2773, 2679, 1242, 1175, 2642, 4175, 2762, 1436, 1245, 2018, 440, 1615, 2696, 1324, 1667, 1675, 1089, 1695, 1117, 1556, 1227, 972]. problems [1748, 2624, 1707, 1822, 695, 988, 1364, 2733, 1735, 995, 1199, 1760, 2660, 585, 2861, 2860, 2865, 1224, 2630, 1689, 369, 2676, 2629, 2621, 2114, 339, 1979, 1013, 964, 1743, 2176, 2645, 2820, 1620, 965, 2663, 2868, 947, 1742, 3772, 2818, 2746, 2791, 2674, 1660, 2735, 1709, 384, 1651, 1476, 2910, 3002, 445, 1725, 2682, 2775, 1302, 4233, 2809, 2926, 497, 2695, 1246, 1244, 1989, 2822, 2793, 410, 1011, 1572, 2626, 2738, 2787, 2646, 3596, 3639, 381, 18, 2627, 1982, 474, 1726, 312, 2859, 2887, 1977, 3873, 1125, 1339, 1275, 1755, 593, 1692, 2650, 2726, 495, 1741, 1697, 689]. problems [2598, 1226, 2171, 1483, 2681, 1984, 3777, 1713, 1193, 1611, 454, 149, 1632, 382, 487, 651]. Procedure [2388, 2235, 771, 760, 525, 382, 1154, 1752]. Procedures [346, 1520, 1448, 852, 2246, 864, 741, 1199, 460, 441, 1200, 405, 538]. Proceedings [4378, 830, 237]. Process [377, 4250, 476, 2925, 1032, 2060, 1564, 3464, 3088, 937, 3520, 4060, 3391, 3111, 3090, 2830, 3124, 3241, 3346, 3991, 3510, 4034, 967, 1766, 3434, 3955, 978, 127]. process-based [2830, 3991]. Processes [71, 2144, 631, 378, 4173, 116, 840, 135, 640, 521, 3326, 3087, 2933, 3204, 3089, 3684, 3520, 4369, 3391, 243, 3109, 4133, 2831, 3270, 483, 261, 333, 155]. Processing [354, 1203, 4126, 551, 639, 4265, 1730, 2256, 401, 2380, 1853, 3740, 3739, 2443, 4292, 402, 4277, 3369, 3374, 2828, 2903, 2893, 3015, 3686, 2640, 4017, 3749, 3187, 588, 3150, 3370, 3872, 3337, 3811, 3846, 3072, 3346, 3195, 2951, 43, 3245, 3800, 3373, 3418, 304, 575, 552, 801, 541, 353]. Processor [1203, 311, 1042, 1006, 2136, 1072, 1356, 4129, 2001, 1147]. processor-array [4129]. processor-sharing [1356]. Processors [4126, 1134, 2479, 2914, 1988]. produced [2877]. Produces [777]. Producing [4170]. Product [4174, 2348, 1437, 2496, 2429, 2464, 4306, 2024, 4210, 928, 1498, 1950, 4215, 4216, 853, 4217, 3431, 1712, 2542, 2543, 2768, 3967, 1738, 2714, 3051, 3254]. production [3281]. Products [1786, 2337, 556, 913, 1683, 2774, 3833, 4099].

Produkt [556, 853, 1950].

Produkt-Quadraturformeln [1950]. Produkte [4099]. Produkten [1683]. Produktion [252]. Produktionsplanung [92]. Prof [605]. Prof. [106, 605]. Professor [1793, 1554, 1223, 1508]. proficiency [2924]. PROFIL [2044]. PROFIL/BIAS [2044]. Profile [3455, 3938]. Profiles [2570]. profiling [3893, 3190]. profit [3448, 4018]. profit-aware [4018]. Program [826, 2045, 311, 1111, 386, 2385, 1875, 1969, 3239, 4227, 1015, 4133, 591, 1136, 1636, 1407]. Programm [386]. Programmablaufpläne [629]. Programme [347, 1016, 568, 81, 82]. Programmed [1336, 1788, 569]. Programmen [357, 81, 82]. programmes [3764]. Programmier [492]. Programmier-Handbuch [492]. Programmieranleitung [817]. Programmieren [515, 288, 829, 510]. Programmierlogik [629]. Programmiersprache [602, 730]. Programmiersprachen [728]. Programmiersysteme [349]. Programmiertechniken [490]. Programmierter [41]. Programmierung [603, 319, 426, 305, 235, 1940, 97, 582, 511, 577, 855, 194, 657]. Programming [420, 419, 3240, 533, 107, 1915, 1187, 889, 2590, 99, 855, 140, 1940, 2148, 1939, 157, 752, 120, 4172, 4306, 2242, 64, 279, 627, 700, 4341, 1829, 1192, 2072, 2559, 1816, 2284, 1239, 204, 213, 351, 1205, 1969, 1245, 3037, 3751, 2867, 1382, 2938, 3021, 1375, 558, 3072, 3566, 3022, 2814, 2822, 2638, 1992, 629, 3219, 391, 792, 388, 514, 430, 193, 281, 13, 253, 644, 177, 373, 352]. Programming-ALGOL [514]. Programs [2403, 111, 2280, 2588, 1782, 81, 82, 2175, 1745, 568, 3499, 4131, 1213, 2798, 1016]. Progress [48, 205, 3648]. Project [83, 1165, 29, 1473, 3283, 3198]. Projected [4336, 2617, 1230, 1752]. Projection [1604, 1567, 1906, 2084, 753, 451, 2733, 2643, 1661, 2732, 2726]. projections [1095, 1248]. projects [4081, 3546]. Projektion [451]. Promoted [3639]. prone [3257]. Proof [1625, 525, 2321, 3260, 694, 3436, 3602, 912, 165]. Proofs [4337, 1788, 1841, 1926, 2607]. propagating [3588]. propagation [3588, 4004, 2909, 3600, 3446, 3502, 3098, 3889, 3586, 3168]. propagators [2619]. Properties [4242, 4315, 494, 1580, 4270, 2456, 4214, 3447, 1366, 4223, 407]. Property [2260, 1819, 2268, 1102, 451, 1295]. Proportional [2295, 3132]. proportionality [1748]. Proposal [475]. Proposing [3253]. propositions [342]. Proppe [745]. prospects [4006, 3732, 246]. protection [4016, 3659, 3851, 3295, 3487, 4011]. proteins [3292]. protocol [3365, 3641, 3298, 3054, 3107, 2986, 2952, 3657, 3578, 2976, 3884, 2936, 3812, 3704, 3019, 3023, 3635, 2840, 2984, 2989, 3515]. protocols [3664, 3661, 2999, 4040]. prototype [2777, 3638]. prototyping [3453]. Protter [4364]. provable [3805]. Provably [3806, 3635]. provenance [3255, 3619, 3216]. provider [3518]. providers [3668]. Providing [1573]. Proving [771]. provision [3656, 2956]. provisioning [3272, 3070, 3000, 4066, 3599, 3505, 3096, 3142, 3140]. Prozeduren [346, 741]. Prozedurschachtelungen [1154]. Prozesse [4369, 284, 271, 187]. Prozessen [356]. Prozessorzuteilung [1042, 1147]. Prozeßrechner [702]. Prozeßstauungen [1032]. Prozessumschaltungen [808]. prs [1625]. Prüfbare [363]. pruned [3403]. pruning [3177]. Pseudo [844, 475, 1520, 926, 2174, 382, 1843, 2106, 2442, 4384, 2608, 637]. pseudo-gradient [382]. Pseudo-Gradientenverfahren [382]. Pseudo-Inverses [844]. Pseudo-Lévy [2442]. Pseudo-Random [475, 926, 1843, 2106, 2174, 4384].

Pseudo-Recursive [1520]. Pseudoinverse [860, 2182, 1737]. Pseudolinear [1315]. Pseudorandom [1638, 1991]. Pseudospectra [2203]. Pseudozufallszahlenfolgen [637]. Psi [1648, 842]. PSO [4046, 4069, 3704, 3121, 3942]. PSO-based [3704]. Psychologie [482]. psychometric [3098]. PTL [3231]. PTL-CFS [3231]. Public [2885, 3066, 3547, 3572, 3592, 4075]. Publications [67]. publicly [3297]. publishing [3799]. Pudlatz [730]. Pufferspeicherung [412]. Pugachev [149]. Pure [1967, 2403, 35, 36, 682, 40, 42]. Purpose [1837]. Pylon [4296]. Pyramid [4288, 2310]. Pyramidal [4270]. Python [2851].

Q [4074, 3640]. Q-learning [4074, 3640]. QoE [3107, 3656, 3649, 3650]. QoE-aware [3650]. QoE/QoS [3656]. QoPC [3390]. Qos [3897, 3655, 3107, 2854, 3159, 3523, 3110, 3654, 3399, 3787, 3687, 3390, 3656, 3669, 4041, 3458, 3508, 3649, 3322, 3149, 3653, 3509, 3206]. QoS-aware [3653, 2854, 3159, 3787, 3687, 3669]. Qos-based [3897, 3399, 3206]. QQLAOA [4074]. QR [2740]. Quadrate [195, 431, 797]. Quadratic [4174, 875, 2544, 2590, 2491, 855, 2148, 4354, 4306, 2282, 1948, 4341, 1829, 55, 2072, 2559, 1649, 1816, 1933, 1614, 1799, 1902, 2907, 2605, 1375, 1308, 2140, 1053, 1992, 2614]. Quadratic-Programming [4306]. Quadratically [887, 1245, 1422]. quadratische [887, 855, 1308]. quadratischen [1933, 1053]. quadratischer [1948, 55]. Quadratur [2010, 983, 850]. Quadrature [2494, 2430, 918, 2023, 708, 2429, 144, 993, 2093, 850, 1565, 2240, 2037, 1471, 1014, 607, 563, 716, 1585, 667, 395, 1485, 1738, 1091, 1764, 1335, 1392, 242]. quadrature-type [1585]. Quadraturen [708]. Quadratures [1623, 837, 1876, 905, 1150, 983, 1101, 948]. Quadraturformeln [993, 471, 1950]. Quadraturkonvergenz [607]. Quadraturverfahren [563, 395, 242]. Quadratwurzelfunktion [370]. Quadric [2445]. Quadrics [2130]. Quadrilateral [2204, 1996, 2620, 2615]. Qualitative [789]. qualitativen [789]. Quality [2439, 2446, 3371, 3505, 4042, 3488, 3333, 3033, 3370, 3337, 4028, 3048, 3988, 3372, 3249, 3546, 3476, 3457, 3637]. quality-adjustable [3048]. Quality-of-service [4042]. Quantifiability [3743]. quantification [3044, 3250]. Quantifier [4144]. quantify [3503, 3838, 3190]. Quantifying [3847]. Quantitative [2351, 182, 3405]. Quantities [4106]. Quantized [1825, 3968]. Quantum [4094, 2484, 3763, 2691, 4093, 4084, 4074, 3683, 4071, 3486]. Quantumized [3398]. Quartic [1623, 2450, 4254, 2791, 863, 2708]. Quartics [1815]. Quasi [2366, 1930, 2085, 1941, 2456, 1605, 1124, 1962, 1034, 2708, 2881, 1097, 1212, 1467, 1470, 1096]. Quasi-Block-Tridiagonal [2366]. Quasi-Interpolants [2456]. Quasi-Interpolating [1962, 2708]. quasi-interpolierenden [1962]. quasi-interpretations [1034]. Quasi-Newton [1930, 2085, 1605, 1097, 1212, 1467, 1470, 1096]. Quasi-Random [1124]. quasi-static [2881]. Quasi-Steady [1941]. Quasigroups [2423]. Quasilinear [2102, 1506, 912, 1315, 1850, 1887, 849, 711, 2367, 1525, 975, 986, 1013, 1208, 1011, 1697, 131]. Quasilinear-Implicit [2102]. quasilineare [1013, 131]. quasilinearen [986, 711]. quasilinearer [975, 912, 849, 1011]. Quasilinearization [2]. Quasioptimal [901]. quasistationare [1941]. Queen [2927]. Quellen [1376, 1683]. Queries [1495, 2930, 3150, 3370, 3872, 3388, 3157]. Query [2792, 3744, 3430, 3900, 3417].

Query-based [2792]. Querying

[4213, 3067, 3208]. questions [2716]. queue [2751]. Queueing [886, 2136, 274, 1644, 1041, 1461, 274]. Queues [333, 341, 1751, 2672, 3243, 3323]. queuing [3563, 3929]. quick [1632, 1136]. quick-sort [1136]. Quickcent [3921]. Quicksort [1015]. Quindiagonal [1109]. Quintic [4260, 523]. Quintics [1815]. Quo [3276]. Quotient [239, 2788, 1984]. Quotientenautomat [239]. QZ [1613].

R [421, 48, 792, 826, 331, 731, 225, 35, 238, 271, 418, 211, 575, 10, 552, 174, 625, 288, 253, 532, 188, 702, 487, 229, 257, 187, 320, 280, 404, 800, 1508, 115, 802, 391, 2892]. R-transform [2892]. R. [392, 416, 319, 325, 211, 512, 361, 575, 486, 119, 323, 121, 287, 362, 378, 488, 547, 656, 1136, 1326, 32, 229, 154, 219, 205, 142, 213, 280, 300, 376, 546, 758, 538, 326, 655]. Radau [563, 2353]. Radau-Typ [563]. Radau-type [563]. Radial [4251, 2574, 2547, 2060, 4331, 2715]. Radiation [1573, 1604]. Radiative [2164, 2056, 2645]. Rading [73]. radio [3626, 3665, 3309, 3217, 3277]. Radiosity [2079]. Radius [4200, 1444, 4339, 2018, 586, 638]. radix [2040]. Radon [2063]. Radtschik [676]. RAFM [3350, 3738]. rainfall [2867]. rainfall-runoff [2867]. Rall [125]. Ralston [233, 334, 546]. Randbedingung [722]. Randbedingungen [619]. Randeigenwertproblemen [679]. Randelement [2010]. Randelement-kollokation [2010]. Randern [1942]. Randmaximumsätze [767]. Random [475, 1330, 926, 1417, 1568, 1851, 738, 1923, 2862, 494, 1871, 2181, 845, 2688, 821, 1124, 1637, 2192, 1159, 4198, 1843, 2106, 1705, 2903, 4384, 3968, 1274, 1249, 1363, 1463, 2174, 3844, 2694, 3053, 637, 3403].

Randomized [2105, 3284]. Randomly [1946]. Randomness [1386, 2686]. Randsteuerung [987]. Randwertaufgaben [760, 275]. Randwertprobleme [297, 1934, 849, 994, 1013, 1999, 689]. Randwertproblemen [410, 1107, 384]. Randwertproblems [1045]. Range [1801, 1316, 1468, 466, 2374, 920, 2035, 1894, 2930, 1455, 979, 2819, 1155, 3394, 3872, 2622, 1127]. Ranges [4101, 990]. Rank [2408, 2467, 2464, 2613, 2542, 2543, 583, 1476, 2779]. rank-deficient [1476]. ranking [3452, 3338, 3721, 3014, 3033, 3823, 3051, 3154, 3147, 3624]. Rannacher [2660]. Raphson [1844, 2041]. Rapid [4088, 3387, 2799]. rapidly [2800]. RAS [1158]. RAS-algorithm [1158]. Raster [4107, 1670]. Rasters [4107]. Rate [4351, 2274, 1920, 606, 4348, 3101, 3335, 445, 3039, 2016, 1685, 2183, 296]. rated [3546]. rater [3012]. Rates [2493, 1115]. Rating [2277]. ratio [2692, 3601, 3312, 2754]. Rational [1786, 4245, 1475, 660, 846, 4316, 4251, 4301, 4208, 2130, 4115, 920, 1614, 1799, 1873, 1083, 4220, 110, 1792, 1836, 1675, 1019, 2643, 1545, 457, 1747, 19, 1490, 1806, 2662, 2658, 1407, 2719]. rationalen [1545, 19]. rationaler [457, 110]. Rationally [2154]. Raum [895]. Räume [118, 711, 4105]. Räumen [440, 409, 433, 468, 469, 632, 526, 724]. Raumes [711, 895]. Raviart [2722]. Ray [2224, 4167]. Ray-Shooting [2224]. Rayleigh [1984]. RDF [3695, 3157, 3328]. Re [1126, 2707, 3883, 3420]. re-detection [3420]. Re-Entrant [1126]. re-identification [3883]. re-weighted [2707]. reachability [4130]. Reaction [2433, 4338, 2525, 2478, 2371, 2318, 2746, 2626]. Reaction-Diffusion [2433, 2525, 2318, 2478, 2746, 2626]. Reaction-Diffusion-Transport [2371]. reactive [3022]. read [3257]. read/write [3257]. reader [3083]. READEX [3171]. readiness [3668]. Readings [353]. Real [1360, 823, 1255, 4183, 1827, 4144, 4145, 1468, 4101, 3194, 2077, 863, 4135, 739, 1882, 2361, 1198, 3682, 692, 3984, 4017, 3359, 4184, 393, 2612, 1140, 3764, 3842, 3460, 2960, 3367, 3420, 3302, 1938]. Real-Time [1882, 3194, 3682, 3984, 4017, 3764, 3842, 3460, 3302, 1938]. Realistic [2022]. reality [2797, 3560]. realizability [2717]. realization [1629, 2821]. Realizations [2384]. Realized [1234]. Realizing [3867, 3910]. Reallocation [3204]. Reasonability [3021]. Reasoning [3025]. ReCAP [3255]. receiver [3143]. receiving [3824]. rechargeable [3773, 3535, 3807]. Rechenanlagen [200, 10, 4105, 234, 349, 799, 4128, 97, 272, 615, 84]. Rechenaufwandes [1008]. Rechenautomaten [74, 478]. Rechenberg [74]. Rechenstudie [1946]. Rechensysteme [531]. Rechentechnik [255]. Rechenzeit [1082]. Rechenzeitvergleich [437]. Rechnen [4105, 862]. Rechnerarithmetik [4104]. Rechteckgittern [1921, 1953]. rechter [789]. recipe [3902]. Reciprocal [922]. Recognition [1133, 2257, 4280, 4267, 726, 4215, 4258, 22, 3744, 3863, 3526, 1007, 3141, 3990, 4379, 2892, 2990, 3997, 3975, 4134, 3582, 2802, 3147, 3377, 3286, 3212, 3426, 3501, 122]. recommendation [3714, 3689, 3906, 3496, 3902, 3124, 3937, 3146, 3233, 3797, 3475]. Recommendations [3412, 3408, 3693, 4013]. recommender [3043, 3545, 3364, 3903, 3203, 2957]. recommending [3278]. Recompression [2500]. reconfigurable [2218, 2001, 2926, 3795]. reconfiguration [3533, 2967]. Reconstructing [3859, 1248]. Reconstruction [4280, 4206, 4322, 4304, 2027, 1594, 1095, 3778, 4020, 3291]. Record [2152]. recorded [2954]. Recording [240]. records [3422, 1322]. recover [3465]. Recoverable [3678]. Recovery [2522, 2339, 3180, 3579, 3344, 3023, 2900, 3414]. Rectangle [2294, 1643]. Rectangular [1550, 1921, 1953]. Recurrence [1792, 1836, 663, 1302, 1780]. recurrences [1542, 1582]. recurrent [2752, 3935]. Recursion [77]. recursions [1685, 2183]. Recursive [1520, 1851, 2012, 726, 373, 2271, 909, 1313, 1570]. Recursively [1346]. Reduce [1505, 3145]. Reduced [1625, 2493, 1522, 1574, 385, 495]. Reducible [477]. Reducing [2076, 857, 3794, 3874, 3859, 917]. Reduction [438, 4298, 1901, 2376, 2508, 2560, 1932, 2237, 4311, 3573, 3763, 2770, 3170, 847, 3401, 3811, 1208, 3038, 3663, 3489, 3639, 1206, 3819, 2736, 1032]. Reduktion [1932]. Redundancy [1946, 1017, 3351, 2996, 3652, 3176]. Redundant [2281, 2351, 1495, 435, 524]. Redundanz [1946]. reduzierter [495]. Reduzierung [1032]. reefs [3530]. reellem [4135]. reellen [823, 863]. reeller [4101, 393]. reentrant [2621]. refactoring [3139]. Reference [2379, 2615]. Referencing [1131]. Refined [2424, 2579, 2865]. Refinement [2146, 2235, 2098, 2253, 2374, 2029, 2462, 2254, 1822, 1900, 1305, 3260, 2005, 2178]. Refining [1598, 3895]. Reflecting [2438]. reflective [3447]. reformulation [2698]. regard [384]. Regarding [942, 3647, 2896, 2895]. Regel [1269]. Regelkreistheorie [624]. Regeln [123]. regelprozesse [319]. Regelsysteme [325]. Regelsystemen [130, 161, 115]. Regelungen [810]. Regelungssysteme [60, 733]. Regelungstechnik [65, 284]. Regelungstechniker [489]. regenerative [3976]. Region [4287, 2147, 2119, 769, 920, 4291, 1288, 1602, 1918, 2034, 1459, 3581, 1669, 2889, 2789]. Regions [2048, 4307, 3745, 3914, 112, 4234].

Register [655, 3257]. Registration [2374, 2489, 3998, 3986, 3385]. registry [3951]. Regression [820, 2385, 174, 2471, 1593, 2707, 2828, 1004, 3403, 1349, 4048, 3173, 1144, 1257, 1341]. regressions [3144]. Regula [595, 526, 503, 1234, 1447, 595]. Regula-falsi [1234, 1447]. Regula-falsi- [595]. Regula-falsi-Typ [1234]. Regular [1337, 876, 1472, 1241, 1865]. regulären [876]. regularisation [2890]. Regularisierung [1960]. regularities [2901]. Regularity [4359, 2054]. Regularization [1717, 1960, 1777, 2539, 1665, 1629, 1982]. Regularized [2490, 2563, 2489, 2589, 1755]. rein [682]. reinforcement [3857, 3530, 4019, 3893, 4078, 3928, 3832, 3894, 3994, 3411, 3774, 3944, 3970, 3867, 3650, 3868, 3760, 4040, 4008, 3509, 3063]. Reinitialisierung [1968]. reinitialization [2849]. Rejection [2540, 2307, 2317, 1952, 1751]. rejuvenation [2971]. rekurrenter [663]. Rekursionsformel [77]. rekursive [2012]. Related [2496, 4188, 192, 213, 2633, 3172, 1980, 4369, 3699]. Relatedness [1119]. relating [2716]. Relation [4293, 108, 79, 1659]. Relational [4283, 1265, 1495, 4285, 3400, 3067]. relationaler [1265]. Relationen [663]. Relations [2269, 210, 2242, 2128, 1792, 1836, 663]. Relationship [2341, 3287, 3327]. relationships [3721]. relationstheoretischer [224]. relativ [1128]. Relative [1128, 1835, 1511, 1197, 1669, 1847]. Relative-Optimal [1128]. Relatively [2361]. Relaxation [1478, 1163, 2240, 2074, 307, 1564, 3751, 1325, 1260, 452, 1977, 1248, 1469]. Relaxationen [2009]. Relaxations [2009]. Relaxationsparameter [1260].

Relaxationsverfahrens [307]. relaxed [3323]. Relaxtionverfahrens [108]. relay [2953]. Release [2476, 4030]. Relevance [3797]. Relevance-aware [3797]. relevant [3021, 4138]. Reliability [3964, 1791, 3367, 2492, 3626, 3665, 2776, 3924, 3368, 3047]. Reliable [416, 1917, 2560, 196, 2108, 2482, 1263, 4057, 2916, 3438, 3638, 4239, 1407, 1718, 3609]. reliably [3497]. relief [272]. Reliefdarstellungen [272]. rellen [692]. Remainder [4147, 893, 1021]. Remark [942, 1284, 572, 1530, 1805, 1405, 1048, 1369, 24]. Remarkable [2347]. Remarks [1389, 315, 2205, 1641, 1277, 2182, 879, 1904, 781, 1215, 607, 693, 1945]. Remembering [3451]. Remez [1304, 1377, 2006, 1566]. Remez-type [1566]. Remote [138, 3643, 2985, 3072, 3732, 3418, 3231, 3385]. Removing [4306]. renameability [873]. rendezvous [3497, 2997]. renewable [3132]. Renewal [1631]. Rentrop [1430]. Rentrop-methods [1430]. Reordering [2057, 2973]. Rep [4312]. Repacking [1947]. Repair [565, 2709, 480]. repairable [3086]. repairing [3184, 3017]. repairmen [3218]. Repeated [1434, 4252, 3133]. Repertorium [429]. replacement [4130, 3459]. replication [3263]. replications [3755, 3754]. Replies [3634]. reply [1845]. Report [372, 391, 543, 539, 72, 45, 178, 143, 184, 171, 220, 232, 234, 492, 173, 14, 330, 331, 185, 221, 225, 16, 35, 41, 186, 179, 200, 238, 169, 285, 156, 578, 36, 74, 147, 126, 201, 148, 137, 181, 149, 536, 65, 138, 337, 226, 118, 145, 283, 99, 215, 235, 302, 214, 180, 218, 140, 355, 756, 157, 122, 328, 195, 222, 100, 144, 139, 120, 119, 193, 224, 230, 151, 323, 281, 329, 301, 236, 97, 101, 103, 94, 64, 194, 153, 174, 303, 327, 287, 488, 146, 190, 191, 197, 75]. Report [104, 321, 253, 532, 198, 188, 513, 98, 34, 39, 237, 582, 30, 95, 31, 33, 38, 43, 229, 116, 745, 192, 96, 141, 154, 251, 63, 47, 117, 216, 227, 155, 219, 152, 124, 176, 203, 187, 142, 217, 196, 260, 254, 135, 177, 127, 150, 199, 204, 228, 324, 247, 269, 334, 282, 248, 252, 250, 256, 182, 183, 231, 284, 322, 538, 202, 326, 175, 158, 223, 1539, 249, 59]. repost [3477]. repräsentative [1318]. repräsentativen [757]. Representation [4295, 959, 4204, 2260, 786, 2543, 2453, 4256, 893, 2473, 1180, 4328, 4281, 2602, 2743, 1675, 2797, 3375, 3904, 3379, 463, 3731, 43, 3267, 3905, 958, 1395]. Representations [2488, 36, 787, 4270, 4210, 1420, 4257, 2379, 4379, 1114, 4229, 272, 1056, 2719, 184]. representative [3264, 1318, 3457]. representative-based [3457]. represented [1494]. Representing [4213]. Reproducibility [3255]. reproduction [1]. reputation [3821, 3655, 3012, 3449]. requester [3014]. requirement [3594]. requirements [3834, 1079, 3920, 3509]. resampling [3825]. Research [539, 1145, 70, 67, 68, 426, 427, 650, 151, 648, 649, 627, 700, 581, 104, 744, 628, 540, 530, 534, 582, 651, 248, 672, 3376, 3979, 3026, 3506, 3648, 3583, 51, 3371, 3589, 3833, 2759, 48, 327, 47, 205]. Residual [4165, 2000, 4361, 2482, 1008, 4233, 1571, 1733]. Residual-based [2482]. Residuals [2108]. Residue [1959]. Residuum [2000, 1008]. resilience [3492, 2919]. resiliency [3240]. resilient [3490, 3493, 3838, 3491]. Resnet [3799]. Resolution [2495, 825, 3384, 3148, 1241, 2927, 3988, 3720]. resolutions [3027]. Resource [1170, 3403, 3351, 3857, 3530, 3070, 3615, 3834, 3940, 3901, 3104, 2992, 3684, 3832, 3419, 3964, 4072, 3193, 4066, 3642, 3881, 4063, 3053, 3030, 3960, 3433, 3195, 3149, 3106, 4082, 4021, 3077, 3562, 3283, 3839, 4064, 3095, 3483, 3602, 3637]. Resource-efficient [3403]. Resources [2151, 4012, 3296, 3496, 3329, 2846, 3190, 3482]. Respect [845, 1945]. Respecting [1871, 3226]. Response [1500, 3923, 3122, 1042]. responsibility [3840]. Restart [1968]. RESTful [3910]. Restglied [893]. Restgliedes [1021]. Restklassenzahlensystemen [1959]. Restoration [4267, 2638, 3387, 2636]. restoring [3845]. Restricted [2035, 932, 4365, 592]. restrictions [3902]. Result [4169, 2323, 1980, 4373, 4233, 1326, 1728]. Results [4179, 4350, 1512, 2142, 2278, 2179, 1289, 1450, 210, 2216, 79, 1774, 4170, 2054, 2350, 1073, 3417, 1196, 2936, 2879, 1708, 2222]. resume [1204]. Retarded [1768, 1776]. RETGCN [3985]. rethinking [3698]. retinal [2754]. Retracted [3189]. Retraction [3738, 3058]. Retrieval [2081, 3400, 2979, 3075]. Return [4351, 2796]. Reupprecht [236]. reuse [3968, 3930, 3099, 3454, 3978]. Reusing [2990, 3434]. Reusken [2088]. revenue [3782]. Reversal [2479]. Reverse [2604, 2213, 2461, 2761]. Reversible [940, 2078, 2995]. Revidierte [461]. Review [89, 304, 246, 335, 353, 354, 374, 375, 392, 420, 421, 419, 422, 542, 544, 545, 673, 814, 44, 134, 268, 602, 509, 490, 792, 817, 533, 531, 483, 73, 416, 603, 105, 70, 87, 26, 67, 68, 27, 28, 102, 69, 286, 289, 271, 319, 325, 429, 418, 426, 427, 551, 577, 579, 580, 646, 639, 12, 388, 485, 484, 675, 674, 652, 811, 85, 15, 46, 71, 261, 262, 424, 377, 60, 727, 732, 305, 347, 349, 114, 211, 350, 360, 358, 514, 430, 512, 515, 757, 698, 654, 264, 299, 361, 575]. Review [659, 799, 431, 486, 90, 336, 401, 641, 650, 796, 291, 357, 648, 649, 797, 406, 815, 10, 121, 552, 13, 88, 92, 11, 290, 279, 362, 378, 623, 627, 576, 549, 625, 547, 701, 700, 656, 626, 581, 670, 671, 729, 743, 746, 42, 288, 812, 66, 425, 744, 794, 333, 628, 540, 702, 685, 487, 530, 534, 657, 651, 644, 653, 816, 338, 29, 37, 32, 61, 332, 263, 356, 363, 677, 642, 647, 389, 481, 510, 172, 390, 86, 40, 212, 257, 265, 699, 809, 818, 170, 205, 266, 267, 91, 93, 125, 213, 233, 255]. Review [270, 320, 258, 280, 300, 351, 373, 402, 352, 404, 405, 376, 604, 537, 546, 643, 759, 758, 813, 800, 801, 793, 62, 676, 115, 136, 403, 482, 428, 462, 489, 491, 493, 535, 645, 802, 810, 795, 359, 423, 417, 798, 348, 548, 624, 550, 541, 672, 733, 629, 686, 728, 730, 511, 640, 658, 655, 123, 3392, 826, 3148, 731, 3789, 3924, 3828, 3897, 3816, 3379, 3428, 3303, 829, 832, 827, 831, 833, 828, 830, 3700, 3792]. reviews [898, 3210, 3051, 884]. revised [1022]. Revising [3111]. Revisited [2392, 2284, 4330, 2725, 1845]. revocable [3308]. revolution [3469]. Rew [3927]. Rew-LSTM [3927]. Rewards [840]. reweight [3927]. reweight-long [3927]. rewriting [3870, 4131]. Rheingans [515]. Riccati [2413]. Rich [257]. Richards [188, 2875]. Richardson [611, 127, 1452, 611, 21, 1474, 2775, 1452]. Richardson-Algorithmus [21]. Richardson-Extrapolation [611, 1452, 611]. Richter [797]. Riemann [143, 1632, 1269]. Riemann-Problems [1632]. Riemannschen [1269]. Riesenfeld [2607]. Right [4293, 1858, 789, 1279]. right-hand [1279]. Rigorous [2022]. Rigorously [2221]. Ring [2120, 3324, 2842, 2725]. ring-structured [2842]. RISC [3778]. RISC-V [3778]. Risiko [197]. Risk [2386, 3408, 3122, 3415]. Risset [532]. Ritout [320]. Rittel [647]. Ritz [836, 1660, 520]. Ritz-Galerkin [836]. Ritz-method [520]. Ritz-Verfahren [520]. RL [3070]. Rndom [149]. RNe2Vec [3477]. RNN [3386]. RNN-based [3386]. Road [3554, 3474, 3452, 3872]. Robbins [219]. Robert [508]. Robin [258, 1132]. robot [3813, 3480]. robots [3317, 2603]. Robust [4335, 2095, 2319, 2453, 3969, 4271, 1854, 2116, 2459, 2300, 3984, 3281, 2746, 3736, 2802, 2840, 3025, 3725, 3286, 3905, 1914]. Robuste [1914]. Robustness [2036, 2075, 3890, 3503, 2626]. rock [3553]. Roemer [950]. ROI [3824]. Role [623, 4215, 2898, 3296, 3329]. rolling [2608]. Romberg [680, 634, 680, 634]. Romberg-Integration [680, 634, 680, 634]. Root [4332, 1360, 1765, 859, 240, 4356, 370, 955, 1844, 1268, 2132, 4363, 1862, 2190, 871, 2145, 1384, 1443, 1352]. Root-Finding [4363, 1384]. rooted [1217, 1326]. rootfinding [4226]. Rooting [1106]. Roots [2592, 4350, 4099, 2515, 721, 1198, 943, 692, 1222, 1005, 379, 1422, 1718]. Rosenbrock [150, 831, 2102, 1263, 1474, 1775, 1785]. Rosenbrock-Type [2102, 1785]. Rosenfeld [727]. Rosser [1269]. Rosserschen [1269]. rotating [2877]. Rotation [2346, 314]. Rotational [2213]. Rotations [57, 2705, 1613]. Rothe [1045]. Rotheverfahren [1045]. Rottmann [190, 288]. Rough [2520, 3551, 3235]. Rougier [320]. Round [913, 627, 709, 1132, 2690, 2947]. Round-Off [913, 627, 709]. round-trip [2947]. Rounded [4096, 4123]. Rounding [1448, 556, 2024, 4357, 1306, 1379, 1498, 4121, 155, 4122, 4107, 1545, 1070, 2662, 2658, 752]. Roundings [4111]. Roundoff [1496, 2659, 1197, 781]. Route [791, 930, 3408, 461, 2945]. router [3810, 2749]. router-level [2749]. routers [3040]. Routes [1113, 2002]. Routh [1247]. routine [596, 1354]. Routing [2271, 3641, 3537, 3107, 2986, 2952, 3657, 3523, 3040, 4044, 3578, 3854, 2999, 3884, 3300, 3243, 3356, 2936, 3812, 4057, 2947, 3427, 3470, 3673, 3136, 3217, 4042, 3456, 4077, 3206]. Row [1968, 244, 847, 554, 1710, 1813]. Rows [2366]. RPG [288]. RPL [3656, 4052, 3527]. RPL-based [4052]. RSA [2689]. RSS [3384]. RSSI [3898]. RTOS [3035]. RUCIB [3836]. Rudeanu [327]. Rudolf [605, 1223]. Rudwick [542]. Rule [1648, 2392, 1927, 1951, 130, 161, 1458, 1493, 1662, 1269, 3836, 3920, 2792, 2744]. rule-based [3836]. Ruled [4301]. Rules [1963, 1771, 2023, 1565, 1040, 2765, 2729, 1433, 2599, 716, 2904, 1392, 3168, 2697]. Rumor [3634, 3989]. Run [2061, 1974, 2899]. Run-Length [2061, 1974]. run.time [2898]. Rundreise [461, 278]. Rundung [1545, 4107]. Rundungsfehler [556, 484, 752]. rundungsgenaue [1070]. RUNGE [364, 207, 1967, 1954, 1533, 467, 888, 129, 1492, 995, 1339, 1358, 1526, 1964, 2748, 2785, 1910, 2179, 1532, 1839, 1831, 1954, 1768, 1833, 1988, 1830, 2384, 1588, 2091, 364, 467, 888, 1491, 2166, 207, 1628, 1617, 923, 2065, 631, 691, 1747, 1057, 2172, 1989, 1654, 1726, 1339, 1526, 1677, 1964, 1770, 1083, 725, 1678, 2293]. RUNGE-KUTTA-Formeln [364]. RUNGE-KUTTA-Verfahren [207]. Running [2556, 2327]. runoff [2867]. Runtime [4001, 3781, 2961, 2830]. Russian [811]. Rutishauser [389, 481, 404, 529]. RWP [3892]. Ryder [493].

S [509, 483, 271, 12, 149, 138, 305, 167, 264, 90, 336, 144, 641, 281, 13, 194, 327, 338, 96, 389, 481, 212, 187, 233, 228, 334, 546, 62, 548, 1145, 1700]. S. [545, 26, 578, 811, 138, 756, 641, 623, 75, 98, 34, 39, 196, 248, 676, 231, 482]. S.I [3674]. SaaS [2757, 2980]. Saaty [323]. Sabin [2035]. Saddle [2320, 2793, 2681]. SADM [3495]. SADM-SDNC [3495]. Safe [1735, 2048, 3609]. safety [3408, 3021]. safety-relevant [3021]. Saite [987]. Salesman [2066, 461, 278]. Salomaa [659]. salp [3669]. Samarskii [42]. Samelson [10, 389, 481]. Sammet [533]. Sammlung [46]. Sample [820, 1116, 1071]. Sampled [116, 550, 2972]. Sampled-date [550]. Sampling [788, 1216, 2062, 2167, 966, 1779, 1746, 2523, 1159, 3744, 1233, 1734, 3384, 1129, 3835, 583, 3510]. Sampson [826]. Sandwich [1920, 1214]. Sandwich-Algorithmus [1920]. Sansone [35]. SAR [2256]. SAT [3870, 3466, 3409]. SAT-based [3409]. satellite [4044]. satisfaction [2820]. Satisfiability [1802]. satisfied [500]. SATS [3709]. Saturated [2302, 3664]. Satz [607, 369, 38]. Sátze [342]. Sauer [238, 418, 119, 10, 287, 362, 488, 547, 656, 238, 508]. Savas [483]. save [2813]. saving [3804, 4008, 3423]. Saxon [305]. Scalability [3645, 3112, 3741, 4038]. Scalable [2696, 3164, 2152, 3263, 3648, 3048, 2972, 2651, 3954, 3783, 4087, 2922, 3457, 3679]. Scalar [4174, 1168, 1712, 2910, 3002, 2554]. Scale [2413, 2033, 3649, 4261, 3780, 3170, 2636, 3547, 3117, 2931, 3890, 2946, 3218, 3337, 3669, 4043, 2998, 2835, 4055, 3921, 2843, 3965, 3366, 3705, 3389, 3986, 3091, 2994, 3420]. scale-free [3890, 3921]. Scale-Invariant [4261]. Scaled [1828, 4354]. Scales [2155, 1629]. Scaling [2270, 2409, 1605, 2020, 456, 2839, 776, 3293, 3149]. scaling-invariante [776]. scans [4082, 3737]. Scatter [1232]. Scattered [2035, 2229, 4218, 2547, 2715]. Scattering [2518, 2375, 2519, 2520, 2521]. scenario [2763, 3449]. scenarios [3266, 3898, 3482]. Scene [2379]. Scenes [1519, 4213]. Schaack [800]. Schalen [1063]. Schaltalgebra [756, 321]. Schaltkreis [202]. Schaltkreis- [202]. Schaltkreise [133]. Schaltkreisen [146]. Schaltungen [136]. Schaltwerk [816, 202]. Schaltwerk- [816]. Schaltwerk-Theorie [202]. Schärf [832]. schedule [3773, 3307, 3807]. scheduled [3101]. scheduler [3001, 3181]. schedulers [2804, 3094]. Schedules [1178, 2479]. Scheduling [1178, 3183, 2357, 1837, 2232, 4012, 2414, 2470, 2540, 2105, 3359, 1881, 2250, 2307, 4191, 1042, 2420, 1853, 2391, 2476, 3842, 2157, 1072, 2002, 1882, 3625, 2151, 1935, 2826, 3188, 3199, 3611, 3398, 3220, 2217, 3197, 3223, 3945, 3594, 4027, 1165, 1979, 3706, 3550, 3129, 3504, 3822, 3035, 3964, 3194, 3603, 3774, 3896, 3522, 4041, 4074, 3761, 3950, 3803, 4051, 3368, 3121, 3151, 3078, 3794, 3191, 3760, 2846, 3942, 3553, 3790, 3448, 3367, 3612, 3283, 3535, 2932, 3027, 3084, 3886, 3707]. Schema [3327, 2964]. Schemata [882, 1172]. Scheme [1794, 2233, 4204, 1833, 2372, 2152, 1834, 2119, 1379, 1385, 2327, 2367, 1761, 3643, 3863, 2591, 3471, 3806, 2630, 3184, 3129, 2985, 3337, 3356, 1693, 2595, 3862, 2017, 2682, 3340, 2745, 1535, 4034, 2718, 3604, 3027, 3102, 3805, 3820, 4075, 4002, 3539, 3709, 4011, 3308]. Schemes [2341, 2575, 2257, 1506, 1789, 2196, 2166, 748, 2280, 2338, 2340, 2332, 1887, 1406, 1888, 2126, 1996, 310, 2125, 1064, 2753, 1822, 2857, 1224, 1715, 2674, 3428, 445, 1640]. Schieferdecker [812]. Schilling [14]. Schlaepfer [658]. Schleifen [1970]. Schleifen-Invarianten [1970]. Schmidt [1995, 1995, 937, 1719, 25]. Schmidt-Verfahren [1995]. Schmitt [198]. Schneider [235]. Schnelle [37, 1462, 566]. schnellen [314, 805, 998]. schneller [1632, 2013]. Schnittebenenverfahren [904]. Schnurr [2597]. school [3138]. Schopf [38]. Schranke [1070]. Schranken [466, 442, 839, 164, 275, 634, 1936]. Schrankentreue [1236]. Schrankenzahlen [273, 773]. Schriftenreihe [579, 66]. Schrittverfahren [988]. Schrittweiten [467, 888, 364]. Schrittweiten-Kontrolle [467, 888, 364]. Schrittweitenalgorithmus [1951]. Schrittweitenfunktionen [1027]. Schrittweitensteuerungen [1349]. Schroder [4362, 1809]. Schroder-like [4362, 1809]. Schrödinger [2570, 2784, 2611, 2355, 1328, 1820]. Schrödinger-KdV [2784]. Schroeppel [1362]. Schulstundenplänen [224]. Schultze [581]. Schulz [1600, 1656, 1641, 935]. Schulz-Methods [935]. Schulz-Typ [935]. Schuppe [551]. Schur [2314, 2505]. Schütte [156].

Schuwalowa [509]. schwach [619]. schwachen [554]. schwachsingulären [241]. Schwarz [211, 265, 404, 428, 2275, 2103, 2627, 2356]. schwingenden [987]. Schwingungen [971]. Schwingungsgleichung [385]. Science [375, 419, 455, 3755, 3754, 137, 3616, 3397, 3118, 2759, 126]. sciences [185, 71, 623, 251, 229]. Scientific [1634, 4186, 4373, 3887, 2653, 2654, 3255, 3121, 3094, 3794, 3216, 4239, 3091, 3792]. scientists [3674]. scores [2965]. scoring [3624, 3290]. Scott [2667]. Screen [2103, 2129, 2935]. Screening [3556, 3730]. screens [1182]. Scripte [675]. Sculpting [4325]. SDN [3857, 4046, 3875, 3442, 4021, 3919]. SDNC [3495]. SDWSN [3671]. SeaDoc [2956]. Seamless [3500, 3445]. Search [2409, 1724, 2202, 1653, 1583, 2484, 2927, 1854, 1875, 1787, 1124, 1774, 3639, 3887, 596, 3422, 3666, 3177, 2978, 3444, 1293]. Searches [2270]. Searching [1668, 1232, 1113, 1246]. seasonal [3281]. Secant [955, 1373, 1230, 934]. Second [2099, 1468, 1506, 870, 838, 908, 88, 1932, 1550, 2239, 2465, 2111, 1578, 1161, 2633, 2727, 1684, 1061, 1715, 470, 444, 1535, 1528, 2638, 1358, 3562, 1201]. Second-Order [1506, 838, 1578, 2638, 1358]. secondary [3940, 1545]. secret [2694, 4011]. section [2717, 4014]. Sectionally [1549]. Sections [1032]. sector [1186, 3235]. Sectors [4316]. Secure [3507, 3912, 3422, 1024, 3655, 3746, 2976, 3356, 3642, 4086, 3456, 3635, 3653, 4071, 3102, 3805, 3483, 2994]. secured [3513]. Securing [3977, 4062, 3337]. security [3160, 4006, 3298, 3410, 3893, 3662, 3572, 3495, 3816, 3485, 3109, 3508, 3266, 3044, 3999, 4031, 3544, 3519, 3142, 2760, 3509]. security-aware [3109]. seeds [3598]. Segment [2474, 2530]. Segmentation [4288, 4287, 4286, 4276, 2461, 3728, 3581, 3288]. Segmented [1252, 1303, 2374, 2223, 1173]. segments [3452]. Seibt [643]. Seidel [1961, 2049]. Seidel-Algorithmen [1961]. Seidel-Algorithms [1961]. Seiffart [653]. sein [1947]. seine [490, 1254]. seismic [4167]. Seite [789]. selbstanpassende [319]. selbsttätiger [60]. select [3146]. Selected [388, 1319, 2958, 3320, 3052, 3105, 3246]. selecting [3932]. Selection [3159, 1583, 1495, 2379, 4056, 3655, 3730, 3014, 847, 3421, 3897, 1650, 3669, 4041, 4035, 3497, 1079, 4005, 3309, 3898, 3080, 1318, 3868, 3330, 3582, 3962, 3028, 3524, 3025, 3232, 3531, 3144]. selective [3992, 4052]. selectivity [3917]. Self [3621, 2570, 2243, 4172, 2207, 2378, 3701, 3017, 3781, 3780, 1756, 1821, 2919, 2834, 3110, 1086, 3672, 3636, 3016, 4028, 3667, 3867, 2758, 2967, 3622, 3480, 3747, 3531, 3883, 2956, 3905]. self-adaptation [3747]. self-adaptive [3781, 3780, 2919, 2834, 3867, 3480, 2956]. self-adjusted [3531]. self-advising [3672]. self-attention [3883]. self-attentions [3905]. self-calibrated [4028]. Self-healing [3621, 3110, 2758]. self-learning [3667]. self-organization [3672]. self-organizing [3016]. self-reconfiguration [2967]. Self-repairing [3017]. Self-similar [2570]. Self-Stabilizing [2243, 2207, 2378, 3701]. self-tracked [3622]. self-tuning [1086]. Self-validating [4172, 1756]. self-verified [1821]. selfadaptive [4168]. Selfcomplementary [869]. selfdual [1751]. selfish [3821]. Semantic [4279, 3363, 3073, 3461, 3581, 3848, 3577, 2979, 3329, 3482, 3157, 2925]. Semantically [3910]. Semanticizing [1265]. semantics [2949, 3795, 3233]. semantisierung [1265]. Semi [2294, 2470, 2549, 2250, 4341, 1872, 2360, 390, 1413, 2593, 2589, 1138, 2679, 365, 1852, 1940, 2861, 3983, 3414, 3885, 3860, 365]. Semi- [2294]. semi-definite [2861].

Semi-Discrete [1138]. semi-greedy [3983]. Semi-Groups [390]. Semi-Implicit [1413, 2549]. Semi-Infinite [1940, 4341]. semi-infiniten [1940]. Semi-linear [1872]. semi-logarithmic [1852]. Semi-Online [2470, 2593, 2679]. Semi-smooth [2589]. semi-supervised [3885, 3860]. Semi-Thue-Systemen [365]. semi-Thue-systems [365]. Semi-Unstructured [2360]. Semiconductor [2100, 2004]. Semidefinite [2280, 2588]. Semigroup [2137]. Semigroups [763, 1114]. Semilinear [2497, 1999, 2304]. semiorder [54]. Sengupta [2597]. sensing [3570, 3117, 3525, 3072, 3418, 3231, 3385, 3700]. Sensitive [1133, 2325, 3625, 3081, 4027, 3889, 3460, 3278, 312, 3383]. Sensitivitätsanalyse [628]. Sensitivity [2824, 3574, 3969, 3190]. Sensitivity-based [2824]. Sensor [3200, 3525, 3559, 3365, 3552, 3221, 3802, 3353, 2997, 3773, 3661, 3570, 3578, 3129, 2688, 3275, 3300, 3569, 3401, 3242, 2920, 3048, 2876, 3179, 3497, 2844, 3427, 3926, 2987, 3019, 3845, 3929, 3535, 3776, 3807, 3063]. sensors [3526, 3020, 3554]. sensory [3137]. Sentence [29]. sentiment [3488, 3979]. sentiment-based [3488]. Seoratschek [545]. Separable [468, 409, 433, 2542]. separablem [409, 433]. separablen [468]. Separated [851]. Separating [2463]. separator [1865]. Sequence [4288, 3995, 3050, 3241, 1701, 2954, 3289]. Sequences [2269, 4315, 2144, 2185, 4147, 448, 2557, 2767, 1334, 1367, 664, 1405, 1693, 637, 1261, 2614]. Sequencing [2380, 1040]. Sequential [1336, 299, 133, 1297, 1416, 655, 1230, 592, 4013]. Sequential-Parallel [1297]. Sequentialtests [592]. sequentieller [133]. serendipity [3364]. serendipity-oriented [3364]. Serial [1658]. Series [354, 420, 542, 35, 646, 36, 2167, 401, 1291, 42, 910, 1427, 163, 572, 40, 1040, 2442, 2743, 1675, 3712, 1481, 1090, 2783, 4, 3967, 3382, 991, 714, 3332, 3381, 4064, 249, 172]. series-based [3712]. Series-Parallel [1040]. Server [2414, 2136, 3110, 3983, 3795, 4018, 3722]. serverless [3987, 3961, 4007, 4061]. servers [3263, 3394, 3719, 3248, 3342, 4023, 3429]. Service [3159, 3014, 3113, 341, 3947, 3655, 1644, 3180, 3272, 3866, 3244, 2915, 2834, 3473, 3953, 3399, 4076, 3907, 2912, 3079, 3897, 3687, 2911, 3200, 3560, 3669, 3109, 3599, 3566, 3505, 3568, 3096, 3124, 3336, 3080, 4042, 3518, 3138, 3668, 3971, 2900, 2924, 3815, 3186, 3252, 3025, 3095, 3211, 3457, 2994, 3637, 3955]. service-based [3109, 2994]. Service-oriented [3947, 3866, 2911, 3096, 3138, 2924]. Service-requester-centered [3014]. Services [3294, 3488, 3013, 3528, 3481, 2977, 3746, 2757, 3471, 3788, 3117, 3910, 3118, 2925, 3334, 3667, 3988, 3589, 3311, 3566, 4063, 3120, 3508, 3568, 3080, 3314, 3075, 3278, 3222, 3436, 3140, 3268, 544]. serving [3987]. servoing [3990]. Sesitivitätsanalysen [511]. session [84, 3797]. session-based [3797]. Set [1167, 2081, 784, 1188, 2153, 2329, 1365, 3069, 1415, 4343, 2492, 1935, 2849, 456, 2633, 1621, 2788, 3859, 3178, 2820, 2635, 1516, 2637, 3551, 3235, 3874, 1483, 1191, 2878, 1935]. Set-based [3069]. Set-Oriented [2081]. Set-Up [1935]. Set-up-Zeiten [1935]. set-ups [456]. Set-Valued [1167, 2820]. Sets [4110, 4111, 4349, 1783, 210, 1724, 2256, 2420, 869, 2516, 2587, 1267, 636, 1222, 1695, 3210, 2652, 3215, 1293, 1181, 2882, 2902, 2964, 593, 1191]. setting [3297]. seventh [364]. seventh-order [364]. Several [2549, 4181, 1855, 3000, 2704, 53]. severity [4035, 3415]. SG [3983]. SG-MOACO [3983]. SGM [3610]. Shading [2122]. Shafer [2821]. Shan [96]. Shape [4296, 2117, 4248, 4300, 4319, 4303, 4306, 2122, 4217, 4260, 1799, 2460, 2743, 1370, 2628]. Shape-Failures [4306]. Shape-Preserving [4319, 4260]. Shaped [4331, 2713, 2739]. sharding [4038]. share [3782]. Shared [2503, 1442, 2938, 3123, 3190, 3623]. Sharing [2136, 3133, 3134, 3746, 2912, 1321, 3622, 3494, 3544, 4075, 1356, 4011, 3637]. sharing-based [2912]. Sharp [2606, 2093, 4164, 2887]. Shaw [419]. Shell [4313, 2907]. Shells [1063]. Sherman [120]. shield [3638]. shielding [2855]. Shift [655]. Shifted [2411]. Shileiko [100]. Shkil [641]. Shock [696]. Shooting [2173, 2211, 924, 900, 925, 2224, 1994, 1663, 924, 971]. Shooting-Verfahren [924, 971]. Shop [2357, 1985, 1294, 1837, 2414, 2391, 1688, 1979]. Shops [2198]. Short [2551, 2324, 2325, 3927, 2645, 3828, 3021, 3710, 2150]. short-characteristics [2645]. short-circuit [3021]. Shortest [1846, 1810, 2201, 2281, 278, 1331, 807, 4365, 3305, 2767, 598, 3479, 1620, 1681, 1412, 1487, 2751, 461]. Shorthand [4121]. shot [3728, 4029]. shrink [2871]. Shubert [1616]. Shubik [192]. Shulman [552]. sicheren [1024]. Side [1858, 878, 585, 3841, 3258, 3342, 1279]. Sided [894, 1137, 956, 4367, 1377, 693]. Sides [2294, 789]. siebenter [364]. Sieht [28, 38]. Sigla [1188]. sign [2710, 3074]. Signal [2637, 4004, 3544, 3386, 541]. Signals [2564]. signatures [2691, 2693]. signcryption [3702, 3806, 3156]. Signed [3771]. significance [3588]. Significant [1342, 1371]. Signorini [2214]. signs [1493, 3732]. Similar [1853, 1972, 2570]. similarities [3164, 703, 3323]. Similarity [3721, 3666, 3998, 3189, 3978, 3157, 3158]. similarity-based [3998]. Simple [1625, 159, 1851, 2669, 1818, 1523, 921, 1781, 2165, 1463, 1633, 609, 1344, 2647, 1753, 557, 1422, 1272, 1718, 1696]. simpler [527]. Simplex [1595, 22, 596, 1022]. Simplex-Method [1595]. Simplexmatrix [22]. Simplification [4140]. Simplified [2565, 2309, 1354]. Simplifying [2065]. simply [112]. simply-connected [112]. Simulafed [4031]. Simulated [2577, 2326, 1657, 4031, 3236]. Simulating [3262]. Simulation [2576, 1568, 80, 228, 502, 560, 802, 672, 3301, 609, 2858, 2864, 2866, 1693, 243, 2970, 3742, 1603, 3540, 2880, 4052, 2852, 609]. Simulations [2569, 2415, 2909, 3238, 2857, 2862, 3091, 3574]. Simulator [4127, 3302]. Simulators [4127]. simultanen [1949]. Simultaneous [1821, 4159, 1812, 2515, 1549, 1368, 1550, 1949, 2132, 4346, 4362, 1018, 4222, 1179]. Simultaneously [4350, 1766]. sine [3876]. Single [4288, 2278, 2414, 1331, 162, 2136, 1371, 871, 1018, 1075, 4222, 923, 1098, 1092, 2209, 3632, 2751, 1757, 3074]. Single-Machine [2278]. Single-Precision [1371]. single-sign-on [3074]. single-step [1075, 4222, 1098, 1757]. Singly [1768, 4365]. Singly-Implicit [1768]. Singular [2555, 2494, 1612, 2023, 2264, 2537, 2010, 1673, 2329, 2330, 2038, 1944, 2185, 2298, 241, 1850, 1527, 1775, 1818, 2485, 2524, 925, 2037, 2165, 368, 1909, 1908, 1270, 2293, 2350, 2785, 1472, 2727, 2680, 979, 1540, 4163, 1659, 1698, 1805, 1238, 470, 1422, 1966, 1764, 1697, 2786, 1696]. Singuläre [145]. singularen [1944]. singularer [2010, 1540, 470, 368]. Singularities [2382, 1419, 2309, 4364]. Singularity [1832, 1660, 2718]. singularity-free [2718]. Singularly [2055, 2525, 2338, 2402, 2478, 2339, 1536, 2746, 2595, 2775]. sink [3200, 3497, 3776]. Sintering [1942]. Sinterproblems [1942]. sinusoidal [3627]. SIoT [3838]. SIS [2801]. sites [3161, 3293]. Situation [3141, 3296, 3864]. situation-aware [3141, 3864]. situations [3673]. Six [639, 3177]. Size [1951, 1931, 2593, 316, 4168, 3289]. size/variable [4168]. Sized [2109]. sizes [1002]. SJL [3022]. Skalierung [456].

Skeleton [2305]. Skeletons [2325]. sketch [2717]. sketches [3417]. Skew [1858]. Skew-Symmetric [1858]. Skin [3728]. sky [2760]. Skyline [801, 3370]. Slabs [1943]. Slave [2235]. sleep [3223, 3251]. sleep-scheduling [3223]. SLI [1914]. SLI-Arithmetik [1914]. slicing [3239]. slicing-based [3239]. Sliding [2535, 4048]. slight [456]. slip [2732]. Slope [1251]. Slopes [1866]. slot [1509]. slow [3484]. slow-start [3484]. Slowly [2528]. Small [4120, 2367, 2759, 4047]. Smart [3931, 3513, 3137, 3638, 3688, 3447, 2953, 3730, 3912, 3657, 4015, 3117, 3473, 3974, 3636, 3811, 3336, 3136, 3071, 3564, 3489, 3147, 4065, 3972, 3476, 3424, 3604, 3805, 3483, 3212, 3396, 3472, 3501, 3517]. smart-relay-based [2953]. SmartBox [2934]. SMARTGREENS [3131]. smartphone [3577, 3020]. smartphones [3700]. smells [3918]. Smers [189]. Smirnow [40]. Smith [174, 2508]. Smolyak [2248]. Smooth [4313, 4208, 4304, 2507, 2131, 1534, 4261, 371, 1608, 1942, 2564, 2886, 2580, 2589]. Smoothed [2111]. Smoother [2319]. Smoothers [2320]. Smoothing [1437, 2485, 4308, 4347, 1902, 4219, 2477, 1380, 1054, 1056, 1682]. smuggler [3845]. snake [3945]. SNIFF [1086]. SOA [3244]. SOA-based [3244]. Sobolev [39]. Social [3360, 3470, 192, 3655, 3588, 3565, 3647, 3840, 3598, 3618, 3338, 3557, 185, 3876, 3767, 3161, 3646, 3830, 3752, 3473, 3502, 3894, 3585, 3098, 3601, 3631, 3996, 3313, 3586, 3849, 3366, 2904, 3692, 3424, 3436, 3892, 4013, 3709, 3700, 3989]. social-networks [3313]. Society [104, 3397]. soCloud [3096]. soft [3240, 3756, 3574]. soft-errors [3240]. soft-fairness [3756]. Software [2917, 3061, 1410, 3416, 3393, 952, 2971, 4124, 3478, 3240, 3264, 3490, 2776, 2918, 3431, 3018, 4094, 1748, 4093, 3178, 3421, 3697, 3572, 3495, 3163, 3021, 3569, 4084, 2668, 3454, 4035, 3303, 3508, 3493, 4167, 3078, 4119, 3071, 3146, 3185, 4136, 4168, 3201, 3011, 3784, 3692, 1733, 3211, 3268, 2923]. software-controlled [3490, 3493]. software-defined [3495]. Software-driven [3416]. soil [3759]. solar [3880, 3135]. Soleymani [2896]. Solid [4259, 3331, 2907, 2717]. solid-shell [2907]. Solids [4201]. Solodovnikow [259]. Solution [844, 2173, 454, 465, 4349, 109, 317, 277, 2532, 2083, 2509, 2497, 2555, 778, 2585, 276, 1631, 2433, 904, 1397, 824, 2570, 1281, 1390, 1051, 2180, 2440, 1610, 1254, 411, 1320, 2413, 1524, 2103, 838, 957, 2359, 1573, 1867, 1790, 1126, 1185, 1314, 2208, 1315, 2478, 1590, 1775, 1567, 900, 4343, 1174, 1030, 849, 2276, 2189, 2022, 1120, 1391, 1000, 2186, 1267, 925, 394, 4120, 4164, 1890, 1824, 860, 1036, 1195, 1435, 1504, 247, 877, 1908, 865, 825, 1270, 2056, 930, 2442, 1525, 517, 3180, 1472, 2642, 1066, 2664, 2851, 1324, 1667, 4224, 2700, 1632, 1431, 3901]. solution [396, 988, 2698, 1760, 2784, 1547, 4163, 937, 2631, 1108, 2397, 1585, 2663, 2868, 3939, 2652, 1749, 1700, 1214, 382, 1208, 399, 1709, 3439, 1359, 1983, 3589, 470, 1092, 1516, 1225, 497, 1162, 1266, 410, 1067, 1011, 1328, 4235, 18, 474, 1528, 2220, 414, 2661, 2846, 1426, 1125, 459, 2889, 495, 3509, 446, 984, 1424, 2807, 902, 855, 112, 866]. Solution-Dependent [1775]. Solutions [4110, 4179, 4333, 2154, 599, 1141, 2561, 1300, 2267, 2313, 313, 4337, 427, 2161, 2536, 2251, 1797, 2205, 4355, 1624, 1604, 1999, 1415, 912, 1635, 1290, 4173, 2511, 2053, 1425, 4361, 4345, 1926, 2051, 606, 707, 2396, 1958, 176, 1259, 4171, 1175, 1149, 2436, 2680, 4094, 989, 1540, 1735, 1689, 339, 1013, 1661, 4229, 1050, 1742, 1762, 2747, 2735, 2783, 450, 3605, 3663, 1276, 503, 1107, 1741, 2598, 1082, 1483, 91]. Solvability [1239, 1279]. Solvable [2066, 2345]. Solve [2588, 1375, 689]. solved [2818]. solvent [2790]. Solver [2275, 2057, 2394, 1948, 2116, 2477, 2613, 2645, 3437, 1718]. Solvers [2387, 2131, 1948, 1823, 2354, 2807, 2810, 2883, 2114, 2667, 974, 1280]. Solving [2233, 1546, 3870, 2575, 1910, 875, 2567, 1808, 1994, 2120, 1047, 1450, 1551, 2264, 1767, 1835, 2481, 1499, 787, 1968, 4172, 2820, 1917, 2306, 241, 919, 4359, 1832, 1200, 1432, 2064, 2400, 2560, 1605, 1787, 1887, 933, 4233, 2371, 2084, 1774, 2042, 1342, 3639, 2214, 2239, 2315, 1737, 1969, 2474, 2013, 1800, 2070, 1250, 2782, 1745, 2670, 2778, 1323, 704, 1721, 1482, 1117, 1556, 1340, 1423, 3751, 2806, 584, 1061, 1074, 1743, 2176, 1698, 1428, 965, 1579, 2781, 2894, 1693, 2668, 1651, 2721, 1476, 2844, 3050, 1603, 2771, 500, 1982, 413, 2895, 1755, 1096, 1156, 2726]. solving [1083, 2779, 3777, 1611]. Some [1251, 2497, 2335, 1786, 1064, 315, 427, 2031, 1998, 1930, 1488, 210, 1869, 1491, 79, 21, 1285, 1196, 1499, 957, 787, 1277, 2380, 2515, 1548, 944, 4114, 1904, 4154, 1949, 719, 1774, 2886, 1504, 2054, 892, 1772, 1601, 1544, 620, 2150, 2018, 4223, 1347, 2720, 4226, 2382, 1542, 1582, 964, 1209, 2652, 1071, 4160, 436, 1702, 1764, 1392, 495, 1235, 54, 1939, 1269, 879, 782, 1884]. Sommerfeld [260]. Sonar [4126, 4126]. Sonar-Feldverarbeitung [4126]. SONET [2725]. SOR [1260, 1564, 784, 4358, 1260, 2016]. SOR-like [4358]. SOR-methods [1260]. SOR-Verfahren [1260]. sort [1136]. Sortieren [1023]. Sorting [1441, 1386, 1088, 2828, 1231, 2665, 1408, 1023]. Source [2520, 1331, 2987, 3471, 3943, 4081, 3169, 2751, 3634]. Source-Replies [3634]. Sources [2524, 1376, 1683, 3382, 3450]. sowie [78, 238, 418]. Sozialwissenschaften [185]. Soziologie [482]. Space [2269, 836, 1128, 1810, 1947, 2011, 2549, 2538, 2451, 2453, 4112, 2395, 895, 2887, 783, 2113, 3651, 2636, 3743, 1822, 2630, 3177, 1711, 2888]. space-ground-integrated [3651]. Space-time [2887, 2630]. Spaces [4297, 468, 469, 2566, 526, 724, 4105, 4113, 318, 711, 4123, 1502, 440, 985, 3547, 409, 433, 632, 2006, 613, 1754, 1485, 1566, 3864, 934, 1847, 3116, 72]. Spacetime [1420]. Spaltenänderung [244]. Spanning [2201, 848, 1552, 963]. Spannungs [1941]. Spannungs-induzierter [1941]. Spark [3549, 3774, 3556]. Sparse [2426, 1255, 2628, 2430, 1390, 2046, 635, 735, 2112, 2057, 2364, 2427, 1961, 2226, 2245, 2287, 2330, 2321, 1432, 1605, 1373, 2371, 4345, 2126, 2562, 2428, 308, 718, 2584, 3384, 4079, 1467, 717, 1556, 847, 2699, 2673, 2611, 2630, 2390, 1620, 1200, 1651, 2721, 3731, 2945, 630, 1096, 3291]. Sparse-Grid [2330]. Sparse-Matrix [2371]. Sparsity [3130, 4005]. Späth [833]. Spatial [4262, 3418, 3384, 3666, 3905]. Spatial-feature [3418]. spatial-temporal [3384, 3666, 3905]. spatio [2963]. spatio-temporal [2963]. spatiotemporal [3418, 3922]. SPBench [3749]. Special [3256, 3162, 1864, 3481, 208, 3042, 3616, 3170, 3320, 3464, 3490, 3024, 2120, 3034, 4351, 954, 1109, 3052, 1116, 3229, 1818, 3717, 1373, 3153, 3105, 927, 3605, 2559, 3779, 2160, 3899, 636, 1008, 3373, 2928, 3294, 2013, 3261, 4382, 3674, 1745, 3866, 185, 3315, 3397, 3410, 3018, 3087, 988, 1165, 3118, 2968, 2911, 1579, 2983, 1017, 1071, 384, 1220, 3003, 3209, 2796, 689, 3381, 2509, 2601, 2706, 3718]. specialization [3434]. Specific [4217, 3684, 3290]. Specification [3432, 1410, 2816, 3044, 3781, 4136, 2923]. Specified [4257, 2747]. specifying [3755, 3754]. spectra [3843]. Spectral [1794, 2352, 4336, 4337, 2046, 249, 4339, 2312, 2580, 586, 638, 3249, 984]. Spectrum [2590, 682]. Speech [732, 3369, 4133, 2802, 3377, 3620]. Speed [311, 739, 953, 3744, 307, 583, 384, 3273]. Speed-up [311, 384]. Speed-Ups [953]. speeding [4034]. Speicher [1023]. Speicherung [2039]. speieller [689]. Spektraldarstellungen [984]. Spektralradius [586, 638]. Spektrum [682]. Sperner [596]. spezieller [208, 569, 368, 1008, 2013]. Sphere [2134, 1437, 2574, 1608, 1210, 1566]. Spheres [2375]. Spherical [1440, 2459, 199]. spiders [2623]. Spiele [347]. Spiess [515]. spiking [3238]. spillover [3237]. Spine [2460, 1003]. Splat [2602]. Spline [1003, 1518, 2545, 2544, 1771, 4248, 2449, 2450, 4254, 1395, 4306, 722, 1548, 785, 2406, 383, 833, 4311, 1025, 4246, 1202, 4249, 2006, 2791, 1354, 1220, 1225, 1636, 1806, 2718, 1975]. Spline-Algorithmen [833]. Spline-Interpolation [785]. Splinefunktionen [722, 1225]. Splines [959, 1456, 2228, 1962, 4203, 872, 936, 1596, 1437, 1623, 4319, 1353, 1559, 2451, 4211, 908, 2035, 2225, 1933, 4260, 1614, 1799, 1902, 2328, 574, 522, 523, 1608, 2104, 4310, 1547, 2708, 1060, 1350, 1054, 2021, 1053, 1873, 1682, 1976, 1736, 959, 1962]. Split [4321, 3362]. Splitting [1925, 2049, 1696, 1589, 1138]. SPP [3991]. Sprache [1265, 568, 569, 30, 31, 141, 686]. Sprachen [17, 698, 342, 312]. Sprachkunde [29, 30, 37, 31, 33, 32, 38, 43, 158]. sprachliche [30]. Sprachübersätzung [37]. spread [3614, 3313]. spreaders [3338, 3617, 3312, 3871]. Springer [16, 3361, 2911]. Spurious [1300]. SQL [3067, 3357]. Square [1106, 1128, 1668, 370, 1844, 1862, 721, 1222, 1676, 1069]. Squares [844, 4252, 2011, 2297, 1835, 2249, 2369, 1787, 1029, 1036, 2133, 2302, 209, 2389, 2170, 2707, 3802, 1707, 1547, 1651, 1476, 2715, 1986, 1737, 973, 2764]. squaring [1765]. SSD [3968]. ST [3799]. ST-Resnet [3799]. St. [646]. stabile [1084, 682, 1332]. stabilisation [2733]. Stabilised [2336]. Stabilität [915, 916, 981, 1945, 1526, 115, 690, 706]. Stabilitätsatzes [456]. Stabilitätsproblem [1043]. Stabilitätstheorie [403]. Stability [875, 774, 2436, 2102, 4178, 1830, 2158, 1834, 2297, 1199, 1043, 4253, 2119, 2487, 2735, 309, 2324, 2312, 562, 2087, 981, 633, 1945, 310, 4278, 1880, 3588, 1044, 456, 1684, 982, 1981, 1822, 923, 2808, 1747, 4234, 2823, 1708, 1526, 999, 4001, 1430, 1201, 690, 706, 915, 916]. Stabilization [2337, 2034]. Stabilized [2506, 2480, 2795, 2656, 2704, 2732, 3352]. Stabilizing [2243, 2678, 2207, 2378, 3701]. Stable [1066, 1724, 761, 682, 2424, 2594, 1084, 1077, 1332, 2358, 1002, 1407]. Stabley [290]. STABULUS [1724]. stacked [3737]. Staff [204]. stage [2785, 4225, 3439, 2802, 3077, 3820]. Stahlknecht [151]. stakeholder [3988]. Stammfunktionen [578]. Stamps [1587]. Stand [435]. Stand-by [435]. Standard [4183, 460, 4184, 441, 2814, 942]. Standardfunktionen [460, 441]. Standardization [622]. Stapelverarbeitung [588]. Star [2770, 2316]. stark [851]. Starke [673, 746]. Starobin [576]. Start [2278, 3961, 4007, 3484]. Starting [2048, 1844, 1700]. starts [3412]. starvation [3675]. starvation-freedom [3675]. State [374, 5, 900, 3331, 2589, 3026, 3802, 3107, 3263, 617, 987, 2888, 3155, 4137, 3874, 3383]. State-constrained [2589]. state-machine [3263]. state-of-the-art [3026]. state-space [2888]. statements [297, 78]. states [3731]. static [2930, 3266, 2881]. Stationären [840, 1]. Stationary [2372, 1, 2795, 2870, 700]. Statistical [147, 137, 494, 101, 94, 538, 2106, 209, 1233, 1, 3712, 3394, 243, 3376]. Statistics [1639, 251, 185, 238, 1980, 1780, 3799, 90]. Statistik [185, 238, 576, 15, 757]. Statistische [181, 326]. statistischen [1, 173]. statistischer [209]. status [4020, 2759]. Steady [3802, 1941, 2506, 2869]. Steady-state [3802]. stealing [2951]. steam [2864]. steel [3350, 3738]. Steep [2173]. steering [987]. Stefan [1824, 2878]. Steffensen [887]. Steffensen-Verfahrens [887]. Stegun [134, 69]. steifer [1339]. Steigungsbegriffes [1234]. Steinberg [491]. Steinbuch [236]. Steiner [2333, 3879]. stencil [3049]. Step [2388, 1668, 1776, 2693, 790, 869, 2432, 2049, 1951, 449, 1168, 1138, 1018, 1075, 4222, 1684, 1098, 1849, 1700, 1398, 1328, 4168, 3660, 1002, 1125, 1452, 1757, 3827, 3498]. Step-Cycle [790]. Step-Cycles [869]. Step-out [2693]. step-sizes [1002]. Steplength [1166]. Steps [955, 2078, 1770, 1380, 1659, 1752]. Stepsize [1796, 467, 888, 1491, 1027, 1388, 364, 1349]. Stereo [4268, 3287]. Stetige [1967]. stets [887, 823]. Stetter [798, 1793, 1534]. Steuerung [952]. Steuerungen [534]. Stichprobenparameter [820]. Stiefel [389, 481, 404]. Stiff [1769, 1804, 1795, 2353, 750, 2276, 2432, 1534, 2293, 2350, 1458, 1761, 1615, 1748, 4157, 1263, 2032, 1474, 1398, 1607, 1339, 1160]. Stiffness [1796]. Stimulated [904]. STL [4001]. Stochasstische [262]. Stochastic [2180, 2569, 71, 261, 536, 2264, 4015, 480, 2568, 5, 333, 435, 2428, 3690, 524, 1843, 640, 1376, 1683, 3324, 590, 949, 2862, 3959, 1344, 243, 3341, 715, 834, 3843, 3991, 2823, 697]. Stochastische [671, 743, 284]. stochastischen [590, 949, 650, 697]. stochastischer [1376, 1683]. stock [3405, 3411]. Stoer [674]. Stoffers [298]. Stokes [2115, 2506, 2795, 2301, 2344, 2656, 2704, 2732, 2354, 2589]. Stopping [2577, 1075, 1076]. Storage [1232, 2371, 1385, 3086, 3648, 3192, 3130, 2969, 3948, 3048, 2916, 3222, 1322, 3076, 3308, 3922]. storage-centric [3086]. store [3930]. Storey [831]. Storing [1587, 3123]. storm [3603]. Störungen [1945, 1]. Störungsrechnung [300]. Strangschale [1943]. Strategies [3400, 2578, 1450, 621, 2256, 2049, 3492, 2893, 3954, 4022, 3923, 4090, 4037, 3000, 3928, 3890, 3411, 2967, 3152]. Strategy [2095, 1832, 3602, 3744, 3202, 2767, 1615, 3165, 776, 3444, 3525, 2916, 4051, 3254, 3898, 3868, 4008, 3768, 3978, 3498, 3237]. Strategy-proof [3602]. Strationovich [378]. Stray [476]. stream [4017, 3749, 3529, 2694, 3793, 3997, 3747]. Streaming [2489, 3417, 3603, 3650, 2836, 3450, 3277]. Streamline [2349]. Streamline-Diffusion [2349]. streams [3419, 3297, 3922]. Strecher [817]. Streckenkomplexe [234]. street [4015, 3972]. streng [783]. Strengen [1042, 1147]. Stress [1941, 3350, 3738]. Stress-Induced [1941]. Strict [1042, 1571, 1147, 1672]. Strictly [1178, 783, 1060, 1236, 1726]. String [2061, 3870, 2218, 987]. Strings [2061, 1974]. Strip [2168, 2328]. strips [1541]. Stromungen [1941]. Strong [2433, 2207, 1239, 1708, 4285, 3102]. Strongly [210, 2023, 941, 1527, 1380, 1103]. Structural [183, 3185, 3073]. Structure [1723, 1183, 2464, 4127, 4286, 4102, 3532, 436, 1373, 4213, 2197, 681, 1761, 17, 440, 3721, 3161, 3646, 507, 2866, 3930, 2751, 3815, 3703, 3871, 3757, 4054, 3291, 3158, 594, 612, 895]. Structure-guided [3532]. Structured [792, 3392, 2842, 2613, 4005, 2840, 1345]. Structures [1804, 1133, 1825, 786, 4250, 2097, 1420, 1594, 862, 598, 3067]. Struktur [17, 440, 594, 612, 895]. strukturalistischer [38]. Strukturen [436, 862]. Strukturnumwandlungen [234]. Stuart [430, 193]. stückweise [637]. Studenplanerstellung [397]. student [3524, 1229]. Studer [1136]. Studien [658]. Studienbücher [580]. studies [3755, 3754, 655]. Study [420, 1634, 4186, 2548, 1773, 1946, 1960, 2452, 3375, 333, 4083, 1006, 2901, 3911, 3626, 440, 1532, 2939, 3062, 3086, 3670, 4094, 3880, 1981, 4093, 3541, 3698, 2869, 4081, 3503, 3847, 3266, 4052, 3833, 3314, 3794, 2737, 1677, 3327]. Studying [3336]. Sturm [1740, 2053, 1244, 340, 1895]. Stützstellen [837]. style [3783]. Sub [4279, 3764]. Sub-Hypergraph [4279]. subareas [3679]. Subcolorings [2398]. Subdivision [2447, 2452, 2455, 4309, 2712, 2713, 2604, 2745, 2716]. Subdomain [1823]. subgradient [1650]. Subgraph [4283, 1592, 2739]. subgraph-shaped [2739]. subgraphs [298, 3705, 3065]. subgraphs-density [3705]. subgrid [2720]. Subgroup [806]. Subject [4155, 1, 950, 1650]. Submatrices [2025, 1627]. submatrix [2747]. subroutine [1323]. Subsemigroups [763]. Subsets [687, 1674, 1262, 1215]. Subspace [2399, 2274, 2354]. Subspaces [1598, 1558]. substitution [2744]. Substructuring [2358]. Subsurface [2302]. subtitling [3764]. subtree [1048]. success [3546]. Successive [1477, 1478, 1604, 1661, 152, 1770, 967]. Suche [741, 737]. Suchowitzki [676]. Sufficient [2263, 2193]. suffix [3935]. Suggestions [980]. suicidal [3618]. suitability [614]. suite [3064]. Sum [4312, 1959, 554, 3297, 2600]. Sum-CRCW [2600]. Summanden [805]. summarization [3069]. Summation [2586, 4098, 954, 2574, 2546, 4143, 1306, 989]. Summen [805]. summer [3138]. Sums [206, 805, 2528]. Sup [2395, 2534]. superblock [2973]. Supercomputer [1899]. Superconvergence [1099, 2309, 2437, 2697]. Supergraph [2308]. Superlinear [1300, 1348, 1164]. Superlinearly [4338, 1874]. Supervised [3488, 3885, 3860, 3762]. supplier [3551]. supply [3915, 3539]. support [3199, 2940, 3461, 3785, 3495, 4048, 3275, 3960, 3151, 4014]. Supported [2134, 2651, 1990]. supporting [2898, 3533, 3196, 2912, 2831, 3455, 2792, 3436]. Surface [2375, 4241, 4244, 2010, 4302, 4321, 4322, 2451, 2453, 4269, 4324, 1519, 2213, 4328, 4220, 3554, 2135, 1630]. Surfaces [2518, 1434, 2446, 4313, 4200, 958, 4245, 4246, 4204, 4247, 4314, 4301, 4206, 2609, 4208, 4253, 4303, 4209, 1419, 2452, 4306, 1944, 4323, 2520, 4325, 4257, 2050, 4219, 2458, 4261, 2460, 4330, 4263, 2442, 2602, 1025, 2605, 2702, 2790, 1016, 2713, 2604, 1001, 2711, 2718, 1026, 1055, 2603]. surveillance [2931, 3020, 3073]. Survey [4335, 3043, 3565, 3469, 3987, 3432, 2891, 3326, 3661, 3961, 3104, 2999, 3505, 3166, 2758, 3668, 3544, 3710, 3644, 1692, 2760, 2975, 3380, 3206, 3112]. survivability [3923]. Survival [2800]. Susan [654]. susceptibility [3288]. susceptibility-weighted [3288]. sustainability [2939]. Sustainable [2950, 3162, 3165, 4064]. SVD [2512]. SVFLDetector [3882]. SVLM [3445]. SVM [3395]. SW [2926]. Swarm [3953, 2927, 3193, 3094, 3541, 3334, 3669, 3205, 3389, 3531, 3236, 3425, 3486, 4002, 3777]. swarm-based [3334, 3389]. sweep [1313, 1494]. sweep-plane [1313, 1494]. Sweschnikow [336, 576]. Switching [421, 422, 2211, 1796, 928, 435]. Switching-Method [2211]. Sylvester [2015]. symbiosis [3751]. symbol [1031, 614]. Symbolic [2045, 4370, 4371, 1521, 1944]. Symbolmanipulation [1944, 614]. symbols [1]. Symm [1457]. Symm-Wilkinson [1457]. Symmetric [4349, 1255, 1569, 958, 1409, 1831, 1954, 735, 2423, 1109, 914, 4268, 1828, 1858, 4343, 7, 449, 1297, 867, 244, 1018, 2731, 2141, 717, 2806, 1676, 1402, 2793, 1454, 2859, 2681]. symmetric-triangular [2793]. Symmetriegruppen [1931]. Symmetries [1480, 2396, 1716, 1553]. symmetrisch [914]. Symmetrische [1018, 1954]. symmetrischen [1402, 867, 244]. Symmetrischer [1409, 958, 7, 404, 244]. Symmetry [1872, 1931, 2239, 2315, 2584, 3437, 2139]. Symplectic [1888, 2125]. Symposium [830]. symptoms [3638]. Synchro [3764]. Synchro-Sub [3764]. synchronisation [3764]. Synchronisationsoperationen [1170]. Synchronization [1170, 3207, 3746]. synchronization-free [3207]. synchronized [3499]. synchronous [3753, 3022]. synergy [3566]. Syntactic [17]. Syntaktische [17]. Syntax [111]. Synthesis [476, 2521, 183, 655, 3976, 3016, 3391, 2771]. Synthesizing [3088]. System [542, 544, 1145, 289, 1592, 1730, 337, 4353, 2057, 1188, 619, 565, 4127, 588, 1867, 516, 2453, 103, 290, 4255, 2478, 1818, 1887, 1030, 3039, 1425, 540, 1523, 435, 501, 2562, 274, 341, 2272, 524, 204, 2212, 478, 1416, 952, 1356, 3183, 3730, 3641, 2826, 3474, 3440, 3387, 3299, 3734, 2810, 3199, 4079, 2934, 3533, 3581, 3817, 3545, 3702, 3654, 480, 2862, 3591, 2991, 3727, 3119, 2969, 3990, 2961, 3555, 2892, 3636, 3122, 3597, 1017, 3756, 2595, 3225, 2877, 3761, 3458, 1041, 1204, 3967, 3260, 4133, 3736, 3382, 3336, 3146, 4134, 2744, 3393, 4065, 3724]. system [2904, 3142, 3620, 4138, 3076, 3233, 4001, 2956, 3609, 2764, 3047, 1145, 3503]. system-based [3761]. system-level [3817]. System/360 [490, 290, 204]. Systemanalyse [647]. Systematic [2352, 771, 4083, 3148, 3670, 4094, 3789, 3203, 3303, 3663, 3833, 3700, 3792, 3501]. Systematische [114]. Systeme [619, 501, 4135, 450, 2013, 831]. Systemen [1254, 338, 707, 984, 365, 1024, 452]. Systemparametern [1945]. Systemprogrammierung [812]. Systems [2341, 335, 542, 465, 4174, 2532, 1804, 778, 2366, 1409, 1910, 875, 731, 2120, 1506, 1586, 1254, 1146, 2166, 1797, 2411, 2464, 4100, 2475, 2573, 1499, 4128, 2281, 2351, 1968, 650, 1858, 648, 649, 1917, 1879, 2131, 1853, 2000, 765, 627, 700, 2503, 1635, 1432, 2064, 2400, 1775, 1567, 1605, 4343, 854, 886, 768, 750, 1373, 4116, 2172, 1006, 2157, 2189, 2371, 990, 628, 851, 2516, 1774, 582, 651, 1959, 2042, 4120, 1342, 1890, 2126, 7, 3779, 1036, 4135, 707, 117, 1072, 1606, 1898, 2273, 2149, 1882, 130, 161, 789, 672, 2343, 808, 2013, 4152, 1250]. Systems [1024, 3043, 1018, 1110, 3780, 1472, 2769, 3162, 1761, 2731, 3682, 3611, 3042, 1066, 2825, 3220, 2778, 521, 3912, 3301, 1509, 3134, 3490, 2919, 2763, 3675, 2918, 3174, 3018, 3714, 365, 3223, 3034, 3059, 3009, 1482, 3012, 2700, 1540, 1340, 3347, 3891, 3648, 3901, 4015, 3783, 1852, 1699, 2806, 585, 584, 3538, 3104, 3789, 4087, 4045, 3590, 4026, 3579, 3753, 2652, 2925, 3344, 1050, 3218, 4231, 3364, 2221, 3001, 3181, 4028, 3903, 3993, 2668, 1321, 3896, 2001, 1754, 3542, 3193, 3740, 3739, 3742, 384, 4038, 3307, 473, 1516, 2721, 450, 744, 3303, 3584, 3775, 2957, 3867, 3534, 3153, 2835, 3257]. systems [3022, 3265, 554, 4130, 3166, 452, 3567, 3044, 3325, 3247, 3151, 3493, 2929, 3683, 3663, 3003, 3336, 3182, 3975, 2758, 1276, 2837, 3195, 2824, 2823, 1566, 2882, 1706, 4235, 3216, 3185, 1454, 1528, 414, 340, 2661, 1664, 1977, 2901, 4036, 4238, 976, 3135, 3607, 3438, 1757, 3155, 2840, 1461, 3491, 1083, 3874, 3084, 3798, 984, 2928, 3482, 2971, 1356, 2994, 2974, 2981, 2807, 4089, 3115, 732, 188, 116, 266, 247, 183, 550, 828]. systems/paradigm [3336]. Systolic [2352, 1814, 1692, 1642, 1487]. Szabö [287, 362, 488, 656]. Sze [422]. Sze-Tsen [422]. Szyperski [643].

T [539, 48, 71, 486, 153, 402]. T.

[45, 377, 299, 253, 332, 177]. Tabellen [26, 156]. Table [2062, 1232, 604, 3041, 3930, 3327]. Tables [220, 386, 2188, 88, 4141, 212, 199, 258, 69, 147, 101, 604]. Tabu [1724, 1653, 2927]. Tabulator [668, 564, 668]. Tabulator-Turing [564, 668]. Tabulator-Turingmaschine [668, 564]. tackling [2843]. tactics [4080]. Tafeln [26].

Tafelwerke [156]. Tagesspeicherkraftwerkes [1938]. tagging [3329, 3426]. Taguchi [3776]. Taguchi-based-GWO [3776]. Tagung [105]. Tagungsbericht [59]. Takács [261, 262]. tamper [3414]. tandem [2803, 1041]. Tangent [2450, 2453]. Tangential [766, 2231, 2453]. tangible [2935]. tangierenden [766]. tape [502, 560, 412]. Tapeless [681]. Tardiness [1882]. target [3861]. Tarski [263]. Task [2187, 3760, 3768, 4067, 3188, 3199, 3220, 3940, 3945, 3706, 3765, 3684, 3504, 3035, 3896, 3522, 4041, 4074, 3862, 3950, 3803, 3633, 3519, 4065, 3077, 3025, 3839, 4089]. task-scheduling [3199]. task-specific [3684]. Tasks [1881, 3625, 3183, 3101, 4027, 3000, 3359, 3439, 3307, 3842, 3932, 2846, 3448, 3367, 4002, 2956]. Tau [2216, 1391, 1266, 1728, 1709]. Taub [623]. Taut [1618]. taxonomies [3615]. taxonomy [3326, 3104, 3733, 3360, 3169, 3206]. Taylor [1646]. Tchebycheff [500]. TCP [3273]. TDMA [1509]. TDMA-systems [1509]. teachers [238]. teaching [3687, 3525]. teaching-learning-based [3687]. Tearing [2435]. Technical [551, 77, 3544]. Technik [105, 646, 1994, 961, 300, 548]. Technique [2146, 276, 2322, 1994, 1724, 2329, 2330, 2033, 2321, 2560, 2170, 3834, 3721, 1556, 989, 2673, 3570, 3348, 1474, 3810, 3811, 3938, 3818, 3307, 2775, 47, 669, 961, 3582, 3250, 3414, 4021, 3551, 2954, 3268]. technique-based [3834]. Techniques [375, 2545, 1939, 1499, 858, 1653, 2123, 2291, 1432, 4083, 4292, 2499, 2214, 150, 373, 1969, 2113, 3043, 3392, 3170, 2883, 3008, 3714, 3726, 4161, 3104, 3344, 3595, 1017, 2817, 2625, 3336, 3195, 3146, 3794, 3142, 2754, 3235, 3482, 731, 153, 228]. Technische [330, 576]. technisehen [356]. technologies [3306, 2943, 3431, 3018, 4094, 3167, 3733, 3153, 3736, 2979, 3071]. Technology [419, 532, 3083, 3833, 3952]. TEDA [3851]. Teichroew [126]. Teil [185, 409, 433, 151, 294, 306, 362, 488, 915, 916, 37, 201]. Teilgraphen [298]. Teillistensperrung [1032]. Teilnehmer [531, 84]. Teilnehmer-Rechenanlagen [84]. Teilnehmer-Rechensysteme [531]. teilweise [837]. telecom [3762]. telecommunication [3235]. Telecommunications [2157, 2273]. telehealth [3079]. Teletraffic [830]. template [2814]. templates [3141, 3072, 2814]. Temple [369]. Templescher [369]. temporal [2963, 3384, 3682, 3230, 3666, 3520, 2946, 3328, 3385, 3905, 3777, 3989]. temporal-difference [3777]. Ten [220]. tenancy [3957]. Tension [4211, 2456, 1054]. Tensor [2348, 1437, 2496, 2429, 2464, 4306, 4210, 4216, 2714, 2695, 2890]. Tensor-Product [2496, 4306]. ter [78, 988, 877, 1125]. Term [2440, 893, 572, 1021, 3927, 3996, 1685, 2183, 3420]. Termination [2409]. Terms [2012, 4143, 4142, 805, 2528]. Ternary [2346]. Tesselations [4214]. Test [2045, 111, 2031, 4178, 2162, 2175, 3232, 1539, 3239, 2686, 1627, 1213, 1664, 1733, 3934, 822, 899]. testbed [2936]. Testing [3301, 3167, 2530, 2917, 2737]. Testmatrizen [822, 899]. Tests [2236, 2469, 1387, 2654, 950, 2597, 592, 1978]. Tetrahedral [2439, 2098, 2628]. Teubner [580]. text [3549, 3375, 3378]. texts [3979]. textual [4035]. Texture [4286]. Th [323, 745, 78, 2145, 988, 1443, 877, 1125, 791, 930]. Th. [814, 649, 228, 223]. Thaler [266]. Theia [3128]. Their [419, 1826, 268, 1355, 467, 1491, 79, 1849, 378, 1501, 1658, 333, 4141, 2907, 2341, 1044, 143, 2727, 2436, 982, 1210, 2531, 3520, 2454, 4129, 1627, 623, 2130, 1095, 3967, 2830, 2794, 751, 614, 718, 1741, 1601]. them [3177]. Theodore [112]. Theodorsen [112].

Theorem [2592, 771, 485, 2309, 2521, 868, 456, 985, 607, 1286, 1507, 597, 1759, 962]. Theorems [4334, 762, 595, 1555]. Theoretic [178, 2257, 6, 3494]. Theoretical [4268, 567, 4277, 964, 342, 1073, 4378]. Theoretically [2651]. Theoretiche [424]. theoretische [260]. Theorie [4096, 271, 429, 264, 49, 336, 1269, 4104, 744, 344, 187, 970, 615, 789, 326, 581, 202]. Theory [2386, 335, 421, 422, 4096, 143, 184, 107, 416, 4372, 1323, 1477, 285, 289, 1901, 148, 137, 149, 261, 2261, 1928, 49, 659, 4128, 770, 641, 378, 4104, 4190, 66, 540, 2579, 192, 344, 47, 227, 4275, 209, 789, 620, 3821, 3641, 4376, 517, 6, 3170, 113, 1269, 3098, 597, 3929, 970, 4374, 615, 3235, 495, 1193, 3261, 225, 299, 2022, 142]. theory-based [3641]. Theory-Controlled [2261]. thermal [2914, 3135]. thermoelastic [2872]. Thermoelasticity [1790]. Thermoluminescence [2848]. Thermoviscoelasticity [1666]. Things [3298, 3294, 3976, 4085, 3698, 4026, 3993, 4042, 3424, 3449, 3821, 3641, 4006, 3413, 3481, 3410, 3901, 3060, 3275, 3656, 3372, 3229, 3511, 3311, 4086, 3137, 3527, 3851, 3295, 3117]. things-edge [3993]. Thinning [4266, 3310]. Third [1794, 88, 2315, 2396]. Third-order [2396]. Thomas [2366]. those [4138]. Thrashing [1006]. THRD [3030]. thread [3748, 2973]. thread-level-parallelism [2973]. threaded [3000]. threads [3745]. Three [2138, 2228, 1598, 1045, 2314, 2122, 2291, 2480, 524, 4278, 2743, 2909, 2649, 1743, 2176, 965, 1038, 1272]. Three-Dimensional [2314, 2291, 1045, 2743, 2909]. Three-field [2480]. Three-Partitioning [2138]. Threshold [3030, 3100, 3725]. Threshold-based [3030, 3100]. thresholding [2995]. throughput [3205, 3957]. Thue [365, 365]. tier [3948, 4003, 3640]. Tight [2559]. Tightening [2048, 2806]. Tighter [2357]. Tikhonov [42, 1665, 1629, 1982].

Tikhonov-regularization [1629]. tiled [3207]. Time [2388, 1294, 2278, 1509, 3712, 1851, 2376, 249, 2155, 2292, 2105, 386, 1500, 2338, 1042, 1938, 700, 2520, 2443, 3889, 3257, 1587, 851, 2240, 3241, 437, 4135, 1898, 1882, 2151, 1460, 2074, 2479, 1147, 2068, 3306, 1644, 3682, 3611, 1231, 1264, 3984, 1822, 4017, 1199, 3359, 2611, 2630, 3974, 1123, 3897, 4048, 3390, 3658, 2746, 3194, 1321, 3844, 2001, 3215, 3764, 1204, 3967, 2899, 3842, 2682, 3460, 3382, 2823, 3443, 3942, 2887, 3332, 1048, 3367, 2741, 2742, 3874, 3381, 4064, 1082, 3957, 2764, 2779, 3420, 3302]. Time-Accurate [2338]. time-aware [3897, 3390]. Time-Conditions [1147]. Time-Dependent [2376, 2682, 2741, 2742, 2764]. Time-efficient [3257]. time-homogeneous [1199]. Time-Lags [2278]. Time-minimal [2479]. Time-Parallel [2074]. Time-sensitive [3889]. time-series [3382]. time-sharing [1321]. Time-slot [1509]. time-triggered [3611]. time-varying [3957, 2779]. Times [2380, 1853, 2476, 2443, 1935, 1216]. TimesLap [3995]. Timetable [399]. Timetable-Problem [399]. timetables [397, 400]. Timman [331]. Titan [1899]. Toeplitz [1948, 1948]. together [3177]. tolerance [1748]. Tolerant [1897, 3611, 3040, 3263, 3571, 3356]. tolerating [3040]. Tomography [2225]. Tool [2216, 3009, 2768, 3128, 3778]. Toolchain [4084]. Tools [4250, 3392, 3026, 3148, 2891, 3266, 3071]. top [3208]. top-k [3208]. topic [3228, 3375]. Topics [4382, 815]. Topological [4096, 472, 1171, 2888]. Topologies [2271, 3432]. topologischen [4096]. topology [3890, 3841, 3362, 2749]. TOPSIS [3969, 3633]. torsion [2710]. torus [3317, 2135]. TOSCA [3432]. Total [207, 2249, 2593, 863, 2696]. totalen [207]. Totally [2423, 1834, 2549, 863]. touch [2935]. Tourenproblemen [649]. Tower [1732]. Towers [1570]. Trace [2045, 1521, 3127, 3394, 4030]. traceroute [2949]. Tracing [2487, 4167]. Track [2008]. tracked [3622]. Tracking [2258, 1312, 2909, 2616, 3736, 3638, 3420]. Tracts [419, 16]. trade [4059, 3637]. trade-off [3637]. trade-offs [4059]. Tradeoffs [4268]. trading [3405, 3411, 3483]. traditional [4020]. traditioneller [38]. Traffic [696, 1113, 3985, 3474, 3664, 3452, 4046, 3657, 3223, 559, 2944, 3974, 3924, 3927, 3630, 3636, 3274, 2948, 3875, 3484, 3078, 80, 3864, 474, 4089, 4128]. traffic-aware [3223, 4089]. tragen [2039]. train [3609]. train-centric [3609]. trained [4080]. Training [2490, 3405, 3741, 3750, 3859, 2990, 3960, 3937, 3898, 4034, 3685]. trait [3098]. Trajectories [2258, 3577]. trajectory [3196, 3813, 4088, 3922]. trans [3621]. trans-cloud [3621]. Transaction [2028, 3954]. Transaction-like [2028]. transactional [3041, 2977, 3675, 3056]. transactions [2932]. Transcendental [21, 4149]. Transfer [2776, 618, 2164, 1966, 2056, 2311, 2769, 2645, 3443, 3934, 4029, 342]. Transferoperatoren [1966]. transferring [3858]. Transfinite [4205]. Transform [2431, 2177, 4266, 1427, 227, 2611, 584, 1108, 2892, 314, 3287, 2001, 2995, 2063, 1052, 3865, 1911, 280]. Transformation [2439, 2508, 2485, 265, 478, 664, 669, 1462, 1114, 1462]. Transformationendarstellungen [1114]. Transformations [4283, 2181, 1637, 1481, 3177]. Transformed [2317]. transformer [4072]. transformer-based [4072]. transformers [4070]. Transforms [1346, 719, 998, 132, 2987]. Transient [3354, 2701]. Transients [257]. Transistoren [267]. transition [715]. transitional [2858]. Transitive [897, 892, 3852, 1383]. Translation [1177, 858, 806, 869, 3374, 3376, 881]. Transmission [2519, 2524, 3107, 2933, 4003, 2881, 3679, 1713]. transonic [2864]. transparency [2842]. transparent [3747]. Transport [107, 904, 2378, 2371, 22, 2850, 3, 2910, 3002, 250]. Transport-Betriebswirtschaftslehre [250]. transportation [1117, 3014, 965, 3939, 3482]. Transportenetze [347]. transportproblem [3]. Transportproblemen [22]. Transportproblems [904, 965]. Transputer [2043]. Transversals [1846]. transzendenter [21]. Trapezoidal [1648]. Traveling [2066, 2887]. Travelling [461, 278]. traversal [3810]. Treated [1045]. Treatise [172]. Treatment [1943, 1046, 1043, 2355, 1501, 994, 1192, 2539, 789, 1081, 1089, 559, 972, 975, 986, 1738, 973, 1226, 1339]. Tree [1033, 1329, 1105, 4187, 1460, 1510, 4262, 3230, 1428, 3958, 4091, 3016, 2966, 3146, 2821, 961, 3759, 3879, 4089, 3634]. tree-based [3759]. Tree-Like [4187]. Trees [1330, 2201, 1773, 848, 1442, 2333, 1658, 4213, 1137, 2346, 1104, 1552, 2723, 3992, 963, 1217, 1148, 1326, 3331]. treewidth [2657]. Trends [4006, 3784, 3506, 2799]. Triangle [4251, 1411, 434, 1676, 1039]. Triangle-Free [1411]. Triangular [2289, 4205, 4318, 2127, 2298, 4344, 2042, 2459, 1237, 2616, 2721, 2793, 2718, 2681]. Triangulate [4323]. Triangulated [4209, 1885, 2711]. Triangulation [1885, 4212, 4329, 2462, 3874]. Triangulationen [1933]. Triangulations [4317, 1870, 2640]. Tricomi [429, 1648, 842]. Tridiagonal [1569, 2366, 1521, 2120, 1679, 573, 2013, 2671]. tridiagonaler [2013]. trie [3680]. tries [1327]. trig [1005]. triggered [3611]. triggers [3406]. Trigonometric [1998, 1346, 780, 172, 945, 1102, 2248, 1900, 1295, 1671].

Trigonometrically [4257]. trigonometrische [1998]. trigonometrischen [945, 1102]. trigonometrischer [780]. Trim [4256]. Trimmed [4247, 4308, 4263]. Trinocular [4268]. Trinomials [2562]. trip [2947]. Tripel [1972]. Triple [2578, 1467, 3152]. Triplex [273]. Triplex-Algol [273]. Trommeln [1153]. Truncated [1819]. Truncation [1057, 1654, 1338]. truncations [1218]. Trust [3202, 2147, 3642, 3868, 1602, 1918, 2034, 3655, 3298, 3893, 3545, 3502, 2833, 4086, 3838, 3962, 3449, 2994, 2789, 4011]. trust-based [2833, 3838, 3962]. Trust-driven [3868]. Trust-Region [1602, 2789]. trusted [3136, 3851]. trusting [3586]. trustworthy [3589]. truth [3878]. TSCH [3929, 3623]. Tschebysche [382]. Tschebysche-Problems [382]. Tschebyscheff [766, 578, 370, 500, 110, 504, 693]. Tschebyscheff-Approximation [500, 504, 693]. Tschebyscheff-Approximationen [370]. Tschebyscheff-Verfahrens [766]. Tschebyscheffsche [52]. Tsen [422]. TTI [3006]. Tukey [2352]. tuning [1086, 3172, 3003, 3171]. Tuple [896, 557]. tuply [1217, 1326]. Turan [691, 905]. turbines [2864]. TurboLock [3041]. Turek [2660]. Turing [502, 609, 502, 560, 564, 668]. Turingmaschine [609, 668, 564]. Turingmaschinen [560]. Turna [3778]. Turning [1850, 1906, 1343, 1243, 1697]. Tutorial [3590]. TV [3764]. tweet [3699]. twin [4045]. twin-driven [4045]. Twitter [2942]. Two [1591, 2357, 2504, 2267, 2548, 2275, 1294, 2308, 2567, 1808, 2414, 2025, 4279, 754, 2540, 556, 1722, 1400, 2672, 1588, 411, 1643, 4225, 1112, 2071, 1788, 2250, 2506, 2368, 2420, 763, 2391, 2306, 2314, 2291, 1904, 2209, 1246, 1841, 1626, 1863, 1328, 2716, 274, 341, 956, 4367, 4170, 2198, 2802, 217, 371, 449, 2751, 2593, 561, 2356, 1770, 2462, 2607, 2367, 1416, 2603, 2728, 1394, 1179, 1688, 1763, 4224, 741, 2631, 3907, 3808, 1742, 2616, 4003, 4160, 3427, 1162, 1685, 1701, 1056, 2002, 3660, 3077, 3873, 560, 1439, 1012, 4067, 737, 1093, 1950, 839, 574]. Two- [2314, 2291]. two-archive [3873]. Two-Dimensional [1722, 1400, 1643, 217, 2504, 371, 2603, 4224, 2616, 1056, 574]. Two-grid [2567]. Two-Level [2275, 2368, 2356, 2506, 2209, 2751, 4160, 3427]. Two-Machine [1294, 2414]. Two-Parameter [1863]. two-phase [3907]. Two-Point [1808, 1588, 1788, 1841, 4170, 1742, 1162]. two-sequence [1701]. Two-Sided [956, 4367]. Two-stage [4225, 2802, 3077]. Two-Step [449, 1328, 3660]. two-tape [560]. two-term [1685]. two-tier [4003]. two-yield [2631]. Typ [1933, 1234, 1008, 563, 935]. Type [2286, 2096, 824, 2102, 1789, 2384, 2147, 1941, 4317, 2382, 2525, 691, 750, 1785, 4363, 905, 1820, 747, 2106, 1969, 1008, 753, 2707, 1471, 2826, 3230, 1418, 4223, 2933, 563, 2733, 2006, 716, 937, 1238, 1585, 1574, 2869, 243, 3764, 3554, 1738, 1446, 4233, 1566, 1759, 18, 1575, 3393, 1555, 3757, 1678, 1215, 1563, 3426, 1703]. Type- [4317]. type-diversity [3757]. Types [763, 3360, 524, 1394, 3228, 3106, 4067]. Typesetting [246]. Typs [747]. Typus [1941, 18]. Tzschach [357].

U [1129, 333]. U. [234]. UAV

[3984, 3846, 4063, 3671, 3820, 4009, 4002]. UAV-assisted [3846]. UAV-mounted [3820]. UB [3723]. UB-H [3723]. Über- [452]. Überblick [37]. Überdeckungsprobleme [835]. überlauffreie [456]. Überschätzung [990]. Übersetzung [33, 881]. Übersicht [699]. Übertragung [342]. Ubiquitous [2957, 3447, 3763, 3018, 2983, 3203, 2966, 4002, 3501, 3455]. Übungen [376]. Uebe [648]. UHPr [3455]. Uhr [122]. Ultra [3609, 4028]. Ultra-safe [3609]. ultra-wide [4028]. ULV [2535]. Umformen [478]. UML [3683]. UML2PROV [3619]. Umpacken [1947]. Umpalumpa [3800]. Umwandlung [81, 82]. unabhängiger [863]. Unambiguity [896]. unbalanced [3723, 3143, 3819]. unbalanced-hierarchical [3723]. Unbekannten [1917]. unbetwerteten [710]. Unbounded [2154, 2359, 1288]. Unboxing [3816]. Uncertain [2794, 3528, 3281, 2824]. uncertainties [2823]. Uncertainty [891, 453, 3015]. Unconditional [562]. Unconstrained [2085, 2033, 1310, 2072, 487, 1951, 2059, 1027, 1918, 1467, 1636]. under-relaxation [452]. Underdetermined [2189, 1036]. Underflow [1284, 1498]. undersampling [3484]. understand [3406, 3585]. Understanding [732, 3333, 2200, 3647, 4133, 3322]. underwater [3353, 3300, 3247]. undirected [2921]. undiscounted [967]. uneigentlicher [850]. unendlich [558]. unendlichdimensionalen [613]. Unendliche [652]. unendliehen [499]. Unger [102, 418, 324]. Ungleichun [456]. Ungleichungssysteme [772]. Unified [4332, 4221, 2752, 1485, 1096]. Uniform [2232, 2540, 1353, 1871, 2307, 1583, 3753, 2340, 2332, 2263, 2593, 3438, 1705, 2217, 3748, 2397, 2682, 2787, 1697]. Uniformly [1895, 1637, 4198, 2343, 1557, 664, 1693]. unilateral [2676]. Unimodal [4242, 1417]. Union [1544]. unions [4228]. uniprocessor [3194]. Uniqueness [2268, 762, 78, 504, 457, 1190]. Unit [1203, 1294, 525, 2495, 2298, 1396, 2151, 943, 1210]. Unit-clause [525]. Unit-Time [1294, 2151]. unitären [469]. Unitary [469]. Units [863, 2828, 2903, 3187]. unity [1370]. Univariate [1538, 2328, 1505, 4230, 1845, 1670, 1714]. Universal [1187, 1189, 609, 2007]. universelle [609]. University [386]. unknown [3704, 4088]. unknowns [2727, 459]. unmanned [3976, 3925]. Unorganised [4323]. Unraveled [2347]. Unraveling [3894]. unreliable [3817]. Unrestricted [4118]. unrestringierte [1951]. unrestringierten [1027, 1918]. unsaturated [1609]. unslotted [3664]. Unstable [1397, 1999]. unsteady [2857, 2869]. unstetiger [789]. Unstructured [2576, 2234, 2086, 1519, 2360, 2113, 2969]. unsupervised [3693, 3516, 3762]. Unsymmetric [1432, 2064]. unsymmetrical [1325]. unterbestimmten [1036]. untere [839]. Unterhalbgruppenverbände [763]. Unternchmensleitung [282]. Unternehmensforschung [85, 511, 658, 51]. Unterrelaxation [452]. Unterricht [41]. Untersuchung [1960, 1024]. Untersuchungen [578, 207, 953, 130, 161, 789]. Untersuchungsmethoden [336]. unterworfen [1]. unvollständigen [344]. unvollständiger [438]. unwesentlich [567]. upcoming [2814]. Update [1968, 2932, 3168]. Updated [2108, 1286]. updates [2642, 1230]. updating [1022]. upon [2114, 3842]. Upper [2110, 1573, 1396, 1740]. Ups [953, 456]. upwind [2595, 1535, 2646]. ur [66, 32]. urban [3452, 3408, 3864]. Usability [4103, 3160, 3421, 3336]. Usage [204, 3149, 3769, 3327, 3423]. Use [89, 2743, 891, 1145, 4246, 77, 1551, 1134, 467, 2413, 858, 4128, 990, 2225, 496, 620, 3530, 1988, 51, 3371, 1015, 1136, 1982, 3548]. Used [2175]. Useful [1881, 4154]. user [3911, 3407, 2709, 3643, 3161, 3698, 3902, 3433, 2831, 3282, 2845, 3058, 2792, 4013, 3613, 2940]. user-driven [2831].

User-Level [2940]. user-preference [3613]. Users [3647, 3272, 3973, 3123, 3620, 3436]. Uses [436]. Using [1360, 4174, 3741, 1868, 2762, 2045, 2134, 3968, 2576, 4315, 3037, 2440, 1410, 4251, 760, 1680, 4301, 1045, 2991, 4322, 2181, 1653, 1860, 2451, 3029, 1784, 1900, 2249, 2507, 1314, 1853, 1944, 2185, 4323, 2122, 2332, 2521, 3303, 1823, 1919, 1587, 4308, 1427, 1523, 1959, 2087, 4328, 1799, 2165, 2192, 2580, 2237, 2463, 4056, 3573, 3641, 3369, 3745, 3840, 3857, 2752, 4004, 3992, 3272, 2705, 3264, 2934, 2642, 3432, 2628, 3230, 3301, 2943, 3600, 1765, 3767, 3712, 3581, 3471, 3526, 2819, 3880, 3545, 3891, 3782, 1735, 3343, 2673, 3928, 3359, 3726, 2944, 3060, 3421, 3538, 2643, 3931, 3959, 3549]. using [3980, 3933, 3255, 3953, 3727, 4007, 3828, 4026, 3495, 3897, 2688, 3906, 3463, 2868, 3419, 3958, 4091, 3500, 2985, 3569, 3274, 4060, 2873, 3964, 2818, 3242, 3098, 3631, 4028, 3349, 3837, 3996, 3970, 3525, 3358, 3973, 3742, 2998, 3428, 4038, 3497, 3967, 3775, 1738, 3847, 3340, 3248, 3260, 3888, 3051, 4233, 3831, 2814, 3039, 4005, 1350, 3926, 3736, 2935, 3151, 3997, 2995, 2837, 3217, 3173, 3249, 7, 3639, 1528, 4082, 3681, 3582, 3760, 3065, 3519, 3737, 3865, 3929, 3245, 3942, 3414, 3790, 4021, 2904, 3135, 3562, 3916, 3720, 3692, 3283, 2718, 2614, 1976, 3235, 3232, 3292, 4064, 2603]. using [3073, 3383, 3807, 2942, 3046, 3168, 3385, 3934, 3144, 3219]. Utilising [2979]. utility [3028]. Utilization [1500, 3530, 3459]. utilizing [4000, 3296, 3827].

V [375, 455, 72, 70, 271, 138, 698, 575, 100, 797, 532, 40, 212, 187, 217, 213, 183, 259, 541, 418, 1900, 3778]. V. [388, 149, 671, 40, 212, 793, 322, 493]. V2V [3665, 3812]. V2V2I [4077]. vaccine [3647]. vadis [3276]. Vakuumtechnik [330]. Valach [328]. Validated [4235]. validating [1756, 4172, 4231]. Validation [4376, 2047, 4165, 2126, 2598, 2918, 3015, 2817]. Validity [1625, 3682]. Valuation [2568]. Value [2055, 2173, 2233, 1840, 599, 1934, 1141, 1300, 1397, 4337, 1963, 1808, 918, 1771, 1254, 1588, 760, 1045, 1788, 819, 1673, 4286, 2370, 1314, 2093, 1999, 912, 1465, 1315, 2291, 1501, 1290, 1174, 849, 994, 2276, 2365, 1841, 1926, 2052, 2339, 2553, 4170, 295, 2165, 1958, 1151, 877, 1536, 2367, 4171, 753, 1502, 1578, 1296, 297, 2785, 1644, 1455, 1161, 1242, 1175, 1436, 1615, 1324, 1667, 1211, 1150, 1662, 1748, 988, 1364, 995, 1013, 1271, 2663, 1609, 1742, 2791, 2735, 3988, 384, 1351, 445, 1725, 2682, 2809, 1162, 1989, 410, 1011, 1572, 18, 53, 1726, 1977, 1107, 1125]. value [1339, 1275, 1755, 1392, 3551, 1156, 1741, 689, 2171, 1483, 2786, 3762, 3922, 1193, 1611]. value/quality [3988]. Valued [1167, 555, 741, 737, 2820, 1256]. Values [1801, 2564, 1316, 1468, 2298, 1844, 464, 604, 1770, 209, 275, 2486, 4078, 1155, 3889, 2622, 1127, 53, 593]. VAMPIRE [3686]. VANET [3608, 3523, 3102]. VANETs [3702, 3806, 3340]. Vanishing [2091]. Variable [2426, 736, 2366, 2337, 4319, 2246, 1381, 1164, 2296, 2037, 2078, 2465, 2109, 1388, 3479, 1561, 1140, 272, 715, 2627, 4168, 1311, 4077, 3144, 1349]. variable-length [3479]. Variable-Metric [736]. Variable-Order [2337]. Variable-Sized [2109]. Variables [1591, 2548, 1893, 1917, 4181, 900, 920, 2361, 1705, 1459, 1249, 1363, 1229, 53, 1800]. Variably [2302]. Variance [2237]. variant [2689, 1665]. Variate [1417, 1568, 1851, 1370, 2542]. Variates [2192, 1952, 1463]. Variation [2195, 47]. Variational [1868, 1997, 2313, 4208, 4302, 4254, 2538, 2635, 34, 4347, 4273, 2197, 2070, 2651, 1364, 2677, 4070, 947]. Variations [225, 1854]. Variationsungleichungen [1997]. Varieties [4316]. variety [2926]. Various [2043, 567, 3911, 3228, 3982, 3563]. variously [3323]. variously-relaxed [3323].

Varying [1478, 3957, 2779]. Vector [2228, 1134, 3743, 3909, 1712, 3495, 4048, 613, 3841, 3189, 4130, 3151, 669, 4126]. vectorization [3008]. vectorized [3686]. vectorizing [3064]. Vectors [635, 25, 2266, 1637, 308, 3720]. vehicle [3581, 4066, 3605]. Vehicles [3884, 3976, 3925, 3657, 3523, 3974, 4001, 3302, 3812, 4077, 3475]. vehicular [3626, 2952, 3702, 3806, 3356, 3458, 3862]. vein [3288]. Vektoren [25]. Vektorprozessoren [4126]. Vektorräunen [613]. Vektortransformation [669]. Vel [1384]. Velocities [2344]. Velocity [2856]. Velocity-pressure [2856]. Velte [2354]. venous [2754]. Verallgemeinerte [1128, 1954, 988, 240, 770]. verallgemeinerten [1234]. verallgemeinerter [314, 912]. Verallgemeinerung [456, 519, 78, 6]. Veränderlichen [272, 53]. Veränderungen [807]. Verarbeitung [1914, 252]. verbal [950]. Verbandsstrukturen [4102]. Verbandstheoretische [567]. Verbesserte [1098]. Verbesserung [1131, 1545, 442]. Verbindung [669]. verbose [3299]. Vereinen [4096]. vereinfachtes [527]. Verfahren [1, 823, 6, 599, 1915, 1995, 836, 889, 741, 737, 129, 207, 760, 1940, 1084, 370, 1968, 617, 701, 241, 385, 1260, 1503, 497, 591, 55, 805, 1949, 582, 18, 527, 22, 296, 413, 606, 935, 956, 4367, 1234, 414, 1113, 973, 1059, 1964, 182, 1918, 1954, 586, 638, 914, 595, 924, 971, 520]. Verfahrens [766, 682, 384, 783, 1655, 1861, 887]. Verfahrenzur [432]. Vergleich [861]. Vergröberung [917]. Verhalten [130, 161]. VERICOMP [2810]. verifiable [3340, 3297, 4011]. Verification [4179, 2161, 2493, 4169, 2511, 4345, 4238, 2323, 2918, 4017, 4373, 4233, 1015, 3044, 1136, 4001]. Verifications [4333]. Verified [4177, 2142, 4224, 2199, 2014, 4176, 2823, 4170, 2810, 1821, 2915, 2819, 4229, 2821, 4237, 2807]. Verify [2268, 2510]. Verkehrsablaufes [80]. Verkehrsflußmodells [559]. Verkehrsnetzen [1113]. Verkehrstheorie [4128]. Verknüpfungen [4096]. Vernon [289]. verschiedenen [851]. verschiedener [567]. Versicherungsmathematik [677]. Version [1641, 2129, 1348, 2356, 456, 3675, 2669, 1286, 1616]. versioned [3130]. Versions [1600, 3469, 1656, 3246]. Versuchspläne [697]. Versuchsund [114]. Versus [2224, 3537, 3546]. verteilten [1952]. Verteilung [316, 556, 637]. Vertex [2153, 3345, 3068, 1865, 1439]. vertex-centric [3345]. vertical [3858, 3882]. Vertices [4312, 906, 1357]. Verwandtschaft [1119]. Verwendung [440, 33, 1234, 694]. Verwerfung [1952]. Very [942, 1835, 2774]. Verzogerung [1967]. Verzweigungsprobleme [1226]. vessel [2754]. Via [1952, 2648, 3384, 6, 2, 1615, 1477, 3893, 3848, 2166, 3852, 3894, 3858, 1032, 4043, 4041, 3708, 3867, 4063, 272, 1774, 1737, 2572, 2277, 2796, 3874, 3285, 4075, 3482, 3291]. viable [3062]. VIB [4177]. vibrating [987]. Vibration [289]. Vibrations [4177, 2227]. vibratory [2935]. Vidal [550]. Video [4284, 3107, 3079, 3958, 4091, 3650, 3731, 3073]. vielen [558, 805]. Vier [312]. vierdimensionale [634]. Vierecksgittern [1996]. vierter [467, 975, 986]. View [2523, 2946, 3139]. Views [4278, 3163, 3520]. VIFL [4026]. Vilenkin [152]. Violation [942]. Virtual [3304, 3124, 3640, 2838, 3164, 2756, 3575, 3441, 3928, 3980, 3933, 3595, 3085, 3031, 4010, 3888, 3459, 3125, 3100, 3769, 3190, 3423, 3869, 961]. virtualization [3306, 3167, 3505]. virtualized [3070, 3834, 3095]. Virtuelle [961]. virtuellem [1023]. virus [3724]. viscoelastic [506]. Viscous [1942]. Visible [1519]. Vision [3581, 4272, 4273, 4070, 4378, 3250, 2829].

Vision-based [3581]. viskosen [1942]. Visual [2446, 4265, 2947, 3691, 3029, 3990]. Visualization [4294, 2415, 2194, 2941, 2711, 2884]. visualizer [2948]. Visualizing [2944, 2945, 2943, 2949]. Visually [1815, 3517]. vital [3732, 3871, 3262]. VITAL-compliant [3262]. Vitäsek [135]. Viterbi [3187]. VLDNet [3581]. VM [3445, 3000, 3818, 4010, 3991, 3518, 4021]. vmBBProfiler [3190]. VMs [3359]. VNE [3509]. VNF [3834, 3957]. Vo5G [3485]. VoD [3075]. Vogel [291]. Vogelius [2667]. Vogl [153]. Vol [354, 375, 68, 16, 35, 426, 646, 36, 650, 648, 649, 627, 700, 581, 104, 744, 628, 540, 530, 534, 582, 651, 389, 481, 40, 125, 672]. Vol. [42]. Volatilities [2568]. Voll [1961]. Voll- [1961]. vollautomatische [162]. Vollkugel [1754]. vollständigen [298]. Vollständigkeit [1461]. vollstetiger [409, 433]. Volterra [1967, 1282, 1967, 2555, 2727, 1323, 1081, 824, 774, 1157, 1099, 1667, 1684, 1830, 870, 1046, 1777, 1359, 1983, 1115, 948, 2406, 2240, 2485, 2661, 1201]. Volterraschen [1046]. Volume [2096, 2233, 2279, 2576, 4315, 2292, 2568, 2029, 43, 233, 2140, 3556, 2646, 238]. Volumes [4312, 1876, 2890]. Volumetric [4276]. Vom [286, 105, 563, 1941, 33, 935, 1234, 1008]. Vonhof [415]. vor [1032]. Vorgabeexperimente [590]. vorgebbaren [1062]. vorgegebenen [593]. vorgeschriebenen [837]. Vorhersage [1376, 1683]. Vorlesungen [216, 260]. vorliegended [862]. Voronoi [2838, 2610]. Vorteil [1947]. Vote [2283]. voting [3992, 4045]. voting-GA [4045]. vulnerabilities [3264, 3266]. vulnerability [3931, 4026, 4035, 3265].

W [335, 45, 792, 826, 14, 27, 551, 579, 15, 377, 305, 355, 358, 1554, 236, 13, 685, 582, 1884, 363, 481, 142, 373, 535, 538, 624, 672, 658]. W. [415, 531, 73, 221, 639, 515, 281, 291, 406, 64, 146, 685, 1884, 363, 647, 136, 795, 733, 511]. Wacker [794, 1859, 1937]. Wahl [1058]. Wahrscheinlichkeitsrechnung [576]. Wahrscheinlichkeitstheorie [46]. Waiting [1204, 3259, 1321]. Walker [265, 548]. walking [3582]. walks [3053]. WANETs [3795]. war [3322, 3415]. warehouse [3480]. Wärmeleitungsgleichung [1045]. Wärmeleitungsprobleme [467]. warming [4007]. warning [3382]. Warteschlangen [274]. Warteschlangensystemen [1461]. Wartesystem [274]. Was [219]. water [1609]. watermarking [2995, 3544, 3414, 4071]. Wave [2376, 2563, 2466, 4364, 3701, 385, 1208, 3926, 1630]. Waveform [2074, 1977]. Wavefront [1856]. Wavelet [2115, 2537, 2551, 2702, 2790, 2455, 4327, 2036, 2613, 2987, 3865]. Wavelet-Based [2455]. Wavelets [2447, 4317, 3227, 1900, 1990, 3271]. Waves [696, 3291]. Way [4343, 3359, 2796]. Ways [1772]. WBANs [3962]. WCET [3499]. Weak [2483, 768, 554, 1445]. Weakly [2555, 2264, 619, 241, 2485, 2727, 1764, 1741]. weapons [3691]. wearable [3526]. Weather [3880, 127]. web [2915, 3062, 3059, 3910, 3033, 3897, 2978, 3560, 3153, 3265, 3248, 3266, 3329, 2974, 2943, 3159, 3060, 3033, 2925, 3363, 3278, 2884, 2900]. web-service [2915]. Weber [733, 3286]. Weber-face [3286]. WebRTC [3333, 3079]. WebRTC-based [3079]. weed [3281]. Weganzahlen [960]. Wege [293, 222, 848]. Wegen [791, 930]. Weglängen [807]. Weibull [1631, 2848]. Weierstrass [1019, 4232, 4358, 2513]. Weierstrass-like [2513]. Weight [4245, 2482, 2003, 3040, 3022, 993]. Weight-Optimal [993]. Weighted [2313, 2363, 2206, 2038, 3175, 3080, 2389, 2707, 3230, 2729, 3959, 3852, 3831, 3825, 3245, 2614, 3558, 2739, 3288]. weighting [667]. Weights [144, 2505, 2069, 1433, 1091]. Weimar [455, 200, 455]. Weinstein [682, 682]. Weiszfeld [1219]. Weiterentwicklung [893]. welches [777]. Well [2434, 2345, 78, 3465]. Well-Solvable [2345]. Weltbild [31]. wenigen [1917]. Wenke [580]. Wentzel [194]. Werkzeug [745]. Werner [675, 800, 1543]. Wertebereich [466]. Wertebereiches [4101]. Wertebereichs [990]. Wertemengen [593]. wesentlich [1236]. wesentlich-nichtnegativer [1236]. Westermayer [629, 629]. wet [2864]. whale [3854]. wheels [2629, 3642]. Which [867, 1404, 445, 4138]. whose [1]. Wickes [406]. wide [4028]. Widely [851]. width [3997, 1311]. Wiedergabe [1]. Wiener [2689]. Wiley [542]. Wilf [233, 334, 546, 1014]. Wilf-quadrature [1014]. Wilfert [65]. Wilkins [802]. Wilkinson [484, 155, 1457]. Wilks [90]. Willard [1896, 1905]. Wilson [103, 1574, 1348, 1575]. Wilson-type [1574, 1575]. Window [2535, 4048]. Winkler [270]. Winograd [196]. wireframe [2717]. Wireless [3661, 3649, 3055, 3559, 3365, 2953, 3552, 3221, 3802, 2997, 3773, 3570, 3578, 3924, 3129, 2999, 2688, 3275, 3300, 3569, 3242, 2920, 3846, 3497, 3251, 3427, 3508, 3309, 3019, 3845, 3217, 3929, 3476, 3535, 3776, 3807, 3063, 2989, 3200, 3206]. WirelessHART [3083]. Wirkungsweise [702]. Wirtachaftsalgebra [580]. Wirtschaft [26, 238, 418]. Wirtschafts [185]. Wirtschafts-Sozialwissenschaften [185]. Wirtschaftswissenschaft [482]. Wise [565, 3723]. Wiss [488]. Wissenschaften [674, 362, 389, 481, 390]. within [943, 2800]. without [3240, 2828, 1097, 1599, 2491, 4267, 955, 4269, 2221, 1687, 413, 1272, 1697]. WITS [3212]. Witzgall [674]. WK [2271]. WK-Recursive [2271]. Wolf [195, 431, 198]. Wolff [677]. Wolodin [576]. Word [909, 3380]. Words [2082]. work [3674]. workers [3932]. workflow [3196, 3550, 3822, 3761, 3368, 3195, 3216, 3826, 3519, 3942, 3790, 3612, 3140, 3027, 3886, 3707, 3219]. workflows [3887, 3402, 4012, 3594, 3000, 3255, 3141, 3121, 3794, 3025, 3095]. Working [1897]. Workload [3625, 3966, 3355, 4022, 3712, 3394, 3538, 3591, 3995, 4024, 3719, 3950, 4029]. workload-aware [3538, 3719]. workloads [4069, 3800, 3957]. Workshop [2509, 245, 4380, 4378, 4379]. world [2800]. Worst [2224, 2109, 2729, 2016]. Worst-Case [2224, 2109]. Wörterbuch [41]. Worthäufigkeitsforsehung [32]. Wosnik [124, 267]. WoT [3528]. wrapper [3660]. Wrapping [2076, 2221, 4234, 3874]. write [3968, 3257]. Wroblewski [539]. WSN [3489]. WSNs [3253, 3623, 3515]. Wurzelmenge [807]. Wurzeln [4099]. Wurzelortskurven [240]. WVSN [3811]. Wymore [335].

X [3785]. X-IoT [3785]. XBridge [2978]. XBridge-Mobile [2978]. XDN [3339]. XII [511]. Xing [3161]. XIV [658]. XL [3949]. XML [2978, 3208, 3158]. XR [3951]. XSC [2815, 4239]. Xu [2548].

Ya [138, 152]. years [1793, 1896, 1554, 1508]. yield [1455, 2631, 3350, 3738]. YOLO [3474, 4009, 4047]. YOLO-based [3474]. yolov7 [3984]. YOLOv8 [3990]. York [830]. young [3349].

Zahlen [556, 55, 862, 2039, 50, 566]. Zahlendarstellung [314]. Zahlensysteme [1852]. Zahlkörpern [863]. Zassenhaus [2191]. Zehnder [357]. Zeichen [1]. Zeichnen [1016]. Zeilen [244]. Zeilen- [244].

Zeilenkorrektur [1968]. Zeilenkorrektur-ABS- [1968]. Zeilensummenkriterium [554]. Zeitbedingungen [1042, 1147]. Zeiten [1935]. Zeitgewinns [953]. Zeitkonstanten [851]. Zeitparameter [4135]. zelbestimmung [1949]. Zemanek [304, 801]. Zener [391]. zentrische [1308]. Zerlegbaren [477]. Zerlegung [365, 1020]. Zernike [2613]. Zero [823, 4357, 162, 2072, 1816, 1205, 3893, 1657]. Zero-One [2072, 1816, 1205, 1657]. Zeroes [1680, 1365, 432, 20, 1449, 166]. Zeros [160, 890, 616, 4145, 1812, 366, 705, 1889, 1549, 2265, 2247, 2422, 458, 2077, 1368, 1268, 1550, 1809, 1949, 2244, 4346, 4362, 2514, 2513, 921, 4222, 1179, 1213, 1664, 1766, 1235, 1191]. Zertifikat [167, 168, 345, 8, 58]. zeta [1269]. Zetafunktion [1269]. Zhang [2779]. Zhen [1805]. Zhidkov [12]. Zhurina [199]. Zielke [141]. Zierer [43, 43]. Ziffernrechner [98]. Zilinskas [1845]. Zipf [2765]. Zopf [567]. Zopfgruppe [567]. zufallig [1946]. Zufallsfunktionen [336]. Zufallsrichtungen [1923]. Zufallszahlen [1952]. Zugriff [1131]. Zuhe [2622]. zum [439, 106, 605, 512, 760, 1016, 1020, 662, 570, 410, 22, 270, 252, 130, 161, 478, 554, 4107]. Zuordnungsproblem [2008]. Zuordnungsprobleme [593]. Zur [1038, 1, 465, 823, 890, 599, 1170, 692, 710, 740, 1915, 4126, 438, 889, 1994, 741, 737, 1254, 1632, 586, 638, 65, 396, 988, 366, 386, 1265, 1960, 1515, 569, 370, 1968, 397, 1917, 1269, 1972, 983, 617, 608, 1999, 382, 693, 379, 241, 385, 960, 879, 933, 1503, 1030, 1079, 112, 497, 55, 851, 1000, 805, 1949, 163, 1011, 29, 30, 37, 31, 33, 38, 43, 1920, 18, 1070, 697, 614, 956, 4367, 707, 1256, 809, 1931, 414, 980, 1113, 973, 953, 1059, 694, 182, 833, 495, 789, 689, 791, 930, 952, 984, 658]. zur [1918, 2013, 158, 1024, 766, 1037, 1110, 1957, 945, 1021, 1943, 521, 1089, 872, 409, 433, 632, 1540, 972, 611, 855, 859, 1852, 663, 975, 49, 1438, 4101, 1944, 724, 850, 457, 1154, 571, 1402, 866, 344, 895, 1945, 1063, 903, 4108, 615, 1131, 1226]. Zürich [579]. Zurmühl [154]. zusammenhängende [4369]. zusammenhängenden [740]. zusammenhängender [112]. zusätzlichen [3]. zusaunmenhängenden [617]. Zustand [987, 840]. Zustandsreduktion [438]. zuverlassige [1917]. zwei [741, 763, 274, 737, 1950, 839, 560]. Zwei-Band [560]. Zweidimensionale [371, 1139]. zweidimensionalen [1942]. zweier [556]. zweiseitigen [956, 4367]. Zweiter [1932, 838, 908, 444, 1358]. zweitzer [470]. Zwikker [268]. zwischen [461, 278]. Zyklen [741, 737]. Zypkin [264, 338].

References

Albrecht:1966:VIZ

[1] R. Albrecht and W. Werner. Ein Verfahren zur Identifizierung von Zeichen, deren Wiedergabe stationären statistischen Størungen unterworfen ist. (German) [A method for identification of symbols whose reproduction is subject to stationary statistical perturbations]. Computing, 1(1):1–7, March 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bellman:1966:ACQ

[2] R. Bellman, R. Kalaba, and R. Sridhar. Adaptive control via quasilinearization and differential approximation. Computing, 1(1):8–17, March 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Meissl:1966:TBZ

[3] P. Meissl. Über das transportproblem mit beliebigen zusätzlichen linearen nebenbedingungen. (German) [On the transport problem with arbitrary additional linear constraints]. Computing, 1(1):18–38, March 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Meinardus:1966:AAE

[4] G. Meinardus. Über die Approximation asymptotischer Entwicklungen I. (German) [On the approximation of asymptotic series I]. Computing, 1(1):39–49, March 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kuich:1966:BSM

[5] W. Kuich and K. Walk. Blockstochastic matrices and associated finite-state languages. Computing, 1(1):50–61, March 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Apostolatos:1966:VFV

[6] N. Apostolatos. Verallgemeinerung funktionentheoretischer Verfahren für Analogrechner auf allgemeinere Funktionentheorien. (German) [Generalization of a function-theoretic method for analog computers via ordinary function theory]. Computing, 1(1):62–76, March 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rutishauser:1966:ALG

[7] H. Rutishauser. Algorithmus 1: Lineares Gleichungssystem mit symmetrischer positiv-definiter Bandmatrix nach Cholesky. (German) [Algorithm 1: Linear equation systems with symmetric positive definite band matrix using Cholesky]. Computing, 1(1):77–78, March 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See certification [8].

Nake:1966:ZAC

[8] F. Nake. Zertifikat zu Algorithmus 1, Cholbd. (German) [Certification of algorithm 1, Cholbd]. Computing, 1(1): 79, March 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [7].

Gunzenhauser:1966:IKI

[9] R. Gunzenhäuser. Mitteilung/Note: IFIP-Kongreß 65 — Interdata 65 — ein Oberblick. (German) [IFIP Congress 65 – Interdata 65 — an overview]. Computing, 1(1):80–87, March 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Knodel:1966:FBJ

[10] W. Knödel. Book review: F. L. Bauer, J. Heinhold, K. Samelson, und R. Sauer, Moderne Rechenanlagen (Leitfäden der angewandten Mathematik und Mechanik: Band 5). Computing, 1(1):88, March 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1966:FMA

[11] H. Kremser. Book review: F. M. Arscott, Periodic Differential Equations. Computing, 1(1):88, March 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Collatz:1966:BPZ

[12] L. Collatz. Book review: I. S. Berezin and N. P. Zhidkov, Computing Methods. Computing, 1(1):89, March 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1966:BRDa

[13] H. Kremser. Book review: D. D. McCracken und W. S. Dorn, Numerical Methods and Fortran Programming. Computing, 1(1):90, March 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Beilner:1966:BRH

[14] H. Beilner. Book report: H. P. Künzi und W. Schilling, Einführung in die elektronische Datenverabeitung. Computing, 1(1):91, March 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eberl:1966:BRJ

[15] W. Eberl. Book review: J. Heinhold und K. W. Gaede, Ingenieur-Statistik. Computing, 1(1):91, March 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1966:BRGa

[16] E. Bukovics. Book report: G. Meinardus, Approximation von Funktionen und ihre numerische Behandlung (Springer Tracts in Natural Philosophy, Ergebnisse der angewandten Mathematik: Vol 4). Computing, 1(1): 92, March 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bauer:1966:SSF

[17] F. L. Bauer. Syntaktische Struktur formaler und natürlicher Sprachen. (German) [Syntactic structure formulas and natural languages]. Computing, 1(2):93–104, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sauer:1966:NVL

[18] R. Sauer. Numerische Verfahren zur Lösung n-dimensionaler Anfangswertprobleme von hyperbolischen Typus. (German) [Numerical method for solution of n-dimensional initial-value problems of hyperbolic type]. Computing, 1(2):105–114, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Meinardus:1966:IRA

[19] G. Meinardus. Invarianz bei rationalen Approximationen. (German) [Invariance of rational approximations]. Computing, 1(2):115–118, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Heinhold:1966:NEA

[20] J. Heinhold and U. Kulisch. Über die Nullstellen der ersten Ableitung von Besselfunktionen. (German) [On the zeroes of the first derivation of the Bessel function]. Computing, 1 (2):119–126, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Filippi:1966:BEE

[21] S. Filippi. Die Berechnung einiger elementarer transzendenter Funktionen mit Hilfe des Richardson-Algorithmus. (German) [The computation of some elementary transcendental functions by means of the Richardson algorithm]. Computing, 1(2):127–132, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmid:1966:VET

[22] K. Schmid. Ein Verfahren zum Erkennen von Transportproblemen aus der Simplexmatrix. (German) [A method for recognition of the transport problem through the simplex matrix]. Computing, 1(2):133–145, June 1966. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Niethammer:1966:IKA

[23] W. Niethammer. Iterationsverfahren bei der konformen Abbildung. (German) [Iterative method for conformal maps]. Computing, 1(2):146–153, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Knapp:1966:BID

[24] H. Knapp. Bemerkung zu iterationsverfahren bei Differentialgleichungen. (German) [Remark on an iterative method for differential equations]. Computing, 1(2):154–158, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rutishauser:1966:AOV

[25] H. Rutishauser. Algorithmus 2: Orthonormierung von Vektoren nach E. Schmidt. (German) [Algorithm 2: Orthonormalization of vectors by E. Schmidt]. Computing, 1(2):159–161, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See certification [58].

Brusberg:1966:BRE

[26] H. Brusberg. Book review: S. Eilon, Tafeln and Tabellen für Wirtschaft und Industrie. Computing, 1(2):162, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1966:BRD

[27] E. Bukovics. Book review: D. K. Faddejew und W. N. Faddejewa, Numerische Methoden der linearen Algebra. Computing, 1(2):163, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1966:BRHa

[28] E. Bukovics. Book review: H. Frank (Herausgeber), Lehrmaschinen in kybernetischer und pädagogisscher Sieht, Band 2. Computing, 1(2):163, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Roos:1966:BRG

[29] P. Roos. Book review: E., von Glaserfeld, A Project for Automatic Sentence Analysis (Beitrage zur Sprachkunde und Informationsverarbeitung, Heft 4, 38–46). Computing, 1(2):164, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Roos:1966:BRHa

[30] P. Roos. Book report: A. Hoppe, Der sprachliche Formulierungsprozeß in den Funktionsebenen der Sprache (Beiträge zur Sprachkunde und Informationsverarbeitung, Heft 4, 7–32). Computing, 1(2):165, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Roos:1966:BRK

[31] P. Roos. Book report: Koschmieder, E., Sprache und Weltbild (Beiträge zur Sprachkunde und Informationsverabeitung, Heft 3, 8–18). Computing, 1(2):165, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Roos:1966:BRR

[32] P. Roos. Book review: R. Hartmann, Bedeutung und Entwicklung der Worthäufigkeitsforsehung (Beiträge ur Sprachkunde und Informationsverarbeitung, Heft 4, 47–57). Computing, 1(2):165, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Roos:1966:BRL

[33] P. Roos. Book report: L. Lippmann, Praktische Probleme bei der Verwendung eines Elektronenrechners für die Übersetzung vom Englischen ins Deutsche (Beiträge zur Sprachkunde und Informationsverabeitung, Heft 3, 62–75). Computing, 1(2):166, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Parkus:1966:BRG

[34] H. Parkus. Book report: S. G. Mikhlin, Variational Methods in Mathematical Physics. Computing, 1(3):166, September 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1966:BRGb

[35] E. Bukovics. Book report: G. Sansone und R. Conti, Nonlinear Differential Equations (International Series of Monographs in Pure and Applied mathematics: Vol. 67). Computing, 1(2):167, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dirl:1966:BRM

[36] R. Dirl. Book report: M. A. Naimark, Linear Representations of the Lorentz Group (International Series of Monographs in Pure and Applied Mathematics: Vol. 63). Computing, 1(2):167, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Roos:1966:BRHb

[37] P. Roos. Book review: H. Schnelle, Maschinelle Sprachübersätzung – ein kritischer Überblick, 1. und 1. Teil (Beiträge zur Sprachkunde und Informationsverarbeitung, Heft 3, 41–61, Heft 4, 58–67). Computing, 1(2):168, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Roos:1966:BRS

[38] P. Roos. Book report: A. Schopf, Grammatische Kategorie und Satz in traditioneller und strukturalistischer Sieht (Beiträge zur Sprachkunde und Informationsverarbeitung, Heft 3, 19– 40). Computing, 1(2):168, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Parkus:1966:BRS

[39] H. Parkus. Book report: S. L. Sobolev, Partial Differential Equations of mathematical Physics, 3 Aufl. Computing, 1(2):169, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Scholz:1966:BRV

[40] H. Scholz. Book review: V. J. Smirnow, A Course of Higher Mathematics. Vol. I–V (International Series of Monographs in Pure and Applied Mathematics, Vol. 57 to 62). Computing, 1(2):169, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1966:BRW

[41] E. Bukovics. Book report: Wörterbuch Programmierter Unterricht. Computing, 1(2):170, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kummer:1966:BMB

[42] W. Kummer. Book review: B. M. Budak, A. A. Samarskii and A. N. Tikhonov, A Collection of Problems on Mathematical Physics (International Series of Monographs on PUre and Applied mathematics, Vol. 52). Computing, 1(2):170, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Roos:1966:BRZ

[43] P. Roos. Book report: E. Zierer, Ein Model zur Darstellung von diskontinuierlichen Konstituenten (Beiträge zur Sprachkunde und Informationsverarbeitung, Heft 4, 33–37). (German) [E. Zierer, A model for representation of discontinuous constituents (Contributions to linguistics and data processing, volume 4, 33–37)]. Computing, 1(2):170, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Baron:1966:BGA

[44] G. Baron. Book review: L. Baumgartner, Gruppentheorie, 4 Aufl. Computing, 1(2):171, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Baron:1966:BRW

[45] G. Baron. Book report: T. W. McRae, The Impact of Computers on Accounting. Computing, 1(2):171, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eberl:1966:HBW

[46] W. Eberl. Book review: H. Bauer, Wahrscheinlichkeitstheorie und Grundzüge der Maßtheorie, Band I (Sammlung Göschen: Band 1216/1216a). Computing, 1(2):171, June 1966. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Scholz:1966:BRG

[47] H. Scholz. Book report: L. G. Johnson, Theory and technique of Variation Research. Computing, 1(2):171, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Anonymous:1966:DBH

[48] Anonymous. D. B. Hertz and R. T. Eddison (Editors), Progress in Operations Research. Computing, 1(2):172, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Handler:1966:TEA

[49] W. Händler. Zur Theorie endlicher Automaten. (German) [On the theory of finite automata]. Computing, 1(3): 173–181, September 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schonhage:1966:MGZ

[50] A. Schönhage. Multiplikation großer Zahlen. (German) [Multiplication of large numbers]. Computing, 1(3):

182–196, September 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Heinhold:1966:AAU

[51] J. Heinhold. Die Anwendung des Analogrechners in der Unternehmensforschung. (German) [The use of analog computer in operations research]. Computing, 1(3):197–213, September 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brosowski:1966:TAD

[52] B. Brosowski. Über Tschebyscheffsche Approximationen durch asymptotisch konvexe Funktionenfamilien. (German) [On Chebyshev approximation with asymptotically convex function families]. Computing, 1(3): 214–223, September 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmidt:1966:EFF

[53] J. W. Schmidt and H.-F. Trinkaus. Extremwertermittlung mit Funktionswerten bei Funktionen von mehreren Veränderlichen. (German) [Determination of the extremum value by means of function values for functions of several variables]. Computing, 1(3): 224–232, September 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Collatz:1966:EAB

[54] L. Collatz. Einige abstrakte Begriffe in der numerischen Mathematik (Anwendungen der Halbordnung). (German) [Some abstract notions in numerical mathematics. (applications et semiorder)]. Computing, 1(3):233–255, September 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ockert:1966:VPG

[55] R. Ockert. Ein Verfahren zur Primfaktorzerlegung grosser Zahlen mit Hilfe binärer quadratischer Formen. (German) [A method for prime factorization of large numbers with the aid of binary quadratic forms]. Computing, 1 (3):256–263, September 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Braess:1966:DMG

[56] Dietrich Braess. Über Dämpfung bei Minimalisierungsverfahren. (German) [On damping in minimization methods]. Computing, 1(3):264–272, September 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hadeler:1966:MGJ

[57] K.-P. Hadeler. Über Matrixnormen, Gerschgorinkreise und Jacobirotation. (German) [On matrix norms, Gershgorin disks, and Jacobi rotations]. Computing, 1(3):273–280, September 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Nake:1966:ZAO

[58] F. Nake. Zertifikat zu Algorithmus 2 ornach. (German) [Certification of algorithm 2]. Computing, 1(3):281, September 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [25].

Hadeler:1966:T

[59] K.-P. Hadeler. Mitteilung/Note: Tagungsbericht. (German) [Conference report]. Computing, 1(3):282, September 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fasol:1966:BRF

[60] K. H. Fasol. Book review: F. H. Effertz und F. Kolberg, Einführung in die Dynamik selbsttätiger Regelungssysteme. Computing, 1(3):283, September 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Roos:1966:FBH

[61] P. Roos. Book review: E. F. Bechenbach (Herausgeber), Applied Combinatorial Mathematics. Computing, 1 (3):283, September 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Thumb:1966:BRG

[62] N. Thumb. Book review: A. S. Goldman, maintainability. Computing, 1 (3):284, September 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Scholz:1966:BRD

[63] H. Scholz. Book report: D. Laugwitz, Ingenieurmathematik, Bank I, II und III (B. I. Hochschultaschenbücher: Nr. 59, 60 und 61). Computing, 1 (3):285, September 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1966:BRW

[64] H. Kremser. Book report: W. Prager, Introduction to FORTRAN. Programming and Numerical Methods. Computing, 1(3):286, September 1966. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fasol:1966:BRH

[65] K. H. Fasol. Book report: H. Kindler and H. Buchta und H. H. Wilfert, Aufgabensammlung zur Regelungstechnik. Computing, 1(3):286, September 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Meinardus:1966:PBJ

[66] G. Meinardus. Book review: P. L. Butzer und J. Korevaar (Herausgeber), On Approximation Theory – Über Approximationstheorie (International Schriftenreihe ur Numerischen Mathematik: Band 5). Computing, 1(3):287, September 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brusberg:1966:CBO

[67] H. Brusberg. Book review: A Comprehensive Bibliography on Operations Research 1957–1958 (Publications in Operations Research: No. 8). Computing, 1(3):288, September 1966. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brusberg:1966:JHB

[68] H. Brusberg. Book review: J. H. Batchelor, Operations Research, Vol. 1 and 2, 1959. Computing, 1(3):289, September 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1966:MAS

[69] E. Bukovics. Book review: M. Abramowitz und I. A. Stegun (Herausgeber), Handbook of Mathematical Functions With Formulas, Graphs and Mathematical Tables. Computing, 1(3):289, September 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brusberg:1966:BRB

[70] H. Brusberg. Book review: B. V. Dean (Herausgeber), Operations in Research and Development. Computing, 1(3):290, September 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eberl:1966:JBE

[71] W. Eberl. Book review: N. T. J. Bailey, The Elements of Stochastic Processes with Applications to the Natural Sciences. Computing, 1(3):290, September 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Baron:1966:BRV

[72] G. Baron. Book report: L. V. Kantorovich und G. P. Akilov, Functional Analysis in Normed Spaces. Computing, 1(3):291, September 1966. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bijleveld:1966:ARD

[73] R. Albrecht. Book review: W. J. Bijleveld, Automatic Rading of Digits. Computing, 1(3):291, September 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eberl:1966:BRP

[74] W. Eberl. Book report: P. Rechenberg, Grundzüge digitaler Rechenautomaten. Computing, 1(3):292, September 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Meinardus:1966:BRB

[75] M. Meinardus. Book report: S. B. Norkin (Herausgeber), The Elements of Computational Mathematics. Computing, 1(3):292, September 1966. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hajek:1966:GMA

[76] P. Hajek, I. Havel, and M. Chyth. The GUHA method of automatic hypotheses determination. Computing, 1 (4):293–308, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bundke:1966:NRL

[77] W. Bundke. Eine neue Rekursionsformel für die Legendre-Polynome und ihre Ableitungen für maschinentechnische Anwendungen. (German) [A new recursion formula for Legendre polynomials and its application to machinetechnical use]. Computing, 1(4): 309–315, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Apostolatos:1966:VBG

[78] N. Apostolatos. Verallgemeinerung des Begriffes einer gewöhnlichen Differentialgleichung n-ter Ordnung in Komplexen sowie Existenz- und Eindeutigkeitsaussagen. (German) [Generalization of the concept of an ordinary differential equation of nth order in complexes, as well as existence and uniqueness statements]. Computing, 1 (4):316–326, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Feichtinger:1966:SRR

[79] G. Feichtinger. Some results on the relation between automat and their automorphism groups. Computing, 1(4): 327–340, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ress:1966:DSV

[80] H. Reß. Digitale Simulation des Verkehrsablaufes auf einer landstraße. (German) [Digital simulation of highway traffic flow]. Computing, 1(4): 341–353, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Theissen:1966:AUBa

[81] E. Theissen. Automatische Umwandlung von BELL-Programmen in ALGOL-Programme. (German) [Automatic conversion from BELL programs to ALGOL programs]. Computing, 1 (4):354–357, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Theissen:1966:AUBb

[82] E. Theissen. Algorithmus 3: Umwandlung von BELL-Programmen in ALGOL-Programme. (German) [Algorithm 3: Conversion from BELL programs to ALGOL programs]. Computing, 1(4):358, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Anonymous:1966:MNI

[83] Anonymous. Mitteilung/Note: International Congress for Project Planning by Network Analysis (INTERNET). Computing, 1(4):368, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zemanek:1966:MNN

[84] H. Zemanek. Mitteilung/Note: NTG-Diskussionsitzung: Teilnehmer-Rechenanlagen. (German) [Note: Discussion session: Participants’ computers]. Computing, 1(4):368, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dremser:1966:BRH

[85] H. Dremser. Book review: H. Brusberg, Der Entwicklungsstand der Unternehmensforschung mit besonderer Berücksichtigung der Bundesrepublik Deutschland. Computing, 1(4):370, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Scholz:1966:BRP

[86] H. Scholz. Book review: P. Calligaert, Principles of Computation. Computing, 1(4):370, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brusberg:1966:BRD

[87] H. Brusberg. Book review: D. D. Mc-Cracken, A Guide to FORTRAN IV. Computing, 1(4):371, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1966:BRH

[88] H. Kremser. Book review: H. E. Fettis and J. C. Caslin, Tables of Elliptic Integrals of the First, Second and Third Kind. Computing, 1(4):371, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Adler:1966:BRC

[89] H. Adler. Book review: C. P. Gilbert, The Design and Use of Electronic Analogue Computers. Computing, 1 (4):372, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Heinhold:1966:BRG

[90] J. Heinhold. Book review: I. Guttman und S. Wilks, Introductory Engineering Statistics. Computing, 1(4):372, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1966:BRJa

[91] H. J. Stetter. Book review: J. Gruenberger und G. Jaffray, Problems for Computer Solutions. Computing, 1 (4):372, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1966:BRK

[92] H. Kremser. Book review: K.-H. Kaps, Die Bedeutung der lagerhaltung für die Produktionsplanung in Industriebetrieben. Computing, 1(4):373, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1966:BRJb

[93] H. J. Stetter. Book review: J. Hawgood, Numerical Methods in ALGOL. Computing, 1(4):373, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1966:BRJ

[94] H. Kremser. Book report: J. Mandel, The Statistical Analysis of Experimental Data. Computing, 1(4):374, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Roos:1966:BRHc

[95] P. Roos. Book report: H. Meschkowski, Einführung in die moderne Mathematik. Computing, 1(4):374, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmid:1966:BRS

[96] E. Schmid. Book report: Shan S. Kuo, Numerical Methods and Computers. Computing, 1(4):374, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1966:BRDb

[97] H. Kremser. Book report: D. Müller, Programmierung elektronischer Rechenanlagen. Computing, 1 (4):375, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Oliva:1966:BRJ

[98] P. Oliva. Book report: S. John Murphy, Elektronische Ziffernrechner. Computing, 1(4):375, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fischer:1966:BRK

[99] K. Fischer. Book report: K. Nicot, Elementary Programming and Algol. Computing, 1(4):376, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Heinhold:1966:BRV

[100] J. Heinhold. Book report: A. V. Shileiko, Differential analysers. Computing, 1(4):377, December 1966. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1966:BRDc

[101] H. Kremser. Book report: D. B. Owen, Handbook of Statistical Tables. Computing, 1(4):377, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1966:BRHb

[102] E. Bukovics. Book review: H. Unger, Elektronische Datenverarbeitungsanlagen und Automatentheorie. Computing, 1(4):378, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1966:BRG

[103] H. Kremser. Book report: I. G. Wilson und M. E. Wilson, Information, Computers, and System Design. Computing, 1(4):378, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Meissl:1966:BRO

[104] P. Meissl. Book report: Operations Research. The Journal of the Operations Research Society of America. Vol 12, Number 6. Computing, 1 (4):379, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brusberg:1966:BDT

[105] H. Brusberg. Book review: Betriebssysteme, für Datenverarbeitungsanlagen. Tagung vom 29, Juni 1965 in Haus der Technik, Essen. Computing, 1(4):380, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1967:PIG

[106] E. Bukovics and W. Knödel. Prof. Inzinger zum 60, Geburtstag. (German) [Prof. Inzinger’s 60th birthday]. Computing, 2(1):1–4, March 1967. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bellman:1967:DPI

[107] R. Bellman, H. Gagiwada, and R. Kalaba. Dynamic programming and an inverse problem in transport theory. Computing, 2(1):5–16, March 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Apostolatos:1967:KRN

[108] N. Apostolatos and U. Kulisch. Über die Konvergenz des Relaxtionverfahrens bei nicht-negativen und diagonaldominanten Matrizen. (German) [On the convergence of relation methods for nonnegative and diagonallydominant matrices]. Computing, 2 (1):17–24, March 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Allgower:1967:FAA

[109] E. Allgower and R. Guenther. A functional analytic approach to the numerical solution of nonlinear elliptic equations. Computing, 2(1):25–33, March 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Werner:1967:BNR

[110] H. Werner. Die Bedeutung der Normalität bei rationaler Tschebyscheff Approximation. (German) [The meaning of normality for rational Chebyshev approximation]. Computing, 2(1):35– 52, March 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Burkhardt:1967:GTP

[111] W. H. Burkhardt. Generating test programs from syntax. Computing, 2(1):53–73, March 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Motzfeld:1967:BKA

[112] H. Motzfeld and H. Kietzer. Beitrag zur konformen Abbildung einfach zusammenhängender Gebiete mit Hilfe einer elektrischen Analogie. Lösung der Integralgleichung von Theodorsen. (German) [Contribution to the conformal mapping of simply-connected regions with the aid of an electrical analogy. Solution of Theodore’s integral equations]. Computing, 2(1):74–78, March 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Collatz:1967:NNA

[113] L. Collatz and G. Neinardus. Mitteilung/Note: Numerische Analysis, ins besondere Approximationstheorie. (German) [Numerical analysis, especially approximation theory]. Computing, 2(1):79–80, March 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fleischmann:1967:BRAa

[114] A. Fleischmann. Book review: A. Adam, Systematische Datenverarbeitung bei der Auswertung von Versuchsund Beobachtungsergebnissen. Computing, 2(1):81, March 1967. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Troch:1967:BRMa

[115] I. Troch. Book review: M. A. Aiserman and F. R. Gantmacher, Die absolute Stabilität von Regelsystemen. Computing, 2(1):81, March 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Scacel:1967:BRJ

[116] F. Scacel. Book report: A. J. Monroe, Digital Processes for Sampled Data Systems. Computing, 2(1):82, March 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Scholz:1967:BRB

[117] H. Scholz. Book report: B. E. Margulis, Systems of Linear Equations. Computing, 2(1):82, March 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fieber:1967:BRH

[118] H. Fieber. Book report: H. Meschkowski, Hilbertsche Räume mit Kernfunktionen. Computing, 2(1):83, March 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hornich:1967:BRR

[119] H. Hornich. Book report: R. Sauer, Ingenieurmathematik, Bank II: Differentialgleichungen und Funktionentheorie. Computing, 2(1):83, March 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hoffman:1967:BRP

[120] H.-J. Hoffman. Book report: Philip M. Sherman, Programming and Coding the IBM 709-, 7090-, 7094-Computers. Computing, 2(1):84, March 1967. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Knodel:1967:BRR

[121] W. Knödel. Book review: R. Baumann (Herausgeber), Manual der ALCOR-Gruppe. Computing, 2(1):84, March 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gunzenhauser:1967:BRU

[122] R. Gunzenhäuser. Book report: L. Uhr (Herausgeber), Pattern Recognition. Computing, 2(1):85, March 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zemanek:1967:BRA

[123] H. Zemanek. Book review: A. Adam, (Herausgeber), Messen und Regeln in der Betriebswirtschaft. Computing, 2(1):85, March 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schwarzenberger:1967:BRJ

[124] F. Schwarzenberger. Book report: J. Wosnik (Herausgeber), Fernwirkteehnik (Nachrichtentechnische Fachberichte, Band 30). Computing, 2(1): 86, March 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1967:BRBa

[125] H. J. Stetter. Book review: L. B. Rall, Error in Digital Computation, Vol. I. Computing, 2(1):86, March 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eberl:1967:BRDb

[126] W. Eberl. Book report: D. Teichroew, An Introduction to the Management Science. Computing, 2(1):87, March 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1967:BRF

[127] H. J. Stetter. Book report: L. F. Richardson, Weather Prediction by Numerical Process. Computing, 2(1): 87, March 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Apostolatos:1967:GM

[128] N. Apostolatos and U. Kulisch. Grundlagen Einer Maschinenintervallarithmetik. (German) [Foundations of a machine interval arithmetic]. Computing, 2(2):89–104, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Filippi:1967:ERK

[129] S. Filippi and C. Stimberg. Explizite Runge–Kutta-Verfahren für partielle hyperbolische Differentialgleichungen. (German) [Explicit Runge– Kutta method for hyperbolic partial differential equations]. Computing, 2(2):105–119, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Troch:1967:UVAa

[130] I. Troch. Untersuchungen zum Verhalten von algebraischen Nichtlinearitäten in Regelsystemen. I. (German) [Investigation of the behavior of algebraic nonlinearities in rule systems. I]. Computing, 2(2):120–138, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hotz:1967:QAG

[131] G. Hotz. Quasilineare Automaten. (German) [Quasilinear automata]. Computing, 2(2):139–152, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wing:1967:EMN

[132] O. Wing. An efficient method of numerical inversion of Laplace transforms. Computing, 2(2):153–166, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wallach:1967:AES

[133] Y. Wallach. Algorithmus 4: Entwurf sequentieller Schaltkreise nach der Methode Huffman-Mealy. (German) [Algorithm 4: The design of sequential circuits by the Huffman-Mealy method]. Computing, 2(2):167–168, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Baron:1967:BRM

[134] G. Baron. Book review: M. Abramowitz and I. A. Stegun, Handbook of Mathematical Functions. Computing, 2(2):169, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1967:BRBb

[135] H. J. Stetter. Book report: I. Babuška, M. Präger und Vitäsek, Numerical Processes in Differential Equations. Computing, 2(2):169, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Troch:1967:BRW

[136] I. Troch. Book review: W. Ammon, Schaltungen der Analogrechentechnik. Computing, 2(2):169, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eberl:1967:BRKa

[137] W. Eberl. Book report: K. A. Brownlee, Statistical Theory and Methodology in Science and Engineering. Computing, 2(2):170, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fasol:1967:BRG

[138] K. H. Fasol. Book report: S. A. Ginzburg, I. Ya. Lekhtman und V. S. Malov, Fundamentals of Automation and Remote Control. Computing, 2(2):171, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Heinrich:1967:BRP

[139] H. Heinrich. Book report: P. Henrici, Elements of Numerical Analysis. Computing, 2(2):171, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fleischmann:1967:BRK

[140] A. Fleischmann. Book report: K. E. Iverson, A Programming Language. Computing, 2(2):172, June 1967. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmid:1967:BRK

[141] E. Schmid. Book report: I. O. Kerner und G. Zielke, Einführung in die algorithmische Sprache ALGOL. Computing, 2(2):172, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Skacel:1967:BRRb

[142] F. Skacel. Book report: R. W. P. King, Fundamental Electromagnetic Theory. Computing, 2(2):172, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Baron:1967:BRFa

[143] G. Baron. Book report: F. Klein, On Riemann’s Theory of Algebraic Functions and their Integrals. Computing, 2(2):173, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Heinrich:1967:BRK

[144] H. Heinrich. Book report: A. S. Kronrod, Nodes and Weights of Quadrature Formulas. Computing, 2(2):173, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fieber:1967:BRM

[145] H. Fieber. Book report: N. I. Muschelischwili, Singuläre Integralgleichungen. Computing, 2(2):174, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kudielka:1967:BRW

[146] V. Kudielka. Book report: W. Michel, Der logische Entwurf von Schaltkreisen. Computing, 2(2):174, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eberl:1967:BRD

[147] W. Eberl. Book report: D. B. Owen, Handbook of Statistical Tables. Computing, 2(2):175, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eberl:1967:BRJ

[148] W. Eberl. Book report: J. von Neumann und O. Morgenstern, Theory of Games and Economic Behavior. Computing, 2(2):175, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eberl:1967:BRV

[149] W. Eberl. Book report: V. S. Pugachev, Theory of Rndom functions in its Application to Control Problems. Computing, 2(2):176, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1967:BRH

[150] H. J. Stetter. Book report: H. H. Rosenbrock, Computational Techniques for Chemical Engineers. Computing, 2(2):176, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Junginger:1967:BRP

[151] W. Junginger. Book report: P. Stahlknecht, Operations Research, Teil I: Input-Output-Modelle – Optimierungsverfahren. Computing, 2(2): 177, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Scholz:1967:BRY

[152] H. Scholz. Book report: N. Ya Vilenkin, Successive Approximation. Computing, 2(2):177, June 1967. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1967:BRL

[153] K. Kremser. Book report: A. Lavi und T. P. Vogl (herausgeber), Recent Advance in Optimization Techniques. Computing, 2(2):178, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmid:1967:BRR

[154] E. Schmid. Book report: R. Zurmühl, Praktische Mathematik für Ingenieure und Physiker. Computing, 2(2):178, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Scholz:1967:BRJb

[155] H. Scholz. Book report: J. H. Wilkinson, Rounding Errors in Algebraic Processes. Computing, 2(2):178, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Collatz:1967:BRK

[156] L. Collatz. Book report: K. Schütte (herausgeber), Index mathematischer Tafelwerke und Tabellen aus allen Gebieten der Naturwissenschaften. Computing, 2(2):179, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gunzenhauser:1967:BRF

[157] R. Gunzenhäuser. Book report: L. Fox (Herausgeber) Advances in Programming and Non-Numerical Computations. Computing, 2(2):180, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zemanek:1967:BRHb

[158] H. Zemanek. Book report: H. Marchl (Herausgeber), Beiträge zur Sprachkunde und Informationsverarbeitung. Computing, 2(2):180, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Apostolatos:1967:AEI

[159] N. Apostolatos and U. Kulisch. Approximation der erweiterten Intervallarithmetik durch die einfache Maschinenintervallarithmetik. (German) [Approximation of extended interval arithmetic with simple machine interval arithmetic]. Computing, 2(3):181–194, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Apostolatos:1967:ENG

[160] N. Apostolatos, U. Kulisch, and K. Nickel. Ein Einschliessungsverfahren für Nullstellen. (German) [An interval method for zeros]. Computing, 2(3):195–201, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Troch:1967:UVAb

[161] I. Troch. Untersuchungen zum Verhalten von algebraischen Nichtlinearitäten in Regelsystemen. II. (German) [Investigation of the behavior of algebraic nonlinearities in rule systems. II]. Computing, 2(3):202–231, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Nickel:1967:VBE

[162] K. Nickel. Die vollautomatische Berechnung einer einfachen Nullstelle von F (T) = 0 einschließlich einer Fehlerabschätzung. (German) [The completely automatic calculation of a single zero of F (T) = 0 including an error bound]. Computing, 2(3): 232–245, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pittnauer:1967:MBG

[163] F. Pittnauer. Eine Methode zur Bestimmung günstiger Konvergenzfaktoren für asymptotische Entwicklungen. (German) [A method for determination of beneficial convergence factors for asymptotic series]. Computing, 2 (3):246–256, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1967:ASN

[164] F. Stetter. Ableitungsfreie Schranken für die numerische Integration. (German) [Derivative-free bounds for numerical integration]. Computing, 2(3): 257–262, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Veenker:1967:BP

[165] G. Veenker. Beweisalgorithmen für die Prädikatenlogik. (German) [Proof algorithms for predicate logic]. Computing, 2(3):263–283, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Nickel:1967:ANP

[166] K. Nickel. Algorithmus 5: Die Nullsstellen eines Polynoms. (German) [Algorithm 5: Polynomial zeroes]. Computing, 2(3):284–288, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See certification [167, 168].

Fischer:1967:ZAP

[167] K. Fischer. I. Zertifikat zu Algorithmus 5, [Polroot, K. Nickel, Computing 2 (1967), S. 284–288]. (German) [I. Certification of algorithm 5]. Computing, 2(3):289, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [166].

Nickel:1967:ZA

[168] K. Nickel. 2. Zertifikat zu Algorithmus 5. (German) [2. Certification of algorithm 5]. Computing, 2(3):290, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [166].

Bukovics:1967:BRM

[169] E. Bukovics. Book report: M. Athans und F. Falb, Optimal Control. Computing, 2(3):291, September 1967. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Skacel:1967:BRM

[170] F. Skacel. Book review: M. A. Arrib, Brains, Machines and Mathematics. Computing, 2(3):291, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Baron:1967:BRHa

[171] G. Baron. Book report: H. J. Bartsch, Mathematische Formeln. Computing, 2(3):292, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schnabl:1967:BRK

[172] R. Schnabl. Book review: N. K. Bary, A Treatise on Trigonometric Series. Computing, 2(3):292, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bednarczyk:1967:BRJ

[173] H. Bednarczyk. Book report: A. J. Chintschin, Mathematische Grundlagen der statistischen Mechanik. Computing, 2(3):293, September 1967. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1967:BRR

[174] H. Kremser. Book report: N. R. Draper und H. Smith, Applied Regression Analysis. Computing, 2(3):293, September 1967. CODEN CMPTA2.

ISSN 0010-485X (print), 1436-5057 (electronic).

Zemanek:1967:BRH

[175] H. Zemanek. Book report: H. Billing (Herausgeber), Lornende Automaten. Computing, 2(3):293, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Skacel:1967:BRD

[176] F. Skacel. Book report: E. Durand, Solutions numérique des équations algébriques. Computing, 2(3):294, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1967:BRE

[177] H. J. Stetter. Book report: T. Ekman und C. E. Fröberg, Introduction to ALGOL-Programming. Computing, 2(3):294, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Baron:1967:BRB

[178] G. Baron. Book report: B. Higman, Applied Group-Theoretic and Matrix Methods. Computing, 2(3):295, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1967:BRHa

[179] E. Bukovics. Book report: H. Kneser, Funktionentheorie. Computing, 2 (3):295, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fleischmann:1967:BRB

[180] A. Fleischmann. Book report: B. A. Galler, The Language of Computers. Computing, 2(3):295, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eberl:1967:BRKb

[181] W. Eberl. Book report: E. Kreyszig, Statistische Methoden und ihre Anwendungen. Computing, 2(3):296, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Troch:1967:BRG

[182] I. Troch. Book report: G. Maess, Quantitative Verfahren zur Bestimmung periodischer Lösungen autonomer nichtlinearer Differentialgleichungen. Computing, 2(3):296, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Troch:1967:BRMb

[183] I. Troch. Book report: M. V. Meerov, Structural Synthesis of Hiagh-Accuracy Automatic Control Systems. Computing, 2(3):297, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Baron:1967:BRFb

[184] G. Baron. Book report: F. D. Murnaghan, The Theory of Group Representations. Computing, 2(3):298, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bosse:1967:BRJ

[185] L. Bosse. Book report: J. Pfanzagi, Allgemeine Methodenlehre der Statistik, Teil 1: Elementare Methodenunter besonderer Berücksichtigung der Anwendungen in den Wirtschafts-Sozialwissenschaften. (German) [J. Pfanzagi, General methods of statistics, Part 1: elementary methods with (special) emphasis on applications in the economic- and social sciences]. Computing, 2(3):298, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1967:BRG

[186] E. Bukovics. Book report: G. Polya, Von Lösen mathematischer Aufgaben. Einsicht und Entdocken. Lernen und Lehren. Computing, 2(3):299, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Skacel:1967:BRP

[187] F. Skacel. Book report: L. S. Pontrjagin, V. G. Boltjanskij, R. V. Gamkrelidze und E. F. Miscenko, Mathematische Theorie optimaler Prozesse. Computing, 2(3):299, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Neuhold:1967:BRK

[188] E. Neuhold. Book report: K. R. Richards, Electronic Digital Systems. Computing, 2(3):300, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Holecek:1967:HSM

[189] K. Holecek. H. Smers, Das machinelle Lochkartenverfahren. Computing, 2 (3):301, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lehner:1967:BRK

[190] L. Lehner. Book report: K. Rottmann, Mathematische Formelsammlung. Computing, 2(3):301, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lehner:1967:BRM

[191] L. Lehner. Book report: Mathematische Funktionstafeln. Computing, 2 (3):301, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmetterer:1967:BRM

[192] L. Schmetterer. Book report: M. Shubik (Herausgeber), Game Theory and Related Approaches to Social Behavior. Computing, 2(3):301, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Junginger:1967:BRF

[193] W. Junginger. Book report: F. Stuart. Introductory Computer Programming. Computing, 2(3):302, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1967:BRJ

[194] H. Kremser. Book report: J. S. Wentzel, Elemente der dynamisehen Programmierung. Computing, 2 (3):302, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hauer:1967:BRF

[195] F. Hauer. Book report: F. Wolf, Ausgleichsrechnung nach der Methode der kleinsten Quadrate. Computing, 2(3):303, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Skacel:1967:BRW

[196] F. Skacel. Book report: S. Winograd und J. D. Cowan, Reliable Computation in the Prescence of Noise. Computing, 2(3):303, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lichtblau:1967:BRG

[197] K. Lichtblau. Book report: G. Friedrichs (Herausgeber), Automation, Risiko und Chance. Computing, 2 (3):304, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Neuhold:1967:BRF

[198] E. Neuhold. Book report: F. Wolf und A. Schmitt, Modelle lernender Automaten. Computing, 2(3):304, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1967:BRM

[199] H. J. Stetter. Book report: M. I. Zhurina und L. N. Karmazina, Tables and Formulae for the Spherical Functions p^{m} − \frac{1}{2} + τ(Z). Computing, 2(3):304, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1967:BRHb

[200] E. Bukovics. Book report: H. Matazke (Herausgeber), Berichte des III. Internationalen Kolloquiums über Anwendungen der Mathematik in den Ingenieurwissenschaften zu Weimar. Anwendungen elektronischer Rechenanlagen im Bauwesen. Computing, 2 (3):305, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eberl:1967:BRE

[201] W. Eberl. Book report: Einführung in die automatische Datenverabeitung,

1. und 2. Teil. Computing, 2(3):306, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Walk:1967:BRC

[202] K. Walk. Book report: 2. Colloquium über Schaltkreis- und Schaltwerk-Theorie. Computing, 2(3):306, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Skacel:1967:BRL

[203] F. Skacel. Book report: A. A. Ljapunow (herausgeber), Probleme der Kybernetik, Band 2. Computing, 2 (3):307, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1967:BRO

[204] H. J. Stetter. Book report: A. Opler (Herausgeber), Programming the IBM System/360 by the Staff of the Computer Usage Company. Computing, 2(3):307, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Skacel:1967:BRR

[205] F. Skacel. Book review: R. L. Ackoff (Herausgeber), Progress in Operations Research. Computing, 2(3):308, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Braess:1967:AEG

[206] Dietrich Braess. Approximation mit exponentialsummen. (German) [Approximation with exponential sums]. Computing, 2(4):309–321, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Filippi:1967:UTD

[207] S. Filippi and C. Stimberg. Untersuchungen über den totalen Diskretisierungsfehler bei RUNGE-KUTTA-Verfahren für partielle, hyperbolische Differentialgleichungen. (German) [Investigation of the total discretization errors in Runge–Kutta methods for partial differential equations]. Computing, 2 (4):322–331, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Baron:1967:BAS

[208] G. Baron. Bestimmung der Automorphismengruppe spezieller Klassen von Graphen. (German) [Determination of the automorphism group special classes for graphs]. Computing, 2 (4):332–335, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Vachenauer:1967:ELA

[209] P. Vachenauer. Ein Estimationsproblem der linearen Approximationstheorie beim Ausgleich statistischer Messwerte. (German) [An estimation problem of linear approximation theory by least squares statistical mass values]. Computing, 2(4):336–352, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Deussen:1967:SRS

[210] P. Deussen. Some results on the sets of congruence relations in a finite strongly connected automation. Computing, 2 (4):353–367, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fleischmann:1967:BRR

[211] A. Fleischmann. Book review: R. A.- R Amer und H. R. Schwarz, Contributions to the Approximation Problem of Electrical Filters. Computing, 2(4):368, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Scholz:1967:BRV

[212] H. Scholz. Book review: V. S., Aizenshtadt und V. I. Krylov and A. S. Metel’skii, Tables of Laguerre Polynomials and Functions. Computing, 2(4):368, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1967:BRRa

[213] H. J. Stetter. Book review: R. V. Andree, Computer Programming and Related Mathematics. Computing, 2 (4):368, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fleischmann:1967:BRAb

[214] A. Fleischmann. Book report: A. Angermann. Computing, 2(4):369, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fischer:1967:BRC

[215] K. Fischer. Book report: N. L. Collins und D. Michie, Machine Intelligence 1. Computing, 2(4):370, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Scholz:1967:BRG

[216] H. Scholz. Book report: G. Böhme, Mathematik, Vorlesungen für ingenieurschulen. Computing, 2(4):370, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Skacel:1967:BRV

[217] F. Skacel. Book report: L. V. Bewley, Two-Dimensional Fields in Electrical Engineering. Computing, 2(4):370, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fleischmann:1967:BRF

[218] A. Fleischmann. Book report: F. Erwe, Differential- und Integralrechnung I und II. Computing, 2(4):371, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Scholz:1967:BRR

[219] H. Scholz. Book report: R. Courant und H. Robbins, Was ist Mathematik?

Computing, 2(4):371, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Baron:1967:BRHb

[220] G. Baron. Book report: H. E. Fettis und J. C. Caslin, Ten Place Tables of the Jacobian Elliptic functions. Computing, 2(4):372, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brusberg:1967:BRW

[221] H. Brusberg. Book report: W. Frank, Einführung in die Optimierungsrechnung. Computing, 2(4):372, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hauer:1967:BRG

[222] F. Hauer. Book report: E. Gigas, Physikalisch-Geodätische Meßverfahren (Neue Wege der Dreiecksund Höhenmessung). Computing, 2(4):372, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zemanek:1967:BRT

[223] H. Zemanek. Book report: Th. Fromme, Der Äquivalenzkalkül. Computing, 2(4):372, December 1967. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Junginger:1967:BRH

[224] W. Junginger. Book report: H. J. Genrich, Die automatische Aufstellung von Schulstundenplänen auf relationstheoretischer Grundlage. Computing, 2(4):373, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovices:1967:BRM

[225] E. Bukovices. Book report: M. R. Hestenes, Calculus of Variations and Optimal Control Theory. Computing, 2(4):374, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fieber:1967:BRG

[226] H. Fieber. Book report: E. A. Guillemin, Mathematische Methoden des ingenieurs. Computing, 2(4):374, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Scholz:1967:BRJ

[227] H. Scholz. Book report: E. J. Jury, Theory and Application of the z-Transform Method. Computing, 2 (4):374, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1967:BRT

[228] H. J. Stetter. Book report: Th. H. Naylor, J. L. Balintfy, D. S. Burdick und Kong Chu, Computer Simulation Techniques. Computing, 2(4):375, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Roos:1967:BRR

[229] P. Roos. Book report: R. R. Korfhage, Logic and Algorithms. With applications to the computer and information sciences. Computing, 2(4):376, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Junginger:1967:BRK

[230] W. Junginger. Book report: K. C. Parton, The Digital Computer. Computing, 2(4):377, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Troch:1967:BRMc

[231] I. Troch. Book report: S. Miller, An Introduction to the Calculus of Finite Differences and Difference Equations. Computing, 2(4):377, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Baron:1967:BRJ

[232] G. Baron. Book report: J. Piehler, Einführung in die dynamische Optimierung. Computing, 2(4):378, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1967:BRRb

[233] H. J. Stetter. Book review: A. Ralston und H. S. Wilf (Herausgeber), Mathematical Methods for Digital Computers, Volume 2. Computing, 2(4):378, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Baron:1967:BRU

[234] G. Baron. Book report: U. Pape, Strukturnumwandlungen alphanumerischer Streckenkomplexe und ihre Ausführung mit digitalen Rechenanlagen. Computing, 2(4):379, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fischer:1967:BRH

[235] K. Fischer. Book report: H. J. Schneider und D. Jurksen, Programmierung von Datenverarbeitungsanlagen. Computing, 2(4):379, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kraus:1967:BRK

[236] G. Kraus. Book report: K. Steinbuch und W. Reupprecht, Nachriehtentechnik. Eine einführende Darstellung. Computing, 2(4):380, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Parkus:1967:BRH

[237] H. Parkus. Book report: H. Görtler (Herausgeber), Applied Mechanics. proceedings of the eleventh international congress of Applied Mechanics. Computing, 2(4):380, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1967:BRHc

[238] E. Bukovics. Book report: H. Behnke, G. Bertram und R. Sauer (Herausgeber), Grundzüge der Mathematik für Lehrer an Gymnasiem sowie für Mathematiker in Industrie und Wirtschaft. Band IV: Praktische Methoden und Anwendungen der Mathematik (Geometrie und Statistik). (German) [H. Behnke, G. Bertram, and R. Sauer, editors. basics of mathematics for teachers at gymnasium and mathematicians in industry and business. Volume IV: Practical methods and applications of mathematics (geometry and statistics)]. Computing, 2(4):396, December 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Feichtinger:1968:QNE

[239] G. Feichtinger. Der Quotientenautomat nach einem direkten Faktor einer Automorphismengruppe. (German) [The quotient automaton for a direct factor of an automorphism group]. Computing, 3(1):1–8, March 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gabler:1968:VWA

[240] W. Gabler. Verallgemeinerte Wurzelortskurven. Automatische digitale Berechnung und Aufzeichnung. (German) [Generalized root location curves. automatic digital computation and recording]. Computing, 3(1):9–21, March 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kussmaul:1968:FNV

[241] R. Kußmaul and P. Werner. Fehlerabschätzung für ein numerisches Verfahren zur Auflösung linearer Integralgleichungen mit schwachsingulären Kernen. (German) [Error estimates for a numerical method for solving linear integral equations with weakly singular kernels]. Computing, 3(1):22–46, March 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Nickel:1968:QF

[242] K. Nickel. Quadraturverfahren mit Fehlerschranken. (German) [Quadrature method with error bounds]. Computing, 3(1):47–64, March 1968. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Levert:1968:MSD

[243] C. Levert and J. van Galen. Mathematical and statistical details on the simulation of Markoff-type stochastic processes on an electronic computer. Computing, 3(1):65–73, March 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zielke:1968:AAI

[244] G. Zielke. Algorithmus 6: Änderung der Iversen einer symmetrischen Matrix bei symmetrischer Zeilen- un Spaltenänderung. (German) [Algorithm 6: Iversen interchanges of a symmetric matrix with symmetric row and column interchanges]. Computing, 3(1):76–77, March 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Collatz:1968:NAO

[245] L. Collatz and W. Wetterling. Mitteilungen: Arbeitstagung. Optimierungsaufgaben, 11.–16. Juni 1967, Oberwolfach. (German) [Notes: Workshop on Optimization Problems, 11– 16 June 1967, Oberwolfach]. Computing, 3(1):78–79, March 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Albrecht:1968:BRM

[246] R. Albrecht. Book review: M. P. Barnett, Computer Typesetting. Experiments and Prospects. Computing, 3(1):80, March 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1968:BRGa

[247] H. J. Stetter. Book report: G. Forsythe and C. B. Moler, Computer Solution of Linear Algebraic Systems. Computing, 3(1):80, March 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Thumb:1968:BRHa

[248] N. Thumb. Book report: S. Hellfors, Management-Datenverarbeitung-Operations Research. Die Auswirkung moderner Datenverarbeitungstechniken auf die betriebliche Leitungstätigkeit. Computing, 3(1):81, March 1968. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eberl:1968:BRB

[249] W. Eberl. Book report: B. Harris (Herausgeber), Spectral Analysis of Time Series. Computing, 3(1):82, March 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Thumb:1968:BRI

[250] N. Thumb. Book report: I. L. Illetschko, Transport-Betriebswirtschaftslehre. Computing, 3(1):82, March 1968. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schnabl:1968:BRJ

[251] R. Schnabl. Book report: N. L. Johnson and F. C. Leone, Statistics and Experimental Design: In Engineering and Physical Sciences. Computing, 3(1):83, March 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Thumb:1968:BRHb

[252] N. Thumb. Book report: H. Kaufmann, Informations-Verarbeitung und Automatisierung. Einführende Betrachtungen zum Problem der Automatisierung in Büro und Produktion. Computing, 3(1):83, March 1968. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Nake:1968:BRR

[253] F. Nake. Book report: T. R. McCalla, Introduction to Numerical Methods and FORTRAN Programming. Computing, 3(1):84, March 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Skacel:1968:BRL

[254] F. Skacel. Book report: A. A. Ljapunow (Herausgeber), Probleme der Kybernetik. Computing, 3(1):84, March 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1968:BRGb

[255] H. J. Stetter. Bool Review: G. Paulin, Kleines Lexikon der Rechentechnik und Datenverarbeitung. Computing, 3(1):85, March 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Thumb:1968:BRK

[256] N. Thumb. Book report: K. L. Pedell, Elektronische Informationsverabeitung und Fertigungssteuerung. Computing, 3(1):85, March 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Scholz:1968:BRG

[257] H. Scholz. Book review: G. R. Rich, Hydraulie Transients. Computing, 3(1):86, March 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1968:BRRa

[258] H. J. Stetter. Book review: L. Robin (Herausgeber), Tables Numériques. Des Fonctions associées de Legendre. Computing, 3(1):86, March 1968. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Troch:1968:BMV

[259] I. Troch. Book Mark: V. V. Solodovnikow (Herausgeber), Automatic Conrol and Computer Engineering. Computing, 3(1):86, March 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Skacel:1968:BRB

[260] F. Skacel. Book report: B. Bopp, und J. Meixner, Sommerfeld: Vorlesungen über theoretische Physik. Elektrodynamik. Computing, 3(1):87, March 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eberl:1968:BRTa

[261] W. Eberl. Book review: L. Takács, Combinatorial Methods in the Theory of Stochastic Processes. Computing, 3(1):88, March 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eberl:1968:BRTb

[262] W. Eberl. Book review: L. Takács, Stochasstische I’rozesse, Aufgaben und Lösungen. Computing, 3(1):88, March 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Roos:1968:BRT

[263] P. Roos. Book review: A. Tarski, Einführung in die mathematische Logik. Computing, 3(1):88, March 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hahn:1968:BRJ

[264] W. Hahn. Book review: J. S. Zypkin, Theorie der linearen Implussysteme. Computing, 3(1):89, March 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Scholz:1968:BRM

[265] H. Scholz. Book review: M. Walker, The Schwarz–Christoffel Transformation and its Applications. Computing, 3(1):89, March 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Skacel:1968:BRG

[266] F. Skacel. Book review: G. J. Thaler and M. P. Pastel, Analysis and Design of Nonlinear Feedback Control Systems. Computing, 3(1):89, March 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Skacel:1968:BRJ

[267] F. Skacel. Book review: J. Wosnik (Herausgeber), Transistoren bei großer Aussteuerung. Computing, 3(1):90, March 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Baron:1968:BRC

[268] G. Baron. Book review: C. Zwikker, The Advanced Geometry of Plane Curves and Their Applications. Computing, 3(1):91, March 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1968:BRHa

[269] H. J. Stetter. Book report: H. E. Fettis and J. C. Caslin (Herausgeber), Elliptic Functions for Complex Arguments. Computing, 3(1):91, March 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1968:BRHb

[270] H. J. Stetter. Book review: H. Winkler, Anleitung zum praktischen Gebrauch von PL/I. Computing, 3(1):91, March 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1968:BRP

[271] E. Bukovics. Book review: L. S. Pontrjagin, V. G. Boltjanskij, R. V. Gamkrelidze und E. F. Miscenko, Mathematische Theorie optimaler Prozesse. Computing, 3(1):92, March 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Muller:1968:HHP

[272] P. F. Müller and E. Geller. Über die Herstellung von Höhenkarten und perspektivischen Reliefdarstellungen von Funktionen einer komplexen Veränderlichen mittels elektronischer Rechenanlagen. (German) [On the manufacturing of contour maps and perspective-relief representations of functions of a complex variable via electronic computers]. Computing, 3(2): 93–98, June 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wippermann:1968:DST

[273] H.-W. Wippermann. Definition von Schrankenzahlen in Triplex-Algol. (German) [Definition of number bounds in Triplex-Algol]. Computing, 3(2):99–109, June 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schassberger:1968:WZP

[274] R. Schassberger. Ein Wartesystem mit zwei parallelen Warteschlangen. (German) [A Queueing system with two parallel queueing]. Computing, 3(2):110–124, June 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wetterling:1968:LOS

[275] W. Wetterling. Lokal optimale Schranken bei Randwertaufgaben. (German) [Locally optimal bounds for boundary values]. Computing, 3(2):125–130, June 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bellman:1968:NTN

[276] R. Bellman. A new technique for the numerical solution of Fredholm integral equations. Computing, 3(2):131–138, June 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Allgower:1968:NSH

[277] E. Allgower and R. Guenther. On the numerical solution of higher order nonlinear parabolic equations. Computing, 3(2):139–150, June 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See correction [317].

Knodel:1968:AKR

[278] W. Knödel. Algorithmus 7: Kürzeste Rundreise zwischen n Orten (Travelling Salesman Problem). (German) [Algorithm 7: Shortest circuit between n points]. Computing, 3(2):151–156, June 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1968:BRF

[279] H. Kremser. Book review: F. Bates and M. L. Douglas, PL/1 Programming Language/One. Computing, 3(2):157, June 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1968:BRRb

[280] H. J. Stetter. Book review: R. Bellman, R. E. Klaba and J. A. Loskett, Numerical Inversion of the Laplace Transform. Computing, 3(2):157, June 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Junginger:1968:BRWa

[281] W. Junginger. Book report: W. S. Dorn and H. J. Greenberg, Mathematics and Computing with FORTRAN Programming. Computing, 3(2):158, June 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Thumb:1968:BRD

[282] N. Thumb. Book report: A. Diemer, Die automatisierte elektronische Datenverarbeitung und ihre Bedeutung für die Unternchmensleitung. Computing, 3(2):158, June 1968. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fieber:1968:BRG

[283] H. Fieber. Book report: G. E. Franks, Mathematical Modeling in Chemical Engineering. Computing, 3(2):159, June 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Troch:1968:BRP

[284] I. Troch. Book report: P. M. E. M. Van Der Grinten, Stochastische Prozesse in der Meß- und Regelungstechnik. Computing, 3(2):159, June 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1968:BRGa

[285] E. Bukovics. Book report: G. Meinardus, Approximation of Functions: Theory and Numerical Methods. Computing, 3(2):160, June 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1968:BRGb

[286] E. Bukovics. Book review: G. Pölya, Vom Lösen mathematischer Aufgaben. Einsicht und Entdeckung, Lernen und Lehren. Computing, 3(2):160, June 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1968:BRR

[287] H. Kremser. Book report: R. Sauer un I. Szabö (herausgeber), Mathematische Hilfsmittel des Ingenieurs. Computing, 3(2):161, June 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lorenz:1968:BRH

[288] H. Lorenz. Book review: H. R. Rottmann, IBM/360 Modell 20, Programmieren leicht gemacht mit RPG. Computing, 3(2):161, June 1968. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1968:BRJ

[289] E. Bukovics. Book review: J. B. Vernon, Linear Vibration and Control System Theory. With Computer Applications. Computing, 3(2):163, June 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1968:BRD

[290] H. Kremser. Book review: D. H. Stabley, System/360 Assembler Language. Computing, 3(2):163, June 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Junginger:1968:BRWb

[291] W. Junginger. Book review: W. Vogel, Lineares Optimieren. Computing, 3(2):164, June 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Merz:1968:PNI

[292] G. Merz. Padésche Näherungsbrüche und Iterationsverfahren höherer Ordnung. (German) [Padé approximation fractions and higher order iteration methods]. Computing, 3(3):165–183, September 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Albrecht:1968:BMW

[293] R. Albrecht. Bestimmung minimaler Wege in endlichen, gerichteten, bewerteten Graphen. (German) [Determination of minimum paths, in finite, directed, evaluated graphs]. Computing, 3(3):184–193, September 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Klemm:1968:APTa

[294] U. Klemm. Ein algorithmisches Planaritätskriterium (Teil I). (German) [An algorithmic planarity criterion (Part I)]. Computing, 3(3):194–204, September 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Shampine:1968:CMN

[295] L. F. Shampine. Contraction mappings for nonlinear boundary value problems. Computing, 3(3):205–214, September 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmidt:1968:AVH

[296] J. W. Schmidt and H. Schwetlick. Ableitungsfreie Verfahren mit höherer Konvergenzgeschwindigheit. (German) [Methods with rate of convergence not involving derivatives]. Computing, 3 (3):215–226, September 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ade:1968:IGE

[297] H. Ade. Iterative Gewinnung von Existenzaussagen für Randwertprobleme mit gewöhnlichen Differentialgleichungen 2. Ordnung. (German) [Iterative method of obtaining existence statements for boundary value problems with ordinary differential equations of 2nd order]. Computing, 3(3): 227–238, September 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Knodel:1968:ABA

[298] W. Knödel. Algorithmus 8: Bestimmung aller maximalen, vollständigen Teilgraphen eines Graphen G nach Stoffers. (German) [Algorithm 8: Determination of all maximal complete subgraphs of a graph G]. Computing, 3 (3):239–240, September 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See certification [345].

Handler:1968:BRB

[299] W. Händler. Book review: T. L. Booth, Sequential Machines and Automata Theory. Computing, 3(3):241, September 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1968:BRRc

[300] H. J. Stetter. Book review: R. Bellman, Methoden der Störungsrechnung in Mathematik, Physik und Technik. Computing, 3(3):241, September 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kolbe:1968:BRG

[301] O. Kolbe. Book report: G. A. Chebotarev, Analytical and Numerical Methods of Celestial Mechanics. Computing, 3(3):242, September 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fischer:1968:BRD

[302] K. Fischer. Book report: E. Dale and D. Michie (Herausgeber), Machine Intelligence 2. Computing, 3 (3):243, September 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1968:BRJ

[303] H. Kremser. Book report: J. Peters, Einführung in die allgemeine Informationstheorie. Computing, 3(3):243, September 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Albrecht:1968:BRH

[304] R. Albrecht. Book review: H. Zemanek, Alphabets and Codes 1967 for Information Processing. Computing, 3(3):244, September 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Firneis:1968:BRJ

[305] F. Firneis. Book review: A. J. Saxon und W. S. Plette, Die Programmierung der IBM 1401. Computing, 3(3):244, September 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Klemm:1968:APTb

[306] U. Klemm. Ein algorithmisches Planaritätskriterium (Teil II). (German) [An algorithmic planarity criterion (Part II)]. Computing, 3(4):245–257, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Alefeld:1968:AKA

[307] G. Alefeld. Über die asymptotische Konvergenzgeschwindigkeit des allgemeinen Relaxationsverfahrens bei nichtnegativen Matrizen. (German) [On the asymptotic convergence speed of the general relaxation method with non-negative matrices]. Computing, 3 (4):258–267, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tewarson:1968:OSV

[308] R. P. Tewarson. On the orthonormalization of sparse vectors. Computing, 3 (4):268–279, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Liniger:1968:CSL

[309] W. Liniger. A criterion for A-stability of linear multistep integration formulae. Computing, 3(4):280–285, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1968:IAS

[310] H. J. Stetter. Improved absolute stability of predictor-corrector schemes. Computing, 3(4):286–296, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Burkhardt:1968:APS

[311] W. H. Burkhardt. Automation of program speed-up on parallel-processor computers. Computing, 3(4):297–310, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schnorr:1968:VEK

[312] C. P. Schnorr. Vier Entscheidbarkeitsprobleme für kontextsensitive Sprachen. (German) [Four decidability problems for context-sensitive languages]. Computing, 3(4):311–317, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brannen:1968:NIS

[313] J. P. Brannen and R. E. Haymond. Non-iterative solutions to functional equations. Computing, 3(4):318–333, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Konig:1968:AAV

[314] E. v. König. Anwendung einer Art verallgemeinerter Zahlendarstellung bei der schnellen Fouriertransformation. (German) [Application of a kind of generalized numerical rotation to the fast Fourier transform]. Computing, 3 (4):334–350, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brauer:1968:SRP

[315] W. Brauer. Some remarks on a paper of G. Feichtinger. Computing, 3 (4):351–353, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Knodel:1968:VBG

[316] W. Knödel. Über die Verteilung der Binärziffern einer gemessenen Große. (German) [On the distribution of binary numbers of measured size]. Computing, 3(4):354–361, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Allgower:1968:CPN

[317] E. Allgower and R. Guenther. Correction to the paper: “On the Numerical Solution of Higher Order Nonlinear Parabolic Equations”. Computing, 3(4):362–364, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [277].

Scherer:1968:COS

[318] K. Scherer. Conference on abstract spaces and approximation. Computing, 3(4):365, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1968:BRRa

[319] E. Bukovics. Book review: R. Bellman, Dynamische Programmierung und selbstanpassende regelprozesse. Computing, 3(4):366, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1968:BRJ

[320] H. J. Stetter. Book review: J. Bertin, R. Ritout und J. G. Rougier, L’exploitation Partagée des Calculateurs. Computing, 3(4):366, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Muller:1968:BRH

[321] W. Müller. Book report: H. Bühler, Einführung in die Schaltalgebra. Computing, 3(4):367, December 1968. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Troch:1968:BRV

[322] I. Troch. Book report: V. Borsky and J. Matyäs, Computation by Electronic Analogue Computers. Computing, 3(4):367, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Junginger:1968:BRR

[323] W. Junginger. Book report: R. G. Busacker und Th. L. Saaty, Endliche Graphen und Netzwerke. Computing, 3(4):368, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1968:BRC

[324] H. J. Stetter. Book report: L. Collatz, G. Meinardus und H. Unger (Herausgeber), Funktionalanalysis, Approximationstheorie, Numerische Mathematik. Computing, 3(4):368, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1968:BRRb

[325] E. Bukovics. Book review: R. Drenick, Optimierung linearer Regelsysteme. Computing, 3(4):369, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wehrmann:1968:BRR

[326] W. Wehrmann. Book report: R. M. Fano, Informationsübertragung, Eine Statistische Theorie der Nachrichtenübertragung. Computing, 3(4):369, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1968:BRP

[327] H. Kremser. Book report: P. L. Hammer, and S. Rudeanu, Boolean Methods in Operations Research. Computing, 3(4):370, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Handler:1968:BRJ

[328] W. Händler. Book report: J. Klir and M. Valach, Cybernetic Modelling. Computing, 3(4):371, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Knodel:1968:BRJ

[329] W. Knödel. Book report: J. F. Harat, Computer Approximations. Computing, 3(4):371, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bertele:1968:BRM

[330] H. Bertele. Book report: M. Knoll, und J. Eichmeier, Technische Elektronik. I. Band: Grundlagen und Vakuumtechnik. Computing, 3(4):372, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bokovics:1968:BRK

[331] E. Bokovics. Book report: L. Kuipers und R. Timman, Handbuch der Mathematik. Computing, 3(4):373, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Roos:1968:BRM

[332] P. Roos. Book review: T. Menes, Kybernetische Maschinen. Computing, 3(4):373, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Meisslk:1968:BRU

[333] P. Meissl. Book review: N. U. Prabhu, Queues and Inventories. A Study of Their Basic Stochastic Processes. Computing, 3(4):374, December 1968. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1968:BRRd

[334] H. J. Stetter. Book report: A. Ralston und H. S. Wilf, Mathematische Methoden für Digitalrechner. Computing, 3(4):374, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Albrecht:1968:BRW

[335] R. Albrecht. Book review: A. W. Wymore, A Mathematical Theory of Systems Engineering: The Elements. Computing, 3(4):375, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Heinhold:1968:BRS

[336] J. Heinhold. Book review: A. S. Sweschnikow, Untersuchungsmethoden der Theorie der Zufallsfunktionen mit praktischen Anwendungen. Computing, 3(4):375, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fasol:1968:BRK

[337] K. J. Fasol. Book report: K. C. Garner, Introduction to Control System Performance Measurements. Computing, 3(4):376, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Risak:1968:BRJ

[338] R. Risak. Book review: J. S. Zypkin, Adaption und Lernen in automatischen Systemen. Ergebnisse, Probleme Perspektiven. Computing, 3(4):376, December 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Henze:1969:CMN

[339] D. Henze. Über die Charakterisierung von Minimallösungen bei nichtlinearen Approximationsproblemen. (German) [On the characterization of minimal solutions to nonlinear approximation problems]. Computing, 4(1):1–9, March 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Scott:1969:IIC

[340] M. R. Scott, L. F. Shampine, and G. M. Wing. Invariant imbedding and the calculation of eigenvalues for Sturm–Liouville systems. Computing, 4(1):10–23, March 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schassberger:1969:SST

[341] R. Schassberger. A service system with two parallel queues. Computing, 4(1):24–29, March 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hotz:1969:UAS

[342] G. Hotz. Übertragung automatentheoretishcer Sátze auf Chomsky-Sprachen. (German) [Transfer of automatontheoretical propositions to Chomsky languages]. Computing, 4(1):30–42, March 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ratschek:1969:EIG

[343] H. Ratschek. Über einige intervallarithmetische Grundbegriffe. (German) [On an interval arithmetic concept]. Computing, 4(1):43–55, March 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmitt:1969:TNU

[344] Alfred Schmitt. Zur Theorie der nichtdeterministischen und unvollständigen Automaten. (German) [Toward a theory of nondeterministic and incomplete automata]. Computing, 4(1):56–74, March 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Knodel:1969:ZA

[345] W. Knödel. Zertifikat zu Algorithmus 8. (German) [Certification of algorithm 8]. Computing, 4(1):75, March 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [298].

Albrecht:1969:AAP

[346] R. Albrecht and E. Visotschnig. Algorithmus 9. ALGOL-Prozeduren zu den modifizierten Algorithmen nach Minty und Moore. (German) [Algorithm 9. ALGOL-procedures to the modified algorithms according to Minty and Moore]. Computing, 4(1):76–81, March 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fischer:1969:BRC

[347] K. Fischer. Book review: C. Berge and A. Ghouila-Houri, Programme, Spiele, Transportenetze. Computing, 4(1):82, March 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wehrmann:1969:BRC

[348] W. Wehrmann. Book review: A. A. Charkewitsch. Computing, 4(1):82, March 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fischer:1969:BRD

[349] K. Fischer. Book review: D. N. Chorafas, Programmiersysteme für elektronische Rechenanlagen. Computing, 4(1):83, March 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Florian:1969:BRJ

[350] H. Florian. Book review: J. C. Cluley, Electronic Computers. Computing, 4(1):83, March 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1969:BRC

[351] H. J. Stetter. Book review: C. B. Germain, Programming the IBM 360. Computing, 4(1):84, March 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1969:BRFb

[352] H. J. Stetter. Book review: F. H. George, An Introduction to Computer Programming. Computing, 4(1):84, March 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Albrecht:1969:BRDa

[353] R. Albrecht. Book review: D. G. Hays, (Hg), Readings in Automatic language Processing. Computing, 4(1):85, March 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Albrecht:1969:BRDb

[354] R. Albrecht. Book review: D. G. Hays, (Hg), Introduction to Computational Linguistics (Series of Mathematical Linguistics and Automatic Language Processing, Vol. 2). Computing, 4(1):85, March 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Florian:1969:BRH

[355] H. Florian. Book report: H. W. Gschwind, Design of Digital Computers. Computing, 4(1):85, March 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Roos:1969:BRH

[356] P. Roos. Book review: H. Hotes, Digitalrechner in technisehen Prozessen. Computing, 4(1):86, March 1969. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Junginger:1969:BRH

[357] W. Junginger. Book review: H. P. Künzi. H. G. Tzschach und C. A. Zehnder, Numerische Methoden der mathematischen Optimierung mit ALGOL- und FORTRAN-Programmen (Leitfäden der angewandten Mathematik und Maechanik, Bd. 8). Computing, 4(1):87, March 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Florian:1969:BRP

[358] H. Florian. Book review: P. W. Melentjew, Näherungsmethoden. Computing, 4(1):88, March 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Unger:1969:BRD

[359] C. Unger. Book review: D. D. Mc-Cracken FORTRAN with Engineering Applications. Computing, 4(1):88, March 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Florian:1969:BRM

[360] H. Florian. Book review: M. P. Moyle, Introduction to Computers for Engineers. Computing, 4(1):89, March 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Handler:1969:BRR

[361] W. Händler. Book review: R. J. Nelson introduction to automata. Computing, 4(1):89, March 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1969:BRRa

[362] H. Kremser. Book review: R. Sauer und I Szabö (Hg.), Mathematische Hilfsmittel des Ingenieurs, Teil III (Die Grundlehren der mathematischen Wissenschaften in Einzeldarstellungen, Bd 141). Computing, 4(1):91, March 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Roos:1969:BRW

[363] P. Roos. Book review: W. W. Peterson, Prüfbare und korrigierbare Codes. Computing, 4(1):91, March 1969. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fehlberg:1969:KRK

[364] E. Fehlberg. Klassische RUNGE-KUTTA-Formeln fünfter und siebenter Ordnung mit Schrittweiten-Kontrolle. (German) [Classical fifth- and seventhorder Runge–Kutta formulas with stepsize control]. Computing, 4(2):93–106, June 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Claus:1969:ZST

[365] V. Claus and H. Walter. Zerlegung von Semi-Thue-Systemen. (German) [Decompositions of semi-Thue-systems]. Computing, 4(2):107–124, June 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fischer:1969:MEN

[366] H. Fischer. Eine Methode zur Einschliessung der Nullstellen von Polynomen mit komplexen Koeffizienten. (German) [A method for bracketing of zeros of a polynomial with complex coefficients]. Computing, 4(2):125–138, June 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tewarson:1969:CGI

[367] R. P. Tewarson. On computing generalized inverses. Computing, 4(2):139– 152, June 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stark:1969:ASS

[368] E. L. Stark. Über die Approximationsmaße spezieller singulärer Integrale. (German) [On the measure of approximation of particular singular integrals]. Computing, 4(2):153–159, June 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hadeler:1969:TSN

[369] K. P. Hadeler. Ein Templescher Satz fuer nichtlineare Eigenwertaufgaben. (German) [Temple bounds for nonlinear eigenvalue problems]. Computing, 4(2):160–167, June 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Holzwarth:1969:VBB

[370] A. Holzwarth. Ein Verfahren zur Bestimmung bester Tschebyscheff-Approx-imationen der Quadratwurzelfunktion. (German) A Method for Determination of Best Chebyshev Approximations to the Square Root Function. Computing, 4(2):168–177, June 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Spath:1969:AZG

[371] H. Späth. Algorithmus 10: Zweidimensionale glatte Interpolation. (German) [Algorithm 10: Two-dimensional smooth interpolation]. Computing, 4(2):178–182, June 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Albrecht:1969:BRC

[372] R. Albrecht. Book report: Ch. H. Davidson and E. C. Koenig, Computers: Introduction to Computers and Applied Computing Concepts. Computing, 4(2):183, June 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1969:BRD

[373] H. J. Stetter. Book review: D. W. Barron, Recursive Techniques in Programming. Computing, 4(2):183, June 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Albrecht:1969:BRF

[374] R. Albrecht. Book review: F. C. Hennie, Finite-State Models for Logical Machines. Computing, 4(2):184, June 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Albrecht:1969:BRG

[375] R. Albrecht. Book review: A. G. and V. G. Lapa, Cybernetics and Forecasting Techniques (Modern Analytic and Computational Methods in Science Mathematics, Vol. 8). Computing, 4(2):184, June 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1969:BRR

[376] H. J. Stetter. Book review: R. Herschel, ALGOL: – Übungen. Computing, 4(2):184, June 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eitel:1969:BRH

[377] H. Eitel. Book review: T. H. Lee, G. E. Adams and W. M. Gaines, Computer Process Control. Computing, 4(2):185, June 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1969:BRRb

[378] H. Kremser. Book review: R. L. Strationovich, Conditional MARKOV Processes and Their Application to the Theory of Optimal Control. Computing, 4(2):186, June 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Krawczyk:1969:NAB

[379] R. Krawczyk. Newton-Algorithmen zur Bestimmung von Nullstellen mit Fehlerschranken. (German) [Newton algorithms for evaluation of roots with error bounds]. Computing, 4(3): 187–201, September 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fettis:1969:NCJ

[380] H. E. Fettis. Note on the computation of Jacobi’s norm and its inverse. Computing, 4(3):202–206, September 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sarman:1969:BIE

[381] S. Šarman. Bemerkungen zu inversen Eigenwertprolemen. (German) [Notes to inverse eigenvalue problems]. Computing, 4(3):207–215, September 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Krabs:1969:PGL

[382] W. Krabs. Ein Pseudo-Gradientenverfahren zur Lösung des diskreten linearen Tschebysche-Problems. (German) [A pseudo-gradient procedure for the solution of the discrete linear Chebyshev approximation problem]. Computing, 4 (3):216–224, September 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Spath:1969:ESI

[383] H. Späth. Exponential spline interpolation. Computing, 4(3):225–233, September 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Marsal:1969:KIL

[384] D. Marsal. Konvergenzbeschleunigte Iteration von linearen Gleichungssystemen bei Divergenz des klassischen Verfahrens unter besonderer Berücksichtigung von Randwertproblemen. (German) [Convergence speedup iteration of linear systems of equations on divergence of the classical method with special regard to boundary value problems]. Computing, 4(3): 234–245, September 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kussmaul:1969:NVL

[385] R. Kussmaul. Ein numerisches Verfahren zur Lösung des Neumannschen Außenraumproblems für die Helmholtzsche Schwingungsgleichung. (German) [A numerical method for the exterior Neumann problem for the reduced wave equation]. Computing, 4(3):246–273, September 1969.

CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fischer:1969:PEH

[386] K. Fischer. Mitteilung/Note: Ein Programm zur Erstellung von Hochschulstundenplänen. (German) [Note: A program for creating university time tables]. Computing, 4(3):274, September 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Heinhold:1969:NAI

[387] J. Heinhold. Notes: AICA-IFIP Conference on Hybrid Computation (6th AICA Congress). Computing, 4(3):275, September 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Collatz:1969:BRV

[388] L. Collatz. Book review: V. Bracken and G. P. McCormick, Selected Applications of Nonlinear Programming. Computing, 4(3):276, September 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmid:1969:BRF

[389] E. Schmid. Book review: F. L. Bauer and A. S. Householder and F. J. Olver and H. Rutishauser and K. Samelson and E. Stiefel, Handbook for Automatic computation, Vol. I Die Grundlehren der mathematischen Wissenschaften in Einzeldarstellungen mit besonderer Berücksichtigung der Anwendungsgebiete, Band 135. Computing, 4(3):276, September 1969. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schnabl:1969:BRP

[390] R. Schnabl. Book review: P. L. Butzer and H. Erens, Semi-Groups of Operators and Approximation (Grundlehren der mathematischen Wissenschaften, Bd 145). Computing, 4(3):277, September 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Albrecht:1969:BRRa

[391] R. Albrecht. Book report: R J. Duffin and E. L. Peterson und C., M. Zener, Geometric Programming. Computing, 4(3):278, September 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Albrecht:1969:BRRb

[392] R. Albrecht. Book review: R. Gunzenhäuser, Nicht-numerische Informationsverarbeitung. Computing, 4 (3):278, September 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Krawczyk:1969:FRE

[393] R. Krawczyk. Fehlerabschátzung reeller Eigenwerte un Eigenvektoren von Matrizen. (German) [Error estimates for real eigenvalues and eigenvectors of matrices]. Computing, 4 (4):281–293, December 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rokne:1969:AEA

[394] J. Rokne and P. Lancaster. Automatic errorbounds for the approximate solution of equations. Computing, 4 (4):294–303, December 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Locher:1969:FQC

[395] F. Locher. Fehlerabschátzuang für das Quadraturverfahren von Clenshaw und Curtis. (German) [Error estimates for the quadrature method of Clenshaw and Curtis]. Computing, 4(4): 304–315, December 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Feilmeier:1969:MLG

[396] M. Feilmeier. Methode zur Lösung gewisser linearer Operatorgleichungen. (German) [A method of determining the solution of certain linear operator equations]. Computing, 4(4): 316–328, December 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Junginger:1969:EBE

[397] W. Junginger. Erfahrungen beun Einsartz eines Computers zur Studenplanerstellung. (German) [Experiences in construction of timetables by computer]. Computing, 4(4):329–344, December 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schwetlick:1969:LPO

[398] H. Schwetlick. Lineare positive Operatoren und Fehlerabschätzungen bei Operatorgleichungen. (German) [Linear positive operators and error estimates to operator equations]. Computing, 4(4):345–358, December 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lazak:1969:HAA

[399] D. Lazak. A heuristic approach to algorithms intended for the solution of the timetable-problem. Computing, 4 (4):359–377, December 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lazak:1969:ADG

[400] D. Lazak. Algorithmus 11. Dienstplanerstellung. (German) [Algorithm 11. construction of timetables]. Computing, 4(4):378–384, December 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hoffmann:1969:BRJ

[401] H. J. Hoffmann. Book review: J. M. Foster, List Processing (Computer Monographs Series, No. 1). Computing, 4(4):385, December 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1969:BRF

[402] H. J. Stetter. Book review: E. F. Bradley and O. T. Denmead, The Collection and Processing of Field Data. Computing, 4(4):385, December 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Troch:1969:BRH

[403] I. Troch. Book review: H. Leipholz, Stabilitätstheorie (Leitfaden der Angewandten Mathematik, Bd. 10). Computing, 4(4):385, December 1969. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1969:BRH

[404] H. J. Stetter. Book review: H. R. Schwarz, H. Rutishauser und E. Stiefel, Numerik symmetrischer Matrizen (Leitfaden der angewandten mathematik und Mechanik, Bd. 11).

Computing, 4(4):386, December 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1969:BRP

[405] H. J. Stetter. Book review: Procédures ALGOL en Analyse Numérique. Computing, 4(4):386, December 1969. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Klar:1969:BRW

[406] R. Klar. Book review: W. E. Wickes, Logic Design with Integrated Circuits. Computing, 4(4):387, December 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wilson:1970:CPD

[407] M. W. Wilson. Convergence properties of discrete analogs of orthogonal polynomials. Computing, 5(1):1–5, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Spaniol:1970:DIG

[408] O. Spaniol. Die Distributivität in der Intervallarithmetik. (German) [Distributivity in interval arithmetic]. Computing, 5(1):6–16, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dirschmid:1970:EEVa

[409] H. Dirschmid. Zur Einschließung der Eigenwerte vollstetiger positiver Operatoren in separablem Hilbert-Räumen (Teil I). (German) [Enclosing the eigenvalues of completely continuous positive operators in separable Hilbert spaces. Part I]. Computing, 5(1):17–26, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Petry:1970:ILR

[410] W. Petry. Ein Iterationsverfahren zum Lösen von Randwertproblemen nichtlinearer Differentialgleichungen. (German) [An iteration method for the solution of boundary value problems of nonlinear differential equations]. Computing, 5(1):27–44, March 1970. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Even:1970:DSD

[411] R. K. Even and Y. Wallach. On the direct solution of Dirichlet’s problem in two dimensions. Computing, 5(1):45–56, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Vollmar:1970:APG

[412] R. Vollmar. Über einen Automaten mit Pufferspeicherung. (German) [On an automaton with buffer-tape]. Computing, 5(1):57–70, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmidt:1970:BVA

[413] J. W. Schmidt and D. Leder. Bleitungsfreie Verfahren Ohne Auflösung linearer Gleichungen. (German) [Derivationfree methods without solving linear equations]. Computing, 5(1):71–81, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schwetlick:1970:AAV

[414] H. Schwetlick. Algorithmus 12. Ein abeitungsfreies Verfahren zur Lösung endlich-dimensionaler nichtlinearer Gleichungssysteme. (German) [Algorithm 12. A discrete method for the solution of finite-dimensional systems of nonlinear equations]. Computing, 5(1):82–87, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Anonymous:1970:WV

[415] Anonymous. W. Vonhof. Computing, 5(1):88, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bohme:1970:BRR

[416] J. F. Böhme. Book review: R. G. Gallager, Information Theory and Reliable Communications. Computing, 5(1):89, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Unger:1970:BRP

[417] C. Unger. Book review: E. P. Billeter, Der praktische Einsatz von Datenverarbeitungssystemen. Computing, 5(1):89, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1970:BRHa

[418] E. Bukovics. Book review: H. B. Behnke, L. Collatz, R. Sauer und H. Unger (Hg.), Grundzüge der Mathematik, für Lehrer an Gymnasien sowie für Mathematiker in Industrie und Wirtschaft. Band V. Computing, 5(1): 90, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Albrecht:1970:BRM

[419] R. Albrecht. Book review: M. I. Halpern und C. J. Shaw, Annual Review in Automatic Programming 5 (International Tracts in Computer Science and Technology and Their Application). Computing, 5(1):91, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Albrecht:1970:BRB

[420] R. Albrecht. Book review: B. Higman, A Comparative Study of Programming Languages (Computer Monograph Series). Computing, 5(1):92, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Albrecht:1970:BRF

[421] R. Albrecht. Book review: F, J, Hill and G. R. Peterson, Introduction to Switching Theory and Logical Design. Computing, 5(1):92, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Albrecht:1970:BRS

[422] R. Albrecht. Book review: Sze-Tsen Hu, Mathematical Theory of Switching Circuits and Automata. Computing, 5(1):92, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Unger:1970:BRI

[423] C. Unger. Book review: IFIP-Fachwörterbuch der Informationsverarbeitung. Computing, 5(1):92, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eier:1970:BRH

[424] R. Eier. Book review: H. Kaden, Theoretiche Grundlagen der Datenübertragung. Computing, 5(1):93, March 1970. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Meissl:1970:BRP

[425] P. Meissl. Book review: P. Kreis, COBOL-Praxis. Computing, 5(1):93, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1970:BRHb

[426] E. Bukovics. Book review: H. P. Künzi, O. Müller und E. Nieveelt, Einführungskursus in die dynamische Programmierung (Lecture Notes in Operations Research and Mathematical Economics, Vol. 6). Computing, 5(1): 94, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1970:BRJ

[427] E. Bukovics. Book review: J. B. Plant, Some Iterative Solutions in Optimal Control Research Monograph, No. 44. Computing, 5(1):94, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Troch:1970:BRH

[428] I. Troch. Book review: H. Schwarz, Mehrfachregelungen. Computing, 5(1): 94, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1970:BRF

[429] E. Bukovics. Book review: F. G. Tricomi Repertorium der Theorie der Differentialgleichungen. Computing, 5(1): 95, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Florian:1970:BRF

[430] H. Florian. Book review: F. Stuart, Fortran Programming. Computing, 5(1):95, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hauer:1970:BRH

[431] F. Hauer. Book review: H. Wolf, Ausgleichsrechnung nach der Methode der kleinsten Quadrate. Computing, 5(1):96, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bauhuber:1970:DVB

[432] F. Bauhuber. Direkte Verfahrenzur Berechnung der Nullstellen von Polynomen. (German) [Direct methods for evaluation of zeroes of polynomials]. Computing, 5(2):97–118, June 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dirschmid:1970:EEVb

[433] H. Dirschmid. Zur Einschließung der Eigenwerte vollstetiger positiver Operatoren in separablem Hilbert-Räumen (Teil II). (German) [Enclosing the eigenvalues of completely continuous positive operators in separable Hilbert spaces. Part II]. Computing, 5(2):119– 127, June 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ehlich:1970:MDG

[434] H. Ehlich and E. Pittnauer. Mechanisierung von Dreieckskonstruktionen. (German) [Mechanization of triangle constructions]. Computing, 5(2):128– 135, June 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Prakash:1970:SBS

[435] S. Prakash and S. Kumar. Stochastic behaviour of a stand-by redundant complex system with imperfect switching over devices. Computing, 5(2):136–143, June 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). With 1 figure.

Mayer:1970:AMS

[436] O. Mayer. Algebraische und metrische Strukturen in der Intervallrechnung und einige Anwendungen. (German) [Algebraic and metric structure in interval computation and some uses]. Computing, 5(2):144–162, June 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmid:1970:RDG

[437] E. Schmid. Rechenzeitvergleich bei Digitalrechnern. (German) [Computing time comparisons of digital computers]. Computing, 5(2):163–177, June 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brauer:1970:BZU

[438] W. Brauer. Eine Bemerkung zur Zustandsreduktion unvollständiger Automaten. (German) [A comment on condition reduction for incomplete automata]. Computing, 5(2):178–184, June 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Anonymous:1970:LCG

[439] Anonymous. Lothar Collatz zum 60. Geburtstag. (German) [On lothar collatz’s 60th birthday]. Computing, 5 (3):185, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bohl:1970:MSR

[440] E. Bohl. Die metrische Struktur von Räumen mit allgemeinerem Abstandsbegriff und ihre Verwendung bei Behandlung nichtlinearer Probleme. (German) [The metric structure of spaces with generalized distance function and its application to the study of nonlinear problems]. Computing, 5(3): 189–199, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Krawczyk:1970:IAS

[441] R. Krawczyk. Intervallmäßerung Auswertung von Standardfunktionen in ALGOL-60. (German) [ALGOL-60 procedures evaluating standard functions in interval analysis]. Computing, 5(3):200, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Krawczyk:1970:VSE

[442] R. Krawczyk. Verbesserung von Schranken für Eigenwerte und Eigenvektoren von Matrizen. (German) [Improvement of bounds for matrix eigenvalues and eigenvectors]. Computing, 5 (3):200–206, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Babuska:1970:FEMa

[443] I. Babuška. The finite element method for elliptic equations with discontinuous coefficients. Computing, 5(3): 207–213, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Natterer:1970:BNI

[444] F. Natterer. Berechnung der Norm der Inversen eines Differential-operators zweiter Ordnung. (German) [Computation of the norm of the inverse of a second order differential operator]. Computing, 5(3):214–220, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Monien:1970:KDP

[445] B. Monien. Über die Konvergenzordnung von Differenzenverfahren, die parabolische Anfangsrandwertaufgaben approximieren. (German) [On the rate of convergence of differenceschemes which approximate parabolic initial-value problems]. Computing, 5 (3):221–245, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Werner:1970:EPL

[446] J. Werner. Einschließungassätze für periodische Lösungen der Liénardschen Differentialgleichung. (German) [Limits for the periodic solution of the Lienard differential equation]. Computing, 5 (3):246–252, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Laasonen:1970:IMI

[447] P. Laasonen. An imbedding method of iteration with global convergence. Computing, 5(3):253–258, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Opfer:1970:MPL

[448] G. Opfer. On monotonicity preserving linear extrapolation sequences. Computing, 5(3):259–266, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1970:STS

[449] H. J. Stetter. Symmetric two-step algorithms for ordinary differential equations. Computing, 5(3):267–280, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Meis:1970:ILN

[450] Th. Meis and W. Törnig. Iterative Lösung nichtlinearer Gleichungssysteme und Diskretisierungsverfahren bei elliptischen Differentialgleichungen. (German) [Iterative solutions for systems of non-linear equations and discretisation of elliptic differential equations]. Computing, 5(3):281–294, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brosowski:1970:FMP

[451] B. Brosowski. Über eine Fixpunkteigenschaft der metrischen Projektion. (German) [On a fixed-point property of the metric projection]. Computing, 5 (3):295–302, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Niethammer:1970:ULG

[452] W. Niethammer. Über- und Unterrelaxation bei linearen Gleichungssystemen. (German) [Over- and underrelaxation applied to systems of linear equations]. Computing, 5(3):303–311, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

vanderSluis:1970:DUP

[453] A. van der Sluis. Domains of uncertainty for perturbed operator equations. Computing, 5(3):312–323, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Albrecht:1970:ALO

[454] R. Albrecht and B. Buchberger. Algorithmus 13, Lösung eines Optimum-Mix Problems. (German) [Algorithm 13, solution of optimal mixing problems]. Computing, 5(3):324–331, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Anonymous:1970:MVI

[455] Anonymous. Mitteilung: V. Internationaler Kongreß über Anwendungen der Mathematik in den Ingenieurwissenschaften in Weimar (V. Ikm). (German) [Note: Fifth International Congress on Application of Mathematics in Engineering Science in Weimar]. Computing, 5(3):332, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Behringer:1970:BFI

[456] F. A. Behringer. Eine bedingte Form der Integral-Ungleichun von Gronwal, leichte Verallgemeinerung eins Stabilitätsatzes von Perron und überlauffreie Skalierung von Analogrechenschaltungen. (German) [A conditional version of the integral inequality of Gronwall, a slight generalization of a stability theorem of Perron, and overflow-free scaling of analogue computer set-ups]. Computing, 5(4): 333–348, December 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lamprecht:1970:MAN

[457] G. Lamprecht. Zur Mehrdeutigkeit bei der Approximation in der L_{p}-Norm mit Hilfe rationaler Funktionen. (German) [On the non-uniqueness of the best rational LI_{p}-approximation]. Computing, 5(4):349–355, December 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Krawczyk:1970:ENH

[458] R. Krawczyk. Einschliessung von Nullstellen mit Hilfe einer Intervallarithmetik. (German) [Bracketing of zeros with the help of interval arithmetic]. Computing, 5(4):356–370, December 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tewarson:1970:MSI

[459] R. P. Tewarson. Minimax solution of n + 1 inconsistent linear equations in n unknowns. Computing, 5(4): 371–376, December 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Herzberger:1970:IAS

[460] J. Herzberger. Intervallmäßige Auswertung von Standardfunktionen in Algol-60. (German) [ALGOL-60 procedures evaluating standard functions in interval analysis]. Computing, 5(4): 377–384, December 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hausch:1970:RFA

[461] E. Hausch and W. Knödel. Revidierte Fassung von Algorithmus 7. Kürzeste Rundreise zwischen n Orten (Travelling Salesman Problem). (German) [Shortest route for n points (travelling salesman problem)]. Computing, 5 (4):385–392, December 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Troch:1970:BRKa

[462] I. Troch. Book review: K. H. Fasol, Frequenzkennlinien. Computing, 5(4):394, December 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Nuding:1970:COR

[463] E. Nuding and I. Kahlert-Warmbold. A computer-oriented representation of matrices. Computing, 6(1–2):1–8, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Nickel:1970:FNP

[464] K. Nickel. Fehlerschranken zu Näherungswerten von Polynomwurzeln. (German) [Error bounds for approximate values of polynomials]. Computing, 6(1–2):9–27, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Alefeld:1970:AAA

[465] G. Alefeld and J. Herzberger. Algol-60 Algorithmen zur Auflösung linearer Gleichungssysteme mit Fehlererfassung. (German) [Algol-60 algorithms for solution of linear systems of equation with error control]. Computing, 6(1–2):28–34, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dussel:1970:BSW

[466] R. Dussel and B. Schmitt. Die Berechnung von Schranken für den Wertebereich eines Polynoms in einem Intervall. (German) [The calculation of bounds for the range of a polynomial in an interval]. Computing, 6 (1–2):35–60, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fehlberg:1970:KRK

[467] E. Fehlberg. Klassische Runge–Kutta-Formeln vierter und niedrigerer Ordnung mit Schrittweiten-Kontrolle und ihre Anwendung auf Wärmeleitungsprobleme. (German) [Classical Runge–Kutta formulas of fourth and lower order with stepsize control and their use for heat problems]. Computing, 6(1–2):61–71, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dirschmid:1970:EAG

[468] H. Dirschmid. Ein Einschließungssatz für den absolut grössten Eigenwert normaler Operatoren in separablen Hilbert-Räumen. (German) [An inclusion estimate for the absolutely largest eigenvalue of normal operators in separable Hilbert spaces]. Computing, 6 (1–2):72–75, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dirschmid:1970:EKE

[469] H. Dirschmid. Ein Einschließung der Komponenten des erster Eigenvektors normaler Operatoren in unitären Räumen. (German) [An inclusion of the components of the first eigenvectors of normal operators in unitary spaces]. Computing, 6(1–2):76–79, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Marsal:1970:EEL

[470] D. Marsal. Elementare Eingabelung der Lösungsfunktion und der Eigenwerte inhomogener gewöhnlicher und singulärer Integralgleichungen zweitzer. (German) [Elementary determination of eigenvalue bounds and of functions bounding the solution of inhomogeneous ordinary and singular integral equations of the second kind with fixed limits of integration]. Computing, 6(1–2):80–93, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmidt:1970:QGI

[471] K. Schmidt and P. Zinterhof. Über Quadraturformeln auf T^{ω}. (German) [On the integral of functions on T^{ω}]. Computing, 6(1–2):94–96, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fischer:1970:TIM

[472] W. L. Fischer. Topological invariants of matrices. Computing, 6(1–2):97–103, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mayer:1970:KKI

[473] O. Mayer. Über eine Klasse komplexer Intervallgleichungssysteme mit iterationsfähiger Gestalt. (German) [On a class of systems of complex interval equations evaluable with iteration methods]. Computing, 6(1–2):104–106, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmidt:1970:KMK

[474] G. Schmidt. Ein kontinuierliches Modell für Kraftfahrzeugsströmungen. (German) [A continuous model for the solution of traffic flow problems]. Computing, 6(1–2):107–120, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). Check pages??

Ahrens:1970:PRN

[475] J. H. Ahrens, U. Dieter, and A. Grube. Pseudo-random numbers. A new proposal for the choice of multiplicators. Computing, 6(1–2):121–138, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Golinski:1970:ICS

[476] J. Golinski. Investigation of a certain stray process applied to the optimum synthesis problems. Computing, 6(1–2):139–160, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Alefeld:1970:MZO

[477] G. Alefeld. Über die aus monoton Zerlegbaren Operatoren Gebildeten Iterationsverfahren. (German) [On iterative methods derived from monotonically reducible operators]. Computing, 6(1–2):161–172, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Vogel:1970:SFU

[478] E. Vogel. Ein System zum formalen Umformen mathematischer Ausdrücke mit Hilfe von Rechenautomaten. (German) [A system for computer-aided formal transformation of mathematical expressions]. Computing, 6(1– 2):173–190, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Culik:1970:AA

[479] Karel Culik II. Asynchronous automata. Computing, 6(1–2):191–199, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Govil:1970:SBC

[480] A. K. Govil and S. Kumar. Stochastic behaviour of a complex system under priority repair. Computing, 6 (1–2):200–213, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmid:1970:BRF

[481] E. Schmid. Book review: F. L. Bauer and A. S. Householder and F. W. J. Oliver and H. Rutishauser and K. Samelson and E. Stiefel (Hg.), Handbook for Automatic Computation, Vol. I, Part b (Die Grundlehren der mathematischen Wissenschaften in Einzeldarstellungen mit besonderer Berücksichtigung der Anwendungsgebiete: Band 137). Computing, 6(1–2): 214, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Troch:1970:BRG

[482] I. Troch. Book review: S. Goldberg, Differenzengleichungen und ihre Anwendung in Wirtschaftswissenschaft, Psychologie und Soziologie. Computing, 6(1–2):214, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bertele:1970:BRS

[483] H. Bertele. Book review: E. S. Savas, Computer Control of Industrial Processes. Computing, 6(1–2):215, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Collatz:1970:BRJ

[484] L. Collatz. Book review: J. H. Wilkinson, Rundungsfehler. Computing, 6(1–2):215, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Collatz:1970:BRB

[485] L. Collatz. Book review: B. Dehib and O. Gebrucg (Hg), Constructive Aspects of the Fundamental Theorem of Algebra. Computing, 6(1–2):216, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Havelec:1970:BRR

[486] L. Havelec. Book review: R. T. Kimaber, Automation in Libraries. Computing, 6(1–2):217, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pirzl:1970:BRJ

[487] I. Pirzl. Book review: J. Kowalik und M. R. Osborne, Methods for Unconstrained Optimization Problems. Computing, 6(1–2):217, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1970:BRR

[488] H. Kremser. Book report: R. Sauer und I. Szabó (Hg.), Mathematische Hilfsmittel des Ingenieurs, Teil II (Grundlehren der math. Wiss. in Einzeldarstellungen: Bankd 140). Computing, 6(1–2):218, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Troch:1970:BRKb

[489] I. Troch.  Book review:  K.

Göldner, Mathematische Grundlagen für Regelungstechniker. Computing, 6(1–2):218, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Barth:1970:BRO

[490] W. Barth. Book review: A. Opler (Hg.), Das IBM-System/360 und seine Programmiertechniken. Computing, 6(1–2):219, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Troch:1970:BRS

[491] I. Troch. Book review: N. Steinberg, Analogiques et Hybrides. Computing, 6(1–2):219, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Barth:1970:BRC

[492] W. Barth. Book report: Clarence B. Germain, Das Programmier-Handbuch der IBM/360. Computing, 6(1–2):220, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Troch:1970:BRV

[493] I. Troch. Book review: V. Paschkis und F. L. Ryder, Direct Analog Computers. Computing, 6(1–2):220, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gustavson:1970:FRN

[494] F. G. Gustavson and W. Liniger. A fast random number generator with good statistical properties. Computing, 6(3– 4):221–226, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Visotschnig:1970:RMA

[495] E. Visotschnig. Ein reduzierter Matrix-Algorithmus zur Lösung einiger graphentheoretischer Probleme. (German) [A reduced matrix algorithm for the solution of some problems in the theory of graphs]. Computing, 6(3– 4):227–240, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tewarson:1970:UGI

[496] R. P. Tewarson. On the use of generalized inverses in function minimization. Computing, 6(3–4):241–248, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Neumann:1970:VLG

[497] K. Neumann. Ein Verfahren zur Lösung gewisser nichtlinearer Kontrollprobleme. (German) [A method for the solution of certain nonlinear control problems]. Computing, 6(3–4): 249–263, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Babuska:1970:FEMb

[498] I. Babuška. Finite element method for domains with corners. Computing, 6(3– 4):264–273, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kress:1970:IEU

[499] R. Kress. Interpolation auf einem unendliehen Intervall. (German) [Interpolation on an infinite interval]. Computing, 6(3–4):274–288, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schaback:1970:LLT

[500] R. Schaback and D. Braess. Eine Lösungsmethode für die lineare Tschebyscheff-Approximation bei nicht erfüllter Haarscher Bedingung. (German) [A method for solving the linear Tchebycheff approximation problem when the Haar condition is not satisfied].

Computing, 6(3–4):289–294, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ratschek:1970:BSI

[501] H. Ratschek. Die binären Systeme der Intervallarithmetik. (German) [The binary system of interval arithmetic]. Computing, 6(3–4):295–308, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stoss:1970:BSK

[502] Hanns-Jörg Stoss. k-Band-Simulation von k-Kopf-Turing-Maschinen. (German) [k-tape simulation of k-head Turing machines]. Computing, 6(3–4): 309–317, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmidt:1970:ELH

[503] J. W. Schmidt and H. Leonhardt. Eingrenzung von Lösungen mit Hilfe der Regula falsi. (German) [Inclusion of solutions by regula falsi]. Computing, 6(3–4):318–329, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Henze:1970:ENT

[504] D. Henze. Über Eindeutigkeitsfragen bei nichtlinearer D- un Tschebyscheff-Approximation. (German) [On uniqueness in nonlinear L^{1}- and Chebyshevapproximation]. Computing, 6(3–4): 330–338, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

DeOliveira:1970:IEP

[505] G. N. De Oliveira. Inverse eigenvalue problems for complex matrices. Computing, 6(3–4):339–341, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Habetler:1970:FDM

[506] G. J. Habetler and R. L. Schiffman. A finite difference method for analyzing the compression of poro- viscoelastic media. Computing, 6(3–4): 342–348, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Elgot:1970:CAS

[507] Calvin C. Elgot. The common algebraic structure of exit-automata and machines. Computing, 6(3–4):349–370, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1970:RS

[508] H. J. Stetter. Robert Sauer. Computing, 6(3–4):371–372, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Barth:1970:BRB

[509] W. Barth. Book review: B. P. Demidowitsch and I. A. Maron und E. S. Schuwalowa, Numerische Methoden der Analysis. Computing, 6 (3–4):373, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmid:1970:BRG

[510] E. Schmid. Book review: G. Bayer, Einführung in das Programmieren (De Gruyter Lehrbuch IV). Computing, 6 (3–4):373, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wolff:1970:BRW

[511] K.-H. Wolff. Book review: W. Dinkelbach, Sesitivitätsanalysen und parametrische Programmierung (Ökonometrie und Unternehmensforschung, Band XII). Computing, 6(3–4):374, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Florian:1970:BRR

[512] H. Florian. Book review: R. Herschel, Anleitung zum praktischen Gebrauch von ALGOL 60, 3. Aufl. Computing, 6(3–4):375, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Oehme:1970:BRJ

[513] F. Oehme. Book report: J. K. Hawkins, Circuit Design of Digital Computers. Computing, 6(3–4):375, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Florian:1970:BRD

[514] H. Florian. Book review: D. J. Malcolme-Lawes, Programming-ALGOL. Computing, 6(3–4):376, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Florian:1970:BRW

[515] H. Florian. Book review: W. E. Spiess und F. G. Rheingans, Einführung in das Programmieren in FORTRAN. Computing, 6(3–4):376, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Jeng:1971:DCA

[516] D.-T. Jeng. Direct computational approaches to a magnetohydrodynamic model system. Computing, 7(1–2):1– 12, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Allen:1971:NSC

[517] R. C. Allen, Jr. On the numerical solution of a certain class of nonlinear Cauchy problems arising from integral equation theory. Computing, 7(1–2):13–16, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

McLeod:1971:NA

[518] J. B. McLeod. A note on the ϵalgorithm. Computing, 7(1–2):17–24, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Petry:1971:VNI

[519] W. Petry. Eine Verallgemeinerung des Newtonschen Iterationsverfahrens. (German) [A generalization of the Newton-iteration method]. Computing, 7(1–2):25–45, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schock:1971:MRV

[520] E. Schock. Modifizierte Ritz-Verfahren. (German) [The modified Ritz-method]. Computing, 7(1–2):46–52, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bohl:1971:ING

[521] E. Bohl. Zur Iteration bei nichtlinearen Gleichungssystemen. (German) [On iterative processes for nonlinear systems]. Computing, 7(1–2):53–64, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Spath:1971:NCH

[522] H. Späth. The numerical calculation of high degree Lidstone splines with equidistant knots by blockunderrelaxation. Computing, 7(1–2):65–74, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Spath:1971:NCQ

[523] H. Späth. The numerical calculation of quintic splines with equidistant knots by blockunderrelaxation. Computing, 7(1–2):75–82, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Srivastava:1971:SBP

[524] S. S. Srivastava, R. C. Garg, N. K. Malhotra, and S. Kumar. Stochastic behaviour of a parallel redundant complex system with three types of components. Computing, 7(1–2):83–90, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hoffmann:1971:UCP

[525] G.-R. Hoffmann and G. Veenker. The unit-clause proof procedure with equality. Computing, 7(1–2):91–105, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hofmann:1971:RFB

[526] W. Hofmann. Die Regula falsi in Banach-Räumen. (German) [The regula false in Banach spaces]. Computing, 7(1–2):106–112, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schindler:1971:VNV

[527] S. Schindler. Ein vereinfachtes Newtonsches Verfahren für Differentialgleichungen. (German) [A simpler Newton’s method for differential equations]. Computing, 7(1–2):113–126, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wegler:1971:A

[528] B. Wegler. Algorithmus 14. (German) [Algorithm 14]. Computing, 7 (1–2):127–128, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bauer:1971:HR

[529] F. L. Bauer. Heinz Rutishauser (10.11.1970). Computing, 7(1–2): 129–130, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/BF02279949.

Pirzl:1971:BRC

[530] I. Pirzl. Book review: Computing Methods in Optimization Problems (Lecture Notes in Operations Research and Mathematical Economics: Vol. 14). Computing, 7(1–2):131, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Beilner:1971:BRW

[531] H. Beilner. Book review: W. Händler, Teilnehmer-Rechensysteme. Computing, 7(1–2):133, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Nake:1971:BRM

[532] F. Nake. Book report: M. V. Mathews with collaboration of J. E. Miller, F. R. Moore J. R. Pierce, J. C. Risset, The Technology of Computer Music. Computing, 7(1–2):134, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Beilner:1971:BRJ

[533] H. Beilner. Book review: J. E. Sammet, Programming Languages: History and Fundamentals. Computing, 7(1–2):135, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pirzl:1971:BRHa

[534] I. Pirzl. Book review: H. Bauer und K. Neumann, Berechnung optimaler Steuerungen, Lecture Notes in Operations Research and Mathematical Economics: Vol. 17. Computing, 7(1–2):136, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Troch:1971:BRG

[535] I. Troch. Book review: G. A. Bekey und W. J. Karplus, Hybrid Computation. Computing, 7(1–2):136, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eberl:1971:BRL

[536] W. Eberl. Book report: E. Lukacs, Stochastic Convergence (Health Mathematical Monographs). Computing, 7(1–2):137, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1971:BRJ

[537] H. J. Stetter. Book review: J. G. Hayes, Numerical Approximations to Functions and Data. Computing, 7(1–2): 138, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Uberhuber:1971:BRP

[538] Ch. Überhuber. Book report: P.-R. Lohnes and W. W. Cooley, Introduction to Statistical Procedures. Computing, 7(1–2):138, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Albrecht:1971:BRK

[539] R. Albrecht. Book report: A. Kusko und T. Wroblewski, Computer-Aided Design of Magnetic Circuits (Research Monograph No. 55). Computing, 7 (1–2):139, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Osterreicher:1971:BRP

[540] F. Österreicher. Book review: L. P. Hyvärinen, Information Theory for System Engineers (Lectures Notes in Operations Research and Mathematical Economics: Vol. 5). Computing, 7(1–2):139, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Weinmann:1971:BRV

[541] A. Weinmann. Book review: A. V. (Hg.), Papers on digital signal processing. Computing, 7(1–2):139, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Albrecht:1971:BRB

[542] R. Albrecht. Book review: B. H. Rudwick, System Analysis for Effective Planning: Principles and Cases (Wiley Series on Systems Engineering and Analysis). Computing, 7(1–2):140, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Albrecht:1971:BRH

[543] R. Albrecht. Book report: H. B. Mann, (Hg.), Error Correcting Codes. Computing, 7(3–4):140, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Albrecht:1971:BRI

[544] R. Albrecht. Book review: The International Directory of Computer and Information System Services. Computing, 7(1–2):140, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Albrecht:1971:BRS

[545] R. Albrecht. Book review: S. Seoratschek, Einführung in die Datenverabeitung. Computing, 7(1–2):141, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1971:BRR

[546] H. J. Stetter. Book review: R. Ralston und H. S. Wilf, Mathematische Methoden für Digitalrechner II. Computing, 7(1–2):141, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1971:BRR

[547] H. Kremser. Book review: R. Sauer, Ingenieur-Mathenatuj. Computing, 7 (1–2):142, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Weinmann:1971:BRB

[548] A. Weinmann. Book review: B. S. Walker, Einführung in die Computer-Technik. Computing, 7(1–2):142, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1971:BRK

[549] K. Kremser. Book review: K. M. Brauer, Binäre Optimierung. Computing, 7(1–2):143, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Weinmann:1971:BRP

[550] A. Weinmann. Book review: P. Vidal, Non-linear Sampled-date Systems. Computing, 7(1–2):143, March 1971.

CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1971:BRB

[551] E. Bukovics. Book review: E. Bürger und W. Schuppe, Four-language Technical Dictionary of Data Processing, Computers and Office Machines. Computing, 7(1–2):144, March 1971. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Koepf:1971:BRR

[552] G. Koepf. Book review: A. R. Shulman, Optical Data Processing. Computing, 7(1–2):144, March 1971. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rokne:1971:FEM

[553] J. Rokne. Fehlererfassung bei Eigenwertproblemen von Matrizen. (German) [Error control of the matrix eigenvalue problem]. Computing, 7(3– 4):145–152, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Muller:1971:SZN

[554] Karl Hans Müller. Zum schwachen Zeilensummenkriterium bei nichtlinearen Gleichungssystemen. (German) [On the weak row sum criterion for nonlinear systems of equations]. Computing, 7(3–4):153–171, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ratschek:1971:AIF

[555] H. Ratschek and Gunnar Schröder. Über die Ableitung von intervallwertigen Funktionen. (German) [On the derivatives of interval-valued functions]. Computing, 7(3–4):172–187, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ecker:1971:VRB

[556] K. Ecker. Verteilung der Rundungsfehler beim Produkt zweier Zahlen auf einer Festkommamaschine. (German) [Distribution of rounding errors in products of two numbers in a fixedpoint machine]. Computing, 7(3–4): 188–193, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kuich:1971:IAS

[557] W. Kuich and H. Maurer. On the inherent ambiguity of simple tuple languages. Computing, 7(3–4):194–203, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Krabs:1971:NOU

[558] W. Krabs. Niehtlineare Optimierung mit unendlich vielen Nebenbedingungen. (German) [Nonlinear programming with infinitely many constraints]. Computing, 7(3–4):204–214, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dutter:1971:NBV

[559] R. Dutter and H. Zackor. Numerische Behandlung des Verkehrsflußmodells von Prigogine. (German) [Numerical treatment of the traffic flow model established by Prigogine]. Computing, 7(3–4):215–221, September 1971.

CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stoss:1971:ZBS

[560] H. J. Stoss. Zwei-Band Simulation von Turingmaschinen. (German) [A twotape simulation of Turing machines]. Computing, 7(3–4):222–235, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tewarson:1971:TDM

[561] R. P. Tewarson. On two direct methods for computing generalized inverses. Computing, 7(3–4):236–239, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Odeh:1971:NUF

[562] F. Odeh and W. Liniger. A note on unconditional fixed-h stability of linear multistep formulae. Computing, 7(3– 4):240–253, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Esser:1971:KQR

[563] H. Esser. Konvergenz von Quadraturverfahren vom Radau-Typ. (German) [Convergence of quadrature formulae of Radau-type]. Computing, 7(3–4): 254–263, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Weicker:1971:TTK

[564] R. Weicker. Tabulator-Turingmaschine und Komplexität. (German) [Tabulator-Turing machine and complexity]. Computing, 7(3–4):264–274, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See amendment [668].

Govil:1971:CSP

[565] A. K. Govil. Complex system pointwise availability with priority repair. Computing, 7(3–4):275–280, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schonhage:1971:SMG

[566] A. Schönhage and V. Strassen. Schnelle Multiplikation großer Zahlen. (German) [Fast multiplication of large numbers]. Computing, 7(3–4):281–292, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). Check pages??

Langmaack:1971:VEK

[567] H. Langmaack. Verbandstheoretische Einbettung von Klassen unwesentlich verschiedener Ableitungen in die Zopfgruppe. (German) [Group theoretical embedding of a class of insignificant various derivatives in the Zopf group]. Computing, 7(3–4):293–310, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hieber:1971:HSH

[568] L. Hieber and H. Rzehak. HYGOL, eine Sprache für hybride Programme. (German) [HYGOL, a language for hybrid programs]. Computing, 7(3–4): 311–323, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hieber:1971:SFS

[569] L. Hieber. Eine Sprache zur Formulierung spezieller Computergesteuerter Lehrprogramme. (German) [A language for formulating computerprogrammed instructions]. Computing, 7(3–4):324–332, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Niegel:1971:AAE

[570] W. Niegel. Ein Algorithmus zum aufbau einer endlichen Gruppe. (German) [An algorithm for the generation of a finite group]. Computing, 7(3–4): 333–343, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Meis:1971:DNE

[571] Th. Meis. Zur Diskretisierung nichtlinearer elliptischer Differentialgleichungen. (German) [On the application of finite difference methods to nonlinear elliptic differential equations]. Computing, 7(3–4):344–352, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pittnauer:1971:BMA

[572] F. Pittnauer. Eine Bemerkung über das Minimalglied asymptotischer Potenzreihen. (German) [A remark on the minimal term of asymptotic potential series]. Computing, 7(3–4):353–356, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wallach:1971:AIB

[573] Y. Wallach and A. Barlevi. Algorithm 15. Inversion of a blockwise tridiagonal matrix. Computing, 7(3–4):

357–363, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Spath:1971:ATD

[574] H. Späth. Algorithm 16. Two-Dimensional exponential splines. Computing, 7(3–4):364–369, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Handler:1971:BRR

[575] W. Händler. Book review: R. R. Arnold and H. C. Hill und A. V. Nichols, Modern Data Processing. Computing, 7(3–4):370, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1971:BRJ

[576] H. Kremser. Book review: I. J. Diner, M. P. Ganin, L. B. Komarow, A. A. Sweschnikow, K. B. Starobin und B. G. Wolodin, Wahrscheinlichkeitsrechnung und mathematische Statistik in Aufgaben (Mathematik für Technische Hochschulen: Band 9). Computing, 7 (3–4):370, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1971:BRG

[577] E. Bukovics. Book review: G. Hadley, Nichtlineare und dynamische Programmierung. Computing, 7(3– 4):371, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Collatz:1971:BRF

[578] L. Collatz. Book report: S. Filippi, Untersuchungen über die Fourier-Tschebyscheff-Approximation von Stammfunktionen. Computing, 7 (3–4):371, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1971:BRH

[579] E. Bukovics. Book review: H. P. Künzi und W. Krelle, Einführung in die mathematische Optimierung (Schriftenreihe des Betriebswissenschaftlichen Institutes der ETH, Zürich, Band 2). Computing, 7(3–4):372, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1971:BRJ

[580] E. Bukovics. Book review: A. Jäger und K. Wenke, Lineare Wirtachaftsalgebra (Teubner Studienbücher). Computing, 7(3–4):372, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kuich:1971:BRS

[581] W. Kuich. Book review: E. Schultze, Einführung in die mathematischen Grundlagen der Informations-theorie (Lecture Notes in Operations Research and Mathematical Economics: Vol. 9). Computing, 7(3–4):373, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pirzl:1971:BRHb

[582] I. Pirzl. Book report: H. P. Künzi und W. Krelle, Nichtlineare Programmierung: Neuere Verfahren (Lecture Notes in Operations Research and Mathematical Systems: Vol. 16). Computing, 7(3–4):373, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Maghsoodloo:1971:IHS

[583] S. Maghsoodloo. An investigation by high speed sampling of the frequency distribution of rank correlation. Computing, 8(1–2):1–12, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Guenther:1971:TMS

[584] Ronald Guenther. A transform method for solving certain systems of linear algebraic equations. Computing, 8 (1–2):13–18, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gessner:1971:AGD

[585] P. Gessner. Approximationsprobleme mit gewöhnlichen Differentialgleichungssystemen als Nebenbedingung. (German) [Approximation problems with systems of ordinary differential equations as side condition]. Computing, 8(1–2):19–31, March 1971. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Elsner:1971:VBS

[586] L. Elsner. Verfahren zur Berechnung des Spektralradius nichtnegativer irreduzibler Matrizen. (German) [Methods for computing the spectral radius of nonnegative irreducible matrices]. Computing, 8(1–2):32–39, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wenzel:1971:MMC

[587] A. Wenzel and Th. Scheidemandel. Multivariate Modelle in der chemischen Industrie. (German) [Multivariate models in chemical industry]. Computing, 8(1–2):40–61, March 1971. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hieber:1971:HSS

[588] L. Hieber. Ein Hybrid-System mit Stapelverarbeitung. (German) [A hybrid system with batch processing]. Computing, 8(1–2):62–73, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schutt:1971:GCA

[589] D. Schütt. A generalized concept of automaton automorphisms. Computing, 8(1–2):74–78, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bohme:1971:DVS

[590] J. F. Böhme. Diagnostische Vorgabeexperimente mit stochastischen Automaten. (German) [Diagnosing preset experiments with stochastic automata]. Computing, 8(1–2):79–98, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Noltemeier:1971:BBV

[591] H. Noltemeier. Ein Branch-and-Bound-Verfahren-Generator. (German) [A branch-and-bound program generator]. Computing, 8(1–2):99–106, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Scheiber:1971:BOA

[592] V. Scheiber. Berechnung der OC- und ASN-Funktionen geschlossener Sequentialtests bei Binomialverteilung. (German) [The computing of OC- and ASNfunctions of restricted sequential tests in binomial distributions]. Computing, 8(1–2):107–112, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tinhofer:1971:MZN

[593] G. Tinhofer. Mehrdimensionale Zuordnungsprobleme und Netzwerkflüsse mit vorgegebenen Wertemengen. (German) [Optimal assignment problems and network flows with given sets of values]. Computing, 8(1–2):113–120, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pirzl:1971:ONSa

[594] I. Pirzl. Optimierung unter Nebenbedingungen, Struktur einer Klasse von Algorithmen, I. (German) [Optimization under constraints. Structure of a class of algorithms. I]. Computing, 8 (1–2):121–142, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hofmann:1971:MRF

[595] W. Hofmann. Monotoniesätze für Regula-falsi- und Newton-Verfahren. (German) [Monotonicity theorems for Regula falsi and Newton’s method]. Computing, 8(1–2):143–156, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Allgower:1971:SRS

[596] E. L. Allgower and C. L. Keller. A search routine for a Sperner simplex. Computing, 8(1–2):157–165, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pirzl:1971:BEO

[597] I. Pirzl. Ein Beitrag zu Existenzkriterien in der Optimierungstheorie. (German) [On the existence theorem in the optimization theory]. Computing, 8 (1–2):166–170, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Braess:1971:BKP

[598] D. Braess. Die Bestimmung kürzester Pfade in Graphen und passende Datenstrukturen. (German) [The computation of the shortest path in graphs and appropriate data structures]. Computing, 8(1–2):171–181, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Avenhaus:1971:VEL

[599] J. Avenhaus. Ein Verfahren zur Einschliessung der Lösung des Anfangswertproblems. (German) [A method for bracketing solutions of initial value problems]. Computing, 8(1–2):182–190, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tinhofer:1971:FAN

[600] G. Tinhofer. Flüsse in antisymmetrischen Netzwerken. (German) [Flows in antisymmetric networks]. Computing, 8(1–2):191–199, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Spath:1971:BAG

[601] H. Späth. Berichtigung zu Algorithmus 10. (German) [Correction to Algorithm 10]. Computing, 8(1–2):200–201, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Baron:1971:BRG

[602] W. Baron. Book review: G. Lamprecht, Einführung in die Programmiersprache FORTRAN IV. Computing, 8(1–2):202, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brown:1971:BRM

[603] W. Brown. Book review: A. M. Bork, Programmierung und Benutzung des Computersystems IBM 1130. Computing, 8(1–2):202, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1971:BRH

[604] H. J. Stetter. Book review: H. E. Fettis und J. C. Caslin, A Table of the Complete Elliptic Integral of the First Kind for Complex Values of the Modulus. III. Auxiliary Tables. Computing, 8(1–2):202, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1971:NHP

[605] E. Bukovics. Notes: Hern Prof. Dr. Rudolf Inzinger zum 65, Geburtstag. (German) [Prof. Dr. Rudolf Inzinger’ 65th birthday]. Computing, 8(3–4): 203–207, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmidt:1971:ELN

[606] J. W. Schmidt. Eingrenzung von Lösungen nichtlinearer Gleichungen durch Verfahren mit höherer Konvergenzgeschwindigkeit. (German) [Enclosure of solutions of nonlinear equations by a method with higher convergence rate]. Computing, 8(3–4): 208–215, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Esser:1971:BES

[607] Henning Esser. Bemerkungen zu einem Satz von G. Freud über Quadraturkonvergenz. (German) [Remarks on a theorem of G. Freud concerning the quadrature convergence]. Computing, 8(3– 4):216–220, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Konigshofer:1971:MEI

[608] F. Königshofer. Das modifizierte Ergänzungsverfahren zur Invertierung von Matrizen. (German) [A modified method of completing for inversion of matrices]. Computing, 8(3–4):221–240, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Feldmann:1971:EEU

[609] H. Feldmann, H. Oberquelle, and C.-P. Ortlieb. Eine einfache universelle Turingmaschine in ALGOL 60 Simulation. (German) [A simple universal Turing machine in ALGOL 60 simulation]. Computing, 8(3–4):241–249, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Meinardus:1971:PCG

[610] G. Meinardus. Über ein Problem von L. Collatz. (German) [On a problem by L. Collatz]. Computing, 8(3–4): 250–254, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Engels:1971:ARE

[611] H. Engels. Zur Anwendung der Richardson-Extrapolation auf die numerische Differentiation. (German) [The application of the Richardsonextrapolation to numerical differentiation]. Computing, 8(3–4):255–271, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pirzl:1971:ONSb

[612] I. Pirzl. Optimierung unter Nebenbedingungen, Struktur einer Klasse von Algorithmen, II. (German) [Optimization under constraints. Structure of a class of algorithms. II]. Computing, 8(3–4):272–283, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lempio:1971:LOU

[613] F. Lempio. Lineare Optimierung in unendlichdimensionalen Vektorräunen. (German) [Linear optimization in infinitely dimensional vector spaces]. Computing, 8(3–4):284–290, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schindler:1971:IFG

[614] S. Schindler. Iterationsalgorithmen fuřr gewöhnliche Differentialgleichungen und ihre Eignung zur Parallelarbeit mittels Symbolmanipulation. (German) [Iterative algorithm for ordinary differential equation and their suitability for symbol manipulation].

Computing, 8(3–4):291–308, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Unger:1971:TDE

[615] C. Unger. Zur Theorie des Dumps an elektronischen Rechenanlagen. (German) [On the theory of dumping on computers]. Computing, 8(3–4): 309–319, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Barth:1971:NIG

[616] W. Barth. Nullstellenbestimmung mit der Intervallrechnung. (German) [Determination of zeros by interval computation]. Computing, 8(3–4):320–328, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Knodel:1971:VFI

[617] W. Knödel. Ein Verfahren zur Feststellung der Isomorphie von endlichen, zusaunmenhängenden Graphen. (German) [A method to state the isomorphism of finite, connected graphs]. Computing, 8(3–4):329–334, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dierieck:1971:PCA

[618] C. Dierieck and Y. Kamp. Polynomial Chebyshev approximation of a complex transfer function. Computing, 8(3–4): 335–342, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gorenflo:1971:DMA

[619] R. Gorenflo. Differenzenschemata monotoner Art für schwach gekoppelte Systeme parabolischer Differentialgleichungen mit gemischten Randbedingungen. (German) [Monotone difference formulas for weakly coupled system of parabolic differential equations with mixed boundary conditions]. Computing, 8(3–4):343–362, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zimmer:1971:SCA

[620] H. G. Zimmer. Some computational aspects of, and the use of computers in, algebraic number theory. Computing, 8(3–4):363–381, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Douglas:1971:ECE

[621] A. Douglas. Examples concerning efficient strategies for Gaussian elimination. Computing, 8(3–4):382–394, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Anonymous:1971:NNM

[622] Anonymous. Notes: New major contributions to international standardization by ECMA. Computing, 8 (3–4):395, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1971:BRF

[623] H. Kremser. Book review: S. Fernbach und A. Taub (Hg.), Computers and their Role in the Physical Sciences. Computing, 8(3–4):396, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Weinmann:1971:BRD

[624] A. Weinmann. Book review: D. Bell und A. W. J. Griffin(Hg), Regelkreistheorie und Datenverarbeitung. Computing, 8(3–4):396, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1971:BRM

[625] H. Kremser. Book review: M. A. R. Gunston practical matrix algebra. Computing, 8(3–4):397, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kuich:1971:BRH

[626] W. Kuich. Book review: H. Hermes, Aufzählbarkeit, Entscheidbarkeit, Berechenbarkeit, 2 Aufl. Computing, 8 (3–4):397, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1971:BRH

[627] H. Kremser. Book review: H. Müller-Merbach, On Round-Off Errors in Linear Programming (Lecture Notes in Operations Research and Mathematical Systems: Vol. 37). Computing, 8 (3–4):398, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Noltemeier:1971:BRS

[628] H. Noltemeier. Book review: Sensitivitätsanalyse bei diskreten linearen Optimierungsproblemen (Lecture Notes in Operations Research and Mathematical Systems: Vol. 30). Computing, 8(3–4):398, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Weiss:1971:BRH

[629] R. Weiss. Book review: H. Westermayer, Programmierlogik, Programmablaufpläne. (German) [H. Westermayer, logic of programming]. Computing, 8(3–4):398, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tewarson:1972:GEM

[630] R. P. Tewarson. On the Gaussian elimination method for inverting sparse matrices. Computing, 9(1):1–7, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kastlunger:1972:RKP

[631] K. H. Kastlunger and G. Wanner. Runge–Kutta processes with multiple nodes. Computing, 9(1):9–24, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dirschmid:1972:BEL

[632] H.-J. Dirschmid. Zur Berechnung der Eigenvektoren linearer Operatoren in endlichdimensionalen Räumen. (German) [The computation of eigenvectors of linear operators in finite dimensional spaces]. Computing, 9(1):25–36, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Riha:1972:OSP

[633] W. Riha. Optimal stability polynomials. Computing, 9(1):37–43, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Antes:1972:OVR

[634] H. Antes. Ober die vierdimensionale Romberg-Integration mit Schranken. (German) [On the four-dimensional Romberg-integration with bounds]. Computing, 9(1):45–52, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chen:1972:FWS

[635] Y. T. Chen and R. P. Tewarson. On the fill-in when sparse vectors are orthonormalized. Computing, 9(1):53–56, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tzschach:1972:CCH

[636] H. Tzschach. Construction of the convex hull for special sets. Computing, 9(1):57–61, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schincke:1972:EPS

[637] E. Schincke and J. Höpfner. Erzeugungen von Pseudozufallszahlenfolgen mit stückweise polygonaler Verteilung. (German) [Generation of pseudo random number sequences with piecemeal polynomial distribution]. Computing, 9(1):63–68, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Elsner:1972:VBS

[638] L. Elsner. Verfahren zur Berechnung des Spektralradius nichtnegativer irreduzibler Matrizen II. (German) [Methods for computing the spectral radius of a non-negative irreducible matrix. II]. Computing, 9(1):69–73, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1972:BRW

[639] E. Bukovics. Book review: W. E. Clason, Elsevier’s Dictionary of Computers, Automatic Control and Data Processing, in Six Languages. Computing, 9(1):75, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wolff:1972:BRC

[640] K. H. Wolff. Book review: C. L. Chiang, Introduction to Stochastic Processes In Biostatistics. Computing, 9(1):75, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hornich:1972:BRF

[641] H. Hornich. Book review: S. S. Feshchenko, N. I. Shkil’ und L. D. Nikolenko, Asymptotic Methods in the Theory of Linear Differential Equations. Computing, 9(1):76, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schauer:1972:BRH

[642] H. Schauer. Book review: H. Futh, EDV-Organisation. Computing, 9(1): 76, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1972:BRGa

[643] H. J. Stetter. Book review: E. Grochla, N. Szyperski und D. Seibt, Ausbildung und Fortbildung in der automatisierten Datenverarbeitung. Computing, 9(1):76, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pirzl:1972:BRJ

[644] I. Pirzl. Book review: J. Abadie (Hg.), Integer and Nonlinear Programming. Computing, 9(1):77, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Troch:1972:BRD

[645] I. Troch. Book review: D. E. Hyndman, Analog and Hybrid Computing. Computing, 9(1):77, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bukovics:1972:BRS

[646] E. Bukovics. Book review: St. Fenyö, Moderne mathematische Methoden in der Technik, Band 2 (International Series of Numerical Mathematics, Vol. 11). Computing, 9(1):78, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schauer:1972:BRW

[647] H. Schauer. Book review: W. Kunz und H. Rittel, Systemanalyse und Informationsverabeitung in der Forschung. Computing, 9(1):78, March 1972. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Junginger:1972:BRG

[648] W. Junginger. Book review: G. Uebe, Optimale Fahrpläne (Lecture Notes in Operations Research and Mathematical Systems: Vol. 20). Computing, 9(1):79, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Junginger:1972:BRT

[649] W. Junginger. Book review: Th. M. Liebling, Graphentheorie in Plunungsund Tourenproblemen (Lecture Notes in Operations Research and Mathematical Systems: Vol. 21). Computing, 9(1):79, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hummel:1972:BRG

[650] H. Hummel. Book review: G. Feichtinger, Lernprozesse in stochastischen Automaten (Lecture Notes in Operations Research and Mathematical Systems: Vol. 24). Computing, 9(1):80, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pirzl:1972:BRH

[651] I. Pirzl. Book review: I. H. Mufti, Computational Methods in Optimal Control Problems (Lecture Notes in Operations Research and Mathematical Systems: Vol. 17). Computing, 9(1):80, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dobler:1972:BRM

[652] K. U. Dobler. Book review: M. L. Minsky, Endliche und Unendliche Maschinen. Computing, 9(1):81, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pirzl:1972:BRK

[653] I. Pirzl. Book review: K. Manteuff und E. Seiffart, Einführung in die lineare Algebra und lineare Optimierung. Computing, 9(1):81, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gunzenhauser:1972:BRS

[654] R. Gunzenhäuser. Book review: Susan Markle, Gute Lernschritte. Computing, 9(1):82, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Yoeli:1972:BRR

[655] M. Yoeli. Book review: R. L. Martin, Studies in Feedback-Shift-Register Synthesis of Sequential Machines. Computing, 9(1):82, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1972:BRR

[656] H. Kremser. Book review: R. Sauer und I. Szabó (Hg.), Mathematische Hilfsmittel des Ingenieurs. Computing, 9(1):83, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pirzl:1972:BRG

[657] I. Pirzl. Book review: G. L. Nernhauser, Einführung in die Praxis der dynamischen Programmierung. Computing, 9(1):84, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wolff:1972:BRN

[658] K. H. Wolff. Book review: E. Nievergelt, O. Müller, F. E. Schlaepfer und W. H. Landis, Praktische Studien zur Unternehmensforschung (Ökonometrie und Unternehmensforschung, Band XIV). Computing, 9(1):84, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Handler:1972:BRS

[659] W. Händler. Book review: A. Salomaa, Theory of Automata. Computing, 9(1):85, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dunham:1972:BMR

[660] C. B. Dunham. Best mean rational approximation. Computing, 9(2):87–93, June 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

deOliviera:1972:MIA

[661] G. N. de Oliviera. Matrix inequalities and the additive inverse eigenvalue problem. Computing, 9(2):95– 100, June 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Luhrs:1972:EKP

[662] J. G. Lührs. Ein Einschließungssatz zum Knapsack-Problem. (German) [An inclusion concerning to the knapsackproblem]. Computing, 9(2):101–105, June 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gautschi:1972:NRR

[663] W. Gautschi. Zur Numerik rekurrenter Relationen. (German) [Numerical aspects of recurrence relations]. Computing, 9(2):107–126, June 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hlawka:1972:OTG

[664] E. Hlawka and R. Mück. Ober eine transformation von gleichverteilten Folgen II. (German) [On a transformation of uniformly distributed sequences. II]. Computing, 9(2):127–138, June 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tinhofer:1972:BKE

[665] G. Tinhofer. Über die Bestimmung von Kernen in endlichen Graphen. (German) [On the definition of kernels in finite graphs]. Computing, 9(2):139–147, June 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rokne:1972:ECL

[666] J. Rokne. Explicit calculation of the Lagrangian interval interpolating polynomial. Computing, 9(2):149–157, June 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kussmaul:1972:CCQ

[667] R. Kussmaul. Clenshaw–Curtis quadrature with a weighting function. Computing, 9(2):159–164, June 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Weicker:1972:BTT

[668] R. Weicker. Berichtigung zu: Tabulator-Turingmaschine und Komplexität. (German) [Amendment to: Tabulator-Turing machine and complexity]. Computing, 9(2):165–167, June 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [564].

Schwanhausser:1972:MVN

[669] G. Schwanhäußer. Mitteilung: Verbindung einer Netzplantechnik mit der Vektortransformation. (German) [Note: Combination of a network technique with vector transformation]. Computing, 9(2):169–170, June 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kuich:1972:BRM

[670] W. Kuich. Book review: M. Gross und A. Lentin, Mathematische Linguistik. Computing, 9(2):171, June 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kuich:1972:BRV

[671] W. Kuich. Book review: V. Claus, Stochastische Automaten. Computing, 9(2):171, June 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Weinmann:1972:BRK

[672] A. Weinmann. Book review: K. Bauknecht und W. Nef (Hg.), Digitale Simulation (Lecture Notes in Operations Research and Mathematical Systems: Vol. 51). Computing, 9(2):171, June 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Albrecht:1972:BRP

[673] R. Albrecht. Book review: P. H. Starke, Abstrakte Automaten. Computing, 9(2):172, June 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Collatz:1972:BRJ

[674] L. Collatz. Book review: J. Stoer und C. Witzgall, Convexity and Optimization in Finite Dimension I (Die Grundlehren der mathematischen Wissenschaften: Band 163). Computing, 9(2):172, June 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Collatz:1972:BRH

[675] L. Collatz. Book review: H. Werner, Praktische Mathematik (Mathematics Scripte 1). Computing, 9(2):173, June 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Thumb:1972:BRS

[676] N. Thumb. Book review: S. I. Suchowitzki und I. A. Radtschik, Mathematische Methoden der Netzplantechnik. Computing, 9(2):173, June 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rybarz:1972:BRK

[677] J. Rybarz. Book review: K.-H. Wolff, Versicherungsmathematik. Computing, 9(2):174, June 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Voith:1972:CGI

[678] R. P. Voith, W. G. Vogt, and M. H. Mickle. On the computation of the generalized inverse by classical minimization. Computing, 9(3):175–187, September 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Werner:1972:MMP

[679] B. Werner. Min-Max-Prinzipien für nichtlinear auftretende Eigenwerte bei klassischen und natürlichen Randeigenwertproblemen. (German) [Min-max-principles for the eigenvalues of a nonlinear eigenvalue problem]. Computing, 9(3):189–201, September 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Antes:1972:ORI

[680] H. Antes. Ober die Romberg-Integration in n Dimensionen. (German) [On the Romberg-integration in n dimensions]. Computing, 9(3):203–209, September 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wagner:1972:BAM

[681] E. G. Wagner. Bounded action machines. II: The basic structure of tapeless computers. Computing, 9(3): 211–232, September 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Linden:1972:SEK

[682] Hj. Linden. Über stabile Eigenwerte und die Konvergenz des Weinstein– AronszajnBazley-Fox-Verfahrens bei nicht notwendig rein diskretem Spektrum. (German) [On stable eigenvalues and the convergence of the Weinstein– AronszajnBazley-Fox methods for a not necessarily pure discrete spectrum]. Computing, 9(3):233–244, September 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chen:1972:OCP

[683] Y. T. Chen and R. P. Tewarson. On the optimal choice of pivots for the Gaussian elimination. Computing, 9 (3):245–250, September 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dorfler:1972:BMM

[684] W. Dörfler and J. Mühlbacher. Bestimmung eines maximalen Matching in beliebigen Graphen. (German) [Determination of a maximal matching in arbitrary graphs]. Computing, 9(3): 251–257, September 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Paul:1972:BRW

[685] M. Paul. Book review: W. W. Black, An Introduction to On-Line Computers. Computing, 9(3):259, September 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Weiss:1972:BRJ

[686] R. Weiss. Book review: J. Dörr und G. Hotz (Hg.) Automatentheorie und formale Sprache. Computing, 9(3):260, September 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dunham:1972:AAF

[687] C. B. Dunham. Approximation by alternating families on subsets. Computing, 9(4):261–265, December 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wynn:1972:CAM

[688] P. Wynn. Convergence acceleration by a method of intercalation. Computing, 9(4):267–273, December 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wacker:1972:ILS

[689] H. Wacker. Ein Iterationsverfahren zur Lösung speieller nichtlinearer Randwertprobleme. (German) [An iteration method to solve special nonlinear boundary value problems]. Computing, 9(4):275–291, December 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fleissner:1972:SLOa

[690] P. Fleissner. Stabilität linearer ökonometrischer Modelle I. (German) [Stability of linear econometric models I]. Computing, 9(4):293–315, December 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kastlunger:1972:TTI

[691] K. Kastlunger and G. Wanner. On Turan type implicit Runge–Kutta methods. Computing, 9(4):317–325, December 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Barth:1972:ABA

[692] W. Barth. Ein Algorithmus zur Berechnung aller rellen Nullstellen in einem Interval. (German) [An algorithm to complete all real roots in an interval]. Computing, 9(4):327–333, December 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Krabs:1972:BET

[693] W. Krabs. Bemerkungen zur einseitigen Tschebyscheff-Approximation. (German) [Remarks on one-sided Chebyshev approximation]. Computing, 9(4):335–342, December 1972. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Spremann:1972:FBM

[694] K. Spremann. Ein funktionalanalytischer Beweis des Maximumprinzips von Pontrjagin und dessen Verwendung zur Herleitung der Politikiteration von Howard. (German) [A proof of Pontrjagin’s maximum principle by methods of functional analysis and its application to deduce Howard’s policy iteration]. Computing, 9(4):343–353, December 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Feilmeier:1972:NAE

[695] M. Feilmeier. Numerische Aspekte bei der Einbettung nichtlinearer Probleme. (German) [Numerical aspects in the imbedding of nonlinear problems]. Computing, 9(4):355–364, December 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmidt:1972:SWC

[696] G. Schmidt. Shock waves of a continuous model of traffic flow. Computing, 9(4):365–381, December 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Scheiber:1972:VSO

[697] V. Scheiber. Versuchspläne zur stochastischen Optimierung. (German) [Experimental designs for stochastic optimization]. Computing, 9(4): 383–399, December 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fortelny:1972:BRG

[698] N. Fortelny. Book review: G. Hotz und V. Claus, Automatentheorie und Formale Sprachen. Computing, 9(4):401, December 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schulz:1972:BRC

[699] A. Schulz. Book review: C. O. Mertin, Datenerfassung – eine Übersicht. Computing, 9(4):401, December 1972. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1972:BRK

[700] H. Kremser. Book review: K. Hinderer, Foundations of Non-stationary Dynamic Programming with Discrete Time Parameter (Lecture Notes in Operations Research and Mathematical Systems: Vol. 33). Computing, 9 (4):402, December 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1972:BRB

[701] H. Kremser. Book review: B. Brosowski und E. Martensen (Hg.), Methoden und Verfahren der mathematischen Physik. Computing, 9 (4):402, December 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Paul:1972:BRK

[702] M. Paul. Book review: K. Anke, H. Kaltenecker und R. Oetker, Prozeßrechner. Wirkungsweise und Einsatz. Computing, 9(4):403, December 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Strom:1972:MNL

[703] Torsten Ström. Minimization of norms and logarithmic norms by diagonal similarities. Computing, 10(1–2):1–7, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Burgmeier:1972:MOB

[704] J. W. Burgmeier and M. R. Scott. A method for obtaining bounds on eigenvalues and eigenfunctions by solving non-homogeneous integral equations. Computing, 10(1–2):9–22, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Herzberger:1972:NNB

[705] J. Herzberger. Über die Nullstellenbestimmung bei näherungsweise Berechneten Funktionen. (German) [On the determination of zeros of approximated functions]. Computing, 10(1–2):23–31, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fleissner:1972:SLOb

[706] P. Fleissner and K. Hietler. Stabilität linearer ökonometrischer Modelle. II. (German) [Stability of linear econometric models. II]. Computing, 10(1–2):33–62, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schneider:1972:NEP

[707] K. R. Schneider. Näherungsverfahren zur Ermittlung periodischer Lösungen von Systemen nichtlinearer periodischer Differentialgleichungen. (German) [Approximation method for determination of periodic solutions of systems of nonlinear periodic differential equations]. Computing, 10(1–2):63–82, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Engels:1972:AGQ

[708] H. Engels. Über allgemeine Gauss’sche Quadraturen. (German) [On ordinary Gauss quadrature]. Computing, 10 (1–2):83–95, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Miller:1972:APR

[709] W. Miller. Automatic a priori roundoff analysis, I. Computing, 10(1–2):97– 106, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bauersfeld:1972:ABA

[710] G. Bauersfeld, Ch. Essmann, and H. Lóhle. Algorithmus zur Bestimmung von Abständen in endlichen, gerichteten und unbetwerteten Graphen. (German) [An algorithm to compute distances in finite, directed and nonvalued graphs]. Computing, 10(1–2):107– 109, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schroder:1972:CQR

[711] Gunnar Schröder. Charakterisierung des quasilinearen Raumes I(R) und Klassifizierung der quasilinearen Räume der Dimension 1 und 2. (German) [Characterization of quasilinear spaces I(R) and classification of quasilinear spaces of dimension 1 and 2]. Computing, 10(1–2):111–120, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Knauer:1972:NPGa

[712] B. Knauer. Normalformen planarer Graphen. I.. (German) [The normal form of planar graphs. I]. Computing, 10(1–2):121–136, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Knauer:1972:NPGb

[713] B. Knauer. Normalformen planarer Graphen. II. (German) [The normal form of planar graphs. II]. Computing, 10(1–2):137–152, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sieveking:1972:ADP

[714] M. Sieveking. An algorithm for division of power series. Computing, 10 (1–2):153–156, March 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Paredaens:1973:FSA

[715] J. J. Paredaens. Finite stochastic automata with variable transition probabilities. Computing, 11(1):1–20, March 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gustafson:1973:CFI

[716] S.-Å. A. Gustafson and Ingrid Melinder. Computing Fourier integrals by means of near-optimal quadrature rules of the Lagrangian type. Computing, 11(1):21–26, March 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Cheng:1973:NMB

[717] K. Y. Cheng. Note on minimizing the bandwidth of sparse, symmetric matrices. Computing, 11(1):27–30, March 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tewarson:1973:DFS

[718] R. P. Tewarson and K. Y. Cheng. A desirable form for sparse matrices when computing their inverse in factored forms. Computing, 11(1):31–38, March 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Piessens:1973:ISI

[719] R. Piessens and A. Haegemans. Inversion of some irrational Laplace transforms. Computing, 11(1):39–43, March 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tinhofer:1973:KAA

[720] G. Tinhofer. Der Out-Of-Kilter-Algorithmus für antisymmetrische Netzwerke. (German) [The out-of-kilter-algorithm for antisymmetric networks]. Computing, 11(1):45–50, March 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Yohe:1973:IBS

[721] J. M. Yohe. Interval bounds for square roots and cube roots. Computing, 11(1):51–57, March 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kilberth:1973:RKS

[722] K. Kilberth. Eine Randbedingung für kubische Splinefunktionen. (German) [A boundary condition for cubic spline functions]. Computing, 11(1):59–67, March 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Findler:1973:CMC

[723] N. V. Findler and E. Goit. A computerized matchmaker that is capable of learning. Computing, 11(1):69–72, March 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kracht:1973:ILR

[724] M. Kracht and Gunnar Schröder. Zur Intervallrechnung in linearen Räumen. (German) [Toward interval calculation in linear spaces]. Computing, 11(1):73– 79, March 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wanner:1973:RKM

[725] G. Wanner. Runge–Kutta-methods with expansion in even powers of h. Computing, 11(1):81–85, March 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kaminger:1973:GRP

[726] F. P. Kaminger. Generation, recognition and parsing of context-free languages by means of recursive graphs. Computing, 11(1):87–96, March 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Findler:1973:BRF

[727] N. V. Findler. Book review: F. Nake and A. Rosenfeld (Hrsg.), Graphic Languages. Computing, 11(1):97, March 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Will:1973:BRHa

[728] H. Will. Book review: H. Langmaack und M. Paul (Hrsg.), 1. Fachtagung über Programmiersprachen. Computing, 11(1):97, March 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kuich:1973:BRJ

[729] W. Kuich. Book review: J. Loeckx, Computability and Decidability. Computing, 11(1):98, March 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Will:1973:BRHb

[730] H. Will. Book review: H. Kamp und H. Pudlatz, Einführung in die Programmiersprache PL/1. Computing, 11(1):98, March 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brockhaus:1973:BRC

[731] M. Brockhaus. Book review: C. A. R. Hoare und R. H. Perrot (Hrsg.), Operating Systems Techniques. Computing, 11(1):99, March 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Findler:1973:BRN

[732] N. V. Findler. Book review: A. Newell, et al., Speech Understanding Systems. Computing, 11(1):99, March 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Weinmann:1973:BRW

[733] A. Weinmann. Book review: W. Weber, Adaptive Regelungssysteme. Computing, 11(1):99, March 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Anonymous:1973:AA

[734] Anonymous. Ankündigungen/Announcements. Computing, 11(1):101–102, March 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Cheng:1973:MBS

[735] K. Y. Cheng. Minimizing the bandwidth of sparse symmetric matrices. Computing, 11(2):103–110, June 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Adachi:1973:CVM

[736] N. Adachi. On the convergence of variable-metric methods. Computing, 11(2):111–123, June 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Domschke:1973:ZVS

[737] W. Domschke. Zwei Verfahren zur Suche negative Zyklen in bewerteten Digraphen. (German) [Two algorithms to detect negative cycles in a valued digraph]. Computing, 11(2):125–136, June 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dieter:1973:CMG

[738] U. Dieter and J. H. Ahrens. A combinatorial method for the generation of normally distributed random numbers. Computing, 11(2):137–146, June 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Singhal:1973:ASR

[739] K. Singhal and J. Vlach. Accuracy and speed of real and complex interpolation. Computing, 11(2):147–158, June 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bauersfeld:1973:AAF

[740] G. Bauersfeld, Ch. Essmann, and H. Löhle. Algorithmen: Algorithmus zur Feststellung der Isomorphie von endlichen, zusammenhängenden Graphen. (German) [An algorithm to find isomorphism to finite and connected graphs]. Computing, 11(2):159– 168, June 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Domschke:1973:AAP

[741] W. Domschke. Algorithmen: ALGOL-Prozeduren für zwei Verfahren zur Suche negativer Zyklen in bewerteten Diagraphen. (German) [ALGOL-procedures for two methods for detecting negative cycles in valued digraphs]. Computing, 11(2):169–173, June 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Spath:1973:ACO

[742] H. Späth. Algorithmen: Clustering of one-dimensional ordered data. Computing, 11(2):175–177, June 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kuich:1973:BRK

[743] W. Kuich. Book review: K. H. Bóhling und G. Dittrich, Endliche Stochastische Automaten. Computing, 11(2):179, June 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Meissl:1973:BRP

[744] P. Meissl. Book review: P. B. Hagelschurer, Theorie der linearen Dekomposition (Lecture Notes in Operations Research and mathematical systems: Vol. 58). Computing, 11(2):179, June 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Scheiber:1973:BRC

[745] V. Scheiber. Book report: C. Th., Ehlers, N. Hollberg und A Proppe (hrsg.), Computer: Werkzeug der Medizin. Computing, 11(2):179, June 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kuich:1973:BRP

[746] W. Kuich. Book review: P. H. Starke, Abstract Automata. Computing, 11(2):180, June 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Strassen:1973:BPA

[747] Volker Strassen. Berechnungen in partiellen Algebren endlichen Typs. (German) [Computational complexity in partial algebras of finite type]. Computing, 11(3):181–196, September 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gerber:1973:NDS

[748] P. Dean Gerber and Willard L. Miranker. Nonlinear difference schemes for linear partial differential equations. Computing, 11(3):197–211, September 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bundke:1973:MDA

[749] Werner Bundke. Multiprogramming für DV-Anlagen mittlerer Leistung. (German) [Multiprogramming for computers of average performance]. Computing, 11(3):213–220, September 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Miranker:1973:NMB

[750] Willard L. Miranker. Numerical methods of boundary layer type for stiff systems of differential equations. Computing, 11(3):221–234, September 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Perl:1973:AGA

[751] Jürgen Perl. Anwendung von Graphenalgorithmen auf allgemeinere Problemklassen. (German) [Graph algorithms and their application to more general problem classes]. Computing, 11(3):235–247, September 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hahn:1973:RBA

[752] Willi Hahn. Rundungsfehler beim “All-Integer Integer Programming Algorithm” von Gomory. (German) [Rounding error in Gomory’s “all-integer integer programming algorithm”]. Computing, 11(3):249–254, September 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wittenbrink:1973:HOP

[753] K. A. Wittenbrink. High order projection methods of moment-and collocation-type for nonlinear boundary value problems. Computing, 11 (3):255–274, September 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Domschke:1973:TNA

[754] Wolfgang Domschke. Two new algorithms for minimal cost flow problems. Computing, 11(3):275–285, September 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hairer:1973:MMM

[755] E. Hairer and G. Wanner. Multistepmultistage-multiderivative methods for ordinary differential equations. Computing, 11(3):287–303, September 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Florian:1973:BRD

[756] H. Florian. Book report: S. Dworatschek, unter Mitarbeit von Gehring, H. Schaltalgebra und digitale Grundschaltungen. Computing, 11(3):305, September 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Florian:1973:BRP

[757] H. Florian. Book review: E. P. Billeter, Grundlagen der repräsentativen Statistik. Computing, 11(3):305, September 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1973:BRR

[758] J. Stetter. Book review: R. D. Grigorieff, Numerik gewóhnlicher Differentialgleichungen. Band 1. Computing, 11(3):305, September 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1973:BRJ

[759] J. Stetter. Book review: J. D. Lambert, Computational Methods in Ordinary Differential Equations. Computing, 11 (3):306, September 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Forster:1973:FGV

[760] P. Forster. Fehlerabschätzungen zum Galerkin-Verfahren für nicht lineare Randwertaufgaben mit Polynomen als Ansatzfunktionen. (German) [Error bounds for the Galerkin procedure for nonlinear boundary value problems using polynomials as coordinate functions]. Computing, 11(4):307–313, December 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Joubert:1973:SED

[761] G. R. Joubert. Stable explicit difference approximation of the one-dimensional diffusion equation. Computing, 11(4): 315–325, December 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Spiess:1973:UTN

[762] Jürgen Spiess. Uniqueness theorems for nonlinear L_{2}-approximation problems. Computing, 11(4):327–335, December 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Jurgensen:1973:BUZ

[763] Helmut Jürgensen and Peter Wick. Bestimmung der Unterhalbgruppenverbände für zwei Klassen endlicher Halbgruppen. (German) [Computing the lattices of subsemigroups for two types of finite semigroups]. Computing, 11(4):337–351, December 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kreifelts:1973:OBG

[764] Thomas Kreifelts. Optimale Basiswahl für eine Gleitkomma-Arithmetik. (German) [Optimal choice of basis for a floating-point arithmetic]. Computing, 11(4):353–363, December 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See correction [883].

Krawczyk:1973:FNG

[765] R. Krawczyk. Fehlerabschätzung bei nichtlinearen Gleichungssystemen. (German) [Error estimation of nonlinear systems of equations]. Computing, 11(4):365–377, December 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Alefeld:1973:KVT

[766] Goetz Alefeld. Zur Konvergenz des Verfahrens der tangierenden Hyperbeln und des Tschebyscheff-Verfahrens bei konvexen Abbildungen. (German) [Convergence of Chevyshev’s method and the method of tangential hyperbolas for convex mappings]. Computing, 11(4):379–390, December 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tornig:1973:MRD

[767] W. Törnig. Monotonie- und Randmaximumsätze bei Diskretisierungen des Dirichletproblems allgemeiner nichtlinearer elliptischer Differentialgleichungen. (German) [Monotonicity and maximum principle for discretizations of the Dirichlet-problem of general nonlinear elliptic differential equations]. Computing, 11(4):391–401, December 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Micchelli:1974:CCA

[768] C. A. Micchelli. Characterization of Chebyshev approximation by weak Markhoff systems. Computing, 12(1): 1–8, March 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ichida:1974:NII

[769] K. Ichida and T. Kiyono. Numerical integration in the irregular region. Computing, 12(1):9–15, March 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hoffmann:1974:VDI

[770] K.-H. Hoffmann and I. Kolumbán. Verallgemeinerte Differenzierbarkeitsbegriffe und ihre Anwendung in der Optimierungstheorie. (German) [Generalized differentiability and its application in optimization theory]. Computing, 12(1):17–41, March 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bibel:1974:AST

[771] W. Bibel. An approach to a systematic theorem proving procedure in firstorder logic. Computing, 12(1):43–55, March 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eckhardt:1974:ILU

[772] U. Eckhardt. Ein Iterationsverfahren für lineare Ungleichungssysteme. (German) [An iterative method for linear inequalities]. Computing, 12(1):57–66, March 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fischer:1974:HAS

[773] H. Fischer. Hypernormbälle als abstrakte Schrankenzahlen. (German) [Hypernormballs as abstract number bounds]. Computing, 12(1):67–73, March 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brunner:1974:SNM

[774] H. Brunner and J. D. Lambert. Stability of numerical methods for Volterra integro-differential equations. Computing, 12(1):75–89, March 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kameda:1974:AMM

[775] T. Kameda and I. Munro. O(|V | · |E|) algorithm for maximum matching of graphs. Computing, 12(1):91–98, March 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hebgen:1974:SIP

[776] Michael Hebgen. Eine scalinginvariante Pivotsuche für Intervallmatrizen. (German) [A scaling invariant pivoting strategy for interval matrices]. Computing, 12(2):99–106, June 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hebgen:1974:IWO

[777] Michael Hebgen. Ein Iterationsverfahren, welches die optimale Intervall-Einschließung des Inversen eines M-Matrixintervalls liefert. (German) [An iterative method that produces the best possible interval inclusion of the inverse of a M-matrix interval]. Computing, 12(2):107–115, June 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Barth:1974:OLI

[778] W. Barth and E. Nuding. Optimale Lösung von Intervallgleichungssystemen. (German) [Optimal solution of systems of interval equations]. Computing, 12(2):117–125, June 1974. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Esser:1974:KMG

[779] H. Esser and K. Scherer. Konvergenzordnungen von ein- und Mehrschrittverfahren bei gewöhnlichen Differentialgleichungen. (German) [Orders of convergence of one and multistep methods for ordinary differential equations]. Computing, 12(2):127–143, June 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gaier:1974:AAT

[780] D. Gaier. Ableitungsfreie Abschätzungen bei trigonometrischer Interpolation und Konjugierten-Bestimmung. (German) [Derivative-free approximation by trigonometric interpolation and conjugate determination]. Computing, 12(2):145–148, June 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Miller:1974:RCR

[781] W. Miller. Remarks on the complexity of roundoff analysis. Computing, 12(2):149–161, June 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mohn:1974:FGN

[782] K. H. Mohn. Fehlerabschätzungen für gewisse Näherungskoeffizienten. (German) [Some error estimations for certain approximate coefficients]. Computing, 12(2):163–165, June 1974. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Werner:1974:KDF

[783] Jochen Werner. Über die Konvergenz des Davidon-Fletcher-Powell-Verfahrens für streng konvexe Minimierungsaufgaben im Hilbertraum. (German) [On the convergence of the Davidon-Fletcher-Powell-method for strictly convex minimization problems in Hilbert space]. Computing, 12(2):167–176, June 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ehrlich:1974:PBS

[784] L. W. Ehrlich. Point and block SOR applied to a coupled set of difference equations. Computing, 12(3): 181–194, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Merz:1974:EFS

[785] G. Merz. Erzeugende Funktionen bei Spline-Interpolation mit äquidistanten Knoten. (German) [Generating functions in spline interpolation with equidistant knots]. Computing, 12(3): 195–201, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Cremers:1974:EDL

[786] A. B. Cremers. Über eine explizite Darstellung von Listenstrukturen. (German) [On an explicit representation of list structures]. Computing, 12(3):203–213, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Herzberger:1974:MIN

[787] J. Herzberger. Über Matrixdarstellungen für Iterationsverfahren bei nichtlinearen Gleichungen. (German) [On matrix representations of some iterative methods for solving nonlinear equations]. Computing, 12(3):215–222, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ahrens:1974:CMS

[788] J. H. Ahrens and Ulrich Dieter. Computer methods for sampling from gamma, beta, Poisson and binomial distributions. Computing, 12(3): 223–246, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See note [1129].

Vogl:1974:UQB

[789] F. Vogl. Untersuchungen zur qualitativen Behandlung von linearen Differentialgleichungssystemen mit unstetiger rechter Seite mit Hilfe der Theorie der Distributionen. (German) [Qualitative treatment of systems of linear differential equations with discontinuous right hand sides by means of the theory of distribution]. Computing, 12(3):247–267, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mossige:1974:ASC

[790] S. Mossige. Algorithm 19: Stepcycle generation. Computing, 12(3): 269–272, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Weigand:1974:ALA

[791] M. M. Weigand. Algorithmus 20. Ein leistungsfähiger Algorithmus zur Bestimmung von k-kürzesten Wegen in einem Graphen. (German) [Algorithm 20. A powerful algorithm for the k-th best route in a graph]. Computing, 12 (3):273–283, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Barth:1974:BRJ

[792] W. Barth. Book review: O. J. Dahl, W. Dijkstra und C. A. R. Hoare, Structured Programming. Computing, 12 (3):285, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1974:BRV

[793] H. J. Stetter. Book review: V. Briatenberg, Gehirngespinste. Computing, 12(3):285, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Meissl:1974:BRP

[794] P. Meissl. Book review: P. Gessner und H. Wacker, Dynamische Optimierung. Computing, 12(3):286, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Troch:1974:BRW

[795] I. Troch. Book review: W. Frisch, Analogrechnen in der Kernreaktortechnik. Computing, 12(3):286, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Jeram:1974:BRM

[796] H. Jeram. Book review: M. J. A. Hoffmann, Datenfernverarbeitung. Computing, 12(3):287, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Junginger:1974:BRHb

[797] E. Junginger. Book review: H. Richter und V. Mammitzsch, Methode der kleinsten Quadrate. Computing, 12 (3):287, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Unger:1974:BRJ

[798] H. Unger. Book review: J. J. Stetter, Discretization for Ordinary Differential Equations. Computing, 12 (3):288, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Handler:1974:BRH

[799] W. Händler. Book review: H. Kunsemüller, Digitale Rechenanlagen. Computing, 12(3):289, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1974:BRHa

[800] J. J. Stetter. Book review: H. Werner und R. Schaack, PraktischeMathematik II. Computing, 12 (3):289, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1974:BRHb

[801] H. J. Stetter. Book review: H. Zemanek (Hrsg.), The Skyline of Information Processing. Computing, 12 (3):290, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Troch:1974:BRB

[802] I. Troch. Book review: B. R. Wilkins, Analogue and Iterative Methods in Computation, Simulation, and Control. Computing, 12(3):290, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Levi:1974:GIH

[803] G. Levi. Graph isomorphism: a heuristic edge-partitioning-oriented algorithm. Computing, 12(4):291–313, December 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Merle:1974:CED

[804] G. Merle and H. Späth. Computational experiences with discrete L_{P} - approximation. Computing, 12(4): 315–321, December 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pasedach:1974:VSA

[805] K. Pasedach. Ein Verfahren zur schnellen Approximation von Summen mit vielen Summanden. (German) [Method for the fast approximation of sums with many terms]. Computing, 12 (4):323–331, December 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mossige:1974:ACT

[806] Svein Mossige. An algorithm for computation of translation invariant groups and the subgroup lattice. Computing, 12(4):333–355, December 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pape:1974:NVK

[807] U. Pape. Netzwerk-Veränderungen und Korrektur kürzester Weglängen von einer Wurzelmenge zu allen anderen Knoten. (German) [Changes in networks and adjustment of the lengths of shortest paths]. Computing, 12(4): 357–362, December 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wettstein:1974:PBG

[808] H. Wettstein. Prozessumschaltungen in Betriebssystemen. (German) [Dispatching in operating systems]. Computing, 12(4):363–382, December 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schulz:1974:BRD

[809] A. Schulz. Book review: E. Dotzauer, Praktikum zur angewandten Informatik. Computing, 12(4):383, December 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Troch:1974:BRF

[810] I. Troch. Book review: O. Fóllinger, Nichtlineare Regelungen. Computing, 12(4):383, December 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dorfler:1974:BRH

[811] W. Dórfler. Book review: S. H. Gould, Russian for the Mathematician. Computing, 12(4):384, December 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lorenz:1974:BRK

[812] H. Lorenz. Book review: K. Ganzhorn, E. Schieferdecker und A. Endres (Hrsg.), Systemprogrammierung. Computing, 12(4):384, December 1974. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1974:BRDa

[813] H. J. Stetter. Book review: D. Greenspan, Discrete Models. Computing, 12(4):384, December 1974. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Albrecht:1974:BRT

[814] R. Albrecht. Book review: Th. Lutz und H. Klimesch, Die Datenbank im Informationssystem. Computing, 12 (4):385, December 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Klima:1974:BRK

[815] J. Klima. Book review: Ken-ichi Kojima, Mathematical Topics in Population Genetics. Computing, 12(4):385, December 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pressler:1974:BRC

[816] J. Pressler. Book review: C. Hackl, Schaltwerk- und Automatentheorie. Computing, 12(4):385, December 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Barth:1974:BRK

[817] W. Barth. Book review: K. H. Müller und I. Strecher, FORTRAN, Programmieranleitung, 2 Aufl. Computing, 12(4):386, December 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schulz:1974:BRP

[818] A. Schulz. Book review: P. Mertens, (Hrsg.), Angewandte Informatik. Computing, 12(4):386, December 1974. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hairer:1974:BGG

[819] E. Hairer and G. Wanner. On the Butcher group and general multivalue methods. Computing, 13(1):1–15, March 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kasack:1974:OLR

[820] U. Kasack. Orthogonale lineare Regression: Stichprobenparameter. (German) [Orthogonal linear regression: Sample parameters]. Computing, 13(1):17–20, March 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lunow:1974:GRN

[821] W. Lünow. On the generation of random numbers with at choice distribution. Computing, 13(1):21–31, March 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zielke:1974:TMK

[822] G. Zielke. Testmatrizen mit Maximaler Konditionszahl. (German) [Test matrices with maximal condition number]. Computing, 13(1):33–54, March 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Alefeld:1974:SKV

[823] G. Alefeld. Stets Konvergente Verfahren höherer Ordnung zur Berechnung von reellen Nullstellen. (German) [Always convergent higher order methods for the computation of a real zero]. Computing, 13(1):55–65, March 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brunner:1974:ASF

[824] H. Brunner. On the approximate solution of first-kind integral equations of Volterra type. Computing, 13(1):67–79, March 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tewarson:1974:GIR

[825] R. P. Tewarson and P. Narain. Generalized inverses and resolution in the solution of linear equations. Computing, 13(1):81–88, March 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Barth:1974:BRR

[826] W. Barth. Book review: A. R. Brown and W. A. Sampson: Program Debugging. Computing, 13(1):89–??, March 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1974:BRDb

[827] H. J. Stetter. Book review: D. H. Moore, Heaviside Operational Calculus. Computing, 13(1):89–??, March 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Urbanek:1974:BRJ

[828] F. J. Urbanek. Book review: J. Chinal: Design Methods for Digital Systems. Computing, 13(1):89–??, March 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schauer:1974:BRG

[829] H. Schauer. Book review: G. Niemeyer, Einführung in das Programmieren in Assembler. Computing, 13(1):90, March 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Will:1974:BRJ

[830] H. Will. Book review: J. Fox (Hrsg.), Proceedings of the Symposium on Computer-Communications Networks and Teletraffic, New York, 4.-6. April 1972. Computing, 13(1):90– ??, March 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Troch:1974:BRH

[831] I. Troch. Book review: H. H. Rosenbrock und C. Storey, Mathematik Dynamischer Systeme. Computing, 13(1): 91, March 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schauer:1974:BRJ

[832] H. Schauer. Book review: J. Schärf, BASIC für Anfänger. Computing, 13(1):91–??, March 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Uberhuber:1974:BRH

[833] Ch. W. Überhuber. Book review: H. Späth, Spline-Algorithmen zur Konstruction glatter Kurven under Flächen. Computing, 13(1):92, March 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Paredaens:1974:GDS

[834] J. J. Paredaens. A general definition of stochastic automata. Computing, 13(2):93–105, June 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Perl:1974:GUG

[835] J. Perl. Graphenalogrithmen und Überdeckungsprobleme. (German) [Graph algorithms and covering problems]. Computing, 13(2):107–113, June 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Braunleder:1974:RGV

[836] B. Braunleder. Ritz–Galerkin Verfahren für normale Operatoren im Hilbertraum. (German) [Ritz-galerkin methods for normal operators in Hilbert space]. Computing, 13(2):115– 120, June 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Engels:1974:PIT

[837] H. Engels. Positive Interpolationsquadraturen mit teilweise vorgeschriebenen Stützstellen. (German) [Positive interpolarity quadratures with partially preassigned nodes]. Computing, 13(2):121–141, June 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Heidt:1974:NLL

[838] M. Heidt. Die numerische Lösung linearer Differentialgleichungen zweiter Ordnung mit konstanten Koeffizienten. (German) [The numerical solution of linear second-order differential equations with constant coefficients]. Computing, 13(2):143–154, June 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schnorr:1974:ZLU

[839] C. P. Schnorr. Zwei lineare untere Schranken für die Komplexität Boolescher Funktionen. (German) [Two linear lower bounds on the complexity of Boolean functions]. Computing, 13(2):155–171, June 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Spremann:1974:BMI

[840] K. Spremann and P. Gessner. Bewertete Markovprozesse im Stationären Zustand — Ein neuer Algorithmus mit Beispeil. (German) [Ergodic Markov processes with rewards — a new algorithm]. Computing, 13(2):173–181, June 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Piessens:1974:AAI

[841] R. Piessens and Ann Haegemans. Algorithm for the automatic integration of highly oscillatory functions. Computing, 13(2):183–193, June 1974. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wimp:1974:CTP

[842] J. Wimp. On the computation of Tricomi’s Psi function. Computing, 13(3– 4):195–203, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dunham:1974:CA

[843] Ch. B. Dunham. Chebyshev approximation by A + B^{∗}Φ(CX). Computing, 13(3–4):205–213, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Abdelmalek:1974:SLL

[844] N. N. Abdelmalek. On the solution of the linear least squares problems and pseudo-inverses. Computing, 13(3–4): 215–228, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ichida:1974:IMF

[845] K. Ichida and T. Kiyono. Interpolation of multivariable functions with respect to random points. Computing, 13(3– 4):229–233, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dunham:1974:FLM

[846] Ch. B. Dunham. Failure of Loeb’s method for rational L_{1} approximation.

Computing, 13(3–4):235–237, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Duff:1974:PSR

[847] Iain S. Duff. Pivot selection and row ordering in Givens reduction on sparse matrices. Computing, 13(3–4): 239–248, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Jung:1974:BMW

[848] H. A. Jung. Über die Bestimmung minimaler Wege und Gerüste in Graphen. (German) [On the determination of minimal paths and spanning trees in graphs]. Computing, 13(3–4):249–252, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mittelmann:1974:ALG

[849] H. D. Mittelmann. Die Approximation der Lösungen Gemischter Randwertprobleme quasilinearer elliptischer Differentialgleichungen. (German) [Approximate solution of mixed boundary value problems for quasilinear elliptic differential equations]. Computing, 13(3–4):253–265, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kress:1974:QUI

[850] R. Kress. Zur Quadratur uneigentlicher Integrale bei Analytischen Funktionen. (German) [On the quadrature of improper integrals for analytic functions]. Computing, 13(3–4):267–277, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Oesterhelt:1974:MNI

[851] G. Oesterhelt. Mehrschrittverfahren zur numerischen Integration von Differentialgleichungssystemen mit stark verschiedenen Zeitkonstanten. (German) [Multistep methods for the numerical integration of systems of differential equations with widely separated time constants]. Computing, 13(3–4): 279–298, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hauenschild:1974:AKK

[852] M. Hauenschild. Arithmetiken für komplexe Kreise. (German) [Arithmetic procedures for complex circles]. Computing, 13(3–4):299–312, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ratschek:1974:PIG

[853] H. Ratschek. Über das Produkt von Intervallpolynomen. (German) [About the product of interval polynomials]. Computing, 13(3–4):313–325, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Meister:1974:IDO

[854] G. Meister. Über die Integration von Differentialgleichungssystemen: 1. Ordnung mit exponentiell angepassten numerischen Methoden. (German) [On the integration of systems of first order differential equations: 1. exponentially fitted numerical methods]. Computing, 13(3–4):327–352, September 1974.

CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fischer:1974:LKE

[855] F. D. Fischer. Zur Lösung des Kontaktproblems elastischer Körper mit ausgedehnter Kontaktfläche durch quadratische Programmierung. (German) [Solution of the contact problem of elastic bodies with extended contact areas by quadratic programming]. Computing, 13(3–4):353–384, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Knodel:1975:EB

[856] W. Knödel. Erich Bukovics: 1921– 1975. Computing, 14(1–2):1–3, March 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rokne:1975:RDI

[857] Jon Rokne. Reducing the degree of an interval polynomial. Computing, 14(1–2):5–14, March 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hecht:1975:LTT

[858] V. Hecht and H. Rzehak. Languages and translation techniques for extensive use of parallel operations. Computing, 14(1–2):15–27, March 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Forst:1975:BFN

[859] W. Forst. Zur Bestimmung der Frobeniuswurzel nichtnegativer Matrizen. (German) [Determination of the Frobenius root of nonnegative matrices]. Computing, 14(1–2):29–35, March 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sautter:1975:PFP

[860] W. Sautter. A Posteriori-Fehlerabschätzungen für die Pseudoinverse und die Lösung minimaler Länge. (German) [A posteriori error bounds for the pseudoinverse and the minimal-length solution]. Computing, 14(1–2):37–44, March 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wisskirchen:1975:VII

[861] P. Wisskirchen. Vergleich intervallarithmetischer Iterationsverfahren. (German) [Comparison of interval arithmetic iteration methods]. Computing, 14(1–2):45–49, March 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ullrich:1975:BNR

[862] Ch. Ullrich. Über die beim numerischen Rechnen mit komplexen Zahlen und intervallen vorliegended mathematischen Strukturen. (German) [On the mathematical structures occurring in numerical computations with complex numbers and intervals]. Computing, 14 (1–2):51–65, March 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pohst:1975:BUE

[863] M. Pohst. Berechnung unabhängiger Einheiten und Klassenzahlen in total reellen biquadratischen Zahlkörpern. (German) [Computation of independent units and class numbers in totally real quartic fields]. Computing, 14(1–2):67–78, March 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ritter:1975:APM

[864] K. Ritter. Accelerating procedures for methods of conjugate directions. Computing, 14(1–2):79–105, March 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Takasawa:1975:ASO

[865] Y. Takasawa and H. Yanai. Approximate solution of optimal allocation problem by a nomograph. Computing, 14(1–2):107–117, March 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Neumann:1975:LNK

[866] K. Neumann. Zur Lösung nichtlinearer Kontrollprobleme mit Hilfe eines Gradientenverfahrens. (German) [Solution of nonlinear control problems with a gradient method]. Computing, 14 (1–2):119–130, March 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ziegler:1975:IIA

[867] B. Ziegler. Inversion einer im allgemeinen nicht positiv definiten symmetrischen Matrix. (German) [Inversion of a matrix which is symmetric but in general not positive definite]. Computing, 14(1–2):131–139, March 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Thieler:1975:ABF

[868] P. Thieler. Eine Anwendung des Brouwerschen Fixpunktsatzes in der Intervallarithmetik der Matrizen. (German) [An application of the Brouwer fixed point theorem in matrix interval arithmetic]. Computing, 14(1– 2):141–147, March 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mossige:1975:TSP

[869] S. Mossige. Translation sets of permutations: Enumeration of selfcomplementary step-cycles. Computing, 14 (1–2):149–152, March 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

ElTom:1975:EAV

[870] M. E. A. El Tom. Efficient algorithms for Volterra integral equations of the second kind. Computing, 14 (1–2):153–166, March 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

VandeVel:1975:MCR

[871] H. Van de Vel. Method for computing a root of a single nonlinear equation, including its multiplicity. Computing, 14(1–2):167–171, March 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Delvos:1975:KSH

[872] F. J. Delvos and H. W. Kösters. Zur Konstruktion von M-Splines höheren Grades. (German) [On the construction of higher degree M-splines]. Computing, 14(1–2):173–182, March 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

vanEmdeBoas:1975:NRH

[873] P. van Emde Boas. The nonrenameability of honesty classes. Computing, 14(1–2):183–193, March 1975.

CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

James:1975:AAD

[874] K. R. James and W. Riha. Algorithm 24. algorithm for deriving the chromatic polynomial of a graph. Computing, 14(1–2):195–203, March 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See remark [942].

Brezinski:1975:NSQ

[875] C. Brezinski. Numerical stability of a quadratic method for solving systems of non linear equations. Computing, 14 (3):205–211, September 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ottmann:1975:RGD

[876] Th. Ottmann. Mit regulären Grundbegriffen definierbare Prädikate. (German) [On predicates definable by regular concepts]. Computing, 14(3): 213–223, September 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Strehmel:1975:NLA

[877] K. Strehmel and V. Hoffmann. Numerische Lösung von Anfangwertaufgaben gewöhnlicher Differentialgleichungen n-ter Ordnung. (German) [Numerical solution of initial value problems for ordinary differential equations of n-th order]. Computing, 14 (3):225–234, September 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Knauff:1975:OAN

[878] W. Knauff. Optimal Approximation mit Nebenbedingungen an lineare Funktionale auf H^{2}(E_{ρ}) und L^{2}(E_{ρ}). (German) [Optimal approximation with side conditions to linear functionals on H^{2}(E_{ρ}) and L^{2}(E_{ρ})]. Computing, 14(3):235–250, September 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mackenroth:1975:EBD

[879] U. Mackenroth. Einige Bemerkungen zur Dualität in der konvexen Optimierung. (German) [Some remarks on duality in convex optimization]. Computing, 14(3):251–259, September 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Attmann:1975:NEF

[880] Josef Attmann and Willi Hahn. Numerische Erfahrungen mit der Filtermethode von Balas in der linearen binären Optimierung. (German) [Numerical experience with the filter method of Balas]. Computing, 14(3): 261–270, September 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Weller:1975:UKG

[881] Th. Weller. Übersetzung von Kommandosprachen. (German) [Translation of command languages]. Computing, 14 (3):271–284, September 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Cherniavsky:1975:FIL

[882] J. C. Cherniavsky. Function iteration logics and flowchart schemata. Computing, 14(3):285–312, September 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See erratum [1172].

Kreifelts:1975:OBF

[883] T. Kreifelts. Optimale Basiswahl für eine Gleitkomma-Arithmetik (Berichtigung) (German) Optimal Basis Choice for a Floating-Point Arithmetic (Correction). Computing, 14(3):313–314, September 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [764].

Lorez:1975:BGB

[884] H. Lorez, F. J. Urbanek, H. Will, R. Weiss, W. Baron, et al. Buchbesprechungen (German) Book Reviews. Computing, 14(3):315–322, September 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [764].

Kulisch:1975:FIF

[885] U. Kulisch. Formalization and implementation of floating-point arithmetics. Computing, 14(4):323–348, December 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Meister:1975:QAG

[886] Bernd Meister. Queueing analysis of general loop systems. Computing, 14 (4):349–365, December 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Esser:1975:SQK

[887] H. Esser. Eine stets quadratische konvergente Modifikation des Steffensen-Verfahrens. (German) [An always quadratically convergent modifications of the Steffensen method]. Computing, 14(4):367–369, December 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fehlberg:1975:KRK

[888] Erwin Fehlberg. Klassische Runge– Kutta–Nyström-Formeln mit Schrittweiten-Kontrolle für Differentialgleichungen d^{2}x/dt^{2} = f(t, x, dx/dt). (German) [Classical Runge–Kutta-nystrom formula with stepsize control for differential equations d^{2}x/dt^{2} = f(t, x, dx/dt)]. Computing, 14(4): 371–387, December 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Burkard:1975:AVG

[889] Rainer E. Burkard and Herbert Enge. Algorithmus 25. Verfahren zur Gemischt-Ganzzahligen, konvexen Optimierung. (German) [Method for mixed integer convex programming]. Computing, 14(4):389–396, December 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Apostolatos:1975:MNF

[890] N. Apostolatos. Eine Methode zur Nullstellenbestimmung von Funktionen. (German) [A method to compute the zeros of functions]. Computing, 15(1):1–10, March 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bereanu:1975:UCP

[891] B. Bereanu. On the use of computers in planning under conditions of uncertainty. Computing, 15(1):11–32, March 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Syslo:1975:CES

[892] M. M. Syslo and J. Dzikiewicz. Computational experiences with some transitive closure algorithms. Computing, 15(1):33–39, March 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Neuhaus:1975:WAK

[893] W. Neuhaus and S. Schottländer. Die Weiterentwicklung der Aireyschen Konvergenzfaktoren für das Exponentialintegral zu einer Darstellung mit Restglied. (German) [Development of Airey’s converging factors of the exponential integral to a representation with remainder term]. Computing, 15(1):41–52, March 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Das:1975:OSD

[894] P. C. Das. One-sided directional derivative and applications. Computing, 15(1):53–65, March 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schroder:1975:SRH

[895] G. Schröder. Zur Struktur des Raumes der Hypernormbälle über einem linearen Raum. (German) [Structure of the space of hypernormballs on a linear space]. Computing, 15(1):67–70, March 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hojka:1975:UTL

[896] N. Hojka. On the unambiguity of ntuple languages. Computing, 15(1):71– 74, March 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dzikiewicz:1975:AAF

[897] J. Dzikiewicz. Algorithmus 26. an algorithm for finding the transitive closure of a digraph. Computing, 15(1):75–79, March 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Anonymous:1975:BR

[898] Anonymous. Book reviews. Computing, 15(1):81, 169, 287, 365, March 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zielke:1975:TFP

[899] Gerhard Zielke. Testmatrizen mit freien Parametern. (German) [Test matrices with free parameters]. Computing, 15(2):87–103, June 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Maurer:1975:AMS

[900] H. Maurer and W. Gillessen. Application of multiple shooting to the numerical solution of optimal control problems with bounded state variables. Computing, 15(2):105–126, June 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

King:1975:QFE

[901] J. Thomas King. Quasioptimal finite element method for elliptic interface problems. Computing, 15(2):127–135, June 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Feilmeier:1975:HKI

[902] M. Feilmeier. Hermitesche Kollokation bei Integralgleichungen. (German) [Solution of integral equations by Hermitecollocation]. Computing, 15(2):137– 146, June 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tinhofer:1975:BAE

[903] G. Tinhofer. Zur Bestimmung der Automorphismen eines endlichen Graphen. (German) [On the determination of the automorphisms of a finite graph]. Computing, 15(2):147–156, June 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brass:1975:ASL

[904] D. Bräss. Anregungen aus dem Schnittebenenverfahren für die Lösung des allgemeinen Transportproblems. (German) [Numerical method for the solution of the general transport problem, stimulated by the cutting plane method]. Computing, 15(2):157–167, June 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Riess:1975:GTQ

[905] R. D. Riess. Gauss-Turan quadratures of Chebyshev type and error formulae. Computing, 15(3):173–179, September 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Altherr:1975:AEA

[906] W. Altherr. Algorithm for enumerating all vertices of a convex polyhedron. Computing, 15(3):181–193, September 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Albrecht:1975:KMA

[907] R. Albrecht and P. Vachenauer. Über konsistente Mehrschrittverfahren für das abstrakte Cauchy-Problem. (German) [On consistent multistep methods for the abstract Cauchy-problem]. Computing, 15(3):195–203, September 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kosters:1975:BSO

[908] H. W. Kösters and K. H. Schlosser. Berechnung von L-Splines zu Operatoren zweiter Ordnung mit konstanten Koeffizienten. (German) [Computation of L-splines corresponding to operators of second order with constant coefficients]. Computing, 15(3): 205–216, September 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

vonHenke:1975:PRW

[909] F. W. von Henke, K. Indermark, G. Rose, and K. Weihrauch. On primitive recursive word functions. Computing, 15(3):217–234, September 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Locher:1975:KCE

[910] F. Locher. Konvergenzbeschleunigung von Cebysev-Entwicklungen. (German) [Accelerating the convergence of Chebyshev series]. Computing, 15(3): 235–246, September 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sommerhalder:1975:NAN

[911] R. Sommerhalder. On the number of automorphisms of non-cyclic automata. Computing, 15(3):247–250, September 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kreth:1975:NEV

[912] H. Kreth. Der Nachweis der Existenz verallgemeinerter Lössungen quasilinearer Anfangswertaufgaben mittels Differenzenverfahren. (German) [Proof for the existence of generalized solutions for quasilinear initial-value problems by finite-difference methods]. Computing, 15(3):251–261, September 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Goodman:1975:REP

[913] R. Goodman and A. Feldstein. Roundoff error in products. Computing, 15 (3):263–273, September 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

GarciaRuffinatti:1975:VBE

[914] J. Garcia Ruffinatti. Verfahren für die Bestimmung von Eigenwerten und Eigenverktoren von symmetrisch nicht-Hermiteschen Matrizen. (German) [Method for the determination of eigenvalues and eigenvectors of symmetric nonHermitian matrices]. Computing, 15(3):275–286, September 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Nickel:1975:SKNa

[915] K. Nickel. Über die Stabilität und Konvergenz numerischer Algorithmen: Teil 1. (German) [Stability and convergence of numeric algorithms: Part 1]. Computing, 15(4):291–309, December 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Nickel:1975:SKNb

[916] K. Nickel. Über die Stabilität und Konvergenz numerischer Algorithmen: Teil 2. (German) [Stability and convergence of numeric algorithms: Part 2]. Computing, 15(4):311–328, December 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Appelt:1975:VIG

[917] W. Appelt. Vergröberung von Intervallpolynomen. (German) [Reducing the degree of interval polynomials]. Computing, 15(4):329–346, December 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chawla:1975:QFC

[918] M. M. Chawla and N. Jayarajan. Quadrature formulas for Cauchy principal value integrals. Computing, 15 (4):347–355, December 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kydes:1975:IMS

[919] A. S. Kydes and R. P. Tewarson. Iterative method for solving partitioned linear equations. Computing, 15(4): 357–363, December 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Moore:1976:CRR

[920] R. E. Moore. On computing the range of a rational function of n variables over a bounded region. Computing, 16(1–2):1–15, March 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rokne:1976:ESZ

[921] Jon Rokne. Errorbounds for simple zeros of λ-matrices. Computing, 16 (1–2):17–27, March 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hegedus:1976:MCR

[922] C. J. Hegedus. Method for calculating the reciprocal coefficients of function matrices. Computing, 16(1–2):29–37, March 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fuchs:1976:SRK

[923] F. Fuchs. A-stability of Runge– Kutta methods with single and multiple nodes. Computing, 16(1–2):39–48, March 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hofmann:1976:KSV

[924] Wolf Hofmann and Heinrich Voss. Über die Konvergenz von Shooting-Verfahren. (German) [On the convergence of shooting methods]. Computing, 16(1–2):49–60, March 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rentrop:1976:NSS

[925] P. Rentrop. Numerical solution of the singular Ginzburg–Landau equations by multiple shooting. Computing, 16(1–2):61–67, March 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chamayou:1976:DAG

[926] J. M. F. Chamayou. On a direct algorithm for the generation of log-normal pseudo-random numbers. Computing, 16(1–2):69–76, March 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Osyczka:1976:AOS

[927] Andrzej Osyczka. An algorithm of optimization for a special class of networks. Computing, 16(1–2):77–97, March 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mehlhorn:1976:MSC

[928] Kurt Mehlhorn and Zvi Galil. Monotone switching circuits and Boolean matrix product. Computing, 16(1– 2):99–111, March 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Nagl:1976:FLL

[929] M. Nagl. Formal languages of labelled graphs. Computing, 16(1–2):113–137, March 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Weigand:1976:ANA

[930] Michael M. Weigand. Algorithmus 27. Ein neuer Algorithmus zur Bestimmung von k-kürzesten Wegen in einem Graphen. (German) [Algorithm 27. A new algorithm for the solution of the k-th best route problem]. Computing, 16(1–2):139–151, March 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

James:1976:AAG

[931] K. R. James and W. Riha. Algorithm 28. algorithm for generating graphs of a given partition. Computing, 16 (1–2):153–161, March 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Riha:1976:AEA

[932] W. Riha and K. R. James. Algorithm 29. efficient algorithms for doubly and multiply restricted partitions. Computing, 16(1–2):163–168, March 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Menzel:1976:OEL

[933] R. Menzel and H. Schwetlick. Über einen Ordnungsbegriff bei Einbettungsalgorithmen zur Lösung nichtlinearer Gleichungen. (German) [On a notion of order in the case of continuation methods for solving nonlinear equations]. Computing, 16(3):187–199, September 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Solak:1976:SMB

[934] W. Solak and M. Strus. The secant method in Banach spaces. Computing, 16(3):201–209, September 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmidt:1976:MEI

[935] J. W. Schmidt. Monotone Einschliessung von Inversen positiver Elemente durch Verfahren vom Schulz-Typ. (German) [Monotonous enclosing of inverses of positive elements by Schulz-methods]. Computing, 16(3):

211–219, September 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Delvos:1976:PS

[936] F.-J. Delvos and H.-W. Kösters. On periodic M-splines. Computing, 16(3): 221–229, September 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Henderson:1976:STP

[937] D. S. Henderson and C. J. Wright. A Schmidt type process for the solution of Ax = b. Computing, 16(3): 231–237, September 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kulisch:1976:FIF

[938] U. Kulisch and G. Bohlender. Formalization and implementation of floatingpoint matrix operations. Computing, 16(3):239–261, September 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Papadimitriou:1976:NCB

[939] Ch. H. Papadimitriou. The NPcompleteness of the bandwidth minimization problem. Computing, 16(3): 263–270, September 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Durre:1976:FRC

[940] K. Dürre. Fast reversible coding and indexing of colorations. Computing, 16 (3):271–279, September 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ginsburg:1976:SEC

[941] S. Ginsburg and H. A. Maurer. On strongly equivalent context-free grammar forms. Computing, 16(3):281–290, September 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Einarsson:1976:RAR

[942] B. Einarsson. Remark on algorithm 24, regarding a very common violation of the Fortran Standard. Computing, 16 (3):291–292, September 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [874].

Anderson:1976:CAD

[943] P. G. Anderson, M. R. Garey, and L. E. Heindel. Computational aspects of deciding if all roots of a polynomial lie within the unit circle. Computing, 16 (4):293–304, December 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Majumder:1976:SAI

[944] K. L. Majumder and G. P. Bhattacharjee. Some algorithms for interval interpolating polynomial. Computing, 16 (4):305–317, December 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Baron:1976:TI

[945] W. Baron. Zur trigonometrischen Interpolation. (German) [On trigonometric interpolation]. Computing, 16(4): 319–328, December 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ichida:1976:CFP

[946] K. Ichida, T. Kiyono, and F. Yoshimoto. Curve fitting by a piecewise cubic polynomial. Computing, 16(4): 329–338, December 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

King:1976:BFE

[947] J. Th. King and St.M. Serbin. Boundary flux estimates for elliptic problems by the perturbed variational method. Computing, 16(4):339–347, December 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Nevanlinna:1976:PQV

[948] O. Nevanlinna. Positive quadratures for Volterra equations. Computing, 16 (4):349–357, December 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fahrmeir:1976:ASD

[949] L. Fahrmeir. Approximation von stochastischen Differentialgleichungen auf Digital- und Hybridrechnern. (German) [Approximation of stochastic differential equations on digital and hybrid computers]. Computing, 16(4): 359–371, December 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Findler:1976:AAE

[950] N. V. Findler and W. Rhomberg. On an approach to the automatic evaluation of the subjects’ verbal behaviour in Roemer’s inkblot tests. Computing, 16 (4):373–378, December 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Spath:1976:CA

[951] H. Späth. L_{1} cluster analysis. Computing, 16(4):379–387, December 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wettstein:1976:IHS

[952] H. Wettstein. Ein integriertes Hardware-Software-System zur Steuerung der Ein/Ausgabe. (German) [An integrated hardware-software-system for the control of input/output]. Computing, 17(1):1–21, March 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Spiess:1976:UZD

[953] J. Spiess. Untersuchungen des Zeitgewinns durch neue Algorithmen zur Matrix-Multiplikation. (German) [Investigations of speed-ups by new algorithms for matrix multiplication]. Computing, 17(1):23–36, March 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Deuflhard:1976:ASC

[954] P. Deuflhard. On algorithms for the summation of certain special functions. Computing, 17(1):37–48, March 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

King:1976:MSS

[955] R. F. King. Methods without secant steps for finding a bracketed root. Computing, 17(1):49–57, March 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmidt:1976:VZA

[956] J. W. Schmidt and W. Nauber. Über Verfahren zur zweiseitigen Approximation inverser Elemente. (German) [On methods for the two-sided approximation of inverse elements]. Computing, 17(1):59–67, March 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See correction [4367].

Hertling:1976:NSS

[957] J. Hertling. On the numerical solution of some algebraic integral equations. Computing, 17(1):69–77, March 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bohm:1976:DKS

[958] W. Böhm. Darstellung und Korrektur symmetrischer Kurven und Flächen auf EDV-Anlagen. (German) [Representation and correction of symmetric curves and surfaces]. Computing, 17(1):79–85, March 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bohm:1976:PKB

[959] W. Böhm. Parameterdarstellung kubischer und bikubischer Splines. (German) [Parametric representation of cubic and bicubic splines]. Computing, 17(1):87–92, March 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lohse:1976:ABW

[960] D. Lohse. Ein Algorithmus zur Bestimmung von Weganzahlen in Netzen. (German) [An algorithm for the computation of the number of paths in a network]. Computing, 17(1):93–96, March 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Seehusen:1976:VBT

[961] J. Seehusen. Virtuelle Z-Baum-Technik. (German) [Virtual Z-tree technique]. Computing, 17(2):97–104, June 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stork:1976:TLD

[962] H. G. Stork. A theorem on linear Diophantine equations and the pagingcomplexity of loop-chains. Computing, 17(2):105–113, June 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kameda:1976:MDS

[963] T. Kameda. On maximally distant spanning trees of a graph. Computing, 17(2):115–119, June 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Horst:1976:GTA

[964] R. Horst. Ein graphentheoretischer Algorithmus für einige Flußprobleme in Netzwerken mit Kantengewinnen. (German) [A graph-theoretical algorithm for some flow problems through networks with gains]. Computing, 17(2):121–127, June 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Junginger:1976:NEL

[965] W. Junginger. Numerische Erfahrungen bei der Lösung des dreidimensionalen Transportproblems. (German)

[Numerical experiences in solving three dimensional transportation problems]. Computing, 17(2):129–145, June 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fishman:1976:SPD

[966] G. S. Fishman. Sampling from the Poisson distribution on a computer. Computing, 17(2):147–156, June 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wal:1976:SAA

[967] J. van der Wal. A successive approximation algorithm for an undiscounted Markov decision process. Computing, 17(2):157–162, June 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pottinger:1976:PGP

[968] P. Pottinger. Polynomoperatoren in C^{k}[a, b]. (German) [Polynomial operators in C^{k}[a, b]]. Computing, 17(2):163– 167, June 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

DeMey:1976:IEM

[969] G. De Mey. An integral equation method for the numerical calculation of ion drift and diffusion in evaporated dielectrics. Computing, 17(2):169–176, June 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Taubert:1976:ETG

[970] K. Taubert. Eine Erweiterung der Theorie von G. Dahlquist. (German) [An extension of the theory given by G. Dahlquist]. Computing, 17(2):177–185, June 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hofmann:1976:SMN

[971] W. Hofmann and H. Voss. Shooting-Verfahren bei nichtlinearen, autonomen Schwingungen. (German) [Shooting methods for nonlinear autonomous oscillations]. Computing, 17(2):187–191, June 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eckhardt:1976:NBI

[972] U. Eckhardt. Zur numerischen Behandlung inkorrekt gestellter Aufgaben. (German) [On the numerical treatment of improperly posed problems]. Computing, 17(3):193–206, September 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Spellucci:1976:ATL

[973] P. Spellucci. Ein Verfahren zur Behandlung von Ausgleichsaufgaben mit Intervallkoeffizienten. (German) [An algorithm for the treatment of the linear least squares problem with interval coefficients]. Computing, 17(3): 207–218, September 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sameh:1976:PPB

[974] A. H. Sameh, S. C. Chen, and D. J. Kuck. Parallel Poisson and biharmonic solvers. Computing, 17(3): 219–230, September 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gentzch:1976:DQE

[975] W. Gentzch. Zur Diskretisierung quasilinearer elliptischer Differentialgleichungen vierter Ordnung. (German) [The numerical treatment of quasilinear elliptic differential equations of the fourth order by the method of finite differences]. Computing, 17 (3):231–246, September 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Strehmel:1976:MED

[976] K. Strehmel. Mehrschrittverfahren mit Exponentialanpassung für Differentialgleichungssysteme 1. Ordnung. (German) [Multistep methods for systems of first order differential equations with exponential adaptation]. Computing, 17(3):247–260, September 1976. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

XXX:1976:XXX

[977] ????  ????  Computing, 17(3):

256–279, September 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [1369].

vanderWal:1976:NAM

[978] J. van der Wal. Note on algorithm 21 (Markov process). Computing, 17 (3):261–264, September 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

DeDoncker:1976:ACI

[979] E. De Doncker and R. Piessens. Automatic computation of integrals with singular integrand, over a finite or an infinite range. Computing, 17(3): 265–279, September 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Seybold:1977:BKC

[980] H. Seybold. Beitrag zur konstruktiven Computer-Geometrie. (German) [Suggestions for the application of a plotter in descriptive geometry]. Computing, 17(4):281–287, December 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Reimer:1977:AFO

[981] M. Reimer. Auswertungsalgorithmen fast-optimaler numerischer Stabilität für Polynome (German) Algorithms of Near-optimal Numerical Stability for the Evaluation of Polynomials. Computing, 17(4):289–296, December 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Elsaesser:1977:KLI

[982] U. Elsaesser and W. Niethammer. Kriterien für die Liapunovstabilität und ihre numerische Anwendung. (German) [Criteria for the Liapunov stability and their numerical application]. Computing, 17(4):297–308, December 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Knauff:1977:FQA

[983] W. Knauff. Fehlernormen zur Quadratur analytischer Funktionen. (German) [Error norms of quadratures for analytic functions]. Computing, 17 (4):309–322, December 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wolf:1977:KDL

[984] G. Wolf. Konstruktion von Differenzenverfahren zur Lösung von hyperbolischen Systemen partieller Differentialgleichungen erster Ordnung mit Hilfe von Spektraldarstellungen. (German) [Construction of difference methods for the solution of hyperbolic systems of first order partial differential equations by spectral decompositions]. Computing, 17(4):323–334, December 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Butzer:1977:AFO

[985] P. L. Butzer, W. Dickmeis, Hu. Jansen, and R. J. Nessel. Alternative forms with orders of the Lax equivalence theorem in Banach spaces. Computing, 17 (4):335–342, December 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gentzsch:1977:PMN

[986] W. Gentzsch. Penalty-Methode bei der numerischen Behandlung von quasilinearen elliptischen Differentialgleichungen vierter und höherer Ordnung. (German) [Penalty-method in the numerical treatment of quasilinear elliptic differential equations of fourth and higher order]. Computing, 17(4):343–350, December 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Krabs:1977:ERS

[987] W. Krabs. Über die einseitige Randsteuerung einer schwingenden Saite in einen Zustand minimaler Energie. (German) [On steering at one end a vibrating string into a state of minimal energy]. Computing, 17(4):351–359, December 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Filippi:1977:VSO

[988] S. Filippi. Verallgemeinerte k-Schrittverfahren der Ordnung p = 3k − m + 2 und der Ordnung p = 2k − m + 1 zur numerischen Lösung von Anfangswertaufgaben bei Differentialgleichungen m-ter Ordnung der Form y^{(m)} = f(x, y). (German) [Generalized multistep methods of order p = 3k−m+2 and of order p = 2k−m+1 for the numerical solution of initial value problems for the special m-th order differential equation y^{(m)} = f(x, y)]. Computing, 17(4):361–372, December 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Deuflhard:1977:STM

[989] P. Deuflhard. A summation technique for minimal solutions of linear homogeneous difference equations. Computing, 18(1):1–13, March 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Nickel:1977:UWF

[990] K. Nickel. Die Überschätzung des Wertebereichs einer Funktion in der Intervallrechnung mit Anwendungen auf lineare Gleichungssysteme. (German) [The overestimation of function ranges in interval computation with use of linear equation systems]. Computing, 18(1):15–36, March 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rietsch:1977:FSK

[991] E. Rietsch. A Fourier series kernel based on Chebyshev polynomials. Computing, 18(1):37–50, March 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wagner:1977:AAA

[992] H.-J. Wagner. An algorithm for asymptotic approximation of Laplace integrals. Computing, 18(1):51–58, March 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Knauff:1977:GQA

[993] W. Knauff. Gewichtsoptimale Quadraturformeln bei analytischen Funktionen. (German) [Weight-optimal quadrature formulas for analytic functions]. Computing, 18(1):59–66, March 1977. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mittelmann:1977:NBN

[994] H. D. Mittelmann. Numerische Behandlung nichtlinearer Randwertprobleme mit finiten Elementen. (German) [Numerical treatment of nonlinear boundary value problems with finite elements]. Computing, 18(1):67– 77, March 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gekeler:1977:GRK

[995] E. Gekeler. Galerkin–Runge–Kutta methods and hyperbolic initial boundary value problems. Computing, 18 (1):79–88, March 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mossige:1977:AGC

[996] S. Mossige. An algorithm for Gray codes. Computing, 18(1):89–92, March 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Goodman:1977:EGD

[997] R. Goodman and A. Feldstein. Effect of guard digits and normalization options on floating point multiplication. Computing, 18(2):93–106, June 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schwarz:1977:EDS

[998] H. R. Schwarz. Elementare Darstellung der schnellen Fouriertransformation. (German) [Elementary formulation of fast Fourier transforms]. Computing, 18(2):107–116, June 1977. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Trigiante:1977:ASD

[999] D. Trigiante. Asymptotic stability and discretization on an infinite interval. Computing, 18(2):117–129, June 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ostermann:1977:HDA

[1000] A. Ostermann. Hermitesche Differenzenverfahren zur Approximation der Lösung von Differentialgleichungen der Form u_{xy} = f(x, y, u, u_{x}, u_{y}). (German) [Hermitian difference method for approximation of the solution of differential equations of the form u_{xy} = f(x, y, u, u_{x}, u_{y})]. Computing, 18(2):131–139, June 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mittelmann:1977:ACS

[1001] H. D. Mittelmann. On the approximation of capillary surfaces in a gravitational field. Computing, 18(2):141–148, June 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Spellucci:1977:NSC

[1002] P. Spellucci. Numerically stable computation of step-sizes for descent methods. The nonconvex case. Computing, 18(2):149–160, June 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bohm:1977:KBS

[1003] W. Bohm. Über die Konstruktion von B-Spline-Kurven. (German) [On calculating of B-spine curves]. Computing, 18(2):161–166, June 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dutter:1977:AHE

[1004] R. Dutter. Algorithms for the Huber estimator in multiple regression. Computing, 18(2):167–176, June 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hajek:1977:NRE

[1005] O. Hajek. On the number of roots of exp-trig polynomials. Computing, 18(2):177–183, June 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Nehmer:1977:ESP

[1006] J. Nehmer. Experimental study of processor thrashing in multiprocessor systems. Computing, 18(3):185–197, September 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Golumbic:1977:CCG

[1007] M. C. Golumbic. The complexity of comparability graph recognition and coloring. Computing, 18(3):199–208, September 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wacker:1977:MRG

[1008] Hj. Wacker. Minimierung des Rechenaufwandes bei Globalisierungen spezieller Iterationsverfahren vom Typ minimales Residuum. (German) [Minimization of the computational cost by globalization of special interation methods with minimal residual type]. Computing, 18(3):209–224, September 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rokne:1977:BIP

[1009] J. Rokne. Bounds for an interval polynomial. Computing, 18(3):225–240, September 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chawla:1977:EEN

[1010] M. M. Chawla and R. Kress. Estimating errors of numerical approximation for periodic analytic functions. Computing, 18(3):241–248, September 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Puke:1977:KIM

[1011] V. Puke. Konvergenz impliziter Mehrschrittverfahren zur numerischen Lösung quasilinearer Anfangswertaufgaben. (German) [Convergence of implicit multi-level methods for the numerical solution of quasilinear initialvalue problems]. Computing, 18(3): 249–262, September 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Watson:1977:TMD

[1012] G. A. Watson. On two methods for discrete L_{p} approximation. Computing, 18 (3):263–266, September 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hohn:1977:MMP

[1013] W. Hohn and W. Törnig. Ein Maximum-Minimum-Prinzip für Lösungen von finite-Element- Approximationen für quasilineare elliptische Randwertprobleme. (German) [A maximumminimum-principle for finite element solutions of quasilinear elliptic boundary value problems]. Computing, 18 (3):267–270, September 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Engels:1977:WQ

[1014] H. Engels and U. Eckhardt. Wilfquadrature. Computing, 18(3):271– 279, September 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Neuhold:1977:UHM

[1015] E. J. Neuhold and R. Studer. The use of Hoare’s method of program verification for the Quicksort algorithm. Computing, 18(4):281–293, December 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See coments [1136].

Haubitz:1977:PZA

[1016] I. Haubitz. Programme zum Zeichnen von allgemeinen Flächenstücken. (German) [Programs drawing pieces of surfaces]. Computing, 18(4):295–315, December 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lam:1977:RTF

[1017] D. K. Lam and A. Wouk. Redundancy techniques and fast algorithms for a special large linear system. Computing, 18(4):317–327, December 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Alefeld:1977:SEL

[1018] G. Alefeld. Das Symmetrische Einzelschrittverfahren Bei Linearen Gleichungen Mit Intervallen Als Koeffizienten. (German) [The symmetric single step method for systems of simultaneous linear equations with intervals as coefficients]. Computing, 18 (4):329–340, December 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eckhardt:1977:RAW

[1019] U. Eckhardt. A rational approximation to Weierstrass’ elliptic function. II. The lemniscatic case. Computing, 18 (4):341–349, December 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hinteregger:1977:ZKG

[1020] J. Hinteregger and G. Tinhofer.

Zerlegung der Knotenmengen von Graphen zum Nachweis der Isomorphie. (German) [Node partitioning for graph isomorphism]. Computing, 18 (4):351–359, December 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Berg:1977:ARA

[1021] L. Berg. Zur Abschätzung des Restgliedes in der asymtotischen Entwicklung des Exponentialintegrals. (German) [On the estimation of the remainder term in the asymptotic expansion of the exponential integral]. Computing, 18(4):361–363, December 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mandl:1977:NOU

[1022] C. Mandl. Number of operations for updating the elimination form of the basis-inverse of the revised simplex algorithm. Computing, 18(4): 365–366, December 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Waldschmidt:1977:SVS

[1023] H. Waldschmidt. Sortieren bei virtuellem Speicher. (German) [Sorting in a paging environment]. Computing, 19(1):1–14, March 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

vorderBruck:1977:AMU

[1024] H. vor der Brück. Abstrakte Modelle zur Untersuchung von sicheren Systemen. (German) [Abstract model for investigation of secure systems]. Computing, 19(1):15–28, March 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bohm:1977:CBS

[1025] W. Böhm. Cubic B-spline curves, and surfaces in computer aided geometric design. Computing, 19(1):29–34, March 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wagner:1977:CNA

[1026] H.-J. Wagner. A contribution to the numerical approximation of minimal surfaces. Computing, 19(1):35–58, March 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Warth:1977:ESU

[1027] W. Warth and J. Werner. Effiziente Schrittweitenfunktionen bei unrestringierten Optimierungsaufgaben. (German) [Efficient stepsize functions for unconstrained optimization problems]. Computing, 19(1):59–72, March 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bhattacharjee:1977:CHI

[1028] G. P. Bhattacharjee and K. L. Majumder. On the computation of Hermite interval interpolating polynomials. Computing, 19(1):73–83, March 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Richter:1977:PLS

[1029] G. R. Richter and W. L. Steiger. Polynomial least squares approximations with ill-conditioned bases. Computing, 19(1):85–88, March 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Miklosko:1977:MLL

[1030] F. Miklosko. Über eine Methode zur Lösung eines linearen Gleichungssystems. (German) [On a method for solution of a linear equation system]. Computing, 19(1):89–93, March 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Muller:1977:OSP

[1031] H. Müller. A one-symbol printing automaton escaping from every labyrinth. Computing, 19(2):95–110, June 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lieflander:1977:RPV

[1032] G. Liefländer and G. Nees. Reduzierung von Prozeßstauungen vor kritischen Abschnitten mittels Teillistensperrung. (German) [Reduction of process congestion for critical sections via partial list blocking]. Computing, 19(2):111–127, June 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Culik:1977:TCG

[1033] Karel Culik II and H. A. Maurer. Tree controlled grammars. Computing, 19(2):129–139, June 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ginsburg:1977:QIG

[1034] S. Ginsburg and H. A. Maurer. On quasi-interpretations of grammar forms. Computing, 19(2):141–147, June 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zlamal:1977:MIC

[1035] J. Zlamal. A measure for illconditioning of matrices in interval arithmetic. Computing, 19(2):149–155, June 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schek:1977:LSL

[1036] H.-J. Schek and R. Eggensperger.

Least-Squares-Lösungen und Dämpfung bei unterbestimmten nichtlinearen Gleichungssystemen. (German) [Leastsquares solution and damping by underdetermined nonlinear equation systems]. Computing, 19(2):157–170, June 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Alefeld:1977:IEP

[1037] G. Alefeld. Zur Iterativen Einschliessung Von Positiven Inversen. (German) [On the approximation of positive inverse elements]. Computing, 19(2):171–174, June 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmidt:1977:KKE

[1038] J. W. Schmidt. Ein Kubisch Konvergentes Einschliessungsverfahren Zur Inversion Positiver Elemente. (German) [A monotone order three method for the inversion of positive elements]. Computing, 19(2):175–178, June 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Haegemans:1977:AAI

[1039] A. Haegemans. An algorithm for the automatic integration over a triangle. Computing, 19(2):179–187, June 1977.

CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Smith:1978:ISP

[1040] B. J. Smith and H. F. Jordan. Implications of series-parallel sequencing rules. Computing, 19(3):189–201, September 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Meister:1978:TQS

[1041] B. Meister. A tandem queueing system with priorities. Computing, 19 (3):203–208, September 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Henn:1978:APS

[1042] R. Henn. Antwortzeitgesteuerte Prozessorzuteilung unter Strengen Zeitbedingungen. (German) [Response time driven processor scheduling with strict deadlines]. Computing, 19(3): 209–220, September 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Glaser:1978:SNB

[1043] J. Glaser. Das Stabilitätsproblem bei der numerischen Behandlung von Differentialgleichungen. (German) [The stability problem in the numerical treatment of differential equations]. Computing, 19(3):221–231, September 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Albrecht:1978:EOS

[1044] P. Albrecht. Explicit, optimal stability functionals and their application to cyclic discretization methods. Computing, 19(3):233–249, September 1978.

CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gerdes:1978:LAR

[1045] W. Gerdes. Die Lösung des Anfangs-Randwertproblems für die Wärmeleitungsgleichung im R^{3} mit einer Integralgleichungsmethode nach dem Rotheverfahren. (German) [Initial boundary value problem for the threedimensional heat equation treated with Rothe’s line method using integral equations]. Computing, 19(3):251–268, September 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Esser:1978:NBV

[1046] R. Esser. Numerische Behandlung einer Volterraschen Integralgleichung. (German) [Numerical treatment of a Volterra integral equation]. Computing, 19(4):269–284, December 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Derigs:1978:APM

[1047] U. Derigs and U. Zimmermann. Augmenting path method for solving linear bottleneck assignment problems. Computing, 19(4):285–295, December 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stadel:1978:RTC

[1048] M. Stadel. A remark on the time complexity of the subtree problem. Computing, 19(4):297–302, December 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Thieler:1978:CEE

[1049] P. Thieler. On componentwise error estimates for inverse matrices. Computing, 19(4):303–312, December 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kioustelidis:1978:AEE

[1050] J. B. Kioustelidis. Algorithmic error estimation for approximate solutions of nonlinear systems of equations. Computing, 19(4):313–320, December 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Concus:1978:NSN

[1051] P. Concus, G. H. Golub, and D. P. O’Leary. Numerical solution of nonlinear elliptic partial differential equations by a generalized conjugate gradient method. Computing, 19(4): 321–339, December 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schwarz:1978:FFT

[1052] H. R. Schwarz. The fast Fourier transform for general order. Computing, 19 (4):341–350, December 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmidt:1978:KQI

[1053] J. W. Schmidt and H. Mettke. Konvergenz von quadratischen interpolationsund flächenabgleichssplines. (German) [Convergence of quadratic interpolatory and histo splines]. Computing, 19 (4):351–363, December 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pruess:1978:ACS

[1054] S. Pruess. An algorithm for computing smoothing splines in tension. Computing, 19(4):365–373, December 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wagner:1978:CON

[1055] H.-J. Wagner. Consideration of obstacles in the numerical approximation of minimal surfaces. Computing, 19 (4):375–379, December 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schultheis:1978:ANS

[1056] H. Schultheis and R. Schultheis. An algorithm for non-smoothing contour representations of two-dimensional arrays. Computing, 19(4):381–387, December 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Merluzzi:1978:RKI

[1057] P. Merluzzi and C. Brosilow. Runge– Kutta integration algorithms with built-in estimates of the accumulated truncation error. Computing, 20(1):1– 16, March 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Scheu:1978:WPB

[1058] G. Scheu. Über Eine Wahl Des Parameters Beim Parallelenverfahren (Parallel-Chord-Method). (German) [On the choice of the parameter in the parallel-chord-method]. Computing, 20(1):17–26, March 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sprekels:1978:VIE

[1059] J. Sprekels and H. Voss. Ein Verfahren zur iterativen Einschliessung des positiven Eigenvektors einer irreduziblen, nichtnegativen Matrix. (German) [A method for iterative bracketing of positive eigenvectors of an irreducible nonnegative matrix]. Computing, 20(1):27–34, March 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kioustelidis:1978:ASC

[1060] J. B. Kioustelidis and K. J. Spyropoulos. L_{1} approximations of strictly convex functions by means of first degree splines. Computing, 20(1):35–45, March 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hackbusch:1978:MCS

[1061] W. Hackbusch. On a method of characteristics for solving a hyperbolic equation of second order. Computing, 20(1):47–60, March 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Koppe:1978:AAP

[1062] R. Koppe. Automatische Abbildung eines planaren Graphen in die Ebene mit beliebig vorgebbaren Örtern der Knotenbilder. (German) [Automatic mapping of a planar graph in the plane with any arbitrarily-prescribed knot positions]. Computing, 20(1):61–73, March 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stephan:1978:ASH

[1063] E. Stephan and V. Weissgerber. Zur Approximation von Schalen mit hybriden Elementen. (German) [On the approximation of shells with hybrid elements]. Computing, 20(1):75–94, March 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Beth:1978:SFS

[1064] Th. Beth and K. S. Vijayan. Some filing schemes based on finite groups of automorphisms. Computing, 20(2):95–108, June 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Levy:1978:LGG

[1065] L. S. Levy and Kang Yueh. On labelled graph grammars. Computing, 20(2):109–125, June 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Berg:1978:SSU

[1066] L. Berg. Stable solution of instable systems of linear equations with a band matrix. Computing, 20(2):127–137, June 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Proskurowski:1978:NSE

[1067] W. Proskurowski. On the numerical solution of the eigenvalue problem of the Laplace operator by a capacitance matrix method. Computing, 20(2):139–151, June 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hansen:1978:IFN

[1068] E. Hansen. Interval forms of Newton’s method. Computing, 20(2):153–163, June 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rokne:1978:PLS

[1069] J. Rokne. Polynomial least square interval approximation. Computing, 20(2):165–176, June 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schaback:1978:RFM

[1070] R. Schaback. Eine rundungsgenaue Formel zur maschinellen Berechnung der Prager-Oettli-Schranke. (German) [An exact rounding formula of Prager-Oettli-Schranke for machine computation]. Computing, 20(2):177–182, June 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lux:1978:SMS

[1071] I. Lux. A special method to sample some probability density functions. Computing, 20(2):183–188, June 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Scholz:1978:SMP

[1072] K. Scholz. On scheduling multiprocessor systems with algebraic objectives. Computing, 20(3):189–205, September 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Frank:1978:IDC

[1073] R. Frank and C. W. Überhuber. Iterated defect correction for differential equations. I. Theoretical results. Computing, 20(3):207–228, September 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hackbusch:1978:NMS

[1074] W. Hackbusch. A numerical method for solving parabolic equations with opposite orientations. Computing, 20(3): 229–240, September 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Appelt:1978:AIG

[1075] W. Appelt. Ein Abbruchkriterium für das intervallarithmetische Gesamtund Einzelschrittverfahren. (German) [A stopping criterion for interval arithmetic multi- and single-step methods]. Computing, 20(3):241–255, September 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wildenauer:1978:ANA

[1076] P. Wildenauer. Ein Abbrechkriterium für einen numerischen Algorithmus. (German) [A stopping criterion for a numerical algorithm]. Computing, 20 (3):257–265, September 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mihelcic:1978:SCC

[1077] M. Mihelčić. A(α)-stable cyclic composite multistep methods of order 5. Computing, 20(3):267–272, September 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1978:MOP

[1078] H. J. Stetter. On the maximal order in PC-codes. Computing, 20(3): 273–278, September 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mittermeir:1978:KMO

[1079] R. Mittermeir. Komplexitätserfordernisse von Modellen zur optimalen Indexselektion für Datenbanken. (German) [Complexity requirements of a model of optimal index selection for databases]. Computing, 20(4):279–290, December 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hackbusch:1978:MMA

[1080] W. Hackbusch. On the multigrid method applied to difference equations. Computing, 20(4):291–306, December 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Branca:1978:NVE

[1081] H. W. Branca. The nonlinear Volterra equation of Abel’s kind and its numerical treatment. Computing, 20(4): 307–324, December 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wildenauer:1978:MRB

[1082] P. Wildenauer. Minimierung der Rechenzeit bei der Berechnung von Lösungen nichtlinearer Integralgleichungen. (German) [Minimizing the computing time for solutions of nonlinear integral equations]. Computing, 20 (4):325–331, December 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wambecq:1978:RRK

[1083] A. Wambecq. Rational Runge–Kutta methods for solving systems of ordinary differential equations. Computing, 20 (4):333–342, December 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Grigorieff:1978:SVH

[1084] R. D. Grigorieff and J. Schroll. Über A(α)-stabile Verfahren höher Konsistenzordnung. (German) [On A(α)- stable methods with high order of consistency]. Computing, 20(4):343–350, December 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Jackiewicz:1978:CMM

[1085] Z. Jackiewicz and M. Kwapisz. On the convergence of multistep methods for the Cauchy problem for ordinary differential equations. Computing, 20 (4):351–361, December 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Garribba:1978:ASE

[1086] S. Garribba, L. Quartapelle, and G. Reina. Algorithm 36. SNIFF: Efficient self-tuning algorithm for numerical integration. Computing, 20(4): 363–375, December 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bayer:1978:PDM

[1087] R. Bayer and T. Haerder. Preplanning of disk merges. Computing, 21(1):1–16, March 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bayer:1978:PMP

[1088] R. Bayer and T. Haerder. A performance model for preplanned disk sorting. Computing, 21(1):17–35, March 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Cromme:1978:PBL

[1089] L. Cromme. Zur praktischen Behandlung linearer diskreter Approximationsprobleme in der Maximumsnorm. (German) [On the practical treatment of linear discrete approximation problems in the maximum norm]. Computing, 21(1):37–52, March 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gustafson:1978:CAG

[1090] S.-Å. Gustafson. Convergence acceleration on a general class of power series. Computing, 21(1):53–69, March 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Paulik:1978:DWM

[1091] A. Paulik. On the determination of the weights in multidimensional numerical quadrature. Computing, 21(1):71–79, March 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Martello:1978:AAS

[1092] S. Martello and P. Toth. Algorithm 37. Algorithm for the solution of the 0-1 single knapsack problem. Computing, 21(1):81–86, March 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gustafson:1978:ATC

[1093] S.-Å. Gustafson. Algorithm 38. Two computer codes for convergence acceleration. Computing, 21(1):87–91, March 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rivlin:1979:AC

[1094] T. J. Rivlin. Approximation by circles.

Computing, 21(2):93–104, June 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Marti:1979:CDA

[1095] J. T. Marti. On the convergence of the discrete ART algorithm for the reconstruction of digital pictures from their projections. Computing, 21(2):105–111, June 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tewarson:1979:UDQ

[1096] R. P. Tewarson. A unified derivation of quasi-Newton methods for solving nonsparse and sparse nonlinear equations. Computing, 21(2):113–125, June 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brauninger:1979:QNM

[1097] J. Bräuninger. A quasi-Newton method for minimization under linear constraints without evaluating any derivatives. Computing, 21(2):127–141, June 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Heidtmann:1979:VEG

[1098] K. Heidtmann. Verbesserte Einzelschrittverfahren. (German) [Improved single-step methods]. Computing, 21(2):143–150, June 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brunner:1979:SCM

[1099] H. Brunner. Superconvergence of collocation methods for Volterra integral equations of the first kind. Computing, 21(2):151–157, June 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See note [1157].

Rokne:1979:NBA

[1100] J. Rokne. A note on the Bernstein algorithm for bounds for interval polynomials. Computing, 21(2):159–170, June 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lether:1979:PCN

[1101] F. Lether, D. Wilhelmsen, and R. Frazier. On the positivity of Cotes numbers for interpolatory quadratures with Jacobi abscissas. Computing, 21(2):171–175, June 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bitzer:1979:MAT

[1102] G. Bitzer. Über eine Minimaleigenschaft der äquidistanten trigonometrischen Interpolation in 2n + 1 Knoten. (German) [On a minimal property of the equidistant trigonometric interpolation in 2n+1 knots]. Computing, 21(2):177– 182, June 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Martello:1979:AFM

[1103] S. Martello. An algorithm for finding a minimal equivalent graph of a strongly connected digraph. Computing, 21(3): 183–194, September 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Vaishnavi:1979:HBT

[1104] V. K. Vaishnavi, H. P. Kriegel, and D. Wood. Height balanced 2-3 trees. Computing, 21(3):195–211, September 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Paun:1979:GCT

[1105] Gh. Păun. On the generative capacity of tree controlled grammars. Computing, 21(3):213–220, September 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Alt:1979:SRD

[1106] H. Alt. Square rooting is as difficult as multiplication. Computing, 21(3): 221–232, September 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stiller:1979:KNL

[1107] Silke Stiller. Konvergenzverhalten bei der numerischen Lösung von Randwertproblemen mittels eines Doppelschichtpotential-Ansatzes. (German) [Convergence behavior of numerical solutions of the boundary value problem with a double layer potential estimate]. Computing, 21(3):233–243, September 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Huntley:1979:AND

[1108] E. Huntley and A. S. I. Zinober. Applications of numerical double Laplace transform algorithms to the solution of linear partial differential equations. Computing, 21(3):245–258, September 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Evans:1979:FSS

[1109] D. J. Evans and A. Hadjidimos. On the factorization of special symmetric periodic and non-periodic quindiagonal matrices. Computing, 21(3): 259–266, September 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Alefeld:1979:KIL

[1110] G. Alefeld. Zur Konvergenz des Iterationsverfahrens bei linearen Gleichungssystemen. (German) [On the convergence of the iteration method for linear systems of equations]. Computing, 21(3):267–272, September 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Clarke:1979:PIF

[1111] E. M. Clarke. Program invariants as fixed points. Computing, 21(4): 273–294, December 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gerhards:1979:CDN

[1112] L. Gerhards and W. Lindenberg. Clique detection for nondirected graphs: Two new algorithms. Computing, 21(4):295–322, December 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sonntag:1979:VAV

[1113] H. Sonntag. Ein Verfahren zur Alternativroutensuche in Verkehrsnetzen. (German) [Algorithm for searching alternative routes in traffic networks]. Computing, 21(4):323–331, December 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Jurgensen:1979:TEA

[1114] H. Jürgensen. Transformationendarstellungen endlicher abstrakt präsentierter Halbgruppen. (German) [Transformation representations of finite abstractly presented semigroups]. Computing, 21 (4):333–342, December 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

McKee:1979:BCR

[1115] S. McKee. Best convergence rates of linear multistep methods for Volterra first kind equations. Computing, 21 (4):343–358, December 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lux:1979:ASM

[1116] I. Lux. Another special method to sample probability density functions. Computing, 21(4):359–364, December 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Derigs:1979:APM

[1117] U. Derigs and U. Zimmermann. An augmenting path method for solving linear bottleneck transportation problems. Computing, 22(1):1–15, March 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hamacher:1979:NIM

[1118] H. Hamacher. Numerical investigations on the maximal flow algorithm of Karzanov. Computing, 22(1):17–29, March 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Walter:1979:VKG

[1119] H. K. G. Walter. Die Verwandtschaft kontextfreier Grammatiken. (German) [On the relatedness of context-free grammars]. Computing, 22(1):31–58, March 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

OLeary:1979:CGA

[1120] Dianne P. O’Leary. Conjugate gradient algorithms in the solution of optimization problems for nonlinear elliptic partial differential equations. Computing, 22(1):59–77, March 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gutknecht:1979:FAC

[1121] M. H. Gutknecht. Fast algorithms for the conjugate periodic function. Computing, 22(1):79–91, March 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bjorstad:1979:ANA

[1122] Petter Bjørstad and Jorge Nocedal. Analysis of a new algorithm for onedimensional minimization. Computing, 22(1):93–100, March 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Heller:1979:MPA

[1123] D. Heller. Minimal parallelism for associative computations under time constraints. Computing, 22(2):101–118, June 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Niederreiter:1979:OFQ

[1124] H. Niederreiter and Kevin McCurley.

Optimization of functions by quasirandom search methods. Computing, 22(2):119–123, June 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Strehmel:1979:NLA

[1125] K. Strehmel and C. Peper. Numerische Lösung von Anfangswertaufgaben gewöhnlicher Differentialgleichungen n-ter Ordnung mittels Einschrittverfahren. (German) [Numerical solution of initial value problems of nth-order ordinary differential equations with one-step methods]. Computing, 22(2):125–139, June 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kaspar:1979:NBP

[1126] W. Kaspar and R. Remke. Die numerische Behandlung der Poisson-Gleichung auf einem Gebiet mit einspringenden Ecken. (German) [On the numerical solution of the Poisson equation on a domain with re-entrant corners]. Computing, 22(2):141–151, June 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rokne:1979:RVC

[1127] J. Rokne. The range of values of a complex polynomial over a complex interval. Computing, 22(2):153–169, June 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brombeer:1979:VAR

[1128] R. Brombeer. Verallgemeinerte Ausgleichsformeln und relativ optimale Formeln in Hilberträumen. (German) [Generalized least-square approximate formulas and relative-optimal formulas in Hilbert space]. Computing, 22(2):171–182, June 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Apostolopoulos:1979:IAN

[1129] N. Apostolopoulos and G. Schuff. Initializing algorithms: a note to the article: “Computer methods for sampling from gamma, beta, Poisson and binomial distributions” [Computing 12 (1974), no. 3, 223–246; MR 52 #15949] by J. H. Ahrens and U. Dieter. Computing, 22(2):185–189, June 1979. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [788].

Trivedi:1979:AP

[1130] K. S. Trivedi. An analysis of prepaging. Computing, 22(3):191–210, September 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Waldschmidt:1979:VLI

[1131] H. Waldschmidt. Zur Verbesserung der Lokalität im Zugriff auf Felder. (German) [Improvement of the locality in referencing data]. Computing, 22(3): 211–224, September 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brown:1979:RRD

[1132] Th. Brown. M/G/1 round robin discipline. Computing, 22(3):225–241, September 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Barth:1979:FRC

[1133] G. Barth. Fast recognition of contextsensitive structures. Computing, 22(3): 243–256, September 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dubois:1979:AIM

[1134] P. F. Dubois, A. Greenbaum, and G. H.

Rodrigue. Approximating the inverse of a matrix for use in iterative algorithms on vector processors. Computing, 22(3):257–268, September 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

VanderCruyssen:1979:LDE

[1135] P. Van der Cruyssen. Linear difference equations and generalized continued fractions. Computing, 22(3): 269–278, September 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pritchard:1979:CPJ

[1136] P. Pritchard. Comments on a paper by E. J. Neuhold and R. Studer: “The use of Hoare’s method of program verification for the quick-sort algorithm” [Computing 18 (1977), no. 4, 281–293; MR 56 #13753]. Computing, 22(3): 279–282, September 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [1015].

Ottmann:1979:OSH

[1137] Th. Ottmann, H. W. Six, and D. Wood.

One-sided k-height-balanced trees. Computing, 22(4):283–290, December 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

vanderHouwen:1979:OSS

[1138] P. J. van der Houwen and J. G. Verwer. One-step splitting methods for semidiscrete parabolic equations. Computing, 22(4):291–309, December 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Delvos:1979:BZL

[1139] F. J. Delvos and H. Posdorf. Boolesche zweidimensionale Lagrange-Interpolation. (German) [Boolean bivariate Lagrange interpolation]. Computing, 22(4):311–323, December 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Markov:1979:CIF

[1140] S. Markov. Calculus for interval functions of a real variable. Computing, 22 (4):325–337, December 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bauch:1979:IIS

[1141] H. Bauch. On the iterative inclusion of solutions in initial-value problems for ordinary differential equations. Computing, 22(4):339–354, December 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Reichmann:1979:ABI

[1142] K. Reichmann. Abbruch beim Intervall-Gauß-Algorithmus. (German) [Breaking down of the interval Gauss algorithm]. Computing, 22(4):355–361, December 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Vitoria:1979:GIC

[1143] J. Vitória. Generalization of inequalities by Carmichael and Mason and by Parodi. Computing, 22(4):363–365, December 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Spath:1979:ACL

[1144] H. Späth. Algorithm 39. Clusterwise linear regression. Computing, 22(4): 367–373, December 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See correction [1257].

Bingulac:1979:LAS

[1145] S. P. Bingulac and M. A. Farias. L-A-S (Linear Algebra and System) language and its use in control system education and research. Computing, 23(1):1–23, March 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Engel:1979:DSC

[1146] A. B. Engel. Dynamical systems characterized by automata. Computing, 23(1):25–29, March 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ziegler:1979:PSZ

[1147] B. Ziegler. Prozessorzuteilung unter strengen Zeitbedingungen. (German) [Processor allocation under strict timeconditions]. Computing, 23(1):31–41, March 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ottmann:1979:CBA

[1148] Th. Ottmann, H.-W. Six, and D. Wood. On the correspondence between AVL trees and brother trees. Computing, 23(1):43–54, March 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bellen:1979:CMN

[1149] A. Bellen. The collocation method for the numerical approximation of the periodic solutions of functional differential equations. Computing, 23(1):55– 66, March 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chawla:1979:CQC

[1150] M. M. Chawla and Sheo Kumar. Convergence of quadratures for Cauchy principal value integrals. Computing, 23(1):67–72, March 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sprekels:1979:BDE

[1151] J. Sprekels and H. Voss. Bounds and domains of existence for hyperbolic initial value problems. Computing, 23(1):73–83, March 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ichida:1979:IAM

[1152] K. Ichida and Y. Fujii. An interval arithmetic method for global optimization. Computing, 23(1):85–97, March 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bolch:1979:EMP

[1153] G. Bolch and B. Schmidt. Ein erweitertes Modell für Paging-Trommeln. (German) [An expanded model for paging drums]. Computing, 23(2):99–118, June 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lippe:1979:EPG

[1154] W. M. Lippe and F. Simon. Zur Elimination von Prozedurschachtelungen. (German) [On the elimination of procedure nestings]. Computing, 23(2):119–138, June 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Grassmann:1979:RVC

[1155] E. Grassmann and J. Rokne. The range of values of a circular complex polynomial over a circular complex interval. Computing, 23(2):139–169, June 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Vajtersic:1979:FAS

[1156] M. Vajteršic. A fast algorithm for solving the first biharmonic boundary value problem. Computing, 23(2):171–178, June 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brunner:1979:NCM

[1157] H. Brunner. A note on collocation methods for Volterra integral equations of the first kind. Computing, 23(2):179–187, June 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [1099].

Bachem:1979:RA

[1158] A. Bachem and B. Korte. On the RASalgorithm. Computing, 23(2):189–198, June 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tadikamalla:1979:RSG

[1159] P. R. Tadikamalla. Random sampling from the generalized gamma distribution. Computing, 23(2):199–203, June 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Uberhuber:1979:IDC

[1160] C. W. Überhuber. Implementation of defect correction methods for stiff differential equations. Computing, 23(3): 205–232, September 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Allgower:1979:GMI

[1161] E. L. Allgower, St. F. McCormick, and D. V. Pryor. A general mesh independence principle for Newton’s method applied to second order boundary value problems. Computing, 23 (3):233–246, September 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Nickel:1979:CPB

[1162] K. Nickel. The construction of a priori bounds for the solution of a twopoint boundary value problem with finite elements. I. Computing, 23(3): 247–265, September 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Miranker:1979:HR

[1163] W. L. Miranker. Hierarchical relaxation. Computing, 23(3):267–285, September 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ritter:1979:LSC

[1164] K. Ritter. Local and superlinear convergence of a class of variable metric methods. Computing, 23(3):287–297, September 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fix:1979:PSS

[1165] W. Fix and K. Neumann. Project scheduling by special GERT networks. Computing, 23(3):299–308, September 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Murray:1979:SAM

[1166] W. Murray and M. L. Overton.

Steplength algorithms for minimizing a class of nondifferentiable functions. Computing, 23(4):309–331, December 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ansorge:1979:SVC

[1167] R. Ansorge and K. Taubert. Setvalued collectively compact operators and applications to set-valued differential equations. Computing, 23(4): 333–343, December 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Werschulz:1979:CCO

[1168] A. G. Werschulz. Computational complexity of one-step methods for a scalar autonomous differential equation. Computing, 23(4):345–355, December 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lipski:1979:MPG

[1169] W. Lipski, Jr. More on permutation generation methods. Computing, 23 (4):357–365, December 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ballin:1979:OBA

[1170] W. Ballin. Operationen zur Betriebsmittelsteuerung Aufbaünd auf Synchronisationsoperationen. (German) [Operations for resource management based on synchronization operations]. Computing, 23(4):367–380, December 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Glinert:1979:AAI

[1171] E. P. Glinert and E. Katz. Algorithm 40. An algorithm for the integral homology of certain topological groups. Computing, 23(4):381–391, December 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Cherniavsky:1979:EFI

[1172] J. C. Cherniavsky. Erratum to: “Function iteration logics and flowchart schemata” [Computing 14 (1975), no. 3, 285–312; MR 53 #14962]. Computing, 23(4):393–394, December 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [882].

Kioustelidis:1980:OSA

[1173] J. B. Kioustelidis. Optimal segmented approximations. Computing, 24(1):1–8, March 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Meyn:1980:CSB

[1174] K. H. Meyn. Convergent solution bounds for linear elliptic boundary value problems of monotone kind. Computing, 24(1):9–19, March 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bauch:1980:IIS

[1175] H. Bauch. On the iterative inclusion of solutions in characteristic initial value problems with hyperbolic differential equation z_{st} = f(s, t, z). Computing, 24(1):21–32, March 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dimitrova:1980:DEI

[1176] Neli Dimitrova. Über die Distributivgesetze der erweiterten Intervallarithmetik. (German) [On the distributive law of extended interval arithmetic]. Computing, 24(1):33–49, March 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Akritas:1980:CAT

[1177] A. G. Akritas and S. D. Danielopoulos.

On the complexity of algorithms for the translation of polynomials. Computing, 24(1):51–60, March 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Abdel-Wahab:1980:SOS

[1178] H. M. Abdel-Wahab and T. Kameda. On strictly optimal schedules for the cumulative cost-optimal scheduling problem. Computing, 24(1):61–86, March 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Berg:1980:SCT

[1179] L. Berg. On the simultaneous calculation of two zeros. Computing, 24(1):87–91, March 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ratschek:1980:RIO

[1180] H. Ratschek. Representation of interval operations by coordinates. Computing, 24(2–3):93–96, June 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Nickel:1980:ACS

[1181] K. Nickel. Arithmetic of complex sets. Computing, 24(2–3):97–105, June 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Klatte:1980:CBS

[1182] R. Klatte and Ch. Ullrich. On conversions between screens. Computing, 24(2–3):107–114, June 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Claudio:1980:CSC

[1183] D. M. Claudio. Contributions to the structure of computer arithmetic. Computing, 24(2–3):115–118, June 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Krawczyk:1980:IEI

[1184] R. Krawczyk. Interval extensions and interval iterations. Computing, 24 (2–3):119–129, June 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kaucher:1980:GIM

[1185] E. Kaucher and S. M. Rump. Generalized iteration methods for bounds of the solution of fixed point operatorequations. Computing, 24(2–3):131– 137, June 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Klatte:1980:CSA

[1186] R. Klatte and Ch. Ullrich. Complex sector arithmetic. Computing, 24 (2–3):139–148, June 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bohlender:1980:EUC

[1187] G. Bohlender. Embedding universal computer arithmetic in higher programming languages. Computing, 24 (2–3):149–160, June 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gardenes:1980:FSI

[1188] E. Gardenes and A. Trepat. Fundamentals of Sigla, an interval computing system over the completed set of intervals. Computing, 24(2–3):161–179, June 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gruner:1980:IUC

[1189] K. Grüner. Implementation of universal computer arithmetic with optimal accuracy. Computing, 24(2–3):181–193, June 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Raith:1980:EUI

[1190] M. Raith. Existence and uniqueness of inclusion isotonic centered ball extensions. Computing, 24(2–3):195–202, June 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

WolffvonGudenberg:1980:DMS

[1191] J. Wolff von Gudenberg. Determination of minimum sets of the set of zeros of a function. Computing, 24 (2–3):203–212, June 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Olschlagel:1980:IAT

[1192] D. Ölschlägel and H. Süsse. Interval analytic treatment of convex programming problems. Computing, 24 (2–3):213–225, June 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

deHoog:1980:ATB

[1193] F. R. de Hoog and R. Weiss. An approximation theory for boundary value problems on infinite intervals. Computing, 24(2–3):227–239, June 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Raptis:1980:EFM

[1194] A. D. Raptis. Exponential-fitting methods for the numerical integration of the fourth-order differential equation y^{IV} + f · y = g. Computing, 24(2–3):241–250, June 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmid:1980:NSN

[1195] H. J. Schmid. On the numerical solution of non-linear equations characterizing minimal cubature formulae. Computing, 24(2–3):251–257, June 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hadjidimos:1980:SBR

[1196] A. Hadjidimos. Some basic results on M-matrices in connection with the accelerated overrelaxation (AOR) method. Computing, 24(2–3):259–268, June 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Larson:1980:ARE

[1197] J. L. Larson and A. H. Sameh. Algorithms for roundoff error analysis — a relative error approach. Computing, 24 (4):275–297, December 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Akritas:1980:FEA

[1198] A. G. Akritas. The fastest exact algorithms for the isolation of the real roots of a polynomial equation. Computing, 24(4):299–313, December 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gekeler:1980:SGM

[1199] E. Gekeler. Stability of Galerkin multistep procedures in time-homogeneous hyperbolic problems. Computing, 24 (4):315–324, December 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lipitakis:1980:SNL

[1200] E. A. Lipitakis and D. J. Evans. Solving non-linear elliptic difference equations by extendable sparse factorization procedures. Computing, 24(4): 325–339, December 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

vanderHouwen:1980:SMF

[1201] P. J. van der Houwen and P. H. M. Wolkenfelt. On the stability of multistep formulas for Volterra integral equations of the second kind. Computing, 24(4):341–347, December 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dierckx:1980:AAA

[1202] P. Dierckx. Algorithm/Algorithmus 42. An algorithm for cubic spline fitting with convexity constraints. Computing, 24(4):349–371, December 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Albrecht:1980:CMP

[1203] R. F. Albrecht. Concept of a multiprocessor parallel processing unit. Computing, 25(1):1–16, March 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Meister:1980:WTP

[1204] B. Meister. Waiting time in a preemptive resume system with compound-Poisson input. Computing, 25(1):17– 28, March 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Toth:1980:DPA

[1205] P. Toth. Dynamic programming algorithms for the zero-one knapsack problem. Computing, 25(1):29–45, March 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schwarze:1980:AHR

[1206] J. Schwarze. An algorithm for hierarchical reduction and decomposition of a directed graph. Computing, 25(1):47–57, March 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brunner:1980:PLO

[1207] J. K. Brunner. Piecewise linear optimization. Computing, 25(1):59–76, March 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lavery:1980:NSC

[1208] J. E. Lavery. Numerical solution of a class of quasilinear hyperbolic equations by reduction to the wave equation. Computing, 25(1):77–88, March 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hsiao:1980:GCM

[1209] G. C. Hsiao, P. Kopp, and W. L. Wendland. A Galerkin collocation method for some integral equations of the first kind. Computing, 25(2):89–130, June 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Freeden:1980:IFU

[1210] W. Freeden. On integral formulas of the (unit) sphere and their application to numerical computation of integrals. Computing, 25(2):131–146, June 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Caprani:1980:MVF

[1211] O. Caprani and K. Madsen. Mean value forms in interval analysis. Computing, 25(2):147–154, June 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brauninger:1980:QNM

[1212] J. Bräuninger. A quasi-Newton method with Cholesky factorization. Computing, 25(2):155–162, June 1980. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ohnaka:1980:TPN

[1213] K. Ohnaka, Y. Isomoto, and S. Makinouchi. On a test of programs for numerical zeros of polynomials. Computing, 25(2):163–174, June 1980. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Koberl:1980:SNE

[1214] D. Köberl. The solution of nonlinear equations by the computation of fixed points with a modification of the sandwich method. Computing, 25(2):175–179, June 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Werner:1980:RNT

[1215] H. Werner. Remarks on Newton type multivariate interpolation for subsets of grids. Computing, 25(2):181–191, June 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ahrens:1980:SBP

[1216] J. H. Ahrens and U. Dieter. Sampling from binomial and Poisson distributions: a method with bounded computation times. Computing, 25(3): 193–208, September 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kemp:1980:AHT

[1217] R. Kemp. The average height of rtuply rooted planted plane trees. Computing, 25(3):209–232, September 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [1326].

Weistroffer:1980:CON

[1218] H. R. Weistroffer. Constrained optimization with normed moving truncations penalty-functions. Computing, 25 (3):233–241, September 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Voss:1980:LCG

[1219] H. Voß and U. Eckhardt. Linear convergence of generalized Weiszfeld’s method. Computing, 25(3):243–251, September 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Multhei:1980:EAS

[1220] H. N. Mülthei and B. Schorr. Error analysis for a special X-spline. Computing, 25(3):253–267, September 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Spellucci:1980:ABA

[1221] P. Spellucci. An approach to backward analysis for linear and nonlinear iterative methods. Computing, 25 (3):269–282, September 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Burmeister:1980:OIE

[1222] W. Burmeister. Optimal interval enclosing of certain sets of matrices, with application to monotone enclosing of square roots. Computing, 25 (3):283–295, September 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1980:PRI

[1223] Hans J. Stetter. Professor Rudolf Inzinger. Computing, 25(3):297–298, September 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gorenflo:1980:CDS

[1224] R. Gorenflo and Marianne Niedack. Conservative difference schemes for diffusion problems with boundary and interface conditions. Computing, 25(4): 299–316, December 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Multhei:1980:NLG

[1225] H. N. Mülthei. Numerische Lösung gewöhnlicher Differentialgleichungen mit Splinefunktionen. (German) [Numerical solution of ordinary differential equations with spline functions]. Computing, 25(4):317–335, December 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Weber:1980:NBG

[1226] H. Weber. Zur numerischen Behandlung gestörter Verzweigungsprobleme bei gewöhnlichen Differentialgleichungen. (German) [On the numerical treatment of perturbed bifurcation problems for ordinary differential equations]. Computing, 25(4):337–349, December 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Donnelly:1980:DII

[1227] J. D. P. Donnelly. Discrete invariant imbedding applied to free boundary problems. Computing, 25(4):351–361, December 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Osyczka:1980:MNO

[1228] A. Osyczka. Multicriterion network optimization problem. Computing, 25 (4):363–368, December 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kinderman:1980:NMG

[1229] A. J. Kinderman and J. F. Monahan.

New methods for generating Student’s t and gamma variables. Computing, 25 (4):369–377, December 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Martinez:1980:CSS

[1230] J. M. Martínez and T. L. Lopes. Combination of the sequential secant method and Broyden’s method with projected updates. Computing, 25(4): 379–386, December 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Devroye:1981:ATB

[1231] L. Devroye and T. Klincsek. Average time behavior of distributive sorting algorithms. Computing, 26(1):1–7, March 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Muehlbacher:1981:FTS

[1232] J. R. Muehlbacher. Full table scatter storage parallel searching. Computing, 26(1):9–18, March 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ahrens:1981:CME

[1233] J. H. Ahrens and K. D. Kohrt. Computer methods for efficient sampling from largely arbitrary statistical distributions. Computing, 26(1):19–31, March 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schneider:1981:MED

[1234] N. Schneider. Monotone Einschließung durch Verfahren vom Regula-falsi-Typ unter Verwendung eines verallgemeinerten Steigungsbegriffes. (German) [Monotone inclusion by regulafalsi methods realized by a generalized difference operator]. Computing, 26(1):33–44, March 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wolfe:1981:CSM

[1235] M. A. Wolfe. On the convergence of some methods for determining zeros of order-convex operators. Computing, 26(1):45–56, March 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Nuding:1981:SBE

[1236] E. Nuding. Schrankentreue Berechnung der Exponentialfunktion wesentlichnichtnegativer Matrizen. (German) [Barrier faithful computation of the exponential function of strictly nonnegative matrices]. Computing, 26(1):57–66, March 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Filipponi:1981:ACF

[1237] P. Filipponi. An algorithm for computing functions of triangular matrices. Computing, 26(1):67–71, March 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ioakimidis:1981:NIF

[1238] N. I. Ioakimidis. On the natural interpolation formula for Cauchy-type singular integral equations of the first kind. Computing, 26(1):73–77, March 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rohn:1981:SSI

[1239] J. Rohn. Strong solvability of interval linear programming problems. Computing, 26(1):79–82, March 1981.

CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Veselinov:1981:DFD

[1240] V. M. Veselinov. Derivatives of functions by derivatives of linear positive operators. Computing, 26(1):83–86, March 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Justen:1981:NRR

[1241] K. Justen. A note on regular resolution. Computing, 26(1):87–89, March 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bank:1981:AMM

[1242] R. E. Bank and A. H. Sherman. An adaptive, multilevel method for elliptic boundary value problems. Computing, 26(2):91–105, June 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ponisch:1981:CTP

[1243] G. Pönisch and H. Schwetlick. Computing turning points of curves implicitly defined by nonlinear equations depending on a parameter. Computing, 26(2):107–121, June 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Paine:1981:CFD

[1244] J. W. Paine, F. R. de Hoog, and R. S. Anderssen. On the correction of finite difference eigenvalue approximations for Sturm–Liouville problems. Computing, 26(2):123–139, June 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Best:1981:GQC

[1245] M. J. Best, J. Bräuninger, K. Ritter, and S. M. Robinson. A globally and quadratically convergent algorithm for general nonlinear programming problems. Computing, 26(2):141–153, June 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Overmars:1981:TGM

[1246] M. H. Overmars and J. van Leeuwen.

Two general methods for dynamizing decomposable searching problems. Computing, 26(2):155–166, June 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dekker:1981:FMA

[1247] K. Dekker. Formula manipulation in ALGOL 68 and application to Routh’s algorithm. Computing, 26(2):167–187, June 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Trummer:1981:RPP

[1248] M. R. Trummer. Reconstructing pictures from projections: on the convergence of the ART algorithm with relaxation. Computing, 26(3):189–195, September 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See note [1469].

Devroye:1981:CGP

[1249] L. Devroye. The computer generation of Poisson random variables. Computing, 26(3):197–207, September 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

vanNunen:1981:SLS

[1250] J. van Nunen and J. Wessels. Solving linear systems by methods based on a probabilistic interpretation. Computing, 26(3):209–225, September 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Alefeld:1981:BSP

[1251] G. Alefeld. Bounding the slope of polynomial operators and some applications. Computing, 26(3):227–237, September 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kioustelidis:1981:OSP

[1252] J. B. Kioustelidis. Optimal segmented polynomial L_{s}-approximations. Computing, 26(3):239–246, September 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chawla:1981:FFO

[1253] M. M. Chawla and S. R. Sharma. Families of fifth order Nyström methods for y′′ = f(x, y) and intervals of periodicity. Computing, 26(3):247–256, September 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Edenhofer:1981:ICL

[1254] J. Edenhofer and G. Schmitz. Ein implizites Charakteristikenverfahren zur Lösung von Anfangsrandwertaufgaben bei hyperbolischen Systemen und seine Konvergenz. (German) [An implicit method of characteristics for the solution of hyperbolic initial-boundary-value problems of hyperbolic systems and its convergence]. Computing, 26 (3):257–264, September 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Berg:1981:RFS

[1255] L. Berg. On real factorizations of symmetric circulant sparse matrices. Computing, 26(3):265–270, September 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schroder:1981:BDI

[1256] G. Schröder. Bemerkung zur Differentiation von intervallwertigen Funktionen. (German) [On the differentiation of interval-valued functions]. Computing, 26(3):271–274, September 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Spath:1981:CAC

[1257] H. Späth. Correction to: “Algorithm 39: clusterwise linear regression” [Computing 22 (1979), no. 4, 367– 373]. Computing, 26(3):275, September 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [1144].

Bohlender:1981:FCN

[1258] G. Bohlender, E. Kaucher, R. Klatte, U. Kulisch, W. L. Miranker, Ch. Ullrich, and J. Wolff v. Gundenberg. Fortran for contemporary numerical computation. Computing, 26(4):277–314, December 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Weber:1981:ADN

[1259] H. Weber and W. Werner. On the accurate determination of nonisolated solutions of nonlinear equations. Computing, 26(4):315–326, December 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Merten:1981:AOR

[1260] K. Merten. Asymptotisch optimale Relaxationsparameter bei nichtlinearen SOR-Verfahren (Newton-SOR-Verfahren). (German) [Asymptotic optimal relaxation parameters for nonlinear SOR-methods (Newton-SOR-methods)]. Computing, 26(4):327–332, December 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmidt:1981:OCS

[1261] J. W. Schmidt. On the R-order of coupled sequences. Computing, 26(4): 333–342, December 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bartusch:1981:AGA

[1262] M. Bartusch. An algorithm for generating all maximal independent subsets of posets. Computing, 26(4): 343–354, December 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gottwald:1981:RRI

[1263] B. A. Gottwald and G. Wanner. A reliable Rosenbrock integrator for stiff differential equations. Computing, 26 (4):355–360, December 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Devroye:1981:NLE

[1264] L. Devroye and G. T. Toussaint. A note on linear expected time algorithms for finding convex hulls. Computing, 26 (4):361–366, December 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fohmann:1981:DSU

[1265] L. H. Fohmann. Ein Durchbruch zur “semantisierung” relationaler Datenmodelle? Das Datenmodell der Sprache LEGOL-2. (German) [Breakthrough to “semanticizing” of relational data models? the LEGOL-2 data model]. Computing, 27(1):1–14, March 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ortiz:1981:OAT

[1266] E. L. Ortiz and H. Samara. An operational approach to the tau method for the numerical solution of nonlinear differential equations. Computing, 27(1):15–25, March 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Reemtsen:1981:LSA

[1267] R. Reemtsen. On level sets and an approximation problem for the numerical solution of a free boundary problem. Computing, 27(1):27–35, March 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Petkovic:1981:GRI

[1268] M. S. Petković. On a generalisation of the root iterations for polynomial complex zeros in circular interval arithmetic. Computing, 27(1):37–55, March 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Karkoschka:1981:EAR

[1269] E. Karkoschka and P. Werner. Einige Ausnahmen zur Rosserschen Regel in der Theorie der Riemannschen Zetafunktion. (German) [Some failures of Rosser’s rule in the theory of the Riemann zeta function]. Computing, 27(1):57–69, March 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tsamasphyros:1981:ECD

[1270] G. Tsamasphyros and P. S. Theocaris. Equivalence and convergence of direct and indirect methods for the numerical solution of singular integral equations. Computing, 27(1):71–80, March 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ioakimidis:1981:NED

[1271] N. I. Ioakimidis. On the numerical evaluation of derivatives of Cauchy principal value integrals. Computing, 27(1):81–88, March 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Truszczynski:1981:SAF

[1272] M. Truszczyński. A simple algorithm for finding a cycle of length greater than three and without diagonals. Computing, 27(1):89–91, March 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Martello:1981:HAM

[1273] S. Martello and P. Toth. Heuristic algorithms for the multiple knapsack problem. Computing, 27(2):93–112, June 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Deak:1981:EMR

[1274] I. Deák. An economical method for random number generation and a normal generator. Computing, 27(2):113–121, June 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Taubert:1981:CMM

[1275] K. Taubert. Converging multistep methods for initial value problems involving multivalued maps. Computing, 27(2):123–136, June 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Qi:1981:IBS

[1276] L. Qi. Interval boxes of solutions of nonlinear systems. Computing, 27(2):137–144, June 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hohn:1981:SRD

[1277] W. Höhn and H.-D. Mittelmann. Some remarks on the discrete maximumprinciple for finite elements of higher order. Computing, 27(2):145–154, June 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Yanai:1981:IMO

[1278] H. Yanai, M. Ozawa, and S. Kaneko.

Interpolation methods in one dimensional optimization. Computing, 27 (2):155–163, June 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Struwe:1981:SOD

[1279] M. Struwe. On the solvability of ordinary differential equations with only partially continuous right-hand side by difference methods. Computing, 27(2):165–167, June 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1981:ECO

[1280] H. J. Stetter and E. Weinmüller. On the error control in ODE solvers with local extrapolation. Computing, 27 (2):169–177, June 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Carpaneto:1981:ASB

[1281] G. Carpaneto and P. Toth. Algorithm for the solution of the bottleneck assignment problem. Computing, 27(2):179–187, June 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Andrade:1981:CLM

[1282] Célia Andrade, Neide Bertoldi Franco, and S. McKee. Convergence of linear multistep methods for Volterra first kind equations with k(t, t) ≡ 0. Computing, 27(3):189–204, September 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Alt:1981:CBC

[1283] H. Alt and J. van Leeuwen. The complexity of basic complex operations. Computing, 27(3):205–215, September 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Miller:1981:RGU

[1284] W. Miller. A remark on gradual underflow. Computing, 27(3):217–225, September 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Goodman:1981:SME

[1285] R. Goodman. Some models of error in floating point multiplication. Computing, 27(3):227–236, September 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Miel:1981:UVK

[1286] G. Miel. An updated version of the Kantorovich theorem for Newton’s method. Computing, 27(3):237–244, September 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Baboolal:1981:CEA

[1287] S. Baboolal and G. A. Watson. Computational experience with an algorithm for discrete L_{1} approximation. Computing, 27(3):245–252, September 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Robinson:1981:AAC

[1288] I. Robinson and Elise de Doncker.

Algorithm 45/Algorithmus 45: Automatic computation of improper integrals over a bounded or unbounded planar region. Computing, 27(3): 253–284, September 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [977, 1369].

Brunner:1981:PLO

[1289] J. K. Brunner. Piecewise linear optimization: Duality and numerical results. Computing, 27(4):285–298, December 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mattheij:1981:EES

[1290] R. M. M. Mattheij. Estimates for the errors in the solutions of linear boundary value problems, due to perturbations. Computing, 27(4):299–318, December 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kolbig:1981:PFS

[1291] K. S. Kölbig and F. Schwarz. On positive function series. Computing, 27 (4):319–337, December 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rokne:1981:CFI

[1292] J. Rokne. The centered form for interval polynomials. Computing, 27(4): 339–348, December 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Loukakis:1981:DFS

[1293] E. Loukakis and C. Tsouros. A depth first search algorithm to generate the family of maximal independent sets of a graph lexicographically. Computing, 27 (4):349–366, December 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brucker:1981:MML

[1294] P. Brucker. Minimizing maximum lateness in a two-machine unit-time job shop. Computing, 27(4):367–370, December 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sundermann:1981:MPT

[1295] B. Sündermann. On a minimal property of trigonometric interpolation at equidistant nodes. Computing, 27(4): 371–372, December 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Acharya:1981:NDC

[1296] B. P. Acharya and R. N. Das. Numerical determination of Cauchy principal value integrals. Computing, 27 (4):373–378, December 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wallach:1982:ASP

[1297] Y. Wallach. Alternating sequentialparallel calculation of eigenvalues for symmetric matrices. Computing, 28(1): 1–16, March 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rokne:1982:CCC

[1298] J. Rokne and T. Wu. The circular complex centered form. Computing, 28(1):17–30, March 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See note [1372].

Kaucher:1982:MFP

[1299] E. Kaucher and S. M. Rump. Emethods for fixed point equations f(x) = x. Computing, 28(1):31–42, March 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Beyn:1982:SSD

[1300] W. J. Beyn and J. Lorenz. Spurious solutions for discrete superlinear boundary value problems. Computing, 28(1):43–51, March 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Blum:1982:FEM

[1301] H. Blum and M. Dobrowolski. On finite element methods for elliptic equations on domains with corners. Computing, 28(1):53–63, March 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Muller:1982:RDP

[1302] K.-H. Müller. A recurrence for Dirichlet problems. Computing, 28(1):65–68, March 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kioustelidis:1982:MEO

[1303] J. B. Kioustelidis. Monotonicity of the error of optimal segmented approximations. Computing, 28(1):69–74, March 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dunham:1982:RAC

[1304] C. B. Dunham. Remez algorithm for Chebyshev approximation with interpolation. Computing, 28(1):75–78, March 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Monch:1982:IRA

[1305] W. Mönch. Iterative refinement of approximations to a generalized inverse of a matrix. Computing, 28(1):79–87, March 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Leuprecht:1982:PAR

[1306] H. Leuprecht and W. Oberaigner. Parallel algorithms for the rounding exact summation of floating point numbers. Computing, 28(2):89–104, June 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ratschek:1982:LIE

[1307] H. Ratschek and W. Sauer. Linear interval equations. Computing, 28(2):105–115, June 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Krawczyk:1982:ZFI

[1308] R. Krawczyk and K. Nickel. Die zentrische Form in der Intervallarithmetik, ihre quadratische Konvergenz und ihre Inklusionsisotonie. (German) [The centered form in interval arithmetic: quadratic convergence and inclusion isotonicity]. Computing, 28 (2):117–137, June 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hornung:1982:DMP

[1309] R. Hornung. Discrete minimax problem — algorithms and numerical comparisons. Computing, 28(2):139–154, June 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Luksan:1982:NCM

[1310] L. Lukšan. New combined method for unconstrained minimization. Computing, 28(2):155–169, June 1982. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Waldvogel:1982:BMV

[1311] P. Waldvogel. Bisection for Ax = λBx with matrices of variable band width. Computing, 28(2):171–180, June 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Donnelly:1982:DII

[1312] J. D. P. Donnelly. Discrete invariant imbedding and front-tracking. Computing, 28(2):181–188, June 1982. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bieri:1982:RSP

[1313] H. Bieri and W. Nef. A recursive sweepplane algorithm, determining all cells of a finite division of R^{d}. Computing, 28 (3):189–198, September 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Keller:1982:NSI

[1314] G. Keller. Numerical solution of initialvalue problems by collocation methods using generalized piecewise functions. Computing, 28(3):199–211, September 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lavery:1982:SQH

[1315] J. E. Lavery. Solution of quasilinear hyperbolic initial-boundary-value problems by the method of pseudolinear equations. Computing, 28(3): 213–223, September 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Asaithambi:1982:CRV

[1316] N. S. Asaithambi, Shen Zuhe, and R. E. Moore. On computing the range of values. Computing, 28(3): 225–237, September 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rokne:1982:OCB

[1317] J. Rokne. Optimal computation of the Bernstein algorithm for the bound of an interval polynomial. Computing, 28 (3):239–246, September 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Reuss:1982:RAK

[1318] W. J. Reuß. Eine repräsentative Auswahl von Knotenmatrizen zu Klassen isomorpher Graphen. (German) [A representative selection of matrices describing classes of isomorphic graphs]. Computing, 28(3):247–255, September 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Slominski:1982:PAC

[1319] L. Slominski. Probabilistic analysis of combinatorial algorithms: a bibliography with selected annotations. Computing, 28(3):257–267, September 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fayard:1982:AAA

[1320] D. Fayard and G. Plateau. Algorithm / Algorithmus 47 an algorithm for the solution of the 0-1 knapsack problem. Computing, 28(3):269–287, September 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Leitsch:1982:FIW

[1321] A. Leitsch. Fiducial intervals for the waiting time in batch and timesharing systems. Computing, 28(4): 289–303, December 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Vaquero:1982:OAA

[1322] A. Vaquero and J. M. Troya. Optimization and approximation algorithm for placement of records on linear storage devices. Computing, 28(4): 305–316, December 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bownds:1982:TPS

[1323] J. M. Bownds. Theory and performance of a subroutine for solving Volterra integral equations. Computing, 28(4):317–332, December 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Breitenecker:1982:SLB

[1324] F. Breitenecker. On the solution of linear boundary value problems by extended invariant imbedding. Computing, 28(4):333–343, December 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gipser:1982:GCC

[1325] M. Gipser. On global convergence of coordinate relaxation in the case of an unsymmetrical, diagonally dominant Jacobian. Computing, 28(4): 345–361, December 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Prodinger:1982:NRR

[1326] H. Prodinger. A note on a result of R. Kemp on r-tuply rooted planted plane trees. Computing, 28(4):363–366, December 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [1217].

Devroye:1982:NAD

[1327] L. Devroye. A note on the average depth of tries. Computing, 28 (4):367–371, December 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Raptis:1982:TSM

[1328] A. D. Raptis. Two-step methods for the numerical solution of the Schrödinger equation. Computing, 28(4):373–378, December 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ginsburg:1982:TAG

[1329] S. Ginsburg and O. Mayer. Tree acceptors and grammar forms. Computing, 29(1):1–9, March 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bagchi:1982:AIR

[1330] A. Bagchi and E. M. Reingold. Aspects of insertion in random trees. Computing, 29(1):11–29, March 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mateti:1982:PAS

[1331] P. Mateti and N. Deo. Parallel algorithms for the single source shortest path problem. Computing, 29(1):31–49, March 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Multhei:1982:SKM

[1332] H. N. Mülthei. A-stabile Kollokationsverfahren mit mehrfachen Knoten. (German) [A-stable collocation methods with multiple nodes]. Computing, 29(1):51–61, March 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sisser:1982:IIH

[1333] F. S. Sisser. Inverting an interval Hessian of a factorable function. Computing, 29(1):63–72, March 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Burmeister:1982:OCS

[1334] W. Burmeister and J. W. Schmidt. On the R-order of coupled sequences. II. Computing, 29(1):73–81, March 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sottas:1982:PQF

[1335] G. Sottas. On the positivity of quadrature formulas with Jacobi abscissas. Computing, 29(1):83–88, March 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bunke:1982:GPS

[1336] H. Bunke. On the generative power of sequential and parallel programmed graph grammars. Computing, 29(2):89– 112, June 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hamacher:1982:DGF

[1337] H. Hamacher. Decomposition of group flows in regular matroids. Computing, 29(2):113–133, June 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Boese:1982:PET

[1338] G. Boese. An a priori estimate for the truncation error of a continued fraction expansion to the Gaussian error function. Computing, 29(2):135–152, June 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Strehmel:1982:BSA

[1339] K. Strehmel and R. Weiner. Behandlung steifer Anfangswertprobleme gewöhnlicher Differentialgleichungen mit adaptiven Runge–Kutta-Methoden. (German) [Treatment of stiff initial value problems for ordinary differential equations with adaptive Runge–Kutta methods]. Computing, 29(2):153–165, June 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Evans:1982:PJM

[1340] D. J. Evans and N. M. Missirlis. On the preconditioned Jacobi method for solving large linear systems. Computing, 29(2):167–173, June 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Spath:1982:FAC

[1341] H. Späth. A fast algorithm for clusterwise linear regression. Computing, 29 (3):175–181, September 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rump:1982:SNS

[1342] S. M. Rump. Solving nonlinear systems with least significant bit accuracy. Computing, 29(3):183–200, September 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Melhem:1982:CMD

[1343] R. G. Melhem and W. C. Rheinboldt. A comparison of methods for determining turning points of nonlinear equations. Computing, 29(3):201–226, September 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hoffmann:1982:CSS

[1344] U. Hoffmann. A class of simple stochastic online bin packing algorithms. Computing, 29(3):227–239, September 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Widjaja:1982:ESA

[1345] H. Widjaja. An effective structured approach to finding optimal partitions of networks. Computing, 29(3):241–262, September 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dobes:1982:FFT

[1346] K. Dobes. Fast Fourier transforms with recursively generated trigonometric functions. Computing, 29(3): 263–276, September 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ecker:1982:MGE

[1347] A. Ecker. Über die mathematischen Grundlagen einiger Chiffrierverfahren. (German) [On the mathematical foundations of some cryptosystems]. Computing, 29(4):277–287, December 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kleinmichel:1982:GSC

[1348] H. Kleinmichel, C. Richter, and K. Schönefeld. On the global and superlinear convergence of a discretized version of Wilson’s method. Computing, 29(4):289–307, December 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Haussler:1982:VSH

[1349] W. M. Häußler. Variable Schrittweitensteuerungen für die Homotopiemethode bei adäquaten nichtlinearen Ausgleichsproblemen. (German) [Variable stepsize control for the homotopy method applied to adequate nonlinear regression]. Computing, 29(4):309–326, December 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Neuman:1982:CCF

[1350] E. Neuman. Calculation of complex Fourier coefficients using natural splines. Computing, 29(4):327–336, December 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Monegato:1982:NEO

[1351] G. Monegato. The numerical evaluation of one-dimensional Cauchy principal value integrals. Computing, 29 (4):337–354, December 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Popovski:1982:NKM

[1352] D. B. Popovski. A note on King’s method F for finding a bracketed root. Computing, 29(4):355–359, December 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gfrerer:1982:UCI

[1353] H. Gfrerer. Uniform convergence of interpolation by cubic splines. Computing, 29(4):361–364, December 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lee:1982:SBS

[1354] E. T. Y. Lee. A simplified B-spline computation routine. Computing, 29 (4):365–371, December 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See comments [1548].

Danielopoulos:1982:CEP

[1355] S. D. Danielopoulos. On the cost of evaluating polynomials and their derivatives. Computing, 29(4):373–380, December 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Yassouridis:1983:MAE

[1356] A. Yassouridis and R. Koller. Mehrstufige Auswahldisziplin in einem M/G/1-Processor-sharing-System mit Prioritäten. (German) [Multilevel processorsharing algorithm for M/G/1 systems with priorities]. Computing, 30(1):1– 18, March 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ebert:1983:OVB

[1357] J. Ebert. st-ordering the vertices of biconnected graphs. Computing, 30(1):19–33, March 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Strehmel:1983:ANR

[1358] K. Strehmel and R. Weiner. Adaptive Nyström–Runge–Kutta-Methoden für gewöhnliche Differentialgleichungssysteme zweiter Ordnung. (German) [Adaptive Nystrom–Runge–Kutta methods for second-order ordinary differential equations]. Computing, 30(1):35– 47, March 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Makroglou:1983:BBM

[1359] A. Makroglou. A block-by-block method for the numerical solution of Volterra delay integro-differential equations. Computing, 30(1):49–62, March 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Akritas:1983:EAP

[1360] A. G. Akritas and King H. Ng. Exact algorithms for polynomial real root approximation using continued fractions. Computing, 30(1):63–76, March 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gold:1983:MPA

[1361] R. Gold and J. Sattler. Modifikationen des Pollard-Algorithmus. (German) [Modifications of the Pollard algorithm]. Computing, 30(1):77–89, March 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sattler:1983:EFM

[1362] J. Sattler and C.-P. Schnorr. Ein Effizienzvergleich der Faktorisierungsverfahren von Morrison-Brillhart und Schroeppel. (German) [An efficiency comparison of the integer factoring algorithms of Morrison-Brillhart and Schroeppel]. Computing, 30(2):91–110, June 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Devroye:1983:MIR

[1363] L. Devroye. Moment inequalities for random variables in computational geometry. Computing, 30(2):111–119, June 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Frank:1983:AID

[1364] R. Frank, J. Hertling, and J. P. Monnet. The application of iterated defect correction to variational methods for elliptic boundary value problems. Computing, 30(2):121–135, June 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hoenders:1983:CEN

[1365] B. J. Hoenders and C. H. Slump. On the calculation of the exact number of zeroes of a set of equations. Computing, 30(2):137–147, June 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bialas:1983:CPH

[1366] S. Białas. On certain properties of Hurwitz determinants for interval polynomials. Computing, 30(2):149–155, June 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Burmeister:1983:OCS

[1367] W. Burmeister and J. W. Schmidt. On the R-order of coupled sequences. III. Computing, 30(2):157–169, June 1983.

CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Milovanovic:1983:COM

[1368] G. V. Milovanović and M. S. Petković.

On the convergence order of a modified method for simultaneous finding polynomial zeros. Computing, 30(2):171–178, June 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Yserentant:1983:RNC

[1369] H. Yserentant. A remark on the numerical computation of improper integrals. Computing, 30(2):179–183, June 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [1288].

Best:1983:NGV

[1370] D. J. Best. A note on gamma variate generators with shape parameter less than unity. Computing, 30(2):185–188, June 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rump:1983:LSB

[1371] S. M. Rump and H. Böhm. Least significant bit evaluation of arithmetic expressions in single-precision. Computing, 30(3):189–199, September 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rokne:1983:NCC

[1372] J. Rokne and T. Wu. A note on the circle complex centered form. Computing, 30(3):201–211, September 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See

Monch:1983:SMS

[1373] W. Mönch. Secant methods for sparse systems of nonlinear equations with a special structure. Computing, 30(3): 213–223, September 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gfrerer:1983:GCA

[1374] H. Gfrerer, J. Guddat, and Hj. Wacker. A globally convergent algorithm based on imbedding and parametric optimization. Computing, 30(3):225–252, September 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Korner:1983:EBB

[1375] F. Körner. An efficient branch and bound algorithm to solve the quadratic integer programming problem. Computing, 30(3):253–260, September 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Apostolatos:1983:AOV

[1376] N. Apostolatos and V. Vlassopoulos. Approximation der optimalen Vorhersage der Ausgabe stochastischer Quellen. (German) [Approximation of the optimal prediction of the output of stochastic sources]. Computing, 30 (3):261–273, September 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dunham:1983:OSR

[1377] C. B. Dunham. The one-sided Remez algorithm. Computing, 30(3): 275–278, September 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Acharya:1983:AEM

[1378] B. P. Acharya and R. N. Das. Approximate evaluation of multiple complex integrals of analytic functions. Computing, 30(3):279–283, September 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Muller:1983:REA

[1379] K.-H. Müller. Rounding error analysis of Horner’s scheme. Computing, 30 (4):285–303, December 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Yserentant:1983:CMM

[1380] H. Yserentant. On the convergence of multi-level methods for strongly nonuniform families of grids and any number of smoothing steps per level. Computing, 30(4):305–313, December 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Luksan:1983:VMM

[1381] L. Lukšan. Variable metric methods for linearly constrained nonlinear minimax approximation. Computing, 30 (4):315–334, December 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hock:1983:CPE

[1382] W. Hock and K. Schittkowski. A comparative performance evaluation of 27 nonlinear programming codes. Computing, 30(4):335–358, December 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmitz:1983:ITC

[1383] L. Schmitz. An improved transitive closure algorithm. Computing, 30(4): 359–371, December 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

King:1983:IVV

[1384] R. F. King. Improving the Van de Vel root-finding method. Computing, 30 (4):373–378, December 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Otoo:1983:SSE

[1385] E. J. Otoo and T. H. Merrett. Storage scheme for extendible arrays. Computing, 31(1):1–9, March 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Doberkat:1983:CME

[1386] E. E. Doberkat. Continuous models that are equivalent to randomness for the analysis of many sorting algorithms. Computing, 31(1):11–31, March 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Papworth:1983:ETF

[1387] D. G. Papworth. Exact tests of fit for a Poisson distribution. Computing, 31(1):33–45, March 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zlatev:1983:CCG

[1388] Z. Zlatev. Consistency and convergence of general linear multistep variable stepsize variable formula methods. Computing, 31(1):47–67, March 1983.

CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ansorge:1983:RCI

[1389] R. Ansorge and K. Taubert. Remarks on composite integration methods for ordinary differential equations. Computing, 31(1):69–82, March 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Carpaneto:1983:ASA

[1390] G. Carpaneto and P. Toth. Algorithm for the solution of the assignment problem for sparse matrices. Computing, 31(1):83–94, March 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ortiz:1983:NSD

[1391] E. L. Ortiz and H. Samara. Numerical solution of differential eigenvalue problems with an operational approach to the tau method. Computing, 31(2):95–103, June 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tsamasphyros:1983:CSQ

[1392] G. Tsamasphyros and P. S. Theocaris.

On the convergence of some quadrature rules for Cauchy principal-value and finite-part integrals. Computing, 31(2):105–114, June 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rosenberg:1983:BP

[1393] A. L. Rosenberg and I. H. Sudborough.

Bandwidth and pebbling. Computing, 31(2):115–139, June 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bathelt:1983:CTP

[1394] P. Bathelt and F. Hofmann. A comparison of two Petri net types. Computing, 31(2):141–148, June 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hering:1983:DBK

[1395] L. Hering. Darstellung von Bézier-Kurven als B-Spline-Kurven. (German) [Representation of Bézier curves as Bspline curves]. Computing, 31(2):149– 153, June 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Prochazka:1983:CMU

[1396] W. Prochazka. Conformal mapping of the unit circle or of the upper half plane onto a polygon. Computing, 31(2):155–172, June 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Breitenecker:1983:CAI

[1397] F. Breitenecker. On the choice of an appropriate imbedding algorithm for the solution of unstable linear boundary value problems. Computing, 31 (2):173–176, June 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lautsch:1983:ISP

[1398] M. Lautsch. An implicit off-step point method for the integration of stiff differential equations. Computing, 31(2):177–183, June 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

WolffvonGudenberg:1983:CMA

[1399] J. Wolff von Gudenberg. Computing z^{y} with maximum accuracy. Computing, 31(2):185–189, June 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

El-Gindy:1983:ATD

[1400] H. El-Gindy, D. Avis, and G. Toussaint. Applications of a twodimensional hidden-line algorithm to other geometric problems. Computing, 31(3):191–202, September 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Korchemkin:1983:HPA

[1401] M. B. Korchemkin. Heuristic partitioning algorithm for a packaging problem. Computing, 31(3):203–209, September 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Munzel:1983:CBS

[1402] G. Münzel and G. Renner. Zur Charakterisierung und Berechnung von symmetrischen Kubaturformeln. (German) [On the characterization and computation of symmetric cubature formulas]. Computing, 31(3):211–230, September 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Haussler:1983:LCA

[1403] W. M. Häussler. A local convergence analysis for the Gauss–Newton and Levenberg-Morrison-Marquardt algorithms. Computing, 31(3):231–244, September 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Krawczyk:1983:IOF

[1404] R. Krawczyk. Interval operators of a function of which the Lipschitz matrix is an interval M-matrix. Computing, 31 (3):245–253, September 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Krawczyk:1983:RAC

[1405] R. Krawczyk. A remark about the convergence of interval sequences. Computing, 31(3):255–259, September 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Qin:1983:DSD

[1406] Meng Zhao Qin. Difference schemes for the dispersive equation. Computing, 31 (3):261–267, September 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Werner:1983:ARN

[1407] H. Werner. Algorithm 51. A reliable and numerically stable program for rational interpolation of Lagrange data. Computing, 31(3):269–286, September 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

vanDam:1983:ABD

[1408] W. B. van Dam, J. B. G. Frenk, and A. H. G. Rinnooy Kan. The asymptotic behaviour of a distributive sorting method. Computing, 31 (4):287–303, December 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bolch:1983:ALS

[1409] G. Bolch. Approximation von Leistungsgrössen Symmetrischer Mehrprozessorsysteme. (German) [Performance measure approximation for symmetric multiprocessor systems]. Computing, 31(4):305–315, December 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Engels:1983:SSU

[1410] G. Engels, R. Gall, M. Nagl, and W. Schaefer. Software specification using graph grammars. Computing, 31 (4):317–346, December 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Maas:1983:DIN

[1411] C. Maas. Determining the interval number of a triangle-free graph. Computing, 31(4):347–354, December 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Monien:1983:CDS

[1412] B. Monien. The complexity of determining a shortest cycle of even length. Computing, 31(4):355–369, December 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Strehmel:1983:NCC

[1413] K. Strehmel and R. Weiner. Nonlinear contractivity of a class of semi-implicit multistep methods. Computing, 31 (4):371–381, December 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmidt:1984:EMP

[1414] J. W. Schmidt and H. Schneider. Enclosing methods in perturbed nonlinear operator equations. Computing, 32(1):1–11, March 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Krawczyk:1984:III

[1415] R. Krawczyk. Interval iterations for including a set of solutions. Computing, 32(1):13–31, March 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wallach:1984:TME

[1416] Y. Wallach. On two more eigenvalue methods for an alternating sequential parallel system. Computing, 32(1):33–41, March 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Devroye:1984:RVG

[1417] L. Devroye. Random variate generation for unimodal and monotone densities. Computing, 32(1):43–68, March 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bozzini:1984:AML

[1418] M. Bozzini, F. De Tisi, and L. Lenarduzzi. An approximation method of the local type. Application to a problem of heart potential mapping. Computing, 32(1):69–80, March 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hewgill:1984:CSC

[1419] D. E. Hewgill. Computing surfaces of constant mean curvature with singularities. Computing, 32(1):81–92, March 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Miranker:1984:SRC

[1420] W. L. Miranker and A. Winkler. Spacetime representations of computational structures. Computing, 32(2):93–114, June 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Frank:1984:AEE

[1421] R. Frank, F. Macsek, and C. W.

Überhuber. Asymptotic error expansions for defect correction iterates. Computing, 32(2):115–125, June 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Menzel:1984:QCM

[1422] R. Menzel and G. Pönisch. A quadratically convergent method for computing simple singular roots and its application to determining simple bifurcation points. Computing, 32(2):127–138, June 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Evans:1984:PIM

[1423] D. J. Evans. On preconditioned iterative methods for solving (A−λB)x = 0. Computing, 32(2):139–152, June 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zietak:1984:SLM

[1424] K. Zietak. The l_{p}-solution of the linear matrix equation AX + Y B = C. Computing, 32(2):153–162, June 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Neta:1984:HOM

[1425] B. Neta and H. D. Victory, Jr. A higher order method for determining nonisolated solutions of a system of nonlinear equations. Computing, 32(2):163–166, June 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sideridis:1984:NSM

[1426] A. B. Sideridis. A numerical solution of the membrane eigenvalue problem. Computing, 32(2):167–176, June 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Piessens:1984:CFT

[1427] R. Piessens and Maria Branders. Computation of Fourier transform integrals using Chebyshev series expansions. Computing, 32(2):177–186, June 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Jovanovic:1984:NAS

[1428] A. Jovanović and D. Danilović. A new algorithm for solving the tree isomorphism problem. Computing, 32(3): 187–198, September 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gfrerer:1984:GCD

[1429] H. Gfrerer. Globally convergent decomposition methods for nonconvex optimization problems. Computing, 32(3): 199–227, September 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

vanVeldhuizen:1984:SKR

[1430] M. van Veldhuizen. D-stability and Kaps–Rentrop-methods. Computing, 32(3):229–237, September 1984. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Evans:1984:GEM

[1431] D. J. Evans and A. R. Abdullah. The group explicit method for the solution of Burgers’ equation. Computing, 32 (3):239–253, September 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lipitakis:1984:GEL

[1432] E. A. Lipitakis. Generalized extended to the limit sparse factorization techniques for solving unsymmetric finite element systems. Computing, 32(3): 255–270, September 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dagnino:1984:CNW

[1433] Catterina Dagnino and Cinzia Fiorentino.

Computation of nodes and weights of extended Gaussian rules. Computing, 32(3):271–278, September 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Atkinson:1984:DBR

[1434] H. H. Atkinson, I. Gargantini, and M. V. S. Ramanath. Determination of the 3D border by repeated elimination of internal surfaces. Computing, 32 (4):279–295, December 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Scholz:1984:NSO

[1435] R. Scholz. Numerical solution of the obstacle problem by the penalty method. Computing, 32(4):297–306, December 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bellen:1984:CMB

[1436] A. Bellen and M. Zennaro. A collocation method for boundary value problems of differential equations with functional arguments. Computing, 32 (4):307–318, December 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dierckx:1984:ASD

[1437] P. Dierckx. Algorithms for smoothing data on the sphere with tensor product splines. Computing, 32(4): 319–342, December 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Haverkamp:1984:KAI

[1438] R. Haverkamp. Zur Konvergenz der Ableitungen von Interpolationspolynomen. (German) [Convergence of derivatives of interpolating polynomials]. Computing, 32(4):343–355, December 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Toussaint:1984:OAC

[1439] G. T. Toussaint. An optimal algorithm for computing the minimum vertex distance between two crossing convex polygons. Computing, 32(4): 357–364, December 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Freeden:1984:ECS

[1440] W. Freeden and R. Reuter. Exact computation of spherical harmonics. Computing, 32(4):365–378, December 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Akl:1984:OPA

[1441] S. G. Akl. Optimal parallel algorithms for computing convex hulls and for sorting. Computing, 33(1):1–11, March 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lausen:1984:ICC

[1442] G. Lausen. Integrated concurrency control in shared B-trees. Computing, 33(1):13–26, March 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Petkovic:1984:RCA

[1443] Ljiljana Petković and M. Petković. On the kth root in circular arithmetic. Computing, 33(1):27–35, March 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Burmeister:1984:DCR

[1444] W. Burmeister and J. W. Schmidt.

Determination of the cone radius for positive concave operators. Computing, 33(1):37–49, March 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ypma:1984:DNL

[1445] T. J. Ypma. Difference Newton-like methods under weak continuity conditions. Computing, 33(1):51–64, March 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Moret:1984:NNT

[1446] I. Moret. A note on Newton type iterative methods. Computing, 33(1):65–73, March 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dietze:1984:ERF

[1447] S. Dietze. An example for the regulafalsi method with an asymptotic cycle. Computing, 33(1):75–81, March 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brugner:1984:REA

[1448] G. Brügner. Rounding error analysis of interpolation procedures. Computing, 33(1):83–87, March 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Knuth:1984:ABZ

[1449] D. E. Knuth. An algorithm for Brownian zeroes. Computing, 33(1):89–94, March 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Derigs:1984:ASS

[1450] U. Derigs. Alternate strategies for solving bottleneck assignment problems — analysis and computational results. Computing, 33(2):95–106, June 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Frank:1984:IDC

[1451] R. Frank, F. Macsek, and C. W. Überhuber. Iterated defect correction for differential equations: Part II: Numerical experiments. Computing, 33(2):107–129, June 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Strehmel:1984:LFM

[1452] K. Strehmel and R. Weiner. Lokale Fehlerschätzung mittels modifizierter Richardson-Extrapolation in linear impliziten Einschrittverfahren. (German) [Estimation of the local error by a modified Richardson extrapolation in linear implicit one-step methods]. Computing, 33(2):131–140, June 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Jahn:1984:MFI

[1453] K.-U. Jahn. Maximale Fixpunkte von Intervallfunktionen. (German) [Maximal fixed points of interval functions]. Computing, 33(2):141–151, June 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schwandt:1984:SII

[1454] H. Schwandt. A symmetric iterative interval method for systems of nonlinear equations. Computing, 33(2):153–164, June 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Alefeld:1984:CFM

[1455] G. Alefeld. The centered form and the mean value form—a necessary condition that they yield the range. Computing, 33(2):165–169, June 1984. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bohm:1984:EES

[1456] W. Böhm. Efficient evaluation of splines. Computing, 33(2):171–177, June 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Yamamoto:1984:SWM

[1457] T. Yamamoto. The Symm-Wilkinson method for improving an approximate eigenvalue and its associated eigenvector. Computing, 33(2):179–184, June 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

vanVeldhuizen:1984:AEG

[1458] M. van Veldhuizen. Asymptotic expansions of the global error for the implicit midpoint rule (stiff case). Computing, 33(2):185–192, June 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Czech:1984:RAA

[1459] Z. Czech. A region analysis algorithm for the live variables problem. Computing, 33(3–4):193–210, September 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tinhofer:1984:LTT

[1460] G. Tinhofer and H. Schreck. Linear time tree codes. Computing, 33(3–4): 211–225, September 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tzschach:1984:VWG

[1461] H. Tzschach. M-Vollständigkeit von Warteschlangensystemen. (German) [M-completeness of queueing systems]. Computing, 33(3–4):227–235, September 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

West:1984:SMT

[1462] G. West. Eine schnelle Mellin-Transformation. (German) [A fast Mellin transformation]. Computing, 33 (3–4):237–245, September 1984. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Devroye:1984:SAG

[1463] L. Devroye. A simple algorithm for generating random variates with a logconcave density. Computing, 33(3–4): 247–257, September 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Anderssen:1984:FDM

[1464] R. S. Anderssen and F. R. de Hoog. Finite difference methods for the numerical differentiation of non-exact data. Computing, 33(3–4):259–267, September 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lamp:1984:GCI

[1465] U. Lamp, T. Schleicher, E. Stephan, and W. L. Wendland. Galerkin collocation for an improved boundary element method for a plane mixed boundary value problem. Computing, 33(3– 4):269–296, September 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Verwer:1984:CML

[1466] J. G. Verwer and J. M. Sanz-Serna.

Convergence of method of lines approximations to partial differential equations. Computing, 33(3–4):297–313, September 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chen:1984:QNM

[1467] Dao Qi Chen and R. P. Tewarson. A quasi-Newton method with sparse triple factorization for unconstrained minimization. Computing, 33(3–4): 315–329, September 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Cornelius:1984:CRV

[1468] H. Cornelius and R. Lohner. Computing the range of values of real functions with accuracy higher than second order. Computing, 33(3–4): 331–347, September 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Trummer:1984:NAR

[1469] M. R. Trummer. A note on the ART of relaxation. Computing, 33(3–4): 349–352, September 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [1248].

Martinez:1984:QNM

[1470] J. M. Martínez. A quasi-Newton method with modification of one column per iteration. Computing, 33(3– 4):353–362, September 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Akrivis:1984:DQF

[1471] G. Akrivis and K.-J. Förster. On the definiteness of quadrature formulae of Clenshaw–Curtis type. Computing, 33(3–4):363–366, September 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ansorge:1984:CBC

[1472] R. Ansorge. Connections between the Cimmino-method and the Kaczmarzmethod for the solution of singular and regular systems of equations. Computing, 33(3–4):367–375, September 1984. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Neumann:1985:EPN

[1473] K. Neumann. EOR project networks.

Computing, 34(1):1–15, March 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kaps:1985:RMS

[1474] P. Kaps, S. W. H. Poon, and T. D.

Bui. Rosenbrock methods for stiff ODEs: a comparison of Richardson extrapolation and embedding technique. Computing, 34(1):17–40, March 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Cuyt:1985:MRI

[1475] Annie A. M. Cuyt and Brigitte M. Verdonk. Multivariate rational interpolation. Computing, 34(1):41–61, March 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Menzel:1985:SNL

[1476] R. Menzel. On solving nonlinear least-squares problems in case of rankdeficient Jacobians. Computing, 34 (1):63–72, March 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brewster:1985:GCN

[1477] M. E. Brewster and R. Kannan. Global convergence of nonlinear successive overrelaxation via linear theory. Computing, 34(1):73–79, March 1985. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brewster:1985:VRP

[1478] M. E. Brewster and R. Kannan. Varying relaxation parameters in nonlinear successive overrelaxation. Computing, 34(1):81–85, March 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Cohen:1985:BAC

[1479] A. M. Cohen. A bounding approach to calculating α^{1/p}. Computing, 34(1):87– 89, March 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schwarz:1985:ADS

[1480] F. Schwarz. Automatically determining symmetries of partial differential equations. Computing, 34(2):91–106, June 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See addendum [1553].

Gabutti:1985:ACG

[1481] B. Gabutti. An algorithm for computing generalized Euler’s transformations of series. Computing, 34(2):107–116, June 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Demmel:1985:IAS

[1482] J. W. Demmel and F. Krückeberg. An interval algorithm for solving systems of linear equations to prespecified accuracy. Computing, 34(2):117–129, June 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wildenauer:1985:NMA

[1483] P. Wildenauer. A new method for automatical computation of error bounds for the set of all solutions of nonlinear boundary value problems. Computing, 34(2):131–154, June 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Anwar:1985:AAP

[1484] M. N. Anwar and M. N. El Tarazi.

Asynchronous algorithms for Poisson’s equation with nonlinear boundary conditions. Computing, 34(2):155–168, June 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Merschen:1985:MQF

[1485] B. A. Merschen. Minimal quadrature formulae for the spaces H_{2}^{R} and L_{2}^{R}. A unified interpolatory approach. Computing, 34(2):169–177, June 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Shampine:1985:LEE

[1486] L. F. Shampine. Local error estimation by doubling. Computing, 34(2):179–190, June 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rote:1985:SAA

[1487] Günter Rote. A systolic array algorithm for the algebraic path problem (shortest paths; matrix inversion). Computing, 34(3):191–219, September 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Christiansen:1985:MSF

[1488] Søren Christiansen. Modifications of some first kind integral equations with logarithmic kernel to improve numerical conditioning. Computing, 34(3): 221–242, September 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Krawczyk:1985:CFI

[1489] R. Krawczyk. Centered forms and interval operators. Computing, 34(3): 243–259, September 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rokne:1985:LCE

[1490] J. Rokne. A low complexity explicit rational centered form. Computing, 34 (3):261–263, September 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fehlberg:1985:SON

[1491] E. Fehlberg. Some old and new Runge– Kutta formulas with stepsize control and their error coefficients. Computing, 34(3):265–270, September 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Filippi:1985:RKN

[1492] S. Filippi and J. Gräf. Ein Runge– Kutta–Nyström-Formelpaar der Ordnung 11 (12) für Differentialgleichungen der Form y′′ = f(x, y). (German) [A Runge–Kutta–Nystrom formulapair of order 11 (12) for differential equations of the form y′′ = f(x, y)]. Computing, 34(3):271–282, September 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Akritas:1985:CRS

[1493] A. G. Akritas and S. D. Danielopoulos. A converse rule of signs for polynomials. Computing, 34(3):283–286, September 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bieri:1985:SPA

[1494] H. Bieri and W. Nef. A sweep-plane algorithm for computing the Eulercharacteristic of polyhedra represented in Boolean form. Computing, 34(4): 287–302, December 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Majster-Cederbaum:1985:ERO

[1495] Mila E. Majster-Cederbaum. Elimination of redundant operations in relational queries with general selection operators. Computing, 34(4):303–323, December 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Barlow:1985:RED

[1496] J. L. Barlow and E. H. Bareiss. On roundoff error distributions in floating point and logarithmic arithmetic. Computing, 34(4):325–347, December 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Barlow:1985:PEA

[1497] J. L. Barlow and E. H. Bareiss. Probabilistic error analysis of Gaussian elimination in floating point and logarithmic arithmetic. Computing, 34(4): 349–364, December 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Neumaier:1985:IPR

[1498] A. Neumaier. Inner product rounding error analysis in the presence of underflow. Computing, 34(4):365–373, December 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hahn:1985:STS

[1499] W. Hahn, K. Mohr, and U. Schauer. Some techniques for solving linear equation systems with guarantee. Computing, 34(4):375–379, December 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Heiss:1985:OUU

[1500] H.-U. Heiss and G. Totzauer. Optimizing utilization under response time constraints. Computing, 35(1):1–12, March 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Marz:1985:IVP

[1501] Roswitha März. On initial value problems in differential-algebraic equations and their numerical treatment. Computing, 35(1):13–37, March 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Yserentant:1985:HBF

[1502] Harry Yserentant. Hierarchical bases of finite-element spaces in the discretization of nonsymmetric elliptic boundary value problems. Computing, 35(1):39– 49, March 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Merz:1985:MNV

[1503] J. Merz. Ein modifiziertes Newton-Verfahren zur Lösung des Hochrechnungsproblems nach dem Prinzip des minimalen Informationsverlustes. (German) [A modified Newton algorithm for the adjustment of microdata with minimum information loss]. Computing, 35(1):51–61, March 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Shearer:1985:SAS

[1504] J. M. Shearer and M. A. Wolfe. Some algorithms for the solution of a class of nonlinear algebraic equations. Computing, 35(1):63–72, March 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wright:1985:EIA

[1505] F. J. Wright and G. Dangelmayr. Explicit iterative algorithms to reduce a univariate catastrophe to normal form. Computing, 35(1):73–83, March 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dai:1985:SOE

[1506] J. Dai. Second-order explicit characteristic difference schemes for quasilinear hyperbolic systems. Computing, 35(1):85–91, March 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pan:1985:TLE

[1507] Xiaoshu Pan and Hua Dai. A theorem for locating eigenvalues. Computing, 35(1):93–96, March 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See note [1584].

Stetter:1985:PRA

[1508] Hans J. Stetter. Professor R. Albrecht, 60 years. Computing, 35(2):97–98, June 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Burkard:1985:TSA

[1509] Rainer E. Burkard. Time-slot assignment for TDMA-systems. Computing, 35(2):99–112, June 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tinhofer:1985:CCP

[1510] G. Tinhofer and H. Schreck. Computing the characteristic polynomial of a tree. Computing, 35(2):113–125, June 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Goodman:1985:REF

[1511] R. H. Goodman, A. Feldstein, and J. Bustoz. Relative error in floatingpoint multiplication. Computing, 35 (2):127–139, June 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Auzinger:1985:AAR

[1512] W. Auzinger and H. J. Stetter. Accurate arithmetic results for decimal data on non-decimal computers. Computing, 35(2):141–151, June 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mascagni:1985:AIA

[1513] M. Mascagni and W. L. Miranker. Arithmetically improved algorithmic performance. Computing, 35(2):153– 175, June 1985. CODEN CMPTA2.

ISSN 0010-485X (print), 1436-5057 (electronic).

Alefeld:1985:HOC

[1514] G. Alefeld and R. Lohner. On higher order centered forms. Computing, 35(2):177–184, June 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Herzberger:1985:IEI

[1515] J. Herzberger. Über ein Iterationsverfahren zur Einschließung der inversen Matrix. (German) [On an iteration method for the inclusion of the inverse matrix]. Computing, 35(2):185–188, June 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mayer:1985:ESS

[1516] G. Mayer. Enclosing the solution set of linear systems with inaccurate data by iterative methods based on incomplete LU-decompositions. Computing, 35(2):189–206, June 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rokne:1985:IIL

[1517] J. Rokne. Including iterations for the lambda-matrix eigenproblem. Computing, 35(2):207–218, June 1985. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Burmeister:1985:CSI

[1518] W. Burmeister, W. Hess, and J. W.

Schmidt. Convex spline interpolants with minimal curvature. Computing, 35(2):219–229, June 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mueller:1985:VSC

[1519] H. Mueller, A. Schmitt, and St. Abramowski. Visible surface calculation for complex unstructured polygonal scenes. Computing, 35(3–4): 231–246, September 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Armbruster:1985:PP

[1520] D. Armbruster and R. Oechsle. Pseudorecursive procedures. Computing, 35 (3–4):247–256, September 1985. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chu:1985:SCT

[1521] M. T. Chu. Symbolic calculation of the trace of the power of a tridiagonal matrix. Computing, 35(3–4): 257–268, September 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Afflerbach:1985:CMR

[1522] L. Afflerbach and H. Grothe. Calculation of Minkowski-reduced lattice bases. Computing, 35(3–4):269–276, September 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ponisch:1985:CSB

[1523] Gerd Pönisch. Computing simple bifurcation points using a minimally extended system of nonlinear equations. Computing, 35(3–4):277–294, September 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Grossmann:1985:SDO

[1524] C. Grossmann and A. A. Kaplan.

On the solution of discretized obstacle problems by an adapted penalty method. Computing, 35(3–4):295–306, September 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

AbouEl-Seoud:1985:AMA

[1525] M. S. Abou El-Seoud. An adaptive multigrid algorithm for the numerical solution of quasilinear potential equations. Computing, 35(3–4): 307–324, September 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Strehmel:1985:NSP

[1526] K. Strehmel and R. Weiner. Nichtlineare Stabilität und Phasenuntersuchung adaptiver Nyström-Runge– Kutta-Methoden. (German) [Nonlinear stability and phase analysis for adaptive Nyström-Runge–Kutta methods]. Computing, 35(3–4):325–344, September 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Linz:1985:ACC

[1527] P. Linz. On the approximate computation of certain strongly singular integrals. Computing, 35(3–4):345–353, September 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schwandt:1985:AKL

[1528] H. Schwandt. Accelerating Krawczyklike interval algorithms for the solution of nonlinear systems of equations by using second derivatives. Computing, 35(3–4):355–367, September 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Shampine:1985:BEC

[1529] L. F. Shampine. Better extrapolation codes. Computing, 35(3–4): 369–374, September 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hake:1985:RFM

[1530] J.-Fr. Hake. A remark on Frank matrices. Computing, 35(3–4):375–379, September 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chazelle:1986:CPP

[1531] B. M. Chazelle and D. T. Lee. On a circle placement problem. Computing, 36(1–2):1–16, March 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Burrage:1986:SCR

[1532] K. Burrage, W. H. Hundsdorfer, and J. G. Verwer. A study of B-convergence of Runge–Kutta methods. Computing, 36(1–2):17–34, March 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dekker:1986:OCG

[1533] K. Dekker, J. F. B. M. Kraaijevanger, and M. N. Spijker. The order of B-convergence of the Gaussian Runge–Kutta method. Computing, 36 (1–2):35–41, March 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Scholz:1986:SGE

[1534] S. Scholz. On Stetter’s global error estimation in the smooth phase of stiff differential equations. Computing, 36 (1–2):43–55, March 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Roos:1986:SOM

[1535] H.-G. Roos. A second order monotone upwind scheme. Computing, 36 (1–2):57–67, March 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Vulanovic:1986:ESP

[1536] R. Vulanović, D. Herceg, and N. Petrović.

On the extrapolation for a singularly perturbed boundary value problem. Computing, 36(1–2):69–79, March 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Potra:1986:MCN

[1537] F. A. Potra and W. C. Rheinboldt. On the monotone convergence of Newton’s method. Computing, 36(1–2):81–90, March 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pinter:1986:EUA

[1538] J. Pintér. Extended univariate algorithms for n-dimensional global optimization. Computing, 36(1–2):91–103, March 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [1714, 1845].

Zielke:1986:RTM

[1539] G. Zielke. Report on test matrices for generalized inverses. Computing, 36 (1–2):105–162, March 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Drygalla:1986:BLK

[1540] V. Drygalla. Zur Berechnung von Lösungen konsistenter linearer Gleichungssysteme mit singulärer Koeffizientenmatrix. (German) [On computing solutions of consistent linear algebraic systems with a singular matrix]. Computing, 36(1–2):163–168, March 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Krawczyk:1986:LOF

[1541] R. Krawczyk. A Lipschitz operator for function strips. Computing, 36 (1–2):169–174, March 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Herzberger:1986:OSRa

[1542] J. Herzberger. On the R-order of some recurrences with applications to inclusion-methods. Computing, 36(1– 2):175–180, March 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brass:1986:HWG

[1543] Dietrich Bräss. Helmut Werner. (German). Computing, 36(1–2):181–182, March 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Widmayer:1986:SUI

[1544] P. Widmayer, Y. F. Wu, M. D. F. Schlag, and C. K. Wong. On some union and intersection problems for polygons with fixed orientations. Computing, 36(3):183–197, September 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kaiserswerth:1986:VRR

[1545] M. Kaiserswerth. Verbesserung der Rundung bei der rationalen Arithmetik durch die Einführung von Nebennäherungsbrüchen. (German) [Improvement to rounding in rational arithmetic by the introduction of secondary convergents]. Computing, 36 (3):199–204, September 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Arbenz:1986:SNE

[1546] P. Arbenz and W. Gander. Solving nonlinear eigenvalue problems by algorithmic differentiation. Computing, 36 (3):205–215, September 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Heidemann:1986:LAE

[1547] U. Heidemann. Linearer Ausgleich mit Exponentialsplines bei automatischer Bestimmung der Intervallteilungspunkte. (German) [Linear least squares solution with exponential splines with automatic determination of the interval partitioning]. Computing, 36(3):217–227, September 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lee:1986:CSB

[1548] E. T. Y. Lee. Comments on some Bspline algorithms. Computing, 36(3): 229–238, September 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [1354].

Ioakimidis:1986:SDZ

[1549] N. I. Ioakimidis and E. G. Anastasselou. On the simultaneous determination of zeros of analytic or sectionally analytic functions. Computing, 36 (3):239–247, September 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Petkovic:1986:SOM

[1550] Miodrag S. Petković and Lidija V. Stefanović. On a second order method for the simultaneous inclusion of polynomial complex zeros in rectangular arithmetic. Computing, 36(3):249–261, September 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Derigs:1986:UOF

[1551] U. Derigs and A. Metz. On the use of optimal fractional matchings for solving the (integer) matching problem. Computing, 36(3):263–270, September 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Akl:1986:ACO

[1552] S. G. Akl. An adaptive and costoptimal parallel algorithm for minimum spanning trees. Computing, 36 (3):271–277, September 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schwarz:1986:AAD

[1553] F. Schwarz. Addendum to: “Automatically determining symmetries of partial differential equations”. Computing, 36 (3):279–280, September 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [1480].

Gunzenhauser:1986:PWK

[1554] Rul Gunzenhäuser and Hans J. Stetter.

Professor W. Knödel: 60 years. Computing, 36(4):283–284, December 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tinhofer:1986:GIT

[1555] G. Tinhofer. Graph isomorphism and theorems of Birkhoff type. Computing, 36(4):285–300, December 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Derigs:1986:ELT

[1556] U. Derigs and A. Metz. An efficient labeling technique for solving sparse assignment problems. Computing, 36 (4):301–311, December 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Csirik:1986:PAU

[1557] J. Csirik and G. Galambos. An O(n)

bin-packing algorithm for uniformly distributed data. Computing, 36(4): 313–319, December 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Alefeld:1986:IIC

[1558] G. Alefeld and H. Spreuer. Iterative improvement of componentwise error bounds for invariant subspaces belonging to a double or nearly double eigenvalue. Computing, 36(4):321–334, December 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hess:1986:CPI

[1559] W. Hess and J. W. Schmidt. Convexity preserving interpolation with exponential splines. Computing, 36(4): 335–342, December 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kirchgraber:1986:EME

[1560] U. Kirchgraber and G. Pospiech. An extrapolation method for the efficient composition of maps with applications to nonlinear oscillations. Computing, 36(4):343–354, December 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Luksan:1986:CVM

[1561] L. Lukšan. A compact variable metric algorithm for nonlinear minimax approximation. Computing, 36(4): 355–373, December 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Santoro:1986:IAB

[1562] N. Santoro and J. Urrutia. An improved algorithm for Boolean matrix multiplication. Computing, 36(4): 375–382, December 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wimp:1986:ABD

[1563] J. Wimp, R. Kline, and A. Galardi. Algorithms based on difference equations of infinite order and the computation of Laplace-type integrals. Computing, 37(1):1–18, March 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brewster:1986:CPC

[1564] M. E. Brewster and R. Kannan. A computational process for choosing the relaxation parameter in nonlinear SOR. Computing, 37(1):19–29, March 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Monegato:1986:QRP

[1565] G. Monegato and V. Pennacchietti.

Quadrature rules for Prandtl’s integral equation. Computing, 37(1):31–42, March 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Reimer:1986:RTA

[1566] M. Reimer and B. Sündermann. A Remez-type algorithm for the calculation of extremal fundamental systems for polynomial spaces on the sphere. Computing, 37(1):43–58, March 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Martinez:1986:SNS

[1567] José Mario Martínez. Solution of nonlinear systems of equations by an optimal projection method. Computing, 37(1):59–70, March 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Devroye:1986:GMS

[1568] L. Devroye. Grid methods in simulation and random variate generation. Computing, 37(1):71–84, March 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bernstein:1986:PIB

[1569] H. J. Bernstein and M. Goldstein. Parallel implementation of bisection for the calculation of eigenvalues of tridiagonal symmetric matrices. Computing, 37(1):85–91, March 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Er:1986:PER

[1570] M. C. Er. Performance evaluations of recursive and iterative algorithms for the Towers of Hanoi problem. Computing, 37(2):93–102, June 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stummel:1986:SOP

[1571] F. Stummel. Strict optimal a posteriori error and residual bounds for Gaussian elimination in floating-point arithmetic. Computing, 37(2):103–124, June 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Reichel:1986:MPM

[1572] Lothar Reichel. A method for preconditioning matrices arising from linear integral equations for elliptic boundary value problems. Computing, 37 (2):125–136, June 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ishihara:1986:EMI

[1573] K. Ishihara. Explicit monotone iterations providing upper and lower bounds for finite element solution with nonlinear radiation boundary conditions. Computing, 37(2):137–149, June 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ishutkin:1986:GWT

[1574] V. S. Ishutkin and K. Schönefeld. On the globalization of Wilson-type optimization methods by means of generalized reduced gradient methods. Computing, 37(2):151–169, June 1986. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See note [1575].

Schonefeld:1986:NGW

[1575] K. Schönefeld. A note on a globalization of Wilson-type optimization methods. Computing, 37(2):171–178, June 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [1574].

Krawczyk:1986:CIN

[1576] R. Krawczyk. A class of interval-Newton-operators. Computing, 37 (2):179–183, June 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Moret:1986:BAN

[1577] I. Moret. On the behaviour of approximate Newton methods. Computing, 37 (3):185–193, September 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

vanderHouwen:1986:NIS

[1578] P. J. van der Houwen, B. P. Sommeijer, K. Strehmel, and R. Weiner. On the numerical integration of secondorder initial value problems with a periodic forcing function. Computing, 37 (3):195–218, September 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kern:1986:EAS

[1579] W. Kern. An efficient algorithm for solving a special class of LPs. Computing, 37(3):219–226, September 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Krawczyk:1986:PIO

[1580] R. Krawczyk. Properties of interval operators. Computing, 37(3):227–245, September 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rokne:1986:LCD

[1581] J. G. Rokne. Low complexity kdimensional centered forms. Computing, 37(3):247–253, September 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Herzberger:1986:OSRb

[1582] J. Herzberger. On the R-order of some recurrences with applications to inclusion-methods. II. Computing, 37 (3):255–259, September 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hu:1986:SOU

[1583] A. J. Hu. Selection of the optimum uniform partition search. Computing, 37 (3):261–264, September 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Valiaho:1986:NLE

[1584] H. Väliaho. A note on locating eigenvalues. Computing, 37(3):265–267, September 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [1507].

Ioakimidis:1986:ENQ

[1585] N. I. Ioakimidis and E. G. Anastasselou. An elementary noniterative quadrature-type method for the numerical solution of a nonlinear equation. Computing, 37(3):269–275, September 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ecker:1986:CCS

[1586] A. Ecker and G. Poch. Check character systems. Computing, 37(4): 277–301, December 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Novacek:1986:UTS

[1587] A. Novacek. Using time stamps for storing and addressing extendible arrays. Computing, 37(4):303–313, December 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Enright:1986:ECR

[1588] W. H. Enright and P. H. Muir. Efficient classes of Runge–Kutta methods for two-point boundary value problems. Computing, 37(4):315–334, December 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

terMaten:1986:SMF

[1589] E. J. W. ter Maten. Splitting methods for fourth order parabolic partial differential equations. Computing, 37 (4):335–350, December 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lu:1986:SME

[1590] Tong Xing Lu. Solution of the matrix equation AX − XB = C. Computing, 37(4):351–355, December 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Acharya:1986:ADI

[1591] B. P. Acharya and T. Mohapatra. Approximation of double integrals of analytic functions of two complex variables. Computing, 37(4):357–364, December 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Das:1986:APM

[1592] S. R. Das and S. Y. Lee. Appraisal of the performance of the mmsc subgraph generation algorithm on a Cyber System 170/720. Computing, 37 (4):365–370, December 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Spath:1986:CLL

[1593] H. Späth. Clusterwise linear least absolute deviations regression. Computing, 37(4):371–378, December 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Telley:1987:RPS

[1594] H. Telley, Th. M. Liebling, and A. Mocellin. Reconstruction of polycrystalline structures: a new application of combinatorial optimization. Computing, 38(1):1–11, March 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Abel:1987:CPC

[1595] P. Abel. On the choice of the pivot columns of the simplex-method: Gradient criteria. Computing, 38(1):13–21, March 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Delvos:1987:BFP

[1596] F.-J. Delvos. Bernoulli functions and periodic B-splines. Computing, 38(1):23–31, March 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schneid:1987:NAO

[1597] J. Schneid. A new approach to optimal B-convergence. Computing, 38(1):33– 42, March 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Demmel:1987:TMR

[1598] J. W. Demmel. Three methods for refining estimates of invariant subspaces. Computing, 38(1):43–57, March 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Faigle:1987:IOO

[1599] U. Faigle and R. Schrader. Interval orders without odd crowns are defect optimal. Computing, 38(1):59–69, March 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Herzberger:1987:MIVa

[1600] J. Herzberger. On the monotonicity of the interval versions of Schulz’s method. Computing, 38(1):71–74, March 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wang:1987:SMP

[1601] De Ren Wang and Yu Jiang Wu. Some modifications of the parallel Halley iteration method and their convergence. Computing, 38(1):75–87, March 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Weihs:1987:BTM

[1602] C. Weihs, G. Calzolari, and L. Panattoni. Behavior of trust-region methods in FIML-estimation. Computing, 38(2):89–100, June 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ploss:1987:SMS

[1603] H. Ploss. On simulation methods for solving the Boltzmann equation. Computing, 38(2):101–115, June 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ishihara:1987:SOM

[1604] Kazuo Ishihara. Successive overrelaxation method with projection for finite element solutions of nonlinear radiation cooling problems. Computing, 38(2):117–132, June 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Martinez:1987:QNM

[1605] J. M. Martínez. Quasi-Newton methods with factorization scaling for solving sparse nonlinear systems of equations. Computing, 38(2):133–141, June 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schwandt:1987:IMS

[1606] H. Schwandt. Iterative methods for systems of equations with interval coefficients and linear form. Computing, 38(2):143–161, June 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Seifert:1987:CEA

[1607] P. Seifert. Computational experiments with algorithms for stiff ODEs. Computing, 38(2):163–176, June 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Traas:1987:SAD

[1608] C. R. Traas. Smooth approximation of data on the sphere with splines. Computing, 38(2):177–184, June 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kharab:1987:FBV

[1609] A. Kharab and R. B. Guenther. A free boundary value problem for water invading an unsaturated medium. Computing, 38(3):185–207, September 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dobner:1987:BSH

[1610] H.-J. Dobner. Bounds for the solution of hyperbolic problems. Computing, 38 (3):209–218, September 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

deHoog:1987:ASM

[1611] F. R. de Hoog and R. M. M. Mattheij.

An algorithm for solving multipoint boundary value problems. Computing, 38(3):219–234, September 1987.

CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Calio:1987:DIA

[1612] Franca Caliò and Elena Marchetti. Derivation and implementation of an algorithm for singular integrals. Computing, 38(3):235–245, September 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Marchy:1987:MQA

[1613] H. P. Märchy. On a modification of the QZ algorithm with fast Givens rotations. Computing, 38(3):247–259, September 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmidt:1987:PIR

[1614] J. W. Schmidt and W. Heß. Positive interpolation with rational quadratic splines. Computing, 38(3):261–267, September 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bohmer:1987:GSH

[1615] K. Böhmer and T. Römer. Grid strategy and high accuracy via defect corrections for the Kreiss-method for stiff boundary value problems. Computing, 38(3):269–273, September 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Shen:1987:IVS

[1616] Z. Shen and Y. Zhu. An interval version of Shubert’s iterative method for the localization of the global maximum. Computing, 38(3):275–280, September 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fine:1987:LOP

[1617] J. M. Fine. Low order practical Runge– Kutta–Nyström methods. Computing, 38(4):281–297, December 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Davis:1987:ILT

[1618] M. Davis and J. Dowden. Interpolation by a local taut cubic piecewise polynomial. Computing, 38(4): 299–313, December 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Scheidt:1987:DFP

[1619] J. K. Scheidt and C. W. Schelin. Distributions of floating point numbers. Computing, 38(4):315–324, December 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Jonker:1987:SAP

[1620] R. Jonker and A. Volgenant. A shortest augmenting path algorithm for dense and sparse linear assignment problems. Computing, 38(4):325–340, December 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brucker:1987:ASP

[1621] P. Brucker and W. Meyer. Approximation of a set of points by points of a grid. Computing, 38(4):341–345, December 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zinterhof:1987:GLP

[1622] P. Zinterhof. Gratis lattice points for multidimensional integration. Computing, 38(4):347–353, December 1987.

CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

ElTarazi:1987:QSA

[1623] M. N. El Tarazi and S. Sallam. On quartic splines with application to quadratures. Computing, 38(4): 355–361, December 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hristova:1987:NAM

[1624] S. G. Hristova and D. D. Baĭnov. Numeric-analytic method for finding the periodic solutions of nonlinear differential-difference equations with impulses. Computing, 38(4):363–368, December 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Akritas:1987:SPV

[1625] A. G. Akritas. Simple proof of the validity of the reduced prs algorithm. Computing, 38(4):369–372, December 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ponisch:1987:CHP

[1626] G. Pönisch. Computing hysteresis points of nonlinear equations depending on two parameters. Computing, 39(1):1–17, March 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Koltracht:1987:DTH

[1627] I. Koltracht and P. Lancaster. A definiteness test for Hankel matrices and their lower submatrices. Computing, 39(1):19–26, March 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fine:1987:IRK

[1628] J. M. Fine. Interpolants for Runge– Kutta–Nyström methods. Computing, 39(1):27–42, March 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Neubauer:1987:NRO

[1629] A. Neubauer. Numerical realization of an optimal discrepancy principle for Tikhonov-regularization in Hilbert scales. Computing, 39(1):43–55, March 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wobst:1987:GEP

[1630] R. Wobst. The generalized eigenvalue problem and acoustic surface wave computations. Computing, 39(1):57– 69, March 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bilgen:1987:WRE

[1631] Semih Bilgen and Z. Şeydao Deligönül.

Weibull renewal equation solution. Computing, 39(1):71–76, March 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Einfeldt:1987:SAL

[1632] B. Einfeldt. Ein schneller Algorithmus zur Lösung des Riemann-Problems. (German) [A quick algorithm for the solution of the Riemann problem]. Computing, 39(1):77–86, March 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Devroye:1987:SGD

[1633] L. Devroye. A simple generator for discrete log-concave distributions. Computing, 39(1):87–91, March 1987. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bleher:1987:FSF

[1634] J. H. Bleher, S. M. Rump, U. Kulisch, M. Metzger, Ch. Ullrich, and W. Walter. FORTRAN-SC. a study of a Fortran extension for engineering scientific computation with access to ACRITH. Computing, 39(2):93–110, June 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lengauer:1987:ESH

[1635] Th. Lengauer and Ch. Wieners. Efficient solutions of hierarchical systems of linear equations. Computing, 39(2):111–132, June 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmidt:1987:UDP

[1636] J. W. Schmidt. An unconstrained dual program for computing convex C^{1}-spline approximants. Computing, 39(2):133–140, June 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stefanescu:1987:CGR

[1637] S. Ştefănescu and I. Văduva. On computer generation of random vectors by transformations of uniformly distributed vectors. Computing, 39 (2):141–153, June 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mullen:1987:OCP

[1638] G. L. Mullen and H. Niederreiter. Optimal characteristic polynomials for digital multistep pseudorandom numbers.

Computing, 39(2):155–163, June 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Frenk:1987:OSL

[1639] J. B. G. Frenk, M. van Houweninge, and A. H. G. Rinnooy Kan. Order statistics and the linear assignment problem. Computing, 39(2):165–174, June 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schneid:1987:MOO

[1640] J. Schneid. On the maximum order of optimal B-convergence of Lobatto III C schemes. Computing, 39(2):175–181, June 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Herzberger:1987:RIV

[1641] J. Herzberger. Remarks on the interval version of Schulz’s method. Computing, 39(2):183–186, June 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Robert:1987:OSA

[1642] Y. Robert and D. Trystram. An orthogonal systolic array for the algebraic path problem. Computing, 39 (3):187–199, September 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Frenk:1987:HNF

[1643] J. B. G. Frenk and G. Galambos. Hybrid next-fit algorithm for the two-dimensional rectangle bin-packing problem. Computing, 39(3):201–217, September 1987. CODEN CMPTA2.

ISSN 0010-485X (print), 1436-5057 (electronic).

Akyildiz:1987:MVA

[1644] I. F. Akyildiz. Mean value analysis of closed queueing networks with Erlang service time distributions. Computing, 39(3):219–232, September 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Potra:1987:MNL

[1645] F. A. Potra. On a monotone Newtonlike method. Computing, 39(3): 233–246, September 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rokne:1987:ITF

[1646] J. G. Rokne and P. Bao. Interval Taylor forms. Computing, 39(3): 247–259, September 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Krawczyk:1987:CII

[1647] R. Krawczyk. Conditionally isotone interval operators. Computing, 39(3): 261–270, September 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Allasia:1987:NCT

[1648] G. Allasia and R. Besenghi. Numerical computation of Tricomi’s psi function by the trapezoidal rule. Computing, 39 (3):271–279, September 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pardalos:1987:GMI

[1649] P. M. Pardalos, J. H. Glick, and J. B. Rosen. Global minimization of indefinite quadratic problems. Computing, 39(4):281–291, December 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kiwiel:1987:SSM

[1650] K. C. Kiwiel. A subgradient selection method for minimizing convex functions subject to linear constraints. Computing, 39(4):293–305, December 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Martinez:1987:ASS

[1651] J. M. Martínez. An algorithm for solving sparse nonlinear least squares problems. Computing, 39(4):307–325, December 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Grotschel:1987:CPA

[1652] M. Grötschel and O. Holland. A cutting plane algorithm for minimum perfect 2-matchings. Computing, 39(4): 327–344, December 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hertz:1987:UTS

[1653] A. Hertz and D. de Werra. Using tabu search techniques for graph coloring. Computing, 39(4):345–351, December 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Richert:1987:IRK

[1654] W. R. Richert. Implicit Runge–Kutta formulae with built-in estimates of the accumulated truncation error. Computing, 39(4):353–362, December 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Alefeld:1987:KIN

[1655] G. Alefeld. Über die Konvergenzordnung des Intervall-Newton-Verfahrens. (German) [On the order of convergence of the interval-Newton-method]. Computing, 39(4):363–369, December 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Herzberger:1987:MIVb

[1656] J. Herzberger. On the monotonicity of the interval versions of Schulz’s method. II. Computing, 39(4):371–375, December 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Drexl:1988:SAA

[1657] A. Drexl. A simulated annealing approach to the multiconstraint zero-one knapsack problem. Computing, 40(1): 1–8, March 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Meijer:1988:BSA

[1658] H. Meijer and S. G. Akl. Bit serial addition trees and their applications. Computing, 40(1):9–17, March 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hoy:1988:RBN

[1659] Annegret Hoy. A relation between Newton and Gauss–Newton steps for singular nonlinear equations. Computing, 40(1):19–27, March 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Li:1988:NKC

[1660] Z. C. Li and T. D. Bui. A new kind of combinations between the Ritz– Galerkin and finite element methods for singularity problems. Computing, 40(1):29–50, March 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ishihara:1988:SOM

[1661] K. Ishihara. Successive overrelaxation method with projection with finite element solutions applied to the Dirichlet problem of the nonlinear elliptic equation Δu = bu^{2}. Computing, 40(1):51–65, March 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chiodo:1988:CRM

[1662] Claudia Chiodo and Giuliana Criscuolo.

On the convergence of a rule by Monegato for the numerical evaluation of Cauchy principal value integrals. Computing, 40(1):67–74, March 1988. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Meyer:1988:FPM

[1663] P. W. Meyer. A fixed point multiple shooting method. Computing, 40(1):75–83, March 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Shen:1988:NMI

[1664] Zu He Shen and A. Neumaier. A note on Moore’s interval test for zeros of nonlinear systems. Computing, 40(1):85–90, March 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

King:1988:VFD

[1665] J. T. King and A. Neubauer. A variant of finite-dimensional Tikhonov regularization with a posteriori parameter choice. Computing, 40(2):91–109, June 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Jiang:1988:FEM

[1666] S. Jiang. On a finite element method for the equations of one-dimensional nonlinear thermoviscoelasticity. Computing, 40(2):111–124, June 1988. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brunner:1988:ASI

[1667] H. Brunner. The approximate solution of initial-value problems for general Volterra integro-differential equations. Computing, 40(2):125–137, June 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Cools:1988:ASF

[1668] R. Cools and Ann Haegemans. Another step forward in searching for cubature formulae with a minimal number of knots for the square. Computing, 40(2):139–146, June 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Katajainen:1988:RAC

[1669] J. Katajainen. The region approach to computing relative neighbourhood graphs in the L_{p} metric. Computing, 40(2):147–161, June 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Volk:1988:ERE

[1670] W. Volk. An efficient raster evaluation method for univariate polynomials. Computing, 40(2):163–173, June 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Weidner:1988:DKM

[1671] P. Weidner. The Durand–Kerner method for trigonometric and exponential polynomials. Computing, 40 (2):175–179, June 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Spedicato:1988:SBC

[1672] E. Spedicato and W. Burmeister. A strict bound to the condition number of bordered positive definite matrices. Computing, 40(2):181–183, June 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hansen:1988:CSV

[1673] Per Christian Hansen. Computation of the singular value expansion. Computing, 40(3):185–199, September 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Muhlbach:1988:MIG

[1674] G. Mühlbach. On multivariate interpolation by generalized polynomials on subsets of grids. Computing, 40(3): 201–215, September 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Charron:1988:ARA

[1675] R. J. Charron. Adapting rational approximants for Fourier series to data representation problems. Computing, 40(3):217–228, September 1988. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gatermann:1988:CSC

[1676] Karin Gatermann. The construction of symmetric cubature formulas for the square and the triangle. Computing, 40 (3):229–240, September 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Strehmel:1988:SCL

[1677] K. Strehmel, R. Weiner, and I. Dannehl. A study of B-convergence of linearly implicit Runge–Kutta methods. Computing, 40(3):241–253, September 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Weiner:1988:TIC

[1678] R. Weiner and K. Strehmel. A type insensitive code for delay differential equations basing on adaptive and explicit Runge–Kutta interpolation methods. Computing, 40(3): 255–265, September 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kaur:1988:PMT

[1679] D. Kaur and Y. Wallach. Parallel methods for tridiagonal equations. Computing, 40(3):267–272, September 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Frommer:1988:EBZ

[1680] Andreas Frommer and Werner Straub. Error-bounds for zeroes of polynomials using complex circular arithmetic. Computing, 40(3):273–280, September 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mohr:1988:PSP

[1681] Th. Mohr and C. Pasche. A parallel shortest path algorithm. Computing, 40 (4):281–292, December 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Spath:1988:FSP

[1682] H. Späth and J. Meier. Flexible smoothing with periodic cubic splines and fitting with closed curves. Computing, 40(4):293–300, December 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Apostolatos:1988:APG

[1683] N. Apostolatos and V. Vlassopoulos.

Abschätzung von Produkten gitterpositiver Matrizen und optimale Vorhersage stochastischer Quellen. (German) [Approximation of products of gridpositive matrices and optimal prediction of stochastic sources]. Computing, 40(4):301–313, December 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Crisci:1988:SOS

[1684] Maria Rosaria Crisci, Elvira Russo, and Antonia Vecchio. On the stability of the one-step exact collocation method for the second kind Volterra integral equation with degenerate kernel. Computing, 40(4):315–328, December 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Peyerimhoff:1988:RCT

[1685] A. Peyerimhoff, E. Stickel, and E. Wirsing. On the rate of convergence for two-term recursions. Computing, 40(4):329–335, December 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Verlinden:1988:CCF

[1686] P. Verlinden, R. Cools, D. Roose, and Ann Haegemans. The construction of cubature formulae for a family of integrals: a bifurcation problem. Computing, 40(4):337–346, December 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lasser:1988:HCF

[1687] R. H. Lasser and J. Tritschler. How to compute the n-fold convolution power of elementary functions without numerical integration. Computing, 40 (4):347–351, December 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brucker:1988:EAJ

[1688] P. Brucker. An efficient algorithm for the job-shop problem with two jobs. Computing, 40(4):353–359, December 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gwinner:1988:ASL

[1689] J. Gwinner. Acceptable solutions of linear complementarity problems. Computing, 40(4):361–366, December 1988. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Herzberger:1988:MHO

[1690] J. Herzberger. On the monotonicity of higher-order interval iterations for bounding the inverse matrix. Computing, 40(4):367–372, December 1988.

CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Nasraoui:1989:PCB

[1691] A. Nasraoui. Parallel coding of binary images. Computing, 41(1–2):1–17, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Umeo:1989:SAC

[1692] H. Umeo and T. Asano. Systolic algorithms for computational geometry problems—a survey. Computing, 41 (1–2):19–40, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lecot:1989:DSM

[1693] C. Lécot. A direct simulation Monte Carlo scheme and uniformly distributed sequences for solving the Boltzmann equation. Computing, 41 (1–2):41–57, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Al-Zanaidi:1989:MID

[1694] M. Al-Zanaidi and Ch. Grossmann.

Monotone iteration discretization algorithm for BVPs. Computing, 41 (1–2):59–74, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Csendes:1989:IMB

[1695] T. Csendes. An interval method for bounding level sets of parameter estimation problems. Computing, 41 (1–2):75–86, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zhen:1989:SIM

[1696] Mei Zhen. Splitting iteration method for simple singular points and simple bifurcation points. Computing, 41 (1–2):87–96, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See remark [1805].

Vulanovic:1989:UNM

[1697] R. Vulanović. A uniform numerical method for quasilinear singular perturbation problems without turning points. Computing, 41(1–2):97–106, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hoy:1989:EIG

[1698] Annegret Hoy. An efficiently implementable Gauss–Newton-like method for solving singular nonlinear equations. Computing, 41(1–2):107–122, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See remark [1805].

Frommer:1989:IBM

[1699] A. Frommer. Implementing Brown’s method for systems of nonlinear equations with dense banded Jacobian. Computing, 41(1–2):123–129, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kiehl:1989:ISS

[1700] M. Kiehl and Chr. Zenger. An improved starting step of the G-B-S-method for the solution of ordinary differential equations. Computing, 41 (1–2):131–136, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rohn:1989:TSM

[1701] J. Rohn. A two-sequence method for linear interval equations. Computing, 41(1–2):137–140, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Petkovic:1989:SAC

[1702] L. D. Petković and M. S. Petković.

On some applications of circular complex functions. Computing, 41(1– 2):141–148, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zuhe:1989:CCK

[1703] Shen Zuhe. A class of componentwise Krawczyk-Moore type iteration methods. Computing, 41(1–2):149–152, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Herzberger:1989:CIM

[1704] J. Herzberger. On the convergence of an iterative method for bounding the inverses of an interval matrix. Computing, 41(1–2):153–162, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ahrens:1989:HAL

[1705] J. H. Ahrens. How to avoid logarithms in comparisons with uniform random variables. Computing, 41(1– 2):163–166, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rohn:1989:NCN

[1706] J. Rohn. New condition numbers for matrices and linear systems. Computing, 41(1–2):167–169, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Evans:1989:NAG

[1707] D. J. Evans and C. Li. Numerical aspects of the generalized CG-method applied to least squares problems. Computing, 41(1–2):171–178, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sleijpen:1989:SSR

[1708] Gerard L. G. Sleijpen. Strong stability results for the hopscotch method with applications to bending beam equations. Computing, 41(3):179–203, September 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Liu:1989:NSO

[1709] K. M. Liu and E. L. Ortiz. Numerical solution of ordinary and partial functional-differential eigenvalue problems with the Tau method. Computing, 41(3):205–217, September 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Scholz:1989:OBC

[1710] S. Scholz. Order barriers for the Bconvergence of ROW methods. Computing, 41(3):219–235, September 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See corrigendum [1813].

Mackens:1989:NDI

[1711] W. Mackens. Numerical differentiation of implicitly defined space curves. Computing, 41(3):237–260, September 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fischer:1989:FMC

[1712] Herbert Fischer. Fast method to compute the scalar product of gradient and given vector. Computing, 41(3): 261–265, September 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zinn:1989:NMT

[1713] Axel Zinn. A numerical method for transmission problems for the Helmholtz equation. Computing, 41 (3):267–274, September 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zilinskas:1989:NEU

[1714] A. Žilinskas. A note on: “Extended univariate algorithms for n-dimensional global optimization” [Computing 36 (1986), no. 1–2, 91– 103; MR 87j:90113 ] by J. Pintér. Computing, 41(3):275–276, September 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See note [1538, 1845].

Hackbusch:1989:FSO

[1715] W. Hackbusch. On first and second order box schemes. Computing, 41 (4):277–296, December 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pohle:1989:ADD

[1716] Hans-Jürgen Pohle and Thomas Wolf.

Automatic determination of dynamical symmetries of ordinary differential equations. Computing, 41(4): 297–306, December 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bohmer:1989:RCB

[1717] K. Böhmer and Zhen Mei. Regularization and computation of a bifurcation problem with corank 2. Computing, 41 (4):307–316, December 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ying:1989:SRS

[1718] Xingren Ying and I. Norman Katz. A simple reliable solver for all the roots of a nonlinear function in a given domain. Computing, 41(4):317–333, December 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hoffmann:1989:IAG

[1719] Walter Hoffmann. Iterative algorithms for Gram–Schmidt orthogonalization. Computing, 41(4):335–348, December 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dontchev:1989:EED

[1720] A. L. Dontchev and E. M. Farkhi. Error estimates for discretized differential inclusions. Computing, 41(4): 349–358, December 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Cvetkovic:1989:AMS

[1721] Ljiljana Cvetković and Dragoslav Herceg. The AOR method for solving linear interval equations. Computing, 41(4):359–364, December 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eichhorn:1989:IAT

[1722] N. Eichhorn and H. Kiesewetter. IN-TEGER analysis of two-dimensional polygonal objects. Computing, 42(1): 1–16, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bartneck:1989:GDS

[1723] N. Bartneck. A general data structure for image analysis based on a description of connected components. Computing, 42(1):17–34, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Friden:1989:STF

[1724] C. Friden, A. Hertz, and D. de Werra.

STABULUS: a technique for finding stable sets in large graphs with tabu search. Computing, 42(1):35–44, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mroz:1989:DDM

[1725] Monika Mróz. Domain decomposition method for elliptic mixed boundary value problems. Computing, 42(1):45– 59, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schneid:1989:CCR

[1726] J. Schneid. Characterization of Bconvergent Runge–Kutta methods for strictly dissipative initial value problems. Computing, 42(1):61–67, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Alefeld:1989:NCI

[1727] G. Alefeld and F. Potra. A new class of interval methods with higher order of convergence. Computing, 42(1):69–80, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Roos:1989:CRT

[1728] H.-G. Roos and E. Pfeifer. A convergence result for the tau method. Computing, 42(1):81–84, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ostrovsky:1989:LCO

[1729] G. M. Ostrovsky, M. G. Ostrovsky, and T. A. Berezinsky. Linearly constrained optimization. Computing, 42(1):85–89, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dreschler-Fischer:1989:KSI

[1730] L. S. Dreschler-Fischer and H. Faasch. A kernel system for iconic image processing. Computing, 42(2–3):91–108, June 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Shermer:1989:HPP

[1731] T. Shermer. Hiding people in polygons. Computing, 42(2–3):109–131, June 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hinz:1989:IAT

[1732] A. M. Hinz. An iterative algorithm for the Tower of Hanoi with four pegs. Computing, 42(2–3):133–140, June 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wildenauer:1989:AMS

[1733] P. Wildenauer and L. Eisner. Adaptive model for a software test to calculate a residual error forecast. Computing, 42(2–3):141–158, June 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ahrens:1989:AMS

[1734] J. H. Ahrens and U. Dieter. An alias method for sampling from the normal distribution. Computing, 42 (2–3):159–170, June 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Frommer:1989:SBS

[1735] A. Frommer and G. Mayer. Safe bounds for the solutions of nonlinear problems using a parallel multisplitting method. Computing, 42(2–3):171–186, June 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Walz:1989:IGP

[1736] G. Walz. On interpolation by generalized planar splines. I. The polynomial case. Computing, 42(2–3):187–194, June 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Spedicato:1989:SLL

[1737] E. Spedicato and E. Bodon. Solving linear least squares by orthogonal factorization and pseudoinverse computation via the modified Huang algorithm in the ABS class. Computing, 42(2–3):195–205, June 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Micke:1989:TIE

[1738] A. Micke. The treatment of integral equations with discontinuous kernels using product type quadrature formulas. Computing, 42(2–3):207–223, June 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rump:1989:GIC

[1739] S. M. Rump. Guaranteed inclusions for the complex generalized eigenproblem. Computing, 42(2–3):225–238, June 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Marti:1989:NUL

[1740] J. T. Marti. New upper and lower bounds for the eigenvalues of the Sturm–Liouville problem. Computing, 42(2–3):239–243, June 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Voller:1989:ESW

[1741] R. L. Voller. Enclosure of solutions of weakly nonlinear elliptic boundary value problems and their computation. Computing, 42(2–3):245–258, June 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kioustelidis:1989:EBAa

[1742] J. B. Kioustelidis. An error bound for approximate solutions of two-point boundary value problems. Computing, 42(2–3):259–270, June 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Horst:1989:MIC

[1743] R. Horst and N. V. Thoai. Modification, implementation and comparison of three algorithms for globally solving linearly constrained concave minimization problems. Computing, 42 (2–3):271–289, June 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See erratum [2176].

Egecioglu:1989:PHI

[1744] Ö. Eğecioğlu, E. Gallopoulos, and Ç. K. Koç. Parallel Hermite interpolation: an algebraic approach. Computing, 42(4):291–307, December 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bein:1989:ASS

[1745] W. Bein and P. Brucker. An O(n log n)- algorithm for solving a special class of linear programs. Computing, 42 (4):309–313, December 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pflug:1989:SDP

[1746] G. Ch. Pflug. Sampling derivatives of probabilities. Computing, 42(4): 315–328, December 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Liu:1989:NSR

[1747] M. Z. Liu. A note on the stability of rational Runge–Kutta methods. Computing, 42(4):329–339, December 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Enright:1989:ATP

[1748] W. H. Enright and W. L. Seward. Achieving tolerance proportionality in software for stiff initial value problems. Computing, 42(4):341–352, December 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kenney:1989:EEE

[1749] Ch. Kenney, P. Linz, and R. L. C. Wang. Effective error estimates for the numerical solution of Fredholm integral equations. Computing, 42(4): 353–362, December 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Okecha:1989:INE

[1750] G. E. Okecha. Improving the numerical evaluation of certain oscillatory integrals. Computing, 42(4):363–371, December 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Akyildiz:1989:DSN

[1751] I. F. Akyildiz and Horst von Brand. Dual and selfdual networks of queues with rejection blocking. Computing, 43(1):1–11, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Moret:1989:CPB

[1752] I. Moret. A continuation procedure based on projected Newton steps. Computing, 43(1):13–25, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Janovsky:1989:NCS

[1753] Vladimír Janovský. A note on computing simple bifurcation points. Computing, 43(1):27–36, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Linde:1989:NBE

[1754] U. Linde, M. Reimer, and B. Sündermann.

Numerische Berechnung extremaler Fundamentalsysteme für Polynomräume über der Vollkugel. (German) [Calculation of extremal fundamental systems for polynomial spaces over the ball]. Computing, 43(1):37–45, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tautenhahn:1989:FIM

[1755] U. Tautenhahn. A fast iterative method for solving regularized parameter identification problems in elliptic boundary value problem. Computing, 43(1):47–58, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bochev:1989:SVN

[1756] Pavel Bochev and Svetoslav Markov.

A self-validating numerical method for the matrix exponential. Computing, 43(1):59–72, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Thiel:1989:GIN

[1757] Siegfried Thiel. A generalisation of the interval Newton single-step method for nonlinear systems of equations. Computing, 43(1):73–84, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Shen:1989:IMM

[1758] Zu He Shen and M. C. Eiermann. Interval-majorant method and global optimization. Computing, 43(1):85–92, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rohn:1989:FTT

[1759] J. Rohn. A Farkas-type theorem for linear interval equations. Computing, 43(1):93–95, March 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gekeler:1989:NSD

[1760] Eckart W. Gekeler. On the numerical solution of double-periodic elliptic eigenvalue problems. Computing, 43(2):97–114, June 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Auzinger:1989:ESI

[1761] W. Auzinger. On the error structure of the implicit Euler scheme applied to stiff systems of differential equations. Computing, 43(2):115–131, June 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kioustelidis:1989:EBAb

[1762] J. B. Kioustelidis. L_{2}-error bounds for approximate solutions of elliptic partial differential equations with Dirichletboundary conditions. Computing, 43 (2):133–140, June 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Cools:1989:MCF

[1763] Ronald Cools and Hans Joachim Schmid. Minimal cubature formulae of degree 2k −1 for two classical functionals. Computing, 43(2):141–157, June 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sidi:1989:CSN

[1764] Avram Sidi. Comparison of some numerical quadrature formulas for weakly singular periodic Fredholm integral equations. Computing, 43(2):159–170, June 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Clenshaw:1989:RSU

[1765] C. W. Clenshaw and P. R. Turner.

Root squaring using level-index arithmetic. Computing, 43(2):171–185, June 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wang:1989:CAP

[1766] De Ren Wang and Feng Guang Zhao.

Complexity analysis of a process for simultaneously obtaining all zeros of polynomials. Computing, 43(2):187– 197, June 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Geyer:1990:DPM

[1767] T. Geyer and J. Wick. A deterministic particle-method solving the linearized Boltzmann equation. Computing, 43 (3):199–207, September 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Claus:1990:SIR

[1768] H. Claus. Singly-implicit Runge–Kutta methods for retarded and ordinary differential equations. Computing, 43(3): 209–222, September 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Auzinger:1990:AEE

[1769] W. Auzinger, R. Frank, and G. Kirlinger. Asymptotic error expansions for stiff equations: Applications. Computing, 43(3):223–253, September 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tsitouras:1990:RKI

[1770] Ch. Tsitouras and G. Papageorgiou. Runge–Kutta interpolants based on values from two successive integration steps. Computing, 43(3):255–266, September 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dagnino:1990:EOD

[1771] Catterina Dagnino and Elisabetta Santi. On the evaluation of onedimensional Cauchy principal value integrals by rules based on cubic spline interpolation. Computing, 43 (3):267–276, September 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Vassilevski:1990:SWA

[1772] P. S. Vassilevski. On some ways of approximating inverses of banded matrices in connection with deriving preconditioners based on incomplete block factorizations. Computing, 43 (3):277–296, September 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eiselt:1990:CSM

[1773] H. A. Eiselt. A computational study of minimax flow centers in trees. Computing, 43(3):297–307, September 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pinter:1990:SNE

[1774] J. Pintér. Solving nonlinear equation systems via global partition and search: Some experimental results. Computing, 43(4):309–323, December 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lubich:1990:RMD

[1775] Ch. Lubich and M. Roche. Rosenbrock methods for differential-algebraic systems with solution-dependent singular matrix multiplying the derivative. Computing, 43(4):325–342, December 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Jankowski:1990:OSM

[1776] Tadeusz Jankowski. One-step methods for retarded differential equations with parameters. Computing, 43(4): 343–359, December 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hofmann:1990:RVI

[1777] B. Hofmann, R. Hausding, and R. Wolke. Regularization of a Volterra integral equation by linear inequalities. Computing, 43(4):361–375, December 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Al-Zanaidi:1990:DGG

[1778] Mansour Al-Zanaidi and Christian Grossmann. 1D-grid generation by monotone iteration discretization. Computing, 43(4):377–390, December 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Flajolet:1990:AS

[1779] P. Flajolet. On adaptive sampling. Computing, 43(4):391–400, December 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Szpankowski:1990:YAA

[1780] Wojciech Szpankowski and Vernon Rego. Yet another application of a binomial recurrence. Order statistics. Computing, 43(4):401–410, December 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Samuel:1990:CEG

[1781] D. Samuel and G. T. Toussaint. Computing the external geodesic diameter of a simple polygon. Computing, 44(1):1–19, March 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ruhe:1990:OBP

[1782] G. Ruhe and B. Fruhwirth. ϵoptimality for bicriteria programs and its application to minimum cost flows. Computing, 44(1):21–34, March 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Burgmeier:1990:ICM

[1783] P. Burgmeier, K. U. Jahn, and A. G. Plochov. An interval computational method for approximating controllability sets. Computing, 44(1):35–46, March 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ichida:1990:MOU

[1784] K. Ichida and Y. Fujii. Multicriterion optimization using interval analysis. Computing, 44(1):47–57, March 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ostermann:1990:CER

[1785] Alexander Ostermann. Continuous extensions of Rosenbrock-type methods. Computing, 44(1):59–68, March 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Berrut:1990:BFS

[1786] J. P. Berrut. Barycentric formulae for some optimal rational approximants involving Blaschke products. Computing, 44(1):69–82, March 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Martinez:1990:ASN

[1787] José Mario Martínez and Rita Filomena Santos. An algorithm for solving nonlinear least-squares problems with a new curvilinear search. Computing, 44(1):83–90, March 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gohlen:1990:PAE

[1788] M. Göhlen, M. Plum, and J. Schröder.

A programmed algorithm for existence proofs for two-point boundary value problems. Computing, 44(2):91–132, June 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dai:1990:CHM

[1789] J. Dai and H. Ren. Convergence of hybrid MUSCL-type schemes. Computing, 44(2):133–146, June 1990. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Jiang:1990:NSC

[1790] S. Jiang. Numerical solution for the Cauchy problem in nonlinear 1-D-thermoelasticity. Computing, 44 (2):147–158, June 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Popescu:1990:OPR

[1791] I. Popescu and I. Văduva. An optimum plan of reliability control. Computing, 44(2):159–168, June 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Candela:1990:RRRa

[1792] V. Candela and A. Marquina. Recurrence relations for rational cubic methods: I. The Halley method. Computing, 44(2):169–184, June 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Anonymous:1990:PHJ

[1793] Anonymous. Professor H. J. Stetter— 60 years. Computing, 44(3):185–186, September 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Abia:1990:SAF

[1794] L. Abia and J. M. Sanz-Serna. The spectral accuracy of a fully-discrete scheme for a nonlinear third order equation. Computing, 44(3):187–196, September 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Auzinger:1990:NCC

[1795] W. Auzinger, R. Frank, and G. Kirlinger. A note on convergence concepts for stiff problems. Computing, 44 (3):197–208, September 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Butcher:1990:OSS

[1796] J. C. Butcher. Order, stepsize and stiffness switching. Computing, 44(3): 209–220, September 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Frommer:1990:EME

[1797] A. Frommer and G. Mayer. Efficient methods for enclosing solutions of systems of nonlinear equations. Computing, 44(3):221–235, September 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Herzberger:1990:EIA

[1798] J. Herzberger and Lj. Petković. Efficient iterative algorithms for bounding the inverse of a matrix. Computing, 44 (3):237–244, September 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmidt:1990:SPH

[1799] J. W. Schmidt, W. Heß, and Th. Nordheim. Shape preserving histopolation using rational quadratic splines. Computing, 44(3):245–258, September 1990.

CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ypma:1990:SNE

[1800] T. J. Ypma and Y. Q. Shen. Solving N + m nonlinear equations with only m nonlinear variables. Computing, 44 (3):259–271, September 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Alefeld:1990:ARV

[1801] G. Alefeld. On the approximation of the range of values by interval expressions. Computing, 44(3):273–278, September 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hansen:1990:AMS

[1802] Pierre Hansen and Brigitte Jaumard. Algorithms for the maximum satisfiability problem. Computing, 44(4): 279–303, December 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bui:1990:AMG

[1803] T. D. Bui and V. N. Hanh. Automatic mesh generation for finite element analysis. Computing, 44(4): 305–329, December 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Auzinger:1990:ESE

[1804] W. Auzinger. On error structures and extrapolation for stiff systems, with application in the method of lines. Computing, 44(4):331–356, December 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hoy:1990:RCM

[1805] Annegret Hoy. A remark concerning Mei Zhen’s paper on singular nonlinear equations. Computing, 44(4):357–364, December 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [1696, 1698].

Schmidt:1990:CPR

[1806] J. W. Schmidt and M. Sakai. A criterion for the positivity of rational cubic C^{2}-spline interpolants. Computing, 44 (4):365–368, December 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

deHoog:1990:NP

[1807] F. R. de Hoog and R. M. M. Mattheij.

A note on polychotomy. Computing, 45(1):1–15, March 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Cash:1990:IIS

[1808] J. R. Cash and Margaret H. Wright.

Implementation issues in solving nonlinear equations for two-point boundary value problems. Computing, 45 (1):17–37, March 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Petkovic:1990:SLA

[1809] M. S. Petković. Schröder-like algorithms for multiple complex zeros of a polynomial. Computing, 45(1):39–50, March 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Burkard:1990:SPP

[1810] R. E. Burkard, G. Rote, E. Y. Yao, and Z. L. Yu. Shortest polygonal paths in space. Computing, 45(1):51–68, March 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Jahn:1990:EHD

[1811] K. U. Jahn. Evaluation of Hausdorff distances in interval mathematics. Computing, 45(1):69–77, March 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Czopik:1990:SDA

[1812] J. Czopik. The simultaneous determination of all zeros of a polynomial. Computing, 45(1):79–91, March 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Scholz:1990:COB

[1813] S. Scholz. Corrigendum to: “Order barriers for the B-convergence of ROW methods” [Computing 41 (1989), no. 3, 219–235; MR 90b:65146]. Computing, 45(1):93, March 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [1710].

Cappello:1990:SCI

[1814] P. R. Cappello, E. Gallopoulos, and Ç. K. Koç. Systolic computation of interpolating polynomials. Computing, 45(2):95–117, June 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lasser:1990:VCQ

[1815] D. Lasser. Visually continuous quartics and quintics. Computing, 45(2):119– 129, June 1990. CODEN CMPTA2.

ISSN 0010-485X (print), 1436-5057 (electronic).

Pardalos:1990:CAB

[1816] P. M. Pardalos and G. P. Rodgers.

Computational aspects of a branch and bound algorithm for quadratic zero-one programming. Computing, 45(2):131–144, June 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Verlinden:1990:CCF

[1817] P. Verlinden and A. Haegemans. The construction of cubature formulae by continuation. Computing, 45(2):145– 155, June 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mei:1990:SES

[1818] Z. Mei. A special extended system and a Newton-like method for simple singular nonlinear equations. Computing, 45(2):157–167, June 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Deng:1990:TNA

[1819] N. Deng and Z. Chen. Truncated nonlinear ABS algorithm and its convergence property. Computing, 45(2):169– 173, June 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Simos:1990:NTM

[1820] T. E. Simos and A. D. Raptis.

Numerov-type methods with minimal phase-lag for the numerical integration of the one-dimensional Schrödinger equation. Computing, 45(2):175–181, June 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bochev:1990:SSV

[1821] P. Bochev. Simultaneous selfverified computation of exp(A) and ∫_{0}^{1} exp(as) ds. Computing, 45(2):183– 191, June 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ewing:1990:FDS

[1822] R. E. Ewing, R. D. Lazarov, and P. S. Vassilevski. Finite difference schemes on grids with local refinement in time and space for parabolic problems. I. Derivation, stability, and error analysis. Computing, 45(3): 193–215, September 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Meyer:1990:PPC

[1823] A. Meyer. A parallel preconditioned conjugate gradient method using domain decomposition and inexact solvers on each subdomain. Computing, 45(3):217–234, September 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sartoretto:1990:NSO

[1824] F. Sartoretto and R. Spigler. Numerical solution for the one-phase Stefan problem by piecewise constant approximation of the interface. Computing, 45 (3):235–249, September 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Caussignac:1990:FEA

[1825] Ph. Caussignac, B. Zimmermann, and R. Ferro. Finite element approximation of electrostatic potential in one dimensional multilayer structures with quantized electronic charge. Computing, 45 (3):251–264, September 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Argyros:1990:MIP

[1826] I. K. Argyros. A mesh-independence principle for nonlinear operator equations and their discretizations under mild differentiability conditions. Computing, 45(3):265–268, September 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chiccoli:1990:AEI

[1827] C. Chiccoli, S. Lorenzutta, and G. Maino. An algorithm for exponential integrals of real order. Computing, 45(3):269–276, September 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Grzegorski:1990:SNM

[1828] S. M. Grzegórski. On the scaled Newton method for the symmetric eigenvalue problem. Computing, 45(3): 277–282, September 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mraz:1990:LMN

[1829] F. Mráz. Local minimizer of a nonconvex quadratic programming problem. Computing, 45(3):283–289, September 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Crisci:1990:GSA

[1830] M. R. Crisci, Z. Jackiewicz, E. Russo, and A. Vecchio. Global stability analysis of the Runge–Kutta methods for Volterra integral and integrodifferential equations with degenerate kernels. Computing, 45(4):291–300, December 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chan:1990:SRK

[1831] R. P. K. Chan. On symmetric Runge– Kutta methods of high order. Computing, 45(4):301–309, December 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Li:1990:CSM

[1832] Z. C. Li and T. D. Bui. Coupling strategy for matching different methods in solving singularity problems. Computing, 45(4):311–319, December 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Cooper:1990:SII

[1833] G. J. Cooper and R. Vignesvaran. A scheme for the implementation of implicit Runge–Kutta methods. Computing, 45(4):321–332, December 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fabbri:1990:STI

[1834] M. Fabbri. Stability of the totally implicit finite difference scheme for the one-dimensional diffusion equation with a moving boundary. Computing, 45(4):333–343, December 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gluchowska:1990:SLL

[1835] J. Głuchowska and A. Smoktunowicz. Solving the linear least squares problem with very high relative accuracy. Computing, 45(4):345–354, December 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Candela:1990:RRRb

[1836] V. Candela and A. Marquina. Recurrence relations for rational cubic methods. II. The Chebyshev method. Computing, 45(4):355–367, December 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brucker:1990:JSS

[1837] P. Brucker and R. Schlie. Jobshop scheduling with multi-purpose machines. Computing, 45(4):369–375, December 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Voller:1990:MIM

[1838] R. L. Voller. A monotone including multigrid-method. Computing, 45(4): 377–382, December 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Calve:1990:IRK

[1839] M. Calvé and R. Vaillancourt. Interpolants for Runge–Kutta pairs of order four and five. Computing, 45 (4):383–388, December 1990. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ascher:1991:PMB

[1840] U. M. Ascher and S. Y. P. Chan. On parallel methods for boundary value ODEs. Computing, 46(1):1–17, March 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Plum:1991:CAE

[1841] M. Plum. Computer-assisted existence proofs for two-point boundary value problems. Computing, 46(1):19–34, March 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schneider:1991:HIB

[1842] C. Schneider and W. Werner. Hermite interpolation: The barycentric approach. Computing, 46(1):35–51, March 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ugrin-Sparac:1991:SIP

[1843] G. Ugrin-Šparac. Stochastic investigations of pseudo-random number generators. Computing, 46(1):53–65, March 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Montuschi:1991:OAE

[1844] P. Montuschi and M. Mezzalama. Optimal absolute error starting values for Newton–Raphson calculation of square root. Computing, 46(1):67–86, March 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pinter:1991:GCR

[1845] J. Pintér. Global convergence revisited: reply to A. Žilinskas: “A note on: ‘Extended univariate algorithms for n-dimensional global optimization’ ” [Computing 41 (1989), no. 3, 275–276; MR 90b:90128]. Computing, 46(1):87–91, March 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [1538, 1714].

Bhattacharya:1991:CST

[1846] B. Bhattacharya and G. Toussaint.

Computing shortest transversals. Computing, 46(2):93–119, June 1991. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Su:1991:CRN

[1847] Tung-Hsin Su and Ruei Chuan Chang.

On constructing the relative neighborhood graphs in Euclidean kdimensional spaces. Computing, 46 (2):121–130, June 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sakurai:1991:IMA

[1848] T. Sakurai, T. Torii, and H. Sugiura. An iterative method for algebraic equation by Padé approximation. Computing, 46(2):131–141, June 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Huang:1991:MSN

[1849] Zhi Jian Huang. Multi-step nonlinear ABS methods and their efficiency analysis. Computing, 46(2):143–153, June 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lin:1991:NMQ

[1850] P. Lin. A numerical method for quasilinear singular perturbation problems with turning points. Computing, 46(2):155–164, June 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Devroye:1991:ETA

[1851] L. Devroye. Expected time analysis of a simple recursive Poisson random variate generator. Computing, 46(2):165–173, June 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fossmeier:1991:ALH

[1852] R. Fößmeier. Zur Axiomatisierung logarithmischer und halblogarithmischer Zahlensysteme. (German) [On the axiomatization of logarithmic and semilogarithmic number systems]. Computing, 46(2):175–182, June 1991. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kellerer:1991:BNS

[1853] H. Kellerer. Bounds for nonpreemptive scheduling of jobs with similar processing times on multiprocessor systems using the LPT-algorithm. Computing, 46 (3):183–191, September 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Louchard:1991:RVI

[1854] G. Louchard. Robust variations of interpolation search: an asymptotic analysis. Computing, 46(3):193–222, September 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Motta:1991:NNM

[1855] S. Motta and J. Wick. A new numerical method for kinetic equations in several dimensions. Computing, 46 (3):223–232, September 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Axelsson:1991:CWI

[1856] O. Axelsson and G. Lindskog. Constant wavefront iteration methods for nine- and 15-point difference matrices. Computing, 46(3):233–252, September 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hudak:1991:DIH

[1857] D. Hudak. On the determination of the interval hull of linear interval equations. Computing, 46(3):253–263, September 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Jansson:1991:ILS

[1858] C. Jansson. Interval linear systems with symmetric matrices, skewsymmetric matrices and dependencies in the right hand side. Computing, 46 (3):265–274, September 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Engl:1991:HW

[1859] Heinz W. Engl. Hansjörg Wacker (1939–1991). Computing, 46(4): 275–278, December 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Herzberger:1991:EAI

[1860] J. Herzberger. Efficient algorithms for the inclusion of the inverse matrix using error-bounds for hyperpower methods. Computing, 46(4):279–288, December 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Alefeld:1991:DIN

[1861] G. Alefeld. Über das Divergenzverhalten des Intervall-Newton-Verfahrens. (German) [On the divergence behaviour of the interval Newton method]. Computing, 46(4):289–294, December 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sherif:1991:CMI

[1862] N. Sherif. On the computation of a matrix inverse square root. Computing, 46 (4):295–305, December 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ponisch:1991:IMC

[1863] G. Pönisch. An indirect method for computing origins for Hopf bifurcation in two-parameter problems. Computing, 46(4):307–320, December 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Babel:1991:FMC

[1864] L. Babel. Finding maximum cliques in arbitrary and in special graphs. Computing, 46(4):321–341, December 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Muller:1991:VSN

[1865] R. Müller and D. Wagner. α-vertex separator is NP-hard even for 3-regular graphs. Computing, 46(4):343–353, December 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gasparo:1991:PCM

[1866] Maria G. Gasparo and R. Morandi. Piecewise cubic monotone interpolation with assigned slopes. Computing, 46(4):355–365, December 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Jansson:1991:GAC

[1867] C. Jansson. A geometric approach for computing A posteriori error bounds for the solution of a linear system. Computing, 47(1):1–9, March 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Behnke:1991:CGB

[1868] H. Behnke. The calculation of guaranteed bounds for eigenvalues using complementary variational principles. Computing, 47(1):11–27, March 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dinkel:1991:SII

[1869] John J. Dinkel, Marietta Tretter, and Danny Wong. Some implementation issues associated with multidimensional interval Newton methods. Computing, 47(1):29–42, March 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Junger:1991:CCD

[1870] M. Jünger, G. Reinelt, and D. Zepf.

Computing correct Delaunay triangulations. Computing, 47(1):43–49, March 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gutjahr:1991:URG

[1871] W. J. Gutjahr. Uniform random generation of expressions respecting algebraic identities. Computing, 47(1):51– 67, March 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mei:1991:PFA

[1872] Z. Mei. Path following around corank-2-bifurcation points of a semi-linear elliptic problem with symmetry. Computing, 47(1):69–85, March 1991. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmidt:1991:RBS

[1873] J. W. Schmidt. Rational biquadratic C^{1}-splines in S-convex interpolation. Computing, 47(1):87–96, March 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Frommer:1991:OCS

[1874] A. Frommer. Orders of convergence for superlinearly convergent chaotic iterations. Computing, 47(1):97–101, March 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lyness:1991:SPF

[1875] J. N. Lyness and T. Sørevik. A search program for finding optimal integration lattices. Computing, 47(2):103–120, June 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Georg:1991:AQV

[1876] K. Georg, R. Widmann, and Ralf Widmann. Adaptive quadratures over volumes. Computing, 47(2):121–136, June 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Haase:1991:ADDa

[1877] G. Haase, U. Langer, and A. Meyer. The approximate Dirichlet domain decomposition method. Part I: An algebraic approach. Computing, 47(2):137– 151, June 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Haase:1991:ADD

[1878] G. Haase, U. Langer, and A. Meyer.

The approximate Dirichlet domain decomposition method. Part II: Applications to 2nd-order elliptic B.V.P.s. Computing, 47(2):153–167, June 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kearfott:1991:DAE

[1879] R. Baker Kearfott. Decomposition of arithmetic expressions to improve the behavior of interval iteration for nonlinear systems. Computing, 47(2):169– 191, June 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Yalamov:1991:BSG

[1880] P. Y. Yalamov and Plamen Y.

Yalamov. On the backward stability of Gauss–Jordan elimination. Computing, 47(2):193–197, June 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Halang:1992:LAD

[1881] Wolfgang A. Halang. Load adaptive dynamic scheduling of tasks with hard deadlines useful for industrial applications. Computing, 47(3–4):199–213, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stoyenko:1992:OLT

[1882] A. D. Stoyenko and L. Georgiadis. On optimal lateness and tardiness scheduling in real-time systems. Computing, 47(3–4):215–234, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Muller:1992:IPM

[1883] H. Müller. Improvements on printing mice in labyrinths. Computing, 47(3– 4):235–246, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Prodinger:1992:EKP

[1884] H. Prodinger. Einige Kommentare zu einer bin-packing Aufgabe von W. Knödel. (German) [Some comments on a bin-packing problem of W. Knodel]. Computing, 47(3–4):247–254, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kalik:1992:ACM

[1885] K. Kalik and W. Wendland. The approximation of closed manifolds by triangulated manifolds and the triangulation of closed manifolds. Computing, 47(3–4):255–275, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Apel:1992:AIA

[1886] T. Apel and M. Dobrowolski. Anisotropic interpolation with applications to the finite element method. Computing, 47 (3–4):277–293, September 1992. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Meister:1992:AEG

[1887] M. Meister. An asymptotic expansion of the global discretization error of difference schemes for numerically solving a quasilinear parabolic system of differential equations. Computing, 47(3– 4):295–308, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Qin:1992:CHO

[1888] Meng Zhao Qin and Wen Jie Zhu.

Construction of higher order symplectic schemes by composition. Computing, 47(3–4):309–321, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hoenders:1992:DNM

[1889] B. J. Hoenders and C. H. Slump.

On the determination of the number and multiplicity of zeros of a function. Computing, 47(3–4):323–336, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rump:1992:SIL

[1890] S. M. Rump. On the solution of interval linear systems. Computing, 47(3– 4):337–353, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kjurkchiev:1992:GAH

[1891] N. Kjurkchiev and A. Andreev. On the generalization of the Alefeld– Herzberger’s method. Computing, 47 (3–4):355–360, September 1992. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gomes:1992:PIB

[1892] F. A. M. Gomes and J. M. Martínez.

Parallel implementations of Broyden’s method. Computing, 47(3–4):361–366, September 1992. CODEN CMPTA2.

ISSN 0010-485X (print), 1436-5057 (electronic).

Chocholaty:1992:NAN

[1893] P. Chocholatý. A numerical approach to 2k + e nonlinear equations with only k nonlinear variables. Computing, 47(3–4):367–372, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rohn:1992:REI

[1894] J. Rohn and A. Deif. On the range of eigenvalues of an interval matrix. Computing, 47(3–4):373–377, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Shampine:1992:UAS

[1895] L. F. Shampine and G. Kraut. Uniformly accurate Sturm–Liouville eigenvalues. Computing, 47(3–4):379–385, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Anonymous:1992:WMY

[1896] Anonymous. Willard L. Miranker, 60 years. Computing, 48(1):1–3, March 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Handler:1992:MWF

[1897] W. Handler. A multiprocessor working as a fault-tolerant cellular automaton. Computing, 48(1):5–20, March 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Shankar:1992:CEB

[1898] R. Shankar and D. P. Miranker. Credible execution of bounded-time parallel systems with delayed diagnosis. Computing, 48(1):21–37, March 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Miranker:1992:TIS

[1899] G. S. Miranker. Titan III supercomputer architectural overview. Computing, 48(1):39–60, March 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Jia:1992:URE

[1900] R.-Q. Jia and C. A. Micchelli. Using the refinement equation for the construction of pre-wavelets. V. Extensibility of trigonometric polynomials. Computing, 48(1):61–72, March 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Douglas:1992:ATD

[1901] C. C. Douglas and J. Mandel. An abstract theory for the domain reduction method. Computing, 48(1):73–96, March 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmidt:1992:CSH

[1902] J. W. Schmidt. Constrained smoothing of histograms by quadratic splines. Computing, 48(1):97–107, March 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bini:1992:PID

[1903] D. Bini and V. Pan. Practical improvement of the divide-and-conquer eigenvalue algorithms. Computing, 48 (1):109–123, March 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mayer:1992:SRT

[1904] G. Mayer. Some remarks on two interval-arithmetic modifications of the Newton method. Computing, 48 (1):125–128, March 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Willner:1992:WG

[1905] B. E. Willner, Chang Hung-Jen, and J. Foy. Willard graphs. Computing, 48(1):129–130, March 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Moret:1992:PMC

[1906] I. Moret. A projection method for computing turning points of nonlinear equations. Computing, 48(2):133–147, June 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Attili:1992:DMC

[1907] Basem S. Attili. A direct method for the characterization and computation of bifurcation points with corank 2. Computing, 48(2):149–159, June 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Suhartanto:1992:DLS

[1908] Heru Suhartanto and Wayne H. Enright. Detecting and locating a singular point in the numerical solution of IVPs for ODEs. Computing, 48(2):161–175, June 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stolle:1992:NIC

[1909] H. W. Stolle and R. Strauss. On the numerical integration of certain singular integrals. Computing, 48(2):177–189, June 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brasey:1992:HER

[1910] V. Brasey. A half-explicit Runge– Kutta method of order 5 for solving constrained mechanical systems. Computing, 48(2):191–201, June 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Surla:1992:ATD

[1911] D. Surla and M. Racković. The application of ψ-transform for determining a near-optimal path in the presence of polyhedral obstacles. Computing, 48(2):203–212, June 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Muthiyalu:1992:ECM

[1912] N. Muthiyalu and S. Usha. Eigenvalues of centrosymmetric matrices. Computing, 48(2):213–218, June 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Reid:1992:VFG

[1913] John Reid. The advantages of Fortran 90. Computing, 48(3):219–238, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lozier:1992:RPV

[1914] D. W. Lozier and P. R. Turner. Robuste Parallel-Verarbeitung in Gleitkommaund SLI-Arithmetik. (German) [Robust parallel computation in floatingpoint and SLI arithmetic]. Computing, 48(3–4):239–257, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bertocchi:1992:MCV

[1915] M. Bertocchi, L. Brandolini, L. Słomiński, and J. Sobczynska. Eine Monte-Carlo Verfahren zur Lösung von 0-1 Optimierungsproblemen. (German) [A Monte-Carlo approach for 0-1 programming problems]. Computing, 48(3–4): 259–274, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wang:1992:ENK

[1916] G. J. Wang. Erzeugung von NURBS-Kurven durch Einhullende. (German) [Generating NURBS curves by envelopes]. Computing, 48(3–4):275–289, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kalovics:1992:ZML

[1917] F. Kálovics. Eine zuverlassige Methode zur Lösung nichtlinearer Gleichungssysteme mit wenigen Unbekannten. (German) [A reliable method for solving nonlinear systems of equations of few variables]. Computing, 48(3– 4):291–302, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Xiao:1992:NKV

[1918] Y. Xiao and F. Zhou. Ein nichtmonotones Konfidenzbereichs-Verfahren mit krummlinigem Pfad zur unrestringierten Optimierung. (German) [Nonmonotone trust region methods with curvilinear path in unconstrained optimization]. Computing, 48 (3–4):303–317, September 1992. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Munack:1992:GOM

[1919] H. Munack. Über globale Optimierung mittels Intervallarithmetik. (German) [On global optimization using interval arithmetic]. Computing, 48(3–4): 319–336, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rote:1992:KSA

[1920] G. Rote. Die Konvergenzrate des Sandwich-Algorithmus zur Approximation konvexer Funktionen. (German) [The convergence rate of the sandwich algorithm for approximating convex functions]. Computing, 48(3–4): 337–361, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmidt:1992:PMK

[1921] J. W. Schmidt. Positive, monotone und S-konvexe C^{1}-Interpolation auf Rechteckgittern. (German) [Positive, monotone, and S-convex C^{1}- interpolation on rectangular grids]. Computing, 48(3–4):363–371, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Morandi:1992:BFI

[1922] R. Morandi. Blending-Funktionen mit inneren Bedingungen. (German) [A blending function method with interior constraints]. Computing, 48(3–4): 373–380, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fix:1992:AZD

[1923] G. J. Fix and R. Kannan. Ein Algorithmus mit Zufallsrichtungen für ein Durchschnittsproblem. (German) [A random direction algorithm for an intersection problem]. Computing, 48 (3–4):381–385, September 1992. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ding:1992:BCL

[1924] Y. Ding and M. A. Weiss. Best case lower bounds for Heapsort. Computing, 49(1):1–9, March 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Blum:1992:DSA

[1925] H. Blum, S. Lisky, and R. Rannacher. A domain splitting algorithm for parabolic problems. Computing, 49(1):11–23, March 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Plum:1992:NEP

[1926] Michael Plum. Numerical existence proofs and explicit bounds for solutions of nonlinear elliptic boundary value problems. Computing, 49(1):25–44, March 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kastner-Maresch:1992:IMR

[1927] Alois E. Kastner-Maresch. The implicit midpoint rule applied to discontinuous differential equations. Computing, 49(1):45–62, March 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Frommer:1992:TPM

[1928] A. Frommer and G. Mayer. On the theory and practice of multisplitting methods in parallel computation. Computing, 49(1):63–74, March 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Xu:1992:CPB

[1929] Zong Ben Xu and Xing Zhong Shi.

A comparison of point and ball iterations in the contractive mapping case. Computing, 49(1):75–85, March 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chen:1992:CSQ

[1930] X. Chen and T. Yamamoto. On the convergence of some quasi-Newton methods for nonlinear equations with nondifferentiable operators. Computing, 49(1):87–94, March 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schwarz:1992:ABG

[1931] Fritz Schwarz. Ein Algorithmus zur Bestimmung der Große von Symmetriegruppen. (German) [An algorithm for determining the size of symmetry groups]. Computing, 49(2):95–115, June 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Midy:1992:CRL

[1932] P. Midy. Computergestutzte Reduktion von Linearen Differentialgleichungen Zweiter Ordnung. (German) [Computer-assisted reduction of second order linear differential equations]. Computing, 49(2):117–127, June 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Riessinger:1992:IBQ

[1933] Th. Riessinger. Interpolation mit bivariaten quadratischen Splines auf Typ 2-Triangulationen. (German) [Interpolation by bivariate quadratic splines on a four-directional mesh]. Computing, 49(2):129–137, June 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bai:1992:MLK

[1934] Y. Bai. Eine modifizierte Lobatto-Kollokation für lineare Randwertprobleme von Algebro-Differentialgleichungen. (German) [A modified Lobatto collocation for linear boundary value problems of differential-algebraic equations]. Computing, 49(2):139–150, June 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Woginger:1992:PSS

[1935] G. J. Wöginger and Z. Yu. Preemptives Scheduling mit Set-up-Zeiten. (German) [A heuristic for preemptive scheduling with set-up times]. Computing, 49(2):151–158, June 1992. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Behnke:1992:SEP

[1936] H. Behnke. Schranken für Eigenwerte parameterabhangiger Matrizen. (German) [Bounds for eigenvalues of parameter-dependent matrices]. Computing, 49(2):159–167, June 1992. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mattheij:1992:MHW

[1937] R. M. M. Mattheij and H. J. Stetter. In memoriam H.-J. Wacker. Computing, 49(3):169, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kiehl:1992:EOT

[1938] M. Kiehl and O. von Stryk. Echtzeit-Optimierung eines Tagesspeicherkraftwerkes. (German) [Real-time optimization of a hydroelectric power plant]. Computing, 49(3):171–191, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Guddat:1992:EAM

[1939] J. Guddat, W. Römisch, and R. Schultz.

Einige Anwendungen der mathematischen Optimierung bei der optimalen Lastverteilung. (German) [Some applications of mathematical programming techniques in optimal power dispatch]. Computing, 49(3):193–200, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fulop:1992:VSI

[1940] J. Fülöp. Ein Verfahren der semiinfiniten Programmierung für die Approximation von Lastdauerlinien durch Polynome. (German) [A semi-infinite programming method for approximating load duration curves by polynomials]. Computing, 49(3):201–212, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fasano:1992:QAS

[1941] A. Fasano, M. Primicerio, and F. Rosso. Über quasistationare achsensymmetrische Stromungen vom Bingham-Typus mit Spannungsinduzierter Degradation. (German) [On quasi-steady axisymmetric flows of Bingham type with stress-induced degradation]. Computing, 49(3): 213–237, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

vandeVorst:1992:NAZ

[1942] G. A. L. van de Vorst and R. M. M. Mattheij. Numerische Analyse eines zweidimensionalen viskosen Sinterproblems mit nichtglatten Randern. (German) [Numerical analysis of a 2-D viscous sintering problem with nonsmooth boundaries]. Computing, 49 (3):239–263, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Binder:1992:NBA

[1943] A. Binder. Zur numerischen Behandlung des Ausbauchens der Strangschale von Brammen. (German) [On the numerical treatment of bulging of continuously cast slabs]. Computing, 49 (3):265–278, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kieser:1992:NAS

[1944] R. Kieser, C. Schwab, and W. L. Wendland. Zur numerischen Auswertung von singularen Integralen auf gekrummten Flachen mit Hilfe von Symbolmanipulation. (German) [Numerical evaluation of singular and finite-part integrals on curved surfaces using symbolic manipulation]. Computing, 49 (3):279–301, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Soderlind:1992:SAH

[1945] G. Söderlind. Zur Stabilität von Algebrodifferentialgleichungen von höherem Index bei Störungen von Systemparametern. (German) [Remarks on the stability of high-index DAEs with respect to parametric perturbations]. Computing, 49(3):303–314, September 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eiselt:1993:RRZ

[1946] H. A. Eiselt and C.-L. Sandblom. Eine Rechenstudie der Redundanz zufallig erzeugter Polytope. (German) [A computational study of redundancy in randomly generated polytopes]. Computing, 49(4):315–327, December 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Galambos:1993:BBS

[1947] G. Galambos and G. J. Wöginger.

Beim Bounded Space On-line Bin-Packing kann Umpacken von Vorteil sein. (German) [Repacking helps in bounded space on-line bin-packing]. Computing, 49(4):329–338, December 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Linzer:1993:ITS

[1948] E. Linzer and M. Vetterli. Iterative Toeplitz Solver mit lokal quadratischer Konvergenz. (German) [Iterative Toeplitz solvers with local quadratic convergence]. Computing, 49(4):

339–347, December 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Petkovic:1993:OEB

[1949] M. S. Petković, L. V. Stefanović, and Z. M. Marjanović. Über die R-Ordnung einiger beschleunigter Verfahren zur simultanen Polynomwur-zelbestimmung. (German) [On the R-order of some accelerated methods for the simultaneous finding of polynomial zeros]. Computing, 49(4):349–361, December 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

PalamaraOrsi:1993:ZAN

[1950] Annamaria Palamara Orsi. Zwei Algorithmen für die numerische Konstruktion von Produkt-Quadraturformeln. (German) [Two algorithms for the construction of product formulas]. Computing, 49(4):363–372, December 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rohn:1993:SUO

[1951] J. Rohn. Ein Schrittweitenalgorithmus für unrestringierte Optimierung. (German) [A step size rule for unconstrained optimization]. Computing, 49 (4):373–376, December 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zechner:1993:EBV

[1952] H. Zechner and E. Stadlober. Erzeugung von beta-verteilten Zufallszahlen mittels Patchwork-Verwerfung. (German) [Generating beta variates via patchwork rejection]. Computing, 50(1):1–18, March 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmidt:1993:PMC

[1953] Jochen W. Schmidt. Positive, monotone und S-convexe C^{1}-Histopolation auf Rechteckgittern. (German) [Positive, monotone, and S-convex C^{1}- histopolation on rectangular grids]. Computing, 50(1):19–30, March 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chan:1993:VSR

[1954] R. P. K. Chan. Verallgemeinerte symmetrische Runge–Kutta-Verfahren. (German) [Generalized symmetric Runge–Kutta methods]. Computing, 50(1):31–49, March 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Basler:1993:MEE

[1955] Chr. Bäsler and W. Törnig. Über monoton einschließ ende nichtlineare Mehrgitterverfahren und Anwendungen. (German) [On monotone including nonlinear multigrid methods and applications]. Computing, 50(1):51–67, March 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gao:1993:FEI

[1956] J. Gao. Ein C^{2}-finites Element und Interpolation. (German) [A C^{2} finite element and interpolation]. Computing, 50(1):69–76, March 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Balla:1993:BBE

[1957] K. Balla, O. S. Guk, and M. Vicsek. Zur Berechnung der Besselfunktion erster Art. (German) [On the computation of Bessel functions of first kind]. Computing, 50(1):77–85, March 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Simos:1993:GNN

[1958] T. E. Simos and A. B. Sideridis. Genaue numerische Naherungen für Anfangswertprobleme mit periodischen Lösungen. (German) [Accurate numerical approximations to initial value problems with periodical solutions]. Computing, 50(1):87–92, March 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Posch:1993:BKR

[1959] K. C. Posch and R. Posch. Basiserweiterung mit einer Konvolutionssumme in Restklassenzahlensystemen. (German) [Base extension using a convolution sum in residue number systems]. Computing, 50(2):93–104, June 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gellrich:1993:URM

[1960] C. Gellrich and B. Hofmann. Eine Untersuchung zur Regularisierung bei Monotonienebenbedingungen. (German) [A study of regularization by monotonicity]. Computing, 50(2):105– 125, June 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Griebel:1993:MGS

[1961] M. Griebel, C. Zenger, and S. Zimmer. Multilevel Gauß–Seidel-Algorithmen für Voll- und Dunngitterprobleme. (German) [Multilevel gauss–seidelalgorithms for full and sparse grid problems]. Computing, 50(2):127–148, June 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dagnino:1993:NIQ

[1962] C. Dagnino, V. Demichelis, and E. Santi. Numerische Integration mit quasi-interpolierenden Splines. (German) [Numerical integration based on quasi-interpolating splines]. Computing, 50(2):149–163, June 1993. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Calio:1993:KIG

[1963] F. Caliò and E. Marchetti. Komplexe Integrationsformeln von Gauss– Kronrod für gewisse Hauptwertintegrale von Cauchy. (German) [Complex Gauss–Kronrod integration rules for certain Cauchy principal value integrals]. Computing, 50(2):165–173, June 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tirani:1993:PIR

[1964] R. Tirani and T. Romanó. Parallele Interpolation für Runge–Kutta-Verfahren höher Ordnung. (German) [Parallel interpolation of high-order Runge–Kutta methods]. Computing, 50(2):175–184, June 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kunisch:1993:KGM

[1965] K. Kunisch. Über eine Klasse gedampfter Morozov’scher Prinzipien. (German) [On a class of damped Morozov principles]. Computing, 50(3):

185–198, September 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Reusken:1993:MMT

[1966] A. Reusken. Mehrgitterverfahren mit matrixabhangigen Transferoperatoren für ein singular gestortes Problem. (German) [Multigrid with matrixdependent transfer operators for a singular perturbation problem]. Computing, 50(3):199–211, September 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See note [2088].

Baddour:1993:SVR

[1967] N. Baddour and H. Brunner. Stetige Volterra–Runge–Kutta-Methoden für Integralgleichungen mit Verzogerung. (German) [Continuous Volterra–Runge– Kutta methods for integral equations with pure delay]. Computing, 50(3): 213–227, September 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Huang:1993:RZA

[1968] Z. Huang. Reinitialisierung bei Zeilenkorrektur-ABS- Verfahren zur Lösung nichtlinearer Gleichungssysteme. (German) [Restart row update ABS methods for solving systems of nonlinear equations]. Computing, 50 (3):229–239, September 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

VanThoai:1993:KDC

[1969] N. Van Thoai. Kanonische d. c. Optimierungstechniken für eine konvexe Optimierungsaufgabe mit einer multiplikativen Nebenbedingung. (German)

[Canonical d. c. programming techniques for solving a convex program with an additional constraint of multiplicative type]. Computing, 50(3): 241–253, September 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Jahn:1993:SIG

[1970] K.-U. Jahn. Schleifen-Invarianten in Gleitpunktalgorithmen. (German) [Loop invariants in floating point algorithms]. Computing, 50(3):255–264, September 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mao:1993:BFP

[1971] W. Mao. Best-k-fit bin packing. Computing, 50(3):265–270, September 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kellerer:1993:GHP

[1972] H. Kellerer and G. Wöginger. Eine Greedy-Heuristik zur Partitionierung in gleichartige Tripel. (German) [A greedy-heuristic for 3-Partitioning with similar elements]. Computing, 50(3): 271–278, September 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bookstein:1993:HCD

[1973] A. Bookstein and S. T. Klein. Is Huffman coding dead. Computing, 50(4): 279–296, December 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bunke:1993:AMR

[1974] H. Bunke and J. Csirik. An algorithm for matching run-length coded strings.

Computing, 50(4):297–314, December 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Walz:1993:IAS

[1975] Guido Walz. An iterative algorithm for spline interpolation. Computing, 50 (4):315–325, December 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Verlan:1993:EMI

[1976] I. I. Verlan. An explicit method of C^{2} interpolation using splines. Computing, 50(4):327–335, December 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Song:1993:CCW

[1977] Y. Song. Convergence and comparisons of waveform relaxation methods for initial value problems of linear ODE systems. Computing, 50(4): 337–352, December 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Shen:1993:CCT

[1978] Z. Shen and M. A. Wolfe. On certain computable tests and componentwise error bounds. Computing, 50 (4):353–368, December 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Heuser:1993:DJS

[1979] W. K. Heuser. Deterministic job-shop scheduling problems with capacity focus. Computing, 51(1):1–13, March 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kirschenhofer:1993:ROS

[1980] P. Kirschenhofer and H. Prodinger.

A result in order statistics related to probabilistic counting. Computing, 51(1):15–27, March 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Erhel:1993:EDP

[1981] J. Erhel. Experiments with data perturbations to study condition numbers and numerical stability. Computing, 51(1):29–44, March 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Scherzer:1993:UMD

[1982] O. Scherzer. The use of Morozov’s discrepancy principle for Tikhonov regularization for solving nonlinear ill-posed problems. Computing, 51(1):45–60, March 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Makroglou:1993:EBD

[1983] A. Makroglou. Extended backward differentiation formulae in the numerical solution of general Volterra integrodifferential equations. Computing, 51(1):61–77, March 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Yang:1993:CGM

[1984] H. Yang. Conjugate gradient methods for the Rayleigh quotient minimization of generalized eigenvalue problems. Computing, 51(1):79–94, March 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brasel:1993:CHA

[1985] H. Bräsel, T. Tautenhahn, and F. Werner. Constructive heuristic algorithms for the open shop problem. Computing, 51(2):95–110, June 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pehlivanov:1993:CAL

[1986] A. I. Pehlivanov, G. F. Carey, R. D. Lazarov, and Y. Shen. Convergence analysis of least-squares mixed finite elements. Computing, 51(2):111–123, June 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Oswald:1993:BPE

[1987] P. Oswald. On a BPX-preconditioner for P1 elements. Computing, 51 (2):125–133, June 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Cooper:1993:UPP

[1988] G. J. Cooper and R. Vignesvaran. On the use of parallel processors for implicit Runge–Kutta methods. Computing, 51(2):135–150, June 1993. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Papageorgiou:1993:RKP

[1989] G. Papageorgiou, Ch. Tsitouras, and S. N. Papakostas. Runge–Kutta pairs for periodic initial value problems. Computing, 51(2):151–163, June 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schneid:1993:PCD

[1990] J. Schneid and S. Pittner. On the parametrization of the coefficients of dilation equations for compactly supported wavelets. Computing, 51 (2):165–173, June 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eichenauer-Herrmann:1993:EIC

[1991] J. Eichenauer-Herrmann. Explicit inversive congruential pseudorandom numbers: the compound approach. Computing, 51(2):175–182, June 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Valiaho:1993:CCO

[1992] H. Väliaho. Cycling can occur in Mráz’s algorithm for nonconvex quadratic programming. Computing, 51(2):183–184, June 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Boglaev:1993:PIG

[1993] Yu. P. Boglaev. Die parallele Integration für gewohnliche Differentialgleichungen auf der Grundlage von konvolutiven Algorithmen. (German) [Parallel integration of ODEs based on convolution algorithms]. Computing, 51(3– 4):185–207, September 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chartier:1993:PST

[1994] P. Chartier and B. Philippe. Eine parallele “shooting” Technik zur Lösung dissipativer gewohnlicher Differentialgleichungen. (German) [A parallel shooting technique for solving dissipative ODE’s]. Computing, 51(3–4): 209–236, September 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bohmer:1993:NLS

[1995] K. Böhmer. Ein numerisches Liapunov–Schmidt-Verfahren für Operatorgleichungen. (German) [On a numerical Liapunov–Schmidt method for operator equations]. Computing, 51(3– 4):237–269, September 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmidt:1993:BMV

[1996] T. Schmidt. Box-Methoden auf Vierecksgittern. (German) [Box schemes on quadrilateral meshes]. Computing, 51(3–4):271–292, September 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Belsky:1993:MVK

[1997] V. Belsky. Eine Mehrgittermethode für Variationsungleichungen bei Kontaktproblemen. (German) [A multi-grid method for variational inequalities in contact problems]. Computing, 51(3– 4):293–311, September 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Carstensen:1993:IAE

[1998] C. Carstensen and M. S. Petković. Über Intervallmethoden für algebraische, exponentielle und trigonometrische Polynome. (German) [On some interval methods for algebraic, exponential and trigonometric polynomials]. Computing, 51(3–4):313–326, September 1993.

CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Korman:1993:ABI

[1999] P. Korman. Ein Algorithmus zur Berechnung instabiler Lösungen halblinearer Randwertprobleme. (German) [An algorithm for computing unstable solutions of semilinear boundary value problems]. Computing, 51(3– 4):327–334, September 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kosmol:1993:RMA

[2000] P. Kosmol and I. Nissen. Eine das Residuum minimierende Algorithmenklasse für lineare Gleichungen. (German) [A minimal residual class of algorithms for linear systems]. Computing, 51(3–4):335–343, September 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lin:1994:CTH

[2001] S. S. Lin. Constant-time Hough transform on the processor arrays with reconfigurable bus systems. Computing, 52(1):1–15, March 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Shakhlevich:1994:STJ

[2002] N. V. Shakhlevich and Yu. N. Sot·skov.

Scheduling two jobs with fixed and nonfixed routes. Computing, 52(1):17–30, March 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Babel:1994:FAM

[2003] L. Babel. A fast algorithm for the maximum weight clique problem. Computing, 52(1):31–38, March 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wick:1994:NAK

[2004] J. Wick. Numerical approaches to the kinetic semiconductor equation. Computing, 52(1):39–49, March 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Petrova:1994:EIC

[2005] S. I. Petrova. An efficient implementation of certain iterative refinement preconditioners. Computing, 52(1):51–63, March 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Freyburger:1994:RTA

[2006] K. Freyburger. A Remez type algorithm for generalized spline spaces. Computing, 52(1):65–87, March 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hormann:1994:UGD

[2007] W. Hörmann. A universal generator for discrete log-concave distributions. Computing, 52(1):89–96, March 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brucker:1994:TZG

[2008] P. Brucker and L. Nordmann. Das k-Track Zuordnungsproblem. (German) [The k-track assignment problem]. Computing, 52(2):97–122, June 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Poljak:1994:KKR

[2009] S. Poljak and F. Rendl. Knotenund Kanten-Relaxationen beim Max-Cut Problem. (German) [Node and edge relaxations of the max-cut problem]. Computing, 52(2):123–137, June 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hackbusch:1994:NQF

[2010] W. Hackbusch and S. A. Sauter.

Numerische Quadratur fast singularer Oberflachenintegrale aus der Randelement-kollokation. (German) [On numerical cubatures of nearly singular surface integrals arising in BEM collocation]. Computing, 52(2):139– 159, June 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Galligani:1994:FNA

[2011] E. Galligani and A. Laratta. Fehleranalyse des Nullraum-Algorithmus für Ausgleichsprobleme mit linearen Gleichungs-Nebenbedingungen. (German) [Error analysis of null space algorithm for linear equality constrained least squares problems]. Computing, 52(2):161–176, June 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

El-Daou:1994:RDK

[2012] M. K. El-Daou and E. L. Ortiz. Eine rekursive Darstellung von Kollokation mit Hilfe kanonischer Polynome. (German) [A recursive formulation of collocation in terms of canonical polynomials]. Computing, 52(2):177–202, June 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Yan:1994:SAL

[2013] W.-M. Yan and K.-L. Chung. Ein schneller Algorithmus zur Lösung spezieller tridiagonaler Systeme. (German) [A fast algorithm for solving special tridiagonal systems]. Computing, 52(2):203–211, June 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Luther:1994:VIE

[2014] W. J. Luther and W. Otten. Verified inclusion for eigenvalues of the first order PLL equation with general phase detector characteristics. Computing, 52 (3):213–232, September 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Seif:1994:ISE

[2015] N. P. Seif, S. A. Hussein, and A. S. Deif. The interval Sylvester equation. Computing, 52(3):233–244, September 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Oswald:1994:CRS

[2016] P. Oswald. On the convergence rate of SOR: a worst case estimate. Computing, 52(3):245–255, September 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Morandi:1994:DMI

[2017] R. Morandi and C. Conti. A discrete monotonous interpolation scheme. Computing, 52(3):257–267, September 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bhattacharya:1994:OAS

[2018] B. K. Bhattacharya, S. Jadhav, A. Mukhopadhyay, and J.-M. Robert. Optimal algorithms for some intersection radius problems. Computing, 52 (3):269–279, September 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Galambos:1994:LBD

[2019] G. Galambos and A. van Vliet. Lower bounds for 1-, 2- and 3-dimensional on-line bin packing algorithms. Computing, 52(3):281–297, September 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ariyawansa:1994:DCS

[2020] K. A. Ariyawansa. On Davidon’s collinear scaling algorithms for optimization. Computing, 52(3):299–307, September 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sakai:1994:NIF

[2021] M. Sakai and R. A. Usmani. Numerical integration formulas based on iterated cubic splines. Computing, 52 (3):309–314, September 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Neumaier:1994:GRR

[2022] A. Neumaier and Murray Hill. Global, rigorous and realistic bounds for the solution of dissipative differential equations. part 1: Theory. Computing, 52 (4):315–336, December 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Diethelm:1994:MCQ

[2023] K. Diethelm. Modified compound quadrature rules for strongly singular integrals. Computing, 52(4):337–354, December 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kobbelt:1994:FDP

[2024] L. Kobbelt. A fast dot-product algorithm with minimal rounding errors. Computing, 52(4):355–369, December 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Deineko:1994:GAA

[2025] V. Deĭneko, R. Rudolf, and G. J. Wöginger. A general approach to avoiding two by two submatrices. Computing, 52(4):371–388, December 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ganley:1994:HLI

[2026] J. L. Ganley and L. S. Heath. Heuristics for laying out information graphs. Computing, 52(4):389–405, December 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Nowinski:1994:CFM

[2027] W. L. Nowinski. A conflict free memory access in parallel image reconstruction. Computing, 53(1):1–11, March 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eckardt:1994:DCA

[2028] H. Eckardt. Disk cache architectures for transaction-like applications in parallel computers. Computing, 53(1):13–31, March 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lazarov:1994:FVM

[2029] R. D. Lazarov, I. D. Mishev, and P. S.

Vassilevski. Finite volume methods with local refinement for convectiondiffusion problems. Computing, 53 (1):33–57, March 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lirkov:1994:CBF

[2030] I. D. Lirkov, S. D. Margenov, and P. S. Vassilevski. Circulant blockfactorization preconditioners for elliptic problems. Computing, 53(1):59–74, March 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Butcher:1994:SOT

[2031] J. C. Butcher. Some orbital test problems. Computing, 53(1):75–94, March 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hlynsky:1994:EID

[2032] Ya. Hlynsky and Yu. Panchyshyn. Efficient implementation of a 2nd derivative method for stiff ODEs. Computing, 53(1):95–99, March 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kanzow:1994:UOT

[2033] C. Kanzow. An unconstrained optimization technique for large-scale linearly constrained convex minimization problems. Computing, 53(2):101–117, June 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zhou:1994:CNS

[2034] F. Zhou and Y. Xiao. A class of nonmonotone stabilization trust region methods. Computing, 53(2):119–136, June 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mulansky:1994:PSS

[2035] B. Mulansky and J. W. Schmidt. Powell–Sabin splines in range restricted interpolation of scattered data. Computing, 53(2):137–154, June 1994. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rieder:1994:RDC

[2036] A. Rieder and X. Zhou. On the robustness of the damped V -cycle of the wavelet frequency decomposition multigrid method. Computing, 53 (2):155–171, June 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schwab:1994:VOC

[2037] C. Schwab. Variable order composite quadrature of singular and nearly singular integrals. Computing, 53(2):173– 194, June 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hungerbuhler:1994:NAS

[2038] N. Hungerbühler. Numerical analysis of singular weighted integrals. Computing, 53(2):195–203, June 1994. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Benouamer:1994:HST

[2039] M. O. Benouamer, P. Jaillon, D. Michelucci, and J.-M. Moreau. Hash-Speicherung von ,,tragen” Zahlen. (German) [Hashing “lazy” numbers]. Computing, 53(3– 4):205–217, September 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). International Symposium on Scientific Computing, Computer Arithmetic and Validated Numerics (Vienna, 1993).

Merrheim:1994:CEF

[2040] X. Merrheim. The computation of elementary functions in radix 2^{p}. Computing, 53(3–4):219–232, September 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). International Symposium on Scientific Computing, Computer Arithmetic and Validated Numerics (Vienna, 1993).

Schulte:1994:OIA

[2041] M. J. Schulte, J. Omar, and E. E.

Swartzlander, Jr. Optimal initial approximations for the Newton–Raphson division algorithm. Computing, 53(3– 4):233–242, September 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). International Symposium on Scientific Computing, Computer Arithmetic and Validated Numerics (Vienna, 1993).

Reith:1994:CGP

[2042] R. Reith and C. P. Ullrich. Coarsegrain parallelizations of interval algorithms decomposing dense matrices and solving triangular systems on multicomputers. Computing, 53(3–4): 243–257, September 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). International Symposium on Scientific Computing, Computer Arithmetic and Validated Numerics (Vienna, 1993).

Bohlender:1994:IMM

[2043] G. Bohlender, T. Kersten, and R. Trier. Implicit matrix multiplication with maximum accuracy on various transputer networks. Computing, 53(3–4): 259–276, September 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). International Symposium on Scientific Computing, Computer Arithmetic and Validated Numerics (Vienna, 1993).

Knuppel:1994:PBF

[2044] O. Knüppel. PROFIL/BIAS — a fast interval library. Computing, 53(3–4): 277–287, September 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). International Symposium on Scientific Computing, Computer Arithmetic and Validated Numerics (Vienna, 1993).

Berlejung:1994:TEU

[2045] H. M. E. Berlejung. A test environment using symbolic program trace. Computing, 53(3–4):289–299, September 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Carpraux:1994:SPN

[2046] J. F. Carpraux, J. Erhel, and M. Sadkane. Spectral portrait for non-Hermitian large sparse matrices. Computing, 53(3–4):301–310, September 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). International Symposium on Scientific Computing, Computer Arithmetic and Validated Numerics (Vienna, 1993).

Alefeld:1994:NVI

[2047] G. Alefeld, A. Gienger, and G. Mayer.

Numerical validation for an inverse matrix eigenvalue problem. Computing, 53(3–4):311–322, September 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). International Symposium on Scientific Computing, Computer Arithmetic and Validated Numerics (Vienna, 1993).

Hong:1994:SSR

[2048] H. Hong and V. Stahl. Safe starting regions by fixed points and tightening. Computing, 53(3–4):323–335, September 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). International Symposium on Scientific Computing, Computer Arithmetic and Validated Numerics (Vienna, 1993).

Ratz:1994:BSS

[2049] D. Ratz. Box-splitting strategies for the interval Gauss–Seidel step in a global optimization method. Computing, 53(3–4):337–353, September 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). International Symposium on Scientific Computing, Computer Arithmetic and Validated Numerics (Vienna, 1993).

Schramm:1994:IPP

[2050] P. Schramm. Intersection problems of parametric surfaces in CAGD. Computing, 53(3–4):355–364, September 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). International Symposium on Scientific Computing, Computer Arithmetic and Validated Numerics (Vienna, 1993).

Rohn:1994:ESL

[2051] J. Rohn. Enclosing solutions of linear interval equations is NP-hard. Computing, 53(3–4):365–368, September 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). International Symposium on Scientific Computing, Computer Arithmetic and Validated Numerics (Vienna, 1993).

Rihm:1994:CEM

[2052] R. Rihm. On a class of enclosure methods for initial value problems. Computing, 53(3–4):369–377, September 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). International Symposium on Scientific Computing, Computer Arithmetic and Validated Numerics (Vienna, 1993).

Neher:1994:ESI

[2053] M. Neher. Enclosing solutions of an inverse Sturm–Liouville problem with finite data. Computing, 53(3– 4):379–395, September 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). International Symposium on Scientific Computing, Computer Arithmetic and Validated Numerics (Vienna, 1993).

Shi:1994:SRR

[2054] S. Shi and R. B. Kearfott. Some results on the regularity of an interval matrix. Computing, ??(??):??, 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Adam:1995:NFE

[2055] D. Adam, A. Felgenhauer, H.-G. Roos, and M. Stynes. A nonconforming finite element method for a singularly perturbed boundary value problem. Computing, 54(1):1–25, March 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Turek:1995:GMI

[2056] S. Turek. A generalized mean intensity approach for the numerical solution of the radiative transfer equation. Computing, 54(1):27–38, March 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=239&
id_volume=17&id_journal=8.

Gallivan:1995:LOR

[2057] K. Gallivan, P. C. Hansen, Tz. Ostromsky, and Z. Zlatev. A locally optimized reordering algorithm and its application to a parallel sparse linear system solver. Computing, 54(1):39–67, March 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=240&
id_volume=17&id_journal=8.

vanDamme:1995:CIC

[2058] R. van Damme and R.-H. Wang.

Curve interpolation with constrained length. Computing, 54(1):69–81, March 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=241&
id_volume=17&id_journal=8.

Sheng:1995:ICM

[2059] S. Sheng. Interpolation by conic model for unconstrained optimization. Computing, 54(1):83–98, March 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=242&
id_volume=17&id_journal=8.

Steele:1995:RBF

[2060] N. C. Steele, C. R. Reeves, M. Nicholas, and P. J. King. Radial basis function artificial neural networks for the inference process in fuzzy logic based control. Computing, 54(2):99–117, June 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=243&
id_volume=18&id_journal=8.

Chung:1995:FSM

[2061] K.-L. Chung. Fast string matching algorithms for run-length coded strings. Computing, 54(2):119–125, June 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=244&
id_volume=18&id_journal=8.

Ahrens:1995:OTM

[2062] J. H. Ahrens. A one-table method for sampling from continuous and discrete distributions. Computing, 54(2):127– 146, June 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springer.at/springer.py?Page=10&
Key=362&cat=300607/tocs/springer.
py?Page=47&Key=340&cat=3&id_abstract=
245&id_volume=18&id_journal=8.

Revers:1995:NIR

[2063] M. Revers. Numerical integration of the Radon transform on classes E_{s}^{α} in multiple (finite) dimensions. Computing, 54(2):147–165, June 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=246&
id_volume=18&id_journal=8.

Lipitakis:1995:EPI

[2064] E. A. Lipitakis and G. A. Gravvanis. Explicit preconditioned iterative methods for solving large unsymmetric finite element systems. Computing, 54(2):167–183, June 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=247&
id_volume=18&id_journal=8.

Higueras:1995:SAR

[2065] I. Higueras. On simplifying assumptions of Runge–Kutta methods for index 2 differential algebraic problems. Computing, 54(2):185–190, June 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=248&
id_volume=18&id_journal=8.

Burkard:1995:PSC

[2066] R. E. Burkard and V. G. Deĭneko. Polynomially solvable cases of the traveling salesman problem and a new exponential neighborhood. Computing, 54(3):191–211, September 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=249&
id_volume=19&id_journal=8.

Hochstattler:1995:HGF

[2067] W. Hochstättler and G. Tinhofer. Hamiltonicity in graphs with few P_{4}’s. Computing, 54(3):213–225, September 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=250&
id_volume=19&id_journal=8.

vonIssendorff:1995:TLC

[2068] H. von Issendorff. Time and logic: a calculus of binary events. Computing, 54(3):227–240, September 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.

at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=251&
id_volume=19&id_journal=8.

Trier:1995:AWU

[2069] U. Trier. Additive weights under the balanced probability model. Computing, 54(3):241–250, September 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=252&
id_volume=19&id_journal=8.

Zhu:1995:ISL

[2070] S. Q. Zhu. Iterations for solving linear variational inequalities on domains with nonlinear boundaries. Computing, 54(3):251–272, September 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=253&
id_volume=19&id_journal=8.

Godzina:1995:BMT

[2071] G. Godzina. Blending methods for two classical integrals. Computing, 54(3):273–282, September 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=254&
id_volume=19&id_journal=8.

Palubeckis:1995:HBB

[2072] G. Palubeckis. A heuristic-based branch and bound algorithm for unconstrained quadratic zero-one programming. Computing, 54(4):283–301, December 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=255&
                                        id_volume=20&id_journal=8.

Babel:1995:ICG

[2073] L. Babel. Isomorphism of chordal (6, 3) graphs. Computing, 54(4):303–316, December 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=256&
                                        id_volume=20&id_journal=8.

Vandewalle:1995:FMA

[2074] S. Vandewalle and G. Horton. Fourier mode analysis of the multigrid waveform relaxation and time-parallel multigrid methods. Computing, 54 (4):317–330, December 1995. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

                                        http://www.springer.at/springer.
                                        py?Page=10&Key=362&cat=300607/tocs/
                                        springer.py?Page=47&Key=340&cat=
                                        3&id_abstract=257&id_volume=20&
                                        id_journal=8.

Stevenson:1995:RAM

[2075] R. P. Stevenson. Robustness of the additive and multiplicative frequency decomposition multi-level method.

Computing, 54(4):331–346, December 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=258&
id_volume=20&id_journal=8.

Barbarosie:1995:RWE

[2076] C. Barbăroşie. Reducing the wrapping effect. Computing, 54(4):347–357, December 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=259&
id_volume=20&id_journal=8.

Locher:1995:ALA

[2077] F. Locher and M.-R. Skrzipek.

An algorithm for locating all zeros of a real polynomial. Computing, 54(4):359–375, December 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=260&
id_volume=20&id_journal=8.

Stoffer:1995:VSR

[2078] D. Stoffer. Variable steps for reversible integration methods. Computing, 55(1):1–22, March 1995. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

http://www.springer.at/springer.
py?Page=10&Key=362&cat=300607/tocs/
springer.py?Page=47&Key=340&cat=

                                        3&id_abstract=261&id_volume=21&
                                        id_journal=8.

Neumann:1995:MCR

[2079] L. Neumann. Monte Carlo radiosity. Computing, 55(1):23–42, March 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=262&
                                        id_volume=21&id_journal=8.

Hong:1995:BFI

[2080] H. Hong and V. Stahl. Bernstein form is inclusion monotone. Computing, 55(1):43–53, March 1995. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

                                        http://www.springer.at/springer.
                                        py?Page=10&Key=362&cat=300607/tocs/
                                        springer.py?Page=47&Key=340&cat=
                                        3&id_abstract=263&id_volume=21&
                                        id_journal=8.

Cacace:1995:SOR

[2081] F. Cacace, P. Samarati, and L. Tanca. Set-oriented retrieval of complex objects in inheritance hierarchies. Computing, 55(1):55–73, March 1995. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

                                        http://www.springer.at/springer.
                                        py?Page=10&Key=362&cat=300607/tocs/
                                        springer.py?Page=47&Key=340&cat=
                                        3&id_abstract=264&id_volume=21&
                                        id_journal=8.

Bunke:1995:FAM

[2082] H. Bunke. Fast approximate matching of words against a dictionary. Computing, 55(1):75–89, March 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=265&
id_volume=21&id_journal=8.

Andrew:1995:NSN

[2083] A. L. Andrew, K. E. Chu, and P. Lancaster. On the numerical solution of nonlinear eigenvalue problems. Computing, 55(2):91–111, June 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=266&
id_volume=22&id_journal=8.

Pereverzev:1995:OPM

[2084] S. V. Pereverzev. Optimization of projection methods for solving ill-posed problems. Computing, 55(2):113–124, June 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=267&
id_volume=22&id_journal=8.

Chen:1995:PQN

[2085] Z. Chen, P. Fei, and H. Zheng. A parallel quasi-Newton algorithm for unconstrained optimization. Computing, 55(2):125–133, June 1995. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

http://www.springer.at/springer.
py?Page=10&Key=362&cat=300607/tocs/
springer.py?Page=47&Key=340&cat=

                                        3&id_abstract=268&id_volume=22&
                                        id_journal=8.

Hamann:1995:AUG

[2086] B. Hamann, H. J. Thornburg, and G. Hong. Automatic unstructured grid generation based on iterative point insertion. Computing, 55(2):135–161, June 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=269&
                                        id_volume=22&id_journal=8.

Reddien:1995:SNM

[2087] G. W. Reddien. On the stability of numerical methods of Hopf points using backward error analysis. Computing, 55(2):163–180, June 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=270&
                                        id_volume=22&id_journal=8.

Hackbusch:1995:NRL

[2088] W. Hackbusch. A note on Reusken’s lemma. Computing, 55(2):181–189, June 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=271&
                                        id_volume=22&id_journal=8.  See

[1966].

dAmore:1995:OCH

[2089] F. d’Amore, V. H. Nguyen, T. Roos, and P. Widmayer. On optimal cuts of hyperrectangles. Computing, 55(3):191–206, September 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=274&
id_volume=23&id_journal=8.

Klar:1995:CNF

[2090] A. Klar. Computation of nonlinear functionals in particle methods. Computing, 55(3):207–221, September 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=273&
id_volume=23&id_journal=8.

Enright:1995:IRK

[2091] W. H. Enright and M. Hu. Interpolating Runge–Kutta methods for vanishing delay differential equations. Computing, 55(3):223–236, September 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=275&
id_volume=23&id_journal=8.

Bainov:1995:FDM

[2092] D. Baĭnov, E. Minchev, and Z. Kamont. The finite difference method for first order impulsive partial differential-functional equations. Computing, 55(3):237–253, September 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=272&
id_volume=23&id_journal=8.

Kohler:1995:ASE

[2093] P. Köhler. Asymptotically sharp error estimates for modified compound quadrature formulae for Cauchy principal value integrals. Computing, 55(3):255–269, September 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=276&
id_volume=23&id_journal=8.

Becker:1995:AEC

[2094] R. Becker, C. Johnson, and R. Rannacher. Adaptive error control for multigrid finite element methods. Computing, 55(4):271–288, December 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=277&
id_volume=24&id_journal=8.

Dorfler:1995:RAS

[2095] W. Dörfler. A robust adaptive strategy for the non-linear Poisson equation. Computing, 55(4):289–304, December 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=278&
id_volume=24&id_journal=8.

Angermann:1995:BPE

[2096] L. Angermann. Balanced a posteriori error estimates for finite-volume type discretizations of convectiondominated elliptic problems. Computing, 55(4):305–323, December 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=279&
id_volume=24&id_journal=8.

Leinen:1995:DSC

[2097] P. Leinen. Data structures and concepts for adaptive finite element methods. Computing, 55(4):325–354, December 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=280&
id_volume=24&id_journal=8.

Bey:1995:TGR

[2098] J. Bey. Tetrahedral grid refinement.

Computing, 55(4):355–378, December 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=281&
id_volume=24&id_journal=8.

Brass:1995:TAM

[2099] D. Bräss. Towards algebraic multigrid for elliptic problems of second order. Computing, 55(4):379–393, December 1995. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=282&
                                        id_volume=24&id_journal=8.

Coomer:1996:MPM

[2100] R. K. Coomer and I. G. Graham. Massively parallel methods for semiconductor device modelling. Computing, 56(1):1–27, March 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=283&
                                        id_volume=25&id_journal=8.

Bekesi:1996:FGA

[2101] J. Békési, G. Galambos, U. Pferschy, and G. J. Wöginger. The fractional greedy algorithm for data compression. Computing, 56(1):29–46, March 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=284&
                                        id_volume=25&id_journal=8.

Buttner:1996:NSI

[2102] M. Büttner, R. Weiner, and K. Strehmel. A note on stability investigations for Rosenbrock-type methods for quasilinear-implicit differential equations. Computing, 56(1):47–59, March 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=285&
id_volume=25&id_journal=8.

Hahne:1996:SIE

[2103] M. Hahne and E. P. Stephan. Schwarz iterations for the efficient solution of screen problems with boundary elements. Computing, 56(1):61–85, March 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=286&
id_volume=25&id_journal=8.

Usmani:1996:NIF

[2104] R. A. Usmani and M. Sakai. Numerical integration formulas based on iterated cubic splines. II. Computing, 56(1):87– 93, March 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springer.at/springer.py?Page=10&
Key=362&cat=300607/tocs/springer.
py?Page=47&Key=340&cat=3&id_abstract=
287&id_volume=25&id_journal=8.

Faigle:1996:ROA

[2105] U. Faigle, R. Garbe, and W. Kern.

Randomized online algorithms for maximizing busy time interval scheduling. Computing, 56(2):95–104, June 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.

                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=288&
                                        id_volume=26&id_journal=8.

Ugrin-Sparac:1996:PET

[2106] G. Ugrin-Šparac and D. Ugrin-Šparac. On a possible error of type II in statistical evaluation of pseudo-random number generators. Computing, 56(2):105–116, June 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=289&
                                        id_volume=26&id_journal=8.

Ferket:1996:FAL

[2107] P. J. J. Ferket and A. A. Reusken. Further analysis of the local defect correction method. Computing, 56(2):117– 139, June 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

                                        springer.at/springer.py?Page=10&
                                        Key=362&cat=300607/tocs/springer.
                                        py?Page=47&Key=340&cat=3&id_abstract=
                                        290&id_volume=26&id_journal=8.

Sleijpen:1996:RUR

[2108] G. L. G. Sleijpen and H. A. van der Vorst. Reliable updated residuals in hybrid Bi-CG methods. Computing, 56(2):141–163, June 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=291&
                                        id_volume=26&id_journal=8.

Zhang:1996:WCA

[2109] G. Zhang. Worst-case analysis of the FFH algorithm for online variable-sized bin packing. Computing, 56(2):165– 172, June 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springer.at/springer.py?Page=10&
Key=362&cat=300607/tocs/springer.
py?Page=47&Key=340&cat=3&id_abstract=
292&id_volume=26&id_journal=8.

Bodenstorfer:1996:HMM

[2110] B. Bodenstorfer. How many minimal upper bounds of minimal upper bounds. Computing, 56(2):173–178, June 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=293&
id_volume=26&id_journal=8.

Vanek:1996:AMS

[2111] P. Vaněk, J. Mandel, and M. Brezina. Algebraic multigrid by smoothed aggregation for second and fourth order elliptic problems. Computing, 56(3):179–196, September 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=297&
id_volume=27&id_journal=8. Inter-

national GAMM-Workshop on Multilevel Methods (Meisdorf, 1994).

Dornseifer:1996:DED

[2112] T. Dornseifer and C. Pflaum. Discretization of elliptic differential equations on curvilinear bounded domains with sparse grids. Computing, 56(3):197–213, September 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=294&
id_volume=27&id_journal=8. Inter-

national GAMM-Workshop on Multilevel Methods (Meisdorf, 1994).

Xu:1996:ASM

[2113] J. Xu. The auxiliary space method and optimal multigrid preconditioning techniques for unstructured grids. Computing, 56(3):215–235, September 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=295&
id_volume=27&id_journal=8. Inter-

national GAMM-Workshop on Multilevel Methods (Meisdorf, 1994).

Heise:1996:PSL

[2114] B. Heise and M. Kuhn. Parallel solvers for linear and nonlinear exterior magnetic field problems based upon coupled FE/BE formulations. Computing, 56(3):237–258, September 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=296&
id_volume=27&id_journal=8. Inter-

national GAMM-Workshop on Multilevel Methods (Meisdorf, 1994).

Dahmen:1996:WGM

[2115] W. Dahmen, A. Kunoth, and K. Urban. A wavelet Galerkin method for the Stokes equations. Computing, 56(3):259–301, September 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=298&
id_volume=27&id_journal=8. Inter-

national GAMM-Workshop on Multilevel Methods (Meisdorf, 1994).

Reusken:1996:RMS

[2116] A. Reusken. On a robust multigrid solver. Computing, 56(3):303–322, September 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springer.at/springer.py?Page=10&
Key=362&cat=300607/tocs/springer.
py?Page=47&Key=340&cat=3&id_abstract=
299&id_volume=27&id_journal=8. In-

ternational GAMM-Workshop on Multi-level Methods (Meisdorf, 1994).

Conti:1996:PSP

[2117] C. Conti and R. Morandi. Piecewise C^{1}-shape-preserving Hermite interpolation. Computing, 56(4):323–341, December 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=300&
id_volume=28&id_journal=8.

Ferket:1996:FDD

[2118] P. J. J. Ferket and A. A. Reusken. A finite difference discretization method for elliptic problems on composite grids. Computing, 56(4):343–369, December 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=301&
id_volume=28&id_journal=8.

Hong:1996:ERS

[2119] H. Hong. The exact region of stability for MacCormack scheme. Computing, 56(4):371–383, December 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=302&
id_volume=28&id_journal=8.

Chung:1996:PASa

[2120] K.-L. Chung, W.-M. Yan, and J.-G. Wu. A parallel algorithm for solving special tridiagonal systems on ring networks. Computing, 56(4):385–395, December 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=303&
id_volume=28&id_journal=8.

Punnen:1996:FAC

[2121] A. P. Punnen. A fast algorithm for a class of bottleneck problems. Computing, 56(4):397–401, December 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&

cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=304&
id_volume=28&id_journal=8.

Kulick:1996:SSU

[2122] T. Kulick. Shape from shading using three images. Computing, 57(1):1–24, March 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=676&
id_volume=49&id_journal=8.

Hiptmair:1996:MPA

[2123] R. Hiptmair, T. Schiekofer, and B. Wohlmuth. Multilevel preconditioned augmented Lagrangian techniques for 2nd order mixed problems. Computing, 57(1):25–48, March 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=677&
id_volume=49&id_journal=8.

Krause:1996:FAP

[2124] R. Krause and E. Rank. A fast algorithm for point-location in a finite element mesh. Computing, 57(1):49–62, March 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=678&
id_volume=49&id_journal=8.

Tippett:1996:SIS

[2125] M. K. Tippett. Symplectic integration schemes for the ABC flow. Computing, 57(1):63–75, March 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=679&
                                        id_volume=49&id_journal=8.

Rump:1996:IIS

[2126] S. M. Rump. Improved iteration schemes for validation algorithms for dense and sparse nonlinear systems. Computing, 57(1):77–84, March 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=680&
                                        id_volume=49&id_journal=8.

Koc:1996:PAF

[2127] Ç. K. Koç and B. Bakkaloğlu. A parallel algorithm for functions of triangular matrices. Computing, 57(1):85–92, March 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Vanselow:1996:RBF

[2128] R. Vanselow. Relations between FEM and FVM applied to the Poisson equation. Computing, 57(2):93–104, June 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=

47&Key=340&cat=3&id_abstract=1195&
id_volume=90&id_journal=8.

Holm:1996:VBE

[2129] H. Holm, M. Maischak, and E. P.

Stephan. The hp-version of the boundary element method for Helmholtz screen problems. Computing, 57 (2):105–134, June 1996. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

http://www.springer.at/springer.
py?Page=10&Key=362&cat=300607/tocs/
springer.py?Page=47&Key=340&cat=
3&id_abstract=1196&id_volume=90&
id_journal=8.

Lu:1996:RPQ

[2130] W. Lü. Rational parameterization of quadrics and their offsets. Computing, 57(2):135–147, June 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=1197&
id_volume=90&id_journal=8.

Kearfott:1996:IEN

[2131] R. Baker Kearfott. Interval extensions of non-smooth functions for global optimization and nonlinear systems solvers. Computing, 57(2):149–162, June 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=1198&
id_volume=90&id_journal=8.

Petkovic:1996:ICC

[2132] M. S. Petković. On initial conditions for the convergence of simultaneous root finding methods. Computing, 57(2):163–177, June 1996. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

                                        http://www.springer.at/springer.
                                        py?Page=10&Key=362&cat=300607/tocs/
                                        springer.py?Page=47&Key=340&cat=
                                        3&id_abstract=1199&id_volume=90&
                                        id_journal=8.

Spath:1996:LSF

[2133] H. Späth. Least-squares fitting by circles. Computing, 57(2):179–185, June 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=1200&
                                        id_volume=90&id_journal=8.

Brand:1996:AHA

[2134] R. Brand, W. Freeden, and J. Fröhlich. An adaptive hierarchical approximation method on the sphere using axisymmetric locally supported basis functions. Computing, 57(3):187–212, September 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

                                        springer.at/springer.py?Page=10&
                                        Key=362&cat=300607/tocs/springer.
                                        py?Page=47&Key=340&cat=3&id_abstract=
                                        1201&id_volume=91&id_journal=8.

Noskov:1996:MCF

[2135] M. V. Noskov and H. J. Schmid. Minimal cubature formulae of degree 3 for integrals over the surface of the torus.

Computing, 57(3):213–223, September 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=1202&
id_volume=91&id_journal=8.

Rao:1996:QMS

[2136] B. M. Rao and S. Ramakrishnan.

Queueing models for a single server LAN with processor sharing disciplines. Computing, 57(3):225–244, September 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=1203&
id_volume=91&id_journal=8.

Chung:1996:PAS

[2137] K.-L. Chung and Y.-C. Lin. A parametric algorithm for semigroup computation on mesh with buses. Computing, 57(3):245–253, September 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=1204&
id_volume=91&id_journal=8.

Chen:1996:TPC

[2138] S.-P. Chen, Y. He, and E.-Y. Yao.

Three-partitioning containing kernels: Complexity and heuristic. Computing, 57(3):255–271, September 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.

                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=1205&
                                        id_volume=91&id_journal=8.

Schwarz:1996:OOD

[2139] F. Schwarz. On 3^{rd} order ordinary differential equations with maximal symmetry group. Computing, 57(3):273–280, September 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=1206&
                                        id_volume=91&id_journal=8.

Liebau:1996:FVE

[2140] F. Liebau. The finite volume element method with quadratic basis functions. Computing, 57(4):281–299, December 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=1211&
                                        id_volume=92&id_journal=8.

Carstensen:1996:PEE

[2141] C. Carstensen. A posteriori error estimate for the symmetric coupling of finite elements and boundary elements. Computing, 57(4):301–322, December 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=1212&
                                        id_volume=92&id_journal=8.

Berner:1996:NRV

[2142] S. Berner. New results on verified global optimization. Computing, 57(4):323–343, December 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=1213&
id_volume=92&id_journal=8.

Ling:1996:CDO

[2143] Y. Ling. The convex-decomposable operator equation and its monotonic inclusive iteration. Computing, 57(4):345–356, December 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=1214&
id_volume=92&id_journal=8.

Herzberger:1996:OCC

[2144] J. Herzberger and L. Metzner. On the Q-order of convergence for coupled sequences arising in iterative numerical processes. Computing, 57(4):357–363, December 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=1215&
id_volume=92&id_journal=8.

Dubeau:1996:ARA

[2145] F. Dubeau. Algorithms for nth root approximation. Computing, 57(4):365–369, December 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=1216&
id_volume=92&id_journal=8.

Axelsson:1997:ARC

[2146] O. Axelsson and M. Nikolova. Adaptive refinement for convection-diffusion problems based on a defect-correction technique and finite difference method. Computing, 58(1):1–30, March 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

http://www.springer.at/springer.
py?Page=10&Key=362&cat=300607/tocs/
springer.py?Page=47&Key=340&cat=
3&id_abstract=1182&id_volume=87&
id_journal=8.

Elster:1997:MTR

[2147] C. Elster and A. Neumaier. A method of trust region type for minimizing noisy functions. Computing, 58(1):31– 46, March 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springer.at/springer.py?Page=10&
Key=362&cat=300607/tocs/springer.
py?Page=47&Key=340&cat=3&id_abstract=
1183&id_volume=87&id_journal=8.

Galligani:1997:EAA

[2148] E. Galligani and L. Zanni. Error analysis of an algorithm for equalityconstrained quadratic programming problems. Computing, 58(1):47–67, March 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&

cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=1184&
id_volume=87&id_journal=8.

Spedicato:1997:NEN

[2149] E. Spedicato and Z. Huang. Numerical experience with Newton-like methods for nonlinear algebraic systems. Computing, 58(1):69–89, March 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=1185&
id_volume=87&id_journal=8.

vonIssendorff:1997:SGC

[2150] H. von Issendorff. Some generalizations of the calculus of binary events (short communication). Computing, 58(1):91– 94, March 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springer.at/springer.py?Page=10&
Key=362&cat=300607/tocs/springer.
py?Page=47&Key=340&cat=3&id_abstract=
1186&id_volume=87&id_journal=8.

Tautenhahn:1997:UTS

[2151] T. Tautenhahn and G. J. Wöginger.

Unit-time scheduling problems with time dependent resources. Computing, 58(2):97–111, June 1997. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

http://www.springer.at/springer.
py?Page=10&Key=362&cat=300607/tocs/
springer.py?Page=47&Key=340&cat=
3&id_abstract=1217&id_volume=93&
id_journal=8.

Eckardt:1997:SRL

[2152] H. Eckardt. A scalable record locking scheme for parallel file access. Computing, 58(2):113–128, June 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=1218&
                                        id_volume=93&id_journal=8.

Hackbusch:1997:FVS

[2153] W. Hackbusch. On the feedback vertex set problem for a planar graph. Computing, 58(2):129–155, June 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=1219&
                                        id_volume=93&id_journal=8.

Al-Musallam:1997:GGP

[2154] F. Al-Musallam, M. Al-Zanaidi, and C. Grossmann. A grid generator for problems on unbounded intervals with rationally decaying solutions. Computing, 58(2):157–171, June 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=1220&
                                        id_volume=93&id_journal=8.

Engstler:1997:MEM

[2155] Ch. Engstler and Ch. Lubich. Multirate extrapolation methods for differential equations with different time scales. Computing, 58(2):173–185, June 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=1221&
id_volume=93&id_journal=8.

Hansen:1997:PLE

[2156] E. R. Hansen. Preconditioning linearized equations. Computing, 58(2):187–196, June 1997. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

http://www.springer.at/springer.
py?Page=10&Key=362&cat=300607/tocs/
springer.py?Page=47&Key=340&cat=
3&id_abstract=1222&id_volume=93&
id_journal=8.

Nemeth:1997:SPP

[2157] G. Németh, I. Lovrek, and V. Sinkovic.

Scheduling problems in parallel systems for telecommunications. Computing, 58(3):199–223, September 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=1401&
id_volume=107&id_journal=8.

Dekker:1997:SGH

[2158] T. J. Dekker, W. Hoffmann, and K. Potma. Stability of the Gauss– Huard algorithm with partial pivoting. Computing, 58(3):225–244, September 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&

                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=1402&
                                        id_volume=107&id_journal=8.

Lirkov:1997:CBF

[2159] I. D. Lirkov, S. D. Margenov, and L. T. Zikatanov. Circulant block-factorization preconditioning of anisotropic elliptic problems. Computing, 58(3):245–258, September 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=1403&
                                        id_volume=107&id_journal=8.

Schroter:1997:SCN

[2160] T. Schröter. On a special class of nonlinear Fredholm integral equations of the first kind. Computing, 58(3):259–279, September 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=1749&
                                        id_volume=107&id_journal=8.

Chen:1997:VMS

[2161] X. Chen. A verification method for solutions of nonsmooth equations. Computing, 58(3):281–294, September 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=1404&
                                        id_volume=107&id_journal=8.

Ling:1997:MTC

[2162] Y. Ling. On the Moore test for coupled equations. Computing, 58(3):295–301, September 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springer.at/springer.py?Page=10&
Key=362&cat=300607/tocs/springer.
py?Page=47&Key=340&cat=3&id_abstract=
1405&id_volume=107&id_journal=8.

Weiler:1997:PFF

[2163] W. Weiler and G. Wittum. Parallel frequency filtering. Computing, 58(4):303–316, December 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=1479&
id_volume=114&id_journal=8.

Fuhrer:1997:PEC

[2164] C. Führer and G. Kanschat. A posteriori error control in radiative transfer. Computing, 58(4):317–334, December 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=1476&
id_volume=114&id_journal=8.

Shen:1997:CSB

[2165] Y.-Q. Shen. Computation of a simple bifurcation point using one singular value decomposition nearby. Computing, 58(4):335–350, December 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=1478&
id_volume=114&id_journal=8.

Ferretti:1997:HOA

[2166] R. Ferretti. High-order approximations of linear control systems via Runge–Kutta schemes. Computing, 58(4):351–364, December 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=1475&
id_volume=114&id_journal=8.

Famoye:1997:SGL

[2167] F. Famoye. Sampling from the generalized logarithmic series distribution. Computing, 58(4):365–375, December 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=1477&
id_volume=114&id_journal=8.

Schmidt:1997:IDS

[2168] J. W. Schmidt. Interpolation in a derivative strip. Computing, 58 (4):377–389, December 1997. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

http://www.springer.at/springer.
py?Page=10&Key=362&cat=300607/tocs/
springer.py?Page=47&Key=340&cat=
3&id_abstract=1474&id_volume=114&
id_journal=8.

Storoy:1997:MPA

[2169] S. Storøy and T. Sørevik. Massively parallel augmenting path algorithms for the assignment problem. Computing, 59(1):1–16, March 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=1453&
id_volume=112&id_journal=8.

Zitnan:1997:DLS

[2170] P. Z̆itn̆an. Discrete least-squares technique for eigenvalues. part I: The one-dimensional case. Computing, 59(1):17–27, March 1997. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

http://www.springer.at/springer.
py?Page=10&Key=362&cat=300607/tocs/
springer.py?Page=47&Key=340&cat=
3&id_abstract=1452&id_volume=112&
id_journal=8.

Wieners:1997:BLE

[2171] C. Wieners. Bounds for the N lowest eigenvalues of fourth-order boundary value problems. Computing, 59(1):29– 41, March 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springer.at/springer.py?Page=10&
Key=362&cat=300607/tocs/springer.
py?Page=47&Key=340&cat=3&id_abstract=
1449&id_volume=112&id_journal=8.

Murua:1997:PHE

[2172] A. Murua. Partitioned half-explicit Runge–Kutta methods for differentialalgebraic systems of index 2. Computing, 59(1):43–61, March 1997. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

http://www.springer.at/springer.
py?Page=10&Key=362&cat=300607/tocs/
springer.py?Page=47&Key=340&cat=
3&id_abstract=1450&id_volume=112&
id_journal=8.

Adams:1997:CDS

[2173] E. Adams, K. Baumann, and Ch. Grossmann. Comparison of direct to shooting enclosures for an inverse-monotone boundary value problem with locally steep solution. Computing, 59(1):63–83, March 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=1448&
id_volume=112&id_journal=8.

Eichenauer-Herrmann:1997:CCC

[2174] J. Eichenauer-Herrmann and E. Herrmann. Compound cubic congruential pseudo-random numbers. Computing, 59(1):85–90, March 1997. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

http://www.springer.at/springer.
py?Page=10&Key=362&cat=300607/tocs/
springer.py?Page=47&Key=340&cat=
3&id_abstract=1451&id_volume=112&
id_journal=8.

Tuan:1997:CLP

[2175] H. D. Tuan. Can linear programs be used to test global optimization algorithms? Computing, 59(1):91–93, March 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=1454&
id_volume=112&id_journal=8.

Horst:1997:EMI

[2176] R. Horst and N. V. Thoai. Erratum: “Modification, implementation and comparison of three algorithms for globally solving linearly constrained concave minimization problems” [Computing 42 (1989), no. 2–3, 271–289; MR 90j:90119]. Computing, 59(1):94, March 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [1743].

Chen:1997:MAT

[2177] Y.-J. Chen and S.-J. Horng. Medial axis transform on mesh-connected computers with hyperbus broadcasting. Computing, 59(2):95–114, June 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=1536&
id_volume=120&id_journal=8.

Traxler:1997:AAM

[2178] C. T. Traxler. An algorithm for adaptive mesh refinement in n dimensions. Computing, 59(2):115–137, June 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=1537&
id_volume=120&id_journal=8.

Bruder:1997:NRP

[2179] J. Bruder. Numerical results for a parallel linearly-implicit Runge–Kutta method. Computing, 59(2):139–151, June 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=1538&
                                        id_volume=120&id_journal=8.

Denk:1997:AME

[2180] G. Denk and S. Schäffler. Adams methods for the efficient solution of stochastic differential equations with additive noise. Computing, 59(2):153–161, June 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=1539&
                                        id_volume=120&id_journal=8.

Herendi:1997:FGR

[2181] T. Herendi, T. Siegl, and R. F. Tichy. Fast Gaussian random number generation using linear transformations. Computing, 59(2):163–181, June 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=1557&
                                        id_volume=120&id_journal=8.

Huang:1997:RFP

[2182] K. Huang and J. Yu. Remarks on a formula for the pseudoinverse by the modified Huang algorithm. Computing, 59(2):183–185, June 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=1540&
id_volume=120&id_journal=8.

Rocha:1997:RCT

[2183] L. Rocha. On the rate of convergence of 2-term recursions in R^{d}. Computing, 59(3):187–207, September 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=1748&
id_volume=131&id_journal=8.

Ponisch:1997:CMP

[2184] G. Pönisch, U. Schnabel, and H. Schwetlick. Computing multiple pitchfork bifurcation points. Computing, 59(3):209–222, September 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=1646&
id_volume=131&id_journal=8.

Klinger:1997:NIS

[2185] B. Klinger. Numerical integration of singular integrands using low-discrepancy sequences. Computing, 59(3):223–236, September 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=

                                        47&Key=340&cat=3&id_abstract=1647&
                                        id_volume=131&id_journal=8.

Pferschy:1997:SMC

[2186] U. Pferschy. Solution methods and computational investigations for the linear bottleneck assignment problem. Computing, 59(3):237–258, September 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=1648&
                                        id_volume=131&id_journal=8.

DeFalco:1997:APH

[2187] I. De Falco, R. Del Balio, and E. Tarantino. An analysis of parallel heuristics for task allocation in multicomputers. Computing, 59(3):259–275, September 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=1649&
                                        id_volume=131&id_journal=8.

Gruner:1997:CPE

[2188] J. Grüner and K. Zoller. Computing partial expectations from tables. Computing, 59(3):277–284, September 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/springer.py?Page=10&Key=362&
                                        cat=300607/tocs/springer.py?Page=
                                        47&Key=340&cat=3&id_abstract=1650&
                                        id_volume=131&id_journal=8.

Neubert:1997:ASC

[2189] R. Neubert. Approximation of solution curves of underdetermined systems of nonlinear equations. Computing, 59(4):285–306, December 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=2504&
id_volume=212&id_journal=8.

Tilli:1997:PRF

[2190] P. Tilli. Polynomial root finding by means of continuation. Computing, 59(4):307–324, December 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=2505&
id_volume=212&id_journal=8.

Hidber:1997:GCZ

[2191] C. Hidber. A generalisation of the Cantor-Zassenhaus algorithm. Computing, 59(4):325–330, December 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=2506&
id_volume=212&id_journal=8.

Struckmeier:1997:GRV

[2192] J. Struckmeier. Generation of random variates using asymptotic expansions. Computing, 59(4):331–347, December 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=2503&
id_volume=212&id_journal=8.

Stahl:1997:SCN

[2193] V. Stahl. A sufficient condition for non-overestimation in interval arithmetic. Computing, 59(4):349–363, December 1997. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=2507&
id_volume=212&id_journal=8.

Ohlberger:1997:HAV

[2194] M. Ohlberger and M. Rumpf. Hierarchical and adaptive visualization on nested grids. Computing, 59 (4):365–385, December 1997. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

http://www.springer.at/springer.
py?Page=10&Key=362&cat=300607/tocs/
springer.py?Page=47&Key=340&cat=
3&id_abstract=2508&id_volume=212&
id_journal=8.

Scherzer:1998:DHO

[2195] O. Scherzer. Denoising with higher order derivatives of bounded variation and an application to parameter estimation. Computing, 60(1):1–27, March 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=

47&Key=340&cat=3&id_abstract=2509&
id_volume=213&id_journal=8.

DeMarchi:1998:CDI

[2196] S. De Marchi. On computing derivatives for C^{1} interpolating schemes: An optimization. Computing, 60(1):29–53, March 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=2510&
id_volume=213&id_journal=8.

VanKeer:1998:VAM

[2197] R. Van Keer and H. De Schepper. A variational approximation method for 2nd order elliptic eigenvalue problems in a composite structure with nonstandard boundary conditions. Computing, 60(1):55–72, March 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=2511&
id_volume=213&id_journal=8.

Sevastianov:1998:MMP

[2198] S. V. Sevastianov and G. J. Wöginger.

Makespan minimization in preemptive two machine job shops. Computing, 60(1):73–79, March 1998. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

http://www.springer.at/springer.
py?Page=10&Key=362&cat=300607/tocs/
springer.py?Page=47&Key=340&cat=
3&id_abstract=2512&id_volume=213&
id_journal=8.

Dobner:1998:VCL

[2199] H.-J. Dobner and S. Ritter. Verified computation of Lamé functions with high accuracy. Computing, 60(1):81– 89, March 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

                                        springer.at/springer.py?Page=10&
                                        Key=362&cat=300607/tocs/springer.
                                        py?Page=47&Key=340&cat=3&id_abstract=
                                        2513&id_volume=213&id_journal=8.

Petrosino:1998:PIU

[2200] A. Petrosino and E. Tarantino. Parallel image understanding algorithms on MIMD multicomputers. Computing, 60(2):91–107, June 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dudas:1998:STS

[2201] T. Dudás and R. Rudolf. Spanning trees and shortest paths in Monge graphs. Computing, 60(2):109–119, June 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ganley:1998:EEL

[2202] J. L. Ganley and L. S. Heath. An experimental evaluation of local search heuristics for graph partitioning. Computing, 60(2):121–132, June 1998. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lavallee:1998:PLM

[2203] P.-F. Lavallée and M. Sadkane. Pseudospectra of linear matrix pencils by block diagonalization. Computing, 60 (2):133–156, June 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Apel:1998:AIE

[2204] Th. Apel. Anisotropic interpolation error estimates for isoparametric quadrilateral finite elements. Computing, 60(2):157–174, June 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hackbusch:1998:RPE

[2205] W. Hackbusch and J. U. Wappler. Remarks on A posteriori error estimation for inaccurate finite element solutions. Computing, 60(2):175–191, June 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Burkard:1998:LAP

[2206] R. E. Burkard and J. Krarup. A linear algorithm for the pos/neg-weighted 1-Median problem on a cactus. Computing, 60(3):193–215, September 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

http://www.springer.at/journals/
pdf/2731.pdf.

Karaata:1998:SSA

[2207] M. H. Karaata and P. Chaudhuri. A self-stabilizing algorithm for strong fairness. Computing, 60(3): 217–228, September 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springer.at/journals/pdf/
2732.pdf.

Lassas:1998:ECS

[2208] M. Lassas and E. Somersalo. On the existence and convergence of the solution of PML equations. Computing, 60(3):229–241, September 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

http://www.springer.at/journals/
pdf/2733.pdf.

Mund:1998:TLM

[2209] P. Mund, E. P. Stephan, and J. Weiße. Two-level methods for the single layer potential in R^{3}. Computing, 60(3): 243–266, September 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springer.at/journals/pdf/
2734.pdf.

Zhang:1998:PLO

[2210] G. Zhang. Parameterized on-line openend bin packing. Computing, 60(3): 267–273, September 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springer.at/journals/pdf/
2735.pdf.

Bellavia:1998:SMM

[2211] S. Bellavia, M. Macconi, and B. Morini. The switching-method in multiple shooting. Computing, 60(3):275– 283, September 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springer.at/journals/pdf/
2736.pdf.

Verdoscia:1998:HLD

[2212] L. Verdoscia and R. Vaccaro. A high-level dataflow system. Computing, 60(4):285–305, December 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

http://www.springer.at/journals/
pdf/3159.pdf.

Pottmann:1998:RHS

[2213] H. Pottmann and T. Randrup. Rotational and helical surface approximation for reverse engineering. Computing, 60(4):307–322, December 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=3156&
id_volume=289&id_journal=8.

Schoberl:1998:SSP

[2214] J. Schöberl. Solving the Signorini problem on the basis of domain decomposition techniques. Computing, 60(4):323–344, December 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=3160&
id_volume=289&id_journal=8.

Bastian:1998:AMM

[2215] P. Bastian, W. Hackbusch, and G. Wittum. Additive and multiplicative multi-grid — a comparison. Computing, 60(4):345–364, December 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/springer.py?Page=10&Key=362&
cat=300607/tocs/springer.py?Page=
47&Key=340&cat=3&id_abstract=3157&
id_volume=289&id_journal=8.

El-Daou:1998:TMA

[2216] M. K. El-Daou and E. L. Ortiz. The tau method as an analytic tool in the discussion of equivalence results across numerical methods. Computing, 60(4):365–376, December 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

                                        http://www.springer.at/journals/
                                        pdf/3158.pdf.

Burkard:1998:LCA

[2217] R. E. Burkard, Y. He, and H. Kellerer. A linear compound algorithm for uniform machine scheduling. Computing, 61(1):1–9, March 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springer.at/journals/pdf/
                                        3207.pdf.

Chung:1998:PFA

[2218] K.-L. Chung and H.-N. Chen. Parallel finding all initial palindromes and periods of a string on reconfigurable meshes. Computing, 61(1):11–21, March 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://www.springer.
                                        at/journals/pdf/3208.pdf.

Jiang:1998:COG

[2219] X. Jiang and H. Bunke. On the coding of ordered graphs. Computing, 61(1):23–38, March 1998. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

                                        http://www.springer.at/journals/
                                        pdf/3209.pdf.

Schwarz:1998:ASA

[2220] F. Schwarz. Algorithmic solution of Abel’s equation. Computing, 61 (1):39–46, March 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springer.at/journals/pdf/
3210.pdf.

Kuhn:1998:RCO

[2221] W. Kühn. Rigorously computed orbits of dynamical systems without the wrapping effect. Computing, 61 (1):47–67, March 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springer.at/journals/pdf/
3211.pdf.

Weiner:1998:ORK

[2222] R. Weiner and B. A. Schmitt. Order results for Krylov-W -methods. Computing, 61(1):69–89, March 1998. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

http://www.springer.at/journals/
pdf/3212.pdf.

Meinardus:1998:CDS

[2223] G. Meinardus and W. Oettli. On a certain duality for segmented minmax problems. Computing, 61 (2):91–101, June 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springer.at/journals/pdf/
3572.pdf;   http://www.springer.
at/journals/pdf/3812.pdf.

Szirmay-Kalos:1998:WCV

[2224] L. Szirmay-Kalos and G. Márton.

Worst-case versus average case complexity of ray-shooting. Computing, 61(2):103–131, June 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springer.at/journals/pdf/

                                        3574.pdf;   http://www.springer.
                                        at/journals/pdf/3813.pdf.

Richter:1998:UBS

[2225] M. Richter. Use of box splines in computer tomography. Computing, 61(2):133–150, June 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springer.at/journals/pdf/
                                        3573.pdf;   http://www.springer.
                                        at/journals/pdf/3814.pdf.

Griebel:1998:ASG

[2226] M. Griebel. Adaptive sparse grid multilevel methods for elliptic PDEs based on finite differences. Computing, 61(2):151–179, June 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springer.at/journals/pdf/
                                        3571.pdf;   http://www.springer.
                                        at/journals/pdf/3815.pdf.

Cap:1998:NNM

[2227] F. F. Cap. A new numerical method to calculate the vibrations of clamped Kirchhoff-plates of arbitrary form. Computing, 61(2): 181–188, June 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springer.at/journals/pdf/
                                        3570.pdf;   http://www.springer.
                                        at/journals/pdf/3816.pdf.

Chen:1998:FEV

[2228] F. Chen and D. Suter. Fast evaluation of vector splines in three dimensions. Computing, 61(3):189–213, September 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/journals/pdf/3809.pdf.

Rossini:1998:DDS

[2229] M. Rossini. 2D-discontinuity detection from scattered data. Computing, 61(3):215–234, September 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

http://www.springer.at/journals/
pdf/3807.pdf.

Oswald:1998:MN

[2230] P. Oswald. Multilevel norms for H^{−1/2}.

Computing, 61(3):235–255, September 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/journals/pdf/3811.pdf.

Buzdin:1998:TD

[2231] A. Buzdin. Tangential decomposition.

Computing, 61(3):257–276, September 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/journals/pdf/3810.pdf.

Burkard:1998:NMS

[2232] R. E. Burkard and Y. He. A note on MULTIFIT scheduling for uniform machines. Computing, 61(3): 277–283, September 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springer.at/journals/pdf/
3808.pdf.

Anthonissen:1998:FVS

[2233] M. J. H. Anthonissen, B. van ’t Hof, and A. A. Reusken. A finite volume scheme for solving elliptic boundary value problems on composite grids.

Computing, 61(4):285–305, December 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/journals/pdf/4463.pdf.

Globisch:1998:HPU

[2234] G. Globisch and S. V. Nepomnyaschikh. The hierarchial preconditioning on unstructured grids. Computing, 61(4):307–330, December 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

http://www.springer.at/journals/
pdf/4466.pdf.

Axelsson:1998:ASP

[2235] O. Axelsson and M. Nikolova. Avoiding slave points in an adaptive refinement procedure for convectiondiffusion problems in 2D. Computing, 61(4):331–357, December 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

http://www.springer.at/journals/
pdf/4464.pdf.

Famoye:1998:BBT

[2236] F. Famoye. Bootstrap based tests for generalized negative binomial distribution. Computing, 61(4):359–369, December 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.
at/journals/pdf/4465.pdf.

Tuffin:1998:VRO

[2237] B. Tuffin. Variance reduction order using good lattice points in Monte Carlo methods. Computing, 61(4): 371–378, December 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springer.at/journals/pdf/
4467.pdf.

Anonymous:1999:PPV

[2238] Anonymous. Preface. Computing, 62(1):v–??, March 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schwarz:1999:SSO

[2239] F. Schwarz. Solving second order ordinary differential equations with maximal symmetry group. Computing, 62(1):1–10, March 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://

link.springer-ny.com/link/service/
journals/00607/bibs/9062001/90620001.
htm;   http://link.springer-ny.
com/link/service/journals/00607/
papers/9062001/90620001.pdf.

Paone:1999:DTR

[2240] A. Paone and A. Vecchio. Discrete time relaxation based on direct quadrature methods for Volterra integral equations. Computing, 62(1):11–25, March 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/9062001/90620011.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/9062001/90620011.
pdf.

Otto:1999:PEN

[2241] F.-C. Otto and G. Lube. A posteriori estimates for a nonoverlapping domain decomposition method. Computing, 62(1):27–43, March 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/9062001/90620027.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/9062001/90620027.
pdf.

Korzak:1999:ERC

[2242] J. Korzak. Eigenvalue relations and conditions of matrices arising in linear programming. Computing, 62(1):45– 54, March 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://link.

springer-ny.com/link/service/journals/
00607/bibs/9062001/90620045.htm;
 http://link.springer-ny.com/link/
service/journals/00607/papers/9062001/
90620045.pdf.

Chaudhuri:1999:SSA

[2243] P. Chaudhuri. An O(n^{2}) self-stabilizing algorithm for computing bridgeconnected components. Computing, 62(1):55–67, March 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://

link.springer-ny.com/link/service/
journals/00607/bibs/9062001/90620055.
htm;   http://link.springer-ny.
com/link/service/journals/00607/
papers/9062001/90620055.pdf.

Petkovic:1999:HLM

[2244] M. S. Petković. Halley-like method with corrections for the inclusion of polynomial zeros. Computing, 62 (1):69–88, March 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://

link.springer-ny.com/link/service/
journals/00607/bibs/9062001/90620069.

htm;   http://link.springer-ny.
com/link/service/journals/00607/
papers/9062001/90620069.pdf.

Hackbusch:1999:SMA

[2245] W. Hackbusch. A sparse matrix arithmetic based on ⟨-matrices. Part I: Introduction to ⟨-matrices. Computing, 62(2):89–108, June 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://

link.springer-ny.com/link/service/
journals/00607/bibs/9062002/90620089.
htm;   http://link.springer-ny.
com/link/service/journals/00607/
papers/9062002/90620089.pdf.

Jung:1999:VAP

[2246] M. Jung and S. V. Nepomnyaschikh.

Variable additive preconditioning procedures. Computing, 62(2):109–128, June 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/9062002/90620109.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/9062002/90620109.
pdf.

Kravanja:1999:MNM

[2247] P. Kravanja and A. Haegemans. A modification of Newton’s method for analytic mappings having multiple zeros. Computing, 62(2):129–145, June 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/9062002/90620129.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/9062002/90620129.
pdf.

Cools:1999:SCC

[2248] R. Cools, E. Novak, and K. Ritter. Smolyak’s construction of cubature formulas of arbitrary trigonometric degree. Computing, 62(2):147–162, June 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/9062002/90620147.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/9062002/90620147.
                                        pdf.

Jukic:1999:PLO

[2249] D. Jukić, R. Scitovski, and H. Späth. Partial linearization of one class of the nonlinear total least squares problem by using the inverse model function. Computing, 62(2):163–178, June 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/9062002/90620163.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/9062002/90620163.
                                        pdf.

He:1999:SLS

[2250] Y. He and G. Zhang. Semi on-line scheduling on two identical machines. Computing, 62(3):179–187, September 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/9062003/90620179.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/9062003/90620179.
                                        pdf.

DelCorso:1999:FES

[2251] G. M. Del Corso and G. Manzini.

Finding exact solutions to the bandwidth minimization problem. Computing, 62(3):189–203, September 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/9062003/90620189.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/9062003/90620189.
pdf.

Mandel:1999:EOA

[2252] J. Mandel, M. Brezina, and P. Vaněk.

Energy optimization of algebraic multigrid bases. Computing, 62(3):205–228, September 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://link.

springer-ny.com/link/service/journals/
00607/bibs/9062003/90620205.htm;
http://link.springer-ny.com/link/
service/journals/00607/papers/9062003/
90620205.pdf.

Carstensen:1999:CNFa

[2253] C. Carstensen and S. A. Funken. Coupling of nonconforming finite elements and boundary elements I: a posteriori estimates and adaptive meshrefinement. Computing, 62(3):229–241, September 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://link.

springer-ny.com/link/service/journals/
00607/bibs/9062003/90620229.htm;
http://link.springer-ny.com/link/
service/journals/00607/papers/9062003/
90620229.pdf.

Carstensen:1999:CNF

[2254] C. Carstensen and S. A. Funken. Coupling of nonconforming finite elements and boundary elements. II. A posteriori estimates and adaptive meshrefinement. Computing, 62(3):243–259, September 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://link.

springer-ny.com/link/service/journals/
00607/bibs/9062003/90620243.htm;
 http://link.springer-ny.com/link/
service/journals/00607/papers/9062003/
90620243.pdf.

Loupias:1999:MI

[2255] E. Loupias, S. Bres, and J. M. Jolion. From methods to images. Computing, 62(4):265–275, December 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/9062004/90620265.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/9062004/90620265.
pdf.

Goller:1999:PPS

[2256] A. Goller. Parallel processing strategies for large SAR image data sets in a distributed environment. Computing, 62(4):277–291, December 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/9062004/90620277.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/9062004/90620277.
pdf.

Borotschnig:1999:CPP

[2257] H. Borotschnig, L. Paletta, and A. Pinz. A comparison of probabilistic, possibilistic and evidence theoretic fusion schemes for active object recognition. Computing, 62(4):293–319, December 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/9062004/90620293.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/9062004/90620293.
pdf.

Chetverikov:1999:FPT

[2258] D. Chetverikov and J. Verestóy. Feature point tracking for incomplete trajectories. Computing, 62(4):321–338, December 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://link.

springer-ny.com/link/service/journals/
00607/bibs/9062004/90620321.htm;
http://link.springer-ny.com/link/
service/journals/00607/papers/9062004/
90620321.pdf.

Hlavac:1999:APL

[2259] V. Hlaváč and J. Fojtík. Adaptive predictor for lossless image compression. Computing, 62(4):339–354, December 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/9062004/90620339.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/9062004/90620339.
pdf.

Burge:1999:MLP

[2260] M. Burge and W. G. Kropatsch. A minimal line property preserving representation of line images. Computing, 62(4):355–368, December 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/9062004/90620355.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/9062004/90620355.
pdf.

Englert:1999:ACM

[2261] R. Englert. Acquisition of complex model knowledge by domain theory-controlled generalization. Computing, 62(4):369–385, December 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/9062004/90620369.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/9062004/90620369.
pdf.

Brass:1999:MMN

[2262] D. Bräss, M. Dryja, and W. Hackbusch. A multigrid method for nonconforming FE-discretisations with application to non-matching grids. Computing, 63(1): 1–25, March 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://link.

springer-ny.com/link/service/journals/
00607/bibs/9063001/90630001.htm;
 http://link.springer-ny.com/link/
service/journals/00607/papers/9063001/
90630001.pdf.

Roos:1999:SCU

[2263] H.-G. Roos and T. Linß. Sufficient conditions for uniform convergence on layer-adapted grids. Computing, 63(1):27–45, March 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://

link.springer-ny.com/link/service/
journals/00607/bibs/9063001/90630027.
htm;   http://link.springer-ny.
com/link/service/journals/00607/
papers/9063001/90630027.pdf.

Fadrani:1999:PSS

[2264] D. Rostami Varnos Fadrani and K. Maleknejad. Preconditioners for solving stochastic boundary integral equations with weakly singular kernels. Computing, 63(1):47–67, March 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/9063001/90630047.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/9063001/90630047.
pdf.

Kravanja:1999:DFA

[2265] P. Kravanja and M. Van Barel. A derivative-free algorithm for computing zeros of analytic functions. Computing, 63(1):69–91, March 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://

link.springer-ny.com/link/service/
journals/00607/bibs/9063001/90630069.
htm;   http://link.springer-ny.
com/link/service/journals/00607/
papers/9063001/90630069.pdf.

Schatte:1999:CAB

[2266] P. Schatte. Computing the angle between vectors. Computing, 63(1):93–96, March 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/9063001/90630093.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/9063001/90630093.
                                        pdf.

Bjorstad:1999:HPS

[2267] P. E. Bjørstad and B. P. Tjøstheim. High precision solutions of two fourth order eigenvalue problems. Computing, 63(2):97–107, June 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        journals/00607/bibs/9063002/90630097.
                                        htm;   http://link.springer-ny.
                                        com/link/service/journals/00607/
                                        papers/9063002/90630097.pdf.

Nagatou:1999:NMV

[2268] K. Nagatou. A numerical method to verify the elliptic eigenvalue problems including a uniqueness property. Computing, 63(2):109–130, June 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/9063002/90630109.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/9063002/90630109.
                                        pdf.

Argyros:1999:RBF

[2269] I. K. Argyros. Relations between forcing sequences and inexact Newton iterates in Banach space. Computing, 63(2):131–144, June 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/9063002/90630131.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/9063002/90630131.
pdf.

Ariyawansa:1999:GCC

[2270] K. A. Ariyawansa and N. Begashaw.

Global convergence of a class of collinear scaling algorithms with inexact line searches on convex functions. Computing, 63(2):145–169, June 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/9063002/90630145.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/9063002/90630145.
pdf.

Verdoscia:1999:ARA

[2271] L. Verdoscia and R. Vaccaro. An adaptive routing algorithm for WKrecursive topologies. Computing, 63 (2):171–184, June 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://

link.springer-ny.com/link/service/
journals/00607/bibs/9063002/90630171.
htm;   http://link.springer-ny.
com/link/service/journals/00607/
papers/9063002/90630171.pdf.

Seif:1999:IPI

[2272] N. P. Seif, M. A. Hassanein, and A. S. Deif. Inverse problem of the interval linear system of equations. Computing, 63(2):185–200, June 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/9063002/90630185.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/9063002/90630185.
                                        pdf.

Sinkovic:1999:LBD

[2273] V. Sinkovic, I. Lovrek, and G. Németh. Load balancing in distributed parallel systems for telecommunications. Computing, 63(3):201–218, September 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/9063003/90630201.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/9063003/90630201.
                                        pdf.

Oliveira:1999:CRP

[2274] S. Oliveira. On the convergence rate of a preconditioned subspace eigensolver. Computing, 63(3):219–231, September 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/9063003/90630219.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/9063003/90630219.
                                        pdf.

Brezina:1999:BBI

[2275] M. Brezina and P. Vanek. A black-box iterative solver based on a two-level Schwarz method. Computing, 63(3):233–263, September 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/9063003/90630233.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/9063003/90630233.
pdf.

Morini:1999:IMN

[2276] B. Morini and M. Macconi. Inexact methods in the numerical solution of stiff initial value problems. Computing, 63(3):265–281, September 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/9063003/90630265.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/9063003/90630265.
pdf.

Turek:1999:PRF

[2277] S. Turek, Chr. Becker, and A. Runge.

Performance rating via the Feast Indices. Computing, 63(3):283–297, September 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://link.

springer-ny.com/link/service/journals/
00607/bibs/9063003/90630283.htm;
 http://link.springer-ny.com/link/
service/journals/00607/papers/9063003/
90630283.pdf.

Brucker:1999:CRS

[2278] P. Brucker and S. Knust. Complexity results for single-machine problems with positive finish-start timelags. Computing, 63(4):299–316, December 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/

                                        bibs/9063004/90630299.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/9063004/90630299.
                                        pdf.

Burkard:1999:VMO

[2279] R. E. Burkard and M. M. Miatselski. Volume maximization and orthoconvex approximation of orthogons. Computing, 63(4):317–330, December 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/9063004/90630317.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/9063004/90630317.
                                        pdf.

Goemans:1999:SPA

[2280] M. X. Goemans and F. Rendl. Semidefinite programs and association schemes. Computing, 63(4):331–340, December 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://link.

                                        springer-ny.com/link/service/journals/
                                        00607/bibs/9063004/90630331.htm;
                                         http://link.springer-ny.com/link/
                                        service/journals/00607/papers/9063004/
                                        90630331.pdf.

Heuberger:1999:MER

[2281] C. Heuberger. Minimal expansions in redundant number systems and shortest paths in graphs. Computing, 63(4):341–349, December 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/9063004/90630341.htm;   http:
                                        //link.springer-ny.com/link/service/

journals/00607/papers/9063004/90630341.
pdf.

Karisch:1999:DFL

[2282] S. E. Karisch, E. Çela, J. Clausen, and T. Espersen. A dual framework for lower bounds of the quadratic assignment problem based on linearization. Computing, 63(4):351–403, December 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/9063004/90630351.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/9063004/90630351.
pdf.

Kleinschmidt:1999:NHM

[2283] P. Kleinschmidt and C. Rank. New heuristics for the multidimensional vote assignment problem. Computing, 63(4):405–417, December 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/9063004/90630405.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/9063004/90630405.
pdf.

Pferschy:1999:DPR

[2284] U. Pferschy. Dynamic programming revisited: Improving knapsack algorithms. Computing, 63(4):419–430, December 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/9063004/90630419.htm;   http:
//link.springer-ny.com/link/service/

                                        journals/00607/papers/9063004/90630419.
                                        pdf.

Burkard:1999:E

[2285] R. E. Burkard. Editorial. Computing, 63(4):v, December 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        journals/00607/bibs/9063004/90630v00.
                                        htm;   http://link.springer-ny.
                                        com/link/service/journals/00607/
                                        papers/9063004/90630v00.pdf.

Andrisani:2000:IDT

[2286] E. Andrisani and G. Di Lena. An improved dqds type algorithm. Computing, 64(1):1–19, March 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        journals/00607/bibs/0064001/00640001.
                                        htm;   http://link.springer-ny.
                                        com/link/service/journals/00607/
                                        papers/0064001/00640001.pdf.

Hackbusch:2000:SHM

[2287] W. Hackbusch and B. N. Khoromskij. A sparse H-matrix arithmetic. Part II: Application to multi-dimensional problems. Computing, 64(1):21–47, March 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/0064001/00640021.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/0064001/00640021.
                                        pdf.

Lube:2000:NOD

[2288] G. Lube, L. Müller, and F. C. Otto. A non-overlapping domain decomposition method for the advection-diffusion problem. Computing, 64(1):49–68, March 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/0064001/00640049.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/0064001/00640049.
pdf.

Buzdin:2000:IBT

[2289] A. A. Buzdin and D. L. Logashenko. Incomplete block triangular decompositions of order ℓ. Computing, 64(1):69–95, March 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/0064001/00640069.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/0064001/00640069.
pdf.

Hiptmair:2000:MGE

[2290] Ralf Hiptmair. Multilevel gauging for edge elements. Computing, 64(2):97– 122, June 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://link.

springer-ny.com/link/service/journals/
00607/bibs/0064002/00640097.htm;
http://link.springer-ny.com/link/
service/journals/00607/papers/0064002/
00640097.pdf.

LeBorne:2000:OTT

[2291] Sabine Le Borne. Ordering techniques for two- and three-dimensional convection-dominated elliptic boundary value problems. Computing, 64 (2):123–155, June 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://

link.springer-ny.com/link/service/
journals/00607/bibs/0064002/00640123.
htm;   http://link.springer-ny.
com/link/service/journals/00607/
papers/0064002/00640123.pdf.

Ewing:2000:FVE

[2292] R. E. Ewing, R. D. Lazarov, and Y. Lin. Finite volume element approximations of nonlocal in time one-dimensional flows in porous media. Computing, 64(2):157–182, June 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/0064002/00640157.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/0064002/00640157.
pdf.

Xiao:2000:EPR

[2293] A. Xiao and S. Li. Error of partitioned Runge–Kutta methods for multiple stiff singular perturbation problems. Computing, 64(2):183–189, June 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/0064002/00640183.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/0064002/00640183.
pdf.

DeSchepper:2000:FEM

[2294] Hennie De Schepper. Finite element methods for eigenvalue problems on a rectangle with (semi- ) periodic boundary conditions on a pair of adjacent sides. Computing, 64(3):191–206, September 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/0064003/00640191.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/0064003/00640191.
pdf.

Ishiwata:2000:AOC

[2295] Emiko Ishiwata. On the attainable order of collocation methods for the neutral functional-differential equations with proportional delays. Computing, 64(3):207–222, September 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/0064003/00640207.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/0064003/00640207.
pdf.

Sauter:2000:VOP

[2296] Stefan Sauter. Variable order panel clustering. Computing, 64(3):223–261, September 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://link.

springer-ny.com/link/service/journals/
00607/bibs/0064003/00640223.htm;
http://link.springer-ny.com/link/
service/journals/00607/papers/0064003/
00640223.pdf.

Galligani:2000:SDE

[2297] Emanuele Galligani and Luca Zanni.

On the stability of the direct elimination method for equality constrained least squares problems. Computing, 64(3):263–277, September 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/0064003/00640263.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/0064003/00640263.
pdf.

Kosowski:2000:CUT

[2298] Przemyslaw Kosowski and Alicja Smoktunowicz. On constructing unit triangular matrices with prescribed singular values. Computing, 64(3):279–285, September 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/0064003/00640279.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/0064003/00640279.
pdf.

Anonymous:2000:P

[2299] Anonymous. Preface. Computing, 64(4):287, December 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://

link.springer-ny.com/link/service/
journals/00607/bibs/0064004/00640287.
htm;   http://link.springer-ny.
com/link/service/journals/00607/
papers/0064004/00640287.pdf.

Wieners:2000:RMM

[2300] Christian Wieners. Robust multigrid methods for nearly incompressible elasticity. Computing, 64(4):289–306, December 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/

bibs/0064004/00640289.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/0064004/00640289.
pdf.

John:2000:CMM

[2301] V. John and L. Tobiska. A coupled multigrid method for nonconforming finite element discretizations of the 2D-Stokes equation. Computing, 64(4):307–321, December 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/0064004/00640307.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/0064004/00640307.
pdf.

Starke:2000:GNM

[2302] Gerhard Starke. Gauss–Newton multilevel methods for least-squares finite element computations of variably saturated subsurface flow. Computing, 64(4):323–338, December 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/0064004/00640323.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/0064004/00640323.
pdf.

Henn:2000:MAM

[2303] Stefan Henn and Kristian Witsch.

A multigrid approach for minimizing a nonlinear functional for digital image matching. Computing, 64(4):339–348, December 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/0064004/00640339.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/0064004/00640339.
                                        pdf.

Shaidurov:2000:CMA

[2304] V. V. Shaidurov and G. Timmermann. A cascadic multigrid algorithm for semilinear indefinite elliptic problems. Computing, 64(4):349–366, December 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/0064004/00640349.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/0064004/00640349.
                                        pdf.

Tyrtyshnikov:2000:ICA

[2305] E. Tyrtyshnikov. Incomplete cross approximation in the mosaic-skeleton method. Computing, 64(4):367–380, December 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://link.

                                        springer-ny.com/link/service/journals/
                                        00607/bibs/0064004/00640367.htm;
                                         http://link.springer-ny.com/link/
                                        service/journals/00607/papers/0064004/
                                        00640367.pdf.

Kowarschik:2000:CAM

[2306] Markus Kowarschik, Ulrich Rüde, Christian Weiß, and Wolfgang Karl. Cache-aware multigrid methods for solving Poisson’s equation in two dimensions. Computing, 64(4):381–399, December 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://link.

springer-ny.com/link/service/journals/
00607/bibs/0064004/00640381.htm;
 http://link.springer-ny.com/link/
service/journals/00607/papers/0064004/
00640381.pdf.

He:2000:LUM

[2307] Y. He and X. Min. On-line uniform machine scheduling with rejection. Computing, 65(1):1–12, March 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/0065001/00650001.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/0065001/00650001.
pdf.

Bunke:2000:MCS

[2308] Horst Bunke, Xiaoyi Jiang, and Abraham Kandel. On the minimum common supergraph of two graphs. Computing, 65(1):13–25, March 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/0065001/00650013.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/0065001/00650013.
pdf.

Li:2000:GSS

[2309] Zi-Cai Li. Global superconvergence of simplified hybrid combinations for elliptic equations with singularities, I. basic theorem. Computing, 65(1):27– 44, March 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://link.

springer-ny.com/link/service/journals/
00607/bibs/0065001/00650027.htm;

                                         http://link.springer-ny.com/link/
                                        service/journals/00607/papers/0065001/
                                        00650027.pdf.

Jaruszewska-Walczak:2000:NMH

[2310] D. Jaruszewska-Walczak and Z. Kamont. Numerical methods for hyperbolic functional differential problems on the Haar pyramid. Computing, 65(1):45–72, March 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        journals/00607/bibs/0065001/00650045.
                                        htm;   http://link.springer-ny.
                                        com/link/service/journals/00607/
                                        papers/0065001/00650045.pdf.

Wagner:2000:ACM

[2311] C. Wagner. On the algebraic construction of multilevel transfer operators. Computing, 65(1):73–95, March 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/0065001/00650073.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/0065001/00650073.
                                        pdf.

Melenk:2000:SGD

[2312] J. M. Melenk, N. P. Kirchner, and C. Schwab. Spectral Galerkin discretization for hydrodynamic stability problems. Computing, 65(2):97–118, June 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/0065002/00650097.htm;   http:
                                        //link.springer-ny.com/link/service/

journals/00607/papers/0065002/00650097.
pdf.

Blum:2000:WEE

[2313] Heribert Blum and Franz-Theo Suttmeier.

Weighted error estimates for finite element solutions of variational inequalities. Computing, 65(2):119–134, June 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/0065002/00650119.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/0065002/00650119.
pdf.

Kraus:2000:CNA

[2314] J. K. Kraus and C. W. Brand.

Condition numbers of approximate Schur complements in two- and threedimensional discretizations on hierarchically ordered grids. Computing, 65(2):135–154, June 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://

link.springer-ny.com/link/service/
journals/00607/bibs/0065002/00650135.
htm;   http://link.springer-ny.
com/link/service/journals/00607/
papers/0065002/00650135.pdf.

Schwarz:2000:STO

[2315] Fritz Schwarz. Solving third order differential equations with maximal symmetry group. Computing, 65(2):155– 167, June 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://link.

springer-ny.com/link/service/journals/
00607/bibs/0065002/00650155.htm;
 http://link.springer-ny.com/link/

                                        service/journals/00607/papers/0065002/
                                        00650155.pdf.

Thiemard:2000:CBS

[2316] E. Thiémard. Computing bounds for the star discrepancy. Computing, 65(2):169–186, June 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        journals/00607/bibs/0065002/00650169.
                                        htm;   http://link.springer-ny.
                                        com/link/service/journals/00607/
                                        papers/0065002/00650169.pdf.

Leydold:2000:NTD

[2317] J. Leydold. A note on transformed density rejection. Computing, 65(2):187– 192, June 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://link.

                                        springer-ny.com/link/service/journals/
                                        00607/bibs/0065002/00650187.htm;
                                         http://link.springer-ny.com/link/
                                        service/journals/00607/papers/0065002/
                                        00650187.pdf.

Olshanskii:2000:CMM

[2318] Maxim A. Olshanskii and Arnold Reusken. On the convergence of a multigrid method for linear reaction-diffusion problems. Computing, 65(3):193–202, September 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/0065003/00650193.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/0065003/00650193.
                                        pdf.

Johannsen:2000:RSC

[2319] Klaus Johannsen. A robust smoother for convection-diffusion problems with closed characteristics. Computing, 65 (3):203–225, September 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://

link.springer-ny.com/link/service/
journals/00607/bibs/0065003/00650203.
htm;   http://link.springer-ny.
com/link/service/journals/00607/
papers/0065003/00650203.pdf.

Zulehner:2000:CSS

[2320] Walter Zulehner. A class of smoothers for saddle point problems. Computing, 65(3):227–246, September 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/0065003/00650227.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/0065003/00650227.
pdf.

Koster:2000:PCF

[2321] Frank Koster. A proof of the consistency of the finite difference technique on sparse grids. Computing, 65(3):247–261, September 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/0065003/00650247.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/0065003/00650247.
pdf.

Casado:2000:NMT

[2322] L. G. Casado, I. García, and T. Csendes. A new multisection technique in interval methods for global optimization. Computing, 65(3):263–269, September 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/0065003/00650263.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/0065003/00650263.
pdf.

Storck:2000:ANI

[2323] Ulrike Storck. An adaptive numerical integration algorithm with automatic result verification for definite integrals. Computing, 65(3):271–280, September 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/0065003/00650271.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/0065003/00650271.
pdf.

Malyshev:2000:SCN

[2324] A. N. Malyshev. Short communication: a note on the stability of Gauss–Jordan elimination for diagonally dominant matrices. Computing, 65(3):281–284, September 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/0065003/00650281.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/0065003/00650281.
pdf.

Mukhopadhyay:2000:SCO

[2325] A. Mukhopadhyay and S. V. Rao. Short communication: Output-sensitive algorithm for computing <beta>-skeletons. Computing, 65(3):285–289, September 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/0065003/00650285.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/0065003/00650285.
pdf.

Parker:2000:SAF

[2326] J. R. Parker. Simulated annealing for fitting linear combinations of Gaussians to data. Computing, 65(4):291–312, December 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/0065004/00650291.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/0065004/00650291.
pdf.

Pena:2000:MHS

[2327] J. M. Peña and Thomas Sauer.

On the multivariate Horner scheme II: Running error analysis. Computing, 65(4):313–322, December 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/0065004/00650313.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/0065004/00650313.
pdf.

Schmidt:2000:USI

[2328] Jochen W. Schmidt. Univariate strip interpolations by nonlinear parametric splines. Computing, 65(4):323–337, December 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/0065004/00650323.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/0065004/00650323.
                                        pdf.

Havik:2000:ASG

[2329] E. D. Havik, W. Hoffmann, and P. W. Hemker. Application of the over-set grid technique to a model singular perturbation problem. Computing, 65(4):339–356, December 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/0065004/00650339.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/0065004/00650339.
                                        pdf.

Hemker:2000:AAS

[2330] P. W. Hemker. Application of an adaptive sparse-grid technique to a model singular perturbation problem. Computing, 65(4):357–378, December 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/0065004/00650357.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/0065004/00650357.
                                        pdf.

Neuss:2001:HM

[2331] Nicolas Neuss, W. Jäger, and Gabriel Wittum. Homogenization and multigrid. Computing, 66(1):1–26, March 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/1066001/10660001.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/1066001/10660001.
pdf.

Linss:2001:UPC

[2332] Torsten Linß. Uniform pointwise convergence of finite difference schemes using grid equidistribution. Computing, 66(1):27–39, March 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/1066001/10660027.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/1066001/10660027.
pdf.

Li:2001:STG

[2333] Y. Y. Li, K. S. Leung, and C. K. Wong.

Steiner trees in general nonuniform orientations. Computing, 66(1):41–78, March 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/1066001/10660041.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/1066001/10660041.
pdf.

Guruswami:2001:KRP

[2334] Venkatesan Guruswami, C. Pandu Rangan, M. S. Chang, G. J. Chang, and C. K. Wong. The Kr-packing problem. Computing, 66(1):79–89, March 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/1066001/10660079.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/1066001/10660079.
                                        pdf.

Bahmann:2001:SCA

[2335] Helge Bahmann and Peter Schatte. Analysis of some elementary algorithms for prime factorization. Computing, 66(1):91–95, March 2001. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

                                        http://link.springer-ny.com/link/
                                        service/journals/00607/bibs/1066001/
                                        10660091.htm;   http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        papers/1066001/10660091.pdf.

Houston:2001:SHF

[2336] P. Houston and Endre Süli. Stabilised hp-finite element approximation of partial differential equations with nonnegative characteristic form. Computing, 66(2):99–119, June 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/1066002/10660099.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/1066002/10660099.
                                        pdf.

Canuto:2001:MSC

[2337] Claudio Canuto. Multilevel stabilization of convection-diffusion problems by variable-order inner products. Computing, 66(2):121–138, June 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/journals/00607/
bibs/1066002/10660121.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/1066002/10660121.
pdf.

Hemker:2001:HOT

[2338] Pieter W. Hemker, Grigorii I. Shishkin, and Lidia P. Shishkina. High-order time-accurate parallel schemes for parabolic singularly perturbed problems with convection. Computing, 66(2):139–161, June 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://

link.springer-ny.com/link/service/
journals/00607/bibs/1066002/10660139.
htm;   http://link.springer-ny.
com/link/service/journals/00607/
papers/1066002/10660139.pdf.

Roos:2001:GRS

[2339] Hans-Görg Roos and Torsten Linß.

Gradient recovery for singularly perturbed boundary value problems I: One-dimensional convectiondiffusion. Computing, 66(2):163–178, June 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/1066002/10660163.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/1066002/10660163.
pdf.

Kopteva:2001:UPC

[2340] Natalia Kopteva. Uniform pointwise convergence of difference schemes for convection-diffusion problems on layer-adapted meshes. Computing, 66(2):179–197, June 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        journals/00607/bibs/1066002/10660179.
                                        htm;   http://link.springer-ny.
                                        com/link/service/journals/00607/
                                        papers/1066002/10660179.pdf.

Ainsworth:2001:FSN

[2341] M. Ainsworth and W. Dörfler. Fundamental systems of numerical schemes for linear convection-diffusion equations and their relationship to accuracy. Computing, 66(2):199–229, June 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/1066002/10660199.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/1066002/10660199.
                                        pdf.

Hinterberger:2001:MII

[2342] W. Hinterberger and O. Scherzer. Models for image interpolation based on the optical flow. Computing, 66(3): 231–247, September 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        journals/00607/bibs/1066003/10660231.
                                        htm;   http://link.springer-ny.
                                        com/link/service/journals/00607/
                                        papers/1066003/10660231.pdf.

Wesseling:2001:UEN

[2343] P. Wesseling and D. R. van der Heul. Uniformly effective numerical methods for hyperbolic systems. Computing, 66(3):249–267, September 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/journals/00607/
bibs/1066003/10660249.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/1066003/10660249.
pdf.

John:2001:ALA

[2344] Volker John and William J. Layton. Approximating local averages of fluid velocities: The Stokes problem. Computing, 66(3):269–287, September 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/1066003/10660269.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/1066003/10660269.
pdf.

Imreh:2001:NWS

[2345] Cs. Imreh. A new well-solvable class of PNS problems. Computing, 66(3):289–296, September 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/1066003/10660289.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/1066003/10660289.
pdf.

Pallo:2001:RCL

[2346] Jean Marcel Pallo. The rotation < χ >-lattice of ternary trees. Computing, 66(3):297–308, September 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/1066003/10660297.htm;   http:
//link.springer-ny.com/link/service/

                                        journals/00607/papers/1066003/10660297.
                                        pdf.

Cuyt:2001:REC

[2347] Annie Cuyt, Brigitte Verdonk, S. Becuwe, and Peter Kuterna. A remarkable example of catastrophic cancellation unraveled. Computing, 66(3): 309–320, September 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        journals/00607/bibs/1066003/10660309.
                                        htm;   http://link.springer-ny.
                                        com/link/service/journals/00607/
                                        papers/1066003/10660309.pdf.

Borm:2001:TPM

[2348] Steffen Börm. Tensor product multigrid for Maxwell’s equation with aligned anisotropy. Computing, 66(4):321–342, December 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/1066004/10660321.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/1066004/10660321.
                                        pdf.

Matthies:2001:SDM

[2349] G. Matthies and L. Tobiska. The streamline-diffusion method for conforming and nonconforming finite elements of lowest order applied to convection-diffusion problems. Computing, 66(4):343–364, December 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/1066004/10660343.htm;   http:

//link.springer-ny.com/link/service/
journals/00607/papers/1066004/10660343.
pdf.

Xiao:2001:CRO

[2350] Aiguo Xiao, Chengming Huang, and Siqing Gan. Convergence results of one-leg and linear multistep methods for multiply stiff singular perturbation problems. Computing, 66(4):365–375, December 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/1066004/10660365.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/1066004/10660365.
pdf.

Heuberger:2001:MER

[2351] Clemens Heuberger and Helmut Prodinger. On minimal expansions in redundant number systems: Algorithms and quantitative analysis. Computing, 66(4):377–393, December 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/1066004/10660377.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/1066004/10660377.
pdf.

Alwan:2001:SDS

[2352] N. A. S. Alwan. Systematic design of systolic correlators with application to parallel Blackman-Tukey spectral estimation. Computing, 66(4):395–412, December 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/1066004/10660395.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/1066004/10660395.
                                        pdf.

Guglielmi:2001:IRI

[2353] Nicola Guglielmi and Ernst Hairer. Implementing Radau IIA methods for stiff delay differential equations. Computing, 67(1):1–12, March 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/1067001/10670001.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/1067001/10670001.
                                        pdf.

Stoyan:2001:ISS

[2354] Gisbert Stoyan. Iterative Stokes solvers in the harmonic Velte subspace. Computing, 67(1):13–33, March 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/1067001/10670013.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/1067001/10670013.
                                        pdf.

Hackbusch:2001:ECC

[2355] Wolfgang Hackbusch. The efficient computation of certain determinants arising in the treatment of Schrödinger’s equations. Computing, 67(1):35–56, March 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://

                                        link.springer-ny.com/link/service/

journals/00607/bibs/1067001/10670035.
htm;   http://link.springer-ny.
com/link/service/journals/00607/
papers/1067001/10670035.pdf.

Tran:2001:TLA

[2356] T. Tran and E. P. Stephan. Twolevel additive Schwarz preconditioners for the h-p version of the Galerkin boundary element method for 2-d problems. Computing, 67(1):57–82, March 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/1067001/10670057.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/1067001/10670057.
pdf.

Anderson:2001:TBP

[2357] E. J. Anderson, T. S. Jayram, and T. Kimbrel. Tighter bounds on preemptive job shop scheduling with two machines. Computing, 67(1):83–90, March 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/1067001/10670083.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/1067001/10670083.
pdf.

Otto:2001:ISM

[2358] F. C. Otto, G. Lube, and L. Müller.

An iterative substructuring method for div-stable finite element approximations of the Oseen problem. Computing, 67(2):91–117, June 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/journals/00607/
                                        bibs/1067002/10670091.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/1067002/10670091.
                                        pdf.

Hu:2001:SMC

[2359] Q. Hu and D. Yu. A solution method for a certain nonlinear interface problem in unbounded domains. Computing, 67(2):119–140, June 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/1067002/10670119.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/1067002/10670119.
                                        pdf.

Pflaum:2001:SUG

[2360] Christoph Pflaum. Semi-unstructured grids. Computing, 67(2):141–166, June 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/1067002/10670141.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/1067002/10670141.
                                        pdf.

Zhou:2001:RCF

[2361] Qing Zhou. Relatively computable functions of real variables. Computing, 67(2):167–181, June 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        journals/00607/bibs/1067002/10670167.
                                        htm;   http://link.springer-ny.
                                        com/link/service/journals/00607/
                                        papers/1067002/10670167.pdf.

Giebermann:2001:MAB

[2362] Klaus Giebermann. Multilevel approximation of boundary integral operators. Computing, 67(3):183–207, September 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/1067003/10670183.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/1067003/10670183.
pdf.

Bunke:2001:WMP

[2363] Horst Bunke and Simon Günter.

Weighted mean of a pair of graphs. Computing, 67(3):209–224, September 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/1067003/10670209.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/1067003/10670209.
pdf.

Garcke:2001:DMS

[2364] J. Garcke, M. Griebel, and M. Thess.

Data mining with sparse grids. Computing, 67(3):225–253, September 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/1067003/10670225.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/1067003/10670225.
pdf.

Neher:2001:MVF

[2365] M. Neher. The mean value form for complex analytic functions. Computing, 67(3):255–268, September 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/1067003/10670255.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/1067003/10670255.
pdf.

Bieniasz:2001:ETA

[2366] L. K. Bieniasz. Extension of the Thomas algorithm to a class of algebraic linear equation systems involving quasi-block-tridiagonal matrices with isolated block-pentadiagonal rows, assuming variable block dimensions. Computing, 67(4):269–285, December 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/1067004/10670269.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/1067004/10670269.

pdf. See erratum [2419].

Vulanovic:2001:HOS

[2367] Relja Vulanovic. A higher-order scheme for quasilinear boundary value problems with two small parameters. Computing, 67(4):287–303, December 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/1067004/10670287.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/1067004/10670287.
pdf.

Heuer:2001:HAT

[2368] N. Heuer, M. E. Mellado, and E. P. Stephan. hp-adaptive two-level methods for boundary integral equations on curves. Computing, 67(4):305–334, December 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/1067004/10670305.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/1067004/10670305.
pdf.

Laratta:2001:CLM

[2369] A. Laratta and F. Zironi. Computation of Lagrange multipliers for linear least squares problems with equality constraints. Computing, 67(4):335–350, December 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/1067004/10670335.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/1067004/10670335.
pdf.

Jung:2002:FEB

[2370] M. Jung and O. Steinbach. A finite element-boundary element algorithm for inhomogeneous boundary value problems. Computing, 68(1):1–17, March 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/2068001/20680001.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/2068001/20680001.
pdf.

Neuss:2002:NSM

[2371] N. Neuß. A new sparse-matrix storage method for adaptively solving large systems of reaction-diffusion-transport equations. Computing, 68(1):19–36, March 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/2068001/20680019.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/2068001/20680019.
                                        pdf.

Croisille:2002:KBS

[2372] J.-P. Croisille. Keller’s box-scheme for the one-dimensional stationary convection-diffusion equation. Computing, 68(1):37–63, March 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        journals/00607/bibs/2068001/20680037.
                                        htm;   http://link.springer-ny.
                                        com/link/service/journals/00607/
                                        papers/2068001/20680037.pdf.

Luo:2002:HAA

[2373] P. Luo and Q. Lin. High accuracy analysis of the Adini’s nonconforming element. Computing, 68(1):65–79, March 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/journals/00607/
                                        bibs/2068001/20680065.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        journals/00607/papers/2068001/20680065.
                                        pdf.

Kverh:2002:NRM

[2374] B. Kverh and A. Leonardis. A new refinement method for registration of range images based on segmented data. Computing, 68(1):81–96, March 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer-
ny.com/link/service/journals/00607/
bibs/2068001/20680081.htm;   http:
//link.springer-ny.com/link/service/
journals/00607/papers/2068001/20680081.
pdf.

Bajnok:2002:CFF

[2375] B. Bajnok, S. B. Damelin, J. Li, and G. L. Mullen. A constructive finite field method for scattering points on the surface of d-dimensional spheres. Computing, 68(2):97–109, June 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
de/link/service/journals/00607/
bibs/2068002/20680097.htm;   http:
//link.springer.de/link/service/
journals/00607/papers/2068002/20680097.
pdf.

Du:2002:DDM

[2376] Q. Du and D. Yu. A domain decomposition method based on natural boundary reduction for nonlinear time-dependent exterior wave problems. Computing, 68(2):111–129, June 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
de/link/service/journals/00607/
bibs/2068002/20680111.htm;   http:
//link.springer.de/link/service/
journals/00607/papers/2068002/20680111.
pdf.

Vanselow:2002:EEF

[2377] R. Vanselow. Error estimates of a FEM with lumping for parabolic PDEs. Computing, 68(2):131–141, June 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/journals/00607/
                                        bibs/2068002/20680131.htm;   http:
                                        //link.springer.de/link/service/
                                        journals/00607/papers/2068002/20680131.
                                        pdf.

Karaata:2002:DSS

[2378] M. H. Karaata and P. Chaudhuri. A dynamic self-stabilizing algorithm for constructing a transport net. Computing, 68(2):143–161, June 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/journals/00607/
                                        bibs/2068002/20680143.htm;   http:
                                        //link.springer.de/link/service/
                                        journals/00607/papers/2068002/20680143.
                                        pdf.

Werner:2002:SRI

[2379] T. Werner, T. Pajdla, V. Hlavác, A. Leonardis, and M. Matousek. Selection of reference images for imagebased scene representations. Computing, 68(2):163–180, June 2002. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

                                        http://link.springer.de/link/service/
                                        journals/00607/bibs/2068002/20680163.
                                        htm;   http://link.springer.de/
                                        link/service/journals/00607/papers/
                                        2068002/20680163.pdf.

Hoogeveen:2002:SCS

[2380] H. Hoogeveen and G. J. Woeginger. Some comments on sequencing with controllable processing times. Computing, 68(2):181–192, June 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
de/link/service/journals/00607/
bibs/2068002/20680181.htm;   http:
//link.springer.de/link/service/
journals/00607/papers/2068002/20680181.
pdf.

Hackbusch:2002:DIN

[2381] Wolfgang Hackbusch. Direct integration of the Newton potential over cubes. Computing, 68(3):193– 216, September 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.de/link/service/
journals/00607/bibs/2068003/20680193.
htm;   http://link.springer.de/
link/service/journals/00607/papers/
2068003/20680193.pdf.

Heinrich:2002:NTM

[2382] B. Heinrich and K. Pietsch. Nitsche type mortaring for some elliptic problem with corner singularities. Computing, 68(3):217–238, September 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
de/link/service/journals/00607/
bibs/2068003/20680217.htm;   http:
//link.springer.de/link/service/
journals/00607/papers/2068003/20680217.
pdf.

Ishihara:2002:DII

[2383] K. Ishihara. Descent iterations for improving approximate eigenpairs of polynomial eigenvalue problems with general complex matrices. Computing, 68(3):239–254, September 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.

                                        de/link/service/journals/00607/
                                        bibs/2068003/20680239.htm;   http:
                                        //link.springer.de/link/service/
                                        journals/00607/papers/2068003/20680239.
                                        pdf.

DelBuono:2002:RKT

[2384] N. Del Buono, L. Lopez, and C. Mastroserio. Runge Kutta type methods for isodynamical matrix flows: Applications to balanced realizations. Computing, 68(3):255–274, September 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/journals/00607/
                                        bibs/2068003/20680255.htm;   http:
                                        //link.springer.de/link/service/
                                        journals/00607/papers/2068003/20680255.
                                        pdf.

Kaufman:2002:AFP

[2385] E. H. Kaufman, G. D. Taylor, P. W. Mielke, and K. J. Berry. An algorithm and Fortran program for multivariate LAD ℓ (ℓ 1 of ℓ 2) regression. Computing, 68(3):275–287, September 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/journals/00607/
                                        bibs/2068003/20680275.htm;   http:
                                        //link.springer.de/link/service/
                                        journals/00607/papers/2068003/20680275.
                                        pdf.

Albrecher:2002:RTN

[2386] H. Albrecher and R. Kainhofer. Risk theory with a nonlinear dividend barrier. Computing, 68(4):289–311, December 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.

de/link/service/journals/00607/
bibs/2068004/20680289.htm;   http:
//link.springer.de/link/service/
journals/00607/papers/2068004/20680289.
pdf.

John:2002:NNM

[2387] V. John, P. Knobloch, G. Matthies, and L. Tobiska. Non-nested multilevel solvers for finite element discretisations of mixed problems. Computing, 68(4):313–341, December 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
de/link/service/journals/00607/
bibs/2068004/20680313.htm;   http:
//link.springer.de/link/service/
journals/00607/papers/2068004/20680313.
pdf.

Acosta:2002:ATS

[2388] G. Acosta, R. G. Durán, and J. D.

Rossi. An adaptive time step procedure for a parabolic problem with blowup. Computing, 68(4):343–373, December 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
de/link/service/journals/00607/
bibs/2068004/20680343.htm;   http:
//link.springer.de/link/service/
journals/00607/papers/2068004/20680343.
pdf.

Wu:2002:DMW

[2389] X.-Y. Wu, J.-L. Xia, and F. Yang.

A dynamic method for weighted linear least squares problems. Computing, 68(4):375–386, December 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.

                                        de/link/service/journals/00607/
                                        bibs/2068004/20680375.htm;   http:
                                        //link.springer.de/link/service/
                                        journals/00607/papers/2068004/20680375.
                                        pdf.

Hackbusch:2002:DSA

[2390] W. Hackbusch and S. Börm. Datasparse approximation by adaptive H^{2}- matrices. Computing, 69(1):1–35, March 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/journals/00607/
                                        bibs/2069001/20690001.htm;   http:
                                        //link.springer.de/link/service/
                                        journals/00607/papers/2069001/20690001.
                                        pdf.

Kis:2002:CNP

[2391] Tamás Kis. On the complexity of non-preemptive shop scheduling with two jobs. Computing, 69(1):37–49, March 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/journals/00607/
                                        bibs/2069001/20690037.htm;   http:
                                        //link.springer.de/link/service/
                                        journals/00607/papers/2069001/20690037.
                                        pdf.

Ceberio:2002:HRI

[2392] M. Ceberio and L. Granvilliers. Horner’s rule for interval evaluation revisited. Computing, 69(1):51–81, March 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/journals/00607/
                                        bibs/2069001/20690051.htm;   http:
                                        //link.springer.de/link/service/

journals/00607/papers/2069001/20690051.
pdf.

Hormann:2002:NPS

[2393] W. Hörmann. A note on the performance of the “Ahrens Algorithm”. Computing, 69(1):83–89, March 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
de/link/service/journals/00607/
bibs/2069001/20690083.htm;   http:
//link.springer.de/link/service/
journals/00607/papers/2069001/20690083.
pdf.

Khoromskij:2002:EDS

[2394] B. N. Khoromskij and J. M. Melenk. An efficient direct solver for the boundary concentrated FEM in 2D. Computing, 69(2):91–117, June 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
de/link/service/journals/00607/
bibs/2069002/20690091.htm;   http:
//link.springer.de/link/service/
journals/00607/papers/2069002/20690091.
pdf.

Matthies:2002:ISC

[2395] G. Matthies and L. Tobiska. The infsup condition for the mapped Q_{k} − P_{k−1}^{disc} element in arbitrary space dimensions. Computing, 69(2):119–139, June 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
de/link/service/journals/00607/
bibs/2069002/20690119.htm;   http:
//link.springer.de/link/service/
journals/00607/papers/2069002/20690119.
pdf.

Schwarz:2002:ECS

[2396] F. Schwarz. Equivalence classes, symmetries and solutions of linear third-order differential equations. Computing, 69(2):141–162, June 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/journals/00607/
                                        bibs/2069002/20690141.htm;   http:
                                        //link.springer.de/link/service/
                                        journals/00607/papers/2069002/20690141.
                                        pdf.

Ibragimov:2002:NSB

[2397] I. Ibragimov and S. Rjasanow. Numerical solution of the Boltzmann equation on the uniform grid. Computing, 69(2):163–186, June 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/journals/00607/
                                        bibs/2069002/20690163.htm;   http:
                                        //link.springer.de/link/service/
                                        journals/00607/papers/2069002/20690163.
                                        pdf.

Broersma:2002:MAS

[2398] Hajo Broersma, Fedor V. Fomin, Jaroslav Nešetřil, and Gerhard J. Woeginger. More about subcolorings. Computing, 69(3):187–203, November 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Braess:2002:SCM

[2399] D. Braess, P. Deuflhard, and K. Lipnikov. A subspace cascadic multigrid method for mortar elements. Computing, 69(3):205–225, November 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Liu:2002:SPP

[2400] G. Liu, D. Fu, and Z. Shen. On solving periodically perturbed nonconservative systems with a one-parameter embedding. Computing, 69(3):227–238, November 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Alberty:2002:MIF

[2401] J. Alberty, C. Carstensen, S. A.

Funken, and R. Klose. Matlab implementation of the finite element method in elasticity. Computing, 69 (3):239–263, November 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kunert:2002:NEN

[2402] Gerd Kunert. A note on the energy norm for a singularly perturbed model problem. Computing, 69(3): 265–272, November 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Baskiyar:2002:EEP

[2403] S. Baskiyar. Efficient execution of pure object-oriented programs by followup compilation. Computing, 69(4): 273–289, December 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Leobacher:2002:MAI

[2404] G. Leobacher and F. Pillichshammer.

A method for approximate inversion of the hyperbolic CDF. Computing, 69 (4):291–303, December 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

El-Daou:2002:CEB

[2405] M. K. El-Daou. Computable error bounds for coefficients perturbation methods. Computing, 69(4):305–317, December 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Oja:2002:CSC

[2406] Peeter Oja and Darja Saveljeva. Cubic spline collocation for Volterra integral equations. Computing, 69(4): 319–337, December 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pflaum:2002:PDG

[2407] C. Pflaum. Problem dependent generalized prewavelets. Computing, 69 (4):339–352, December 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bebendorf:2003:ALR

[2408] M. Bebendorf and S. Rjasanow. Adaptive low-rank approximation of collocation matrices. Computing, 70(1):1– 24, February 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ariyawansa:2003:NLS

[2409] K. A. Ariyawansa and Wayne L. Tabor. A note on line search termination criteria for collinear scaling algorithms. Computing, 70(1):25–39, February 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Paschos:2003:PAG

[2410] V. Th. Paschos. Polynomial approximation and graph-coloring. Computing, 70(1):41–86, February 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Frommer:2003:BFS

[2411] A. Frommer. BiCGStab(ℓ) for families of shifted linear systems. Computing, 70(2):87–109, April 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Garloff:2003:IIC

[2412] J. Garloff, C. Jansson, and A. P. Smith.

Inclusion isotonicity of convex-concave extensions for polynomials based on Bernstein expansion. Computing, 70 (2):111–119, April 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Grasedyck:2003:SLS

[2413] L. Grasedyck, W. Hackbusch, and B. N. Khoromskij. Solution of large scale algebraic matrix Riccati equations by use of hierarchical matrices. Computing, 70(2):121–165, April 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Cheng:2003:SSS

[2414] T. C. E. Cheng and M. Y. Kovalyov. Scheduling a single server in a two-machine flow shop. Computing, 70(2):167–180, April 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Haasdonk:2003:MVH

[2415] B. Haasdonk, M. Ohlberger, M. Rumpf, A. Schmidt, and K. G. Siebert. Multiresolution visualization of higher order adaptive finite element simulations. Computing, 70(3):181–204, June 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Repin:2003:PEE

[2416] S. Repin, S. Sauter, and A. Smolianski. A posteriori error estimation for the Dirichlet problem with account of the error in the approximation of boundary conditions. Computing, 70(3):205–233, June 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Alefeld:2003:IML

[2417] G. Alefeld and U. Schäfer. Iterative methods for linear complementarity problems with interval data. Computing, 70(3):235–259, June 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

LeBorne:2003:MCD

[2418] Sabine Le Borne. H-matrices for convection-diffusion problems with constant convection. Computing, 70 (3):261–274, June 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bieniasz:2003:E

[2419] L. K. Bieniasz. Erratum. Computing, 70(3):275, June 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [2366].

Imreh:2003:SPT

[2420] Csanád Imreh. Scheduling problems on two sets of identical machines. Computing, 70(4):277–294, August 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Grasedyck:2003:CAM

[2421] Lars Grasedyck and Wolfgang Hackbusch. Construction and arithmetics of H-matrices. Computing, 70(4):295– 334, August 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kravanja:2003:EAD

[2422] P. Kravanja, T. Sakurai, and M. Van Barel. Error analysis of a derivativefree algorithm for computing zeros of holomorphic functions. Computing, 70 (4):335–347, August 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Damm:2003:ETA

[2423] Michael Damm. On the existence of totally anti-symmetric quasigroups of order 4k + 2. Computing, 70(4):349–357, August 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Meyer:2003:SEJ

[2424] Arnd Meyer. Stable evaluation of Jacobian matrices on highly refined finite element meshes. Computing, 70 (4):359–373, August 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bungartz:2003:E

[2425] H.-J. Bungartz. Editorial. Computing, 71(1):??, August 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Achatz:2003:HOS

[2426] S. Achatz. Higher order sparse grid methods for elliptic partial differential equations with variable coefficients. Computing, 71(1):1–15, August 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gradinaru:2003:MDD

[2427] V. Gradinaru and R. Hiptmair. Multigrid for discrete differential forms on sparse grids. Computing, 71(1):17–42, August 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schwab:2003:SFE

[2428] Ch. Schwab and R. A. Todor. Sparse finite elements for stochastic elliptic problems – higher order moments. Computing, 71(1):43–63, August 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gerstner:2003:DAT

[2429] T. Gerstner and M. Griebel. Dimension– adaptive tensor–product quadrature. Computing, 71(1):65–87, August 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bungartz:2003:MQA

[2430] H.-J. Bungartz and S. Dirnstorfer. Multivariate quadrature on adaptive sparse grids. Computing, 71(1):89–114, August 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ahn:2003:FIL

[2431] Jaemin Ahn, Sungkwon Kang, and YongHoon Kwon. A flexible inverse Laplace transform algorithm and its application. Computing, 71(2):115– 131, October 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Novati:2003:EOS

[2432] P. Novati. An explicit one-step method for stiff problems. Computing, 71 (2):133–151, October 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Braianov:2003:NSR

[2433] I. Braianov and L. Vulkov. Numerical solution of a reaction-diffusion elliptic interface problem with strong anisotropy. Computing, 71(2):153–173, October 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Carstensen:2003:AFE

[2434] C. Carstensen and K. Jochimsen.

Adaptive finite element methods for microstructures? numerical experiments for a 2-well benchmark. Computing, 71(2):175–204, October 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Langer:2003:BET

[2435] U. Langer and O. Steinbach. Boundary element tearing and interconnecting methods. Computing, 71(3):205–228, November 2003. CODEN CMPTA2.

ISSN 0010-485X (print), 1436-5057 (electronic).

Brunner:2003:SSD

[2436] Hermann Brunner and Rossana Vermiglio. Stability of solutions of delay functional integro-differential equations and their discretizations. Computing, 71(3):229–245, November 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rui:2003:SMC

[2437] Hongxing Rui. Superconvergence of a mixed covolume method for elliptic problems. Computing, 71(3):247–263, November 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hiptmair:2003:NRB

[2438] R. Hiptmair and A. Schädle. Nonreflecting boundary conditions for Maxwell’s equations. Computing, 71 (3):265–292, November 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Apel:2003:THF

[2439] T. Apel and N. Düvelmeyer. Transformation of hexaedral finite element meshes into tetrahedral meshes according to quality criteria. Computing, 71 (4):293–304, November 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Diethelm:2003:ESM

[2440] K. Diethelm. Efficient solution of multi-term fractional differential equations using P(EC)^{m}E methods. Computing, 71(4):305–319, November 2003.

CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kumar:2003:PCB

[2441] B. V. Rathish Kumar, Alfio Quateroni, Luca Formaggia, et al. On parallel computation of blood flow in human arterial network based on 1-D modelling. Computing, 71(4):321–351, November 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

You:2003:GBS

[2442] L. H. You, Jian J. Zhang, and P. Comninos. Generating blending surfaces with a pseudo-Lévy series solution to fourth order partial differential equations. Computing, 71(4): 353–373, November 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mastrolilli:2003:NMF

[2443] Monaldo Mastrolilli. Notes on max flow time minimization with controllable processing times. Computing, 71 (4):375–386, November 2003. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hahmann:2004:EGM

[2444] Stefanie Hahmann, Guido Brunnett, Gerald Farin, and Ron Goldman. Editorial: Geometric modelling. Computing, 72(1–2):v, April 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Albrecht:2004:APQ

[2445] G. Albrecht. An algorithm for parametric quadric patch construction. Computing, 72(1–2):1–12, April 2004. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Azariadis:2004:PDF

[2446] P. N. Azariadis and N. S. Sapidis. Planar development of free-form surfaces: Quality evaluation and visual inspection. Computing, 72(1–2):13–27, April 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bertram:2004:BLS

[2447] M. Bertram. Biorthogonal loopsubdivision wavelets. Computing, 72 (1–2):29–39, April 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chalmoviansky:2004:FCA

[2448] Pavel Chalmovianský and Bert Jüttler. Fairness criteria for algebraic curves. Computing, 72(1–2):41–51, April 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Goldenthal:2004:SCA

[2449] Rony Goldenthal and Michel Bercovier. Spline curve approximation and design by optimal control over the knots. Computing, 72(1–2):53–64, April 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Grandine:2004:PQS

[2450] Thomas A. Grandine and Thomas A. Hogan. A parametric quartic spline interpolant to position, tangent and curvature. Computing, 72(1–2):65–78, April 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hong:2004:SCU

[2451] Don Hong and Larry L. Schumaker.

Surface compression using a space of C^{1} cubic splines with a hierarchical basis. Computing, 72(1–2):79–92, April 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ivrissimtzis:2004:EPS

[2452] I. P. Ivrissimtzis and H.-P. Seidel. Evolutions of polygons in the study of subdivision surfaces. Computing, 72 (1–2):93–103, April 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Johnstone:2004:BTS

[2453] John K. Johnstone. The Bézier tangential surface system: a robust dual representation of tangent space. Computing, 72(1–2):105–115, April 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Karavelas:2004:BDB

[2454] M. I. Karavelas, P. D. Kaklis, and K. V.

Kostas. Bounding the distance between 2D parametric Bézier curves and their control polygon. Computing, 72 (1–2):117–128, April 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Linsen:2004:WBM

[2455] Lars Linsen, Valerio Pascucci, Mark A.

Duchaineau, Bernd Hamann, and Kenneth I. Joy. Wavelet-based multiresolution with ^{n} \sqrt{2} subdivision. Computing, 72(1–2):129–142, April 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Manni:2004:QIT

[2456] Carla Manni and Francesca Pelosi. Quasi-interpolants with tension properties from and in CAGD. Computing, 72(1–2):143–160, April 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sauvage:2004:LPM

[2457] Basile Sauvage, Stefanie Hahmann, and Georges-Pierre Bonneau. Length preserving multiresolution editing of curves. Computing, 72(1–2):161–170, April 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Seong:2004:CHF

[2458] J.-K. Seong, G. Elber, J. K. Johnstone, and M.-S. Kim. The convex hull of freeform surfaces. Computing, 72(1–2):171–183, April 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sheffer:2004:RSP

[2459] A. Sheffer, C. Gotsman, and N. Dyn. Robust spherical parameterization of triangular meshes. Computing, 72 (1–2):185–193, April 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ugail:2004:SBS

[2460] Hassan Ugail. Spine based shape parameterisation for PDE surfaces. Computing, 72(1–2):195–206, April 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Vanco:2004:DSA

[2461] Marek Vanco and Guido Brunnett. Direct segmentation of algebraic models for reverse engineering. Computing, 72 (1–2):207–220, April 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Vigo:2004:TTM

[2462] Marc Vigo, Nuria Pla, and Dolors Ayala. Two triangulation methods based on edge refinement. Computing, 72 (1–2):221–233, April 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wang:2004:ECD

[2463] Wenping Wang, Yi-King Choi, Bin Chan, Myung-Soo Kim, and Jiaye Wang. Efficient collision detection for moving ellipsoids using separating planes. Computing, 72(1–2):235–246, April 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Grasedyck:2004:ECL

[2464] L. Grasedyck. Existence and computation of low Kronecker-rank approximations for large linear systems of tensor product structure. Computing, 72 (3–4):247–265, June 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tausch:2004:VOF

[2465] J. Tausch. The variable order fast multipole method for boundary integral equations of the second kind. Computing, 72(3–4):267–291, June 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lintner:2004:EPL

[2466] M. Lintner. The eigenvalue problem for the 2D Laplacian in H-matrix arithmetic and application to the heat and wave equation. Computing, 72 (3–4):293–323, June 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Borm:2004:LRA

[2467] Steffen Börm and Lars Grasedyck. Low-rank approximation of integral operators by interpolation. Computing, 72(3–4):325–332, June 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lamichhane:2004:MFE

[2468] B. P. Lamichhane and B. I. Wohlmuth. Mortar finite elements for interface problems. Computing, 72(3–4):333– 348, June 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Frommer:2004:CMM

[2469] A. Frommer, B. Lang, and M. Schnurr. A comparison of the Moore and Miranda existence tests. Computing, 72 (3–4):349–354, June 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dosa:2004:SOA

[2470] G. Dósa and Y. He. Semi-online algorithms for parallel machine scheduling problems. Computing, 72(3–4):355– 363, June 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pnar:2004:FCE

[2471] M. Ç. Pnar. Finite computation of the ℓ_{1} estimator from Huber’s M-estimator in linear regression. Computing, 72 (3–4):365–384, June 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Grasedyck:2004:AGB

[2472] L. Grasedyck, W. Hackbusch, and S. Le Borne. Adaptive geometrically balanced clustering of H-matrices. Computing, 73(1):1–23, July 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Petras:2004:NCI

[2473] K. Petras. Numerical computation of an integral representation for arithmetic-average Asian options. Computing, 73(1):25–39, July 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wang:2004:CAS

[2474] Wenqia Wang. A class of alternating segment Crank–Nicolson methods for solving convection-diffusion equations. Computing, 73(1):41–55, July 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gunji:2004:PPH

[2475] Takayuki Gunji, Sunyoung Kim, Masakazu Kojima, Akiko Takeda, Katsuki Fujisawa, and Tomohiko Mizutani. PHoM — a polyhedral homotopy continuation method for polynomial systems. Computing, 73(1):57–77, July 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Li:2004:LSJ

[2476] R. Li and H.-C. Huang. On-line scheduling for jobs with arbitrary release times. Computing, 73(1):79–97, July 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wien:2004:EMS

[2477] R. Wien, F. J. Gaspar, F. J. Lisbona, and C. W. Osterlee. An efficient multigrid solver based on distributive smoothing for poroelasticity equations. Computing, 73(2):99–119, September 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Linss:2004:ASS

[2478] T. Linß and N. Madden. Accurate solution of a system of coupled singularly perturbed reaction-diffusion equations. Computing, 73(2):121–133, September 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Walther:2004:BNP

[2479] A. Walther. Bounding the number of processors and checkpoints needed in time-minimal parallel reversal schedules. Computing, 73(2): 135–154, September 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rapin:2004:STF

[2480] G. Rapin and G. Lube. A stabilized three-field formulation for advectiondiffusion equations. Computing, 73(2): 155–178, September 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Grigoriev:2004:FSL

[2481] D. Grigoriev and F. Schwarz. Factoring and solving linear partial differential equations. Computing, 73(2): 179–197, September 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Suttmeier:2004:RAW

[2482] F. T. Suttmeier. Reliable approximation of weight factors entering residual-based error bounds for FEdiscretisations. Computing, 73(3):199– 205, October 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hackbusch:2004:HMB

[2483] W. Hackbusch, B. N. Khoromskij, and R. Kriemann. Hierarchical matrices based on a weak admissibility criterion. Computing, 73(3):207–243, October 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Imre:2004:GQD

[2484] S. Imre and F. Balázs. The generalized quantum database search algorithm. Computing, 73(3):245–269, October 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pedas:2004:STP

[2485] A. Pedas and G. Vainikko. Smoothing transformation and piecewise polynomial collocation for weakly singular Volterra integral equations. Computing, 73(3):271–293, October 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Xu:2004:LVD

[2486] Y. Xu, H.-L. Chen, and Q. Zou. Limit values of derivatives of the Cauchy integrals and computation of the logarithmic potentials. Computing, 73(4): 295–327, November 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kim:2004:NSP

[2487] S. Kim and M. Kojima. Numerical stability of path tracing in polyhedral homotopy continuation methods. Computing, 73(4):329–348, November 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Buchholz:2004:CMM

[2488] P. Buchholz and T. Dayar. Comparison of multilevel methods for Kronecker-based Markovian representations. Computing, 73(4):349–371, November 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Strzodka:2004:IRR

[2489] R. Strzodka, M. Droske, and M. Rumpf. Image registration by a regularized gradient flow. A streaming implementation in DX9 graphics hardware. Computing, 73(4):373–389, November 2004. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Burger:2005:RGA

[2490] M. Burger and A. Hofinger. Regularized greedy algorithms for network training with data noise. Computing, 74(1):1–22, February 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=74&issue=1&spage=1.

Fan:2005:QCL

[2491] Jin yan Fan and Ya xiang Yuan.

On the quadratic convergence of the Levenberg–Marquardt method without nonsingularity assumption. Computing, 74(1):23–39, February 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=74&issue=1&spage=23.

Verdonk:2005:BSI

[2492] B. Verdonk, J. Vervloet, and A. Cuyt.

Blending set and interval arithmetic for maximal reliability. Computing, 74(1):41–65, February 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=74&issue=1&spage=41.

Hu:2005:VRC

[2493] Hsin-Yun Hu and Zi-Cai Li. Verification of reduced convergence rates. Computing, 74(1):67–73, February 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0010-485X&volume=74&
issue=1&spage=67.

Borm:2005:HQS

[2494] S. Börm and W. Hackbusch. Hierarchical quadrature for singular integrals. Computing, 74(2):

75–100, March 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=74&issue=2&spage=75.

Hu:2005:NIU

[2495] Q. Hu. Numerical integrations and unit resolution multipliers for domain decomposition methods with nonmatching grids. Computing, 74 (2):101–129, March 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=74&issue=2&spage=101.

Gavrilyuk:2005:HTP

[2496] Ivan P. Gavrilyuk, Wolfgang Hackbusch, and Boris N. Khoromskij. Hierarchical tensor-product approximation to the inverse and related operators for high-dimensional elliptic problems. Computing, 74(2): 131–157, March 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=74&issue=2&spage=131.

Atkinson:2005:NSS

[2497] K. Atkinson and A. Sommariva. On the numerical solution of some semilinear elliptic problems: II. Computing, 74(2):159–175, March 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=74&issue=2&spage=159.

Anonymous:2005:P

[2498] Anonymous.  Preface.  Comput-

ing, 74(3):v–vi, May 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=74&issue=3&spage=v.

Popovic:2005:AMT

[2499] N. Popović and D. Praetorius. Applications of H-matrix techniques in micromagnetics. Computing, 74 (3):177–204, May 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=74&issue=3&spage=177.

Grasedyck:2005:ARM

[2500] L. Grasedyck. Adaptive recompression of H-matrices for BEM. Computing, 74(3):205–223, May 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=74&issue=3&spage=205.

Bebendorf:2005:HLD

[2501] M. Bebendorf. Hierarchical LU decomposition-based preconditioners for BEM. Computing, 74 (3):225–247, May 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=74&issue=3&spage=225.

Borm:2005:AIO

[2502] S. Börm. Approximation of integral operators by H^{∈}-matrices with adaptive bases. Computing, 74 (3):249–271, May 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=74&issue=3&spage=249.

Kriemann:2005:PMA

[2503] R. Kriemann. Parallel H-matrix arithmetics on shared memory systems. Computing, 74(3):273–297, May 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0010-485X&volume=74&
issue=3&spage=273.

Beuchler:2005:DDP

[2504] S. Beuchler. A domain decomposition preconditioner for p-FEM discretizations of two-dimensional elliptic problems. Computing, 74 (4):299–317, June 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=74&issue=4&spage=299.

Kraus:2005:CIW

[2505] J. K. Kraus. Computing interpolation weights in AMG based on multilevel Schur complements. Computing, 74(4):319–335, June 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0010-485X&
volume=74&issue=4&spage=319.

He:2005:TLS

[2506] Yinnian He and Kaitai Li. Twolevel stabilized finite element methods for the steady Navier–Stokes problem. Computing, 74(4):337– 351, June 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0010-485X&volume=74&
issue=4&spage=337.

Karkkainen:2005:DSI

[2507] T. Kärkkäinen, K. Kunisch, and K. Majava. Denoising of smooth images using L^{1}-fitting. Computing, 74(4):353–376, June 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=74&issue=4&spage=353.

Jager:2005:RSN

[2508] G. Jäger. Reduction of Smith normal form transformation matrices. Computing, 74(4):377–388, June 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0010-485X&volume=74&
issue=4&spage=377.

Anonymous:2005:ESI

[2509] Anonymous. Editorial: Special issue on GAMM-Workshop on Guaranteed Error-bounds for the Solution of Nonlinear Problems in Applied Mathematics. Computing, 75(1):1, July 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=75&issue=1&spage=1.

Nakao:2005:NMV

[2510] M. T. Nakao, K. Hashimoto, and Y. Watanabe. A numerical method to verify the invertibility of linear elliptic operators with applications to nonlinear problems. Computing, 75(1):1–14, July 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=75&issue=1&spage=1.

Murashige:2005:NVS

[2511] S. Murashige and S. Oishi. Numerical verification of solutions of Nekrasov’s integral equation. Computing, 75(1):15–25, July 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=75&issue=1&spage=15.

Hari:2005:ASB

[2512] V. Hari. Accelerating the SVD Block–Jacobi method. Computing, 75(1):27–53, July 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=75&issue=1&spage=27.

Petkovic:2005:WLM

[2513] L. D. Petković, M. S. Petković, and D. Milosević. Weierstrass-like methods with corrections for the inclusion of polynomial zeros. Computing, 75(1):55–69, July 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=75&issue=1&spage=55.

Petkovic:2005:DFI

[2514] M. S. Petković and D. Milosević.

Derivative free inclusion methods for polynomial zeros. Computing, 75(1):71–83, July 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=75&issue=1&spage=71.

Iliev:2005:SMS

[2515] A. Iliev and N. Kyurkchiev. Some methods for simultaneous extraction of a part of all multiple roots of algebraic polynomials. Computing, 75(1):85–97, July 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=75&issue=1&spage=85.

Ovseevich:2005:LBA

[2516] A. I. Ovseevich. Limit behavior of attainable sets of linear systems. Computing, 75(1):99–107, July 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0010-485X&volume=75&
                                        issue=1&spage=99.

Potthast:2005:P

[2517] Roland Potthast. Preface. Computing, 75(2–3):109–110, August 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=75&issue=2&spage=109.

Arens:2005:CFM

[2518] T. Arens and N. Grinberg. A complete factorization method for scattering by periodic surfaces. Computing, 75 (2–3):111–132, August 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=75&issue=2&spage=111.

Ikehata:2005:ITS

[2519] M. Ikehata. An inverse transmission scattering problem and the enclosure method. Computing, 75(2– 3):133–156, August 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=75&issue=2&spage=133.

Lines:2005:TDP

[2520] C. D. Lines and S. N. Chandler-Wilde. A time domain point source method for inverse scattering by rough surfaces. Computing, 75(2– 3):157–180, August 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=75&issue=2&spage=157.

Luke:2005:ISI

[2521] D. Russell Luke. Image synthesis for inverse obstacle scattering using the eigenfunction expansion theorem. Computing, 75(2– 3):181–196, August 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=75&issue=2&spage=181.

Nakamura:2005:NRC

[2522] G. Nakamura, S. Siltanen, K. Tanuma, and S. Wang. Numerical recovery of conductivity at the boundary from the localized Dirichlet to Neumann map. Computing, 75(2– 3):197–213, August 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=75&issue=2&spage=197.

Potthast:2005:SPM

[2523] Roland Potthast. Sampling and probe methods — an algorithmical view. Computing, 75(2–3): 215–235, August 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=75&issue=2&spage=215.

Potthast:2005:SSM

[2524] R. Potthast and I. Stratis. The singular sources method for an inverse transmission problem. Computing, 75 (2–3):237–255, August 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=75&issue=2&spage=237.

Heinrich:2005:NTM

[2525] B. Heinrich and K. Pönitz. Nitsche type mortaring for singularly perturbed reaction-diffusion problems. Computing, 75(4):257–279, August 2005. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=75&issue=4&spage=257.

Entacher:2005:BLP

[2526] K. Entacher, T. Schell, and A. Uhl. Bad lattice points. Computing, 75 (4):281–295, August 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=75&issue=4&spage=281. See

erratum [2558].

Sun:2005:NCF

[2527] Weiwei Sun and Jiming Wu. Newton– Cotes formulae for the numerical evaluation of certain hypersingular integrals. Computing, 75(4): 297–309, August 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0010-485X&
volume=75&issue=4&spage=297.

Uzer:2005:CAS

[2528] A. Uzer and T. Ege. On the convergence acceleration of slowly convergent sums involving oscillating terms. Computing, 75(4): 311–318, August 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=75&issue=4&spage=311.

Zhou:2005:DA

[2529] Qing Zhou and Weihao Hu. Decidability in analysis. Computing, 75 (4):319–336, August 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=75&issue=4&spage=319.

Zhu:2005:LSI

[2530] Y.-K. Zhu, J.-H. Yong, and G.- Q. Zheng. Line segment intersection testing. Computing, 75(4): 337–357, August 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=75&issue=4&spage=337.

Gorban:2005:EPG

[2531] A. Gorban and A. Zinovyev. Elastic principal graphs and manifolds and their practical applications. Computing, 75(4):359–379, August 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0010-485X&volume=75&
                                        issue=4&spage=359.

Allgower:2006:SPS

[2532] E. L. Allgower, D. J. Bates, A. J. Sommese, and C. W. Wampler. Solution of polynomial systems derived from differential equations. Computing, 76(1–2):1–10, January 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=76&issue=1&spage=1.

Nkemzi:2006:FFE

[2533] B. Nkemzi. The Fourier finiteelement approximation of the Lamé equations in axisymmetric domains with edges. Computing, 76(1–2): 11–39, January 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=76&issue=1&spage=11.

Knobloch:2006:ISC

[2534] P. Knobloch. On the inf–sup condition for the P_{3}^{mod}/P_{2}^{disc} element. Computing, 76(1–2):41–54, January 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0010-485X&volume=76&
                                        issue=1&spage=41.

Erbay:2006:AAS

[2535] H. Erbay and J. Barlow. An alternative algorithm for a sliding window ULV decomposition. Computing, 76 (1–2):55–66, January 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=76&issue=1&spage=55.

Cheng:2006:ESI

[2536] Sui Sun Cheng, Sukrawan Talwong, and Vichian Laohakosol. Exact solutions of iterative functional differential equations. Computing, 76(1– 2):67–76, January 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=76&issue=1&spage=67.

Gantumur:2006:CSI

[2537] T. Gantumur and R. P. Stevenson. Computation of singular integral operators in wavelet coordinates. Computing, 76(1–2): 77–107, January 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=76&issue=1&spage=77.

Hinterberger:2006:VMS

[2538] W. Hinterberger and O. Scherzer. Variational methods on the space of functions of bounded Hessian for convexification and denoising. Computing, 76 (1–2):109–133, January 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=76&issue=1&spage=109.

Steidl:2006:NDT

[2539] G. Steidl. A note on the dual treatment of higher-order regularization functionals. Computing, 76(1– 2):135–148, January 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=76&issue=1&spage=135.

Dosa:2006:PNP

[2540] G. Dósa and Y. He. Preemptive and non-preemptive on-line algorithms for scheduling with rejection on two uniform machines. Computing, 76(1– 2):149–164, January 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=76&issue=1&spage=149.

Aichholzer:2006:CNC

[2541] O. Aichholzer, F. Aurenhammer, and H. Krasser. On the crossing number of complete graphs. Computing, 76(1– 2):165–176, January 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=76&issue=1&spage=165.

Hackbusch:2006:LRKa

[2542] W. Hackbusch and B. N. Khoromskij. Low-rank Kronecker-product approximation to multi-dimensional nonlocal operators. Part I. Separable approximation of multi-variate functions. Computing, 76(3–4): 177–202, January 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=76&issue=3&spage=177.

Hackbusch:2006:LRKb

[2543] W. Hackbusch and B. N. Khoromskij.

Low-rank Kronecker-product approximation to multi-dimensional nonlocal operators. Part II. HKT representation of certain operators. Computing, 76 (3–4):203–225, January 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=76&issue=3&spage=203.

Christara:2006:OQC

[2544] C. C. Christara and Kit Sun Ng. Optimal quadratic and cubic spline collocation on nonuniform partitions. Computing, 76(3– 4):227–257, January 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=76&issue=3&spage=227.

Christara:2006:ATS

[2545] C. C. Christara and Kit Sun Ng.

Adaptive techniques for spline collocation. Computing, 76(3–4): 259–277, January 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=76&issue=3&spage=259.

Klein:2006:GKB

[2546] A. Klein. A generalized Kahan– Babuška-summation–algorithm. Computing, 76(3–4):279–293, January 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=76&issue=3&spage=279.

Sommariva:2006:NCS

[2547] A. Sommariva and M. Vianello. Numerical cubature on scattered data by radial basis functions. Computing, 76 (3–4):295–310, January 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=76&issue=3&spage=295.

Bos:2006:NSX

[2548] L. Bos, M. Caliari, S. De Marchi, and M. Vianello. A numerical study of the Xu polynomial interpolation formula in two variables. Computing, 76 (3–4):311–324, January 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=76&issue=3&spage=311.

Groisman:2006:TDE

[2549] Pablo Groisman. Totally discrete explicit and semi-implicit Euler methods for a blow-up problem in several space dimensions. Computing, 76(3– 4):325–352, January 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0010-485X&
volume=76&issue=3&spage=325.

Romanovski:2006:FLM

[2550] I. V. Romanovski and L. A. Pilipchuk.

Fractal-like matrices. Computing, 76 (3–4):353–357, January 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=76&issue=3&spage=353.

Hackbusch:2006:AXE

[2551] W. Hackbusch. Approximation of 1/||x − y|| by exponentials for wavelet applications (short communication). Computing, 76(3–4): 359–366, January 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=76&issue=3&spage=359.

Borm:2006:MAL

[2552] S. Börm. H^{2}-matrix arithmetics in linear complexity. Computing, 77 (1):1–28, February 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=77&issue=1&spage=1.

Sauter:2006:CFE

[2553] S. A. Sauter and R. Warnke. Composite finite elements for elliptic boundary value problems with discontinuous coefficients. Computing, 77 (1):29–55, February 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=77&issue=1&spage=29.

Kraus:2006:AMB

[2554] J. K. Kraus and J. Schicho. Algebraic multigrid based on computational molecules, 1: Scalar elliptic problems. Computing, 77(1): 57–75, February 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=77&issue=1&spage=57.

Baratella:2006:NSW

[2555] P. Baratella and A. Palamara Orsi. Numerical solution of weakly singular linear Volterra integro-differential equations. Computing, 77(1):77–96, February 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0010-485X&volume=77&
issue=1&spage=77.

Mainar:2006:REA

[2556] E. Mainar and J. M. Peña. Running error analysis of evaluation algorithms for bivariate polynomials in barycentric Bernstein form. Computing, 77 (1):97–111, February 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=77&issue=1&spage=97.

Rocadas:2006:EFD

[2557] L. Roçadas and J. Schoißengeier. An explicit formula for the L_{2}-discrepancy of (nα)-sequences. Computing, 77(1): 113–128, February 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=77&issue=1&spage=113.

Anonymous:2006:E

[2558] Anonymous. Erratum. Computing, 77(1):129, February 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=77&issue=1&spage=129. See

[2526].

Palubeckis:2006:TLB

[2559] G. Palubeckis. A tight lower bound for a special case of quadratic 0-1 programming. Computing, 77 (2):131–145, April 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=77&issue=2&spage=131.

Markot:2006:RAR

[2560] M. C. Markót and T. Csendes.

A reliable area reduction technique for solving circle packing problems. Computing, 77(2):147–162, April 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0010-485X&volume=77&
issue=2&spage=147.

Behloul:2006:CAP

[2561] D. Behloul and S. S. Cheng. Computation of all polynomial solutions of a class of nonlinear differential equations. Computing, 77 (2):163–177, April 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=77&issue=2&spage=163.

Salem:2006:NSG

[2562] Fatima K. Abu Salem. A new sparse Gaussian elimination algorithm and the Niederreiter linear system for trinomials over F_{2}. Computing, 77 (2):179–203, April 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=77&issue=2&spage=179.

Guo:2006:HGM

[2563] L. Guo and H. Chen. H^{1}-Galerkin mixed finite element method for the regularized long wave equation. Computing, 77(2):205–221, April 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0010-485X&volume=77&
                                        issue=2&spage=205.

Ali:2006:PVH

[2564] I. Ali, S. Amat, and J. C. Trillo. Point values Hermite multiresolution for non-smooth noisy signals. Computing, 77(3):223–236, May 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=77&issue=3&spage=223.

Chan:2006:SAO

[2565] T. M. H. Chan and R. P. K. Chan. A simplified approach to the order conditions of integration methods. Computing, 77(3):237–252, May 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=77&issue=3&spage=237.

Hackbusch:2006:CBE

[2566] W. Hackbusch, M. Löhndorf, and S. A. Sauter. Coarsening of boundaryelement spaces. Computing, 77 (3):253–273, May 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=77&issue=3&spage=253.

Carpentieri:2006:MFT

[2567] B. Carpentieri. A matrix-free two-grid preconditioner for solving boundary integral equations in electromagnetism. Computing, 77 (3):275–296, May 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=77&issue=3&spage=275.

Huang:2006:FFV

[2568] C.-S. Huang, C.-H. Hung, and S. Wang.

A fitted finite volume method for the valuation of options on assets with stochastic volatilities. Computing, 77(3):297–320, May 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=77&issue=3&spage=297.

Douglas:2006:DDD

[2569] C. Douglas, Y. Efendiev, R. Ewing, V. Ginting, and R. Lazarov. Dynamic data driven simulations in stochastic environments. Computing, 77(4):321–333, June 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=77&issue=4&spage=321.

Budd:2006:CSS

[2570] C. Budd, O. Koch, and E. Weinmüller. Computation of self-similar solution profiles for the nonlinear Schrödinger equation. Computing, 77(4):335– 346, June 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0010-485X&volume=77&
issue=4&spage=335.

Shu:2006:AMM

[2571] S. Shu, D. Sun, and J. Xu. An algebraic multigrid method for higher-order finite element discretizations. Computing, 77(4):347–377, June 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0010-485X&volume=77&
issue=4&spage=347.

Sun:2006:IPC

[2572] Yongli Sun and Jianping Yu. Implicitization of parametric curves via Lagrange interpolation. Computing, 77(4):379–386, June 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=77&issue=4&spage=379.

Gunji:2006:PPI

[2573] T. Gunji, S. Kim, K. Fujisawa, and M. Kojima. PHoMpara — parallel implementation of the Polyhedral Homotopy Continuation Method for polynomial systems. Computing, 77 (4):387–411, June 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=77&issue=4&spage=387.

Keiner:2006:FSR

[2574] J. Keiner, S. Kunis, and D. Potts.

Fast summation of radial functions on the sphere. Computing, 78 (1):1–15, August 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=78&issue=1&spage=1.

Boglaev:2006:MIS

[2575] I. Boglaev. Monotone iterates for solving nonlinear monotone difference schemes. Computing, 78(1): 17–30, August 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=78&issue=1&spage=17.

Chen:2006:CSC

[2576] Z. Chen, G. Huan, and H. Wang. Computer simulation of compositional flow using unstructured control volume finite element methods. Computing, 78(1):31–53, August 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=78&issue=1&spage=31.

Albrecht:2006:SCL

[2577] Andreas A. Albrecht. A stopping criterion for logarithmic simulated annealing. Computing, 78 (1):55–79, August 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=78&issue=1&spage=55.

Chen:2006:CIE

[2578] Chin-Yun Chen. Computing interval enclosures for definite integrals by application of triple adaptive strategies. Computing, 78 (1):81–99, August 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=78&issue=1&spage=81. See

erratum [3152].

Sauter:2006:RFE

[2579] S. A. Sauter. A refined finite element convergence theory for highly indefinite Helmholtz problems. Computing, 78 (2):101–115, October 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=78&issue=2&spage=101.

Tomar:2006:SEM

[2580] S. K. Tomar. h-p spectral element method for elliptic problems on non-smooth domains using parallel computers. Computing, 78(2): 117–143, October 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=78&issue=2&spage=117.

Hackbusch:2006:EEC

[2581] W. Hackbusch. On the efficient evaluation of coalescence integrals in population balance models. Computing, 78 (2):145–159, October 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=78&issue=2&spage=145.

Bahig:2006:IGM

[2582] Hatem M. Bahig. Improved generation of minimal addition chains. Computing, 78(2):161–172, October 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=78&issue=2&spage=161.

Sezgin:2006:DLP

[2583] F. Sezgin. Distribution of lattice points. Computing, 78(2): 173–193, October 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=78&issue=2&spage=173.

Yserentant:2006:SGA

[2584] H. Yserentant. Sparse grids, adaptivity, and symmetry. Computing, 78 (3):195–209, November 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=78&issue=3&spage=195.

Baur:2006:FSL

[2585] U. Baur and P. Benner. Factorized solution of Lyapunov equations based on hierarchical matrix arithmetic. Computing, 78(3):211– 234, November 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=78&issue=3&spage=211.

Becirovic:2006:HSA

[2586] A. Bećirović, P. Paule, V. Pillwein, A. Riese, C. Schneider, and J. Schöberl. Hypergeometric summation algorithms for high-order finite elements. Computing, 78(3): 235–249, November 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0010-485X&
volume=78&issue=3&spage=235.

Pryce:2006:IAC

[2587] J. D. Pryce and G. F. Corliss.

Interval arithmetic with containment sets. Computing, 78(3):251– 276, November 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=78&issue=3&spage=251.

Povh:2006:BPM

[2588] J. Povh, F. Rendl, and A. Wiegele.

A boundary point method to solve semidefinite programs. Computing, 78 (3):277–286, November 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=78&issue=3&spage=277.

delosReyes:2006:SSN

[2589] J. C. de los Reyes and K. Kunisch. A semi-smooth Newton method for regularized state-constrained optimal control of the Navier–Stokes equations. Computing, 78(4):287– 309, December 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=78&issue=4&spage=287.

Dostal:2006:OAB

[2590] Z. Dostál. An optimal algorithm for bound and equality constrained quadratic programming problems with bounded spectrum. Computing, 78(4):

311–328, December 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=78&issue=4&spage=311.

Croisille:2006:HBS

[2591] J.-P. Croisille. A Hermitian boxscheme for one-dimensional elliptic equations — application to problems with high contrasts in the ellipticity. Computing, 78(4):329– 353, December 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=78&issue=4&spage=329.

Akritas:2006:INT

[2592] A. Akritas, A. Strzeboński, and P. Vigklas. Implementations of a new theorem for computing bounds for positive roots of polynomials. Computing, 78(4):355–367, December 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0010-485X&volume=78&
issue=4&spage=355.

Tan:2006:SOM

[2593] Z. Y. Tan and S. J. Cao. Semionline machine covering on two uniform machines with known total size. Computing, 78(4):369– 378, December 2006. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0010-485X&
volume=78&issue=4&spage=369.

Baran:2007:GLE

[2594] Á. Baran and G. Stoyan. Gauss– Legendre elements: a stable, higher order non-conforming finite element family. Computing, 79(1): 1–21, February 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=79&issue=1&spage=1.

Linss:2007:AUF

[2595] T. Linß. Analysis of an upwind finitedifference scheme for a system of coupled singularly perturbed convectiondiffusion equations. Computing, 79 (1):23–32, February 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=79&issue=1&spage=23.

Smoktunowicz:2007:NEA

[2596] A. Smoktunowicz, I. Wróbel, and P. Kosowski. A new efficient algorithm for polynomial interpolation. Computing, 79(1):33–52, February 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0010-485X&volume=79&
issue=1&spage=33.

Goldsztejn:2007:CHS

[2597] A. Goldsztejn. Comparison of the Hansen–Sengupta and the Frommer– Lang-schnurr existence tests. Computing, 79(1):53–60, February 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=79&issue=1&spage=53.

Wang:2007:VES

[2598] Z. Wang. Validation and enclosure of solutions of linear complementarity problems. Computing, 79 (1):61–77, February 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=79&issue=1&spage=61.

Dick:2007:CGE

[2599] J. Dick, F. Pillichshammer, and B. J. Waterhouse. The construction of good extensible Korobov rules. Computing, 79(1):79–91, February 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=79&issue=1&spage=79.

Eisenstat:2007:ASC

[2600] S. C. Eisenstat. O(log ∗n) algorithms on a Sum-CRCW PRAM. Computing, 79(1):93–97, February 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=79&issue=1&spage=93.

Hahmann:2007:ESI

[2601] Stefanie Hahmann, Guido Brunnett, Gerald Farin, and Ron Goldman. Editorial: Special issue on Geometric Modeling (Dagstuhl 2005). Computing, 79(2–4):99, April 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=79&issue=2&spage=99.

Ayala:2007:SRP

[2602] D. Ayala, N. Pla, and M. Vigo.

Splat representation of parametric surfaces. Computing, 79(2– 4):101–108, April 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=79&issue=2&spage=101.

Xidias:2007:TDM

[2603] E. K. Xidias, P. N. Azariadis, and N. A. Aspragathos. Twodimensional motion-planning for nonholonomic robots using the bumpsurfaces concept. Computing, 79(2– 4):109–118, April 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=79&issue=2&spage=109.

Keller:2007:RES

[2604] P. Keller, M. Bertram, and H. Hagen. Reverse engineering with subdivision surfaces. Computing, 79(2– 4):119–129, April 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=79&issue=2&spage=119.

Chen:2007:IPQ

[2605] F. Chen, L. Shen, and J. Deng. Implicitization and parametrization of quadratic and cubic surfaces by μ-bases. Computing, 79(2–4): 131–142, April 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=79&issue=2&spage=131.

Degen:2007:SEB

[2606] W. L. F. Degen. Sharp error bounds for piecewise linear interpolation of planar curves. Computing, 79(2– 4):143–151, April 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=79&issue=2&spage=143.

Vouga:2007:TBP

[2607] E. Vouga and R. Goldman. Two blossoming proofs of the Lane– Riesenfeld algorithm. Computing, 79 (2–4):153–162, April 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=79&issue=2&spage=153.

Grandine:2007:CPR

[2608] T. A. Grandine and T. A. Hogan. C^{2} pseudo rolling ball filleting. Computing, 79(2–4):163–173, April 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=79&issue=2&spage=163.

Gravesen:2007:SPN

[2609] J. Gravesen. Surfaces parametrized by the normals. Computing, 79(2– 4):175–183, April 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=79&issue=2&spage=175.

Hagen:2007:FVM

[2610] H. Hagen, M. Schneider, M. Schlemmer, M. Ruby, and I. Scheler. Fast Voronoi modeling. Computing, 79 (2–4):185–196, April 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=79&issue=2&spage=185.

Gradinaru:2007:FTS

[2611] V. Gradinaru. Fourier transform on sparse grids: Code design and the time dependent Schrödinger equation. Computing, 80(1):1–22, May 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=80&issue=1&spage=1.

Lundy:2007:NMA

[2612] T. Lundy and J. Van Buskirk. A new matrix approach to real FFTs and convolutions of length 2^{k}. Computing, 80(1):23–45, May 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=80&issue=1&spage=23.

Fedorov:2007:SLR

[2613] M. V. Fedorov, H.-J. Flad, G. N. Chuev, L. Grasedyck, and B. N. Khoromskij. A structured lowrank wavelet solver for the Ornstein– Zernike integral equation. Computing, 80(1):47–73, May 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=80&issue=1&spage=47.

Vandewoestyne:2007:OQC

[2614] B. Vandewoestyne, R. Cools, and T. Warnock. On obtaining quadratic and cubic error convergence using weighted Kronecker-sequences. Computing, 80(1):75–94, May 2007. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=80&issue=1&spage=75.

Matthies:2007:RMQ

[2615] G. Matthies and F. Schieweck. On the reference mapping for quadrilateral and hexahedral finite elements on multilevel adaptive grids. Computing, 80(2):95–119, June 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=80&issue=2&spage=95.

Liu:2007:EAC

[2616] J. Liu, H. Chen, R. Ewing, and G. Qin. An efficient algorithm for characteristic tracking on two-dimensional triangular meshes. Computing, 80 (2):121–136, June 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=80&issue=2&spage=121.

Hackbusch:2007:FEP

[2617] W. Hackbusch. Fast and exact projected convolution for nonequidistant grids. Computing, 80 (2):137–168, June 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=80&issue=2&spage=137.

Oliveira:2007:AAH

[2618] S. Oliveira and F. Yang. An algebraic approach for H-matrix preconditioners. Computing, 80(2): 169–188, June 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=80&issue=2&spage=169.

Caliari:2007:AED

[2619] M. Caliari. Accurate evaluation of divided differences for polynomial interpolation of exponential propagators. Computing, 80(2): 189–201, June 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=80&issue=2&spage=189.

Heuveline:2007:IQH

[2620] V. Heuveline and F. Schieweck. H^{1}-interpolation on quadrilateral and hexahedral meshes with hanging nodes. Computing, 80(3): 203–220, July 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=80&issue=3&spage=203.

Heinrich:2007:FNM

[2621] B. Heinrich and B. Jung. The Fourier– Nitsche-mortaring for elliptic problems with reentrant edges. Computing, 80(3):221–246, July 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=80&issue=3&spage=221.

Nataraj:2007:AZM

[2622] P. S. V. Nataraj and N. Kubal. On the Asaithambi–Zuhe–Moore algorithm for computing the range of values. Computing, 80(3): 247–253, July 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=80&issue=3&spage=247.

Pagourtzis:2007:PMF

[2623] A. Pagourtzis, K. Potika, and S. Zachos. Path multicoloring with fewer colors in spiders and caterpillars. Computing, 80(3):255–274, July 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=80&issue=3&spage=255.

Epstein:2007:CLB

[2624] L. Epstein and R. van Stee. Calculating lower bounds for caching problems. Computing, 80(3):275– 285, July 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0010-485X&volume=80&
issue=3&spage=275.

Ovall:2007:HMT

[2625] J. S. Ovall. Hierarchical matrix techniques for a domain decomposition algorithm. Computing, 80(4): 287–297, September 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=80&issue=4&spage=287.

Reusken:2007:RMM

[2626] A. Reusken and M. Soemers. On the robustness of a multigrid method for anisotropic reaction-diffusion problems. Computing, 80(4):299–317, September 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0010-485X&volume=80&
issue=4&spage=299.

Scheichl:2007:ASA

[2627] R. Scheichl and E. Vainikko. Additive Schwarz with aggregation-based coarsening for elliptic problems with highly variable coefficients. Computing, 80(4):

319–343, September 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=80&issue=4&spage=319.

Beuchler:2007:SSF

[2628] S. Beuchler and V. Pillwein. Sparse shape functions for tetrahedral p-FEM using integrated Jacobi polynomials. Computing, 80(4):345– 375, September 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=80&issue=4&spage=345.

Hatzl:2007:MPW

[2629] J. Hatzl. Median problems on wheels and cactus graphs. Computing, 80(4): 377–393, September 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=80&issue=4&spage=377.

Griebel:2007:SGS

[2630] M. Griebel and D. Oeltz. A sparse grid space-time discretization scheme for parabolic problems. Computing, 81(1):1–34, October 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=81&issue=1&spage=1.

Hofinger:2007:NST

[2631] A. Hofinger and J. Valdman. Numerical solution of the two-yield elastoplastic minimization problem. Computing, 81(1):35–52, October 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=81&issue=1&spage=35.

Vey:2007:AFD

[2632] S. Vey and A. Voigt. Adaptive full domain covering meshes for parallel finite element computations. Computing, 81(1):53–75, October 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=81&issue=1&spage=53.

Bieniasz:2007:SCF

[2633] L. K. Bieniasz. A set of compact finitedifference approximations to first and second derivatives, related to the extended Numerov method of Chawla on nonuniform grids. Computing, 81 (1):77–89, October 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=81&issue=1&spage=77.

Beyn:2007:NFG

[2634] W.-J. Beyn and J. Rieger. Numerical fixed grid methods for differential inclusions. Computing, 81 (1):91–106, October 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=81&issue=1&spage=91.

Juttler:2007:VPL

[2635] B. Jüttler, H. Pottmann, and O. Scherzer. Variational and PDE level set methods. Computing, 81(2– 3):107–108, November 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=81&issue=2&spage=107.

Burger:2007:EEB

[2636] M. Burger, E. Resmerita, and L. He. Error estimation for Bregman iterations and inverse scale space methods in image restoration. Computing, 81(2– 3):109–135, November 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=81&issue=2&spage=109.

Schnorr:2007:SIA

[2637] C. Schnörr. Signal and image approximation with level-set constraints. Computing, 81(2–3):137– 160, November 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=81&issue=2&spage=137.

Steidl:2007:RMF

[2638] G. Steidl, S. Setzer, B. Popilka, and B. Burgeth. Restoration of matrix fields by second-order cone programming. Computing, 81(2–3): 161–178, November 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=81&issue=2&spage=161.

Burgeth:2007:GAD

[2639] B. Burgeth, S. Didas, L. Florack, and J. Weickert. A generic approach to diffusion filtering of matrixfields. Computing, 81(2–3):179– 197, November 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=81&issue=2&spage=179.

Fisher:2007:ACI

[2640] M. Fisher, B. Springborn, P. Schröder, and A. I. Bobenko. An algorithm for the construction of intrinsic Delaunay triangulations with applications to digital geometry processing. Computing, 81(2–3):199– 213, November 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=81&issue=2&spage=199.

LeBorne:2007:CDD

[2641] S. Le Borne and D. Cook. Construction of a discrete divergence-free basis through orthogonal factorization in H-arithmetic. Computing, 81(4): 215–238, December 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=81&issue=4&spage=215.

Bebendorf:2007:ESN

[2642] M. Bebendorf and Y. Chen. Efficient solution of nonlinear elliptic problems using hierarchical matrices with Broyden updates. Computing, 81(4): 239–257, December 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=81&issue=4&spage=239.

Hackbusch:2007:PMC

[2643] W. Hackbusch and W. Kress. A projection method for the computation of inner eigenvalues using high degree rational operators. Computing, 81(4): 259–268, December 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=81&issue=4&spage=259.

Mandel:2007:BFD

[2644] J. Mandel and B. Sousedík. BDDC and FETI-DP under minimalist assumptions. Computing, 81(4):269– 280, December 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=81&issue=4&spage=269.

Hubner:2007:EAS

[2645] T. Hübner and S. Turek. An efficient and accurate short-characteristics solver for radiative transfer problems. Computing, 81(4):281–296, December 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0010-485X&volume=81&
issue=4&spage=281.

Rui:2007:CUC

[2646] H. Rui. Convergence of an upwind control-volume mixed finite element method for convection-diffusion problems. Computing, 81(4):297– 315, December 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=81&issue=4&spage=297.

Hormann:2007:SGD

[2647] W. Hörmann. A simple generator for the t distribution. Computing, 81(4): 317–322, December 2007. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=81&issue=4&spage=317.

Acosta:2008:AEE

[2648] G. Acosta, Thomas Apel, R. G. Durán, and A. L. Lombardi. Anisotropic error estimates for an interpolant defined via moments. Computing, 82(1):1–9, April 2008. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=82&issue=1&spage=1.

Harbrecht:2008:NMB

[2649] Helmut Harbrecht. A Newton method for Bernoulli’s free boundary problem in three dimensions. Computing, 82(1):11–30, April 2008. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=82&issue=1&spage=11.

Vallejos:2008:MOM

[2650] Michelle Vallejos and A. Borzì. Multigrid optimization methods for linear and bilinear elliptic optimal control problems. Computing, 82 (1):31–52, April 2008. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=82&issue=1&spage=31.

Bouchala:2008:TSS

[2651] Jiri Bouchala, Z. Dostál, and M. Sadowská. Theoretically supported scalable BETI method for variational inequalities. Computing, 82 (1):53–75, April 2008. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=82&issue=1&spage=53.

Kearfott:2008:CSM

[2652] R. Baker Kearfott. A comparison of some methods for bounding connected and disconnected solution sets of interval linear systems. Computing, 82(1):77–102, April 2008. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=82&issue=1&spage=77.

Bastian:2008:GGIa

[2653] P. Bastian, M. Blatt, A. Dedner, C. Engwer, R. Klöfkorn, M. Ohlberger, and O. Sander. A generic grid interface for parallel and adaptive scientific computing. Part I: abstract framework. Computing, 82 (2–3):103–119, July 2008. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=82&issue=2&spage=103.

Bastian:2008:GGIb

[2654] P. Bastian, M. Blatt, A. Dedner, C. Engwer, R. Klöfkorn, M. Ohlberger, and O. Sander. A generic grid interface for parallel and adaptive scientific computing. Part II: implementation and tests in DUNE. Computing, 82(2–3):121–138, July 2008. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=82&issue=2&spage=121.

Of:2008:EAM

[2655] G. Of. An efficient algebraic multigrid preconditioner for a fast multipole boundary element method. Computing, 82(2–3):139–155, July 2008. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=82&issue=2&spage=139.

Li:2008:NLS

[2656] Jian Li and Zhangxin Chen. A new local stabilized nonconforming finite element method for the Stokes equations. Computing, 82 (2–3):157–170, July 2008. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=82&issue=2&spage=157.

Gassner:2008:PDP

[2657] Elisabeth Gassner and Johannes Hatzl. A parity domination problem in graphs with bounded treewidth and distancehereditary graphs. Computing, 82 (2–3):171–187, July 2008. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=82&issue=2&spage=171.

Thill:2008:MPR

[2658] M. Thill. A more precise rounding algorithm for rational numbers. Computing, 82(2–3):189–198, July 2008. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0010-485X&volume=82&
issue=2&spage=189. See erratum

[2662].

Carnicer:2008:REP

[2659] J. M. Carnicer, T. N. T. Goodman, and J. M. Peña. Roundoff errors for polynomial evaluation by a family of formulae. Computing, 82 (2–3):199–215, July 2008. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=82&issue=2&spage=199.

Georgiev:2008:MAR

[2660] I. Georgiev, J. Kraus, and S. Margenov.

Multilevel algorithms for Rannacher– Turek finite element approximation of 3D elliptic problems. Computing, 82 (4):217–239, September 2008. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=82&issue=4&spage=217.

Shakourifar:2008:NSN

[2661] M. Shakourifar and M. Dehghan.

On the numerical solution of nonlinear systems of Volterra integrodifferential equations with delay arguments. Computing, 82(4):241– 260, September 2008. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=82&issue=4&spage=241.

Thill:2008:EMP

[2662] M. Thill. Erratum: a more precise rounding algorithm for rational numbers. Computing, 82(4): 261–262, September 2008. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=82&issue=4&spage=261. See

[2658].

Karageorghis:2008:CME

[2663] Andreas Karageorghis and Yiorgos-Sokratis Smyrlis. Conformal mapping for the efficient MFS solution of Dirichlet boundary value problems. Computing, 83(1):1–24, May 2008. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=83&issue=1&spage=1.

Bieniasz:2008:AHM

[2664] L. K. Bieniasz. An adaptive Huber method with local error control, for the numerical solution of the first kind Abel integral equations. Computing, 83(1):25–39, May 2008. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=83&issue=1&spage=25.

Subramani:2008:CIP

[2665] K. Subramani. Computing inversion pair cardinality through partitionbased sorting. Computing, 83 (1):41–54, May 2008. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=83&issue=1&spage=41.

Mandel:2008:MMB

[2666] Jan Mandel, Bedrich Sousedík, and Clark R. Dohrmann. Multispace and multilevel BDDC. Computing, 83(2– 3):55–85, November 2008. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=83&issue=2&spage=55.

Linke:2008:NNM

[2667] Alexander Linke, Gunar Matthies, and Lutz Tobiska. Non-nested multi-grid solvers for mixed divergence-free Scott– Vogelius discretizations. Computing, 83 (2–3):87–107, November 2008. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=83&issue=2&spage=87.

Lee:2008:HSP

[2668] T. L. Lee, T. Y. Li, and C. H.

Tsai. HOM4PS-2.0: a software package for solving polynomial systems by the polyhedral homotopy continuation method. Computing, 83(2–3): 109–133, November 2008. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=83&issue=2&spage=109.

Ferraz-Leite:2008:SPE

[2669] S. Ferraz-Leite and D. Praetorius. Simple a posteriori error estimators for the h-version of the boundary element method. Computing, 83(4): 135–162, December 2008. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=83&issue=4&spage=135.

Bieniasz:2008:IAH

[2670] L. K. Bieniasz. Initialisation of the adaptive Huber method for solving the first kind Abel integral equation. Computing, 83(4): 163–174, December 2008. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0010-485X&
                                        volume=83&issue=4&spage=163.

Alefeld:2008:EBC

[2671] G. Alefeld and Z. Wang. Error bounds for complementarity problems with tridiagonal nonlinear functions. Computing, 83(4): 175–192, December 2008. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=83&issue=4&spage=175.

Elmasry:2008:TNM

[2672] Amr Elmasry, Claus Jensen, and Jyrki Katajainen. Two new methods for constructing double-ended priority queues from priority queues. Computing, 83 (4):193–204, December 2008. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=83&issue=4&spage=193.

Garcke:2009:FMD

[2673] Jochen Garcke and Markus Hegland. Fitting multidimensional data using gradient penalties and the sparse grid combination technique. Computing, 84(1–2):1–25, April 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=84&issue=1&spage=1.

Lass:2009:IAM

[2674] O. Lass, M. Vallejos, A. Borzi, and C. C. Douglas. Implementation and analysis of multigrid schemes with finite elements for elliptic optimal control problems. Computing, 84 (1–2):27–48, April 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=84&issue=1&spage=27.

Khoromskij:2009:AHM

[2675] B. N. Khoromskij, A. Litvinenko, and H. G. Matthies. Application of hierarchical matrices for computing the Karhunen–Loève expansion. Computing, 84(1–2):49–67, April 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0010-485X&volume=84&
issue=1&spage=49.

Haslinger:2009:FDF

[2676] J. Haslinger, T. Kozubek, and R. Kucera. Fictitious domain formulations of unilateral problems: analysis and algorithms. Computing, 84 (1–2):69–96, April 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=84&issue=1&spage=69.

Jiang:2009:AVD

[2677] L. Jiang and C. C. Douglas. An analysis of 4D variational data assimilation and its application. Computing, 84(1–2):97–120, April 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0010-485X&
                                        volume=84&issue=1&spage=97.

Hadid:2009:SMM

[2678] Rachid Hadid and Mehmet Hakan Karaata. Stabilizing maximum matching in bipartite networks. Computing, 84(1–2):121–138, April 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=84&issue=1&spage=121.

Balogh:2009:ILB

[2679] János Balogh, József Békési, Gábor Galambos, and Mihály Csaba Markót. Improved lower bounds for semi-online bin packing problems. Computing, 84 (1–2):139–148, April 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=84&issue=1&spage=139.

Bruno:2009:HOI

[2680] Oscar P. Bruno, Jeffrey S. Ovall, and Catalin Turc. A high-order integral algorithm for highly singular PDE solutions in Lipschitz domains. Computing, 84(3–4):149–181, June 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=84&issue=3&spage=149.

Wu:2009:GBT

[2681] Shi-Liang Wu, Ting-Zhu Huang, and Cui-Xia Li. Generalized block triangular preconditioner for symmetric saddle point problems. Computing, 84 (3–4):183–208, June 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=84&issue=3&spage=183.

Mukherjee:2009:PUH

[2682] Kaushik Mukherjee and Srinivasan Natesan. Parameter-uniform hybrid numerical scheme for time-dependent convection-dominated initial-boundary-value problems. Computing, 84(3– 4):209–230, June 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=84&issue=3&spage=209.

Ceballos:2009:ACM

[2683] M. Ceballos, J. Núñez, and A. F.

Tenorio. Algorithm to compute the maximal abelian dimension of Lie algebras. Computing, 84(3– 4):231–239, June 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=84&issue=3&spage=231.

Heuberger:2009:E

[2684] C. Heuberger and V. Rijmen. Editorial. Computing, 85(1–2):1–2, June 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0010-485X&volume=85&
issue=1&spage=1.

Klingler:2009:DLF

[2685] Lee C. Klingler, Spyros S. Magliveras, Fred Richman, and Michal Sramka. Discrete logarithms for finite groups. Computing, 85(1– 2):3–19, June 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=85&issue=1&spage=3.

Grosek:2009:NMT

[2686] Otokar Grosek, Milan Vojvoda, and Róbert Krchnavý. A new matrix test for randomness. Computing, 85 (1–2):21–36, June 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=85&issue=1&spage=21.

Gangopadhyay:2009:ANE

[2687] Sugata Gangopadhyay, Deepmala Sharma, Sumanta Sarkar, and Subhamoy Maitra. On affine (non)equivalence of Boolean functions. Computing, 85(1–2):37–55, June 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=85&issue=1&spage=37.

Jaworski:2009:RKP

[2688] J. Jaworski, M. Ren, and K. Rybarczyk. Random key predistribution for wireless sensor networks using deployment knowledge. Computing, 85(1–2):57–76, June 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=85&issue=1&spage=57.

Dujella:2009:VWA

[2689] Andrej Dujella. A variant of Wiener’s attack on RSA. Computing, 85 (1–2):77–83, June 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=85&issue=1&spage=77.

Daemen:2009:CAE

[2690] Joan Daemen, Mario Lamberger, Norbert Pramstaller, Vincent Rijmen, and Frederik Vercauteren. Computational aspects of the expected differential probability of 4-round AES and AES-like ciphers. Computing, 85 (1–2):85–104, June 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=85&issue=1&spage=85.

Buchmann:2009:PQC

[2691] Johannes Buchmann, Richard Lindner, Markus Rückert, and Michael Schneider. Post-quantum cryptography: lattice signatures. Computing, 85(1–2):105–125, June 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=85&issue=1&spage=105.

Csirmaz:2009:IFG

[2692] László Csirmaz and Péter Ligeti. On an infinite family of graphs with information ratio 2 − 1/k. Computing, 85(1–2):127–136, June 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=85&issue=1&spage=127.

Klonowski:2009:SGS

[2693] Marek Klonowski, Lukasz Krzywiecki, Miroslaw Kutylowski, and Anna Lauks. Step-out group signatures. Computing, 85(1–2):137–151, June 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=85&issue=1&spage=137.

Mihaljevic:2009:ASC

[2694] Miodrag J. Mihaljević and Hideki Imai. An approach for stream ciphers design based on joint computing over random and secret data. Computing, 85 (1–2):153–168, June 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=85&issue=1&spage=153.

Oseledets:2009:LAT

[2695] I. V. Oseledets, D. V. Savostyanov, and E. E. Tyrtyshnikov. Linear algebra for tensor problems. Computing, 85(3):169–188, July 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0010-485X&
volume=85&issue=3&spage=169.

Bouchala:2009:STB

[2696] J. Bouchala, Z. Dostál, and M. Sadowská. Scalable Total BETI based algorithm for 3D coercive contact problems of linear elastostatics. Computing, 85(3):189–217, July 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=85&issue=3&spage=189.

Zhang:2009:SCN

[2697] Xiaoping Zhang, Jiming Wu, and Dehao Yu. The superconvergence of composite Newton–Cotes rules for Hadamard finite-part integral on a circle. Computing, 85 (3):219–244, July 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=85&issue=3&spage=219.

Fugenschuh:2009:ERS

[2698] A. Fügenschuh, S. Göttlich, M. Herty, C. Kirchner, and A. Martin. Efficient reformulation and solution of a nonlinear PDE-controlled flow network model. Computing, 85 (3):245–265, July 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=85&issue=3&spage=245.

Feuersanger:2009:PML

[2699] Christian Feuersänger and Michael Griebel. Principal manifold learning by sparse grids. Computing, 85 (4):267–299, August 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=85&issue=4&spage=267.

Dratman:2009:SPS

[2700] Ezequiel Dratman and Guillermo Matera. On the solution of the polynomial systems arising in the discretization of certain ODEs. Computing, 85 (4):301–337, August 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=85&issue=4&spage=301.

Kang:2009:FDP

[2701] Tong Kang and Kwang Ik Kim. Fully discrete potential-based finite element methods for a transient eddy current problem. Computing, 85(4): 339–362, August 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=85&issue=4&spage=339.

Harbrecht:2009:WBM

[2702] Helmut Harbrecht and Maharavo Randrianarivony. Wavelet BEM on molecular surfaces: parametrization and implementation. Computing, 86(1): 1–22, September 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=86&issue=1&spage=1.

Karaa:2009:IAA

[2703] S. Karaa. Improved accuracy for the approximate factorization of parabolic equations. Computing, 86(1):23– 36, September 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=86&issue=1&spage=23.

Li:2009:PSS

[2704] Jian Li, Yinnian He, and Zhangxin Chen. Performance of several stabilized finite element methods for the Stokes equations based on the lowest equal-order pairs. Computing, 86 (1):37–51, September 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=86&issue=1&spage=37.

Bai:2009:MIO

[2705] Zhong-Zhi Bai and Jun-Feng Yin.

Modified incomplete orthogonal factorization methods using Givens rotations. Computing, 86(1):53– 69, September 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=86&issue=1&spage=53.

Hahmann:2009:ESI

[2706] Stefanie Hahmann, Gerald Farin, Jörg Peters, and Wenping Wang. Editorial: Special issue on geometric modeling (Dagstuhl 2008). Computing, 86 (2–3):71–72, October 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=86&issue=2&spage=71.

Aigner:2009:DRG

[2707] Martin Aigner and Bert Jüttler. Distance regression by Gauss–Newtontype methods and iteratively reweighted least-squares. Computing, 86 (2–3):73–87, October 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=86&issue=2&spage=73.

Hering-Bertram:2009:AQI

[2708] Martin Hering-Bertram, Gerd Reis, and Frank Zeilfelder. Adaptive quasi-interpolating quartic splines. Computing, 86(2–3):89–100, October 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0010-485X&volume=86&
                                        issue=2&spage=89.

Brunet:2009:MMH

[2709] Pere Brunet, Antoni Chica, Isabel Navazo, and Àlvar Vinacua. Massive mesh hole repair minimizing user intervention. Computing, 86(2–3): 101–115, October 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=86&issue=2&spage=101.

Ginnis:2009:DMC

[2710] Alexandros I. Ginnis, E. I. Karousos, and P. D. Kaklis. A discrete methodology for controlling the sign of curvature and torsion for NURBS. Computing, 86 (2–3):117–129, October 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=86&issue=2&spage=117.

Moorhead:2009:IVI

[2711] Robert Moorhead, Yanlin Guan, Hans Hagen, Sven Böttger, Natallia Kotava, and Christian Wagner. Illustrative visualization: interrogating triangulated surfaces. Computing, 86(2– 3):131–150, October 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=86&issue=2&spage=131.

Hernandez-Mederos:2009:CSA

[2712] Victoria Hernández-Mederos, Jorge C.

Estrada-Sarlabous, Silvio R. Morales, and Ioannis Ivrissimtzis. Curve subdivision with arc-length control. Computing, 86(2–3):151– 169, October 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=86&issue=2&spage=151.

Karciauskas:2009:LSS

[2713] Kestutis Karciauskas and Jorg Peters. Lens-shaped surfaces and C^{2} subdivision. Computing, 86(2–3): 171–183, October 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=86&issue=2&spage=171.

Mossner:2009:EBP

[2714] Bernhard Mößner and Ulrich Reif. Error bounds for polynomial tensor product interpolation. Computing, 86(2– 3):185–197, October 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=86&issue=2&spage=185.

Nielson:2009:NIE

[2715] Gregory M. Nielson. Normalized implicit eigenvector least squares operators for noisy scattered data: radial basis functions. Computing, 86 (2–3):199–212, October 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=86&issue=2&spage=199.

Sabin:2009:TOQ

[2716] Malcolm Sabin. Two open questions relating to subdivision. Computing, 86 (2–3):213–217, October 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=86&issue=2&spage=213.

Kyratzi:2009:SSA

[2717] Sofia Kyratzi and Nickolas S. Sapidis. From sketch to solid: an algebraic cross-section criterion for the realizability of a wireframe sketch. Computing, 86(2–3): 219–234, October 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=86&issue=2&spage=219.

Tong:2009:LSF

[2718] Wei hua Tong and Tae wan Kim. Local and singularity-free G^{1} triangular spline surfaces using a minimum degree scheme. Computing, 86(2– 3):235–255, October 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=86&issue=2&spage=235.

Zheng:2009:NRC

[2719] Jianmin Zheng. C^{1} NURBS representations of G^{1} composite rational Bézier curves. Computing, 86(2– 3):257–268, October 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=86&issue=2&spage=257.

Graser:2009:DSM

[2720] Carsten Gräser and Oliver Sander.

The dune-subgrid module and some applications. Computing, 86(4): 269–290, November 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=86&issue=4&spage=269.

Mayer:2009:PAS

[2721] Jan Mayer. Parallel algorithms for solving linear systems with sparse triangular matrices. Computing, 86(4): 291–312, November 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=86&issue=4&spage=291.

Marcinkowski:2009:CRM

[2722] Leszek Marcinkowski, Talal Rahman, and Jan Valdman. A 3D Crouzeix–Raviart mortar finite element. Computing, 86(4):313– 330, November 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=86&issue=4&spage=313.

Alizadeh:2009:ICL

[2723] Behrooz Alizadeh, Rainer E. Burkard, and Ulrich Pferschy. Inverse 1-center location problems with edge length augmentation on trees. Computing, 86 (4):331–343, November 2009. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=86&issue=4&spage=331.

Douglas:2010:E

[2724] Craig C. Douglas. Editorial. Computing, 87(1–2):1, March 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=87&issue=1&spage=1.

Epstein:2010:MSA

[2725] Leah Epstein, Asaf Levin, and Betzalel Menahem. Minimization of SONET ADMs in ring networks revisited. Computing, 87 (1–2):3–19, March 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=87&issue=1&spage=3.

Vallejos:2010:MGP

[2726] Michelle Vallejos. MGOPT with gradient projection method for solving bilinear elliptic optimal control problems. Computing, 87(1– 2):21–33, March 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=87&issue=1&spage=21.

Bieniasz:2010:AHM

[2727] Leslaw K. Bieniasz. An adaptive Huber method for weakly singular second kind Volterra integral equations with non-linear dependencies between unknowns and their integrals. Computing, 87(1–2):35–54, March 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=87&issue=1&spage=35.

Ye:2010:TEM

[2728] Jun Ye. Two effective measures of intuitionistic fuzzy entropy. Computing, 87(1–2):55–62, March 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=87&issue=1&spage=55.

Cools:2010:CLR

[2729] Ronald Cools, Frances Y. Kuo, and Dirk Nuyens. Constructing lattice rules based on weighted degree of exactness and worst case error. Computing, 87 (1–2):63–89, March 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=87&issue=1&spage=63.

Cap:2010:E

[2730] Clemens H. Cap. Editorial. Computing, 87(3–4):91, May 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=87&issue=3&spage=91.

Bai:2010:MHI

[2731] Zhong-Zhi Bai, Michele Benzi, and Fang Chen. Modified HSS iteration methods for a class of complex symmetric linear systems. Computing, 87(3–4):93–111, May 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=87&issue=3&spage=93.

Li:2010:PPS

[2732] Yuan Li and Kaitai Li. Pressure projection stabilized finite element method for Navier–Stokes equations with nonlinear slip boundary conditions. Computing, 87(3– 4):113–133, May 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=87&issue=3&spage=113.

Franz:2010:LPS

[2733] Sebastian Franz and Gunar Matthies.

Local projection stabilisation on Stype meshes for convection-diffusion problems with characteristic layers. Computing, 87(3–4):135–167, May 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0010-485X&volume=87&
issue=3&spage=135.

Belhaj:2010:CBF

[2734] Skander Belhaj. Computing the block factorization of complex Hankel matrices. Computing, 87(3– 4):169–186, May 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=87&issue=3&spage=169.

Li:2010:SAM

[2735] Zi-Cai Li, Jin Huang, and Hung-Tsai Huang. Stability analysis of method of fundamental solutions for mixed boundary value problems of Laplace’s equation. Computing, 88 (1–2):1–29, June 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=88&issue=1&spage=1.

Zheng:2010:PRB

[2736] Yujun Zheng and Jinyun Xue. A problem reduction based approach to discrete optimization algorithm design. Computing, 88 (1–2):31–54, June 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=88&issue=1&spage=31.

Salem:2010:ESC

[2737] Fatima K. Abu Salem and Rawan N. Soudah. An empirical study of cache-oblivious polygon indecomposability testing. Computing, 88 (1–2):55–78, June 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=88&issue=1&spage=55.

Rezghi:2010:LBP

[2738] Mansoor Rezghi and S. M. Hosseini. Lanczos based preconditioner for discrete ill-posed problems. Computing, 88(1–2):79–96, June 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=88&issue=1&spage=79.

Zhang:2010:PNW

[2739] Xiaoqin Zhang, Liying Kang, and Yukun Cheng. The pos/negweighted median problem on block graphs with subgraph-shaped customers. Computing, 88(3–4):97– 110, July 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0010-485X&volume=88&
issue=3&spage=97.

Benner:2010:QDF

[2740] Peter Benner and Thomas Mach.

On the QR decomposition of Hmatrices. Computing, 88(3–4): 111–129, July 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=88&issue=3&spage=111.

Torres:2010:FCTa

[2741] Ricardo A. Prato Torres, Ernst P.

Stephan, and Florian Leydecker. A FE/BE coupling for the 3D timedependent eddy current problem. Part I: a priori error estimates. Computing, 88(3–4):131–154, July 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=88&issue=3&spage=131.

Torres:2010:FCTb

[2742] Ricardo A. Prato Torres, Ernst P.

Stephan, and Florian Leydecker. A FE/BE coupling for the 3D timedependent eddy current problem. Part II: a posteriori error estimates and adaptive computations. Computing, 88(3–4):155–172, July 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=88&issue=3&spage=155.

Baxansky:2010:UDF

[2743] Artemy Baxansky and Nahum Kiryati. Use of a double Fourier series for three-dimensional shape representation. Computing, 88(3– 4):173–191, July 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=88&issue=3&spage=173.

Seck-Tuoh-Mora:2010:ECA

[2744] Juan C. Seck-Tuoh-Mora, Genaro J. Martínez, Norberto Hernández-Romero, and Joselito Medina-Marín. Elementary cellular automaton Rule 110 explained as a block substitution system Rule 110 as a block substitution system. Computing, 88(3– 4):193–205, July 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=88&issue=3&spage=193.

Mustafa:2010:MPA

[2745] Ghulam Mustafa and Najma Abdul Rehman. The mask of (2b + 4)-point n-ary subdivision scheme. Computing, 90(1–2):1–14, October 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=90&issue=1&spage=1.

Kumar:2010:HOP

[2746] Mukesh Kumar and S. Chandra Sekhara Rao. High order parameter-robust numerical method for time dependent singularly perturbed reactiondiffusion problems. Computing, 90 (1–2):15–38, October 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=90&issue=1&spage=15.

Li:2010:NSA

[2747] Jiao-Fen Li, Xi-Yan Hu, Xue-Feng Duan, and Lei Zhang. Numerical solutions of AXB = C for mirrorsymmetric matrix X under a specified submatrix constraint. Computing, 90 (1–2):39–56, October 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=90&issue=1&spage=39.

Zhang:2010:EPR

[2748] Chengjian Zhang, Tingting Qin, and Jie Jin. The extended Pouzet– Runge–Kutta methods for nonlinear neutral delay-integro-differential equations. Computing, 90(1–2): 57–71, October 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=90&issue=1&spage=57.

Wang:2010:MGR

[2749] Jian Wang, Yan-Heng Liu, and Yu Jiao. Modeling generation of the router-level topology of an ISP network. Computing, 90(1– 2):73–88, October 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=90&issue=1&spage=73.

Li:2010:FCD

[2750] Wulan Li and Xu Da. Finite central difference/finite element approximations for parabolic integrodifferential equations. Computing, 90 (3–4):89–111, November 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=90&issue=3&spage=89.

Subramani:2010:TLH

[2751] K. Subramani and Kamesh Madduri. Two-level heaps: a new priority queue structure with applications to the single source shortest path problem. Computing, 90(3–4): 113–130, November 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=90&issue=3&spage=113.

Argyros:2010:UAC

[2752] Ioannis K. Argyros and Saïd Hilout. A unified approach for the convergence of certain numerical algorithms, using recurrent functions. Computing, 90(3– 4):131–164, November 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=90&issue=3&spage=131.

Dedner:2010:GIP

[2753] Andreas Dedner, Robert Klöfkorn, Martin Nolte, and Mario Ohlberger. A generic interface for parallel and adaptive discretization schemes: abstraction principles and the Dune– Fem module. Computing, 90(3–4): 165–196, November 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=90&issue=3&spage=165.

Vazquez:2010:IRV

[2754] S. G. Vázquez, N. Barreira, M. G.

Penedo, M. Ortega, and A. Pose-Reino. Improvements in retinal vessel clustering techniques: towards the automatic computation of the arterio venous ratio. Computing, 90(3– 4):197–217, November 2010. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=90&issue=3&spage=197.

Dustdar:2011:E

[2755] S. Dustdar. Editorial. Computing, 91(1):1–2, January 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=91&issue=1&spage=1.

Deng:2011:CDE

[2756] Yuhui Deng and Brandon Pung. Conserving disk energy in virtual machine based environments by amplifying bursts. Computing, 91 (1):3–21, January 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=91&issue=1&spage=3.

Doelitzscher:2011:PCC

[2757] Frank Doelitzscher, Anthony Sulistio, Christoph Reich, Hendrik Kuijs, and David Wolf. Private cloud for collaboration and e-learning services: from IaaS to SaaS. Computing, 91 (1):23–42, January 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=91&issue=1&spage=23.

Psaier:2011:SSH

[2758] Harald Psaier and Schahram Dustdar. A survey on self-healing systems: approaches and systems. Computing, 91(1):43–73, January 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=91&issue=1&spage=43.

Truong:2011:CCS

[2759] Hong-Linh Truong and Schahram Dustdar. Cloud computing for small research groups in computational science and engineering: current status and outlook. Computing, 91 (1):75–91, January 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=91&issue=1&spage=75.

Vaquero:2011:LSS

[2760] Luis M. Vaquero, Luis Rodero-Merino, and Daniel Morán. Locking the sky: a survey on IaaS cloud security. Computing, 91(1): 93–118, January 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=91&issue=1&spage=93.

Sathasivam:2011:LMN

[2761] Saratha Sathasivam and Wan Ahmad Tajuddin Wan Abdullah. Logic mining in neural network: reverse analysis method. Computing, 91(2): 119–133, February 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=91&issue=2&spage=119.

Bella:2011:UNF

[2762] Antonio Bella, Cèsar Ferri, José Hernández-Orallo, and María José Ramírez-Quintana. Using negotiable features for prescription problems. Computing, 91(2):135–168, February 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0010-485X&volume=91&
issue=2&spage=135.

Campos:2011:OAM

[2763] Jordi Campos, Marc Esteva, Maite López-Sánchez, Javier Morales, and Maria Salamó. Organisational adaptation of multi-agent systems in a peerto-peer scenario. Computing, 91(2): 169–215, February 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=91&issue=2&spage=169.

Zhang:2011:PLS

[2764] Jiansong Zhang and Danping Yang. Parallel least-squares finite element method for time-dependent convectiondiffusion system. Computing, 91 (3):217–240, March 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=91&issue=3&spage=217.

Chang:2011:CGR

[2765] Harry M. Chang. Constructing ngram rules for natural language models through exploring the limitation of the Zipf–Mandelbrot law. Computing, 91(3):241–264, March 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=91&issue=3&spage=241.

Clift:2011:COA

[2766] Neill Michael Clift. Calculating optimal addition chains. Computing, 91 (3):265–284, March 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=91&issue=3&spage=265.

Bahig:2011:NSG

[2767] Hatem M. Bahig and Hazem M. Bahig.

A new strategy for generating shortest addition sequences. Computing, 91(3):285–306, March 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=91&issue=3&spage=285.

Kulisch:2011:EDP

[2768] Ulrich Kulisch and Van Snyder. The exact dot product as basic tool for long interval arithmetic. Computing, 91(3):307–313, March 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=91&issue=3&spage=307.

Araujo:2011:PME

[2769] Carlos Araújo, Paulo Maciel, Armin Zimmermann, Ermeson Andrade, Erica Sousa, Gustavo Callou, and Paulo Cunha. Performability modeling of electronic funds transfer systems. Computing, 91(4): 315–334, April 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=91&issue=4&spage=315.

Bahig:2011:SRA

[2770] Hatem M. Bahig. Star reduction among minimal length addition chains. Computing, 91(4): 335–352, April 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=91&issue=4&spage=335.

Sammany:2011:HNG

[2771] M. Sammany, E. Pelican, and T. A. Harak. Hybrid neuro-genetic based method for solving ill-posed inverse problem occurring in synthesis of electromagnetic fields. Computing, 91 (4):353–364, April 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=91&issue=4&spage=353.

Bahig:2011:IME

[2772] Hazem M. Bahig. Integer merging on EREW PRAM. Computing, 91 (4):365–378, April 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=91&issue=4&spage=365.

Bai:2011:BPE

[2773] Zhong-Zhi Bai. Block preconditioners for elliptic PDE-constrained optimization problems. Computing, 91 (4):379–395, April 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0010-485X&
volume=91&issue=4&spage=379.

Kulisch:2011:VFE

[2774] Ulrich Kulisch. Very fast and exact accumulation of products. Computing, 91(4):397–405, April 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=91&issue=4&spage=397.

Mukherjee:2011:RET

[2775] Kaushik Mukherjee and Srinivasan Natesan. Richardson extrapolation technique for singularly perturbed parabolic convection-diffusion problems. Computing, 92(1):1–32, May 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0010-485X&volume=92&
issue=1&spage=1.

Chatterjee:2011:TFM

[2776] S. Chatterjee, S. Nigam, J. B. Singh, and L. N. Upadhyaya. Transfer function modelling in software reliability. Computing, 92(1):33–48, May 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0010-485X&volume=92&
issue=1&spage=33.

Dagher:2011:FAB

[2777] Issam Dagher. Fuzzy ART-based prototype classifier. Computing, 92 (1):49–63, May 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=92&issue=1&spage=49.

Boglaev:2011:MIS

[2778] Igor Boglaev. Monotone iterates for solving coupled systems of nonlinear parabolic equations. Computing, 92(1):65–95, May 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=92&issue=1&spage=65.

Zhang:2011:ZNN

[2779] Yunong Zhang, Yiwen Yang, Ning Tan, and Binghuang Cai. Zhang neural network solving for timevarying full-rank matrix Moore– Penrose inverse. Computing, 92 (2):97–121, June 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=92&issue=2&spage=97.

Jabri:2011:PBP

[2780] Riad Jabri. A predictive bottomup parser. Computing, 92(2): 123–167, June 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=92&issue=2&spage=123.

Khattri:2011:DFA

[2781] Sanjay Kumar Khattri and Torgrim Log. Derivative free algorithm for solving nonlinear equations. Computing, 92(2):169–179, June 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=92&issue=2&spage=169. See

erratum [2894] and comments [2895, 2896].

Allahviranloo:2011:NMS

[2782] T. Allahviranloo, S. Abbasbandy, S. Salahshour, and A. Hakimzadeh. A new method for solving fuzzy linear differential equations. Computing, 92(2):181–197, June 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=92&issue=2&spage=181.

Li:2011:NSE

[2783] Zi-Cai Li, Ming-Gong Lee, and Jeng-Tzong Chen. New series expansions for fundamental solutions of linear elastostatics in 2D. Computing, 92(3):199–224, July 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=92&issue=3&spage=199.

Golbabai:2011:MMN

[2784] A. Golbabai and A. Safdari-Vaighani.

A meshless method for numerical solution of the coupled Schrödinger-KdV equations. Computing, 92 (3):225–242, July 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0010-485X&
                                        volume=92&issue=3&spage=225.

Akanbi:2011:SGE

[2785] Moses Adebowale Akanbi. On 3stage geometric explicit Runge–Kutta method for singular autonomous initial value problems in ordinary differential equations. Computing, 92 (3):243–263, July 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=92&issue=3&spage=243.

Yadani:2011:NPH

[2786] Kenichi Yadani, Koichi Kondo, and Masashi Iwasaki. Numerical performance of hyperplane constrained method and its hybrid method for singular value decomposition. Computing, 92(3):265–283, July 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=92&issue=3&spage=265.

Roos:2011:FEL

[2787] Hans-Görg Roos and Martin Schopf. Finite elements on locally uniform meshes for convection-diffusion problems with boundary layers. Computing, 92(4):285–296, August 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=92&issue=4&spage=285.

Chen:2011:EIN

[2788] Chin-Yun Chen. Extended interval Newton method based on the precise quotient set. Computing, 92 (4):297–315, August 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=92&issue=4&spage=297.

Yuan:2011:BTR

[2789] Gonglin Yuan, Zengxin Wei, and Xiwen Lu. A BFGS trust-region method for nonlinear equations. Computing, 92(4):317–333, August 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=92&issue=4&spage=317.

Harbrecht:2011:WBM

[2790] H. Harbrecht and M. Randrianarivony.

Wavelet BEM on molecular surfaces: solvent excluded surfaces. Computing, 92(4):335–364, August 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=92&issue=4&spage=335.

Lang:2011:QBS

[2791] Feng-Gong Lang and Xiao-Ping Xu. Quartic B-spline collocation method for fifth order boundary value problems. Computing, 92(4): 365–378, August 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0010-485X&
                                        volume=92&issue=4&spage=365.

Song:2011:QBA

[2792] Seung-Jae Song, Eung-Hee Kim, Hong-Gee Kim, and Harshit Kumar. Querybased association rule mining supporting user perspective. Computing, 93 (1):1–25, September 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=93&issue=1&spage=1.

Peng:2011:NPB

[2793] Xiao-Fei Peng, Wen Li, and Shuhuang Xiang. New preconditioners based on symmetric-triangular decomposition for saddle point problems. Computing, 93(1):27–46, September 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=93&issue=1&spage=27.

Park:2011:ULH

[2794] Jin Han Park, Min Gwi Gwak, and Young Chel Kwun. Uncertain linguistic harmonic mean operators and their applications to multiple attribute group decision making. Computing, 93 (1):47–64, September 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=93&issue=1&spage=47.

Jia:2011:CSF

[2795] Hongen Jia, Demin Liu, and Kaitai Li. A characteristic stabilized finite element method for the non-stationary Navier–Stokes equations. Computing, 93(1):65–87, September 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=93&issue=1&spage=65.

Ulieru:2011:TBW

[2796] Mihaela Ulieru. Towards a better way of computing: a return from bits to atoms via physics and biology. Foreword to the first special issue on bioinspired computing. Computing, 93 (2–4):89–90, December 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=93&issue=2&spage=89.

Gershenfeld:2011:ARR

[2797] Neil Gershenfeld. Aligning the representation and reality of computation with asynchronous logic automata. Computing, 93(2–4):91– 102, December 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=93&issue=2&spage=91.

Squillero:2011:AEC

[2798] Giovanni Squillero. Artificial evolution in computer aided design: from the optimization of parameters to the creation of assembly programs. Computing, 93(2–4):103– 120, December 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=93&issue=2&spage=103.

Whitacre:2011:RTI

[2799] James M. Whitacre. Recent trends indicate rapid growth of natureinspired optimization in academia and industry. Computing, 93(2–4): 121–133, December 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=93&issue=2&spage=121.

Whitacre:2011:SFE

[2800] James M. Whitacre. Survival of the flexible: explaining the recent popularity of nature-inspired optimization within a rapidly evolving world. Computing, 93(2–4): 135–146, December 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=93&issue=2&spage=135.

VanMieghem:2011:ISE

[2801] Piet Van Mieghem. The Nintertwined SIS epidemic network model. Computing, 93(2–4):147– 169, December 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=93&issue=2&spage=147.

Shen:2012:TSM

[2802] Haifeng Shen, Gang Liu, and Jun Guo. Two-stage model-based feature compensation for robust speech recognition. Computing, 94(1): 1–20, January 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=94&issue=1&spage=1.

Fisher:2012:HFH

[2803] Nils Fisher, Zhen He, and Mitzi Mc-Carthy. A hybrid filesystem for hard disk drives in tandem with flash memory. Computing, 94(1): 21–68, January 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=94&issue=1&spage=21.

Deng:2012:EDI

[2804] Yuhui Deng, Kai Li, Lingwei Zhang, Ming Fang, and Xinyu Huang. Evaluating disk idle behavior by leveraging disk schedulers. Computing, 94 (1):69–93, January 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=94&issue=1&spage=69.

Revol:2012:E

[2805] Nathalie Revol and Goetz Alefeld. Editorial. Computing, 94 (2–4):95, March 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=94&issue=2&spage=95.

Garloff:2012:PTD

[2806] Jürgen Garloff. Pivot tightening for direct methods for solving symmetric positive definite systems of linear interval equations. Computing, 94 (2–4):97–107, March 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=94&issue=2&spage=97.

Zimmer:2012:SVS

[2807] Michael Zimmer, Walter Krämer, and Evgenija D. Popova. Solvers for the verified solution of parametric linear systems. Computing, 94(2– 4):109–123, March 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=94&issue=2&spage=109.

Griewank:2012:NSA

[2808] Andreas Griewank, Kshitij Kulshreshtha, and Andrea Walther. On the numerical stability of algorithmic differentiation. Computing, 94 (2–4):125–149, March 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=94&issue=2&spage=125.

Nakao:2012:PEI

[2809] Mitsuhiro T. Nakao, Takehiko Kinoshita, and Takuma Kimura. On a posteriori estimates of inverse operators for linear parabolic initialboundary value problems. Computing, 94(2–4):151–162, March 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=94&issue=2&spage=151.

Auer:2012:VSC

[2810] Ekaterina Auer and Andreas Rauh.

VERICOMP: a system to compare and assess verified IVP solvers. Computing, 94(2–4):163–172, March 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=94&issue=2&spage=163.

Araya:2012:IEB

[2811] Ignacio Araya, Bertrand Neveu, and Gilles Trombettoni. An interval extension based on occurrence grouping. Computing, 94(2– 4):173–188, March 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=94&issue=2&spage=173.

Bouissou:2012:GPB

[2812] Olivier Bouissou, Eric Goubault, Jean Goubault-Larrecq, and Sylvie Putot. A generalization of p-boxes to affine arithmetic. Computing, 94 (2–4):189–201, March 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=94&issue=2&spage=189.

Siegel:2012:LAL

[2813] Stefan Siegel and Jürgen Wolff von Gudenberg. A long accumulator like a carry-save adder. Computing, 94 (2–4):203–213, March 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=94&issue=2&spage=203.

Nehmeier:2012:IAU

[2814] Marco Nehmeier. Interval arithmetic using expression templates, template meta programming and the upcoming C++ standard. Computing, 94 (2–4):215–228, March 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=94&issue=2&spage=215.

Kramer:2012:MAP

[2815] Walter Krämer. Multiple/arbitrary precision interval computations in C-XSC. Computing, 94(2–4):229–241, March 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0010-485X&volume=94&
                                        issue=2&spage=229.

Nehmeier:2012:SHI

[2816] M. Nehmeier, S. Siegel, and J. Wolff von Gudenberg. Specification of hardware for interval arithmetic. Computing, 94(2–4):243–255, March 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

                                        http://www.springerlink.com/openurl.

asp?genre=article&issn=0010-485X&
volume=94&issue=2&spage=243.

Nava:2012:TIT

[2817] Jaime Nava and Vladik Kreinovich. Towards interval techniques for model validation. Computing, 94(2–4): 257–269, March 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=94&issue=2&spage=257.

Kubica:2012:CPC

[2818] Bartlomiej Jacek Kubica. A class of problems that can be solved using interval algorithms. Computing, 94(2– 4):271–280, March 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=94&issue=2&spage=271.

Dyllong:2012:CVD

[2819] Eva Dyllong and Stefan Kiel. A comparison of verified distance computation between implicit objects using different arithmetics for range enclosure. Computing, 94(2–4): 281–296, March 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=94&issue=2&spage=281.

Jaulin:2012:SSV

[2820] Luc Jaulin. Solving set-valued constraint satisfaction problems. Computing, 94(2–4):297–311, March 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=94&issue=2&spage=297.

Rebner:2012:VRD

[2821] Gabor Rebner, Ekaterina Auer, and Wolfram Luther. A verified realization of a Dempster–Shafer based fault tree analysis. Computing, 94 (2–4):313–324, March 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=94&issue=2&spage=313.

Patil:2012:GOM

[2822] Bhagyesh V. Patil, P. S. V. Nataraj, and Sharad Bhartiya. Global optimization of mixed-integer nonlinear (polynomial) programming problems: the Bernstein polynomial approach. Computing, 94(2–4): 325–343, March 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=94&issue=2&spage=325.

Rauh:2012:VSA

[2823] Andreas Rauh, Ekaterina Auer, Thomas Dötschel, and Harald Aschemann. Verified stability analysis of continuous-time control systems with bounded parameter uncertainties and stochastic disturbances. Computing, 94 (2–4):345–356, March 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0010-485X&
volume=94&issue=2&spage=345.

Rauh:2012:SBF

[2824] Andreas Rauh, Julia Kersten, Ekaterina Auer, and Harald Aschemann. Sensitivity-based feedforward and feedback control for uncertain systems. Computing, 94(2– 4):357–367, March 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=94&issue=2&spage=357.

Berthe:2012:NDD

[2825] V. Berthé. Numeration and discrete dynamical systems. Computing, 94 (2–4):369–387, March 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=94&issue=2&spage=369.

Al-Khateeb:2012:EMS

[2826] Asef Al-Khateeb, Nur’Aini Abdul Rashid, and Rosni Abdullah. An enhanced meta-scheduling system for grid computing that considers the job type and priority. Computing, 94(5):389–410, May 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=94&issue=5&spage=389.

Mahafzah:2012:OHH

[2827] Basel A. Mahafzah, Azzam Sleit, Nesreen A. Hamad, Elham F. Ahmad, and Tasneem M. Abu-Kabeer.

The OTIS hyper hexa-cell optoelectronic architecture. Computing, 94 (5):411–432, May 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=94&issue=5&spage=411.

Beliakov:2012:CHB

[2828] G. Beliakov, M. Johnstone, and S. Nahavandi. Computing of high breakdown regression estimators without sorting on graphics processing units. Computing, 94(5):433–447, May 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=94&issue=5&spage=433.

deBlasio:2012:BCN

[2829] Gabriel de Blasio, Arminda Moreno-Díaz, and Roberto Moreno-Díaz. Bioinspired computing nets for directionality in vision. Computing, 94 (5):449–462, May 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=94&issue=5&spage=449.

Nowak:2012:PDG

[2830] Alexander Nowak, Tobias Binz, Christoph Fehling, Oliver Kopp, Frank Leymann, et al. Patterndriven green adaptation of processbased applications and their runtime infrastructure. Computing, 94 (6):463–487, June 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=94&issue=6&spage=463.

Rabhi:2012:AFS

[2831] Fethi A. Rabhi, Lawrence Yao, and Adnene Guabtni. ADAGE: a framework for supporting user-driven ad-hoc data analysis processes. Computing, 94(6):489–519, June 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=94&issue=6&spage=489.

Maojo:2012:NDN

[2832] Victor Maojo, Martin Fritts, Fernando Martin-Sanchez, Diana De la Iglesia, Raul E. Cachau, et al. Nanoinformatics: developing new computing applications for nanomedicine. Computing, 94(6):521–539, June 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=94&issue=6&spage=521.

Mirtalaei:2012:TBB

[2833] Monireh Sadat Mirtalaei, Morteza Saberi, Omar Khadeer Hussain, Behzad Ashjari, and Farookh Khadeer Hussain. A trust-based bio-inspired approach for credit lending decisions. Computing, 94(7):541–577, July 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0010-485X&volume=94&
issue=7&spage=541.

Cugola:2012:DDA

[2834] Gianpaolo Cugola, Carlo Ghezzi, and Leandro Sales Pinto. DSOL: a declarative approach to self-adaptive service orchestrations. Computing, 94 (7):579–617, July 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=94&issue=7&spage=579.

Montresor:2012:EDS

[2835] Alberto Montresor, Guszti Eiben, and Maarten van Steen. Extreme distributed systems: from large scale to complexity. Computing, 94(8–10): 619–620, September 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=94&issue=8&spage=619.

Payberah:2012:DOP

[2836] Amir H. Payberah, Jim Dowling, Fatemeh Rahimain, and Seif Haridi. Distributed optimization of P2P live streaming overlays. Computing, 94(8– 10):621–647, September 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=94&issue=8&spage=621.

Qiao:2012:LBP

[2837] Ying Qiao and Gregor v. Bochmann. Load balancing in peer-to-peer systems using a diffusive approach. Computing, 94(8–10):649–678, September 2012. CODEN CMPTA2.

ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0010-485X&volume=94&
issue=8&spage=649.

Almashor:2012:FPP

[2838] Mahathir Almashor and Ibrahim Khalil. Fully peer-to-peer virtual environments with 3D Voronoi diagrams. Computing, 94(8–10):679– 700, September 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=94&issue=8&spage=679.

Calcavecchia:2012:DDP

[2839] Nicolò M. Calcavecchia, Bogdan A.

Caprarescu, Elisabetta Di Nitto, Daniel J. Dubois, and Dana Petcu. DEPAS: a decentralized probabilistic algorithm for auto-scaling. Computing, 94(8–10):701–730, September 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0010-485X&volume=94&
issue=8&spage=701.

Tsatsanifos:2012:NRL

[2840] George Tsatsanifos and Vasilis Samoladas. A novel robust on-line protocol for load-balancing in structured peerto-peer systems. Computing, 94(8–10): 731–762, September 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=94&issue=8&spage=731.

Naranjo:2012:APP

[2841] J. A. M. Naranjo, L. G. Casado, and Márk Jelasity. Asynchronous privacy-preserving iterative computation on peer-to-peer networks. Computing, 94(8–10):763–782, September 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0010-485X&volume=94&
                                        issue=8&spage=763.

Datta:2012:MRS

[2842] Anwitaman Datta. Merging ringstructured overlay indices: toward network-data transparency. Computing, 94(8–10):783–809, September 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0010-485X&volume=94&
                                        issue=8&spage=783.

Pruteanu:2012:ATN

[2843] Andrei Pruteanu and Stefan Dulman. ASH: tackling node mobility in largescale networks. Computing, 94(8–10): 811–832, September 2012. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=94&issue=8&spage=811.

Mohamadi:2013:LAB

[2844] Hosein Mohamadi, Abdul Samad Bin Haji Ismail, and Shaharuddin Salleh. A learning automata-based algorithm for solving coverage problem in directional sensor networks. Computing, 95(1):1–24, January 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-012-0208-x;   http:
//link.springer.com/article/10.
1007/s00607-012-0208-x/;   http:
//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=95&issue=1&spage=1-24.

Saleh:2013:NPM

[2845] Ahmed I. Saleh. New probabilistic models for managing user’s locations in PCS networks. Computing, 95(1):25–66, January 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-012-0210-3;   http:
//link.springer.com/article/10.
1007/s00607-012-0210-3/;   http:
//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=95&issue=1&spage=25-66. See

retraction notice [3058].

Sharifi:2013:PPA

[2846] Mohsen Sharifi, Saeed Shahrivari, and Hadi Salimi. PASTA: a power-aware solution to scheduling of precedenceconstrained tasks on heterogeneous computing resources. Computing, 95 (1):67–88, January 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-012-0212-1;   http:
//link.springer.com/article/10.
1007/s00607-012-0212-1/;   http:
//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=95&issue=1&spage=67-88.

Solin:2013:Pa

[2847] Pavel Solin, Pavel Karban, Sascha Schnepp, and Jaroslav Kruis. Preface. Computing, 95(1s):1–3, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s00607-
                                        013-0320-6.pdf.

Morales-Castillo:2013:TAE

[2848] Javier Morales-Castillo, Jonás Velasco, and F-Javier Almaguer …. Thermoluminescence analysis and estimation of the kinetics parameters by the Weibull distribution in a diamond like carbon film. Computing, 95(1s):5–11, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-013-
                                        0296-2.

Basting:2013:MBF

[2849] Christopher Basting and Dmitri Kuzmin. A minimization-based finite element formulation for interfacepreserving level set reinitialization. Computing, 95(1s):13–25, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00607-012-0259-
                                        z.

Bittl:2013:AFC

[2850] Melanie Bittl and Dmitri Kuzmin. An hp-adaptive flux-corrected transport algorithm for continuous finite elements. Computing, 95(1s):27–48, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00607-012-
0223-y.

Braun:2013:DAN

[2851] Moritz Braun. Different approaches to the numerical solution of the 3D Poisson equation implemented in Python. Computing, 95(1s):49–60, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0300-x.

Starzynski:2013:BSE

[2852] Jacek Starzyński, Robert Szmurlo, and Bartosz Chaber. Bioelectromagnetic simulation for everyone. Computing, 95(1s):61–74, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00607-012-0278-9.pdf.

Chleboun:2013:ONP

[2853] Jan Chleboun and Pavel Solin. On optimal node and polynomial degree distribution in one-dimensional hpfem. Computing, 95(1s):75–88, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-012-
0232-x.

DAgostino:2013:QAB

[2854] Daniele D’Agostino, Antonella Galizia, Andrea Clematis, and Matteo Mangini …. A QoS-aware broker for hybrid clouds. Computing, 95(1s):89–109, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00607-012-
                                        0254-4.

Zdenek:2013:SEC

[2855] Kubík Zdenek, Karban Pavel, Pánek David, and Skála Jirí. On the shielding effectiveness calculation. Computing, 95(1s):111–121, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00607-013-0312-6.

Emans:2013:VPC

[2856] Maximilian Emans and Manfred Liebmann. Velocity-pressure coupling on GPUs. Computing, 95(1s):123–143, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-012-
                                        0228-6.

Furmanek:2013:DUA

[2857] Petr Furmánek and Karel Kozel. Deterministic unsteady and aeroelastic flow simulations with high-order FVM schemes. Computing, 95(1s):145–161, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-013-
                                        0291-7.

Furst:2013:NST

[2858] Jirí Fürst, Jaromír Príhoda, and Petr Straka. Numerical simulation of transitional flows. Computing, 95 (1s):163–182, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00607-012-0266-0.

Solin:2013:IAH

[2859] Pavel Solin and Stefano Giani. An iterative adaptive hp-FEM method for non-symmetric elliptic eigenvalue problems. Computing, 95(1s):183–213, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-012-
0251-7.

Giani:2013:HOA

[2860] Stefano Giani. High-order/hp-adaptive discontinuous Galerkin finite element methods for acoustic problems. Computing, 95(1s):215–234, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL http://link.springer.com/
article/10.1007/s00607-012-0253-
5.

Giani:2013:ECA

[2861] Stefano Giani, Luka Grubisić, and Jeffrey S. Ovall. Error control for hp-adaptive approximations of semidefinite eigenvalue problems. Computing, 95(1s):235–257, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL http://link.springer.com/
article/10.1007/s00607-012-0260-
6.

Gugala:2013:RNG

[2862] Karol Gugala, Aleksandra Świetlicka, Michal Burdajewicz, and Andrzej Rybarczyk. Random number generation system improving simulations of stochastic models of neural cells. Computing, 95(1s):259–275, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s00607-
                                        012-0267-z.pdf.

Hakula:2013:MIH

[2863] Harri Hakula and Tomi Tuominen. Mathematica implementation of the high order finite element method applied to eigenproblems. Computing, 95(1s):277–301, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00607-012-0262-4.

Halama:2013:NST

[2864] Jan Halama and Jaroslav Fort. Numerical simulation of transonic flow of wet steam in nozzles and turbines. Computing, 95(1s):303–318, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-013-
                                        0292-6.

Giani:2013:PEEa

[2865] Stefano Giani and Edward Hall. An a-posteriori error estimate for hpadaptive DG methods for elliptic eigenvalue problems on anisotropically refined meshes. Computing, 95 (1s):319–341, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00607-012-0261-5.

Hasnedlova:2013:NSF

[2866] Jaroslava Hasnedlová, Miloslav Feistauer, Jaromír Horácek, and Adam Kosík …. Numerical simulation of fluidstructure interaction of compressible flow and elastic structure. Computing, 95(1s):343–361, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-012-0240-x.

Havlicek:2013:IBH

[2867] Vojtech Havlícek, Martin Hanel, Petr Máca, Michal Kuráz, and Pavel Pech. Incorporating basic hydrological concepts into genetic programming for rainfall-runoff forecasting. Computing, 95(1s):363–380, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-013-0298-0.

Karban:2013:NSC

[2868] Pavel Karban, Frantisek Mach, Pavel Kus, David Pánek, and Ivo Dolezel. Numerical solution of coupled problems using code Agros2D. Computing, 95(1s):381–408, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-013-0294-4.

Keslerova:2013:NSS

[2869] Radka Keslerová and Karel Kozel.

Numerical study of steady and unsteady flow for power-law type generalized Newtonian fluids. Computing, 95(1s):409–424, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-013-0301-9.

Korous:2013:ATH

[2870] Lukas Korous and Pavel Solin.

An adaptive hp-DG method with dynamically-changing meshes for nonstationary compressible Euler equations. Computing, 95(1s):425–444, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-012-
                                        0257-1.

Dolezel:2013:NMC

[2871] I. Dolezel, V. Kotlan, and B. Ulrych. Numerical modelling of cylindrical induction shrink fits. Computing, 95(1s):445–458, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00607-012-0274-0.

Kotlan:2013:ECN

[2872] V. Kotlan, L. Voracek, and B. Ulrych. Experimental calibration of numerical model of thermoelastic actuator. Computing, 95(1s):459–472, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-013-
                                        0299-z.

Koudela:2013:MLU

[2873] Lukás Koudela, Pavel Karban, Oldrich Turecek, and Ladislav Zuzjak. Modeling of loudspeaker using hp-adaptive methods. Computing, 95(1s):473–485, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-013-
                                        0303-7.

Kropik:2013:HOF

[2874] Petr Kropík, Lenka Sroubová, and Roman Hamar. Higher-order finite element modeling and optimization of actuator with non-linear materials. Computing, 95(1s):487–499, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0302-8.

Kuraz:2013:DDA

[2875] Michal Kuraz, Petr Mayer, Vojtech Havlicek, and Pavel Pech. Domain decomposition adaptivity for the Richards equation model. Computing, 95(1s):501–519, ???? 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-012-0279-8. Special Is-

sue on ESCO2012.

Lev:2013:MME

[2876] Jakub Lev, Petr Mayer, Marie Wohlmuthová, and Václav Prosek. The mathematical model of experimental sensor for material distribution detecting on the conveyor. Computing, 95(1s):521–536, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-012-0273-1.

Mach:2013:OSI

[2877] Frantisek Mach, Pavel Kus, Pavel Karban, and Ivo Dolezel. Optimization of the system for induction heating of nonmagnetic cylindrical billets in rotating magnetic field produced by permanent magnets. Computing, 95 (1s):537–552, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00607-013-0297-1.

denOuden:2013:ALS

[2878] D. den Ouden, A. Segal, F. J. Vermolen, L. Zhao, C. Vuik, and J. Sietsma. Application of the level-set method to a mixed-mode driven Stefan problem in 2D and 3D. Computing, 95(1s):553–572, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00607-012-0247-3.

Porizkova:2013:FCC

[2879] Petra Porízková, Karel Kozel, and Jaromír Horácek. Flows in convergent channel: comparison of numerical results of different mathematical models. Computing, 95(1s):573–585, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-012-
                                        0269-x.

Prokop:2013:NSG

[2880] Vladimír Prokop and Karel Kozel. Numerical simulation of generalized Oldroyd–B and generalized Newtonian fluid flows. Computing, 95(1s):587–597, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-012-
                                        0281-1.

Rapetti:2013:EQS

[2881] Francesca Rapetti. Electromagnetic quasi-static models applied to transmission lines. Computing, 95 (1s):599–616, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-012-0272-2.

Rieger:2013:NCS

[2882] Janosch Rieger. Non-convex systems of sets for numerical analysis. Computing, 95(1s):617–638, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-012-0241-9.

Campos:2013:CHP

[2883] Ricardo Silva Campos, Fernando Otaviano Campos, Johnny Moreira Gomes, Ciro de Barros Barbosa, Marcelo Lobosco, and Rodrigo Weber dos Santos. Comparing high performance techniques for the automatic generation of efficient solvers of cardiac cell models. Computing, 95(1s):639–660, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-012-
0268-y.

Sawicki:2013:EVM

[2884] Bartosz Sawicki and Bartosz Chaber.

Efficient visualization of 3D models by Web browser. Computing, 95 (1s):661–673, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-012-0275-z;   http:
//link.springer.com/content/pdf/
10.1007/s00607-012-0275-z.pdf.

Schnepp:2013:NPC

[2885] Sascha M. Schnepp. NCLab: Public Computing Laboratory. Computing, 95(1s):675–694, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-012-0255-3.

Segeth:2013:SCA

[2886] Karel Segeth. Some computational aspects of smooth approximation. Computing, 95(1s):695–708, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL http://link.springer.com/
article/10.1007/s00607-012-0252-
6.

Solin:2013:STA

[2887] Pavel Solin and Lukas Korous. Spacetime adaptive hp-FEM for problems with traveling sharp fronts. Computing, 95(1s):709–722, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-012-0243-7.

Mayer:2013:SSE

[2888] Daniel Mayer, Josef Hrusak, and Milan Stork. On state-space energy based generalization of Brayton–Moser topological approach to electrical network decomposition. Computing, 95 (1s):723–749, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-012-0280-2.

Tezer-Sezgin:2013:BSM

[2889] M. Tezer-Sezgin and S. Han Aydin. BEM solution of MHD flow in a pipe coupled with magnetic induction of exterior region. Computing, 95 (1s):751–770, ???? 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-012-0270-4. Special Is-

sue on ESCO2012.

Valkonen:2013:GAR

[2890] Tuomo Valkonen and Manfred Liebmann. GPU-accelerated regularisation of large diffusion-tensor volumes. Computing, 95(1s):771–784, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-012-
0277-x.

Benedict:2013:PIP

[2891] Shajulin Benedict. Performance issues and performance analysis tools for HPC cloud applications: a survey. Computing, 95(2):89– 108, February 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-012-0213-0;   http:
//link.springer.com/article/10.
1007/s00607-012-0213-0/;   http:
//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=95&issue=2&spage=89-108.

Khan:2013:HAH

[2892] Zafar Ali Khan and Won Sohn.

A hierarchical abnormal human activity recognition system based on R-transform and kernel discriminant analysis for elderly health care. Computing, 95(2):109–127, February 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-012-

                                        0216-x;   http://link.springer.
                                        com/article/10.1007/s00607-012-
                                        0216-x/;   http://www.springerlink.
                                        com/openurl.asp?genre=article&issn=
                                        0010-485X&volume=95&issue=2&spage=
                                        109-127.

Cugola:2013:DSD

[2893] Gianpaolo Cugola and Alessandro Margara. Deployment strategies for distributed complex event processing. Computing, 95(2):129– 156, February 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00607-012-0217-9;   http:
                                        //link.springer.com/article/10.
                                        1007/s00607-012-0217-9/;   http:
                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0010-485X&
                                        volume=95&issue=2&spage=129-156.

Khattri:2013:EDF

[2894] Sanjay Kumar Khattri and Torgrim Log. Erratum to: Derivative free algorithm for solving nonlinear equations. Computing, 95(2):157–158, February 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-012-
                                        0221-0;   http://link.springer.
                                        com/content/pdf/10.1007/s00607-
                                        012-0221-0;   http://www.springerlink.
                                        com/openurl.asp?genre=article&issn=
                                        0010-485X&volume=95&issue=2&spage=

157-158. See [2781, 2895, 2896].

Soleymani:2013:LER

[2895] Fazlollah Soleymani. Letter to the editor regarding the article by Khattri: derivative free algorithm for solving nonlinear equations. Computing, 95 (2):159–162, February 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-012-0207-y;   http:
//link.springer.com/article/10.
1007/s00607-012-0207-y/;   http:
//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=95&issue=2&spage=159-162.

See [2781, 2894, 2896].

Khattri:2013:LER

[2896] Sanjay K. Khattri. Letter to the Editor regarding the letter by Soleymani. Computing, 95(2):163– 164, February 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00607-012-0256-
2?coverImageUrl=/static-content%
2F0.6144%2Fcovers%2Fjournals_single_
issue%2F165%2F00607_095_002.jpg;
 http://link.springer.com/article/
10.1007/s00607-012-0256-2;   http:
//www.springerlink.com/openurl.
asp?genre=article&issn=0010-485X&
volume=95&issue=2&spage=163-164.

See [2781, 2894, 2895].

Calinescu:2013:E

[2897] Radu Calinescu, Robert France, and Carlo Ghezzi. Editorial. Computing, 95(3):165–166, March 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-012-0238-4;   http:
//link.springer.com/content/pdf/
10.1007/s00607-012-0238-4.pdf. Spe-
cial issue on models@runtime.

Bencomo:2013:RMT

[2898] N. Bencomo, A. Bennaceur, P. Grace, G. Blair, and V. Issarny. The

                                        role of models@run.time in sup-

porting on-the-fly interoperability. Computing, 95(3):167–190, March 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-012-
                                        0224-x;   http://link.springer.
                                        com/content/pdf/10.1007/s00607-

012-0224-x.pdf. Special issue on

                                        models@runtime.

Mocci:2013:DCC

[2899] Andrea Mocci and Mario Sangiorgio. Detecting component changes at run time with behavior models. Computing, 95(3):191–221, March 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-012-
                                        0214-z;   http://link.springer.
                                        com/content/pdf/10.1007/s00607-

012-0214-z.pdf. Special issue on

                                        models@runtime.

Simmonds:2013:MRW

[2900] Jocelyn Simmonds, Shoham Ben-David, and Marsha Chechik. Monitoring and recovery for Web service applications. Computing, 95(3):223–267, March 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-012-
                                        0215-y;   http://link.springer.
                                        com/content/pdf/10.1007/s00607-

012-0215-y.pdf. Special issue on

                                        models@runtime.

Song:2013:SEC

[2901] Jie Song, Tiantian Li, Zhi Wang, and Zhiliang Zhu. Study on energyconsumption regularities of cloud computing systems by a novel evaluation model. Computing, 95(4):269–287, April 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-012-
0218-8.

Sacha:2013:IFI

[2902] Jan Sacha and Alberto Montresor.

Identifying frequent items in distributed data sets. Computing, 95 (4):289–307, April 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-012-0220-1.

Beliakov:2013:EIBa

[2903] Gleb Beliakov, Michael Johnstone, Doug Creighton, and Tim Wilkin. An efficient implementation of Bailey and Borwein’s algorithm for parallel random number generation on graphics processing units. Computing, 95(4):309–326, April 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-012-0234-8. See also

[4384].

Srinivasan:2013:SIS

[2904] Sujatha Srinivasan and Sivakumar Ramakrishnan. A social intelligent system for multi-objective optimization of classification rules using cultural algorithms. Computing, 95(4):327–350, April 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-012-
                                        0246-4.

Solin:2013:Pb

[2905] Pavel Solin, Pavel Karban, Sascha Schnepp, and Jaroslav Kruis. Preface. Computing, 95(5):351–353, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-013-
                                        0322-4;   http://link.springer.
                                        com/content/pdf/10.1007/s00607-

013-0322-4.pdf. Special Issue on ESCO2012.

Butcher:2013:PCO

[2906] John C. Butcher and Pavla Sehnalová. Predictor-corrector Obreshkov pairs. Computing, 95(5):355–371, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00607-012-0258-

0. Special Issue on ESCO2012.

Abed-Meraim:2013:NQS

[2907] Farid Abed-Meraim, Vuong-Dieu Trinh, and Alain Combescure. New quadratic solid-shell elements and their evaluation on linear benchmark problems. Computing, 95(5):373–394, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-012-

0265-1. Special Issue on ESCO2012.

Giani:2013:PEEb

[2908] Stefano Giani. An a posteriori error estimator for hp-adaptive continuous Galerkin methods for photonic crystal applications. Computing, 95 (5):395–414, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-012-0244-6. Special Is-

sue on ESCO2012.

Brambilla:2013:ATC

[2909] Paola Brambilla and Alberto Guardone.

Automatic tracking of corona propagation in three-dimensional simulations of non-normal drop impact on a liquid film. Computing, 95(5):415–424, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-012-

0271-3. Special Issue on ESCO2012.

Moller:2013:AFC

[2910] Matthias Möller. Algebraic flux correction for nonconforming finite element discretizations of scalar transport problems. Computing, 95(5):425–448, May 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-012-

0276-y. Special Issue on ESCO2012. See erratum [3002].

Karastoyanova:2013:SCS

[2911] Dimka Karastoyanova. Springer computing special issue: adaptation in service-oriented and Cloud Computing. Computing, 95(6):449–451, June 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0319-z.

Ivanovic:2013:SBA

[2912] Dragan Ivanović, Manuel Carro, and Manuel V. Hermenegildo. A sharingbased approach to supporting adaptation in service compositions. Computing, 95(6):453–492, June 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00607-012-0230-
                                        z.

Andrikopoulos:2013:HAA

[2913] Vasilios Andrikopoulos, Tobias Binz, Frank Leymann, and Steve Strauch. How to adapt applications for the Cloud environment. Computing, 95 (6):493–535, June 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00607-012-0248-2.

Cebrian:2013:EIC

[2914] Juan M. Cebrián, Daniel Sánchez, Juan L. Aragón, and Stefanos Kaxiras. Efficient inter-core power and thermal balancing for multicore processors. Computing, 95(7):537–566, July 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-012-
                                        0236-6.

Cao:2013:DMV

[2915] Jing Cao and Albert Nymeyer. A design methodology for verified webservice mediators. Computing, 95 (7):567–610, July 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-012-0237-5.

Malik:2013:RCS

[2916] Sana Malik, Babar Nazir, Kalim Qureshi, and Imran Ali Khan. A reliable checkpoint storage strategy for grid. Computing, 95(7):611–632, July 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-012-
0250-8.

Bertolino:2013:SAB

[2917] Antonia Bertolino, Paola Inverardi, and Henry Muccini. Software architecture-based analysis and testing: a look into achievements and future challenges. Computing, 95(8):633–648, August 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0338-9.

Chen:2013:AAA

[2918] DeJiu Chen, Lei Feng, Tahir Naseer Qureshi, Henrik Lönn, and Frank Hagl. An architectural approach to the analysis, verification and validation of software intensive embedded systems. Computing, 95(8):649–688, August 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0314-4.

Camara:2013:ABR

[2919] Javier Cámara, Rogério de Lemos, Marco Vieira, Raquel Almeida, and Rafael Ventura. Architecture-based resilience evaluation for self-adaptive systems. Computing, 95(8):689–722, August 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-013-
                                        0311-7.

Lasheng:2013:EDC

[2920] Yu Lasheng, Li Jie, and Liu Renjie. An effective data collection algorithm for wireless sensor network. Computing, 95 (9):723–738, September 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00607-012-0249-1.

Chang:2013:FLC

[2921] Maw-Shang Chang, Ling-Ju Hung, Chih-Ren Lin, and Ping-Chen Su. Finding large k-clubs in undirected graphs. Computing, 95(9):739– 758, September 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00607-012-0263-3.

Mashayekhi:2013:GDS

[2922] Hoda Mashayekhi, Jafar Habibi, Spyros Voulgaris, and Maarten van Steen. GoSCAN: Decentralized scalable data clustering. Computing, 95(9):759– 784, September 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00607-012-0264-2.

Zhao:2013:DGS

[2923] Yongwang Zhao, Zhuqing Li, Hualei Shen, and Dianfu Ma. Development of global specification for dynamically adaptive software. Computing, 95(9):

785–816, September 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-013-0295-3.

Stubbings:2013:LOO

[2924] Gary Stubbings and Simon Polovina. Levering object-oriented knowledge for service-oriented proficiency. Computing, 95(9):817–835, September 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0304-6.

Khan:2013:PIH

[2925] Wajahat Ali Khan, Maqbool Hussain, Khalid Latif, Muhammad Afzal, and Farooq Ahmad …. Process interoperability in healthcare systems with dynamic Semantic Web services. Computing, 95(9):837–862, September 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-012-
0239-3.

Nambiar:2013:HSC

[2926] Vishnu P. Nambiar, Sathivellu Balakrishnan, Mohamed Khalil-Hani, and M. N. Marsono. HW/SW co-design of reconfigurable hardware-based genetic algorithm in FPGAs applicable to a variety of problems. Computing, 95(9): 863–896, September 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-013-0305-5.

Ktari:2013:EPS

[2927] Raida Ktari and Habib Chabchoub. Essential particle swarm optimization queen with tabu search for MKP resolution. Computing, 95(9): 897–921, September 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00607-013-0316-2.

Xhafa:2013:SIC

[2928] Fatos Xhafa, Leonard Barolli, and Omar Hussain. Special issue on cyber physical systems. Computing, 95(10– 11):923–926, October 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00607-013-0318-0.

Parvin:2013:MCF

[2929] Sazia Parvin, Farookh Khadeer Hussain, Omar Khadeer Hussain, and Thandar Thein …. Multi-cyber framework for availability enhancement of cyber physical systems. Computing, 95(10–11):927–948, October 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-012-
                                        0227-7.

Al-Khalidi:2013:AAS

[2930] Haidar Al-Khalidi, David Taniar, and Maytham Safar. Approximate algorithms for static and continuous range queries in mobile navigation. Computing, 95(10–11):949–976, October 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00607-012-
0219-7.

Jeong:2013:ANB

[2931] Young-Sik Jeong and Jong Hyuk Park.

Adaptive network-based fuzzy inference model on CPS for large scale intelligent and cooperative surveillance. Computing, 95(10–11):977–992, October 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0317-1.

Wang:2013:EFP

[2932] Jian-Tao Wang, Kam-Yiu Lam, Song Han, Sang H. Son, and Aloysius K. Mok. An effective fixed priority co-scheduling algorithm for periodic update and application transactions. Computing, 95(10–11):993–1018, October 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-012-
0242-8.

Enokido:2013:EET

[2933] Tomoya Enokido and Makoto Takizawa. The evaluation of the extended transmission power consumption (etpc) model to perform communication type processes. Computing, 95(10–11): 1019–1037, October 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-012-0222-z.

Barolli:2013:ASE

[2934] Leonard Barolli, Evjola Spaho, Fatos Xhafa, and Muhammad Younas. Application of SmartBox end-device for medical care using JXTA–Overlay P2P system. Computing, 95(10–11): 1039–1051, October 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-012-0233-9.

Nishino:2013:TIE

[2935] Hiroaki Nishino, Yuki Fukakusa, Tsuneo Kagawa, and Kouichi Utsumiya. A tangible information explorer using vibratory touch screen. Computing, 95(10–11):1053–1071, October 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-012-
0226-8.

Kulla:2013:ERM

[2936] Elis Kulla, Makoto Ikeda, Tetsuya Oda, Leonard Barolli, and Fatos Xhafa …. Experimental results from a MANET testbed in outdoor bridge environment considering BATMAN routing protocol. Computing, 95(10–11): 1073–1086, October 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-012-0225-9.

Sousedik:2013:AMB

[2937] Bedrich Sousedík, Jakub Sístek, and Jan Mandel. Adaptive-multilevel BDDC and its parallel implementation. Computing, 95(12):1087–1119, December 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0293-5.

Hoefler:2013:MMN

[2938] Torsten Hoefler, James Dinan, Darius Buntinas, Pavan Balaji, and Brian Barrett …. MPI + MPI: a new hybrid approach to parallel programming with MPI plus shared memory. Computing, 95(12):1121–1136, December 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0324-2.

Callou:2013:ESI

[2939] Gustavo Callou, Paulo Maciel, Dietmar Tutsch, João Ferreira, and Julian Araújo …. Estimating sustainability impact of high dependable data centers: a comparative study between Brazilian and US energy mixes. Computing, 95(12):1137–1170, December 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0328-y.

Bland:2013:EUL

[2940] Wesley Bland, Aurelien Bouteiller, Thomas Herault, and Joshua Hursey …. An evaluation of User-Level Failure Mitigation support in MPI. Computing, 95(12):1171–1184, December 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0331-3.

Heidemann:2014:IV

[2941] John Heidemann and Walter Willinger. Internet visualization. Computing, 96(1):1–2, January 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-013-0284-6;   http:
//link.springer.com/content/pdf/
10.1007/s00607-013-0284-6.pdf.

Yang:2014:MAN

[2942] Chao Yang, Ian Jensen, and Paul Rosen. A multiscale approach to network event identification using geolocated Twitter data. Computing, 96(1):3–13, January 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-013-0285-5.

Braun:2014:FIF

[2943] Lothar Braun, Mario Volke, Johann Schlamp, Alexander von Bodisco, and Georg Carle. Flow-inspector: a framework for visualizing network flow data using current Web technologies. Computing, 96(1):15–26, January 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0286-4.

Glatz:2014:VBN

[2944] Eduard Glatz, Stelios Mavromatidis, Bernhard Ager, and Xenofontas Dimitropoulos. Visualizing big network traffic data using frequent pattern mining and hypergraphs. Computing, 96 (1):27–38, January 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-013-0282-8.

Quan:2014:VSI

[2945] Lin Quan, John Heidemann, and Yuri Pradkin. Visualizing sparse Internet events: network outages and route changes. Computing, 96(1):39–51, January 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0283-7.

King:2014:CVT

[2946] Alistair King, Bradley Huffaker, Alberto Dainotti, and k. c. Claffy. A coordinated view of the temporal evolution of large-scale Internet events. Computing, 96(1):53–65, January 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0288-2.

Lozzo:2014:VDC

[2947] Giordano Da Lozzo, Giuseppe Di Battista, and Claudio Squarcella. Visual discovery of the correlation between BGP routing and round-trip delay active measurements. Computing, 96(1):67–77, January 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-013-0287-3.

Miyamoto:2014:PII

[2948] Daisuke Miyamoto and Takuji Iimura.

PACKTER: implementation of Internet traffic visualizer and extension for network forensics. Computing, 96 (1):79–80, January 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00607-013-0289-
                                        1;   http://link.springer.com/article/
                                        10.1007/s00607-013-0289-1.

Gansner:2014:DAT

[2949] Emden R. Gansner, Balachander Krishnamurthy, and Walter Willinger …. Demo abstract: towards extracting semantics by visualizing large traceroute datasets. Computing, 96(1):81–83, January 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-013-
                                        0290-8.

Yang:2014:SMC

[2950] Yuhang Yang and Maode Ma. Sustainable mobile computing. Computing, 96(2):85–86, February 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00607-013-0309-1;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00607-013-0309-1.pdf.

Rodriguez:2014:EEJ

[2951] Juan Manuel Rodriguez, Cristian Mateos, and Alejandro Zunino. Energyefficient job stealing for CPU-intensive processing in mobile devices. Computing, 96(2):87–117, February 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00607-012-0245-
                                        5.

Chang:2014:EEG

[2952] Jian-Ming Chang, Chin-Feng Lai, Han-Chieh Chao, and Rongbo Zhu.

An energy-efficient geographic routing protocol design in vehicular adhoc network. Computing, 96(2): 119–131, February 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-012-0235-7.

Ahmed:2014:EES

[2953] Mohammad Helal Uddin Ahmed, Choong Seon Hong, and Sungwon Lee. Energy efficient smart-relay-based cooperative MAC for wireless networks. Computing, 96(2):133–161, February 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-012-
0229-5.

Vu:2014:TEB

[2954] Thi Hong Nhan Vu, Yang Koo Lee, and The Duy Bui. A technique for extracting behavioral sequence patterns from GPS recorded data. Computing, 96(3):163–188, March 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-013-0333-1.

Vo:2014:MPM

[2955] Phuong Luu Vo, Tuan Anh Le, Sungwon Lee, Choong Seon Hong, Byeongsik Kim, and Hoyoung Song. mReno: a practical multipath congestion control for communication networks. Computing, 96(3):189–205, March 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0341-1.

Yu:2014:SSA

[2956] Dung-Feng Yu and Hewijin Christine Jiau. SeaDoc: a self-adaptive document link provision system for framework extension tasks. Computing, 96 (3):207–222, March 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00607-013-0344-y.

Mettouris:2014:URS

[2957] Christos Mettouris and George A. Papadopoulos. Ubiquitous recommender systems. Computing, 96(3):223–257, March 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-013-
                                        0351-z.

Traff:2014:SPE

[2958] Jesper Larsson Träff and Siegfried Benkner. Selected papers from EuroMPI 2012. Computing, 96 (4):259–261, April 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00607-013-0335-z.

Didelot:2014:IMC

[2959] Sylvain Didelot, Patrick Carribault, Marc Pérache, and William Jalby. Improving MPI communication overlap with collaborative polling. Computing, 96(4):263–278, April 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00607-013-0327-z.

Schneider:2014:AOP

[2960] Timo Schneider, Robert Gerstenberger, and Torsten Hoefler. Applicationoriented ping-pong benchmarking: how to assess the real communication overheads. Computing, 96(4):279–292, April 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0330-4.

Kamal:2014:IFG

[2961] Humaira Kamal and Alan Wagner. An integrated fine-grain runtime system for MPI. Computing, 96(4):293–309, April 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0329-x.

Anonymous:2014:EN

[2962] Anonymous. Editorial note. Computing, 96(4):311, April 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-013-0380-7;   http:
//link.springer.com/content/pdf/
10.1007/s00607-013-0380-7.pdf.

Abdunabi:2014:EST

[2963] Ramadan Abdunabi, Wuliang Sun, and Indrakshi Ray. Enforcing spatiotemporal access control in mobile applications. Computing, 96(4):313–353, April 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0340-2.

Tamayo:2014:PCG

[2964] Alain Tamayo, Carlos Granell, Laura Díaz, and Joaquín Huerta. Personalised code generation from large schema sets for geospatial mobile applications. Computing, 96(5):355–379, May 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0339-8.

El-Bendary:2014:AIA

[2965] Nashwa El-Bendary, Tai hoon Kim, Aboul Ella Hassanien, and Mohamed Sami. Automatic image annotation approach based on optimization of classes scores. Computing, 96(5):381–402, May 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0342-0.

Nasridinov:2014:DTC

[2966] Aziz Nasridinov, Yangsun Lee, and Young-Ho Park. Decision tree construction on GPU: ubiquitous parallel computing approach. Computing, 96(5):403–413, May 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-013-0343-z.

Qin:2014:PES

[2967] Xiulei Qin, Wei Wang, Wenbo Zhang, Jun Wei, Xin Zhao, Hua Zhong, and Tao Huang. PRESC^{2}: efficient selfreconfiguration of cache strategies for elastic caching platforms. Computing, 96(5):415–451, May 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-013-0365-6.

Hu:2014:ISI

[2968] Jia Hu, Seetharami Seelam, and Laurent Lefevre. Introduction to special issue on high performance computing. Computing, 96(6): 453–454, June 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00607-014-0402-
0;   http://link.springer.com/article/
10.1007/s00607-014-0402-0.

Jiang:2014:NCM

[2969] Wenbin Jiang, Lei Zhang, Xiaofei Liao, Hai Jin, and Yaqiong Peng. A novel clustered MongoDB-based storage system for unstructured data with high availability. Computing, 96(6):455–478, June 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0355-8.

Li:2014:PLB

[2970] Qinjian Li, Chengwen Zhong, Kai Li, Guangyong Zhang, Xiaowei Lu, Qing Zhang, Kaiyong Zhao, and Xiaowen Chu. A parallel lattice Boltzmann method for large eddy simulation on multiple GPUs. Computing, 96(6):479–501, June 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-013-0356-7.

Yang:2014:SRC

[2971] Menghui Yang, Geyong Min, Weikang Yang, and Zituo Li. Software rejuvenation in cluster computing systems with dependency between nodes. Computing, 96(6):503–526, June 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00607-014-0385-
                                        x.

Liu:2014:SAE

[2972] Weijiang Liu, Wenyu Qu, and Zhaobin Liu. Scalable algorithms for estimating flow length distributions from sampled data. Computing, 96(6):527–543, June 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-014-
                                        0386-9.

Ye:2014:PTL

[2973] John Ye, Hui Yan, Honglun Hou, and Tianzhou Chen. Potential thread-level-parallelism exploration with superblock reordering. Computing, 96(6):545–564, June 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-014-
                                        0387-8.

Yu:2014:AMW

[2974] Jian Yu, Quan Z. Sheng, Muhammad Younas, and Elhadi Shakshuki. Advances in mobile web information systems. Computing, 96(7):565–567, July 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-014-

0384-y;   http://link.springer.
com/content/pdf/10.1007/s00607-
014-0384-y.pdf.

Wang:2014:EEL

[2975] Lian Wang, Yong Cui, Ivan Stojmenovic, Xiao Ma, and Jian Song. Energy efficiency on location based applications in mobile cloud computing: a survey. Computing, 96(7):569–585, July 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0334-0.

Isaac:2014:DIP

[2976] Jesús Téllez Isaac and Sherali Zeadally.

Design, implementation, and performance analysis of a secure payment protocol in a payment gateway centric model. Computing, 96(7):587–611, July 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0306-4.

Biswas:2014:PPT

[2977] Debmalya Biswas and Krishnamurthy Vidyasankar. Privacy preserving and transactional advertising for mobile services. Computing, 96(7):613–630, July 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0332-2.

Li:2014:XME

[2978] Jianxin Li, Chengfei Liu, and Jiajie Xu. XBridge-Mobile: efficient XML keyword search on mobile web data. Computing, 96(7):631–650, July 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-013-
                                        0315-3.

Osman:2014:UST

[2979] Taha Osman, Dhavalkumar Thakker, and Gerald Schaefer. Utilising semantic technologies for intelligent indexing and retrieval of digital images. Computing, 96(7):651–668, July 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-013-
                                        0350-0.

Walraven:2014:CPO

[2980] Stefan Walraven, Eddy Truyen, and Wouter Joosen. Comparing PaaS offerings in light of SaaS development. Computing, 96(8):669–724, August 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-013-
                                        0346-9.

Zhou:2014:MPD

[2981] Yu Zhou, Xiaoxing Ma, and Harald Gall. A middleware platform for the dynamic evolution of distributed component-based systems. Computing, 96(8):725–747, August 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00607-014-0396-7.

Das:2014:GAN

[2982] Subhra Kanti Das, Chandan Mazumdar, and Kumardeb Banerjee. GPU accelerated novel particle filtering method. Computing, 96(8):749–773, August 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-014-
0400-2.

Kim:2014:ESI

[2983] Tai hoon Kim, Carlos Ramos, and Sabah Mohammed. Editorial to special issue: computer and network applications for ubiquitous computing. Computing, 96(9):775–776, September 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-014-
0422-9;   http://link.springer.
com/content/pdf/10.1007/s00607-
014-0422-9.pdf.

Villalba:2014:DED

[2984] L. J. García Villalba, A. L. Sandoval Orozco, J. García Matesanz, and T.-H. Kim. E-D2HCP: enhanced distributed dynamic host configuration protocol. Computing, 96(9):777–791, September 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0307-3.

Khan:2014:MEK

[2985] Muhammad Khurram Khan, Saru Kumari, and Mridul K. Gupta. More efficient key-hash based fingerprint remote authentication scheme using mobile device. Computing, 96(9): 793–816, September 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-013-0308-2.

Canas:2014:ARP

[2986] Delfín Rupérez Cañas, Luis Javier García Villalba, Ana Lucila Sandoval Orozco, and Tai-Hoon Kim. Adaptive routing protocol for mobile ad hoc networks. Computing, 96(9):817–827, September 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-013-
                                        0310-8.

Orozco:2014:SIM

[2987] A. L. Sandoval Orozco, D. M. Arenas González, J. Rosales Corripio, L. J. García Villalba, and J. C. Hernandez-Castro. Source identification for mobile devices, based on wavelet transforms combined with sensor imperfections. Computing, 96(9): 829–841, September 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00607-013-0313-5.

Li:2014:IBD

[2988] Fagen Li, Pan Xiong, and Chunhua Jin. Identity-based deniable authentication for ad hoc networks. Computing, 96 (9):843–853, September 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00607-013-0321-5.

Zhang:2014:DBI

[2989] Weizhe Zhang, You Zhang, and Tai-Hoon Kim. Detecting bad information in mobile wireless networks based on the wireless application protocol. Computing, 96(9):855–874, September 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0325-1.

Kong:2014:RTD

[2990] Quan Kong and Takuya Maekawa.

Reusing training data with generative/discriminative hybrid model for practical acceleration-based activity recognition. Computing, 96(9):875– 895, September 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-013-0326-0.

Han:2014:UMB

[2991] Soyeon Caren Han, Hee-Geun Yoon, Byeong Ho Kang, and Seong-Bae Park. Using MCRDR based Agile approach for expert system development. Computing, 96(9):897–908, September 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0336-y.

Hasanzadeh:2014:GRD

[2992] Mohammad Hasanzadeh and Mohammad Reza Meybodi. Grid resource discovery based on distributed learning automata. Computing, 96(9): 909–922, September 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-013-0337-x.

Yang:2014:GE

[2993] Laurence T. Yang, Bernady O. Apduhan, Tadashi Dohi, and Manish Parashar. Guest editorial.

Computing, 96(10):923–924, October 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-014-
0425-6;   http://link.springer.
com/content/pdf/10.1007/s00607-
014-0425-6.pdf.

Yau:2014:ADT

[2994] Stephen S. Yau, Yisheng Yao, and Arun Balaji Buduru. An adaptable distributed trust management framework for large-scale secure service-based systems. Computing, 96(10):925–949, October 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0354-9.

Pun:2014:GIT

[2995] Chi-Man Pun and Ka-Cheng Choi. Generalized integer transform based reversible watermarking algorithm using efficient location map encoding and adaptive thresholding. Computing, 96 (10):951–973, October 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-013-0357-6.

Kanso:2014:CRM

[2996] Ali Kanso, Maria Toeroe, and Ferhat Khendek. Comparing redundancy models for high availability middleware. Computing, 96(10):975–993, October 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0361-x.

Dong:2014:RTF

[2997] Mianxiong Dong, Kaoru Ota, He Li, Suguo Du, Haojin Zhu, and Song Guo. RENDEZVOUS: towards fast event detecting in wireless sensor and actor networks. Computing, 96(10): 995–1010, October 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-013-0364-7.

Martelot:2014:FMS

[2998] Erwan Le Martelot and Chris Hankin.

Fast multi-scale detection of overlapping communities using local criteria. Computing, 96(11):1011–1027, November 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-014-
0401-1.

Jahanshahi:2014:MRP

[2999] Mohsen Jahanshahi and Alireza Talebi Barmi. Multicast routing protocols in wireless mesh networks: a survey. Computing, 96(11):1029–1057, November 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-014-
0403-z.

Frincu:2014:ESV

[3000] Marc E. Frincu, Stéphane Genaud, and Julien Gossa. On the efficiency of several VM provisioning strategies for workflows with multi-threaded tasks on clouds. Computing, 96(11): 1059–1086, November 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00607-014-0410-0.

Kumar:2014:ISM

[3001] Neetesh Kumar and Deo Prakash Vidyarthi. Improved scheduler for multi-core many-core systems. Computing, 96(11):1087–1110, November 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-014-
                                        0420-y.

Moller:2014:EAF

[3002] Matthias Möller. Erratum to: “Algebraic flux correction for nonconforming finite element discretizations of scalar transport problems”. Computing, 96(11):1111–1112, November 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-014-
                                        0389-6;   http://link.springer.
                                        com/content/pdf/10.1007/s00607-
                                        014-0389-6.pdf. See [2910].

Pllana:2014:ICS

[3003] Sabri Pllana and Jacob Barhen. Introduction to the computing special issue: performance portability and tuning for multi-core and many-core computing systems. Computing, 96(12): 1113–1114, December 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00607-013-0379-0;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00607-013-0379-0.pdf.

Bigot:2014:LLC

[3004] Julien Bigot, Zhengxiong Hou, Christian Pérez, and Vincent Pichon. A low level component model easing performance portability of HPC applications. Computing, 96(12):1115–1130, December 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0368-3.

Mushtaq:2014:EHP

[3005] Hamid Mushtaq, Zaid Al-Ars, and Koen Bertels. Efficient and highly portable deterministic multithreading (DetLock). Computing, 96(12): 1131–1147, December 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-013-0370-9.

Ghosh:2014:PCG

[3006] Sayan Ghosh, Terrence Liao, Henri Calandra, and Barbara M. Chapman. Performance of CPU/GPU compiler directives on ISO/TTI kernels. Computing, 96(12):1149–1162, December 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0367-4.

Kim:2014:ECD

[3007] Shin gyu Kim, Hyeonsang Eom, Heon Y. Yeom, and Sang Lyul Min. Energy-centric DVFS controlling method for multi-core platforms. Computing, 96(12):1163–1177, December 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-013-
                                        0369-2.

Cebrian:2014:PEI

[3008] Juan M. Cebrián, Lasse Natvig, and Jan Christian Meyer. Performance and energy impact of parallelization and vectorization techniques in modern microprocessors. Computing, 96(12): 1179–1193, December 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00607-013-0366-5.

Dastgeer:2014:PCT

[3009] Usman Dastgeer, Lu Li, and Christoph Kessler. The PEPPHER composition tool: performance-aware composition for GPU-based systems. Computing, 96(12):1195–1211, December 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-013-
                                        0371-8.

Damiani:2015:E

[3010] Ernesto Damiani, Achim P. Karduck, and Moataz A. Ahmed. Editorial. Computing, 97(1):1–2, January 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-013-
                                        0363-8;   http://link.springer.
                                        com/content/pdf/10.1007/s00607-
                                        013-0363-8.pdf.

Shinozaki:2015:CBC

[3011] Tetsuo Shinozaki, Yukiko Yamamoto, and Setsuo Tsuruta. Context-based counselor agent for software development ecosystem. Computing, 97 (1):3–28, January 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-013-0352-y.

Dikow:2015:IAB

[3012] Heidi Dikow, Omar Hasan, Harald Kosch, Lionel Brunie, and Renaud Sornin. Improving the accuracy of Business-to-Business (B2B) reputation systems through rater expertise prediction. Computing, 97(1):29–49, January 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0345-x.

Ardagna:2015:DCS

[3013] Claudio A. Ardagna, Ravi Jhawar, and Vincenzo Piuri. Dependability certification of services: a model-based approach. Computing, 97(1):51–78, January 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00607-013-
0348-7.

Dong:2015:SRC

[3014] Hai Dong and Farookh Khadeer Hussain. Service-requester-centered service selection and ranking model for digital transportation ecosystems. Computing, 97(1):79–102, January 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/article/10.
1007/s00607-013-0347-8.

Cugola:2015:IUC

[3015] Gianpaolo Cugola, Alessandro Margara, Matteo Matteucci, and Giordano Tamburrelli. Introducing uncertainty in complex event processing: model, implementation, and validation. Computing, 97(2):103–144, February 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00607-014-
                                        0404-y.

Kohek:2015:ISS

[3016] Stefan Kohek and Damjan Strnad. Interactive synthesis of self-organizing tree models on the GPU. Computing, 97(2):145–169, February 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00607-014-0424-
                                        7.

Oggier:2015:SRC

[3017] Frédérique Oggier and Anwitaman Datta. Self-repairing codes. Computing, 97(2):171–201, February 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00607-014-0426-
                                        5.

Choi:2015:ISI

[3018] Lynn Choi and Raimund Kirner. Introduction to the special issue on software technologies for embedded and ubiquitous systems. Computing, 97 (3):203–204, March 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00607-014-0437-2.pdf.

Radi:2015:LLO

[3019] Marjan Radi, Behnam Dezfouli, Kamalrulnizam Abu Bakar, Shukor Abd Razak, and Malrey Lee. LINKORD: link ordering-based data gathering protocol for wireless sensor networks. Computing, 97(3):205–236, March 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lee:2015:ISF

[3020] Sang Hoon Lee, Yunmook Nah, and Lynn Choi. Influenza surveillance and forecast with smartphone sensors. Computing, 97(3):237–259, March 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kandl:2015:RMD

[3021] Susanne Kandl and Sandeep Chandrashekar. Reasonability of MC/ DC for safety-relevant software implemented in programming languages with short-circuit evaluation. Computing, 97(3):261–279, March 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00607-014-0418-5.pdf.

Motika:2015:LWS

[3022] Christian Motika and Reinhard von Hanxleden. Light-weight Synchronous Java (SJL): An approach for programming deterministic reactive systems with Java. Computing, 97(3):281–307, March 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Reshmi:2015:FBA

[3023] T. R. Reshmi and K. Murugan. Filterbased address autoconfiguration protocol (FAACP) for duplicate address detection and recovery in MANETs. Computing, 97(3):309–331, March 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chen:2015:GES

[3024] Jinjun Chen and Surya Nepal. Guest editorial: Special issue on data intensive cloud computing. Computing, 97(4):333–335, April 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s00607-015-0449-6.pdf.

Wang:2015:RTD

[3025] Mingzhong Wang, Liehuang Zhu, and Kotagiri Ramamohanarao. Reasoning task dependencies for robust service selection in data intensive workflows. Computing, 97(4):337–355, April 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Alhamazani:2015:OCC

[3026] Khalid Alhamazani, Rajiv Ranjan, Karan Mitra, Fethi Rabhi, Prem Prakash Jayaraman, Samee Ullah Khan, Adnene Guabtni, and Vasudha Bhatnagar. An overview of the commercial cloud monitoring tools: research dimensions, design issues, and state-of-the-art. Computing, 97 (4):357–377, April 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wang:2015:AEA

[3027] Yang Wang, Menglan Hu, and Kenneth B. Kent. ACS: an effective admission control scheme with deadlock resolutions for workflow scheduling in clouds. Computing, 97(4):379–402, April 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sun:2015:MBL

[3028] Yu Sun, Jianzhong Qi, Rui Zhang, Yueguo Chen, and Xiaoyong Du. MapReduce based location selection algorithm for utility maximization with capacity constraints. Computing, 97 (4):403–423, April 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Huang:2015:UAA

[3029] Mao Lin Huang, Liang Fu Lu, and Xuyun Zhang. Using arced axes in parallel coordinates geometry for high dimensional BigData visual analytics in cloud computing. Computing, 97 (4):425–437, April 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mollamotalebi:2015:TTB

[3030] Mahdi Mollamotalebi, Raheleh Maghami, and Abdul Samad Ismail. THRD: Threshold-based hierarchical resource discovery for grid environments. Computing, 97(5):439–458, May 2015. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lira:2015:DVN

[3031] Victor Lira, Eduardo Tavares, Stenio Fernandes, and Paulo Maciel. Dependable virtual network mapping. Computing, 97(5):459–481, May 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Karimi-Majd:2015:HAI

[3032] Amir-Mohsen Karimi-Majd, Mohammad Fathian, and Babak Amiri. A hybrid artificial immune network for detecting communities in complex networks. Computing, 97(5):483–507, May 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Han:2015:WAQ

[3033] Jingyu Han, Kejia Chen, and Jianing Wang. Web article quality ranking based on web community knowledge. Computing, 97(5):509–537, May 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Daneshtalab:2015:SIC

[3034] Masoud Daneshtalab, Nader Bagherzadeh, and Hamid Sarbazi-Azad. Special issue on on-chip parallel and network-based systems. Computing, 97(6):539–541, June 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s00607-
                                        015-0458-5.pdf.

Kim:2015:DTS

[3035] Sang Cheol Kim and Sunggu Lee. Decentralized task scheduling for a fixed priority multicore embedded RTOS. Computing, 97(6):543–555, June 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sklyarov:2015:DIC

[3036] Valery Sklyarov and Iouliia Skliarova.

Design and implementation of counting networks. Computing, 97(6):557–577, June 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Demiriz:2015:UCP

[3037] Ayhan Demiriz, Nader Bagherzadeh, and Abdulaziz Alhussein. Using constraint programming for the design of network-on-chip architectures. Computing, 97(6):579–592, June 2015. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ozturk:2015:ERN

[3038] Ozcan Ozturk, Ismail Akturk, Ismail Kadayif, and Suleyman Tosun. Energy reduction in 3D NoCs through communication optimization. Computing, 97(6):593–609, June 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Neishaburi:2015:SCF

[3039] M. H. Neishaburi and Zeljko Zilic. System on chip failure rate assessment using the executable model of a system. Computing, 97(6):611–629, June 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ebrahimi:2015:LWF

[3040] Masoumeh Ebrahimi and Masoud Daneshtalab. A light-weight faulttolerant routing algorithm tolerating faulty links and routers. Computing, 97(6):631–648, June 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bavarsad:2015:TIA

[3041] Amir Ghanbari Bavarsad and Ehsan Atoofian. TurboLock: increasing associativity of lock table in transactional memory. Computing, 97(6):649–661, June 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bellatreche:2015:SIC

[3042] Ladjel Bellatreche, Abdelmalek Amine, and Otmane Ait Mohamed. Special issue on contributions of computational intelligence in designing complex information systems. Computing, 97(7):663–666, July 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s00607-015-0463-8.pdf.

Abbas:2015:SCA

[3043] Assad Abbas, Limin Zhang, and Samee U. Khan. A survey on contextaware recommender systems based on computational intelligence techniques. Computing, 97(7):667–690, July 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ouchani:2015:SVQ

[3044] Samir Ouchani and Mourad Debbabi. Specification, verification, and quantification of security in model-based systems. Computing, 97(7):691–711, July 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mohand-Oussaid:2015:FMO

[3045] Linda Mohand-Oussaid, Idir Ait-Sadoune, Yamine Ait-Ameur, and Mohamed Ahmed-Nacer. A formal model for output multimodal HCI. Computing, 97(7):713–740, July 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Yu:2015:EFA

[3046] Shuhao Yu, Shenglong Zhu, Yan Ma, and Demei Mao. Enhancing firefly algorithm using generalized oppositionbased learning. Computing, 97(7):741– 754, July 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zhang:2015:PDA

[3047] Jing Zhang and Hang Lei. A pre-distribution algorithm of component reliability in Internetware system. Computing, 97(7):755–768, July 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lee:2015:SMS

[3048] Dongeun Lee, Junhee Ryu, and Heonshik Shin. Scalable management of storage for massive quality-adjustable sensor data. Computing, 97(8):769– 793, August 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mo:2015:NMM

[3049] Tieqiang Mo and Renfa Li. A new memory mapping mechanism for GPG-PUs’ stencil computation. Computing, 97(8):795–812, August 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mora-Gutierrez:2015:AMM

[3050] Roman Anselmo Mora-Gutiérrez, María E. Lárraga-Ramírez, Eric A. Rincón-García, Antonin Ponsich, and Javier Ramírez-Rodríguez. Adaptation of the method of musical composition for solving the multiple sequence alignment problem. Computing, 97(8):813–842, August 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Najmi:2015:CCA

[3051] Erfan Najmi, Khayyam Hashmi, Zaki Malik, Abdelmounaam Rezgui, and Habib Ullah Khan. CAPRA: a comprehensive approach to product ranking using customer reviews. Computing, 97(8):843–867, August 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gramoli:2015:SIN

[3052] Vincent Gramoli and Rachid Guerraoui. Special issue on NETYS selected papers. Computing, 97(9): 869–870, September 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00607-015-0466-5.pdf.

Millan:2015:IRL

[3053] Víctor M. López Millán, Vicent Cholvi, Luis López, and Antonio Fernández Anta. Improving resource location with locally precomputed partial random walks. Computing, 97(9):871–891, September 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Benkaouz:2015:DPP

[3054] Yahya Benkaouz and Mohammed Erradi. A distributed protocol for privacy preserving aggregation with nonpermanent participants. Computing, 97 (9):893–912, September 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Abid:2015:APP

[3055] Mohamed Riduan Abid and Saâd Biaz.

Airtime ping-pong effect characterization in IEEE 802.11s wireless mesh networks. Computing, 97(9):913–937, September 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Didona:2015:IOL

[3056] Diego Didona, Pascal Felber, Derin Harmanci, Paolo Romano, and Jörg Schenker. Identifying the optimal level of parallelism in transactional memory applications. Computing, 97(9): 939–959, September 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Boutet:2015:FAD

[3057] Antoine Boutet, Davide Frey, Arnaud Jégou, Anne-Marie Kermarrec, and Heverson B. Ribeiro. FreeRec: an anonymous and distributed personalization architecture. Computing, 97 (9):961–980, September 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Saleh:2015:RNN

[3058] Ahmed I. Saleh. Retraction note to:

New probabilistic models for managing user’s locations in PCS networks. Computing, 97(9):981, September 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00607-
015-0469-2.pdf. See [2845].

Daniel:2015:MWI

[3059] Florian Daniel, George A. Papadopoulos, and Philippe Thiran. Mobile web information systems (MobiWIS 2013). Computing, 97(10):983–984, October 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s00607-
                                        014-0423-8.pdf.

Gronli:2015:TNP

[3060] Tor-Morten Grønli, Pardis Pourghomi, and Gheorghita Ghinea. Towards NFC payments using a lightweight architecture for the Web of Things. Computing, 97(10):985–999, October 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Corral:2015:SAP

[3061] Luis Corral, Alberto Sillitti, and Giancarlo Succi. Software assurance practices for mobile applications. Computing, 97(10):1001–1022, October 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chaniotis:2015:NJV

[3062] Ioannis K. Chaniotis, Kyriakos-Ioannis D. Kyriakou, and Nikolaos D. Tselikas. Is Node.js a viable option for building modern web applications? A performance evaluation study. Computing, 97(10):1023–1044, October 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Yau:2015:ARL

[3063] Kok-Lim Alvin Yau, Hock Guan Goh, David Chieng, and Kae Hsiang Kwong. Application of reinforcement learning to wireless sensor networks: models and algorithms. Computing, 97(11): 1045–1075, November 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Cebrian:2015:PVP

[3064] Juan M. Cebrian, Magnus Jahre, and Lasse Natvig. ParVec: vectorizing the PARSEC benchmark suite. Computing, 97(11):1077–1100, November 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Shahrivari:2015:DDF

[3065] Saeed Shahrivari and Saeed Jalili. Distributed discovery of frequent subgraphs of a network using MapReduce. Computing, 97(11):1101–1120, November 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dantas:2015:EBP

[3066] Jamilson Dantas, Rubens Matos, Jean Araujo, and Paulo Maciel. Eucalyptusbased private clouds: availability modeling and comparison to the cost of a public cloud. Computing, 97(11): 1121–1140, November 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fragkoulis:2015:ISR

[3067] Marios Fragkoulis, Diomidis Spinellis, and Panos Louridas. An interactive SQL relational interface for querying main-memory data structures. Computing, 97(12):1141–1164, December 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ma:2015:DAV

[3068] Tinghuai Ma, Yuliang Zhang, Jie Cao, Jian Shen, Meili Tang, Yuan Tian, Abdullah Al-Dhelaan, and Mznah Al-Rodhaan. KDVEM: a k-degree anonymity with vertex and edge modification algorithm. Computing, 97(12): 1165–1184, December 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Khan:2015:SBA

[3069] Kifayat Ullah Khan, Waqas Nawaz, and Young-Koo Lee. Set-based approximate approach for lossless graph summarization. Computing, 97(12): 1185–1207, December 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bahrpeyma:2015:ARB

[3070] Fouad Bahrpeyma, Hassan Haghighi, and Ali Zakerolhosseini. An adaptive RL based approach for dynamic resource provisioning in cloud virtualized data centers. Computing, 97(12): 1209–1234, December 2015. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ranjan:2016:NST

[3071] Rajiv Ranjan, Dimitrios Georgakopoulos, and Lizhe Wang. A note on software tools and technologies for delivering smart media-optimized big data applications in the cloud. Computing, 98(1–2):1–5, January 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00607-015-0471-8.pdf.

Ma:2016:PPT

[3072] Yan Ma, Lajiao Chen, Peng Liu, and Ke Lu. Parallel programming templates for remote sensing image processing on GPU architectures: design and implementation. Computing, 98 (1–2):7–33, January 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Xu:2016:SEC

[3073] Zheng Xu, Lin Mei, Yunhuai Liu, Chuanping Hu, and Lan Chen. Semantic enhanced cloud environment for surveillance data management using video structural description. Computing, 98(1–2):35–54, January 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zhao:2016:ECM

[3074] Jiaqi Zhao, Jie Tao, and Achim Streit.

Enabling collaborative MapReduce on the cloud with a single-sign-on mechanism. Computing, 98(1–2):55–72, January 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rocha:2016:HCP

[3075] Vladimir Rocha, Fabio Kon, Raphael Cobe, and Renata Wassermann. A hybrid cloud-P2P architecture for multimedia information retrieval on VoD services. Computing, 98(1–2):73–92, January 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Yang:2016:CDD

[3076] Chao-Tung Yang, Wen-Chung Shih, Chih-Lin Huang, Fuu-Cheng Jiang, and William Cheng-Chung Chu. On construction of a distributed data storage system in cloud. Computing, 98 (1–2):93–118, January 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Song:2016:TSA

[3077] Biao Song, Mohammad Mehedi Hassan, Atif Alamri, Abdulhameed Alelaiwi, Yuan Tian, Mukaddim Pathan, and Ahmad Almogren. A two-stage approach for task and resource management in multimedia cloud environment. Computing, 98(1–2):119–145, January 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pop:2016:ASA

[3078] Florin Pop, Ciprian Dobre, Dragos Comaneci, and Joanna Kolodziej. Adaptive scheduling algorithm for mediaoptimized traffic management in software defined networks. Computing, 98 (1–2):147–168, January 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Jang-Jaccard:2016:WBV

[3079] Julian Jang-Jaccard, Surya Nepal, Branko Celler, and Bo Yan. WebRTCbased video conferencing service for telehealth. Computing, 98(1–2):169– 193, January 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Qi:2016:WPC

[3080] Lianyong Qi, Wanchun Dou, and Jinjun Chen. Weighted principal component analysis-based service selection method for multimedia services in cloud. Computing, 98(1–2):195– 214, January 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chen:2016:HIA

[3081] Jingying Chen, Nan Luo, Yuanyuan Liu, Leyuan Liu, Kun Zhang, and Joanna Kolodziej. A hybrid intelligence-aided approach to affectsensitive e-learning. Computing, 98(1– 2):215–233, January 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lin:2016:FDA

[3082] Kawuu W. Lin, Sheng-Hao Chung, and Chun-Cheng Lin. A fast and distributed algorithm for mining frequent patterns in congested networks. Computing, 98(3):235–256, March 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Nawaz:2016:PAW

[3083] Faiza Nawaz and Varun Jeoti. Performance assessment of WirelessHART technology for its implementation in dense reader environment. Computing, 98(3):257–277, March 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wang:2016:EAA

[3084] Guohui Wang, Yong Guan, Yi Wang, and Zili Shao. Energy-aware assignment and scheduling for hybrid main memory in embedded systems. Computing, 98(3):279–301, March 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Li:2016:EEM

[3085] Hongjian Li, Guofeng Zhu, Chengyuan Cui, Hong Tang, Yusheng Dou, and Chen He. Energy-efficient migration and consolidation algorithm of virtual machines in data centers for cloud computing. Computing, 98(3):303–317, March 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Datta:2016:SPN

[3086] Anwitaman Datta, Lluis Pamies-Juarez, and Frédérique Oggier. A study of the performance of novel storagecentric repairable codes. Computing, 98(3):319–341, March 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dumas:2016:GES

[3087] Marlon Dumas, Richard Hull, and Fabio Patrizi. Guest editorial: special issue on data and artifact-centric business processes. Computing, 98 (4):343–344, April 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s00607-016-0486-9.pdf.

Eshuis:2016:SDC

[3088] Rik Eshuis and Pieter Van Gorp. Synthesizing data-centric models from business process models. Computing, 98(4):345–373, April 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s00607-015-0442-0.pdf.

Haddar:2016:DCA

[3089] Nahla Haddar, Mohamed Tmar, and Faiez Gargouri. A data-centric approach to manage business processes. Computing, 98(4):375–406, April 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mannhardt:2016:BMP

[3090] Felix Mannhardt, Massimiliano de Leoni, Hajo A. Reijers, and Wil M. P. van der Aalst. Balanced multiperspective checking of process conformance. Computing, 98(4):407–437, April 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00607-
015-0441-1.pdf.

Weiss:2016:ECM

[3091] Andreas Weiß and Dimka Karastoyanova. Enabling coupled multi-scale, multi-field experiments through choreographies of data-driven scientific simulations. Computing, 98(4):439–467, April 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ardagna:2016:E

[3092] Danilo Ardagna and Lutz Schubert.

Editorial. Computing, 98(5):469–470, May 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00607-
015-0443-z.pdf.

Canuto:2016:IPM

[3093] Mauro Canuto and Jordi Guitart. Integrated policy management framework for IaaS cloud middleware. Computing, 98(5):471–494, May 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pacini:2016:MOS

[3094] Elina Pacini, Cristian Mateos, and Carlos García Garino. Multi-objective swarm intelligence schedulers for online scientific clouds. Computing, 98 (5):495–522, May 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wei:2016:PVR

[3095] Yi Wei and M. Brian Blake. Proactive virtualized resource management for service workflows in the cloud. Computing, 98(5):523–538, May 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Paraiso:2016:SSO

[3096] Fawaz Paraiso, Philippe Merle, and Lionel Seinturier. soCloud: a serviceoriented component-based PaaS for managing portability, provisioning, elasticity, and high availability across multiple clouds. Computing, 98(5):539– 565, May 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Vinko:2016:EIB

[3097] Tamás Vinkó and Bea Botyánszki. Empirical investigation of BitTorrent community graphs. Computing, 98(6):567– 582, June 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kumar:2016:PAI

[3098] K. P. Krishna Kumar, Agrima Srivastava, and G. Geethakumari. A psychometric analysis of information propagation in online social networks using latent trait theory. Computing, 98(6):583–607, June 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lopez-Lorca:2016:ARD

[3099] Antonio Lopez-Lorca, Ghassan Beydoun, Rafael Valencia-Garcia, and Rodrigo Martínez-Béjar. Automating the reuse of domain knowledge to improve the modelling outcome from interactions between developers and clients. Computing, 98(6):609–640, June 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Salimian:2016:AFT

[3100] Leili Salimian, Faramarz Safi Esfahani, and Mohammad-Hossein Nadimi-Shahraki. An adaptive fuzzy thresholdbased approach for energy and performance efficient consolidation of virtual machines. Computing, 98(6):641–660, June 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Baskiyar:2016:MEC

[3101] Sanjeev Baskiyar, Chiao-Ching Huang, and Tin-Yau Tam. Minimum energy consumption for rate monotonic scheduled tasks. Computing, 98(6):661–684, June 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wang:2016:LLE

[3102] Mingzhong Wang, Dan Liu, Liehuang Zhu, Yongjun Xu, and Fei Wang. LESPP: lightweight and efficient strong privacy preserving authentication scheme for secure VANET communication. Computing, 98(7):685–708, July 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ferreira:2016:MCD

[3103] Alexandre Mello Ferreira and Barbara Pernici. Managing the complex data center environment: an integrated energy-aware framework. Computing, 98(7):709–749, July 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00607-014-0405-x.pdf.

Hameed:2016:STE

[3104] Abdul Hameed, Alireza Khoshkbarforoushha, Rajiv Ranjan, Prem Prakash Jayaraman, Joanna Kolodziej, Pavan Balaji, Sherali Zeadally, Qutaibah Marwan Malluhi, Nikos Tziritas, Abhinav Vishnu, Samee U. Khan, and Albert Zomaya. A survey and taxonomy on energy efficient resource allocation techniques for cloud computing systems. Computing, 98(7):751–774, July 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Noubir:2016:SIN

[3105] Guevara Noubir and Michel Raynal. Special issue on NETYS’14 selected papers. Computing, 98(8):775, August 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00607-
016-0506-9.pdf.

Santos:2016:FLM

[3106] Agustín Santos, Antonio Fernández Anta, José A. Cuesta, and Luis López Fernández. Fair linking mechanisms for resource allocation with correlated player types. Computing, 98(8):777– 801, August 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Boushaba:2016:IMO

[3107] Abdelali Boushaba, Adil Benabbou, Rachid Benabbou, Azeddine Zahi, and Mohammed Oumsis. An intelligent multipath optimized link state routing protocol for QoS and QoE enhancement of video transmission in MANETs. Computing, 98(8):803–825, August 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Boutet:2016:PPD

[3108] Antoine Boutet, Davide Frey, Rachid Guerraoui, Arnaud Jégou, and Anne-Marie Kermarrec. Privacy-preserving distributed collaborative filtering. Computing, 98(8):827–846, August 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lins:2016:ASB

[3109] Fernando Lins, Julio Damasceno, Robson Medeiros, Erica Sousa, and Nelson Rosa. Automation of servicebased security-aware business processes in the cloud. Computing, 98(9): 847–870, September 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Entrialgo:2016:APM

[3110] Joaquín Entrialgo, Ramón Medrano, Daniel Fernando García, and Javier García. Autonomic power management with self-healing in server clusters under QoS constraints. Computing, 98 (9):871–894, September 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Low:2016:RHC

[3111] W. Z. Low, S. K. L. M. vanden Broucke, M. T. Wynn, A. H. M. ter Hofstede, J. De Weerdt, and W. M. P. van der Aalst. Revising history for costinformed process improvement. Computing, 98(9):895–921, September 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

deJuan-Marin:2016:SAC

[3112] Rubén de Juan-Marín, Hendrik Decker, José Enrique Armendáriz-Íñigo, José M. Bernabéu-Aubán, and Francesc D. Muñoz-Escoí. Scalability approaches for causal multicast: a survey. Computing, 98(9):923–947, September 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Goyal:2016:GSL

[3113] Sudhir Goyal, Seema Bawa, and Bhupinder Singh. Green Service Level Agreement (GSLA) framework for cloud computing. Computing, 98 (9):949–963, September 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dustdar:2016:EAC

[3114] Schahram Dustdar. Editorial: 50th anniversary of Computing. Computing, 98 (10):965–966, October 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00607-016-0515-8.pdf.

vanSteen:2016:BID

[3115] Maarten van Steen and Andrew S.

Tanenbaum. A brief introduction to distributed systems. Computing, 98 (10):967–1009, October 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00607-016-0508-7.pdf.

Tsigkanos:2016:ADC

[3116] Christos Tsigkanos, Timo Kehrer, and Carlo Ghezzi. Architecting dynamic cyber-physical spaces. Computing, 98 (10):1011–1040, October 2016. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Georgakopoulos:2016:ITI

[3117] Dimitrios Georgakopoulos and Prem Prakash Jayaraman. Internet of things: from Internet scale sensing to smart services. Computing, 98(10):1041–1058, October 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Helfert:2016:FSI

[3118] Markus Helfert, Donald Ferguson, and Víctor Méndez Muñoz. Foreword special issue on cloud computing and services science. Computing, 98(11): 1059–1060, November 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00607-016-0519-4.pdf.

Ibrahim:2016:NAC

[3119] Rabha W. Ibrahim and Abdullah Gani. A new algorithm in cloud computing of multi-agent fractional differential economical system. Computing, 98(11): 1061–1074, November 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mitropoulou:2016:PCI

[3120] Persefoni Mitropoulou, Evangelia Filiopoulou, Christos Michalakelis, and Mara Nikolaidou. Pricing cloud IaaS services based on a hedonic price index. Computing, 98(11):1075–1089, November 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Nirmala:2016:CPB

[3121] S. Jaya Nirmala and S. Mary Saira Bhanu. Catfish-PSO based scheduling of scientific workflows in IaaS cloud. Computing, 98(11):1091–1109, November 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kholidy:2016:RMA

[3122] Hisham A. Kholidy, Abdelkarim Erradi, Sherif Abdelwahed, and Fabrizio Baiardi. A risk mitigation approach for autonomous cloud intrusion response system. Computing, 98(11): 1111–1135, November 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Shirvani:2016:SSD

[3123] Hussein Shirvani and Hamed Vahdat-Nejad. Storing shared documents that are customized by users in cloud computing. Computing, 98(11):1137–1151, November 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Park:2016:VCB

[3124] Joonseok Park, Youngmin An, Taejun Kang, and Keunhyuk Yeom. Virtual cloud bank: consumer-centric service recommendation process and architectural perspective for cloud service brokers. Computing, 98(11):1153–1184, November 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Radhakrishnan:2016:ECC

[3125] A. Radhakrishnan and V. Kavitha. Energy conservation in cloud data centers by minimizing virtual machines migration through artificial neural network. Computing, 98(11):1185–1202, November 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Raju:2016:ECC

[3126] B. K. S. P. Kumar Raju and G. Geethakumari. Event correlation in cloud: a forensic perspective. Computing, 98(11):1203–1224, November 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Balliu:2016:BDA

[3127] Alkida Balliu, Dennis Olivetti, Ozalp Babaoglu, Moreno Marzolla, and Alina Sîrbu. A big data analyzer for large trace logs. Computing, 98(12): 1225–1249, December 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Orozco:2016:TTF

[3128] Ana Lucila Sandoval Orozco, Jocelin Rosales Corripio, David Manuel Arenas González, Luis Javier García Villalba, and Julio Hernandez-Castro. Theia: a tool for the forensic analysis of mobile devices pictures. Computing, 98(12):1251–1286, December 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Izadi:2016:ADC

[3129] Davood Izadi, Sara Ghanavati, Jemal Abawajy, and Tutut Herawan. An alternative data collection scheduling scheme in wireless sensor networks. Computing, 98(12):1287–1304, December 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Harshan:2016:SEE

[3130] J. Harshan, Frédérique Oggier, and Anwitaman Datta. Sparsity exploiting erasure coding for distributed storage of versioned data. Computing, 98(12): 1305–1329, December 2016. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Helfert:2017:S

[3131] Markus Helfert, Karl-Heinz Krempels, Cornel Klein, and Brian Donnellan. SMARTGREENS 2015. Computing, 99(1):1–2, January 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00607-016-0525-6.pdf.

Beldiceanu:2017:TEP

[3132] Nicolas Beldiceanu, Bárbara Dumas Feris, Philippe Gravey, Sabbir Hasan, Claude Jard, Thomas Ledoux, Yunbo Li, Didier Lime, Gilles Madi-Wamba, Jean-Marc Menaud, Pascal Morel, Michel Morvan, Marie-Laure Moulinard, Anne-Cécile Orgerie, Jean-Louis Pazat, and et al. Towards energyproportional clouds partially powered by renewable energy. Computing, 99 (1):3–22, January 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

AlSkaif:2017:DPS

[3133] Tarek AlSkaif, Manel Guerrero Zapata, Boris Bellalta, and Anders Nilsson. A distributed power sharing framework among households in microgrids: a repeated game approach. Computing, 99(1):23–37, January 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Cagliero:2017:PCC

[3134] Luca Cagliero, Tania Cerquitelli, Silvia Chiusano, Paolo Garza, and Xin Xiao. Predicting critical conditions in bicycle sharing systems. Computing, 99(1):39–57, January 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Suresh:2017:GDC

[3135] Velamuri Suresh and S. Sreejith. Generation dispatch of combined solar thermal systems using dragonfly algorithm. Computing, 99(1):59–80, January 2017.

CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pugalendhi:2017:FBT

[3136] GaneshKumar Pugalendhi, Durgadevi Velusamy, Anand Paul, and Ku-Jin Kim. Fuzzy-based trusted routing to mitigate packet dropping attack between data aggregation points in smart grid communication network. Computing, 99(1):81–106, January 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ramirez:2017:THS

[3137] Alejandro R. García Ramirez, Israel González-Carrasco, Gustavo Henrique Jasper, Amarilys Lima Lopez, Jose Luis Lopez-Cuadrado, and Angel García-Crespo. Towards human smart cities: Internet of Things for sensory impaired individuals. Computing, 99 (1):107–126, January 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Reisig:2017:ASS

[3138] Wolfgang Reisig. The 9th advanced summer school on service-oriented computing, June 28–July 3, Hersonissos, Crete, Greece. Computing, 99 (2):127, February 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s00607-016-0527-4.pdf.

Zimmermann:2017:ARC

[3139] Olaf Zimmermann. Architectural refactoring for the cloud: a decision-centric view on cloud migration. Computing, 99(2):129–145, February 2017. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Vukojevic-Haupt:2017:DPW

[3140] Karolina Vukojevic-Haupt, Florian Haupt, and Frank Leymann. Ondemand provisioning of workflow middleware and services into the cloud: an overview. Computing, 99(2):147–162, February 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hirmer:2017:SRH

[3141] Pascal Hirmer, Matthias Wieland, Holger Schwarz, Bernhard Mitschang, Uwe Breitenbücher, Santiago Gómez Sáez, and Frank Leymann. Situation recognition and handling based on executing situation templates and situationaware workflows. Computing, 99(2): 163–181, February 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Steimle:2017:EPS

[3142] Frank Steimle, Matthias Wieland, Bernhard Mitschang, Sebastian Wagner, and Frank Leymann. Extended provisioning, security and analysis techniques for the ECHO health data management system. Computing, 99 (2):183–201, February 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kim:2017:IRO

[3143] Taeheung Kim, Byung Do Chung, and Jong-Seok Lee. Incorporating receiver operating characteristics into naive Bayes for unbalanced data classification. Computing, 99(3):203–218, March 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

dePaula:2017:PRV

[3144] Lauro Cássio Martins de Paula, Anderson S. Soares, Telma W. L. Soares, Arlindo R. G. Filho, Clarimar J. Coelho, Alexandre C. B. Delbem, and Wellington S. Martins. Parallel regressions for variable selection using GPU. Computing, 99(3):219–234, March 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Deng:2017:EDB

[3145] Yuhui Deng, Jieming Cai, Wen Jiang, and Xiao Qin. Employing dual-block correlations to reduce the energy consumption of disk drives. Computing, 99(3):235–253, March 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rathore:2017:DTL

[3146] Santosh S. Rathore and Sandeep Kumar. A decision tree logic based recommendation system to select software fault prediction techniques. Computing, 99(3):255–285, March 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Silva:2017:MCF

[3147] Francisco Airton Silva, Paulo Maciel, Eder Santana, Rubens Matos, and Jamilson Dantas. Mobile cloud face recognition based on smart cloud ranking. Computing, 99(3):287–311, March 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Beheshti:2017:SRC

[3148] Seyed-Mehdi-Reza Beheshti, Boualem Benatallah, Srikumar Venugopal, Seung Hwan Ryu, Hamid Reza Motahari-Nezhad, and Wei Wang. A systematic review and comparative analysis of cross-document coreference resolution methods and tools. Computing, 99(4):313–349, April 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sahni:2017:HAA

[3149] Jyoti Sahni and Deo Prakash Vidyarthi. Heterogeneity-aware adaptive auto-scaling heuristic for improved QoS and resource usage in cloud environments. Computing, 99(4):351–381, April 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Huang:2017:EPI

[3150] Yuan-Ko Huang. Efficient processing of intelligent probabilistic collision detection queries. Computing, 99(4):383–404, April 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Park:2017:ASH

[3151] YongWon Park and Sanjeev Baskiyar.

Adaptive scheduling on heterogeneous systems using support vector machine. Computing, 99(4):405–425, April 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chen:2017:ECI

[3152] Chin-Yun Chen. Erratum to: Computing interval enclosures for definite integrals by application of triple adaptive strategies. Computing, 99(4):427, April 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00607-
017-0550-0.pdf. See [2578].

Monfort:2017:SIW

[3153] Valérie Monfort and Tim A. Majchrzak. Special issue on web information systems and technologies. Computing, 99(5):429–430, May 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL http://link.springer.com/
content/pdf/10.1007/s00607-017-
0544-y.pdf.

Perez-Roses:2017:IED

[3154] Hebert Pérez-Rosés and Francesc Sebé. Iterated endorsement deduction and ranking. Computing, 99(5):431–446, May 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tiwary:2017:CDD

[3155] Mayank Tiwary, Kshira Sagar Sahoo, Bibhudatta Sahoo, and Rachita Misra. CPS: a dynamic and distributed pricing policy in cyber foraging systems for fixed state cloudlets. Computing, 99(5):447–463, May 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Li:2017:ESD

[3156] Fagen Li, Bo Liu, and Jiaojiao Hong. An efficient signcryption for data access control in cloud computing. Computing, 99(5):465–479, May 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Yan:2017:RAQ

[3157] Li Yan, Ruizhe Ma, Dazhen Li, and Jingwei Cheng. RDF approximate queries based on semantic similarity. Computing, 99(5):481–491, May 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zhao:2017:MMS

[3158] Zhen Zhao and Zongmin Ma. A methodology for measuring structure similarity of fuzzy XML documents. Computing, 99(5):493–506, May 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dahan:2017:EAB

[3159] Fadl Dahan, Khalil El Hindi, and Ahmed Ghoneim. Enhanced Artificial Bee Colony Algorithm for QoSaware Web service selection problem. Computing, 99(5):507–517, May 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Alarifi:2017:MES

[3160] Abdulrahman Alarifi, Mansour Alsaleh, and Noura Alomar. A model for evaluating the security and usability of e-banking platforms. Computing, 99(5):519–535, May 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dietel:2017:TDH

[3161] Jana-Eva Dietel. They do! How the explicit motive-structure predicts user behavior in (business-) social network sites like Xing or LinkedIn. Computing, 99(5):537–550, May 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ardagna:2017:SIS

[3162] Danilo Ardagna and Cinzia Cappiello. Special issue on sustainable computing systems and applications. Computing, 99(6):551–552, June 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00607-017-0556-7.pdf.

Jagroep:2017:ESA

[3163] Erik Jagroep, Jan Martijn van der Werf, Sjaak Brinkkemper, Leen Blom, and Rob van Vliet. Extending software architecture views with an energy consumption perspective. Computing, 99(6):553–573, June 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00607-016-0502-0.pdf.

Canali:2017:SAV

[3164] Claudia Canali and Riccardo Lancellotti. Scalable and automatic virtual machines placement based on behavioral similarities. Computing, 99 (6):575–595, June 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Guitart:2017:TSD

[3165] Jordi Guitart. Toward sustainable data centers: a comprehensive energy management strategy. Computing, 99(6):597–615, June 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Munoz-Escoi:2017:SEM

[3166] Francesc D. Muñoz-Escoí and José M.

Bernabéu-Aubán. A survey on elasticity management in PaaS systems. Computing, 99(7):617–656, July 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kao:2017:TEF

[3167] Chia Hung Kao. Testing and evaluation framework for virtualization technologies. Computing, 99(7):657–677, July 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zalik:2017:CDN

[3168] Krista Rizman Zalik. Community detection in networks using new update rules for label propagation. Computing, 99(7):679–700, July 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mayer:2017:TCL

[3169] Philip Mayer. A taxonomy of crosslanguage linking mechanisms in open source frameworks. Computing, 99 (7):701–724, July 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Benedict:2017:SIE

[3170] Shajulin Benedict, Michael Gerndt, and Siegfried Benkner. Special issue on energy reduction techniques for exa-scale computing: theory and practice. Computing, 99(8):725–726, August 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00607-
017-0570-9.pdf.

Schuchart:2017:RFA

[3171] Joseph Schuchart, Michael Gerndt, Per Gunnar Kjeldsberg, Michael Lysaght, David Horák, Lubomír Ríha, Andreas Gocht, Mohammed Sourouri, Madhura Kumaraswamy, Anamika Chowdhury, Magnus Jahre, Kai Diethelm, Othman Bouizi, Umbreen Sabir Mian, Jakub Kruzík, and et al. The READEX formalism for automatic tuning for energy efficiency. Computing, 99(8):727–745, August 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00607-
016-0532-7.pdf.

Guillen:2017:EMD

[3172] Carla Guillen, Carmen Navarrete, David Brayford, Wolfram Hesse, and Matthias Brehm. Energy model derivation for the DVFS automatic tuning plugin: tuning energy and power related tuning objectives. Computing, 99(8):747–764, August 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rejitha:2017:EPC

[3173] R. S. Rejitha, Shajulin Benedict, Suja A. Alex, and Shany Infanto. Energy prediction of CUDA application instances using dynamic regression models. Computing, 99(8):765–790, August 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chen:2017:FHS

[3174] Cheng Chen, Jianbin Fang, Tao Tang, and Canqun Yang. LU factorization on heterogeneous systems: an energyefficient approach towards high performance. Computing, 99(8):791–811, August 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Papadopoulos:2017:WCA

[3175] Andreas Papadopoulos, George Pallis, and Marios D. Dikaiakos. Weighted clustering of attributed multi-graphs. Computing, 99(9):813–840, September 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sousa:2017:CIP

[3176] Erica Sousa, Fernando Lins, Eduardo Tavares, and Paulo Maciel. Cloud infrastructure planning considering different redundancy mechanisms. Computing, 99(9):841–864, September 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kelefouras:2017:MPS

[3177] Vasilios Kelefouras. A methodology pruning the search space of six compiler transformations by addressing them together as one problem and by exploiting the hardware architecture details. Computing, 99(9):865–888, September 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gupta:2017:SMD

[3178] Shivani Gupta and Atul Gupta. A set of measures designed to identify overlapped instances in software defect prediction. Computing, 99(9): 889–914, September 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Li:2017:PBA

[3179] Jingwei Li, Adam Bowers, Dan Lin, Peng Jiang, and Wei Jiang. PAS: policy-based assistance in sensor networks. Computing, 99(10):915–928, October 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Andrade:2017:AMA

[3180] Ermeson Andrade, Bruno Nogueira, Rubens Matos, Gustavo Callou, and Paulo Maciel. Availability modeling and analysis of a disaster-recovery-as-a-service solution. Computing, 99 (10):929–954, October 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kumar:2017:GBE

[3181] Neetesh Kumar and Deo Prakash Vidyarthi. A GA based energy aware scheduler for DVFS enabled multicore systems. Computing, 99(10):955–977, October 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Poshtkohi:2017:PIF

[3182] Alireza Poshtkohi, M. B. Ghaznavi-Ghoushchi, and Kamyar Saghafi. The Parvicursor infrastructure to facilitate the design of grid and cloud computing systems. Computing, 99(10): 979–1006, October 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Agrawal:2017:SCM

[3183] Tarun K. Agrawal, Aryabartta Sahu, Manojit Ghose, and R. Sharma. Scheduling chained multiprocessor tasks onto large multiprocessor system. Computing, 99(10):1007–1028, October 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Haytaoglu:2017:NHC

[3184] Elif Haytaoglu and Mehmet Emin Dalkilic. A new hybrid coding scheme: homomorphic minimum bandwidth repairing codes. Computing, 99(11): 1029–1054, November 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Savic:2017:AHS

[3185] Milos Savić, Mirjana Ivanović, and Milos Radovanović. Analysis of high structural class coupling in objectoriented software systems. Computing, 99(11):1055–1079, November 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tran:2017:OOS

[3186] Dai H. Tran, Nguyen H. Tran, Chuan Pham, S. M. Ahsan Kazmi, Eui-Nam Huh, and Choong Seon Hong. OaaS: offload as a service in fog networks. Computing, 99(11):1081–1104, November 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hanif:2017:AVA

[3187] Muhammad Kashif Hanif and Karl-Heinz Zimmermann. Accelerating Viterbi algorithm on graphics processing units. Computing, 99(11): 1105–1123, November 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Atef:2017:LBC

[3188] Asmaa Atef, Tarek Hagras, Yousef B. Mahdy, and Jan Janecek. Lower-bound complexity algorithm for task scheduling on heterogeneous grid. Computing, 99(11):1125–1145, November 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Meghanathan:2017:RAD

[3189] Natarajan Meghanathan. Retracted Article: A distance vector similarity metric for complex networks. Computing, 99(11):1147, November 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00607-
017-0565-6.pdf.

Taheri:2017:VBB

[3190] Javid Taheri, Albert Y. Zomaya, and Andreas Kassler. vmBBProfiler: a black-box profiling approach to quantify sensitivity of virtual machines to shared cloud resources. Computing, 99(12):1149–1177, December 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00607-
017-0552-y.pdf.

Sanjeevi:2017:NSA

[3191] P. Sanjeevi and P. Viswanathan. NUTS scheduling approach for cloud data centers to optimize energy consumption. Computing, 99(12):1179–1205, December 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hameed:2017:PIH

[3192] Muhammad Umar Hameed, Syed Ali Haider, and Burak Kantarci. Performance impacts of hybrid cloud storage. Computing, 99(12):1207–1229, December 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Liu:2017:SOA

[3193] Xi Liu, Xiaolu Zhang, Weidong Li, and Xuejie Zhang. Swarm optimization algorithms applied to multi-resource fair allocation in heterogeneous cloud computing systems. Computing, 99(12): 1231–1255, December 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lee:2017:RTU

[3194] Jinkyu Lee. Real-time uniprocessor scheduling with fewer preemptions. Computing, 99(12):1257–1270, December 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ranjan:2018:NRM

[3195] Rajiv Ranjan, Prem Prakash Jayaraman, Massimo Villari, and Dimitrios Georgakopoulos. A note on resource management techniques and systems for big data workflow processing. Computing, 100(1):1–2, January 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00607-
018-0586-9.pdf.

Han:2018:POT

[3196] Wei Han, Ze Deng, Junde Chu, Jing Zhu, Peng Gao, and Tejal Shah. A parallel online trajectory compression approach for supporting big data workflow. Computing, 100(1):3–20, January 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Convolbo:2018:GTO

[3197] Moïse W. Convolbo, Jerry Chou, Ching-Hsien Hsu, and Yeh Ching Chung. GEODIS: towards the optimization of data locality-aware job scheduling in geo-distributed data centers. Computing, 100(1):21–46, January 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zhang:2018:BEG

[3198] Junqiang Zhang, Chonglong Wu, Yuzhu Wang, Yan Ma, Yong Wu, and Xiaoping Mao. The BIM-enabled geotechnical information management of a construction project. Computing, 100(1):47–63, January 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Aziza:2018:BOD

[3199] Hatem Aziza and Saoussen Krichen. Bi-objective decision support system for task-scheduling based on genetic algorithm in cloud computing. Computing, 100(2):65–91, February 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kashi:2018:LBL

[3200] Saeed Sedighian Kashi. A load balanced location service for location information management of multi-sink Wireless Sensor Networks. Computing, 100(2):93–117, February 2018. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sedaghatbaf:2018:MDE

[3201] Ali Sedaghatbaf and Mohammad Abdollahi Azgomi. A method for dependability evaluation of software architectures. Computing, 100(2):119–150, February 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ahmad:2018:TSI

[3202] Asma’a Ahmad, Robin Doss, Majeed Alajeely, and Sarab F. Al Rubeaai. Trust strategy implementation in Opp-Nets. Computing, 100(2):151–181, February 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Machado:2018:SMA

[3203] Guilherme M. Machado, Vinicius Maran, Lorayne P. Dornelles, Isabela Gasparini, Lucinéia H. Thom, and José Palazzo M. de Oliveira. A systematic mapping on adaptive recommender approaches for ubiquitous environments. Computing, 100(2):183–209, February 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gomez-Folgar:2018:MPA

[3204] F. Gomez-Folgar, G. Indalecio, N. Seoane, T. F. Pena, and A. J. Garcia-Loureiro. MPI-Performance-Aware-Reallocation: method to optimize the mapping of processes applied to a cloud infrastructure. Computing, 100(2):211–226, February 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Murtza:2018:OET

[3205] Shahid Ali Murtza, Ayaz Ahmad, Muhammad Yasir Qadri, Nadia N. Qadri, and Jameel Ahmed. Optimizing energy and throughput for MP-SoCs: an integer particle swarm optimization approach. Computing, 100 (3):227–244, March 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Yessad:2018:QBR

[3206] Nawel Yessad, Mawloud Omar, Abdelkamel Tari, and Abdelmadjid Bouabdallah. QoS-based routing in Wireless Body Area Networks: a survey and taxonomy. Computing, 100 (3):245–275, March 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bielecki:2018:GPS

[3207] Wlodzimierz Bielecki, Marek Palkowski, and Piotr Skotnicki. Generation of parallel synchronization-free tiled code. Computing, 100(3):277–302, March 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ma:2018:ATK

[3208] Zongmin Ma, Ting Li, and Li Yan. An approach of top-k keyword querying for fuzzy XML. Computing, 100 (3):303–330, March 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sheng:2018:GES

[3209] Quan Z. Sheng and Yongrui Qin. Guest editorial: special issue on advanced data mining and applications. Computing, 100(4):331–333, April 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00607-
018-0607-8.pdf.

Duan:2018:MDC

[3210] Lei Duan, Lu Liu, Guozhu Dong, Jyrki Nummenmaa, Tingting Wang, Pan Qin, and Hao Yang. Mining distinguishing customer focus sets from online customer reviews. Computing, 100(4):335–351, April 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Xu:2018:NPS

[3211] Xiaofei Xu, Gianmario Motta, Zhiying Tu, Hanchuan Xu, Zhongjie Wang, and Xianzhi Wang. A new paradigm of software service engineering in big data and big service era. Computing, 100(4):353–368, April 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Yao:2018:WIE

[3212] Lina Yao, Quan Z. Sheng, Boualem Benatallah, Schahram Dustdar, Xianzhi Wang, Ali Shemshadi, and Salil S. Kanhere. WITS: an IoT-endowed computational framework for activity recognition in personalized smart homes. Computing, 100(4):369–385, April 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Al-Maskari:2018:BIL

[3213] Sanad Al-Maskari, Zhuoming Xu, Wenping Guo, Xiaoming Zhao, and Xue Li. Bio-inspired learning approach for electronic nose. Computing, 100 (4):387–402, April 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wang:2018:CDP

[3214] Shuliang Wang, Shaopeng Wang, Hanning Yuan, Qi Li, Jing Geng, and Yang Yu. Clustering by differencing potential of data field. Computing, 100(4):403–419, April 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Liu:2018:DPC

[3215] Qian Liu, Shameek Ghosh, Jinyan Li, Limsoon Wong, and Kotagiri Ramamohanarao. Discovering pan-correlation patterns from time course data sets by efficient mining algorithms. Computing, 100(4):421–437, April 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sarikhani:2018:MPC

[3216] Mehdi Sarikhani and Andrew Wendelborn. Mechanisms for provenance collection in scientific workflow systems. Computing, 100(5):439–472, May 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ramezani:2018:LAM

[3217] Mahlagha Ramezani and Mohsen Jahanshahi. Load-aware multicast routing in multi-radio wireless mesh networks using FCA–CMAC neural network. Computing, 100(5):473–501, May 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kirsal:2018:AMO

[3218] Yonal Kirsal. Analytical modelling and optimization analysis of large-scale communication systems and networks with repairmen policy. Computing, 100(5):503–527, May 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

vanZelst:2018:DWN

[3219] S. J. van Zelst, B. F. van Dongen, W. M. P. van der Aalst, and H. M. W. Verbeek. Discovering workflow nets using integer linear programming. Computing, 100(5):529–556, May 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00607-
017-0582-5.pdf.

Bhatti:2018:LAT

[3220] Muhammad Khurram Bhatti, Isil Oz, Sarah Amin, Maria Mushtaq, Umer Farooq, Konstantin Popov, and Mats Brorsson. Locality-aware task scheduling for homogeneous parallel computing systems. Computing, 100(6):557– 595, June 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Atmani:2018:TBE

[3221] Mouloud Atmani, Djamil Aïssani, and Yassine Hadjadj-Aoul. Towards bandwidth and energy optimization in IEEE 802.15.4 wireless sensor networks. Computing, 100(6):597–620, June 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Torres:2018:HAA

[3222] Elton Torres, Gustavo Callou, and Ermeson Andrade. A hierarchical approach for availability and performance analysis of private cloud storage services. Computing, 100(6):621–644, June 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dabaghi-Zarandi:2018:DTA

[3223] Fahimeh Dabaghi-Zarandi and Zeinab Movahedi. A dynamic traffic-aware energy-efficient algorithm based on sleep-scheduling for autonomous systems. Computing, 100(6):645–665, June 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ponce:2018:AAC

[3224] Hiram Ponce, Miguel González-Mendoza, and Lourdes Martínez-Villaseñor. Advances in applied computational intelligence: MICAI 2016. Computing, 100(7):667–669, July 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s00607-
                                        018-0630-9.pdf.

Lopez-Ortega:2018:IMA

[3225] Omar López-Ortega, Félix Castro-Espinoza, and Obed Pérez-Cortés. An intelligent multi-agent system to create and classify fractal music. Computing, 100(7):671–688, July 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mohammed:2018:FBA

[3226] Assem H. Mohammed, Ahmed M. Gadallah, Hesham A. Hefny, and M. Hazman. Fuzzy based approach for discovering crops plantation knowledge from huge agro-climatic data respecting climate changes. Computing, 100(7):689–713, July 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Herrera-Alcantara:2018:IFP

[3227] Oscar Herrera-Alcántara and Miguel González-Mendoza. Inverse formulas of parameterized orthogonal wavelets. Computing, 100(7):715–739, July 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See correction [3271].

Gomez-Adorno:2018:DEL

[3228] Helena Gómez-Adorno, Juan-Pablo Posadas-Durán, Grigori Sidorov, and David Pinto. Document embeddings learned on various types of n-grams for cross-topic authorship attribution. Computing, 100(7):741–756, July 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Manogaran:2018:SIM

[3229] Gunasekaran Manogaran, Naveen Chilamkurti, and Ching-Hsien Hsu. Special issue on machine learning algorithms for Internet of Things, fog computing and cloud computing. Computing, 100 (8):757–758, August 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00607-018-0644-3.pdf.

Bhuvaneswari:2018:NML

[3230] G. Bhuvaneswari and G. Manikandan.

A novel machine learning framework for diagnosing the type 2 diabetics using temporal fuzzy ant miner decision tree classifier with temporal weighted genetic algorithm. Computing, 100 (8):759–772, August 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Yu:2018:PCB

[3231] Xiaodong Yu and Hongbin Dong. PTL-CFS based deep convolutional neural network model for remote sensing classification. Computing, 100(8):773–785, August 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wang:2018:TCS

[3232] Ke chao Wang, Tian tian Wang, and Xiao hong Su. Test case selection using multi-criteria optimization for effective fault localization. Computing, 100(8):787–808, August 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Yang:2018:NRS

[3233] Qin Yang. A novel recommendation system based on semantics and context awareness. Computing, 100 (8):809–823, August 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chen:2018:PAP

[3234] Yegang Chen. Prediction algorithm of PM2.5 mass concentration based on adaptive BP neural network. Computing, 100(8):825–838, August 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Vijaya:2018:CEF

[3235] J. Vijaya and E. Sivasankar. Computing efficient features using rough set theory combined with ensemble classification techniques to improve the customer churn prediction in telecommunication sector. Computing, 100 (8):839–860, August 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wu:2018:CPS

[3236] Zhuang Wu, Shuo Zhang, and Ting Wang. A cooperative particle swarm optimization with constriction factor based on simulated annealing. Computing, 100(8):861–880, August 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00607-
018-0625-6.pdf.

Zhou:2018:DIS

[3237] Huini Zhou, Xiaoguang Gu, and Li Li.

The dynamic investment strategy of online advertising based on spillover effect in duopoly competition market. Computing, 100(8):881–905, August 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fujita:2018:ECE

[3238] Kazuhisa Fujita, Shun Okuno, and Yoshiki Kashimori. Evaluation of the computational efficacy in GPUaccelerated simulations of spiking neurons. Computing, 100(9):907–926, September 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Feyzi:2018:PSB

[3239] Farid Feyzi and Saeed Parsa. A program slicing-based method for effective detection of coincidentally correct test cases. Computing, 100(9):

927–969, September 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Arasteh:2018:PGI

[3240] Bahman Arasteh and Jalal Najafi. Programming guidelines for improving software resiliency against softerrors without performance overhead. Computing, 100(9):971–1003, September 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Polato:2018:TAS

[3241] Mirko Polato, Alessandro Sperduti, Andrea Burattin, and Massimiliano de Leoni. Time and activity sequence prediction of business process instances. Computing, 100(9): 1005–1031, September 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kulshrestha:2018:EBD

[3242] Jagrati Kulshrestha and Manas Kumar Mishra. Energy balanced data gathering approaches in wireless sensor networks using mixed-hop communication. Computing, 100(10):1033–1058, October 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Konovalov:2018:IRD

[3243] Mikhail Konovalov and Rostislav Razumchik. Improving routing decisions in parallel non-observable queues. Computing, 100(10):1059–1079, October 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Boukadi:2018:FNS

[3244] Khouloud Boukadi, Mouna Rekik, Molka Rekik, and Hanêne Ben-Abdallah. FC4CD: a new SOA-based Focused Crawler for Cloud service Discovery. Computing, 100(10):1081–1107, October 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Singh:2018:CDP

[3245] Prem Kumar Singh. Cloud data processing using granular based weighted concept lattice and Hamming distance. Computing, 100(10):1109–1132, October 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pereira:2018:HEI

[3246] Lucas Pereira and Antonio G. Ravelo-García. Highlights of ES2DE and IWOBI 2017: extended versions of selected best papers. Computing, 100(11):1133–1135, November 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL http://link.springer.com/
content/pdf/10.1007/s00607-018-
0660-3.pdf.

Paiva:2018:OEP

[3247] Luís Tiago Paiva and Fernando A.

C. C. Fontes. Optimal electric power generation with underwater kite systems. Computing, 100(11):1137–1153, November 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Munoz:2018:FCF

[3248] Daniel-Jesus Munoz, Mónica Pinto, and Lidia Fuentes. Finding correlations of features affecting energy consumption and performance of web servers using the HADAS eco-assistant. Computing, 100(11):1155–1173, November 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rodriguez-Esparragon:2018:ASQ

[3249] Dionisio Rodríguez-Esparragón, Javier Marcello, Consuelo Gonzalo-Martín, Ángel García-Pedrero, and Francisco Eugenio. Assessment of the spectral quality of fused images using the CIEDE2000 distance. Computing, 100 (11):1175–1188, November 2018. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sengar:2018:CVB

[3250] Namita Sengar, Malay Kishore Dutta, and Carlos M. Travieso. Computer vision based technique for identification and quantification of powdery mildew disease in cherry leaves. Computing, 100(11):1189–1201, November 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mendonca:2018:PWD

[3251] Fábio Mendonça, Sheikh Shanawaz Mostafa, Fernando Morgado-Dias, Juan L. Navarro-Mesa, Gabriel Juliá-Serdá, and Antonio G. Ravelo-García. A portable wireless device based on oximetry for sleep apnea detection. Computing, 100(11):1203–1219, November 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Verma:2018:DSD

[3252] Rohit Verma and Abhishek Srivastava. A dynamic service description for mobile environments. Computing, 100 (12):1221–1249, December 2018. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Banihashemian:2018:PMC

[3253] Seyed Saber Banihashemian, Fazlollah Adibnia, and Mehdi A. Sarram. Proposing a measurement criterion to evaluate the border problem in localization algorithms in WSNs. Computing, 100(12):1251–1272, December 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pikle:2018:HPI

[3254] Nileshchandra K. Pikle, Shailesh R.

Sathe, and Arvind Y. Vyavahare. High performance iterative elemental product strategy in assembly-free FEM on GPU with improved occupancy. Computing, 100(12):1273–1297, December 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hasham:2018:RSW

[3255] Khawar Hasham and Kamran Munir.

Reproducibility of scientific workflows execution using cloud-aware provenance (ReCAP). Computing, 100(12): 1299–1333, December 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Abdulla:2019:SIN

[3256] Parosh Aziz Abdulla and Carole Delporte. Special issue on NETYS’2016. Computing, 101(1):1–2, January 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL http://link.springer.com/
content/pdf/10.1007/s00607-018-
0672-z.pdf.

Mostefaoui:2019:TER

[3257] Achour Mostéfaoui, Michel Raynal, and Matthieu Roy. Time-efficient read/ write register in crash-prone asynchronous message-passing systems. Computing, 101(1):3–17, January 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Olivares:2019:CKA

[3258] Javier Olivares, Anne-Marie Kermarrec, and Nitin Chiluka. The out-of-core KNN awakens: the light side of computation force on large datasets. Computing, 101(1):19–38, January 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Taubenfeld:2019:WCA

[3259] Gadi Taubenfeld. Waiting in concurrent algorithms. Computing, 101 (1):39–57, January 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mutluergil:2019:MRP

[3260] Suha Orhun Mutluergil and Serdar Tasiran. A mechanized refinement proof of the Chase–Lev deque using a proof system. Computing, 101 (1):59–74, January 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

deLemos:2019:SID

[3261] Rogério de Lemos. Special issue on dependable computing: theory and practice. Computing, 101(2):75–76, February 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s00607-
                                        018-0671-0.pdf.

Tuzov:2019:SEL

[3262] Ilya Tuzov, David de Andrés, and Juan-Carlos Ruiz. Simulating the effects of logic faults in implementationlevel VITAL-compliant models. Computing, 101(2):77–96, February 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eischer:2019:SBF

[3263] Michael Eischer and Tobias Distler.

Scalable Byzantine fault-tolerant statemachine replication on heterogeneous servers. Computing, 101(2):97–118, February 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Barbosa:2019:ERS

[3264] Raul Barbosa, Frederico Cerveira, Luís Gonçalo, and Henrique Madeira. Emulating representative software vulnerabilities using field data. Computing, 101 (2):119–138, February 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Movahedi:2019:CBV

[3265] Yazdan Movahedi, Michel Cukier, Ambrose Andongabo, and Ilir Gashi. Cluster-based vulnerability assessment of operating systems and web browsers. Computing, 101(2):139–160, February 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00607-
018-0663-0.pdf.

Nunes:2019:ESC

[3266] Paulo Nunes, Ibéria Medeiros, José Fonseca, Nuno Neves, Miguel Correia, and Marco Vieira. An empirical study on combining diverse static analysis tools for web security vulnerabilities based on development scenarios. Computing, 101(2):161–185, February 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sheikh:2019:GRL

[3267] Nasrullah Sheikh, Zekarias Kefato, and Alberto Montresor. gat2vec: representation learning for attributed graphs. Computing, 101(3):187–209, March 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Yang:2019:PLB

[3268] Chao-Tung Yang, Shuo-Tsung Chen, Jung-Chun Liu, Yi-Wei Su, Deepak Puthal, and Rajiv Ranjan. A predictive load balancing technique for software defined networked cloud services. Computing, 101(3):211–235, March 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Allahbakhsh:2019:CPF

[3269] Mohammad Allahbakhsh, Saeed Arbabi, Mohammadreza Galavii, Florian Daniel, and Boualem Benatallah. Crowdsourcing planar facility location allocation problems. Computing, 101(3):237–261, March 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sellami:2019:LEN

[3270] Mohamed Sellami, Zakaria Maamar, Sylvain Lefebvre, and Noura Faci. Logs and enterprise networks for overcoming obstacles in business processes. Computing, 101(3):263–288, March 2019.

CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Herrera-Alcantara:2019:CIF

[3271] Oscar Herrera-Alcántara and Miguel González-Mendoza. Correction to: Inverse formulas of parameterized orthogonal wavelets. Computing, 101(3):289, March 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00607-
018-0656-z.pdf. See [3227].

Badidi:2019:PCS

[3272] Elarbi Badidi, Yacine Atif, Quan Z.

Sheng, and Muthucumaru Maheswaran. On personalized cloud service provisioning for mobile users using adaptive and context-aware service composition. Computing, 101 (4):291–318, April 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00607-018-0631-8.pdf.

Tsiknas:2019:PET

[3273] Konstantinos Tsiknas, Konstantinos Rantos, Christos J. Schinas, and Andrew Soilemes. Performance evaluation of TCP–BIAD in high-speed, long-distance networks. Computing, 101(4):319–337, April 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kim:2019:MNT

[3274] Jinoh Kim, Alex Sim, Brian Tierney, Sang Suh, and Ikkyun Kim. Multivariate network traffic analysis using clustered patterns. Computing, 101(4):339–361, April 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kashi:2019:ACH

[3275] Saeed Sedighian Kashi. Area coverage of heterogeneous wireless sensor networks in support of Internet of Things demands. Computing, 101(4):363–385, April 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Galindo:2019:AAF

[3276] José A. Galindo, David Benavides, Pablo Trinidad, Antonio-Manuel Gutiérrez-Fernández, and Antonio Ruiz-Cortés. Automated analysis of feature models: Quo vadis? Computing, 101(5):387–433, May 2019. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Vo:2019:CCH

[3277] Phuong L. Vo and Nguyen H. Tran. Cooperative caching for HTTP-based adaptive streaming contents in cacheenabled radio access networks. Computing, 101(5):435–453, May 2019. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Saeed:2019:LSN

[3278] Saad Saeed, Muhammad Asim, Thar Baker, and Zakaria Maamar. A location-sensitive and network-aware broker for recommending Web services. Computing, 101(5):455–475, May 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zhou:2019:APA

[3279] Lingyun Zhou, Lixin Ding, Maode Ma, and Wan Tang. An accurate partially attracted firefly algorithm. Computing, 101(5):477–493, May 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wong:2019:MOO

[3280] Kelvin Kian Loong Wong, Zhihua Liu, and Quan Zou. Multi-objective optimization and data analysis in informationization. Computing, 101 (6):495–498, June 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00607-019-00718-3.pdf.

Goli:2019:MOI

[3281] Alireza Goli, Erfan Babaee Tirkolaee, Behnam Malmir, Gui-Bin Bian, and Arun Kumar Sangaiah. A multiobjective invasive weed optimization algorithm for robust aggregate production planning under uncertain seasonal demand. Computing, 101(6):499–529, June 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ruan:2019:ELM

[3282] Qunsheng Ruan, Qingfeng Wu, Yingdong Wang, Xiling Liu, and Fengyu Miao. Effective learning model of user classification based on ensemble learning algorithms. Computing, 101(6):531–545, June 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tirkolaee:2019:MOM

[3283] Erfan Babaee Tirkolaee, Alireza Goli, Milad Hematian, Arun Kumar Sangaiah, and Tao Han. Multiobjective multi-mode resource constrained project scheduling problem using Pareto-based algorithms. Computing, 101(6):547–570, June 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Li:2019:IRL

[3284] Ming Li, Changqin Huang, Dianhui Wang, Qintai Hu, Jia Zhu, and Yong Tang. Improved randomized learning algorithms for imbalanced and noisy educational data classification. Computing, 101(6):571–585, June 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wang:2019:LPP

[3285] Xizhe Wang, Pengze Wu, Guang Liu, Qionghao Huang, Xiaoling Hu, and Haijiao Xu. Learning performance prediction via convolutional GRU and explainable neural networks in e-learning environments. Computing, 101(6):587– 604, June 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Yang:2019:AWF

[3286] Chao Yang, Shiqian Wu, Hongping Fang, and Meng Joo Er. Adaptive Weber-face for robust illumination face recognition. Computing, 101(6):605– 619, June 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lai:2019:NNP

[3287] Xiaobo Lai, Xiaomei Xu, Lili Lv, Zihe Huang, Jinyan Zhang, and Peng Huang. A novel non-parametric transform stereo matching method based on mutual relationship. Computing, 101(6):621–635, June 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zhang:2019:DLB

[3288] Xiaodong Zhang, Yiqun Zhang, and Qingmao Hu. Deep learning based vein segmentation from susceptibilityweighted images. Computing, 101 (6):637–652, June 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Xu:2019:NML

[3289] Chenchu Xu, Lin Xu, Heye Zhang, Yanping Zhang, Xiuquan Du, and Shuo Li. A novel machine-learning algorithm to estimate the position and size of myocardial infarction for MRI sequence. Computing, 101(6):653–665, June 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zhang:2019:EEJ

[3290] Weiwei Zhang, Jinglin Zhang, Xiuquan Du, Yanping Zhang, and Shuo Li. An end-to-end joint learning framework of artery-specific coronary calcium scoring in non-contrast cardiac CT. Computing, 101(6):667–678, June 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zhang:2019:ASR

[3291] Hanfei Zhang, Yu Lu, Shiwei Ma, Shuhao Cao, Qingwei Xia, Yanyan Liu, and Haiyan Zhang. Adaptive sparse reconstruction of damage localization via Lamb waves for structure health monitoring. Computing, 101(6):679–692, June 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wang:2019:IHB

[3292] Kuo Wang, Sumei Li, Qing Wang, and Chunping Hou. Identification of hormone-binding proteins using a novel ensemble classifier. Computing, 101(6):693–703, June 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Li:2019:MSM

[3293] Taoying Li, Qian Yin, Runyu Song, Mingyue Gao, and Yan Chen. Multidimensional scaling method for prediction of lysine glycation sites. Computing, 101(6):705–724, June 2019. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Xhafa:2019:SII

[3294] Fatos Xhafa. Special issue “Internet of Things (IoT)-based Services”. Computing, 101(7):725–727, July 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s00607-
                                        019-00719-2.pdf.

Zhang:2019:MAE

[3295] Ping Zhang, Mimoza Durresi, and Arjan Durresi. Multi-access edge computing aided mobility for privacy protection in Internet of Things. Computing, 101(7):729–742, July 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kayes:2019:CSM

[3296] A. S. M. Kayes, Wenny Rahayu, and Tharam Dillon. Critical situation management utilizing IoT-based data resources through dynamic contextual role modeling and activation. Computing, 101(7):743–772, July 2019. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wang:2019:IPV

[3297] Xu An Wang, Yudong Liu, Arun Kumar Sangaiah, and Jindan Zhang. Improved publicly verifiable group sum evaluation over outsourced data streams in IoT setting. Computing, 101(7):773–790, July 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Alshehri:2019:FSP

[3298] Mohammad Dahman Alshehri and Farookh Khadeer Hussain. A fuzzy security protocol for trust management in the Internet of Things (Fuzzy-IoT). Computing, 101(7):791–818, July 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Amato:2019:DCI

[3299] Flora Amato, Giovanni Cozzolino, Antonino Mazzeo, and Francesco Moscato. Detect and correlate information system events through verbose logging messages analysis. Computing, 101(7):819–830, July 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Khan:2019:ERC

[3300] Zahoor Ali Khan, Ghazanfar Latif, Arshad Sher, Imran Usman, Mahmood Ashraf, Manzoor Ilahi, and Nadeem Javaid. Efficient routing for corona based underwater wireless sensor networks. Computing, 101(7):831–856, July 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bosmans:2019:TIS

[3301] Stig Bosmans, Siegfried Mercelis, Joachim Denil, and Peter Hellinckx. Testing IoT systems using a hybrid simulation based testing approach. Computing, 101(7):857–872, July 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

deHoog:2019:TDR

[3302] Jens de Hoog, Arthur Janssens, Siegfried Mercelis, and Peter Hellinckx. Towards a distributed real-time hybrid simulator for autonomous vehicles. Computing, 101(7):873–891, July 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Memeti:2019:UMH

[3303] Suejb Memeti, Sabri Pllana, Alécio Binotto, Joanna Kołodziej, and Ivona Brandic. Using meta-heuristics and machine learning for software optimization of parallel computing systems: a systematic literature review. Computing, 101(8):893–936, August 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s00607-
                                        018-0614-9.pdf.

Lira:2019:VNM

[3304] Victor Lira, Eduardo Tavares, Meuse Oliveira, Jr., Erica Sousa, and Bruno Nogueira. Virtual network mapping considering energy consumption and availability. Computing, 101(8):937– 967, August 2019. CODEN CMPTA2.

ISSN 0010-485X (print), 1436-5057 (electronic).

Arslan:2019:PBI

[3305] Hilal Arslan and Murat Manguoglu. A parallel bio-inspried shortest path algorithm. Computing, 101(8):969–988, August 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ahmed:2019:ILT

[3306] Jawad Ahmed, Aqsa Malik, Muhammad U. Ilyas, and Jalal S. Alowibdi. Instance launch-time analysis of Open-Stack virtualization technologies with control plane network errors. Computing, 101(8):989–1014, August 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Maurya:2019:ENC

[3307] Ashish Kumar Maurya and Anil Kumar Tripathi. ECP: a novel clusteringbased technique to schedule precedence constrained tasks on multiprocessor computing systems. Computing, 101 (8):1015–1039, August 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zhao:2019:ERS

[3308] Yang Zhao, Mao Ren, Songquan Jiang, Guobin Zhu, and Hu Xiong. An efficient and revocable storage CP– ABE scheme in the cloud computing. Computing, 101(8):1041–1065, August 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Parvanak:2019:CLL

[3309] A. R. Parvanak, M. Jahanshahi, and M. Dehghan. A cross-layer learning automata based gateway selection method in multi-radio multi-channel wireless mesh networks. Computing, 101(8):1067–1090, August 2019. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

BenBoudaoud:2019:ICB

[3310] Lynda Ben Boudaoud, Basel Solaiman, and Abdelkamel Tari. Implementation and comparison of binary thinning algorithms on GPU. Computing, 101 (8):1091–1117, August 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Megahed:2019:MSI

[3311] Aly Megahed, Jennifer A. Pazour, Ahmed Nazeem, Samir Tata, and Mohamed Mohamed. Monitoring services in the Internet of Things: an optimization approach. Computing, 101 (8):1119–1145, August 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Samadi:2019:IIS

[3312] Negin Samadi and Asgarali Bouyer. Identifying influential spreaders based on edge ratio and neighborhood diversity measures in complex networks. Computing, 101(8):1147–1175, August 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Muthukumar:2019:DBC

[3313] Sumathi Muthukumar, Senthilkumar Muthukrishnan, and Veeramani Chinnadurai. Dynamic behaviour of competing memes’ spread with alert influence in multiplex social-networks. Computing, 101(8):1177–1197, August 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ren:2019:EBP

[3314] Lingmei Ren, Qingyang Zhang, Weisong Shi, and Yanjun Peng. Edgebased personal computing services: fall detection as a pilot study. Computing, 101(8):1199–1223, August 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bouajjani:2019:ESI

[3315] Ahmed Bouajjani and Hugues Fauconnier. Editorial, special issue of NETYS 2015. Computing, 101(9): 1225–1226, September 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00607-019-00743-2.pdf.

Hamza:2019:CL

[3316] Jad Hamza. On the complexity of linearizability. Computing, 101(9): 1227–1240, September 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Devismes:2019:OTE

[3317] Stéphane Devismes, Anissa Lamani, Franck Petit, and Sébastien Tixeuil. Optimal torus exploration by oblivious robots. Computing, 101(9):1241–1264, September 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Vo-Huu:2019:EEP

[3318] Triet Dang Vo-Huu, Erik-Oliver Blass, and Guevara Noubir. EPiC: efficient privacy-preserving counting for MapReduce. Computing, 101(9):1265– 1286, September 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

ElHammouti:2019:FDL

[3319] Hajar El Hammouti, Loubna Echabbi, and Rachid El Azouzi. A fully distributed learning algorithm for power allocation in heterogeneous networks. Computing, 101(9):1287–1303, September 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Benkaouz:2019:SIN

[3320] Yahya Benkaouz and Mohammed Erradi. Special issue on NETYS’17 selected papers. Computing, 101 (9):1305, September 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00607-019-00742-3.pdf.

Reaz:2019:PEB

[3321] Rezwana Reaz, H. B. Acharya, Ehab S. Elmallah, Jorge A. Cobb, and Mohamed G. Gouda. Policy expressions and the bottom-up design of computing policies. Computing, 101(9): 1307–1326, September 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Omar:2019:UPQ

[3322] Driss Ait Omar, M’Hamed Outanoute, Mohamed Baslam, Mohamed Fakir, and Belaid Bouikhalene. On understanding price–QoS war for competitive market and confused consumers. Computing, 101(9):1327–1348, September 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Talmage:2019:ASA

[3323] Edward Talmage and Jennifer L.

Welch. Anomalies and similarities among consensus numbers of variouslyrelaxed queues. Computing, 101(9): 1349–1368, September 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Avin:2019:CCS

[3324] Chen Avin, Louis Cohen, Mahmoud Parham, and Stefan Schmid. Competitive clustering of stochastic communication patterns on a ring. Computing, 101(9):1369–1390, September 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00607-
018-0666-x.pdf.

Ouhammou:2019:MDE

[3325] Yassine Ouhammou, Ladjel Bellatreche, Mirjana Ivanovic, and Alberto Abelló. Model and data engineering for advanced data-intensive systems and applications. Computing, 101(10): 1391–1395, October 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00607-019-00726-3.pdf.

Cai:2019:DAP

[3326] Simin Cai, Barbara Gallina, Dag Nyström, and Cristina Seceleanu. Data aggregation processes: a survey, a taxonomy, and design guidelines. Computing, 101(10):1397–1429, October 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s00607-
                                        018-0679-5.pdf.

Vassiliadis:2019:SEF

[3327] Panos Vassiliadis, Michail-Romanos Kolozoff, Maria Zerva, and Apostolos V. Zarras. Schema evolution and foreign keys: a study on usage, heartbeat of change and relationship of foreign keys to table activity. Computing, 101(10):1431–1456, October 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Yan:2019:ITR

[3328] Li Yan, Ping Zhao, and Zongmin Ma. Indexing temporal RDF graph. Computing, 101(10):1457–1488, October 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Qassimi:2019:RCT

[3329] Sara Qassimi and El Hassan Abdelwahed. The role of collaborative tagging and ontologies in emerging semantic of web resources. Computing, 101(10): 1489–1511, October 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Savic:2019:FSB

[3330] Milos Savić, Vladimir Kurbalija, Zoran Bosnić, and Mirjana Ivanović. Feature selection based on community detection in feature correlation networks. Computing, 101(10):1513–1538, October 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Roumelis:2019:BLB

[3331] George Roumelis, Athanasios Fevgas, Michael Vassilakopoulos, Antonio Corral, Panayiotis Bozanis, and Yannis Manolopoulos. Bulk-loading and bulkinsertion algorithms for xBR^{+}-trees in solid state drives. Computing, 101(10): 1539–1563, October 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Song:2019:NAM

[3332] Xiaoxiang Song, Yan Guo, Ning Li, and Peng Qian. A novel approach for missing data prediction in coevolving time series. Computing, 101(11): 1565–1584, November 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Garcia:2019:UEQ

[3333] Boni García, Micael Gallego, Francisco Gortázar, and Antonia Bertolino. Understanding and estimating quality of experience in WebRTC applications. Computing, 101(11):1585–1607, November 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kumar:2019:GAC

[3334] Ajay Kumar and Seema Bawa. Generalized Ant Colony Optimizer: swarmbased meta-heuristic algorithm for cloud services execution. Computing, 101(11):1609–1632, November 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chaudhry:2019:BIE

[3335] Rashmi Chaudhry and Shashikala Tapaswi. Bio-inspired energy conserving adaptive power and rate control in MANET. Computing, 101(11): 1633–1659, November 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Poniszewska-Maranda:2019:SUA

[3336] Aneta Poniszewska-Maranda, Daniel Kaczmarek, Natalia Kryvinska, and Fatos Xhafa. Studying usability of AI in the IoT systems/paradigm through embedding NN techniques into mobile smart service system. Computing, 101 (11):1661–1685, November 2019. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kolomvatsos:2019:DPI

[3337] Kostas Kolomvatsos. A distributed, proactive intelligent scheme for securing quality in large scale data processing. Computing, 101(11):1687– 1710, November 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00607-018-0683-9.pdf.

Berahmand:2019:NLM

[3338] Kamal Berahmand, Asgarali Bouyer, and Negin Samadi. A new local and multidimensional ranking measure to detect spreaders in social networks. Computing, 101(11):1711–1733, November 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Corno:2019:XCD

[3339] Fulvio Corno, Luigi De Russis, and Teodoro Montanaro. XDN: crossdevice framework for custom notifications management. Computing, 101 (11):1735–1761, November 2019. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mukherjee:2019:EBV

[3340] Sankar Mukherjee, Daya Sagar Gupta, and G. P. Biswas. An efficient and batch verifiable conditional privacypreserving authentication scheme for VANETs using lattice. Computing, 101 (12):1763–1788, December 2019. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mendonca:2019:AMA

[3341] Júlio Mendonça, Ermeson Andrade, and Ricardo Lima. Assessing mobile applications performance and energy consumption through experiments and stochastic models. Computing, 101(12): 1789–1811, December 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sakellariou:2019:HEK

[3342] Georgios Sakellariou and Anastasios Gounaris. Homomorphically encrypted k-means on cloud-hosted servers with low client-side load. Computing, 101 (12):1813–1836, December 2019. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gao:2019:PDL

[3343] Jun Gao, Yi Lu Murphey, and Honghui Zhu. Personalized detection of lane changing behavior using multisensor data fusion. Computing, 101(12): 1837–1860, December 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kim:2019:ETP

[3344] Dong-Oh Kim, Hong-Yeon Kim, Young-Kyun Kim, and Jeong-Joon Kim. Efficient techniques of parallel recovery for erasure-coding-based distributed file systems. Computing, 101(12):1861–1884, December 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kosmatopoulos:2019:HIV

[3345] Andreas Kosmatopoulos, Anastasios Gounaris, and Kostas Tsichlas. Hinode: implementing a vertex-centric modelling approach to maintaining historical graph data. Computing, 101 (12):1885–1908, December 2019. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Pufahl:2019:BAE

[3346] Luise Pufahl and Mathias Weske. Batch activity: enhancing business process modeling and enactment with batch processing. Computing, 101(12): 1909–1933, December 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ever:2019:PAE

[3347] Enver Ever, Purav Shah, Leonardo Mostarda, Fredrick Omondi, and Orhan Gemikonakli. On the performance, availability and energy consumption modelling of clustered IoT systems. Computing, 101(12):1935– 1970, December 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Jiang:2019:AFT

[3348] Jun Jiang, Liangcai Zeng, Bin Chen, Yang Lu, and Wei Xiong. An accurate and flexible technique for camera calibration. Computing, 101(12): 1971–1988, December 2019. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Li:2020:DDM

[3349] Yujun Li, Meijun Yang, Zhi Liu, Yuefeng Zhao, Dongmei Jiang, Lizhen Cui, and Mingyu Wang. Detection and diagnosis of myocarditis in young patients using ECG analysis based on artificial neural networks. Computing, 102(1):1–18, January 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Long:2020:NPD

[3350] Sifan Long, Ming Zhao, and Jieqiong Song. A novel PCA–DC-Bagging algorithm on yield stress prediction of RAFM steel. Computing, 102(1):19– 42, January 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See retraction [3738].

Goncalves:2020:RAB

[3351] Glauco Estácio Gonçalves, Patricia Takako Endo, Moises Rodrigues, Djamel H. Sadok, Judith Kelner, and Calin Curescu. Resource allocation based on redundancy models for high availability cloud. Computing, 102 (1):43–63, January 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Shang:2020:PSF

[3352] Yueqiang Shang. A parallel stabilized finite element method based on the lowest equal-order elements for incompressible flows. Computing, 102 (1):65–81, January 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chen:2020:DCU

[3353] Mingzhi Chen and Daqi Zhu. Data collection from underwater acoustic sensor networks based on optimization algorithms. Computing, 102(1):83–104, January 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lakhan:2020:TFA

[3354] Abdullah Lakhan and Xiaoping Li. Transient fault aware application partitioning computational offloading algorithm in microservices based mobile cloudlet networks. Computing, 102 (1):105–139, January 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Amiri:2020:NEA

[3355] Maryam Amiri, Leyli Mohammad-Khanli, and Raffaela Mirandola. A new efficient approach for extracting the closed episodes for workload prediction in cloud. Computing, 102 (1):141–200, January 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kulkarni:2020:EBI

[3356] Lalit Kulkarni, Jagdish Bakal, and Urmila Shrawankar. Energy based incentive scheme for secure opportunistic routing in vehicular delay tolerant networks. Computing, 102(1):201– 219, January 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schreiner:2020:BSD

[3357] Geomar A. Schreiner, Denio Duarte, and Ronaldo dos Santos Mello. Bringing SQL databases to key-based NoSQL databases: a canonical approach. Computing, 102(1):221–246, January 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Luo:2020:CAL

[3358] Xianxian Luo, Taisheng Zeng, Wei Zeng, and Jianlong Huang. Comparative analysis on Landsat image enhancement using fractional and integral differential operators. Computing, 102 (1):247–261, January 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ghose:2020:SRT

[3359] Manojit Ghose, Sawinder Kaur, and Aryabartta Sahu. Scheduling real time tasks in an energy-efficient way using VMs with discrete compute capacities. Computing, 102(1):263–294, January 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Koukaras:2020:SMT

[3360] Paraskevas Koukaras, Christos Tjortjis, and Dimitrios Rousidis. Social media types: introducing a data driven taxonomy. Computing, 102(1):295–340, January 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ferguson:2020:PSC

[3361] Donald F. Ferguson, Tim A. Majchrzak, and Ernesto Damiani. Preface: Springer Computing journal.

Computing, 102(2):341–342, February 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00607-
020-00790-0.pdf.

Saatkamp:2020:MFA

[3362] Karoline Saatkamp, Uwe Breitenbücher, Oliver Kopp, and Frank Leymann. Method, formalization, and algorithms to split topology models for distributed cloud application deployments. Computing, 102(2):343–363, February 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Prudhomme:2020:IAI

[3363] Claire Prudhomme, Timo Homburg, Jean-Jacques Ponciano, Frank Boochs, Christophe Cruz, and Ana-Maria Roxin. Interpretation and automatic integration of geospatial data into the semantic Web. Computing, 102(2): 365–391, February 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kotkov:2020:HDS

[3364] Denis Kotkov, Jari Veijalainen, and Shuaiqiang Wang. How does serendipity affect diversity in recommender systems? A serendipity-oriented greedy algorithm. Computing, 102 (2):393–411, February 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00607-018-0687-5.pdf.

Abidi:2020:ECP

[3365] Wided Abidi and Tahar Ezzedine. Effective clustering protocol based on network division for heterogeneous wireless sensor networks. Computing, 102(2):413–425, February 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rui:2020:NSF

[3366] Xiaobin Rui, Xiaodong Yang, Jianping Fan, and Zhixiao Wang. A neighbour scale fixed approach for influence maximization in social networks. Computing, 102(2):427–449, February 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Swain:2020:RAS

[3367] Chinmaya Kumar Swain, Neha Saini, and Aryabartta Sahu. Reliability aware scheduling of bag of real time tasks in cloud environment. Computing, 102 (2):451–475, February 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Nik:2020:CDW

[3368] Samaneh Sadat Mousavi Nik, Mahmoud Naghibzadeh, and Yasser Sedaghat. Cost-driven workflow scheduling on the cloud with deadline and reliability constraints. Computing, 102(2):477–500, February 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Al-Makhadmeh:2020:AHS

[3369] Zafer Al-Makhadmeh and Amr Tolba.

Automatic hate speech detection using killer natural language processing optimizing ensemble deep learning approach. Computing, 102(2):

501–522, February 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Huang:2020:EPN

[3370] Yuan-Ko Huang. Efficient processing of neighboring skyline queries with consideration of distance, quality, and cost. Computing, 102(2):523–550, February 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lytra:2020:QAU

[3371] Ioanna Lytra, Carlos Carrillo, Rafael Capilla, and Uwe Zdun. Quality attributes use in architecture design decision methods: research and practice. Computing, 102(2):551–572, February 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Liu:2020:DQI

[3372] Caihua Liu, Patrick Nitschke, Susan P. Williams, and Didar Zowghi. Data quality and the Internet of Things. Computing, 102(2):573–599, February 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wei:2020:SID

[3373] Wei Wei, Jinsong Wu, and Chunsheng Zhu. Special issue on deep learning for natural language processing. Computing, 102(3):601–603, March 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s00607-
                                        019-00788-3.pdf.

Alarifi:2020:OCA

[3374] Abdulaziz Alarifi and Ayed Alwadain. An optimized cognitive-assisted machine translation approach for natural language processing. Computing, 102(3):605–622, March 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Jiang:2020:STR

[3375] Zilong Jiang, Shu Gao, and Liangchen Chen. Study on text representation method based on deep learning and topic information. Computing, 102 (3):623–642, March 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Xia:2020:RSM

[3376] Ying Xia. Research on statistical machine translation model based on deep neural network. Computing, 102 (3):643–661, March 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Song:2020:ESR

[3377] Zhaojuan Song. English speech recognition based on deep learning with multiple features. Computing, 102 (3):663–682, March 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Xie:2020:CTC

[3378] Jinbao Xie, Yongjin Hou, Yujing Wang, Qingyan Wang, Baiwei Li, Shiwei Lv, and Yury I. Vorotnitsky. Chinese text classification based on attention mechanism and feature-enhanced fusion neural network. Computing, 102 (3):683–700, March 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Liu:2020:IFR

[3379] Wentao Liu, Weipeng Jing, and Yang Li. Incorporating feature representation into BiLSTM for deceptive review detection. Computing, 102(3):701–715, March 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wang:2020:SWE

[3380] Shirui Wang, Wenan Zhou, and Chao Jiang. A survey of word embeddings based on deep learning. Computing, 102(3):717–740, March 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wang:2020:ESI

[3381] Lizhe Wang, Dan Chen, Albert Zomaya, Laurence T. Yang, and Dimitrios Georgakopoulos. Editorial for the special issue on big time series data. Computing, 102(3):741–743, March 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00607-
020-00793-x.pdf.

Phengsuwan:2020:OBD

[3382] Jedsada Phengsuwan, Tejal Shah, Philip James, Dhavalkumar Thakker, Stuart Barr, and Rajiv Ranjan. Ontology-based discovery of time-series data sources for landslide early warning system. Computing, 102(3):745–763, March 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00607-
019-00730-7.pdf.

Xu:2020:TES

[3383] Ruyi Xu, Jingying Chen, Jiaxu Han, Lei Tan, and Luhui Xu. Towards emotion-sensitive learning cognitive state analysis of big data in education: deep learning-based facial expression analysis using ordinal information. Computing, 102(3):765–780, March 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ai:2020:FFC

[3384] Haojun Ai, Kaifeng Tang, Weiyi Huang, Sheng Zhang, and Taizhou Li. Fast fingerprints construction via GPR of high spatial-temporal resolution with sparse RSS sampling in indoor localization. Computing, 102 (3):781–794, March 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zeng:2020:MMM

[3385] Yi Zeng, Zihan Ning, Peng Liu, Peilei Luo, Yi Zhang, and Guojin He. A mosaic method for multi-temporal data registration by using convolutional neural networks for forestry remote sensing applications. Computing, 102 (3):795–811, March 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tu:2020:RBS

[3386] Weiping Tu, Yuhong Yang, Bo Du, Wanzhao Yang, Xiong Zhang, and Jiaxi Zheng. RNN-based signal classification for hybrid audio data compression. Computing, 102(3):813–827, March 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Almalki:2020:DAP

[3387] Faris A. Almalki and Marios C. Angelides. Deployment of an aerial platform system for rapid restoration of communications links after a disaster: a machine learning approach. Computing, 102(4):829–864, April 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s00607-
                                        019-00764-x.pdf.

Kolomvatsos:2020:PMA

[3388] Kostas Kolomvatsos and Christos Anagnostopoulos. A probabilistic model for assigning queries at the edge. Computing, 102(4):865–892, April 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s00607-
                                        019-00767-8.pdf.

Shirjini:2020:MSB

[3389] Mahsa Fozuni Shirjini, Saeed Farzi, and Amin Nikanjam. MDPCluster: a swarm-based community detection algorithm in large-scale graphs. Computing, 102(4):893–922, April 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Keshavarzi:2020:ETA

[3390] Amin Keshavarzi, Abolfazl Toroghi Haghighat, and Mahdi Bohlouli. Enhanced time-aware QoS prediction in multi-cloud: a hybrid k-medoids and lazy learning approach (QoPC). Computing, 102(4):923–949, April 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kunchala:2020:AMC

[3391] Jyothi Kunchala, Jian Yu, Sira Yongchareon, and Chengfei Liu. An approach to merge collaborating processes of an inter-organizational business process for artifact lifecycle synthesis. Computing, 102(4):951–976, April 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Al-Mouhamed:2020:RCO

[3392] Mayez A. Al-Mouhamed, Ayaz H.

Khan, and Nazeeruddin Mohammad. A review of CUDA optimization techniques and tools for structured grid computing. Computing, 102(4):977– 1003, April 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Serrato-Barrera:2020:SSD

[3393] Ricardo Serrato-Barrera, Gustavo Rodríguez-Gómez, Julio César Pérez-Sansalvador, Saúl Pomares-Hernández, Leticia Flores-Pulido, and Antonio Muñoz. Software system design based on patterns for Newton-type methods. Computing, 102(4):1005–1030, April 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gupta:2020:LRD

[3394] Shaifu Gupta and A. D. Dileep. Long range dependence in cloud servers: a statistical analysis based on Google workload trace. Computing, 102(4): 1031–1049, April 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Alokaili:2020:SEN

[3395] Amal Alokaili and Mohamed El Bachir Menai. SVM ensembles for named entity disambiguation. Computing, 102 (4):1051–1076, April 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ye:2020:GEC

[3396] Hui Ye, FangMin Li, ZhiXiong Liu, and XuDong Deng. A green energy consumption policy of Bluetooth mobile devices for smart cities. Computing, 102(4):1077–1091, April 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Cerquitelli:2020:ESI

[3397] Tania Cerquitelli, Silvia Chiusano, and Genoveva Vargas-Solar. Editorial special issue on data analytics for engineering, science and society. Computing, 102(5):1093–1096, May 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s00607-020-00810-z.pdf.

Bhatia:2020:QAL

[3398] Munish Bhatia, Sandeep K. Sood, and Simranpreet Kaur. Quantumized approach of load scheduling in fog computing environment for IoT applications. Computing, 102(5):1097–1115, May 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Haytamy:2020:DLB

[3399] Samar Haytamy and Fatma Omara. A deep learning based framework for optimizing cloud consumer QoS-based service composition. Computing, 102 (5):1117–1137, May 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Andrejev:2020:SAD

[3400] Andrej Andrejev, Kjell Orsborn, and Tore Risch. Strategies for array data retrieval from a relational back-end based on access patterns. Computing, 102 (5):1139–1158, May 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00607-020-00804-x.pdf.

Kniess:2020:DRS

[3401] Janine Kniess and Samuel Oliveira.

Data reduction in sensor networks based on dispersion analysis. Computing, 102(5):1159–1170, May 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Belhajjame:2020:PAE

[3402] Khalid Belhajjame, Noura Faci, Zakaria Maamar, Vanilson Burégio, Edvan Soares, and Mahmoud Barhamgi. On privacy-aware eScience workflows. Computing, 102(5):1171–1185, May 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fawagreh:2020:REF

[3403] Khaled Fawagreh and Mohamed Medhat Gaber. Resource-efficient fast prediction in healthcare data analytics: A pruned Random Forest regression approach. Computing, 102 (5):1187–1198, May 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s00607-019-00785-6.pdf.

Renga:2020:DDE

[3404] Daniela Renga, Daniele Apiletti, Danilo Giordano, Matteo Nisi, Tao Huang, Yang Zhang, Marco Mellia, and Elena Baralis. Data-driven exploratory models of an electric distribution network for fault prediction and diagnosis. Computing, 102(5):1199–1211, May 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Cagliero:2020:TEF

[3405] Luca Cagliero, Paolo Garza, Giuseppe Attanasio, and Elena Baralis. Training ensembles of faceted classification models for quantitative stock trading. Computing, 102(5):1213–1225, May 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chen:2020:MJM

[3406] Liuyan Chen and Lukasz Golab. Microjournal mining to understand mood triggers. Computing, 102(5):1227–1244, May 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Boratto:2020:CUB

[3407] Ludovico Boratto, Matteo Manca, Giuseppe Lugano, and Marián Gogola. Characterizing user behavior in journey planning. Computing, 102(5):1245– 1258, May 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Castells-Graells:2020:CRD

[3408] David Castells-Graells, Christopher Salahub, and Evangelos Pournaras.

On cycling risk and discomfort: urban safety mapping and bike route recommendations. Computing, 102 (5):1259–1274, May 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00607-019-00771-y.pdf.

Jabbour:2020:SBM

[3409] Said Jabbour, Nizar Mhadhbi, Badran Raddaoui, and Lakhdar Sais. SATbased models for overlapping community detection in networks. Computing, 102(5):1275–1299, May 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chang:2020:EAI

[3410] Victor Chang, Víctor Méndez Muñoz, and Muthu Ramachandran. Emerging applications of Internet of Things, Big Data, security, and complexity: special issue on collaboration opportunity for IoTBDS and COMPLEXIS. Computing, 102(6):1301–1304, June 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00607-
020-00811-y.pdf.

Li:2020:ADR

[3411] Yuming Li, Pin Ni, and Victor Chang. Application of deep reinforcement learning in stock trading strategies and stock forecasting. Computing, 102(6):1305–1322, June 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ralph:2020:RCS

[3412] David Ralph, Yunjia Li, Gary Wills, and Nicolas G. Green. Recommendations from cold starts in big data. Computing, 102(6):1323–1344, June 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s00607-
                                        020-00792-y.pdf.

Alsboui:2020:EDI

[3413] Tariq Alsboui, Yongrui Qin, Richard Hill, and Hussain Al-Aqrabi. Enabling distributed intelligence for the Internet of Things with IOTA and mobile agents. Computing, 102(6):1345–1363, June 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sivasubramanian:2020:NSF

[3414] Nandhini Sivasubramanian and Gunaseelan Konganathan. A novel semi fragile watermarking technique for tamper detection and recovery using IWT and DCT. Computing, 102 (6):1365–1384, June 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Standley:2020:FAD

[3415] Vaughn H. Standley, Frank G. Nuño, and Jacob W. Sharpe. Fusing attack detection and severity probabilities: a method for computing minimumrisk war decisions. Computing, 102 (6):1385–1408, June 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ranjan:2020:SDB

[3416] Rajiv Ranjan, Zheng Li, Massimo Villari, Yan Liu, and Dimitrios Georgeakopoulos. Software-driven big data analytics. Computing, 102 (6):1409–1417, June 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00607-020-00822-9.pdf.

Guo:2020:OCB

[3417] Ruixin Guo, Erkang Xue, Feng Zhang, Gansen Zhao, and Guangzhi Qu. Optimizing the confidence bound of countmin sketches to estimate the streaming big data query results more precisely. Computing, 102(6):1419–1445, June 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Xu:2020:SFD

[3418] Dong Xu, Yan Ma, Jining Yan, Peng Liu, and Lajiao Chen. Spatial-feature data cube for spatiotemporal remote sensing data processing and analysis. Computing, 102(6):1447–1461, June 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kaur:2020:CRM

[3419] Navroop Kaur, Sandeep K. Sood, and Prabal Verma. Cloud resource management using 3Vs of Internet of Big Data streams. Computing, 102 (6):1463–1485, June 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zhao:2020:LTR

[3420] Qi Zhao, Boxue Zhang, Wenquan Feng, Zhiying Du, Hong Zhang, and Daniel Sun. Long-term real time object tracking based on multi-scale local correlation filtering and global re-detection. Computing, 102(6):1487–1501, June 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00607-
020-00807-8.pdf.

Gupta:2020:FSE

[3421] Deepak Gupta, Anil K. Ahlawat, Arun Sharma, and Joel J. P. C. Rodrigues. Feature selection and evaluation for software usability model using modified moth-flame optimization. Computing, 102(6):1503–1520, June 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chen:2020:SSE

[3422] Lanxiang Chen, Nan Zhang, Hung-Min Sun, Chin-Chen Chang, Shui Yu, and Kim-Kwang Raymond Choo. Secure search for encrypted personal health records from big data NoSQL databases in cloud. Computing, 102 (6):1521–1545, June 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Yang:2020:IES

[3423] Chao-Tung Yang and Tsung-Yueh Wan. Implementation of an energy saving cloud infrastructure with virtual machine power usage monitoring and live migration on OpenStack. Computing, 102(6):1547–1566, June 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ugljanin:2020:ICT

[3424] Emir Ugljanin, Ejub Kajan, Zakaria Maamar, Muhammad Asim, and Vanilson Burégio. Immersing citizens and things into smart cities: a social machine-based and data artifact-driven approach. Computing, 102(7):1567– 1586, July 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Xia:2020:DMS

[3425] Xuewen Xia, Yichao Tang, Bo Wei, Yinglong Zhang, Ling Gui, and Xiong Li. Dynamic multi-swarm global particle swarm optimization. Computing, 102(7):1587–1626, July 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zamanirad:2020:DET

[3426] Shayan Zamanirad, Boualem Benatallah, Moshe Chai Barukh, Carlos Rodriguez, and Reza Nouri. Dynamic event type recognition and tagging for data-driven insights in lawenforcement. Computing, 102(7):1627– 1651, July 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mosavifard:2020:EAC

[3427] Aram Mosavifard and Hamid Barati.

An energy-aware clustering and twolevel routing method in wireless sensor networks. Computing, 102(7):1653– 1671, July 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Masdari:2020:EOS

[3428] Mohammad Masdari and Hemn Khezri. Efficient offloading schemes using Markovian models: a literature review. Computing, 102(7):1673–1716, July 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wang:2020:EEP

[3429] Qiang Wang, Haoran Cai, Qiang Cao, and Fang Wang. An energy-efficient power management for heterogeneous servers in data centers. Computing, 102(7):1717–1741, July 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Akhtar:2020:CBM

[3430] Usman Akhtar, Anita Sant’Anna, Chang-Ho Jihn, Muhammad Asif Razzaq, Jaehun Bang, and Sungyoung Lee. A cache-based method to improve query performance of linked Open Data cloud. Computing, 102(7):1743–1763, July 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Chimalakonda:2020:FSP

[3431] Sridhar Chimalakonda and Kesav V. Nori. A family of software product lines in educational technologies. Computing, 102(8):1765–1792, August 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bellendorf:2020:SCT

[3432] Julian Bellendorf and Zoltán Ádám Mann. Specification of cloud topologies and orchestration using TOSCA: a survey. Computing, 102(8):1793–1815, August 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Qian:2020:ADA

[3433] Zhongsheng Qian, Xiaojin Wang, Xiangyu Liu, Xiaoxin Xie, and Tao Song. An approach to dynamically assigning cloud resource considering user demand and benefit of cloud platform. Computing, 102(8):1817–1842, August 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Yongchareon:2020:RAC

[3434] Sira Yongchareon, Chengfei Liu, and Xiaohui Zhao. Reusing artifactcentric business process models: a behavioral consistent specialization approach. Computing, 102(8):1843–1879, August 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Li:2020:OKL

[3435] Lu Li and Xin Chen. Optimization of kernel learning algorithm based on parallel architecture. Computing, 102 (8):1881–1907, August 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Valderas:2020:SNS

[3436] Pedro Valderas, Victoria Torres, and Vicente Pelechano. A social network for supporting end users in the composition of services: definition and proof of concept. Computing, 102 (8):1909–1940, August 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kiran:2020:GBM

[3437] Utpal Kiran, Sachin Singh Gautam, and Deepak Sharma. GPU-based matrix-free finite element solver exploiting symmetry of elemental matrices. Computing, 102(9):1941–1965, September 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tang:2020:URB

[3438] Jian Tang, Mikel Larrea, Sergio Arévalo, and Ernesto Jiménez. Uniform reliable broadcast in anonymous distributed systems with fair lossy channels. Computing, 102(9):1967–1999, September 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Liu:2020:FPT

[3439] Chunhong Liu, Liping Dai, Yi Lai, Guibing Lai, and Wentao Mao. Failure prediction of tasks in the cloud at an earlier stage: a solution based on domain information mining. Computing, 102(9):2001–2023, September 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Aljubairy:2020:SEP

[3440] Abdulwahab Aljubairy, Wei Emma Zhang, Ali Shemshadi, Adnan Mahmood, and Quan Z. Sheng. A system for effectively predicting flight delays based on IoT data. Computing, 102(9): 2025–2048, September 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ghasemi:2020:MOL

[3441] Arezoo Ghasemi and Abolfazl Toroghi Haghighat. A multi-objective load balancing algorithm for virtual machine placement in cloud data centers based on machine learning. Computing, 102(9):2049–2072, September 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Priyadarsini:2020:EEL

[3442] Madhukrishna Priyadarsini, Shailesh Kumar, Padmalochan Bera, and Mohammad Ashiqur Rahman. An energyefficient load distribution framework for SDN controllers. Computing, 102 (9):2073–2098, September 2020. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Riahi:2020:CAM

[3443] Ali Riahi, Abdorreza Savadi, and Mahmoud Naghibzadeh. Comparison of analytical and ML-based models for predicting CPU–GPU data transfer time. Computing, 102(9):2099–2116, September 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Liu:2020:GSA

[3444] Jingsen Liu, Yuhao Xing, Yixiang Ma, and Yu Li. Gravitational search algorithm based on multiple adaptive constraint strategy. Computing, 102(10): 2117–2157, October 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dhule:2020:PSP

[3445] Chetan Dhule and Urmila Shrawankar.

POF-SVLM: Pareto optimized framework for seamless VM live migration. Computing, 102(10):2159–2183, October 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Dong:2020:ILP

[3446] Shi Dong. Improved label propagation algorithm for overlapping community detection. Computing, 102(10): 2185–2198, October 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Aguilar:2020:PAU

[3447] Jose Aguilar, M. Jerez, M. Mendonça, and M. Sánchez. Performance analysis of the ubiquitous and emergent properties of an autonomic reflective middleware for smart cities. Computing, 102(10):2199–2228, October 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Swain:2020:CAP

[3448] Chinmaya Kumar Swain, Bhawana Gupta, and Aryabartta Sahu. Constraint aware profit maximization scheduling of tasks in heterogeneous datacenters. Computing, 102(10): 2229–2255, October 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ursino:2020:AET

[3449] Domenico Ursino and Luca Virgili. An approach to evaluate trust and reputation of things in a Multi-IoTs scenario. Computing, 102(10):2257–2298, October 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tu:2020:ISD

[3450] Doan Quang Tu, A. S. M. Kayes, Wenny Rahayu, and Kinh Nguyen. IoT streaming data integration from multiple sources. Computing, 102(10):

2299–2329, October 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Anonymous:2020:RFD

[3451] Anonymous. Remembering Florian Daniel. Computing, 102(11):2331– 2332, November 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Anwar:2020:IRR

[3452] Tarique Anwar, Chengfei Liu, Hai L.

Vu, Md. Saiful Islam, Dongjin Yu, and Nam Hoang. Influence ranking of road segments in urban road traffic networks. Computing, 102(11): 2333–2360, November 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Farooq:2020:EOP

[3453] Umer Farooq and Bander A. Alzahrani.

Exploring and optimizing partitioning of large designs for multi-FPGA based prototyping platforms. Computing, 102(11):2361–2383, November 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Makitalo:2020:OSR

[3454] Niko Mäkitalo, Antero Taivalsaari, Arto Kiviluoto, Tommi Mikkonen, and Rafael Capilla. On opportunistic software reuse. Computing, 102(11): 2385–2408, November 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00607-020-00833-6.pdf.

Satti:2020:UHP

[3455] Fahad Ahmed Satti, Taqdir Ali, Jamil Hussain, Wajahat Ali Khan, Asad Masood Khattak, and Sungyoung Lee. Ubiquitous Health Profile (UHPr): a big data curation platform for supporting health data interoperability. Computing, 102(11):2409–2444, November 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00607-
020-00837-2.pdf.

Sahay:2020:NBB

[3456] Rashmi Sahay, G. Geethakumari, and Barsha Mitra. A novel blockchain based framework to secure IoT–LLNs against routing attacks. Computing, 102(11):2445–2470, November 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Yahyaoui:2020:NSR

[3457] Hamdi Yahyaoui, Hala S. Own, Ahmed Agwa, and Zakaria Maamar. A novel scalable representative-based forecasting approach of service quality. Computing, 102(11):2471–2500, November 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mchergui:2020:QEM

[3458] Abir Mchergui, Tarek Moulahi, and Salem Nasri. QoS evaluation model based on intelligent fuzzy system for vehicular ad hoc networks. Computing, 102(12):2501–2520, December 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Qaiser:2020:NEB

[3459] Hammad ur Rehman Qaiser and Gao Shu. Nash equilibrium based replacement of virtual machines for efficient utilization of cloud data centers. Computing, 102(12):2521–2540, December 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Najmuddin:2020:BBC

[3460] Sayed Najmuddin, Muhammad Asim, Kashif Munir, Thar Baker, Zehua Guo, and Rajiv Ranjan. A BBR-based congestion control for delay-sensitive realtime applications. Computing, 102(12): 2541–2563, December 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Carvalho:2020:THS

[3461] Paulo Carvalho, Solange Rito Lima, Luis Álvarez Sabucedo, Juan M. Santos-Gago, and João Marco C. Silva. Towards a holistic semantic support for context-aware network monitoring. Computing, 102(12):2565–2585, December 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wang:2020:EKG

[3462] Kai Wang, Yu Liu, Xiujuan Xu, and Quan Z. Sheng. Enhancing knowledge graph embedding by composite neighbors for link prediction. Computing, 102(12):2587–2606, December 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kang:2020:HHP

[3463] Homin Kang, Hyuck Chan Kwon, and Duksu Kim. HPMaX: heterogeneous parallel matrix multiplication using CPUs and GPUs. Computing, 102(12): 2607–2631, December 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Burattin:2021:SIB

[3464] Andrea Burattin, Jochen De Weerdt, Boudewijn van Dongen, Jan Claes, and Wil van der Aalst. Special issue on business process intelligence. Computing, 103(1):1–2, January 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00607-020-00856-z.pdf.

Tsoury:2021:HWD

[3465] Arava Tsoury, Pnina Soffer, and Iris Reinhartz-Berger. How well did it recover? Impact-aware conformance checking. Computing, 103(1):3–27, January 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Boltenhagen:2021:OSE

[3466] Mathilde Boltenhagen, Thomas Chatain, and Josep Carmona. Optimized SAT encoding of conformance checking artefacts. Computing, 103(1):29–50, January 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mann:2021:NAF

[3467] Zoltán Ádám Mann. Notions of architecture in fog computing. Computing, 103(1):51–73, January 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00607-020-00848-z.pdf.

Liu:2021:APP

[3468] Duen-Ren Liu, Yi-Kuan Hsu, Hsing-Yu Chen, and Huan-Jian Jau. Air pollution prediction based on factoryaware attentional LSTM neural network. Computing, 103(1):75–98, January 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Alkama:2021:IHR

[3469] Lynda Alkama and Louiza Bouallouche-Medjkoune. IEEE 802.15.4 historical revolution versions: a survey. Computing, 103(1):99–131, January 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Penurkar:2021:SCB

[3470] Milind R. Penurkar and Umesh A.

Deshpande. Social characteristicsbased routing algorithm for a mobile social network. Computing, 103 (1):133–153, January 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Domadiya:2021:IHS

[3471] Nikunj Domadiya and Udai Pratap Rao. Improving healthcare services using source anonymous scheme with privacy preserving distributed healthcare data collection and mining. Computing, 103(1):155–177, January 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zhao:2021:IDM

[3472] Liang Zhao, Jia Hu, Ahmed Al-Dubai, Celimuge Wu, Qi Wang, and Ammar Hawbani. Intelligence-driven mobile networks for smart cities.

Computing, 103(2):179–181, February 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00607-
020-00885-8.pdf.

Guo:2021:ESD

[3473] Yonghong Guo, Lu Liu, John Panneerselvam, and Rongbo Zhu. Efficient service discovery in mobile social networks for smart cities. Computing, 103 (2):183–209, February 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Al-qaness:2021:IYB

[3474] Mohammed A. A. Al-qaness, Aaqif Afzaal Abbasi, Hong Fan, Rehab Ali Ibrahim, Saeed H. Alsamhi, and Ammar Hawbani. An improved YOLO-based road traffic monitoring system. Computing, 103(2):211–230, February 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zhang:2021:MCR

[3475] Tianle Zhang, Liwen Zheng, Yu Jiang, Zhihong Tian, Xiaojiang Du, and Mohsen Guizani. A method of chained recommendation for charging piles in Internet of Vehicles. Computing, 103 (2):231–249, February 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00607-020-00832-7.pdf.

Sun:2021:LBL

[3476] Wei Sun, Pengyu Li, Zhi Liu, Xue Xue, Qiyue Li, Haiyan Zhang, and Junbo Wang. LSTM based link quality confidence interval boundary prediction for wireless communication in smart grid. Computing, 103(2):251–269, February 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Shang:2021:RID

[3477] Jiaxing Shang, Shuo Huang, Dingyang Zhang, Zixuan Peng, Dajiang Liu, Yong Li, and Lexi Xu. RNe2Vec: information diffusion popularity prediction based on repost network embedding. Computing, 103(2):271–289, February 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Almaini:2021:LEA

[3478] Amar Almaini, Ahmed Al-Dubai, Imed Romdhani, Martin Schramm, and Ayoub Alsarhan. Lightweight edge authentication for software defined networks. Computing, 103(2): 291–311, February 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00607-020-00835-4.pdf.

Ghannami:2021:DMD

[3479] Aiman Ghannami, Jing Li, Ammar Hawbani, and Naji Alhusaini. Diversity metrics for direct-coded variablelength chromosome shortest path problem evolutionary algorithms. Computing, 103(2):313–332, February 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Varma:2021:SAN

[3480] Ashwini Kumar Varma, Jyotirmoy Karjee, Debjani Mitra, Hemant Kumar Rath, and Arpan Pal. A selfadaptive network for multi-robot warehouse communication. Computing, 103 (2):333–356, February 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Barolli:2021:SII

[3481] Leonard Barolli, Farookh Hussain, and Makoto Takizawa. Special issue on intelligent edge, fog, cloud and Internet of Things (IoT)-based services. Computing, 103(3):357–360, March 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s00607-
                                        020-00892-9.pdf.

Xhafa:2021:AAF

[3482] Fatos Xhafa, Alhassan Aly, and Angel A. Juan. Allocation of applications to fog resources via semantic clustering techniques: with scenarios from intelligent transportation systems. Computing, 103(3):361–378, March 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Yakubu:2021:BBS

[3483] Bello Musa Yakubu, Majid I. Khan, Nadeem Javaid, and Abid Khan. Blockchain-based secure multi-resource trading model for smart marketplace. Computing, 103(3):379–400, March 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Park:2021:COU

[3484] Seunghyun Park and Hyunhee Park. Combined oversampling and undersampling method based on slow-start algorithm for imbalanced network traffic. Computing, 103(3):401–424, March 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kwon:2021:TBI

[3485] Sungmoon Kwon, Seongmin Park, HyungJin Cho, Youngkwon Park, Dowon Kim, and Kangbin Yim. Towards 5G-based IoT security analysis against Vo5G eavesdropping. Computing, 103(3):425–447, March 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00607-
020-00855-0.pdf.

Xie:2021:DPD

[3486] Yun Xie, Peng Li, Jindan Zhang, and Marek R. Ogiela. Differential privacy distributed learning under chaotic quantum particle swarm optimization. Computing, 103(3):449–472, March 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zhang:2021:INL

[3487] Ping Zhang, Mimoza Durresi, and Arjan Durresi. Internet network location privacy protection with multiaccess edge computing. Computing, 103(3):473–490, March 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Almansor:2021:SES

[3488] Ebtesam Hussain Almansor, Farookh Khadeer Hussain, and Omar Khadeer Hussain. Supervised ensemble sentimentbased framework to measure chatbot quality of services. Computing, 103 (3):491–507, March 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Salim:2021:KPB

[3489] Christian Salim and Nathalie Mitton. K-predictions based data reduction approach in WSN for smart agriculture. Computing, 103(3):509–532, March 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Calinescu:2021:SIR

[3490] Radu Calinescu and Felicita Di Giandomenico. Special issue on resilient software and software-controlled systems. Computing, 103(4):533– 534, April 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https://link.

                                        springer.com/article/10.1007/s00607-
                                        020-00887-6.

Vistbakka:2021:MRC

[3491] Inna Vistbakka and Elena Troubitsyna. Modelling resilient collaborative multi-agent systems. Computing, 103(4):535–557, April 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-020-00861-2.

Andersson:2021:CFR

[3492] Jesper Andersson, Vincenzo Grassi, and Diego Perez-Palacin. A conceptual framework for resilience: fundamental definitions, strategies and metrics. Computing, 103(4): 559–588, April 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-020-00874-x.

Parri:2021:FMD

[3493] Jacopo Parri, Fulvio Patara, and Enrico Vicario. A framework for Model-Driven Engineering of resilient software-controlled systems. Computing, 103(4):589–612, April 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-020-00841-
6.

Sanyal:2021:GTA

[3494] Judhajit Sanyal, Tuhina Samanta, and Mainak Chatterjee. Game theoretic approach for bandwidth sharing in future generation networks. Computing, 103(4):613–640, April 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-020-00872-z.

Jafarian:2021:SSS

[3495] Tohid Jafarian, Mohammad Masdari, and Kambiz Majidzadeh. SADM-SDNC: security anomaly detection and mitigation in software-defined networking using C-support vector classification. Computing, 103(4): 641–673, April 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-020-00866-x.

Machado:2021:AFA

[3496] Guilherme Medeiros Machado, Vinicius Maran, and José Palazzo Moreira de Oliveira. AwARE: a framework for adaptive recommendation of educational resources. Computing, 103 (4):675–705, April 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-00903-3.

Mehto:2021:ORP

[3497] Anjula Mehto, Shashikala Tapaswi, and K. K. Pattanaik. Optimal rendezvous points selection to reliably acquire data from wireless sensor networks using mobile sink. Computing, 103(4):707–733, April 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-00917-x.

Yu:2021:IFA

[3498] Shuhao Yu, Xukun Zuo, and Mingjing Pei. An improved firefly algorithm based on personalized step strategy. Computing, 103(4):735–748, April 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00607-021-00919-
                                        9.

Mittermayr:2021:DWA

[3499] Robert Mittermayr and Johann Blieberger. Deadlock and WCET analysis of barrier-synchronized concurrent programs. Computing, 103 (5):749–770, May 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-017-0555-8.

Kaya:2021:SCO

[3500] Mahir Kaya and Yasemin Çetin-Kaya. Seamless computation offloading for mobile applications using an online learning algorithm. Computing, 103 (5):771–799, May 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-020-00873-y.

doNascimento:2021:CRU

[3501] Leonardo Vianna do Nascimento, Guilherme Medeiros Machado, and José Palazzo M. de Oliveira. Context recognition and ubiquitous computing in smart cities: a systematic mapping. Computing, 103 (5):801–825, May 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-020-00878-7.

Hamzelou:2021:PTM

[3502] Nasrin Hamzelou, Mehrdad Ashtiani, and Raha Sadeghi. A propagation trust model in social networks based on the A* algorithm and multi-criteria decision making. Computing, 103 (5):827–867, May 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00918-w.

Montes-Orozco:2021:MQR

[3503] Edwin Montes-Orozco, Roman Anselmo Mora-Gutiérrez, and Miguel Ángel Gutiérrez-Andrade. Methodology to quantify robustness in networks: case study — Higher Education System in Mexico. Computing, 103 (5):869–893, May 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00909-x.

Kaur:2021:PBT

[3504] Gurleen Kaur and Anju Bala. Prediction based task scheduling approach for floodplain application in cloud environment. Computing, 103 (5):895–916, May 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-00936-8.

Mostafavi:2021:QSP

[3505] Seyedakbar Mostafavi, Vesal Hakami, and Maryam Sanaei. Quality of service provisioning in network function virtualization: a survey. Computing, 103(5):917–991, May 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-00925-x.

Carvalho:2021:ECC

[3506] Gonçalo Carvalho, Bruno Cabral, and Jorge Bernardino. Edge computing: current trends, research challenges and future directions. Computing, 103 (5):993–1023, May 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-020-00896-5.

Bocci:2021:SFO

[3507] Alessandro Bocci, Stefano Forti, and Antonio Brogi. Secure FaaS orchestration in the fog: how far are we? Computing, 103(5): 1025–1056, May 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-00924-y.

Mostefaoui:2021:GEQ

[3508] Ahmed Mostefaoui, Gabriele Gianini, and Geyong Min. Guest editorial: QoS and security in software for wireless and mobile micro-services. Computing, 103(6):1057–1059, June 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00938-6.

Wang:2021:VSN

[3509] Chao Wang, Ranbir Singh Batth, and Lihua Ren. VNE solution for network differentiated QoS and security requirements: from the perspective of deep reinforcement learning. Computing, 103 (6):1061–1083, June 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-020-00883-w.

Sani:2021:IBS

[3510] Mohammadreza Fani Sani, Sebastiaan J. van Zelst, and Wil M. P. van der Aalst. The impact of biased sampling of event logs on the performance of process discovery. Computing, 103 (6):1085–1104, June 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00910-4.

Martins:2021:MPB

[3511] Gabriel Martins, Luis Filipe Kopp, and Claudio Miceli de Farias. A multisensor prediction-based heuristic for the Internet of Things. Computing, 103 (6):1105–1120, June 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-020-00888-5.

Mansour:2021:DMC

[3512] Elio Mansour, Richard Chbeir, and Khouloud Salameh. Data management in connected environments. Computing, 103(6):1121–1142, June 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-020-00884-
9.

Merzoug:2021:SCP

[3513] Mohammed Amine Merzoug, Ahmed Mostefaoui, and Ernesto Damiani. Smart connected parking lots based on secured multimedia IoT devices. Computing, 103(6): 1143–1164, June 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00921-1.

Benslimen:2021:AFP

[3514] Yosra Benslimen, Hichem Sedjelmaci, and Ana-Cristina Manenti. Attacks and failures prediction framework for a collaborative 5G mobile network. Computing, 103(6): 1165–1181, June 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-020-00893-8.

Zroug:2021:HFM

[3515] Siham Zroug, Laid Kahloul, and Karim Djouani. A hierarchical formal method for performance evaluation of WSNs protocol. Computing, 103(6):1183–1208, June 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-020-00898-3.

Siddiqui:2021:AEA

[3516] Abdul Jabbar Siddiqui and Azzedine Boukerche. Adaptive ensembles of autoencoders for unsupervised IoT network intrusion detection. Computing, 103(6):1209–1232, June 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00912-2.

Rodrigo-Salazar:2021:IBC

[3517] Lucas Rodrigo-Salazar, Israel González-Carrasco, and Alejandro Rafael Garcia-Ramirez. An IoT-based contribution to improve mobility of the visually impaired in Smart Cities. Computing, 103 (6):1233–1254, June 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00947-5.

Regaieg:2021:MOO

[3518] Rym Regaieg, Mohamed Koubàa, and Taoufik Aguili. Multi-objective optimization for VM placement in homogeneous and heterogeneous cloud service provider data centers. Computing, 103(6):1255–1279, June 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00915-z.

Shishido:2021:OSC

[3519] Henrique Y. Shishido, Júlio C. Estrella, and Stephan Reiff-Marganiec. Optimizing security and cost of workflow execution using task annotation and genetic-based algorithm. Computing, 103(6):1281–1303, June 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00943-9.

Kallel:2021:GPV

[3520] Slim Kallel, Saoussen Cheikhrouhou, and Mohamed Jmaiel. From generating process views over inter-organizational business processes to achieving their temporal consistency. Computing, 103 (7):1305–1331, July 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00934-w.

Nogueira:2021:FBA

[3521] Bruno Nogueira and Erick Barboza. A FPGA-based accelerated architecture for the continuous GRASP. Computing, 103(7): 1333–1352, July 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-020-00850-5.

Madhura:2021:ILB

[3522] R. Madhura, B. Lydia Elizabeth, and V. Rhymend Uthariaraj. An improved list-based task scheduling algorithm for fog computing environment. Computing, 103(7): 1353–1389, July 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00935-9.

Debnath:2021:FLB

[3523] Arindam Debnath, Habila Basumatary, and Bidyut K. Bhattacharyya. Fuzzy logic-based VANET routing method to increase the QoS by considering the dynamic nature of vehicles. Computing, 103(7):1391–1415, July 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-020-00890-x.

Turabieh:2021:EHH

[3524] Hamza Turabieh, Sana Al Azwari, and Mrim Alnfiai. Enhanced Harris Hawks optimization as a feature selection for the prediction of student performance. Computing, 103 (7):1417–1438, July 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-020-00894-7.

Lu:2021:SNS

[3525] Chao Lu, Xunbo Li, and Xiang Li. Sensor network sensing coverage optimization with improved artificial bee colony algorithm using teaching strategy. Computing, 103 (7):1439–1460, July 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00906-0.

Dua:2021:MIC

[3526] Nidhi Dua, Shiva Nand Singh, and Vijay Bhaskar Semwal. Multi-input CNN-GRU based human activity recognition using wearable sensors. Computing, 103(7):1461–1478, July 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-021-00928-
8.

Verma:2021:ICA

[3527] Abhishek Verma and Virender Ranga. The impact of copycat attack on RPL based 6LoWPAN networks in Internet of Things. Computing, 103 (7):1479–1500, July 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-020-00862-1.

Awad:2021:CWS

[3528] Samir Awad, Abdelhamid Malki, and Mimoun Malki. Composing WoT services with uncertain and correlated data. Computing, 103(7): 1501–1517, July 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-020-00879-6.

Hu:2021:NEE

[3529] Shimin Hu, Simon Fong, and Jinan Fiaidhi. Novel evolutionary-EAC instance-learning-based algorithm for fast data stream mining in assisted living with extreme connectivity. Computing, 103(7): 1519–1543, July 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-020-00899-2.

Asghari:2021:CUC

[3530] Ali Asghari and Mohammad Karim Sohrabi. Combined use of coral reefs optimization and reinforcement learning for improving resource utilization and load balancing in cloud environments. Computing, 103(7): 1545–1567, July 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00920-2.

Wei:2021:NSA

[3531] Bo Wei, Xuan Wang, and Yanrong Huang. Novel self-adjusted particle swarm optimization algorithm for feature selection. Computing, 103(8): 1569–1597, August 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-020-00891-w.

Liao:2021:SGA

[3532] Zihan Liao, Wenxin Liang, and Xin Liu. Structure-guided attributed network embedding with “centroid” enhancement. Computing, 103(8): 1599–1620, August 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00916-y.

Cabrera:2021:CAM

[3533] Oscar Cabrera, Marc Oriol, and Jordi Marco. A context-aware monitoring architecture for supporting system adaptation and reconfiguration. Computing, 103(8): 1621–1655, August 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00923-z.

Mishra:2021:DDA

[3534] Amitabh Mishra, Achraf Cohen, and Norman Wilde. Detection of data anomalies at the edge of pervasive IoT systems. Computing, 103(8): 1657–1675, August 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-00927-9.

Tomar:2021:MAD

[3535] Abhinav Tomar and Prasanta K. Jana. A multi-attribute decision making approach for on-demand charging scheduling in wireless rechargeable sensor networks. Computing, 103(8): 1677–1701, August 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-020-00875-w.

Bose:2021:EEA

[3536] Rajesh Bose, Sandip Roy, and Srabanti Chakraborty. Energy-efficient approach to lower the carbon emissions of data centers. Computing, 103 (8):1703–1721, August 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-020-00889-4.

Asadi:2021:MBE

[3537] Ali Naghash Asadi, Mohammad Abdollahi Azgomi, and Reza Entezari-Maleki. Model-based evaluation of the power versus performance of network routing algorithms. Computing, 103 (8):1723–1746, August 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-020-00882-x.

Gupta:2021:DWA

[3538] Manjari Gupta, Lava Bhargava, and S. Indu. Dynamic workload-aware DVFS for multicore systems using machine learning. Computing, 103(8): 1747–1769, August 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-020-00845-2.

Xu:2021:SIB

[3539] Zhijun Xu, Jun Zhang, and Jihua Zhou.

A scheme for intelligent blockchainbased manufacturing industry supply chain management. Computing, 103 (8):1771–1790, August 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-020-00880-z.

Prasad:2021:FFN

[3540] B. M. Prabhu Prasad, Khyamling Parane, and Basavaraj Talawar. FPGA friendly NoC simulation acceleration framework employing the hard blocks. Computing, 103(8): 1791–1813, August 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-020-00901-x.

Gowri:2021:NSI

[3541] R. Gowri and R. Rathipriya. Nonswarm intelligence algorithms: a case study. Computing, 103(8): 1815–1857, August 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-020-00870-1.

Lindsay:2021:EDC

[3542] Dominic Lindsay, Sukhpal Singh Gill, and Peter Garraghan. The evolution of distributed computing systems: from fundamental to new frontiers. Computing, 103(8): 1859–1878, August 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-020-00900-y.

Anonymous:2021:EF

[3543] Anonymous. Editorial foreword. Computing, 103(9):1879–1882, September 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00607-021-00957-
                                        3.

Sharma:2021:BSD

[3544] N. Sharma, A. Anand, and A. K. Singh. Bio-signal data sharing security through watermarking: a technical survey. Computing, 103(9): 1883–1917, September 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-020-00881-y.

ElYebdri:2021:CAR

[3545] Zeyneb El Yebdri, Sidi Mohammed Benslimane, and Djamal Benslimane. Context-aware recommender system using trust network. Computing, 103(9):1919–1937, September 2021.

CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-020-00876-
9.

Sharma:2021:PVQ

[3546] Vishal Sharma, Kyumin Lee, and Curtis Dyreson. Popularity versus quality: analyzing and predicting the success of highly rated crowdfunded projects on Amazon. Computing, 103(9): 1939–1958, September 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00926-w.

De:2021:AEI

[3547] Suparna De, Wei Wang, and Mansour Naser Alraja. Analysing environmental impact of large-scale events in public spaces with cross-domain multimodal data fusion. Computing, 103(9): 1959–1981, September 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00944-8.

Sfaxi:2021:DIB

[3548] Lilia Sfaxi and Mohamed Mehdi Ben Aissa. Designing and implementing a Big Data benchmark in a financial context: application to a cash management use case. Computing, 103(9): 1983–2005, September 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00933-x.

HajKacem:2021:PTC

[3549] Mohamed Aymen Ben HajKacem, Chiheb-Eddine Ben N’cir, and Nadia Essoussi. A parallel text clustering method using spark and hashing. Computing, 103(9):2007– 2031, September 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-00932-y.

Ijaz:2021:EMO

[3550] Samia Ijaz, Ehsan Ullah Munir, and Omer F. Rana. Energy-makespan optimization of workflow scheduling in fogcloud computing. Computing, 103(9): 2033–2059, September 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-00930-0.

Uddin:2021:NRV

[3551] Jamal Uddin, Rozaida Ghazali, and Ijaz Ali Shoukat. A novel rough value set categorical clustering technique for supplier base management. Computing, 103(9):2061– 2091, September 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-00950-w.

Alsboui:2021:EEM

[3552] Tariq Alsboui, Yongrui Qin, and Hussain Al-Aqrabi. An energy efficient multi-mobile agent itinerary planning approach in wireless sensor networks. Computing, 103(9): 2093–2113, September 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00978-y.

Singhal:2021:JSA

[3553] Saurabh Singhal and Ashish Sharma.

A job scheduling algorithm based on rock hyrax optimization in cloud computing. Computing, 103(9):2115– 2142, September 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00942-w.

Menegazzo:2021:RST

[3554] Jeferson Menegazzo and Aldo von Wangenheim. Road surface type classification based on inertial sensors and machine learning. Computing, 103(10): 2143–2170, October 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00914-0.

Khalid:2021:FML

[3555] Yasir Noman Khalid, Muhammad Aleem, and Muhammad Azhar Iqbal. FusionCL: a machine-learning based approach for OpenCL kernel fusion to increase system performance. Computing, 103(10):2171– 2202, October 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00958-2.

Rodrigues:2021:SHV

[3556] Jairson B. Rodrigues, Germano C. Vasconcelos, and Paulo R. M. Maciel. Screening hardware and volume factors in distributed machine learning algorithms on spark. Computing, 103(10): 2203–2225, October 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00965-3.

Berahmand:2021:MDM

[3557] Kamal Berahmand, Elahe Nasiri, and Saman Forouzandeh. A modified DeepWalk method for link prediction in attributed social network. Computing, 103(10):2227– 2249, October 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00982-2.

Yan:2021:WMC

[3558] Yan Yan, Eyeleko Anselme Herman, and Pengshou Xie. A weighted Kmember clustering algorithm for kanonymization. Computing, 103(10): 2251–2273, October 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00922-0.

Abid:2021:IAD

[3559] Aymen Abid, Salim El Khediri, and Abdennaceur Kachouri. Improved approaches for density-based outlier detection in wireless sensor networks. Computing, 103(10): 2275–2292, October 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00939-5.

Li:2021:CCS

[3560] Liang Li, Qiong Lu, and Yuan Li. CMCSF: a collaborative service framework for mobile web augmented reality base on mobile edge computing. Computing, 103(10): 2293–2318, October 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00952-8.

Wang:2021:EFL

[3561] Qizhao Wang, Qing Li, and Peng Zeng. Efficient federated learning for fault diagnosis in industrial cloudedge computing. Computing, 103(10): 2319–2337, October 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00970-6.

Surya:2021:PRC

[3562] K Surya and V. Mary Anita Rajam. Prediction of resource contention in cloud using second order Markov model. Computing, 103(10): 2339–2360, October 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00967-1.

Roy:2021:OQD

[3563] Arnab Roy, Joseph Lalnunfela Pachuau, and Anish Kumar Saha. An overview of queuing delay and various delay based algorithms in networks. Computing, 103(10):2361– 2399, October 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00973-3.

Rodrigues:2021:PAE

[3564] Laécio Rodrigues, Igor Gonçalves, and Francisco Airton Silva. Performance and availability evaluation of an smart hospital architecture. Computing, 103(10):2401– 2435, October 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00979-x.

Aghaee:2021:SMH

[3565] Zahra Aghaee, Mohammad Mahdi Ghasemi, and Afsaneh Fatemi. A survey on meta-heuristic algorithms for the influence maximization problem in the social networks. Computing, 103(11):2437–2477, November 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-021-00945-
7.

Mezni:2021:TBS

[3566] Haithem Mezni, Mokhtar Sellami, and Faouzi Ben Charrada. Towards big services: a synergy between service computing and parallel programming. Computing, 103(11): 2479–2519, November 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00999-7.

Ntentos:2021:DBC

[3567] Evangelos Ntentos, Uwe Zdun, and Wilhelm Hasselbring. Detector-based component model abstraction for microservice-based systems. Computing, 103(11):2521–2551, November 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-021-01002-
z.

Noor:2021:SCM

[3568] Talal H. Noor. A service classification model for IoT services discovery. Computing, 103(11):2553– 2572, November 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01007-8.

Khan:2021:CAW

[3569] Ahmed Nawaz Khan, Muhammad Adnan Tariq, and Thar Baker. Congestion avoidance in wireless sensor network using software defined network. Computing, 103(11):2573– 2596, November 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01010-z.

Gavel:2021:DFB

[3570] Shashank Gavel, Raghavraju Charitha, and Ajay Singh Raghuvanshi. A data fusion based data aggregation and sensing technique for fault detection in wireless sensor networks. Computing, 103(11):2597– 2618, November 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01011-y.

ElOuadrhiri:2021:FMN

[3571] Ahmed El Ouadrhiri, Basma Nissar, and Imane Rahmouni. Fast-MICDTN: a new decentralized control mechanism for content-centric disruption tolerant networks. Computing, 103(11): 2619–2642, November 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-00940-y.

Hu:2021:BBP

[3572] Qinwen Hu, Muhammad Rizwan Asghar, and Sherali Zeadally. Blockchainbased public ecosystem for auditing security of software applications. Computing, 103(11):2643– 2665, November 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-00954-6.

Abdullah:2021:HAL

[3573] Fatima Abdullah, Dragi Kimovski, and Kashif Munir. Handover authentication latency reduction using mobile edge computing and mobility patterns. Computing, 103(11): 2667–2686, November 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-00969-z.

Yetkin:2021:SCF

[3574] E. Fatih Yetkin and Senol Piskin. Sensitivity of computational fluid dynamics simulations against soft errors. Computing, 103(11):2687– 2709, November 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00976-0.

ElMotaki:2021:PBM

[3575] Saloua El Motaki, Ali Yahyaouy, and Hamid Gualous. A prediction-based model for virtual machine live migration monitoring in a cloud datacenter. Computing, 103(11):2711– 2735, November 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00981-3.

Wang:2021:HMO

[3576] Dongxia Wang, Yongmei Lei, and Jianhui Zhou. Hybrid MPI/OpenMP parallel asynchronous distributed alternating direction method of multipliers. Computing, 103(12):2737– 2762, December 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00968-0.

Ibrahim:2021:GSD

[3577] Ahmed Ibrahim, Heng Zhang, and Simon Harper. From GPS to semantic data: how and why — a framework for enriching smartphone trajectories. Computing, 103(12): 2763–2787, December 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00993-z.

Hajipour:2021:EEE

[3578] Zahra Hajipour and Hamid Barati.

EELRP: energy efficient layered routing protocol in wireless sensor networks. Computing, 103(12):2789– 2809, December 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00996-w.

Jimenez:2021:CAA

[3579] Ernesto Jiménez, José Luis López-Presa, and Marta Patiño-Martínez. Consensus in anonymous asynchronous systems with crash-recovery and omission failures. Computing, 103(12): 2811–2837, December 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01023-8.

Zhou:2021:JOO

[3580] Shuchen Zhou and Waqas Jadoon. Jointly optimizing offloading decision and bandwidth allocation with energy constraint in mobile edge computing environment. Computing, 103(12): 2839–2865, December 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00931-z.

Dewangan:2021:VVB

[3581] Deepak Kumar Dewangan, Satya Prakash Sahu, and Aditi Agrawal. VLD-Net: Vision-based lane region detection network for intelligent vehicle system using semantic segmentation. Computing, 103(12):2867– 2892, December 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00974-2.

Semwal:2021:OFS

[3582] Vijay Bhaskar Semwal, Praveen Lalwani, and Jasroop Singh Chadha. An optimized feature selection using bio-geography optimization technique for human walking activities recognition. Computing, 103(12): 2893–2914, December 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01008-7.

Folstad:2021:FDC

[3583] Asbjørn Følstad, Theo Araujo, and Ewa Luger. Future directions for chatbot research: an interdisciplinary research agenda. Computing, 103(12): 2915–2942, December 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01016-7.

Memeti:2021:OHS

[3584] Suejb Memeti and Sabri Pllana. Optimization of heterogeneous systems with AI planning heuristics and machine learning: a performance and energy aware approach. Computing, 103(12): 2943–2966, December 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01017-6.

Khetarpaul:2021:MLB

[3585] Sonia Khetarpaul. Mining location based social networks to understand the citizen’s check-in patterns. Computing, 103(12):2967– 2993, December 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-01020-x.

Orangi:2021:FEN

[3586] Mina Abbaspour Orangi and Alireza Hashemi Golpayegani. Finding effective nodes to maximize the trusting behavior propagation in social networks. Computing, 103(12):2995– 3016, December 2021. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-00949-3.

Geng:2022:BFD

[3587] DaoQu Geng, HanWen He, and Chang Liu. Bearing fault diagnosis based on improved federated learning algorithm. Computing, 104 (1):1–19, January 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-01019-4.

Aghaalizadeh:2022:ISL

[3588] Saeid Aghaalizadeh, Saeid Taghavi Afshord, and Babak Anari. Improving the stability of label propagation algorithm by propagating from lowsignificance nodes for community detection in social networks. Computing, 104(1):21–42, January 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-01027-4.

Marimuthu:2022:ITC

[3589] M. Marimuthu, J. Akilandeswari, and P. R. Chelliah. Identification of trustworthy cloud services: solution approaches and research directions to build an automated cloud broker. Computing, 104(1): 43–72, January 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01015-8.

Hole:2022:TSA

[3590] Kjell Jørgen Hole. Tutorial on systems with antifragility to downtime. Computing, 104(1):73–93, January 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-020-00895-
6.

Gupta:2022:DNN

[3591] Manjari Gupta, Lava Bhargava, and S. Indu. Deep neural network learning for power limited heterogeneous system with workload classification. Computing, 104(1): 95–122, January 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01018-5.

Li:2022:BCT

[3592] Zheng Li, Pedro Pinacho-Davidson, and Rajiv Ranjan. Bonus computing: towards free-of-charge metacomputing in the public cloud. Computing, 104 (1):123–147, January 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01036-3.

Selvaraj:2022:IAD

[3593] Srinivasan Selvaraj. An incremental approach for detecting distributed deadlocks in the generalized model. Computing, 104(1): 149–168, January 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-00904-2.

Davami:2022:FBA

[3594] Fatemeh Davami, Sahar Adabi, and Amir Masoud Rahmani. Fog-based architecture for scheduling multiple workflows with high availability requirement. Computing, 104(1): 169–208, January 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-00905-1.

Koukaras:2022:IAE

[3595] Paraskevas Koukaras, Christos Tjortjis, and Dimitrios Tzovaras. An interdisciplinary approach on efficient virtual microgrid to virtual microgrid energy balancing incorporating data preprocessing techniques. Computing, 104 (1):209–250, January 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-00929-7.

S:2022:NIM

[3596] Vinod Chandra S. S. and Anand H. S. Nature inspired meta heuristic algorithms for optimization problems. Computing, 104(2):251– 269, February 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00955-5.

Lalwani:2022:CCP

[3597] Praveen Lalwani, Manas Kumar Mishra, and Pratyush Sethi. Customer churn prediction system: a machine learning approach. Computing, 104 (2):271–294, February 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00908-y.

Asmi:2022:GCS

[3598] Khawla Asmi, Dounia Lotfi, and Abdallah Abarda. The greedy coupledseeds expansion method for the overlapping community detection in social networks. Computing, 104(2): 295–313, February 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00948-4.

Melo:2022:MBA

[3599] Carlos Melo, Jean Araujo, and Paulo Maciel. A model-based approach for planning blockchain service provisioning. Computing, 104 (2):315–337, February 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00956-4.

Chin:2022:CDU

[3600] Jia Hou Chin and Kuru Ratnavelu.

Community detection using constrained label propagation algorithm with nodes exemption. Computing, 104 (2):339–358, February 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-00966-2.

Kumar:2022:MMD

[3601] Sanjay Kumar, Dipti Lohia, and B. S. Panda. MDER: modified degree with exclusion ratio algorithm for influence maximisation in social networks. Computing, 104(2): 359–382, February 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-00960-8.

Zhang:2022:SPM

[3602] Jixian Zhang, Laixin Chi, and Weidong Li. Strategy-proof mechanism for online resource allocation in cloud and edge collaboration. Computing, 104 (2):383–412, February 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-00962-6.

Li:2022:DEE

[3603] Hongjian Li, Hongxi Dai, and Yang Zou. Dynamic energy-efficient scheduling for streaming applications in storm. Computing, 104(2): 413–432, February 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-00961-7.

Ullah:2022:OSI

[3604] Israr Ullah, Muhammad Fayaz, and DoHyeun Kim. An optimization scheme for IoT based smart greenhouse climate control with efficient energy consumption. Computing, 104 (2):433–457, February 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00963-5.

Ozturk:2022:SIR

[3605] Ozcan Ozturk, Sabri Pllana, and Kaoutar El Maghraoui. Special issue on recent advances in autonomous vehicle solutions in the digital continuum. Computing, 104 (3):459–460, March 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01024-7.

Vega:2022:CBM

[3606] Augusto Vega, Alper Buyuktosunoglu, and Pradip Bose. Cloud-backed mobile cognition. Computing, 104 (3):461–479, March 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00953-7.

Tabani:2022:ABA

[3607] Hamid Tabani, Roger Pujol, and Francisco J. Cazorla. ADBench: benchmarking autonomous driving systems. Computing, 104(3):481–502, March 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-021-00975-

Bangui:2022:HML

[3608] Hind Bangui, Mouzhi Ge, and Barbora Buhnova. A hybrid machine learning model for intrusion detection in VANET. Computing, 104(3): 503–531, March 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-01001-0.

Zamouche:2022:USR

[3609] Djamila Zamouche, Mohamed Mohammedi, and Mawloud Omar. Ultrasafe and reliable enhanced traincentric communication-based train control system. Computing, 104 (3):533–552, March 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-01009-6.

Ghosh:2022:ESB

[3610] Partha Ghosh, Zaid Alam, and Santanu Phadikar. An efficient SGM based IDS in cloud environment. Computing, 104(3):553–576, March 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-022-01059-4.

Behera:2022:FTT

[3611] Lalatendu Behera. A fault-tolerant time-triggered scheduling algorithm of mixed-criticality systems. Computing, 104(3):577–599, March 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-01026-5.

Taghinezhad-Niar:2022:EEW

[3612] Ahmad Taghinezhad-Niar, Saeid Pashazadeh, and Javid Taheri. Energyefficient workflow scheduling with budget-deadline constraints for cloud. Computing, 104(3):601–625, March 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-021-01030-
9.

Yu:2022:MOO

[3613] Hong Yu, Zhao Fu, and Jie Li.

A multi-objective optimization algorithm based on dynamic userpreference information. Computing, 104(3):627–656, March 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00995-x.

Banerjee:2022:AMS

[3614] Suman Banerjee, Mamata Jenamani, and Dilip Kumar Pratihar. An approximate marginal spread computation approach for the budgeted influence maximization with delay. Computing, 104 (3):657–680, March 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00987-x.

Belgacem:2022:DRA

[3615] Ali Belgacem. Dynamic resource allocation in cloud computing: analysis and taxonomies. Computing, 104 (3):681–710, March 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-01045-2.

Bellatreche:2022:SIA

[3616] Ladjel Bellatreche and A Min Tjoa. Special issue on advances in data, information and knowledge engineering in data science era. Computing, 104(4):711–715, April 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-01032-7.

Lampridis:2022:MHC

[3617] Orestis Lampridis, Dimitra Karanatsiou, and Athena Vakali. MANI-FESTO: a huMAN-centric explaInable approach for FakE news spreaders deTectiOn. Computing, 104 (4):717–739, April 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-01013-w.

BenHassine:2022:NID

[3618] Mohamed Ali Ben Hassine, Safa Abdellatif, and Sadok Ben Yahia. A novel imbalanced data classification approach for suicidal ideation detection on social media. Computing, 104(4):741–765, April 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-00984-0.

Saenz-Adan:2022:ANI

[3619] Carlos Sáenz-Adán, Francisco J. García-Izquierdo, and Luc Moreau. Automated and non-intrusive provenance capture with UML2PROV.

Computing, 104(4):767–788, April 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-021-01012-
x.

Tarek:2022:MGI

[3620] Nayera Tarek, Mariam Abo Mandour, and Sara El-Metwally. Morse glasses: an IoT communication system based on Morse code for users with speech impairments. Computing, 104(4):789–808, April 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00959-1.

Brogi:2022:SHT

[3621] Antonio Brogi, Jose Carrasco, and Jacopo Soldani. Self-healing transcloud applications. Computing, 104 (4):809–833, April 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00977-z.

Reda:2022:HST

[3622] Roberto Reda, Filippo Piccinini, and Antonella Carbonaro. Heterogeneous self-tracked health and fitness data integration and sharing according to a linked open data approach. Computing, 104(4):835–857, April 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00988-w.

Touloum:2022:PMI

[3623] Soraya Touloum, Louiza Bouallouche-Medjkoune, and Aïssani Djamil. Performance modeling of the IEEE 802.15.4e TSCH enabling both shared and dedicated links in industrial WSNs. Computing, 104(4):859–891, April 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-021-00990-
2.

Tripathi:2022:CCF

[3624] Diwakar Tripathi, B. Ramachandra Reddy, and Alok Kumar Shukla. CFR: collaborative feature ranking for improving the performance of credit scoring data classification. Computing, 104(4):893–923, April 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01005-w.

Swain:2022:IAW

[3625] Chinmaya Kumar Swain and Aryabartta Sahu. Interference aware workload scheduling for latency sensitive tasks in cloud environment. Computing, 104(4):925–950, April 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01014-9.

Bahramnejad:2022:FAB

[3626] Somayeh Bahramnejad and Naser Movahhedinia. A fuzzy arithmeticbased analytical reliability assessment framework (FAARAF): case study, cognitive radio vehicular networks with drivers. Computing, 104 (4):951–979, April 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00980-4.

Chen:2022:IGO

[3627] Lei Chen, Yu Tian, and Yunpeng Ma. An improved grasshopper optimization algorithm based on dynamic dual elite learning and sinusoidal mutation. Computing, 104 (5):981–1015, May 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00991-1.

Liu:2022:OBD

[3628] Xiulei Liu, Ruoyu Chen, and Li Duan.

An ontology-based deep belief network model. Computing, 104(5): 1017–1032, May 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01021-w.

Almalki:2022:CMD

[3629] Faris A. Almalki, Abdullah A. Alotaibi, and Marios C. Angelides. Coupling multifunction drones with AI in the fight against the coronavirus pandemic. Computing, 104 (5):1033–1059, May 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01022-9.

Ilyas:2022:MLA

[3630] Muhammad U. Ilyas and Soltan Abed Alharbi. Machine learning approaches to network intrusion detection for contemporary Internet traffic. Computing, 104(5):1061–1076, May 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01050-5.

Kumari:2022:PLE

[3631] Anisha Kumari, Ranjan Kumar Behera, and Satya Prakash Sahoo. Prediction of link evolution using community detection in social network. Computing, 104(5): 1077–1098, May 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01035-4.

Pourabbasi:2022:NMS

[3632] Elmira Pourabbasi, Vahid Majidnezhad, and Yasser Jafari. A novel memorizing single chromosome evolutionary algorithm for detecting communities in complex networks. Computing, 104(5):1099–1122, May 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01033-6.

Shariat:2022:HEH

[3633] Sahar Samsam Shariat and Behrang Barekatain. HATMOG: an enhanced hybrid task assignment algorithm based on AHP-TOPSIS and multi-objective genetic in cloud computing. Computing, 104(5): 1123–1154, May 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01049-y.

Bai:2022:RDB

[3634] Na Bai, Fanrong Meng, and Zhixiao Wang. Rumor detection based on a Source-Replies conversation Tree Convolutional Neural Net. Computing, 104 (5):1155–1171, May 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01034-5.

Shukla:2022:NEB

[3635] Shivangi Shukla and Sankita J. Patel. A novel ECC-based provably secure and privacy-preserving multifactor authentication protocol for cloud computing. Computing, 104 (5):1173–1202, May 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01041-6.

Khan:2022:MLD

[3636] Hameed Khan, Kamal K. Kushwah, and Kishor Kumar Sadasivuni. Machine learning driven intelligent and self adaptive system for traffic management in smart cities. Computing, 104(5):1203–1217, May 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01038-1.

daCosta:2022:IRS

[3637] Luis Antonio L. F. da Costa, Rafael Kunst, and Edison Pignaton de Freitas. Intelligent resource sharing to enable quality of service for network clients: the trade-off between accuracy and complexity. Computing, 104 (5):1219–1231, May 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01042-5.

Tripathy:2022:SCS

[3638] Hrudaya Kumar Tripathy, Sushruta Mishra, Shubham Suman, Anand Nayyar, and Kshira Sagar Sahoo. Smart COVID-shield: an IoT driven reliable and automated prototype model for COVID-19 symptoms tracking. Computing, 104(6): 1233–1254, June 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01039-0.

Samieiyan:2022:SDR

[3639] Behrouz Samieiyan, Poorya MohammadiNasab, Mostafa Abbas Mollaei, Fahimeh Hajizadeh, and Mohammadreza Kangavari. Solving dimension reduction problems for classification using Promoted Crow Search Algorithm (PCSA). Computing, 104 (6):1255–1284, June 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01037-2.

Tran:2022:VMM

[3640] Cong Hung Tran, Thanh Khiet Bui, and Tran Vu Pham. Virtual machine migration policy for multi-tier application in cloud computing based on Qlearning algorithm. Computing, 104 (6):1285–1306, June 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01047-0.

Akbari:2022:EGS

[3641] Mohammad Reza Akbari, Hamid Barati, and Ali Barati. An efficient gray system theory-based routing protocol for energy consumption management in the Internet of Things using fog and cloud computing. Computing, 104(6):1307–1335, June 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01048-z.

Mahmood:2022:TWT

[3642] Adnan Mahmood, Sarah Ali Siddiqui, Quan Z. Sheng, Wei Emma Zhang, Hajime Suzuki, and Wei Ni. Trust on wheels: Towards secure and resource efficient IoV networks. Computing, 104 (6):1337–1358, June 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01040-7.

Chatterjee:2022:EBL

[3643] Uddalak Chatterjee, Sangram Ray, Muhammad Khurram Khan, Mou Dasgupta, and Chien-Ming Chen. An ECC-based lightweight remote user authentication and key management scheme for IoT communication in context of fog computing. Computing, 104 (6):1359–1395, June 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01055-8.

Smolka:2022:EFA

[3644] Sven Smolka and Zoltán Ádám Mann.

Evaluation of fog application placement algorithms: a survey. Computing, 104(6):1397–1423, June 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-01031-8.

Li:2022:SPA

[3645] Yuegang Li, Dongyang Ou, Xin Zhou, Congfeng Jiang, and Christophe Cérin. Scalability and performance analysis of BDPS in clouds. Computing, 104 (6):1425–1460, June 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-022-01056-7.

Du:2022:NSO

[3646] Wei Du, Gang Li, and Xiaochen He. Network structure optimization for social networks by minimizing the average path length. Computing, 104 (6):1461–1480, June 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-022-01061-w.

Almars:2022:UOE

[3647] Abdulqader M. Almars, El-Sayed Atlam, Talal H. Noor, Ghada ELmarhomy, Rasha Alagamy, and Ibrahim Gad. Users opinion and emotion understanding in social media regarding COVID-19 vaccine. Computing, 104(6):1481–1496, June 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00607-022-01062-
                                        9.

Fan:2022:SBS

[3648] Xing Fan, Baoning Niu, and Zhenliang Liu. Scalable blockchain storage systems: research progress and models. Computing, 104(6): 1497–1524, June 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01063-8.

Mostefaoui:2022:QQL

[3649] Ahmed Mostefaoui. QoS and QoE in large scale wireless IoT networks. Computing, 104(7):1525, July 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-022-01088-
z.

Naresh:2022:DRL

[3650] Mandan Naresh, Vikramjeet Das, Paresh Saxena, and Manik Gupta. Deep reinforcement learning based QoE-aware actor-learner architectures for video streaming in IoT environments. Computing, 104(7): 1527–1550, July 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01046-1.

Aljeri:2022:MMS

[3651] Noura Aljeri and Azzedine Boukerche.

Mobility management in space-ground-integrated networks. Computing, 104 (7):1551–1564, July 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01060-x.

Mansour:2022:DRM

[3652] Elio Mansour, Faisal Shahzad, Joe Tekli, and Richard Chbeir. Data redundancy management for leaf-edges in connected environments. Computing, 104(7):1565–1588, July 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-01051-4.

Singh:2022:CBE

[3653] Maninderpal Singh, Gagangeet Singh Aujla, Rasmeet Singh Bali, Ranbir Singh Batth, Amritpal Singh, Sahil Vashisht, and Anish Jindal. CovaDel: a blockchain-enabled secure and QoSaware drone delivery framework for COVID-like pandemics. Computing, 104(7):1589–1613, July 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-022-01064-7.

Fayad:2022:TDM

[3654] Moustafa Fayad, Ahmed Mostefaoui, Samir Chouali, and Salima Benbernou. Toward a design model-oriented methodology to ensure QoS of a cyber-physical healthcare system. Computing, 104(7): 1615–1641, July 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-022-01058-5.

Afzali:2022:EFT

[3655] Mahboubeh Afzali, Hamid Pourmohammadi, and Amin Mohammad Vali Samani. An efficient framework for trust evaluation of secure service selection in fog computing based on QoS, reputation, and social criteria. Computing, 104(7): 1643–1675, July 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01053-w.

Kettouche:2022:DMR

[3656] Souhila Kettouche, Moufida Maimour, and Lakhdar Derdouri. Disjoint multipath RPL for QoE/QoS provision in the Internet of Multimedia Things. Computing, 104(7): 1677–1699, July 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01054-9.

Cui:2022:NCB

[3657] Heqi Cui, Peng Sun, and Azzedine Boukerche. A novel cloud-based traffic aware data routing protocol for smart connected vehicles. Computing, 104 (7):1701–1720, July 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01068-3.

Kong:2022:PFT

[3658] Yongfeng Kong, Zhenhui Huang, and Wanjie Xiong. Parallelism of the finite-time dynamics method based on GPU. Computing, 104(8): 1721–1738, August 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01065-6.

Liu:2022:ADA

[3659] Jianghui Liu and Shengxiang Wang. All-dummy k-anonymous privacy protection algorithm based on location offset. Computing, 104(8): 1739–1751, August 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-022-01067-4.

Shojarazavi:2022:WMB

[3660] Taha Shojarazavi, Hamid Barati, and Ali Barati. A wrapper method based on a modified two-step league championship algorithm for detecting botnets in IoT environments. Computing, 104 (8):1753–1774, August 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-022-01070-9.

ElKhediri:2022:WSN

[3661] Salim El Khediri. Wireless sensor networks: a survey, categorization, main issues, and future orientations for clustering protocols. Computing, 104(8): 1775–1837, August 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-022-01071-8.

El-Zawawy:2022:FMI

[3662] Mohamed A. El-Zawawy, Parvez Faruki, and Mauro Conti. Formal model for inter-component communication and its security in Android. Computing, 104(8):1839– 1865, August 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01069-2.

Pioli:2022:ODR

[3663] Laércio Pioli, Carina F. Dorneles, Douglas D. J. de Macedo, and Mario A. R. Dantas. An overview of data reduction solutions at the edge of IoT systems: a systematic mapping of the literature. Computing, 104(8): 1867–1889, August 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01073-6.

Alkama:2022:PAU

[3664] Lynda Alkama, Louiza Bouallouche-Medjkoune, Mouloud Atmani, and Lina Bachiri. Performance analysis of the unslotted IEEE 802.15.4k MAC protocols under saturated traffic and fading channel conditions. Computing, 104(8):1891– 1922, August 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01074-5.

Bahramnejad:2022:REF

[3665] Somayeh Bahramnejad and Naser Movahhedinia. A reliability estimation framework for cognitive radio V2V communications and an ANN-based model for automating estimations. Computing, 104(8): 1923–1947, August 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01072-7.

Dai:2022:OHS

[3666] Hao Dai, Yang Wang, and Chengzhong Xu. Osprey: a heterogeneous search framework for spatial-temporal similarity. Computing, 104(9):1949– 1975, September 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-022-01075-4.

Li:2022:SLB

[3667] Tianhua Li, Yongcheng Yin, Bo Yang, Jialin Hou, and Kai Zhou. A selflearning bee colony and genetic algorithm hybrid for cloud manufacturing services. Computing, 104(9): 1977–2003, September 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-022-01079-0.

Sanda:2022:ICF

[3668] Pranitha Sanda, Digambar Pawar, and V. Radha. An insight into cloud forensic readiness by leading cloud service providers: a survey. Computing, 104(9):2005– 2030, September 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-022-01077-2.

Li:2022:NES

[3669] Jun Li, Hao Ren, ChenYang Li, and Huiling Chen. A novel and efficient salp swarm algorithm for largescale QoS-aware service composition selection. Computing, 104(9):2031– 2051, September 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01080-7.

DelAlamo:2022:SMS

[3670] Jose M. Del Alamo, Danny S.

Guaman, Boni García, and Ana Diez. A systematic mapping study on automated analysis of privacy policies. Computing, 104(9):2053– 2076, September 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01076-3.

Subaselvi:2022:EEU

[3671] S. Subaselvi and K. Gunaseelan. Energy efficient UAV enabled harvesting with beamforming for clustered SDWSN. Computing, 104(9):2077– 2100, September 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01087-0.

Kalantari:2022:GSO

[3672] Somayeh Kalantari, Eslam Nazemi, and Behrooz Masoumi. Guided self-organization through an entropybased self-advising approach. Computing, 104(9):2101–2136, September 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-022-01082-
5.

Pirzadi:2022:NRM

[3673] Somaye Pirzadi, Mohammad Ali Pourmina, and Seyed Mostafa Safavi-Hemami. A novel routing method in hybrid DTN–MANET networks in the critical situations. Computing, 104(9): 2137–2156, September 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01084-3.

Atig:2022:CJG

[3674] Mohamed Faouzi Atig. The Computing Journal gratefully acknowledges the editorial work of the scientists listed below on the special issue entitled “S.I. : ETYS 2019”. Computing, 104(10):2157, October 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01109-x.

Chaudhary:2022:ASF

[3675] Ved Prakash Chaudhary, Chirag Juyal, Sandeep Kulkarni, Sweta Kumari, and Sathya Peri. Achieving starvationfreedom in multi-version transactional memory systems. Computing, 104(10): 2159–2179, October 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00994-y.

Zennou:2022:CCC

[3676] Rachid Zennou, Ranadeep Biswas, Ahmed Bouajjani, Constantin Enea, and Mohammed Erradi. Checking causal consistency of distributed databases. Computing, 104(10): 2181–2201, October 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00911-3.

Chini:2022:LBN

[3677] Peter Chini, Roland Meyer, and Prakash Saivasan. Liveness in broadcast networks. Computing, 104(10): 2203–2223, October 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00986-y.

Jayanti:2022:RME

[3678] Prasad Jayanti and Anup Joshi.

Recoverable mutual exclusion with abortability. Computing, 104(10): 2225–2252, October 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01105-1.

Yu:2022:SBN

[3679] Bin Yu, Xiaofeng Li, He Zhao, and Tong Zhou. A scalable blockchain network model with transmission paths and neighbor node subareas. Computing, 104(10):2253– 2277, October 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00913-1.

Areias:2022:CLF

[3680] Miguel Areias and Ricardo Rocha.

On the correctness of a lock-free compression-based elastic mechanism for a hash trie design. Computing, 104(10):2279–2305, October 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-022-01085-

Sedrati:2022:IGI

[3681] Anass Sedrati, Aafaf Ouaddah, Abdellatif Mezrioui, and Badr Bellaj. IoT-Gov: an IoT governance framework using the blockchain. Computing, 104(10):2307–2345, October 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00607-022-01086-
                                        1.

Bai:2022:MTV

[3682] Tian Bai, Zhi-Jie Li, Bo Fan, and Jun Liao. Maintaining temporal validity of real-time data in componentbased systems. Computing, 104(11): 2347–2374, November 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-022-01089-y.

Perez-Castillo:2022:DCQ

[3683] Ricardo Pérez-Castillo and Mario Piattini. Design of classical-quantum systems with UML. Computing, 104(11): 2375–2403, November 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-022-01091-4.

Ihde:2022:FME

[3684] Sven Ihde, Luise Pufahl, Maximilian Völker, Asvin Goel, and Mathias Weske. A framework for modeling and executing task-specific resource allocations in business processes. Computing, 104(11):2405– 2429, November 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01093-2.

Xiao:2022:EDT

[3685] Danyang Xiao, Chengang Yang, and Weigang Wu. Efficient DNN training based on backpropagation parallelization. Computing, 104(11): 2431–2451, November 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01094-1.

Daher:2022:EVV

[3686] Ali W. Daher, Enrico Ferrari, Marco Muselli, Hussein Chible, and Daniele D. Caviglia. VAMPIRE: vectorized automated ML pre-processing and postprocessing framework for edge applications. Computing, 104(11):2453– 2487, November 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01096-z.

Jin:2022:HTL

[3687] Hong Jin, Cheng Jiang, Shengping Lv, Haiping He, and Xinting Liao. A hybrid teaching-learning-based optimization algorithm for QoS-aware manufacturing cloud service composition. Computing, 104(11):2489– 2509, November 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01083-4.

Afaq:2022:DCA

[3688] Yasir Afaq and Ankush Manocha.

Dew computing-assisted cognitive intelligence-inspired smart environment for diarrhea prediction. Computing, 104(11):2511–2540, November 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-022-01097-
y.

Huang:2022:IEG

[3689] Haichi Huang, Xuan Tian, Sisi Luo, and Yanli Shi. Item enhanced graph collaborative network for collaborative filtering recommendation. Computing, 104(12):2541– 2556, December 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01099-w.

Shojaee:2022:SMD

[3690] Reza Shojaee and Nasser Yazdani. Stochastic model-driven capacity planning framework for multi-access edge computing. Computing, 104(12): 2557–2579, December 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01102-4.

Haq:2022:OAW

[3691] N. U. Haq, M. M. Fraz, T. S. Hashmi, and M. Shahzad. Orientation aware weapons detection in visual data: a benchmark dataset. Computing, 104(12):2581–2604, December 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-022-01095-
0.

Thirumoorthy:2022:CAS

[3692] K. Thirumoorthy and J. Jerold John Britto. A clustering approach for software defect prediction using hybrid social mimic optimization algorithm. Computing, 104(12):2605– 2633, December 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01100-6.

Kumar:2022:OUA

[3693] Chintoo Kumar and C. Ravindranath Chowdary. OPHAencoder: An unsupervised approach to identify groups in group recommendations. Computing, 104(12):2635– 2657, December 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01103-3.

Toliopoulos:2022:ACM

[3694] Theodoros Toliopoulos and Anastasios Gounaris. Adaptivity in continuous massively parallel distance-based outlier detection. Computing, 104(12): 2659–2684, December 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01101-5.

Kettouch:2022:LEB

[3695] Mohamed Salah Kettouch and Cristina Luca. LinkD: element-based data interlinking of RDF datasets in linked data. Computing, 104(12):2685– 2709, December 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-022-01107-z.

Perez:2022:EC

[3696] Jorge Pérez, Jessica Díaz, Javier Berrocal, Ramón López-Viana, and Ángel González-Prieto. Edge computing. Computing, 104(12):2711– 2747, December 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-022-01104-2.

Helali:2022:ICD

[3697] Leila Helali and Mohamed Nazih Omri. Intelligent and compliant dynamic software license consolidation in cloud environment. Computing, 104(12): 2749–2783, December 2022. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-022-01106-0.

Hacid:2023:ESR

[3698] Hakim Hacid, Ons Al-Shamaileh, and Zakaria Maamar. An exploratory study of rethinking user experience in the internet of things. Computing, 105(1):1–12, January 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-022-01111-3.

Lv:2023:AAM

[3699] Jiandong Lv, Xingang Wang, and Cuiling Shao. AMAE: Adversarial multimodal auto-encoder for crisisrelated tweet analysis. Computing, 105(1):13–28, January 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01098-x.

Zhang:2023:PSS

[3700] Heng Zhang, Ahmed Ibrahim, Bijan Parsia, Ellen Poliakoff, and Simon Harper. Passive social sensing with smartphones: a systematic review. Computing, 105(1): 29–51, January 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01112-2.

Karaata:2023:SSE

[3701] Mehmet Hakan Karaata, Ebrahim Alrashed, and Mohammad Allaho. Selfstabilizing c-wave algorithms for arbitrary networks. Computing, 105 (1):53–88, January 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01110-4.

Elkhalil:2023:PHS

[3702] Ahmed Elkhalil and Jiashu Zhang.

Practical heterogeneous signcryption system for vehicular communication in VANETs. Computing, 105(1): 89–113, January 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01114-0.

Tian:2023:IES

[3703] Yang Tian, Gaofeng Nie, Hui Tian, and Qimei Cui. Impact of endpoint structure attributes on local information algorithms based on link prediction. Computing, 105(1): 115–129, January 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01115-z.

Liu:2023:PBF

[3704] Zhiguo Liu, Jiaojiao Zhang, Lin Wang, Jianxin Feng, Yuanming Ding, and ChangQing Ren. PSO-based feature extraction of unknown protocol data frame. Computing, 105(1): 131–149, January 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01118-w.

Saif:2023:SDB

[3705] Seyed Mojtaba Saif, Mohammad Ebrahim Samie, and Ali Hamzeh. A subgraphsdensity based overlapping community detection algorithm for large-scale complex networks. Computing, 105(1): 151–185, January 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01121-1.

Hosseini:2023:EES

[3706] Entesar Hosseini, Mohsen Nickray, and Shamsollah Ghanbari. Energyefficient scheduling based on task prioritization in mobile fog computing. Computing, 105(1):187– 215, January 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01108-y.

Zeedan:2023:EHM

[3707] Maha Zeedan, Gamal Attiya, and Nawal El-Fishawy. Enhanced hybrid multi-objective workflow scheduling approach based artificial bee colony in cloud computing. Computing, 105 (1):217–247, January 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01116-y.

Malekimajd:2023:MMD

[3708] Marzieh Malekimajd. Meddal:

meeting deadlines and data locality via bin packing in cloud environment. Computing, 105(2): 249–273, February 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01122-0.

Zhang:2023:IMB

[3709] Xinxin Zhang, Min Gao, Li Xu, and Zhaobin Zhou. Influence maximization based on SATS scheme in social networks. Computing, 105(2): 275–292, February 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01125-x.

Sirisuriya:2023:APO

[3710] S. C. M. S. De Sirisuriya, T. G. I.

Fernando, and M. K. A. Ariyaratne. Algorithms for path optimizations: a short survey. Computing, 105(2): 293–319, February 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-022-01126-w.

Debbi:2023:CBD

[3711] Hichem Debbi. Constraint-based debugging in probabilistic model checking. Computing, 105(2):321– 351, February 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-022-01130-0.

Devi:2023:TSB

[3712] K. Lalitha Devi and S. Valli. Time series-based workload prediction using the statistical hybrid model for the cloud environment. Computing, 105 (2):353–374, February 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-022-01129-7.

Ortiz:2023:MCB

[3713] Jesus Ortiz, Victoria Torres, and Pedro Valderas. Microservice compositions based on the choreography of BPMN fragments: facing evolution issues. Computing, 105(2): 375–416, February 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-022-01128-8.

Choudhary:2023:CDA

[3714] Chaitali Choudhary, Inder Singh, and Manoj Kumar. Community detection algorithms for recommendation systems: techniques and metrics. Computing, 105(2): 417–453, February 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01131-z.

Zhong:2023:HEG

[3715] Ying Zhong, Chenze Huang, and Qingbiao Zhou. HaSGP: an effective graph partition method for heterogeneous-aware. Computing, 105 (2):455–481, February 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01132-y.

Dhar:2023:EEA

[3716] Amit Kumar Dhar, Barun Gorain, Kaushik Mondal, Shaswati Patra, and Rishi Ranjan Singh. Edge exploration of anonymous graph by mobile agent with external help. Computing, 105 (2):483–506, February 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01136-8.

Mohammed:2023:SIA

[3717] Sabah Mohammed, Wai Chi Fang, and Carlos Ramos. Special issue on “artificial intelligence in cloud computing”. Computing, 105(3): 507–511, March 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00985-z. See cor-

rection [3718].

Mohammed:2023:CSI

[3718] Sabah Mohammed, Wai Chi Fang, and Carlos Ramos. Correction to:

Special issue on “artificial intelligence in cloud computing”. Computing, 105(3):513, March 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01149-x.  See

[3717].

Lin:2023:AWA

[3719] Weiwei Lin, Yufeng Zhang, Wentai Wu, Simon Fong, Ligang He, and Jia Chang. An adaptive workloadaware power consumption measuring method for servers in cloud data centers. Computing, 105(3): 515–538, March 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-020-00819-4.

Thi:2023:MLU

[3720] Thuy Le Thi, Tuoi Phan Thi, and Tho Quan Thanh. Machine learning using context vectors for object coreference resolution. Computing, 105 (3):539–558, March 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00902-4.

Cha:2023:SRT

[3721] Guang-Ho Cha. Similarity ranking technique exploiting the structure of similarity relationships. Computing, 105(3):559–576, March 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-020-00859-w.

You:2023:NAC

[3722] Deguang You, Weiwei Lin, Fang Shi, Jianzhuo Li, Deyu Qi, and Simon Fong. A novel approach for CPU load prediction of cloud server combining denoising and error correction. Computing, 105(3):577–594, March 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-020-00865-y.

Ihm:2023:UHU

[3723] Sun-Young Ihm, So-Hyun Park, and Young-Ho Park. UB-H: an unbalanced-hierarchical layer binarywise construction method for highdimensional data. Computing, 105 (3):595–616, March 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-020-00871-0.

Sood:2023:IHS

[3724] Sandeep Kumar Sood, Vaishali Sood, Isha Mahajan, and Sahil. An intelligent healthcare system for predicting and preventing dengue virus infection. Computing, 105 (3):617–655, March 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-020-00877-8.

Wang:2023:IRF

[3725] Ruiping Wang, Liangcai Zeng, Shiqian Wu, and Kelvin K. L. Wong. Illumination-robust feature detection based on adaptive threshold function. Computing, 105(3): 657–674, March 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-020-00868-9.

Girish:2023:ADC

[3726] L. Girish and Sridhar K. N. Rao. Anomaly detection in cloud environment using artificial intelligence techniques. Computing, 105(3): 675–688, March 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-00941-x.

Hu:2023:NIL

[3727] Quanyi Hu, Feng Wu, Raymond K. Wong, Richard C. Millham, and Jinan Fiaidhi. A novel indoor localization system using machine learning based on Bluetooth low energy with cloud computing. Computing, 105(3):689–715, March 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-020-00897-4.

Xiao:2023:PMG

[3728] Junsheng Xiao, Huahu Xu, Wei Zhao, Chen Cheng, and HongHao Gao. A prior-mask-guided few-shot learning for skin lesion segmentation. Computing, 105(3):717–739, March 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-00907-z.

Hossain:2023:AML

[3729] M. Shamim Hossain, Josu Bilbao, Diana P. Tobón, and Abdulmotaleb El Saddik. Advances of machine learning in IoT-cloud for healthcare. Computing, 105 (4):741–742, April 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01163-z.

Ahmed:2023:IES

[3730] Imran Ahmed, Gwanggil Jeon, and Abdellah Chehri. An IoT-enabled smart health care system for screening of COVID-19 with multi layers features fusion and selection. Computing, 105(4):743–760, April 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00992-0.

Qayyum:2023:APS

[3731] Abdul Qayyum, Moona Mazher, Aliyu Nuhu, Abdesslam Benzinou, Aamir Saeed Malik, and Imran Razzak. Assessment of physiological states from contactless face video: a sparse representation approach. Computing, 105(4):761–781, April 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01028-3.

Rohmetra:2023:AER

[3732] Honnesh Rohmetra, Navaneeth Raghunath, Pratik Narang, Vinay Chamola, Mohsen Guizani, and Naga Rajiv Lakkaniga. AI-enabled remote monitoring of vital signs for COVID-19: methods, prospects and challenges. Computing, 105(4): 783–809, April 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-00937-7.

Khan:2023:HIT

[3733] Suleman Khan, Muhammad Khurram Khan, and Rizwan Khan. Harnessing intelligent technologies to curb COVID-19 pandemic: taxonomy and open challenges. Computing, 105 (4):811–830, April 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-00983-1.

Arowolo:2023:MLB

[3734] Micheal Olaolu Arowolo, Roseline Oluwaseun Ogundokun, Sanjay Misra, Blessing Dorothy Agboola, and Brij Gupta. Machine learning-based IoT system for COVID-19 epidemics. Computing, 105(4):831–847, April 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-022-01057-6.

Mukherjee:2023:ICB

[3735] Rajendrani Mukherjee, Aurghyadip Kundu, Indrajit Mukherjee, Deepak Gupta, Prayag Tiwari, Ashish Khanna, and Mohammad Shorfuzzaman. IoTcloud based healthcare model for COVID-19 detection: an enhanced k-Nearest Neighbour classifier based approach. Computing, 105(4): 849–869, April 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-00951-9.

Nisar:2023:RTS

[3736] Shibli Nisar, Muhammad Ali Zuhaib, Abasin Ulasyar, and Muhammad Tariq. A robust tracking system for COVID-19 like pandemic using advanced hybrid technologies. Computing, 105(4):871–885, April 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-021-00946-
6.

Shorfuzzaman:2023:IES

[3737] Mohammad Shorfuzzaman. IoTenabled stacked ensemble of deep neural networks for the diagnosis of COVID-19 using chest CT scans. Computing, 105(4):887–908, April 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-021-00971-
5.

Long:2023:RNN

[3738] Sifan Long, Ming Zhao, and Jieqiong Song. Retraction note: A novel PCA– DC-Bagging algorithm on yield stress prediction of RAFM steel. Computing, 105(4):909, April 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01090-5.  See

[3350].

Llanos:2023:PDP

[3739] Diego R. Llanos and Dora B. Heras.

Parallel and distributed processing: advances on architectures and applications of parallel systems. Computing, 105(5):911, May 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01127-9. See cor-

rection [3740].

Llanos:2023:CPD

[3740] Diego R. Llanos and Dora B. Heras. Correction to: Parallel and distributed processing: advances on architectures and applications of parallel systems. Computing, 105(5):913, May 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01162-0.  See

[3739].

Barrachina:2023:UML

[3741] Sergio Barrachina, Adrián Castelló, Mar Catalán, Manuel F. Dolz, and Jose I. Mestre. Using machine learning to model the training scalability of convolutional neural networks on clusters of GPUs. Computing, 105(5):915–934, May 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00997-9.

Mankodi:2023:PPC

[3742] Amit Mankodi, Amit Bhatt, and Bhaskar Chaudhury. Predicting physical computer systems performance and power from simulation systems using machine learning model. Computing, 105(5):935–953, May 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01066-5.

Cook:2023:QCC

[3743] Victor Cook, Christina Peterson, Zachary Painter, and Damian Dechev. Quantifiability: a concurrent correctness condition modeled in vector space. Computing, 105(5): 955–978, May 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01092-3.

Agrawal:2023:MBL

[3744] Ankit Agrawal, Sarsij Tripathi, and Manu Vardhan. Multicore based least confidence query sampling strategy to speed up active learning approach for named entity recognition. Computing, 105(5):979–997, May 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01000-1.

Alcaraz:2023:PNT

[3745] Jordi Alcaraz, Ali TehraniJamsaz, Akash Dutta, Anna Sikora, Ali Jannesari, Joan Sorribes, and Eduardo Cesar. Predicting number of threads using balanced datasets for openMP regions. Computing, 105 (5):999–1017, May 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01081-6.

Carrizales-Espinoza:2023:FFP

[3746] Diana Carrizales-Espinoza, Dante D.

Sanchez-Gallegos, J. L. Gonzalez-Compean, and Jesus Carretero. Fed-Flow: a federated platform to build secure sharing and synchronization services for health dataflows. Computing, 105(5):1019–1037, May 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01044-3.

Vogel:2023:OTS

[3747] Adriano Vogel, Gabriele Mencagli, Dalvan Griebler, Marco Danelutto, and Luiz Gustavo Fernandes. Online and transparent self-adaptation of stream parallel patterns. Computing, 105 (5):1039–1057, May 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00998-8.

Dominico:2023:PLT

[3748] Simone Dominico, Eduardo C. de Almeida, and Marco A. Zanata Alves. On the performance limits of thread placement for array databases in non-uniform memory architectures. Computing, 105 (5):1059–1075, May 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01043-4.

Garcia:2023:SFC

[3749] Adriano Marques Garcia, Dalvan Griebler, Claudio Schepke, and Luiz Gustavo Fernandes. SP-Bench: a framework for creating benchmarks of stream processing applications. Computing, 105 (5):1077–1099, May 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01025-6.

Castello:2023:AIM

[3750] Adrián Castelló, Mar Catalán, Manuel F.

Dolz, Enrique S. Quintana-Ortí, and José Duato. Analyzing the impact of the MPI allreduce in distributed training of convolutional neural networks. Computing, 105(5): 1101–1119, May 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01029-2.

Fallah:2023:SBP

[3751] Mohammad K. Fallah, Mahmood Fazlali, and Masoud Daneshtalab. A symbiosis between population based incremental learning and LPrelaxation based parallel genetic algorithm for solving integer linear programming models. Computing, 105 (5):1121–1139, May 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-01004-x.

Flores-Martin:2023:TDH

[3752] Daniel Flores-Martin, Javier Berrocal, José García-Alonso, and Juan M. Murillo. Towards dynamic and heterogeneous social IoT environments. Computing, 105(6):1141–1164, June 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-022-01113-
1.

Jimenez:2023:UAB

[3753] Ernesto Jiménez, José Luis López-Presa, and Marta Patiño-Martínez.

Uniform atomic broadcast and consensus in fully anonymous synchronous systems with crash failures. Computing, 105(6):1165–1187, June 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-022-01135-
9.

Cruz:2023:MBA

[3754] Margarita Cruz, Beatriz Bernárdez, Amador Durán, Cathy Guevara-Vega, and Antonio Ruiz-Cortés. A modelbased approach for specifying changes in replications of empirical studies in computer science. Computing, 105 (6):1189–1213, June 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01133-x. See cor-

rection [3755].

Cruz:2023:CMB

[3755] Margarita Cruz, Beatriz Bernárdez, Amador Durán, Cathy Guevara-Vega, and Antonio Ruiz-Cortés. Correction to: A model-based approach for specifying changes in replications of empirical studies in computer science. Computing, 105(6): 1215–1216, June 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01143-9.  See

[3754].

Li:2023:EES

[3756] Xingxing Li, Weidong Li, and Xuejie Zhang. Extended efficiency and softfairness multiresource allocation in a cloud computing system. Computing, 105(6):1217–1245, June 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01138-6.

Wang:2023:IIM

[3757] Xiaojie Wang, Wushour Slamu, Abudureheman Kadeer, Sixiu Wang, and Xiaojing Hou. IMDCS: influence maximization with typediversity by leveraging community structure. Computing, 105(6): 1247–1270, June 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01134-w.

Barbosa:2023:DMC

[3758] Mariana S. M. Barbosa, Dianne S. V. Medeiros, and Miguel Elias M. Campista. Disjoint multipath closeness centrality. Computing, 105(6): 1271–1294, June 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01137-7.

Shastry:2023:CDT

[3759] K. Aditya Shastry and H. A. Sanjay. Customized decision tree-based approach for classification of soil on cloud environment. Computing, 105 (6):1295–1336, June 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01144-8.

Shahidani:2023:TSE

[3760] Fatemeh Ramezani Shahidani, Arezoo Ghasemi, Abolfazl Toroghi Haghighat, and Amin Keshavarzi. Task scheduling in edge-fog-cloud architecture: a multi-objective load balancing approach using reinforcement learning algorithm. Computing, 105(6): 1337–1359, June 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-022-01147-5.

Marwa:2023:MAS

[3761] Mokni Marwa, Jalel Eddine Hajlaoui, Yassa Sonia, Mohamed Nazih Omri, and Chelouah Rachid. Multi-agent system-based fuzzy constraints offer negotiation of workflow scheduling in fog–cloud environment. Computing, 105(7):1361–1393, July 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-022-01148-4.

Zhao:2023:CUS

[3762] Yang Zhao, Zhen Shao, Wei Zhao, Jun Han, Qingru Zheng, and Ran Jing. Combining unsupervised and supervised classification for customer value discovery in the telecom industry: a deep learning approach. Computing, 105(7):1395–1417, July 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-023-01150-4.

Bagherian:2023:CQC

[3763] Maryam Bagherian, Sarah Chehade, Ben Whitney, and Ali Passian. Classical and quantum compression for edge computing: the ubiquitous data dimensionality reduction. Computing, 105 (7):1419–1465, July 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01154-0.

Masiello-Ruiz:2023:SSA

[3764] Jose Manuel Masiello-Ruiz, Belen Ruiz-Mezcua, Paloma Martinez, and Israel Gonzalez-Carrasco. Synchro-Sub, an adaptive multi-algorithm framework for real-time subtitling synchronisation of multi-type TV programmes. Computing, 105(7): 1467–1495, July 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01156-y.

Hui:2023:MMM

[3765] Hongwen Hui, Fuhong Lin, Lei Meng, Lei Yang, and Xianwei Zhou. Many-to-many matching based task allocation for dispersed computing. Computing, 105(7):1497–1522, July 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01160-2.

Sahargahi:2023:NDI

[3766] Vahideh Sahargahi, Vahid Majidnezhad, Saeid Taghavi Afshord, and Yasser Jafari. A novel discrete ICO algorithm for influence maximization in complex networks. Computing, 105 (7):1523–1546, July 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-023-01157-x.

Devi:2023:IBS

[3767] Kalyanee Devi and Rohit Tripathi. Identification of best social media influencers using ICIRS model. Computing, 105(7):1547–1569, July 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-023-01159-9.

Wang:2023:TAS

[3768] Haoyu Wang, Jianwei An, and Hao Zhou. Task assignment strategy in LEO-multi-access edge computing based on matching game. Computing, 105(7):1571–1596, July 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-023-01151-3.

Sunil:2023:EEV

[3769] Shilpa Sunil and Sanjeev Patel. Energy-efficient virtual machine placement algorithm based on power usage. Computing, 105(7):1597–1621, July 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00607-023-01152-
                                        2.

Mystakidis:2023:EGF

[3770] Aristeidis Mystakidis, Evangelia Ntozi, Konstantinos Afentoulis, Paraskevas Koukaras, Paschalis Gkaidatzis, Dimosthenis Ioannidis, Christos Tjortjis, and Dimitrios Tzovaras. Energy generation forecasting: elevating performance with machine and deep learning. Computing, 105(8):1623– 1645, August 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01164-y.

Wu:2023:SDA

[3771] Yong Wu, Binjun Wang, Wei Li, and Wenmao Liu. Signed directed attention network. Computing, 105(8): 1647–1671, August 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01158-w.

Kiran:2023:GBF

[3772] Utpal Kiran, Deepak Sharma, and Sachin Singh Gautam. A GPU-based framework for finite element analysis of elastoplastic problems. Computing, 105 (8):1673–1696, August 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01169-7.

Dudyala:2023:NED

[3773] Anil Kumar Dudyala and Dinesh Dash.

A novel efficient on demand charging schedule for rechargeable wireless sensor networks. Computing, 105(8): 1697–1715, August 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01170-0.

Li:2023:EAS

[3774] Hongjian Li, Liang Lu, Wenhu Shi, Gangfan Tan, and Hao Luo. Energy-aware scheduling for spark job based on deep reinforcement learning in cloud. Computing, 105(8): 1717–1743, August 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01171-z.

Mena:2023:THA

[3775] Manel Mena, Javier Criado, Luis Iribarne, Antonio Corral, Richard Chbeir, and Yannis Manolopoulos. Towards high-availability cyberphysical systems using a microservice architecture. Computing, 105(8): 1745–1768, August 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01165-x.

Wang:2023:EAC

[3776] ZhiSheng Wang, Shu-Chuan Chu, JianPo Li, and Jeng-Shyang Pan. An energy-adaptive clustering method based on Taguchi-based-GWO optimizer for wireless sensor networks with a mobile sink. Computing, 105(8): 1769–1793, August 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01168-8.

Zhang:2023:PSO

[3777] Desong Zhang and Guangyu Zhu. Particle swarm optimization based on temporal-difference learning for solving multi-objective optimization problems. Computing, 105(8):1795– 1820, August 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01166-w.

Sahin:2023:TCF

[3778] Veysel Harun Sahin. Turna: a control flow graph reconstruction tool for RISC-V architecture. Computing, 105 (8):1821–1845, August 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01172-y.

Scandurra:2023:SIA

[3779] Patrizia Scandurra, Matteo Camilli, Raffaela Mirandola, and Sungwon Kang. Special issue on architecting dependable multi-paradigm computing systems. Computing, 105 (9):1847, September 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01202-9.

Andersson:2023:ADC

[3780] Jesper Andersson, Mauro Caporuscio, Mirko D’Angelo, and Annalisa Napolitano. Architecting decentralized control in large-scale self-adaptive systems. Computing, 105(9):1849– 1882, September 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01167-9.

Alfonso:2023:MBI

[3781] Iván Alfonso, Kelly Garcés, Harold Castro, and Jordi Cabot. A model-based infrastructure for the specification and runtime execution of self-adaptive IoT architectures. Computing, 105(9):1883– 1906, September 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01145-7.

Ferreira:2023:IRS

[3782] Luís Ferreira, Leopoldo Silva, Francisco Morais, Carlos Manuel Martins, Pedro Miguel Pires, Helena Rodrigues, Paulo Cortez, and André Pilastri. International revenue share fraud prediction on the 5G edge using federated learning. Computing, 105(9): 1907–1932, September 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01174-w.

Filippone:2023:ASS

[3783] Gianluca Filippone, Claudio Pompilio, Marco Autili, and Massimo Tivoli. An architectural style for scalable choreography-based microserviceoriented distributed systems. Computing, 105(9):1933–1956, September 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-022-01139-
5.

Soares:2023:TCE

[3784] Rodrigo C. Soares, Rafael Capilla, Vinicius dos Santos, and Elisa Yumi Nakagawa. Trends in continuous evaluation of software architectures. Computing, 105(9):1957– 1980, September 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01161-1.

Corradini:2023:XIM

[3785] Flavio Corradini, Arianna Fedeli, Fabrizio Fornari, Andrea Polini, Barbara Re, and Luca Ruschioni. X-IoT: a model-driven approach to support IoT application portability across IoT platforms. Computing, 105(9): 1981–2005, September 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01155-z.

Dadashi:2023:DDE

[3786] Morteza Dadashi and Amir Rajabzadeh. DAIP: a delay-efficient and availability-aware IoT application placement in fog environments. Computing, 105(9):2007– 2035, September 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01142-w.

Herrera:2023:CQA

[3787] Juan Luis Herrera, Javier Berrocal, Stefano Forti, Antonio Brogi, and Juan M. Murillo. Continuous QoSaware adaptation of Cloud–IoT application placements. Computing, 105(9): 2037–2059, September 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01153-1.

Fragiadakis:2023:CSC

[3788] George Fragiadakis, Vasiliki Liagkou, Evangelia Filiopoulou, Dimitris Fragkakis, Christos Michalakelis, and Mara Nikolaidou. Cloud services cost comparison: a clustering analysis framework. Computing, 105(10):2061– 2088, October 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01173-x.

Hannan:2023:SLR

[3789] Md. Abdul Hannan, Md. Atik Shahriar, Md Sadek Ferdous, Mohammad Jabed Morshed Chowdhury, and Mohammad Shahriar Rahman. A systematic literature review of blockchain-based e-KYC systems. Computing, 105(10):2089– 2118, October 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01176-8.

Sobhanayak:2023:MMO

[3790] Srichandan Sobhanayak. MOHBA:multiobjective workflow scheduling in cloud computing using hybrid BAT algorithm. Computing, 105(10):2119– 2142, October 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01175-9.

Tan:2023:EFB

[3791] Runnan Tan, Qingfeng Tan, Qin Zhang, Peng Zhang, Yushun Xie, and Zhao Li. Ethereum fraud behavior detection based on graph neural networks. Computing, 105(10): 2143–2170, October 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01177-7.

delaTorre-Lopez:2023:AIA

[3792] José de la Torre-López, Aurora Ramírez, and José Raúl Romero. Artificial intelligence to automate the systematic review of scientific literature. Computing, 105(10):2171– 2194, October 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01181-x.

Nandi:2023:DBF

[3793] Arijit Nandi, Fatos Xhafa, and Rohit Kumar. A Docker-based federated learning framework design and deployment for multi-modal data stream classification. Computing, 105(10): 2195–2229, October 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01179-5.

Saadi:2023:REF

[3794] Youssef Saadi, Soufiane Jounaidi, Said El Kafhali, and Hicham Zougagh. Reducing energy footprint in cloud computing: a study on the impact of clustering techniques and scheduling algorithms for scientific workflows. Computing, 105(10): 2231–2261, October 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01182-w.

Tigane:2023:GBF

[3795] Samir Tigane, Fayçal Guerrouf, and Laid Kahloul. A GSPN-based formalism under infinite-server semantics for reconfigurable WANETs. Computing, 105(10):2263–2291, October 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00607-023-01183-
                                        9.

Zhao:2023:CLE

[3796] Zhen Zhao and Shuo Lin. A crosslinguistic entity alignment method based on graph convolutional neural network and graph attention network. Computing, 105(10):2293– 2310, October 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-023-01178-6.

Zeng:2023:RAG

[3797] Yaohai Zeng, Bo Yang, Xiangchao Wen, and Dongsheng Li. Relevanceaware graph neural network for sessionbased recommendation. Computing, 105(10):2311–2335, October 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00607-023-01185-
                                        7.

Wang:2023:FDA

[3798] Zhongmin Wang, Zhihao Wei, Cong Gao, Yanping Chen, and Fengwei Wang. A framework for data anomaly detection based on iterative optimization in IoT systems. Computing, 105(11):2337– 2362, November 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01186-6.

Yan:2023:SRD

[3799] Yan Yan, Zichao Sun, Adnan Mahmood, Yiming Cong, Fei Xu, and Quan Z. Sheng. ST-Resnet: a deep learning-based privacy preserving differential publishing method for location statistics. Computing, 105(11): 2363–2387, November 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01189-3.

Strelak:2023:UFE

[3800] David Střelák, David Myška, Filip Petrovič, Jan Polák, Jaroslav Ol’ha, and Jiří Filipovič. Umpalumpa: a framework for efficient execution of complex image processing workloads on heterogeneous nodes. Computing, 105(11):2389– 2417, November 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01190-w.

Jin:2023:MIA

[3801] Yong Jin, Weidong Lou, and Yanghua Gao. Membership inference attacks against compression models. Computing, 105(11):2419– 2442, November 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01180-y.

Azarnia:2023:SSA

[3802] Ghanbar Azarnia. Steady-state analysis of diffusion least-mean squares with deficient length over wireless sensor networks. Computing, 105(11): 2443–2458, November 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-023-01187-5.

Moustakas:2023:CAT

[3803] Thanasis Moustakas and Kostas Kolomvatsos. Correlation adaptive task scheduling. Computing, 105(11): 2459–2486, November 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-023-01192-8.

Escobar:2023:DEE

[3804] Juan José Escobar, Francisco Rodríguez, Beatriz Prieto, Dragi Kimovski, Andrés Ortiz, and Miguel Damas. A distributed and energy-efficient KNN for EEG classification with dynamic money-saving policy in heterogeneous clusters. Computing, 105(11):2487– 2510, November 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-023-01193-7.

Wang:2023:PSL

[3805] Cong Wang, Peng Huo, Maode Ma, Tong Zhou, and Yiying Zhang. A provable secure and lightweight ECCbased authenticated key agreement scheme for edge computing infrastructure in smart grid. Computing, 105(11): 2511–2537, November 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01188-4.

Elkhalil:2023:PPO

[3806] Ahmed Elkhalil, Jiashu Zhang, Rashad Elhabob, and Nabeil Eltayieb. POOSC: Provably online/offline signcryption scheme for vehicular communication in VANETs. Computing, 105(11): 2539–2561, November 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01196-4.

Yadav:2023:EEP

[3807] Chandra Bhushan Kumar Yadav and Dinesh Dash. An energy efficient periodic data gathering and charging schedule using MVs in wireless rechargeable sensor networks. Computing, 105(11):2563– 2593, November 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01195-5.

Jiang:2023:FAC

[3808] Tao Jiang, Xiang-Yang Wei, Yue Li, Deng-Shan Wang, and Jin-Yun Yuan. A fast and accurate coupled meshless algorithm for the 2D/ 3D Gross–Pitaevskii equations on two GPUs. Computing, 105(12):2595– 2620, December 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01197-3.

Soltani:2023:DCA

[3809] Seyed Hossein Alavi Soltani, Saeed Jalili, and Mohammad Kazem Sheikh El Eslami. Decentralized control architecture for multi-authoring microgrids. Computing, 105(12):2621– 2646, December 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-023-01201-w.

Katta:2023:NTF

[3810] Monika Katta, T. K. Ramesh, and Juha Plosila. A novel technique for flit traversal in network-on-chip router. Computing, 105(12):2647– 2673, December 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-023-01200-x.

Koteich:2023:IPB

[3811] Jana Koteich, Christian Salim, and Nathalie Mitton. Image processing based data reduction technique in WVSN for smart agriculture. Computing, 105(12):2675– 2698, December 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-023-01198-2.

Laha:2023:FIB

[3812] Moyukh Laha and Raja Datta. Fuzzy intelligence based V2V routing protocol in Internet of Vehicles: a crosslayer approach. Computing, 105(12): 2699–2723, December 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01203-8.

Palacios:2023:EMA

[3813] Rodrigo Henrique Cunha Palácios, João Paulo Scarabelo Bertoncini, Gabriel Henrique Oliveira Uliam, Márcio Mendonça, and Lucas Botoni de Souza. Evaluation of mobile autonomous robot in trajectory optimization. Computing, 105(12):2725– 2745, December 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01205-6.

Chouat:2023:ACI

[3814] Houda Chouat, Imed Abbassi, Mohamed Graiet, and Mario Südholt. Adaptive configuration of IoT applications in the fog infrastructure. Computing, 105(12):2747– 2772, December 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01191-9.

Takagi:2023:DIE

[3815] Shiori Takagi, Shin’ichi Arakawa, and Masayuki Murata. Design, implementation and evaluation of a networkoriented service with environmental adaptability based on core/periphery structure. Computing, 105(12):2773– 2792, December 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01204-7.

Kaliya:2023:UFS

[3816] Neha Kaliya and Digambar Pawar. Unboxing fog security: a review of fog security and authentication mechanisms. Computing, 105(12):2793– 2819, December 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-023-01208-3.

Duarte:2023:MPD

[3817] Elias P. Duarte, Luiz A. Rodrigues, Edson T. Camargo, and Rogério C. Turchetti. The missing piece: a distributed system-level diagnosis model for the implementation of unreliable failure detectors. Computing, 105(12): 2821–2845, December 2023. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-023-01211-8.

Kumar:2024:EPV

[3818] K. Dinesh Kumar and E. Umamaheswari. An efficient proactive VM consolidation technique with improved LSTM network in a cloud environment. Computing, 106(1): 1–28, January 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-023-01214-5.

Wang:2024:FRU

[3819] Zhen-Fei Wang, Pei-Yao Yuan, Zhong-Ya Cao, and Li-Ying Zhang. Feature reduction of unbalanced data classification based on density clustering. Computing, 106(1):29–55, January 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-023-01206-
5.

Wang:2024:EEM

[3820] Zhenqian Wang and Huigui Rong.

An energy-efficient multi-stage alternating optimization scheme for UAV-mounted mobile edge computing networks. Computing, 106(1): 57–80, January 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01210-9.

Abdi:2024:NSN

[3821] Gholam Hossein Abdi, Amir Hossein Refahi Sheikhani, Sohrab Kordrostami, Armin Ghane, and Shahram Babaie. A novel selfish node detection based on reputation and game theory in Internet of Things. Computing, 106(1):81–107, January 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01184-8.

Khaledian:2024:EED

[3822] Navid Khaledian, Keyhan Khamforoosh, Reza Akraminejad, Laith Abualigah, and Danial Javaheri. An energy-efficient and deadlineaware workflow scheduling algorithm in the fog and cloud environment. Computing, 106(1): 109–137, January 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01215-4.

Kanwar:2024:EMN

[3823] Kushal Kanwar, Sakshi Kaushal, Harish Kumar, Gaurav Gupta, and Manju Khari. An efficient method for node ranking in complex networks by hybrid neighbourhood coreness. Computing, 106(1):139– 161, January 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-023-01218-1.

Rodriguez:2024:RMI

[3824] David Rodriguez, Jose M. Del Alamo, Miguel Cozar, and Boni García. ROI: a method for identifying organizations receiving personal data. Computing, 106(1): 163–184, January 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-023-01209-2.

Saglam:2024:LRL

[3825] Fatih Saglam and Mehmet Ali Cengiz. Local resampling for locally weighted Naïve Bayes in imbalanced data. Computing, 106(1): 185–200, January 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-023-01219-0.

Shamsa:2024:DPB

[3826] Zari Shamsa, Ali Rezaee, Sahar Adabi, and Amir Masoud Rahmani. A decentralized prediction-based workflow load balancing architecture for cloud/ fog/IoT environments. Computing, 106 (1):201–239, January 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01216-3.

Tsalikidis:2024:ELF

[3827] Nikos Tsalikidis, Aristeidis Mystakidis, Christos Tjortjis, Paraskevas Koukaras, and Dimosthenis Ioannidis. Energy load forecasting: onestep ahead hybrid model utilizing ensembling. Computing, 106(1): 241–273, January 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01217-2.

Ilmini:2024:DEA

[3828] WMKS Ilmini and TGI Fernando.

Detection and explanation of apparent personality using deep learning: a short review of current approaches and future directions. Computing, 106 (1):275–294, January 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01221-6.

Xu:2024:GHC

[3829] Xin Xu, Tao Ye, Jieying Gao, and Dongxiao Chu. Generalized hurdle count data models based on interpretable machine learning with an application to health care demand. Computing, 106(2):295– 325, February 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01224-3.

Fan:2024:IMS

[3830] Chencheng Fan, Zhixiao Wang, Jian Zhang, Jiayu Zhao, and Xiaobin Rui. Influence maximization in social networks based on discrete Harris Hawks optimization algorithm. Computing, 106(2):327– 351, February 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-023-01207-4.

Nallusamy:2024:PCD

[3831] Kalaichelvi Nallusamy and K. S. Easwarakumar. PERMDEC: community deception in weighted networks using permanence. Computing, 106 (2):353–370, February 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-023-01223-4.

Javaheri:2024:AAB

[3832] Seyed Danial Alizadeh Javaheri, Reza Ghaemi, and Hossein Monshizadeh Naeen. An autonomous architecture based on reinforcement deep neural network for resource allocation in cloud computing. Computing, 106 (2):371–403, February 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-023-01220-7.

Rani:2024:SSB

[3833] Prity Rani, Rohit Kumar Sachan, and Sonal Kukreja. A systematic study on blockchain technology in education: initiatives, products, applications, benefits, challenges and research direction. Computing, 106(2): 405–447, February 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01228-z.

Bellili:2024:GDL

[3834] Asma Bellili and Nadjia Kara. A graphical deep learning techniquebased VNF dependencies for multi resource requirements prediction in virtualized environments. Computing, 106 (2):449–473, February 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01225-2.

Lotito:2024:ESM

[3835] Quintino Francesco Lotito, Federico Musciotto, Federico Battiston, and Alberto Montresor. Exact and sampling methods for mining higher-order motifs in large hypergraphs. Computing, 106 (2):475–494, February 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01230-5.

Morovatian:2024:RNR

[3836] Iman Morovatian, Alireza Basiri, and Samira Rezaei. RUCIB: a novel rule-based classifier based on BRADO algorithm. Computing, 106 (2):495–519, February 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01226-1.

Li:2024:MLP

[3837] Min Li, Shuming Zhou, Dajin Wang, and Gaolin Chen. Missing link prediction using path and community information. Computing, 106(2): 521–555, February 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-023-01229-y.

Sagar:2024:CWQ

[3838] Subhash Sagar, Adnan Mahmood, Quan Z. Sheng, Munazza Zaib, and Farhan Sufyan. Can we quantify trust? Towards a trust-based resilient SIoT network. Computing, 106(2): 557–577, February 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-023-01236-z.

Wang:2024:DGB

[3839] Haoyu Wang and Jianwei An. Dynamic game based task offloading and resource pricing in LEO-multi-access edge computing. Computing, 106(2): 579–606, February 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-023-01234-1.

Almars:2024:CSR

[3840] Abdulqader M. Almars and Khalid M. Alharbi. Corporate social responsibility disclosure prediction using LSTM neural network. Computing, 106(2): 607–625, February 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01239-w.

Li:2024:NOA

[3841] Weidong Li, Chunbo Shi, Yongbo Yu, and Zhe Wang. A novel optimization approach to topology checking of pipeline vector data in browser side. Computing, 106(2): 627–649, February 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01241-2.

Mrabet:2024:SAC

[3842] Faten Mrabet, Walid Karamti, and Adel Mahfoudhi. Scheduling analysis and correction for dependent realtime tasks upon heterogeneous multiprocessor architectures. Computing, 106(3):651–712, March 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01237-y.

Qin:2024:DCL

[3843] Chuandong Qin and Yu Cao. 1-D CNNs with lock-free asynchronous adaptive stochastic gradient descent algorithm for classification of astronomical spectra. Computing, 106 (3):713–739, March 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01240-3.

Liao:2024:ATC

[3844] Mingxue Liao and Pin Lv. On the average time complexity of computation with random partition. Computing, 106(3):741–758, March 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01242-1.

Rajeswari:2024:NID

[3845] G. Rajeswari, R. Arthi, and K. Murugan. Nature-inspired donkey and smuggler algorithm for optimal data gathering in partitioned wireless sensor networks for restoring network connectivity. Computing, 106(3): 759–787, March 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01251-0.

Lyu:2024:UAW

[3846] Ting Lyu, Jianwei An, Meng Li, Feifei Liu, and Haitao Xu. UAV-assisted wireless charging and data processing of power IoT devices. Computing, 106 (3):789–819, March 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01245-y.

Montes-Orozco:2024:QIN

[3847] Edwin Montes-Orozco, Roman-Anselmo Mora-Gutiérrez, Sergio-Gerardo delos Cobos-Silva, Eric-Alfredo Rincón-García, Miguel-Ángel Gutiérrez-Andrade, and Pedro Lara-Velázquez. Quantifying influential nodes in complex networks using optimization and particle dynamics: a comparative study. Computing, 106(3): 821–864, March 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01244-z.

Du:2024:HHG

[3848] Hui Du, Cuntao Ma, Depeng Lu, and Jingrui Liu. HHSE: heterogeneous graph neural network via higher-order semantic enhancement. Computing, 106(3):865–887, March 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01246-x.

Perez-Vereda:2024:DMC

[3849] Alejandro Pérez-Vereda, Carlos Canal, and Ramón Hervás. Distributed mobile CEP for collaborative social computing. Computing, 106 (3):889–913, March 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01254-x.

Sun:2024:MNM

[3850] Xiaochuan Sun, Qi Chen, Mingxiang Hao, Yingqi Li, and Bo Sun. Mconvkgc: a novel multi-channel convolutional model for knowledge graph completion. Computing, 106 (3):915–937, March 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01247-w.

Xie:2024:TTE

[3851] Hui Xie, Jun Zheng, Zhengyuan Zhang, Shengjun Wei, and Changzhen Hu. TEDA: a trusted execution environment-and-blockchain-based data protection architecture for Internet of Things. Computing, 106(3): 939–960, March 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01253-y.

Hasan:2024:CDW

[3852] Ahmadi Hasan and Ahmad Kamal. Community detection of weighted complex networks via transitive closure. Computing, 106(3): 961–979, March 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01249-8.

Guruprasad:2024:PAL

[3853] Guruprasad Kapilesh, Sridhar Dhanush, Venkatesan Poovazhaki Gokula Kannan, and Viswasam Mary Anita Rajam. A priority aware local mutual exclusion algorithm for flying ad hoc networks. Computing, 106 (3):981–1012, March 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01250-1.

Heidari:2024:NEA

[3854] Ehsan Heidari. A novel energyaware method for clustering and routing in IoT based on whale optimization algorithm & Harris Hawks optimization. Computing, 106(3): 1013–1045, March 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01252-z.

Ranjan:2024:AML

[3855] Rajiv Ranjan, Dan Chen, and Prem Prakash Jayaraman. Adaptive machine learning in the IoTedge-cloud continuum. Computing, 106(4):1047, April 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01231-4. See cor-

rection [3856].

Ranjan:2024:CAM

[3856] Rajiv Ranjan, Dan Chen, and Prem Prakash Jayaraman. Correction to: Adaptive machine learning in the IoT-edge-cloud continuum. Computing, 106(4):1049, April 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01274-1.  See

[3855].

Anoushee:2024:IRM

[3857] Milad Anoushee, Mehdi Fartash, and Javad Akbari Torkestani. An intelligent resource management method in SDN based fog computing using reinforcement learning. Computing, 106 (4):1051–1080, April 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01141-x.

Li:2024:AVF

[3858] Yuanzhang Li, Tianchi Sha, Thar Baker, Xiao Yu, Zhiwei Shi, and Sikang Hu. Adaptive vertical federated learning via feature map transferring in mobile edge computing. Computing, 106 (4):1081–1097, April 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01117-x.

Ghaffari:2024:RTD

[3859] Hamidreza Ghaffari and Farzaneh Rafeie. Reconstructing the training data set based on reducing boundary complexity. Computing, 106(4): 1099–1119, April 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01124-y.

Xu:2024:OIB

[3860] Ruyi Xu, Jiaxu Han, and Jingying Chen. Ordinal information based facial expression intensity estimation for emotional interaction: a novel semi-supervised deep learning approach. Computing, 106(4): 1121–1138, April 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01140-y.

Cheng:2024:AML

[3861] Ming Cheng, Yunbing Yan, Yang Han, Hongjun Lei, and Chao Deng. Adaptive machine learning algorithm for human target detection in IoT environment. Computing, 106(4): 1139–1150, April 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01123-z.

Mittal:2024:DTO

[3862] Shilpi Mittal, Rajan Kumar Dudeja, Rasmeet Singh Bali, and Gagangeet Singh Aujla. A distributed task orchestration scheme in collaborative vehicular cloud edge networks. Computing, 106 (4):1151–1175, April 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01119-9.

Chhikara:2024:AFL

[3863] Prateek Chhikara, Rajkumar Tekchandani, and Neeraj Kumar. Adaptive federated learning scheme for recognition of malicious attacks in an IoT network. Computing, 106 (4):1177–1192, April 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01146-6.

Sahil:2024:FCI

[3864] Sahil, Sandeep Kumar Sood, and Victor Chang. Fog-Cloud-IoT centric collaborative framework for machine learning-based situationaware traffic management in urban spaces. Computing, 106(4): 1193–1225, April 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01120-2.

Shoughi:2024:AEC

[3865] Armin Shoughi, Mohammad Bagher Dowlatshahi, Arefeh Amiri, Marjan Kuchaki Rafsanjani, and Ranbir Singh Batth. Automatic ECG classification using discrete wavelet transform and one-dimensional convolutional neural network. Computing, 106(4):1227–1248, April 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01243-0.

Benatallah:2024:ISI

[3866] Boualem Benatallah, Hakim Hacid, Eleana Kafeza, Fabio Martinelli, and Athman Bouguettaya. IC-SOC 2020 special issue on serviceoriented computing. Computing, 106 (4):1249–1250, April 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01222-5.

Metzger:2024:RSA

[3867] Andreas Metzger, Clément Quinton, Zoltán Ádám Mann, Luciano Baresi, and Klaus Pohl. Realizing self-adaptive systems via online reinforcement learning and feature-model-guided exploration. Computing, 106 (4):1251–1272, April 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01052-x.

Rjoub:2024:TDR

[3868] Gaith Rjoub, Omar Abdel Wahab, Jamal Bentahar, and Ahmed Bataineh. Trust-driven reinforcement selection strategy for federated learning on IoT devices. Computing, 106(4): 1273–1295, April 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-022-01078-1.

Yousefi:2024:HEA

[3869] Malek Yousefi and Seyed Morteza Babamir. A hybrid energy-aware algorithm for virtual machine placement in cloud computing. Computing, 106 (5):1297–1320, May 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01280-3.

Battyanyi:2024:SSP

[3870] Péter Battyányi. Solving the SAT problem with the string multiset rewriting calculus. Computing, 106 (5):1321–1334, May 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01258-1.

Tong:2024:IVS

[3871] Tianchi Tong, Qian Dong, Wenying Yuan, and Jinsheng Sun. Identifying vital spreaders in complex networks based on the interpretative structure model and improved Kshell. Computing, 106(5): 1335–1358, May 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01268-z.

Huang:2024:EPA

[3872] Yuan-Ko Huang and Chien-Pang Lee. Efficient processing of all neighboring object group queries with budget range constraint in road networks. Computing, 106(5): 1359–1393, May 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01260-7.

Song:2024:IIB

[3873] Weida Song, Shanxin Zhang, Wenlong Ge, and Wei Wang. An improved indicator-based two-archive algorithm for many-objective optimization problems. Computing, 106(5): 1395–1429, May 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01272-3.

Wan:2024:RWE

[3874] Jian Wan and Luc Jaulin. Reducing the wrapping effect of set computation via Delaunay triangulation for guaranteed state estimation of nonlinear discrete-time systems. Computing, 106(5):1431–1449, May 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01275-0.

Paramasivam:2024:CLA

[3875] Suguna Paramasivam, R. Leela Velusamy, and J. V. Nishaanth. Categorical learning for automated network traffic categorization for future generation networks in SDN. Computing, 106(5):1451–1473, May 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01277-y.

Chen:2024:ESC

[3876] Lei Chen, Linyun Ma, and Lvjie Li. Enhancing sine cosine algorithm based on social learning and elite opposition-based learning. Computing, 106(5):1475–1517, May 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01256-3.

Riahi:2024:MBA

[3877] Ali Riahi, Abdorreza Savadi, and Mahmoud Naghibzadeh. Many-BSP: an analytical performance model for CUDA kernels. Computing, 106 (5):1519–1555, May 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01255-w.

Fang:2024:GTD

[3878] Xiu Susie Fang, Xianzhi Wang, Quan Z. Sheng, and Lina Yao. Generalizing truth discovery by incorporating multi-truth features. Computing, 106(5):1557–1583, May 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01288-9.

Swain:2024:EDD

[3879] Chinmaya Kumar Swain, Ravi Shankar, and Aryabartta Sahu. Edge data distribution as a network Steiner tree estimation in edge computing. Computing, 106(5):1585–1609, May 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01259-0.

Ekinci:2024:CSL

[3880] Ekin Ekinci. A comparative study of LSTM–ED architectures in forecasting day-ahead solar photovoltaic energy using weather data. Computing, 106(5):1611–1632, May 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01266-1.

Malukani:2024:PBM

[3881] Shaifali Malukani and C. K. Bhensdadia. Preference based multi-issue negotiation algorithm (PMINA) for fog resource allocation. Computing, 106 (5):1633–1657, May 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01271-4.

Xu:2024:SDC

[3882] Jiuyun Xu, Yinyue Jiang, Hanfei Fan, and Qiqi Wang. SVFLDetector: a decentralized client detection method for Byzantine problem in vertical federated learning. Computing, 106(5):1659–1679, May 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01262-5.

Yin:2024:PRI

[3883] Kangning Yin, Zhen Ding, Zhihua Dong, Xinhui Ji, Zhipei Wang, Dongsheng Chen, Ye Li, Guangqiang Yin, and Zhiguo Wang. Person re-identification method based on fine-grained feature fusion and selfattention mechanism. Computing, 106 (5):1681–1705, May 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01270-5.

Karyakarte:2024:DOR

[3884] Mandar Karyakarte, Aarti Agarkar, Lalit Kulkarni, Milind Patil, Gajanan Chavhan, and Bipin Sule. Dynamic opportunistic routing protocol for ad-hoc Internet of Vehicles (IoV). Computing, 106(6):1707–1728, June 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01248-9.

Sun:2024:LLG

[3885] Chengcheng Sun, Fanrong Meng, Chenhao Li, Xiaobin Rui, and Zhixiao Wang. LGAT: a light graph attention network focusing on message passing for semi-supervised node classification. Computing, 106(6): 1729–1747, June 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01261-6.

Wang:2024:ECA

[3886] Shuang Wang, Yibing Duan, Yamin Lei, Peng Du, and Yamin Wang. Electricity-cost-aware multi-workflow scheduling in heterogeneous cloud. Computing, 106(6):1749–1775, June 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-024-01264-
3.

Akraminejad:2024:MOC

[3887] Reza Akraminejad, Navid Khaledian, Amin Nazari, and Marcus Voelp. A multi-objective crow search algorithm for optimizing makespan and costs in scientific cloud workflows (CSAMOMC). Computing, 106(6): 1777–1793, June 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01263-4.

Naeen:2024:CDC

[3888] Mohammad Ali Monshizadeh Naeen, Hamid Reza Ghaffari, and Hossein Monshizadeh Naeen. Cloud data center cost management using virtual machine consolidation with an improved artificial feeding birds algorithm. Computing, 106 (6):1795–1823, June 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01267-0.

Mokhtarzadeh:2024:TSP

[3889] Salman Mokhtarzadeh, Behzad Zamani Dehkordi, Mohammad Mosleh, and Ali Barati. Time-sensitive propagation values discount centrality measure. Computing, 106(6): 1825–1843, June 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01265-2.

Khan:2024:ETM

[3890] Zahoor Ali Khan, Muhammad Awais, Turki Ali Alghamdi, and Nadeem Javaid. Employing topology modification strategies in scale-free IoT networks for robustness optimization. Computing, 106(6):1845–1872, June 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/

article/10.1007/s00607-024-01273-
2.

Fahim:2024:EIF

[3891] Tamer E. Fahim, Sherif I. Rabia, Ahmed H. Abd El-Malek, and Waheed K. Zahra. Enhancing information freshness in multi-class mobile edge computing systems using a hybrid discipline. Computing, 106 (6):1873–1911, June 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01278-x.

Xu:2024:IMM

[3892] Zhenyu Xu, Xinxin Zhang, Mingzhi Chen, and Li Xu. Influence maximization in mobile social networks based on RWP–CELF. Computing, 106 (6):1913–1931, June 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01276-z.

Dhanaraj:2024:MMO

[3893] Rajesh Kumar Dhanaraj, Anamika Singh, and Anand Nayyar. Matyas-Meyer Oseas based device profiling for anomaly detection via deep reinforcement learning (MMODPAD-DRL) in zero trust security network. Computing, 106(6):1933–1962, June 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-024-01269-
y.

Jiang:2024:UHS

[3894] Xin Jiang, Hongbo Liu, Liping Yang, Bo Zhang, Tomas E. Ward, and Václav Snášel. Unraveling human social behavior motivations via inverse reinforcement learning-based link prediction. Computing, 106 (6):1963–1986, June 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01279-w.

Lin:2024:CAD

[3895] Yonghui Lin, Li Xu, Wei Lin, and Jiayin Li. Community anomaly detection in attribute networks based on refining context. Computing, 106 (6):1987–2006, June 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01284-z.

Li:2024:OEA

[3896] Cen Li and Liping Chen. Optimization for energy-aware design of task scheduling in heterogeneous distributed systems: a meta-heuristic based approach. Computing, 106 (6):2007–2031, June 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01282-1.

Jawabreh:2024:QBW

[3897] Ezdehar Jawabreh and Adel Taweel. Qos-based web service selection using time-aware collaborative filtering: a literature review. Computing, 106 (6):2033–2058, June 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01283-0.

Pinto:2024:ORS

[3898] Braulio Pinto and Horacio Oliveira.

Online RSSI selection strategy for indoor positioning in low-effort training scenarios. Computing, 106(6): 2059–2077, June 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01285-y.

Sheng:2024:SIL

[3899] Michael Sheng, Taotao Cai, and Adnan Mahmood. Special issue on link prediction in complex networks. Computing, 106(7):2079, July 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01298-7.

Ding:2024:MEQ

[3900] Linlin Ding, Sisi Li, Ji Ma, and Mo Li.

Multi-example query over ontologylabel knowledge graphs. Computing, 106(7):2081–2106, July 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01194-6.

Faraji:2024:SRA

[3901] Fatimah Faraji, Amir Javadpour, Arun Kumar Sangaiah, and Hadi Zavieh. A solution for resource allocation through complex systems in fog computing for the Internet of Things. Computing, 106 (7):2107–2131, July 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01199-1.

Morales-Garzon:2024:LPF

[3902] Andrea Morales-Garzón, Karel Gutiérrez-Batista, and Maria J. Martin-Bautista. Link prediction in food heterogeneous graphs for personalised recipe recommendation based on user interactions and dietary restrictions. Computing, 106(7):2133–2155, July 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-023-01233-2.

Lakshmi:2024:LPA

[3903] T. Jaya Lakshmi and S. Durga Bhavani. Link prediction approach to recommender systems. Computing, 106 (7):2157–2183, July 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-023-01227-0.

Kumar:2024:ERL

[3904] Vishnu Kumar and P. Radha Krishna. An effective representation learning model for link prediction in heterogeneous information networks. Computing, 106(7): 2185–2210, July 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-023-01238-x.

Zeng:2024:ISA

[3905] Zhuo Zeng, Chengliang Wang, Fei Ma, Xusheng Li, and Xinrun Chen. Incorporating self-attentions into robust spatial-temporal graph representation learning against dynamic graph perturbations. Computing, 106 (7):2211–2237, July 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01235-0.

Kammari:2024:CRU

[3906] Monachary Kammari and S. Durga Bhavani. Citation recommendation using modified HITS algorithm. Computing, 106(7):2239–2259, July 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-023-01213-
6.

Hu:2024:TPM

[3907] Qiang Hu, Haoquan Qi, Yanzhe Jia, and Lianen Qu. A two-phase method to optimize service composition in cloud manufacturing. Computing, 106 (7):2261–2291, July 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01286-x.

Ding:2024:IOF

[3908] Chen Ding and GuangYu Zhu. Improved optimal foraging algorithm for global optimization. Computing, 106 (7):2293–2319, July 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01290-1.

Fakhouri:2024:FVI

[3909] Hussam N. Fakhouri, Feras M.

Awaysheh, Sadi Alawadi, Mohannad Alkhalaileh, and Faten Hamad. Four vector intelligent metaheuristic for data optimization. Computing, 106 (7):2321–2359, July 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01287-w.

Gu:2024:SRD

[3910] Haijun Gu, Yingyu Ma, Siqi Wang, Xincheng Chen, and Weihua Su. Semantically realizing discovery and composition for RESTful web services. Computing, 106(7):2361–2387, July 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-024-01289-
8.

Ahmad:2024:EIC

[3911] Amro Al-Said Ahmad, Lamis F. Al-Qora’n, and Ahmad Zayed. Exploring the impact of chaos engineering with various user loads on cloud native applications: an exploratory empirical study. Computing, 106 (7):2389–2425, July 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01292-z.

Bojjagani:2024:SPE

[3912] Sriramulu Bojjagani, Denslin Brabin, Kalai Kumar, Neeraj Kumar Sharma, and Umamaheswararao Batta. Secure privacy-enhanced fast authentication and key management for IoMT-enabled smart healthcare systems. Computing, 106(7):2427–2458, July 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01291-0.

Cheng:2024:DPF

[3913] Shuyan Cheng, Peng Li, Ruchuan Wang, and He Xu. Differentially private federated learning with non-IID data. Computing, 106(7): 2459–2488, July 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01257-2.

Chen:2024:ACP

[3914] Jiaxing Chen, Xiaoqian Liu, Chao Wu, Jiahui Ma, Zhiyuan Cui, and Zhihua Liu. An ant colony path planning optimization based on oppositionbased learning for AUV in irregular regions. Computing, 106(7): 2489–2520, July 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01293-y.

Cai:2024:CDO

[3915] Lu Cai, Yongcai Yan, Zhongming Tang, and Aijun Liu. Collaborative distribution optimization model and algorithm for an intelligent supply chain based on green computing energy management. Computing, 106 (8):2521–2539, August 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00972-4.

Teekaraman:2024:IEM

[3916] Yuvaraja Teekaraman, Ramya Kuppusamy, and V. Indragandhi. Investigations on the effect of micro-grid using improved NFIS–PID with hybrid algorithms. Computing, 106(8):

2541–2559, August 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-01006-9.

Papathanasiou:2024:DMS

[3917] Dimitrios Papathanasiou and Kostas Kolomvatsos. Data management and selectivity in collaborative pervasive edge computing. Computing, 106(8): 2561–2584, August 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01297-8.

Hadj-Kacem:2024:MLL

[3918] Mouna Hadj-Kacem and Nadia Bouassida. Multi-label learning for identifying co-occurring class code smells. Computing, 106(8):2585– 2612, August 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01294-x.

vidhya:2024:NBL

[3919] G. Sri vidhya and R. Nagarajan. A novel bidirectional LSTM model for network intrusion detection in SDN– IoT network. Computing, 106(8): 2613–2642, August 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01295-w.

Nisa:2024:ARM

[3920] Habib Un Nisa, Saif Ur Rehman Khan, Shahid Hussain, and Wen-Li Wang. An association rule miningoriented approach for prioritizing functional requirements. Computing, 106 (8):2643–2674, August 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01296-9.

Plana:2024:QFF

[3921] Francisco Plana, Andrés Abeliuk, and Jorge Pérez. Quickcent: a fast and frugal heuristic for harmonic centrality estimation on scalefree networks. Computing, 106(8): 2675–2705, August 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01303-z.

Zhao:2024:IST

[3922] Xiaofei Zhao, Kam-Yiu Lam, and Tei-Wei Kuo. Indexing spatiotemporal trajectory data streams on keyvalue storage. Computing, 106(8): 2707–2735, August 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01304-y.

Fe:2024:EAD

[3923] Iure Fé, Tuan Anh Nguyen, Mario Di Mauro, Fabio Postiglione, Alex Ramos, André Soares, Eunmi Choi, Dugki Min, Jae Woo Lee, and Francisco Airton Silva. Energy-aware dynamic response and efficient consolidation strategies for disaster survivability of cloud microservices architecture. Computing, 106 (8):2737–2783, August 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01305-x.

Herculano:2024:MAC

[3924] Jorge Herculano, Willians Pereira, Marcelo Guimarães, Reinaldo Cotrim, Alirio de Sá, Flávio Assis, Raimundo Macêdo, and Sérgio Gorender. MAC approaches to communication efficiency and reliability under dynamic network traffic in wireless body area networks: a review. Computing, 106 (8):2785–2809, August 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01307-9.

Aslan:2024:BFB

[3925] Selcuk Aslan. Back-and-Forth (BaF): a new greedy algorithm for geometric path planning of unmanned aerial vehicles. Computing, 106(8): 2811–2849, August 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01309-7.

Niaz:2024:AEN

[3926] Fahim Niaz, Jian Zhang, Muhammad Khalid, Kashif Naseer Qureshi, Yang Zheng, Muhammad Younas, and Naveed Imran. AI enabled: a novel IoT-based fake currency detection using millimeter wave (mmWave) sensor. Computing, 106(8):2851– 2873, August 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01300-2.

Hou:2024:PHB

[3927] Jiangang Hou, Xin Li, Hongji Xu, Chun Wang, Lizhen Cui, Zhi Liu, and Changzhen Hu. Packet headerbased reweight-long short term memory (Rew-LSTM) method for encrypted network traffic classification. Computing, 106(8):2875– 2896, August 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01306-w.

Ghasemi:2024:EVM

[3928] Arezoo Ghasemi, Abolfazl Toroghi Haghighat, and Amin Keshavarzi. Enhancing virtual machine placement efficiency in cloud data centers: a hybrid approach using multi-objective reinforcement learning and clustering strategies. Computing, 106(9): 2897–2922, September 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01311-z.

Shudrenko:2024:MEE

[3929] Yevhenii Shudrenko and Andreas Timm-Giel. Modeling end-to-end delays in TSCH wireless sensor networks using queuing theory and combinatorics. Computing, 106(9):2923– 2947, September 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01313-x.

Javaheri:2024:ECR

[3930] Atiyeh Javaheri, Ali Bohlooli, and Kamal Jamshidi. Enhancing computation reuse efficiency in ICN-based edge computing by modifying content store table structure. Computing, 106(9): 2949–2969, September 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01312-y.

Haddouti:2024:SCA

[3931] Samia El Haddouti, Mohammed Khaldoune, Meryeme Ayache, and Mohamed Dafir Ech-Cherif El Kettani. Smart contracts auditing and multiclassification using machine learning algorithms: an efficient vulnerability detection in Ethereum blockchain. Computing, 106(9):2971– 3003, September 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01314-w.

Razavi:2024:CFA

[3932] Seyyed Javad Bozorg Zadeh Razavi, Haleh Amintoosi, and Mohammad Allahbakhsh. A clarity and fairness aware framework for selecting workers in competitive crowdsourcing tasks. Computing, 106(9):3005– 3030, September 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01316-8.

Haris:2024:OPC

[3933] Raseena M. Haris, Mahmoud Barhamgi, Armstrong Nhlabatsi, and Khaled M. Khan. Optimizing pre-copy live virtual machine migration in cloud computing using machine learning-based prediction model. Computing, 106(9): 3031–3062, September 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01318-6. See cor-

rection [3980].

Zhang:2024:DLB

[3934] Yifu Zhang, Qian Sun, Ji Chen, and Huini Zhou. Deep learningbased classification and application test of multiple crop leaf diseases using transfer learning and the attention mechanism. Computing, 106(9): 3063–3084, September 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01308-8.

Rama-Maneiro:2024:ERG

[3935] Efrén Rama-Maneiro, Juan C. Vidal, Manuel Lama, and Pablo Monteagudo-Lago. Exploiting recurrent graph neural networks for suffix prediction in predictive monitoring. Computing, 106(9): 3085–3111, September 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01315-9.

Sachin:2024:FLD

[3936] D. N. Sachin, B. Annappa, and Sateesh Ambesange. Federated learning for digital healthcare: concepts, applications, frameworks, and challenges. Computing, 106(9):3113– 3150, September 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01317-7.

Paul:2024:DRI

[3937] Agyemang Paul, Yuxuan Wan, Zhefu Wu, Boyu Chen, and Shufeng Gong. Deep recommendation with iteration directional adversarial training. Computing, 106(10):3151– 3174, October 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01326-6.

Kumar:2024:EHT

[3938] Anumukonda Naga Seshu Kumar, Rajesh Kumar Yadav, and Nallanthighal Srinivasa Raghava. Efficient hashing technique for malicious profile detection at hypervisor environment. Computing, 106(10): 3175–3193, October 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01325-7.

Kartli:2024:HAO

[3939] Nermin Kartli, Erkan Bostanci, and Mehmet Serdar Guzel. Heuristic algorithm for an optimal solution of fully fuzzy transportation problem. Computing, 106(10): 3195–3227, October 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01319-5.

Chen:2024:PBD

[3940] Yishan Chen, Xiansong Luo, Peng Liang, Junxiao Han, and Zhonghui Xu. Priority-based DAG task offloading and secondary resource allocation in IoT edge computing environments. Computing, 106(10): 3229–3254, October 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01327-5.

Yue:2024:DAG

[3941] Chunguang Yue, Jinbao Li, Qichen Wang, and Donghuan Zhang. Dynamic attention guider network. Computing, 106(10):3255–3277, October 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-024-01328-
4.

Singh:2024:CTE

[3942] Gyan Singh and Amit K. Chaturvedi.

A cost, time, energy-aware workflow scheduling using adaptive PSO algorithm in a cloud-fog environment. Computing, 106(10):3279– 3308, October 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01322-w.

Jena:2024:PPP

[3943] Riyanka Jena, Priyanka Singh, Manoranjan Mohanty, and Manik Lal Das. PP-PRNU: PRNU-based source camera attribution with privacy-preserving applications. Computing, 106(10): 3309–3333, October 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01330-w.

Li:2024:PPN

[3944] Jiahao Li, Tianhan Gao, and Qingwei Mi. Phasic parallel-network policy: a deep reinforcement learning framework based on action correlation. Computing, 106(10):3335– 3352, October 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01329-3.

Damera:2024:ISO

[3945] Vijay Kumar Damera, G. Vanitha, B. Indira, G. Sirisha, and Ramesh Vatambeti. Improved snake optimizationbased task scheduling in cloud computing. Computing, 106(10):3353– 3385, October 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01323-9.

Anonymous:2024:JNa

[3946] Anonymous. Journal navigation. Computing, 106(10):??, October 2024. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Soldani:2024:SOC

[3947] Jacopo Soldani, George A. Papadopoulos, and Florian Rademacher. Serviceoriented and cloud computing. Computing, 106(11):3387, November 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00607-024-01351-
                                        5.

Khan:2024:CST

[3948] Akif Quddus Khan, Mihhail Matskin, Radu Prodan, Christoph Bussler, Dumitru Roman, and Ahmet Soylu. Cloud storage tier optimization through storage object classification. Computing, 106(11):3389–3418, November 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-024-01281-
2.

Cruz-Filipe:2024:XEL

[3949] Luís Cruz-Filipe, Sofia Kostopoulou, Fabrizio Montesi, and Jonas Vistrup. μXL: explainable lead generation with microservices and hypothetical answers. Computing, 106(11):3419– 3445, November 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01321-x.

Mondal:2024:MOC

[3950] Brototi Mondal and Avishek Choudhury. Multi-objective cuckoo optimizer for task scheduling to balance workload in cloud computing. Computing, 106(11):3447– 3478, November 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01332-8.

Makris:2024:EDD

[3951] Antonios Makris, Evangelos Psomakelis, Ioannis Korontanis, Theodoros Theodoropoulos, Ioannis Kontopoulos, Maria Pateraki, Christos Diou, and Konstantinos Tserpes. Edgedriven Docker registry: facilitating XR application deployment. Computing, 106(11):3479–3501, November 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-024-01310-
0.

Zhang:2024:HPM

[3952] Lei Zhang, Ke Pang, and Jiangtao Xu. High-performance microservice differentiated domain communication technology. Computing, 106(11): 3503–3535, November 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01232-3.

Hashemi:2024:NAS

[3953] Sayed Mohsen Hashemi, Amir Sahafi, Amir Masoud Rahmani, and Mahdi Bohlouli. A new approach for service activation management in fog computing using Cat Swarm Optimization algorithm. Computing, 106(11): 3537–3572, November 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01302-0.

Delzer:2024:ASS

[3954] Ole Delzer, Richard Hobeck, Ingo Weber, Dominik Kaaser, Michael Sober, and Stefan Schulte. Analysis of strategies for scalable transaction creation in blockchains. Computing, 106(11): 3573–3602, November 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01324-8.

delaFuente-Anaya:2024:BPD

[3955] Hector A. de la Fuente-Anaya, Heidy M. Marin-Castro, Miguel Morales-Sandoval, and Jose Juan Garcia-Hernandez. Business process discovery as a service with event log privacy and access control over discovered models. Computing, 106(11): 3603–3625, November 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01299-6.

Alhussayen:2024:PEO

[3956] Asma Alhussayen, Kamal Jambi, Fathy Eassa, and Maher Khemakhem. Performance evaluation of an oraclebased interoperability for permissioned blockchain. Computing, 106(11): 3627–3655, November 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01337-3.

Yue:2024:VPN

[3957] Yi Yue, Shiding Sun, Xuebei Zhang, Ruihua Li, Wencong Yang, Chang Cao, and Xiongyan Tang. VNF placement in NFV-enabled networks: considering time-varying workloads and multitenancy with a throughput optimization heuristic. Computing, 106(11): 3657–3690, November 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01336-4.

Kavitha:2024:AVC

[3958] N. Kavitha, K. Ruba Soundar, R. Karthick, and J. Kohila. Automatic video captioning using tree hierarchical deep convolutional neural network and ASRNN-bi-directional LSTM. Computing, 106(11):3691– 3709, November 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01334-6. See cor-

rection [4091].

Haj:2024:CDM

[3959] Abir El Haj. Community detection in multiplex continous weighted nodes networks using an extension of the stochastic block model. Computing, 106(11):3711–3725, November 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00607-024-01341-
                                        7.

Moustakas:2024:DRA

[3960] Thanasis Moustakas, Athanasios Tziouvaras, and Kostas Kolomvatsos. Data and resource aware incremental ML training in support of pervasive applications. Computing, 106(11):3727– 3753, November 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01338-2.

Ghorbian:2024:SCS

[3961] Mohsen Ghorbian and Mostafa Ghobaei-Arani. A survey on the cold start latency approaches in serverless computing: an optimization-based perspective. Computing, 106(11):3755– 3809, November 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01335-5.

Subbaiah:2024:EAT

[3962] C. Venkata Subbaiah and K. Govinda.

Energy-aware and trust-based cluster head selection in healthcare WBANs with enhanced GWO optimization. Computing, 106(11):3811– 3836, November 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01339-1.

Anonymous:2024:JNb

[3963] Anonymous. Journal navigation. Computing, 106(11):??, November 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Krishnan:2024:RPR

[3964] Ragi Krishnan and Selvam Durairaj.

Reliability and performance of resource efficiency in dynamic optimization scheduling using multi-agent microservice cloud-fog on IoT applications. Computing, 106(12):3837– 3878, December 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01301-1.

Rodriguez:2024:LLM

[3965] David Rodriguez, Ian Yang, Jose M.

Del Alamo, and Norman Sadeh. Large language models: a new approach for privacy policy analysis at scale. Computing, 106(12):3879– 3903, December 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01331-9.

Ali:2024:HDL

[3966] Tassawar Ali, Hikmat Ullah Khan, Fawaz Khaled Alarfaj, and Mohammed AlReshoodi. Hybrid deep learning and evolutionary algorithms for accurate cloud workload prediction. Computing, 106(12):3905– 3944, December 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01340-8.

Mejia:2024:DFS

[3967] Santiago Mejía and Jose Aguilar. A demand forecasting system of product categories defined by their time series using a hybrid approach of ensemble learning with feature engineering. Computing, 106(12):3945– 3965, December 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01320-y.

Cha:2024:URF

[3968] Hyejin Cha, In Kee Kim, and Taeseok Kim. Using a random forest to predict quantized reuse distance in an SSD write buffer. Computing, 106(12): 3967–3986, December 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01343-5.

Kumar:2024:REG

[3969] Mandeep Kumar, Gagandeep Kaur, and Prashant Singh Rana. Robust evaluation of GPU compute instances for HPC and AI in the cloud: a TOPSIS approach with sensitivity, bootstrapping, and non-parametric analysis. Computing, 106(12):3987– 4014, December 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01342-6.

Liu:2024:OMD

[3970] Yingze Liu, Yuanbo Guo, Rajiv Ranjan, and Dan Chen. Optimization of mitigation deployment using deep reinforcement learning over an enhanced ATT&CK. Computing, 106(12):4015– 4038, December 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01344-4.

Sharma:2024:MOS

[3971] Shalini Sharma, Bhupendra Kumar Pathak, and Rajiv Kumar. Multiobjective service composition optimization problem in IoT for agriculture 4.0. Computing, 106(12): 4039–4056, December 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01346-2.

Sujana:2024:FIE

[3972] J. Angela Jennifa Sujana, R. Venitta Raj, and V. K. Raja Priya. Fog intelligence for energy efficient management in smart street lamps. Computing, 106(12):4057–4082, December 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-024-01348-
0.

Mahansaria:2024:CAU

[3973] Divyans Mahansaria and Uttam Kumar Roy. Contextual authentication of users and devices using machine learning. Computing, 106(12): 4083–4107, December 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01333-7.

Haider:2024:MJT

[3974] Syed Ali Haider, Junaid A. Zubairi, and Sahar Idwan. Mapping and just-in-time traffic congestion mitigation for emergency vehicles in smart cities. Computing, 106(12): 4109–4130, December 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01345-3.

Prithi:2024:FRC

[3975] M. Prithi and K. Tamizharasi. Facial recognition and classification for customer information systems: a feature fusion deep learning approach with FFDMLC algorithm. Computing, 106(12):4131– 4165, December 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01349-z.

Almalki:2024:SML

[3976] Faris A. Almalki and Marios C. Angelides. A synthesis of machine learning and internet of things in developing autonomous fleets of heterogeneous unmanned aerial vehicles for enhancing the regenerative farming cycle. Computing, 106(12): 4167–4192, December 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01347-1.

Kumar:2024:SCA

[3977] Ashutosh Kumar and Garima Verma.

Securing cloud access with enhanced attribute-based cryptography. Computing, 106(12):4193– 4207, December 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-023-01212-7.

Wu:2024:DRS

[3978] Feng Wu, Hongwei Lv, Tongrang Fan, Wenbin Zhao, and Jiaqi Wang. A data reuse strategy based on deep learning for high dimensional data’s pattern and instance similarity. Computing, 106(12):4209– 4229, December 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-021-00964-4.

Xiaofeng:2024:RSC

[3979] Weng Xiaofeng, Zhao Jinghua, Jiang Chenxi, and Ji Yiying. Research on sentiment classification of futures predictive texts based on BERT. Computing, 106(12):4231– 4248, December 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-021-00989-9.

Haris:2024:COP

[3980] Raseena M. Haris, Mahmoud Barhamgi, Armstrong Nhlabatsi, and Khaled M. Khan. Correction to: Optimizing pre-copy live virtual machine migration in cloud computing using machine learning-based prediction model. Computing, 106(12): 4249, December 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01350-6.  See

[3933].

Anonymous:2024:JNc

[3981] Anonymous. Journal navigation. Computing, 106(12):??, December 2024. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hamed:2025:DAA

[3982] Suhaib Kh. Hamed, Mohd Juzaiddin Ab Aziz, and Mohd Ridzwan Yaakub. A data augmentation approach based on various GAN models to address class imbalance in fine-grained multimodal fake news datasets. Computing, 107 (1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-025-01413-2.

Havas:2025:SMS

[3983] Shahla Havas, Sadoon Azizi, and Alireza Abdollahpouri. SG-MOACO: a semi-greedy multi-objective ACO method for edge server placement in mobile edge computing. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01400-z.

Du:2025:FYF

[3984] Yan Du, Teng Wu, Zifeng Dai, Hui Xie, Changzhen Hu, and Shengjun Wei. F-yolov7: fast and robust realtime UAV detection. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01406-7.

Zhang:2025:TFP

[3985] Hong Zhang, Gang Yang, Hailiang Yu, and Zan Zheng. Traffic flow prediction based on the RETGCN model. Computing, 107 (1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01402-x.

Tang:2025:EMS

[3986] Zhijie Tang and Ke Meng. Efficient multi-scale 3D point cloud registration. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-024-01401-
y.

Aslani:2025:MLI

[3987] Akram Aslani and Mostafa Ghobaei-Arani. Machine learning inference serving models in serverless computing: a survey. Computing, 107 (1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01377-9.

Li:2025:TMS

[3988] Min Li, Zhiying Tu, Hanchuan Xu, Xiaofei Xu, and Zhongjie Wang. Towards multi-stakeholder value/quality conflict detection, measurement and resolution in composite services. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01408-5.

Zhou:2025:BDT

[3989] Qiao Zhou, Xingpeng Lin, Li Xu, and Yan Sun. Bi-directional temporal graph attention networks for rumor detection in online social networks. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00607-024-01395-
                                        7.

Jiang:2025:CIR

[3990] Shuhai Jiang, Xunan Cao, Cun Li, Kangqian Zhou, and Ming Hu. Courier information recognition based on an improved YOLOv8 visual servoing system. Computing, 107 (1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01411-w.

Rahmani:2025:SSP

[3991] Somayeh Rahmani, Vahid Khajehvand, and Mohsen Torabian. SPP: stochastic process-based placement for VM consolidation in cloud environments. Computing, 107 (1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01412-9.

Azad:2025:NEL

[3992] Mohammad Azad, Tasnemul Hasan Nehal, and Mikhail Moshkov. A novel ensemble learning method using majority based voting of multiple selective decision trees. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01394-8.

Larian:2025:IIT

[3993] Habib Larian and Faramarz Safi-Esfahani. InTec: integrated thingsedge computing: a framework for distributing machine learning pipelines in edge AI systems. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01388-6.

Khanna:2025:GAF

[3994] Abhinav Khanna, Gandikota Anjali, Nilesh Kumar Verma, and K. Jairam Naik. A GRL-aided federated graph reinforcement learning approach for enhanced file caching in mobile edge computing. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01396-6.

Jiang:2025:TMW

[3995] Ying Jiang, Yifa Huang, Zenghui Li, Ziran Zuo, and Chunhong Liu. TimesLap: Mutability workload sequence prediction based on Laplacian Kernel in the cloud. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01407-6.

Li:2025:IEN

[3996] Hao Li, Zhaoning Tian, Xiaohui Li, and Peyman Arebi. Identifying effective nodes in term of the influence maximization on the social IoT networks using deep learningbased node embedding. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01376-w.

Pepsi:2025:IBH

[3997] M. Blessa Binolin Pepsi, N. Senthil Kumar, S. Jeyashree, and M. Subitcha. IoT based human activity recognition on drifted data stream using arbitrary width convolution neural network. Computing, 107 (1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01392-w.

Lin:2025:CSB

[3998] Junjie Lin, Xiaogang Wang, Mingrui Chang, Zhiwei Yin, and Liuhong Zhang. A cosine similaritybased maximal clique point cloud registration algorithm. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01403-w.

Qiqieh:2025:DCB

[3999] Issa Qiqieh, Jafar Alzubi, and Omar Alzubi. DNA cryptography based security framework for health-cloud data. Computing, 107 (1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01393-9.

Ghorbani:2025:AAL

[4000] Mehdi Ghorbani, Navid Khaledian, and Setareh Moazzami. ALBLA: an adaptive load balancing approach in edge-cloud networks utilizing learning automata. Computing, 107 (1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01380-0.

Yousefinejad:2025:RVE

[4001] Fatemeh Yousefinejad and Saeed Jalili.

Runtime verification of electronic stability control system in automated vehicles with STL_{3} formalism. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01375-x.

Xie:2025:BBE

[4002] Hui Xie, Teng He, Shengjun Wei, and Changzhen Hu. Blockchain-based entity access control scheme for ubiquitous UAV swarm tasks. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01381-z.

Liu:2025:OCD

[4003] Zhijing Liu. Optimized cooperative data transmission in two-tier NB–IoT networks. Computing, 107 (1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01389-5.

Assayag:2025:EEI

[4004] Yuri Assayag, Eduardo Souto, Raimundo Barreto, Moises Carvalho, Richard Pazzi, and Horácio Fernandes. Efficient exploration of indoor localization using genetic algorithm and signal propagation model. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01391-x.

Nemati:2025:EFS

[4005] Khadijeh Nemati, Amir Hosein Refahi Sheikhani, Sohrab Kordrostami, and Kamrad Khoshhal Roudposhti. The embedded feature selection method using ANT colony optimization with structured sparsity norms. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01387-7.

Ali:2025:TPC

[4006] Tariq Emad Ali, Faten Imad Ali, Pavle Dakić, and Alwahab Dhulfiqar Zoltan. Trends, prospects, challenges, and security in the healthcare Internet of Things. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01352-4.

Hu:2025:MCS

[4007] Qingmiao Hu, Hongwei Li, and Elaheh Nikougoftar. Mitigating cold start problem in serverless computing using predictive pre-warming with machine learning. Computing, 107 (1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01382-y.

Wan:2025:IES

[4008] Yan Wan, Yujia Zhai, Can Cui, and Dexuan Song. Indoor energysaving strategy optimization based on deep reinforcement learning and DDPG algorithm. Computing, 107 (1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01378-8.

Wang:2025:LYL

[4009] Sicheng Wang, Gang Li, Bin He, Bin Cheng, Yulong Ding, and Wei Li. LUMF-YOLO: a lightweight object detection network integrating UAV motion features. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01379-7.

Mukhopadhyay:2025:CEA

[4010] Nirmalya Mukhopadhyay and Babul P. Tewari. Cost and energy aware migration through dependency analysis of VM components in virtual cloud infrastructure. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01383-x.

Zhang:2025:NCP

[4011] Lei Zhang, Mingzeng Cao, Jing Li, Chenglin Zhang, and Lili He. A novel collaborative privacy protection scheme based on verifiable secret sharing and trust mechanism. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01361-3.

Chen:2025:SEW

[4012] Long Chen, Guangrui Liu, Jinquan Zhang, and Xiaodong Zhang. Scheduling ensemble workflows on hybrid resources in IaaS clouds. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01386-8.

Yao:2025:EPR

[4013] Yuan Yao, Hui Zhan, Ali Noorian, and Mehdi Hazratifard. Enhancing POI recommendations on social media: a sequential approach incorporating LSTM and user feedback. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01385-9.

Zafar:2025:ECD

[4014] Muhammad Mohsin Zafar, Nadeem Javaid, Ifra Shaheen, Nabil Alrajeh, and Sheraz Aslam. Enhancing clinical decision support with explainable deep learning framework for C-section forecasting. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01354-2.

Filho:2025:SMS

[4015] Cleunio França Filho and Eduardo Tavares. Stochastic modeling of smart street lighting systems: maintenance models assessment. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01359-x.

Chen:2025:BDI

[4016] Zhen Chen, Jia Huang, Shengzheng Liu, and Haixia Long. A blockchain and A-DCNN integrated framework for privacy protection and intrusion detection of industrial IoT. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01390-y.

Ganji:2025:MVR

[4017] Binazir Ganji, Ali Rezaee, Sahar Adabi, and Ali Movaghar. Model verification of real-time and distributed stream processing architecture. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01384-w.

Wang:2025:PAS

[4018] Zhongmin Wang, Hanchen Dong, Xiaomin Jin, and Yanping Chen. A profit-aware server deployment approach for edge computing. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01357-z.

Bushehrian:2025:DRL

[4019] Omid Bushehrian and Amir Moazeni. Deep reinforcement learning-based optimal deployment of IoT machine learning jobs in fog computing architecture. Computing, 107 (1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01353-3.

Singhal:2025:BTM

[4020] Ankur Kumar Singhal, Shriansh Manhas, and Anurag Singh. Beyond the traditional models: a network reconstruction based model for predicting and analysing individual health status. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01355-1.

Soltani:2025:OSR

[4021] Mohammad Amin Zare Soltani, Seyed Amin Hosseini Seno, and AmirHossein Mohajerzadeh. Optimizing SDN resource allocation using fuzzy logic and VM mapping technique. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01360-4.

Deng:2025:NON

[4022] Miaolei Deng, Umer Nauman, and Yuhong Zhang. NS-OWACC: natureinspired strategies for optimizing workload allocation in cloud computing. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-024-01367-
x.

Silva:2025:MAE

[4023] Alison Silva and Gustavo Callou. Models for availability evaluation of file servers in private clouds. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01372-0.

Kumar:2025:EKF

[4024] M. R. Naveen Kumar, B. Annappa, and Vishwas Yadav. Efficient Kalman filter based deep learning approaches for workload prediction in cloud and edge environments. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01373-z.

Benjamin:2025:ECI

[4025] Jetty Benjamin and Juby Mathew. Enhancing continuous integration predictions: a hybrid LSTM–GRU deep learning framework with evolved DBSO algorithm. Computing, 107 (1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01370-2.

Issa:2025:VVI

[4026] Wael Issa, Nour Moustafa, Benjamin Turnbull, Nasrin Sohrabi, Zahir Tari, and Albert Y. Zomaya. VIFL: vulnerability identification using federated learning in the internet of things systems. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01365-z.

Deldari:2025:IBB

[4027] Arash Deldari and Alireza Holghinezhad. An IoT-based bagof-tasks scheduling framework for deadline-sensitive applications in fogcloud environment. Computing, 107 (1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01371-1.

Lai:2025:TIC

[4028] Shouliang Lai, Peiling Zhou, Xiyu Yi, and Cong Luo. Towards improving context-aware applications using link quality induction and self-calibrated ultra-wide band localization systems. Computing, 107 (1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01368-w.

Zuo:2025:MCT

[4029] Ziran Zuo, Yifa Huang, Zenghui Li, Ying Jiang, and Chunhong Liu. Mixed contrastive transfer learning for few-shot workload prediction in the cloud. Computing, 107(1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01366-y.

Siavashi:2025:GCD

[4030] Ahmad Siavashi and Mahmoud Momtazpour. GPU cluster dynamics: insights from Alibaba’s 2023 trace release. Computing, 107 (1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01369-9.

Rivas:2025:SEF

[4031] Jose M. Rivas, Carlos Fernandez-Basso, Roberto Morcillo-Jimenez, Juan Paños-Basterra, M. Dolores Ruiz, and Maria J. Martin-Bautista. Simulafed: an enhanced federated simulated environment for privacy and security in health. Computing, 107 (1):??, January 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01364-0.

Anonymous:2025:JNa

[4032] Anonymous. Journal navigation. Computing, 107(1):??, January 2025. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lin:2025:PDC

[4033] Kawuu W. Lin, Ya-Jun Zheng, Ju-Chin Chen, Wei-Chiang Wang, and Chao-Chun Chen. A parallel and distributed C4.5 algorithm in cloud computing environments. Computing, 107(2):??, February 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-025-01415-0.

Shi:2025:OSS

[4034] Lei Shi, Ying Ren, Jing Xu, Yibin Xie, Chen Fang, and Yuqi Fan. An optimal scheme for speeding the training process of the asynchronous federated learning based on model partition. Computing, 107 (2):??, February 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01418-x.

Malhotra:2025:HFS

[4035] Ruchika Malhotra and Vidushi. Hybrid feature selection module for improving performance of software vulnerability severity prediction model on textual dataset. Computing, 107 (2):??, February 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01425-y.

Sousa:2025:AEM

[4036] Francisco Monte Sousa and Gustavo Callou. Analysis of evolutionary multiobjective algorithms for data center electrical systems. Computing, 107(2):??, February 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01416-z.

Feitosa:2025:CPE

[4037] Leonel Feitosa, Vandirleya Barbosa, Arthur Sabino, Luiz Nelson Lima, Iure Fé, Luis Guilherme Silva, Gustavo Callou, Juliana Carvalho, Erico Leão, Tuan Anh Nguyen, Paulo Rego, and Francisco Airton Silva. A comprehensive performance evaluation of container migration strategies. Computing, 107(2):??, February 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.

                                        1007/s00607-025-01423-0. See cor-

rection [4090].

Matani:2025:ISB

[4038] Alemeh Matani, Amir Sahafi, and Ali Broumandnia. Improving scalability in blockchain systems using multilevel sharding based on heterogeneity of network nodes. Computing, 107(2):??, February 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-025-01414-1.

Cuevas:2025:NCO

[4039] Erik Cuevas, Oscar Barba, and Héctor Escobar. A novel cheetah optimizer hybrid approach based on opposition-based learning (OBL) and diversity metrics. Computing, 107 (2):??, February 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01397-5.

Si:2025:MIC

[4040] Xiaokai Si, Yubo Song, Xin Sun, Wen Wang, and Zhongyuan Qin. MARLFuzz: industrial control protocols fuzzing based on multi-agent reinforcement learning. Computing, 107(2):??, February 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-025-01421-2.

Mahesar:2025:ETS

[4041] Abdul Rasheed Mahesar, Xiaoping Li, and Dileep Kumar Sajnani. Enhancing task scheduling and QoS optimization in mobile edge computing via microservice-oriented container selection. Computing, 107 (2):??, February 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01410-x.

Rasheed:2025:QSB

[4042] Rashid Rasheed and Shahzad Sarwar. Quality-of-service based multipath routing for the internet of things. Computing, 107(2):??, February 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-024-01374-
y.

Liu:2025:FDF

[4043] Wei Liu, Kaige Li, Yurong Zheng, Wei She, and Zhao Tian. FedBPM: a decentralized federated meta-method for heterogeneous and complex image classification via multi-scale feature fusion. Computing, 107 (2):??, February 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01405-8.

Elbehiry:2025:IMR

[4044] Eman Adel Elbehiry, Ahmed Fares, Basem M. Elhalawany, and Heba A. TagElDein. Inter-mesh routing algorithms in LEO satellite constellations networks. Computing, 107 (2):??, February 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01409-4.

Hemdan:2025:HVG

[4045] Ezz El-Din Hemdan, Samar M. Zayed, Gamal Attiya, Ayman El-Sayed, and Amged Sayed. Hybrid voting-GA ensemble learning for multiclass fault detection in digital twindriven IIoT systems. Computing, 107(2):??, February 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01399-3.

Azizi:2025:HTE

[4046] Somayeh Azizi, Mohammadreza Soltanaghaei, and Hossein Ghaffarian. Hierarchical traffic engineering with PSO: a path to efficient congestion management in SDN. Computing, 107 (2):??, February 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01363-1.

Xu:2025:EYE

[4047] Xin Xu, Qi Li, Jie Pan, Xingzheng Lu, Hongwei Wei, Mingzheng Sun, and Haoze Zhang. ESOD-YOLO: an enhanced efficient small object detection framework for aerial images. Computing, 107(2):??, February 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-024-01398-4.

Jia:2025:CPF

[4048] Shuting Jia, Zhanglu Tan, and Chao Li. Carbon prices forecasting based on sliding time window and improved support vector regression. Computing, 107(2):??, February 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01404-9.

Anonymous:2025:JNb

[4049] Anonymous. Journal navigation. Computing, 107(2):??, February 2025. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Garcia-Diaz:2025:GHG

[4050] Jesús García-Díaz, José Alejandro Cornejo-Acosta, and Joel Antonio Trejo-Sánchez. A greedy heuristic for graph burning. Computing, 107(3):??, March 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01436-9.

Nie:2025:TES

[4051] Guolei Nie and Elaheh Rezvani. Towards an efficient scheduling strategy based on multi-objective optimization in fog environments. Computing, 107(3):??, March 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01448-5.

Rajasekar:2025:EIS

[4052] V. R. Rajasekar and S. Rajkumar. Evaluating the impact of selective forwarding attacks on RPLbased 6LoWPAN in IoT: a comprehensive simulation study. Computing, 107(3):??, March 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01440-z.

Shafi:2025:DML

[4053] Numan Shafi, Muhammad Abdullah, Waheed Iqbal, and Faisal Bukhari. DIMA: machine learning based dynamic infrastructure management for containerized applications. Computing, 107(3):??, March 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-025-01445-8.

Xie:2025:IAC

[4054] Yuancheng Xie, Zhaoxin Zhang, Haoyang Gao, Ning Li, and Jialu Li. Identifying and analysis the core structure of the Internet. Computing, 107(3):??, March 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-025-01419-w.

Noor:2025:OPP

[4055] Sumaiya Noor, Hamid Hussain Awan, Amber Sarwar Hashmi, Aamir Saeed, Salman Khan, and Salman A. AlQahtani. Optimizing performance of parallel computing platforms for largescale genome data analysis. Computing, 107(3):??, March 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-025-01441-y.

ALFahad:2025:NSU

[4056] Saleh ALFahad, Shameem Puthiya Parambath, Christos Anagnostopoulos, and Kostas Kolomvatsos. Node selection using adversarial expertbased multi-armed bandits in distributed computing. Computing, 107(3):??, March 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01443-w.

Liu:2025:RMP

[4057] Liming Liu, Wenxuan Xu, Yanling Shao, Wenyong Dong, Hairui Xu, Hewen Gao, Pingping Shan, and Junying Guo. A reliable multi-path cluster routing method for ground-air cooperative flying ad-hoc networks. Computing, 107(3):??, March 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01442-x.

Garzon:2025:CIG

[4058] Martha Garzón, Lindsay Álvarez-Pomar, and Sergio Rojas-Galeano. From collective intelligence to global optimisation: an agent-based model approach. Computing, 107(3):??, March 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https://link.

springer.com/article/10.1007/s00607-
025-01429-8.

Maas:2025:ICI

[4059] William Maas, Fábio Rossi, Marcelo Caggiani, Philippe Navaux, and Arthur Lorenzon. Investigating cloud instances to achieve optimal trade-offs between performance-cost efficiency. Computing, 107(3):??, March 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01444-9.

Kotteswari:2025:EEE

[4060] K. Kotteswari, Rajesh Kumar Dhanaraj, Balamurugan Balusamy, Anand Nayyar, and Anupam Kumar Sharma. EELB: an energy-efficient load balancing model for cloud environment using Markov decision process. Computing, 107(3):??, March 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-025-01439-6.

Zahed:2025:EFP

[4061] Atiya Zahed, Mostafa Ghobaei-Arani, and Leila Esmaeili. An efficient function placement approach in serverless edge computing. Computing, 107(3):??, March 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-025-01438-7.

Abo-Soliman:2025:AAL

[4062] Mohamed Abo-Soliman, Eman Shaaban, and Karim Emara. ALLC: autonomous lightweight distributed ledger constructor for securing IoT information. Computing, 107(3):??, March 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https://link.

                                        springer.com/article/10.1007/s00607-
                                        025-01424-z.

Mezni:2025:DCE

[4063] Haithem Mezni, Mokhtar Sellami, Hela Elmannai, and Reem Alkanhel. Daas composition: enhancing UAV delivery services via LSTMbased resource prediction and flight patterns mining. Computing, 107 (3):??, March 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01417-y.

Wang:2025:APR

[4064] Han Wang, Kevin Joji Mathews, Muhammed Golec, Sukhpal Singh Gill, and Steve Uhlig. Amazon-AICloud: proactive resource allocation using Amazon Chronos based time series model for sustainable cloud computing. Computing, 107 (3):??, March 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01435-w.

Sohrabi:2025:FBA

[4065] Sakine Sohrabi, Mehdi Sakhaei-nia, Mohamad Nassiri, and Reza Mohammadi. Fog-based architecture and efficient task offloading methodology in IoT-based applications for smart irrigation system. Computing, 107(3):??, March 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01430-1.

Liu:2025:OMR

[4066] Xi Liu, Jun Liu, Hong Wu, and Jing Dong. An online mechanism for resource provisioning and allocation in vehicle computing. Computing, 107(3):??, March 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01428-9.

Wu:2025:TOB

[4067] Da Wu, Zhuo Li, Yongtao Ma, Kaihua Liu, and Peng Luo. Task offloading based on two types of edge–edge collaboration in mobile edge computing. Computing, 107 (3):??, March 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-025-01420-3.

Malathy:2025:GFL

[4068] N. Malathy, A. Lavanya, S. Pooja Shree, and R. Kumaripriya. Green federated learning empowered drug manufacturing mechanism for the pharmaceutical industry. Computing, 107(3):??, March 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-025-01433-y.

Kumar:2025:HNN

[4069] Jitendra Kumar, Deepika Saxena, Jatinder Kumar, Abhilash Bhadoria, A. Anjali, and Ashutosh Kumar Singh. A hybrid neural network and cooperative PSO model for dynamic cloud workloads prediction. Computing, 107(3):??, March 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00607-025-01434-x.

Katare:2025:AVT

[4070] Dewant Katare, Sam Leroux, Marijn Janssen, and Aaron Yi Ding. Approximating vision transformers for edge: variational inference and mixed-precision for multi-modal data.

Computing, 107(3):??, March 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-025-01427-
w.

Velayatipour:2025:TSQ

[4071] Masoumeh Velayatipour, Mohammad Mosleh, Mohsen Yoosefi Nejad, and Mohammad Kheyrandish. Towards secure quantum communication: a novel quantum audio watermarking based on bipolar echo hiding. Computing, 107(3):??, March 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01422-1.

Kumar:2025:MTB

[4072] Bablu Kumar, Anshul Verma, and Pradeepika Verma. A multivariate transformer-based monitor-analyze-plan-execute (MAPE) autoscaling framework for dynamic resource allocation in cloud environment. Computing, 107(3):??, March 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01426-x.

Anonymous:2025:JNc

[4073] Anonymous. Journal navigation. Computing, 107(3):??, March 2025. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Mahjoub:2025:QTS

[4074] Alireza Mahjoub, Madjid Khalilian, and Javad Mohammadzadeh. QQLAOA: task scheduling with multiobjectives quantum mutation and Qlearning based arithmetic optimizer algorithm in cloud data centers. Computing, 107(4):??, April 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01461-8.

Wang:2025:EDS

[4075] Yanting Wang, Yilin Yuan, Zhaoqian Zhang, and Zichen Li. An efficient data sharing and public auditing scheme for digital content via consortium chain. Computing, 107(4):??, April 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01458-3.

Herrera:2025:MOM

[4076] Juan Luis Herrera, Domenico Scotece, Jaime Galán-Jiménez, Javier Berrocal, Giuseppe Di Modica, and Luca Foschini. MUMIPLOP: optimal multi-core IoT service placement in fog environments. Computing, 107(4):??, April 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01460-9.

Wu:2025:VRA

[4077] Tao Wu, Yang Ou, Qizhao Zhou, and Xi Chen. A V2V2I routing algorithm for Internet of Vehicles based on adaptive variable beacon power. Computing, 107(4):??, April 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01452-9.

Fountas:2025:PMV

[4078] Panagiotis Fountas and Kostas Kolomvatsos. Proactive missing values imputation based on reinforcement learning. Computing, 107(4):??, April 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/
article/10.1007/s00607-025-01462-
7.

Baigi:2025:OSM

[4079] Ahmad Shokrani Baigi, Abdorreza Savadi, and Mahmoud Naghibzadeh. Optimizing sparse matrix partitioning in a heterogeneous CPU–GPU system for high-performance. Computing, 107(4):??, April 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01456-5.

Gonzalez:2025:IMO

[4080] Alexandra González, Joel Castaño, Xavier Franch, and Silverio Martínez-Fernández. Impact of ML optimization tactics on greener pretrained ML models. Computing, 107(4):??, April 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01437-8.

Li:2025:EMO

[4081] Xiaozhou Li, Dario Amoroso d’Aragona, Tomas Cerny, Valentina Lenarduzzi, Davide Taibi, and Andrea Janes. Exploring microservice ownership and organizational coupling in open-source projects: an empirical study. Computing, 107(4):??, April 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01454-7.

Sedaghat:2025:MLA

[4082] Shahrzad Sedaghat, Mahdi Nazernejad, and Mohammad Sadegh Sanie Jahromi. A machine learning approach for improved resource allocation in COVID-19 ICUs using HRCT scans and clinical data. Computing, 107(4):??, April 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01453-8.

Moreschini:2025:ATM

[4083] Sergio Moreschini, Shahrzad Pour, Ivan Lanese, Daniel Balouek, Justus Bogner, Xiaozhou Li, Fabiano Pecorelli, Jacopo Soldani, Eddy Truyen, and Davide Taibi. AI techniques in the microservices life–cycle: a systematic mapping study. Computing, 107(4):??, April 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01432-z.

Kinanen:2025:TFI

[4084] Otso Kinanen, Andrés D. Muñoz-Moller, Vlad Stirbu, Juan M. Murillo, and Tommi Mikkonen. Toolchain for faster iterations in quantum software development. Computing, 107(4):??, April 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01450-x.

Fedila:2025:NDP

[4085] Djahida Ali Fedila, Mohamed Ould Khaoua, and Ahmed Al-Dubai. A new dynamic probabilistic broadcast in IPv6-based low-power and lossy networks for the internet of things. Computing, 107(4):??, April 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01457-4.

Naghibi:2025:DTB

[4086] Maryam Naghibi, Hamid Barati, and Ali Barati. A dynamic trust based clustering method for secure data gathering in Internet of Things. Computing, 107(4):??, April 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01455-6.

Hellal:2025:FAS

[4087] Yassmine Gara Hellal, Lazhar Hamel, and Mohamed Graiet. A formal approach for scalable applications in dynamic and constrained IoT–Cloud systems. Computing, 107(4):??, April 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-024-01358-y.

Xu:2025:RTP

[4088] Chenghong Xu, Wenjie Qi, Zhigang Zhang, Shuaishuai Ge, Xiangkui Zhang, Guojian Wang, and Fei Peng. Rapid trajectory planning under multiple constraints in unknown environments. Computing, 107(4):??, April 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01431-0.

Zolghadri:2025:DTO

[4089] Mohammad Zolghadri, Parvaneh Asghari, Seyed Ebrahim Dashti, and Alireza Hedayati. Dynamic task offloading for IoT–Fog–Cloud systems: a network traffic-aware decision tree approach. Computing, 107(4):??, April 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01449-4.

Feitosa:2025:CCP

[4090] Leonel Feitosa, Vandirleya Barbosa, Arthur Sabino, Luiz Nelson Lima, Iure Fé, Luis Guilherme Silva, Gustavo Callou, Juliana Carvalho, Erico Leão, Tuan Anh Nguyen, Paulo Rego, and Francisco Airton Silva. Correction to: A comprehensive performance evaluation of container migration strategies. Computing, 107(4):??, April 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01447-6.  See

[4037].

Kavitha:2025:CAV

[4091] N. Kavitha, K. Ruba Soundar, R. Karthick, and J. Kohila. Correction: Automatic video captioning using tree hierarchical deep convolutional neural network and ASRNNbi-directional LSTM. Computing, 107(4):??, April 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01446-7.  See

[3958].

Anonymous:2025:JNd

[4092] Anonymous. Journal navigation. Computing, 107(4):??, April 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fernandez-Osuna:2025:EDP

[4093] Miriam Fernández-Osuna, Ricardo Pérez-Castillo, José A. Cruz-Lemus, Michal Baczyk, and Mario Piattini. Exploring design patterns in quantum software: a case study. Computing, 107(5):??, May 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01467-2.

Desdentado:2025:QCS

[4094] Elena Desdentado, Coral Calero, M. Ángeles Moraga, and Félix García. Quantum computing software solutions, technologies, evaluation and limitations: a systematic mapping study. Computing, 107(5):??, May 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https:

//link.springer.com/article/10.
1007/s00607-025-01459-2.

Anonymous:2025:JNe

[4095] Anonymous. Journal navigation. Computing, 107(5):??, May 2025. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Albrecht:1977:GTG

[4096] R. Albrecht. Grundlagen einer Theorie gerundeter algebraischer Verknüpfungen in topologischen Vereinen. (German) [Foundations of a theory of rounded algebraic knotting in topological association]. In und Ulrich Kulisch [4366], pages 1–14. CODEN COSPDM. ISBN 0-387-81410-8. ISSN 0344-8029. LCCN QA 162 G924 1975. Diese Artikel stellen eine Auswahl von Vortragen dar, die auf einer vom 4. bis 8. August 1975 im ‘Mathematischen Forschungsinstitut Oberwolfach’ stattgefundenen Tagung gehalten wurden.

Alefeld:1977:DGA

[4097] G. Alefeld. Über die Durchführbarkeit des Gaussschen Algorithmus bei Gleichungen mit Intervallen als Koeffizienten. (German) [On the feasibility of Gaussian algorithms for equations with interval coefficients]. In und Ulrich Kulisch [4366], pages 15– 19. CODEN COSPDM. ISBN 0-387-81410-8. ISSN 0344-8029. LCCN QA 162 G924 1975. Diese Artikel stellen eine Auswahl von Vortragen dar, die auf einer vom 4. bis 8. August 1975 im ‘Mathematischen Forschungsinstitut Oberwolfach’ stattgefundenen Tagung gehalten wurden.

Bohlender:1977:GSG

[4098] G. Bohlender. Genaue Summation von Gleitkommazahlen. (German) [Accurate summation of floating-point numbers]. In und Ulrich Kulisch [4366], pages 21–32. CODEN COSPDM. ISBN 0-387-81410-8. ISSN 0344-8029. LCCN QA 162 G924 1975. Diese Artikel stellen eine Auswahl von Vortragen dar, die auf einer vom 4. bis 8. August 1975 im ‘Mathematischen Forschungsinstitut Oberwolfach’ stattgefundenen Tagung gehalten wurden.

Bohlender:1977:PWG

[4099] G. Bohlender. Produkte und Wurzeln von Gleitkommazahlen. (German) [Products and roots of floating-point numbers]. In und Ulrich Kulisch [4366], pages 33–46. CODEN COSPDM. ISBN 0-387-81410-8. ISSN 0344-8029. LCCN QA 162 G924 1975. Diese Artikel stellen eine Auswahl von Vortragen dar, die auf einer vom 4. bis 8. August 1975 im ‘Mathematischen Forschungsinstitut Oberwolfach’ stattgefundenen Tagung gehalten wurden.

Gruner:1977:FLG

[4100] K. Grüner. Fehlerschranken für lineare Gleichungssysteme. (German) [Error bounds for linear equation systems]. In und Ulrich Kulisch [4366], pages 47–55. CODEN COSPDM. ISBN 0-387-81410-8. ISSN 0344-8029. LCCN QA 162 G924 1975. Diese Artikel stellen eine Auswahl von Vortragen dar, die auf einer vom 4. bis 8. August 1975 im ‘Mathematischen Forschungsinstitut Oberwolfach’ stattgefundenen Tagung gehalten wurden.

Herzberger:1977:AWR

[4101] J. Herzberger. Zur Approximation des Wertebereiches reeller Funktionen durch Intervallausdrücke. (German) [Toward approximation of ranges of real functions through interval expressions]. In und Ulrich Kulisch [4366], pages 57–64. CODEN COSPDM. ISBN 0-387-81410-8. ISSN 0344-8029. LCCN QA 162 G924 1975. Diese Artikel stellen eine Auswahl von Vortragen dar, die auf einer vom 4. bis 8. August 1975 im ‘Mathematischen Forschungsinstitut Oberwolfach’ stattgefundenen Tagung gehalten wurden.

Kaucher:1977:AEI

[4102] E. Kaucher. Algebraische Erweiterungen der Intervallrechnung unter Erhaltung der ordnungs- und Verbandsstrukturen. (German) [Algebraic extensions of interval computation under conservation of order and association structure]. In und Ulrich Kulisch [4366], pages 65–79. CODEN COSPDM. ISBN 0-387-81410-8. ISSN 0344-8029. LCCN QA 162 G924 1975. Diese Artikel stellen eine Auswahl von Vortragen dar, die auf einer vom 4. bis 8. August 1975 im ‘Mathematischen Forschungsinstitut Oberwolfach’ stattgefundenen Tagung gehalten wurden.

Kaucher:1977:EAE

[4103] E. Kaucher. Über Eigenschaften und Anwendungsmöglichkeiten der erweiterten Intervallrechnung und des Hyperbolischen Fastkörpers über R. (German) [On eigenvalues and usability of extended interval computation and hyperbolic near-fields over R ]. In und Ulrich Kulisch [4366], pages 81– 94. CODEN COSPDM. ISBN 0-387-81410-8. ISSN 0344-8029. LCCN QA 162 G924 1975. Diese Artikel stellen eine Auswahl von Vortragen dar, die auf einer vom 4. bis 8. August 1975 im ‘Mathematischen Forschungsinstitut Oberwolfach’ stattgefundenen Tagung gehalten wurden.

Kulisch:1977:KFA

[4104] U. Kulisch. Ein Konzept für eine allgemeine Theorie der Rechnerarithmetik. (German) [A draft of a general theory of computer arithmetic]. In und Ulrich Kulisch [4366], pages 95– 105. CODEN COSPDM. ISBN 0-387-81410-8. ISSN 0344-8029. LCCN QA 162 G924 1975. Diese Artikel stellen eine Auswahl von Vortragen dar, die auf einer vom 4. bis 8. August 1975 im ‘Mathematischen Forschungsinstitut Oberwolfach’ stattgefundenen Tagung gehalten wurden.

Kulisch:1977:BNR

[4105] U. Kulisch. Über die beim numerischen Rechnen mit Rechenanlagen Auftretenden Räume. (German) [On the spaces encountered in numerical computations on computers]. In und Ulrich Kulisch [4366], pages 107– 119. CODEN COSPDM. ISBN 0-387-81410-8. ISSN 0344-8029. LCCN QA 162 G924 1975. Diese Artikel stellen eine Auswahl von Vortragen dar, die auf einer vom 4. bis 8. August 1975 im ‘Mathematischen Forschungsinstitut Oberwolfach’ stattgefundenen Tagung gehalten wurden.

Ratschek:1977:FPM

[4106] H. Ratschek. Fehlererfassung mit Partiellen Mengen. (German) [Error control with partial quantities]. In und Ulrich Kulisch [4366], pages 121–128. CODEN COSPDM. ISBN 0-387-81410-8. ISSN 0344-8029. LCCN QA 162 G924 1975. Diese Artikel stellen eine Auswahl von Vortragen dar, die auf einer vom 4. bis 8. August 1975 im ‘Mathematischen Forschungsinstitut Oberwolfach’ stattgefundenen Tagung gehalten wurden.

Ullrich:1977:BRM

[4107] Chr. Ullrich. Zum Begriff des Rasters und der minimalen Rundung. (German) [Toward a raster concept and minimal rounding]. In und Ulrich Kulisch [4366], pages 129–134. CODEN COSPDM. ISBN 0-387-81410-8. ISSN 0344-8029. LCCN QA 162 G924 1975. Diese Artikel stellen eine Auswahl von Vortragen dar, die auf einer vom 4. bis 8. August 1975 im ‘Mathematischen Forschungsinstitut Oberwolfach’ stattgefundenen Tagung gehalten wurden.

Ullrich:1977:KKK

[4108] Chr. Ullrich. Zur Konstruktion komplexer Kreisarithmetiken. (German) [Toward construction of complex circular arithmetic]. In und Ulrich Kulisch [4366], pages 135–150. CODEN COSPDM. ISBN 0-387-81410-8. ISSN 0344-8029. LCCN QA 162 G924 1975. Diese Artikel stellen eine Auswahl von Vortragen dar, die auf einer vom 4. bis 8. August 1975 im ‘Mathematischen Forschungsinstitut Oberwolfach’ stattgefundenen Tagung gehalten wurden.

Albrecht:1977:GC

[4109] R. Albrecht and U. Kulisch. Grundlagen Der Computer-Arithmetik. (German) [Foundations of computer arithmetic]. In und Ulrich Kulisch [4366], page ?? CODEN COSPDM. ISBN 0-387-81410-8. ISSN 0344-8029. LCCN QA 162 G924 1975. Diese Artikel stellen eine Auswahl von Vortragen dar, die auf einer vom 4. bis 8. August 1975 im ‘Mathematischen Forschungsinstitut Oberwolfach’ stattgefundenen Tagung gehalten wurden.

Adams:1980:MCB

[4110] Ernst Adams. On methods for the construction of the boundaries of sets of solutions for differential equations or finite-dimensional approximations with input sets. In Alefeld and Grigorieff [4368], pages 1–16. CODEN COSPDM. ISBN 0-387-81566-X. ISSN 0344-8029. LCCN QA297 .F84. In cooperation with R. Albrecht, U. Kulisch, and F. Stummel. “Mainly a collection of the invited lectures which were given during a conference … held in June 5–8, 1979, on the occasion of the centennial of the Technical University of Berlin.”.

Albrecht:1980:RAO

[4111] R. Albrecht. Roundings and approximations in ordered sets. In Alefeld and Grigorieff [4368], pages 17–31. CODEN COSPDM. ISBN 0-387-81566-X. ISSN 0344-8029. LCCN QA297 .F84. In cooperation with R. Albrecht, U. Kulisch, and F. Stummel. “Mainly a collection of the invited lectures which were given during a conference … held in June 5–8, 1979, on the occasion of the centennial of the Technical University of Berlin.”.

Kaucher:1980:IAE

[4112] E. Kaucher. Interval analysis in the extended interval space IR. In Alefeld and Grigorieff [4368], pages 33–49. CO-DEN COSPDM. ISBN 0-387-81566-X. ISSN 0344-8029. LCCN QA297 .F84. In cooperation with R. Albrecht, U. Kulisch, and F. Stummel. “Mainly a collection of the invited lectures which were given during a conference … held in June 5–8, 1979, on the occasion of the centennial of the Technical University of Berlin.”.

Kulisch:1980:AOI

[4113] U. W. Kulisch and W. L. Miranker. Arithmetic operations in interval spaces. In Alefeld and Grigorieff [4368], pages 51–67. CODEN COSPDM. ISBN 0-387-81566-X. ISSN 0344-8029. LCCN QA297 .F84. In cooperation with R. Albrecht, U. Kulisch, and F. Stummel. “Mainly a collection of the invited lectures which were given during a conference … held in June 5–8, 1979, on the occasion of the centennial of the Technical University of Berlin.”.

Markov:1980:SAE

[4114] S. M. Markov. Some applications of extended interval arithmetic to interval iterations. In Alefeld and Grigorieff [4368], pages 69–84. CODEN COSPDM. ISBN 0-387-81566-X. ISSN 0344-8029. LCCN QA297 .F84. In cooperation with R. Albrecht, U. Kulisch, and F. Stummel. “Mainly a collection of the invited lectures which were given during a conference … held in June 5–8, 1979, on the occasion of the centennial of the Technical University of Berlin.”.

Matula:1980:FFP

[4115] D. W. Matula and P. Kornerup. Foundations of finite precision rational arithmetic. In Alefeld and Grigorieff [4368], pages 85–111. CODEN COSPDM. ISBN 0-387-81566-X. ISSN 0344-8029. LCCN QA297 .F84. In cooperation with R. Albrecht, U. Kulisch, and F. Stummel. “Mainly a collection of the invited lectures which were given during a conference … held in June 5–8, 1979, on the occasion of the centennial of the Technical University of Berlin.”.

Moore:1980:IMN

[4116] R. E. Moore. Interval methods for nonlinear systems. In Alefeld and Grigorieff [4368], pages 113–120. CODEN COSPDM. ISBN 0-387-81566-X. ISSN 0344-8029. LCCN QA297 .F84. In cooperation with R. Albrecht, U. Kulisch, and F. Stummel. “Mainly a collection of the invited lectures which were given during a conference … held in June 5–8, 1979, on the occasion of the centennial of the Technical University of Berlin.”.

Oberaigner:1980:AMG

[4117] W. Oberaigner. Algorithms for multiplication with given precision. In Alefeld and Grigorieff [4368], pages 121– 129. CODEN COSPDM. ISBN 0-387-81566-X. ISSN 0344-8029. LCCN QA297 .F84. In cooperation with R. Albrecht, U. Kulisch, and F. Stummel. “Mainly a collection of the invited lectures which were given during a conference … held in June 5–8, 1979, on the occasion of the centennial of the Technical University of Berlin.”.

Olver:1980:UAG

[4118] F. W. J. Olver. Unrestricted algorithms for generating elementary functions. In Alefeld and Grigorieff [4368], pages 131–140. CODEN COSPDM. ISBN 0-387-81566-X. ISSN 0344-8029. LCCN QA297 .F84. In cooperation with R. Albrecht, U. Kulisch, and F. Stummel. “Mainly a collection of the invited lectures which were given during a conference … held in June 5–8, 1979, on the occasion of the centennial of the Technical University of Berlin.”.

Rall:1980:ASA

[4119] Louis B. Rall. Applications of software for automatic differentiation in numerical computation. In Alefeld and Grigorieff [4368], pages 141–156. CODEN COSPDM. ISBN 0-387-81566-X. ISSN 0344-8029. LCCN QA297 .F84. In cooperation with R. Albrecht, U. Kulisch, and F. Stummel. “Mainly a collection of the invited lectures which were given during a conference … held in June 5–8, 1979, on the occasion of the centennial of the Technical University of Berlin.”.

Rump:1980:SBS

[4120] S. M. Rump and E. Kaucher. Small bounds for the solution of systems of linear equations. In Alefeld and Grigorieff [4368], pages 157–164. CODEN COSPDM. ISBN 0-387-81566-X. ISSN 0344-8029. LCCN QA297 .F84. In cooperation with R. Albrecht, U. Kulisch, and F. Stummel. “Mainly a collection of the invited lectures which were given during a conference … held in June 5–8, 1979, on the occasion of the centennial of the Technical University of Berlin.”.

Scherer:1980:SNR

[4121] R. Scherer and K. Zeller. Shorthand notation for rounding errors. In Alefeld and Grigorieff [4368], pages 165– 168. CODEN COSPDM. ISBN 0-387-81566-X. ISSN 0344-8029. LCCN QA297 .F84. In cooperation with R. Albrecht, U. Kulisch, and F. Stummel. “Mainly a collection of the invited lectures which were given during a conference … held in June 5–8, 1979, on the occasion of the centennial of the Technical University of Berlin.”.

Stummel:1980:REA

[4122] F. Stummel. Rounding error analysis of elementary numerical algorithms. In Alefeld and Grigorieff [4368], pages 169–195. CODEN COSPDM. ISBN 0-387-81566-X. ISSN 0344-8029. LCCN QA297 .F84. In cooperation with R. Albrecht, U. Kulisch, and F. Stummel. “Mainly a collection of the invited lectures which were given during a conference … held in June 5–8, 1979, on the occasion of the centennial of the Technical University of Berlin.”.

Ullrich:1980:IMS

[4123] Chr. Ullrich. Iterative methods in the spaces of rounded computations. In Alefeld and Grigorieff [4368], pages 197–209. CODEN COSPDM. ISBN 0-387-81566-X. ISSN 0344-8029. LCCN QA297 .F84. In cooperation with R. Albrecht, U. Kulisch, and F. Stummel. “Mainly a collection of the invited lectures which were given during a conference … held in June 5–8, 1979, on the occasion of the centennial of the Technical University of Berlin.”.

Yohe:1980:PSI

[4124] J. M. Yohe. Portable software for interval arithmetic. In Alefeld and Grigorieff [4368], pages 211–229. CODEN COSPDM. ISBN 0-387-81566-X. ISSN 0344-8029. LCCN QA297 .F84. In cooperation with R. Albrecht, U. Kulisch, and F. Stummel. “Mainly a collection of the invited lectures which were given during a conference … held in June 5–8, 1979, on the occasion of the centennial of the Technical University of Berlin.”.

Beth:1981:GPE

[4125] T. Beth. Geometric parallelism enforces parallel algorithms. In Knödel and Schneider [4369], pages 1–8. CO-DEN COSPDM. ISBN 0-387-81606-2. ISSN 0344-8029. LCCN QA76.6 .P353. Essays in honor of W. Handler.

Bohme:1981:VSF

[4126] J. F. Böhme. Vektorprozessoren zur Sonar-Feldverarbeitung. (German) [Vector processors for sonar array processing]. In Knödel and Schneider [4369], pages 9–24. CODEN COSPDM. ISBN 0-387-81606-2. ISSN 0344-8029. LCCN QA76.6 .P353. Essays in honor of W. Handler.

Hellmold:1981:PSG

[4127] K. U. Hellmold and B. Schmidt. Parallelstruktur des Simulators GPSS-FORTRAN — Entwurf für ein angepaßtes Multiprozessorsystem. (German) [Parallel structure of the simulator GPSS-FORTRAN — a design for a matching multiprocessor system]. In Knödel and Schneider [4369], pages 25–39. CODEN COSPDM. ISBN 0-387-81606-2. ISSN 0344-8029. LCCN QA76.6 .P353. Essays in honor of W. Handler.

Herzog:1981:EMV

[4128] U. Herzog and W. Kleinoder. Einführung in die Methodik der Verkehrstheorie und ihre Anwendung bei Multiprozessor-Rechenanlagen. (German) [Traffic theory and its use for multiprocessor systems]. In Knödel and Schneider [4369], pages 41–64. CODEN COSPDM. ISBN 0-387-81606-2. ISSN 0344-8029. LCCN QA76.6 .P353. Essays in honor of W. Handler.

Klar:1981:HMT

[4129] R. Klar. Hardware measurements and their application on performance evaluation in a processor-array. In Knödel and Schneider [4369], pages 65–88. CO-DEN COSPDM. ISBN 0-387-81606-2. ISSN 0344-8029. LCCN QA76.6 .P353. Essays in honor of W. Handler.

Muller:1981:RPP

[4130] H. Müller. On the reachability problem for persistent vector replacement systems. In Knödel and Schneider [4369], pages 89–104. CODEN COSPDM. ISBN 0-387-81606-2. ISSN 0344-8029. LCCN QA76.6 .P353. Essays in honor of W. Handler.

Nagl:1981:AGR

[4131] M. Nagl. Application of graph rewriting to optimization and parallelization of programs. In Knödel and Schneider [4369], pages 105–124. CODEN COSPDM. ISBN 0-387-81606-2. ISSN 0344-8029. LCCN QA76.6 .P353. Essays in honor of W. Handler.

Nees:1981:CK

[4132] G. Nees. CAD und Kollaterale. (German) [CAD and collaterals]. In Knödel and Schneider [4369], pages 125–139. CODEN COSPDM. ISBN 0-387-81606-2. ISSN 0344-8029. LCCN QA76.6 .P353. Essays in honor of W. Handler.

Niemann:1981:PSP

[4133] H. Niemann and H.-W. Hein. A program system of parallel processes for understanding continuous speech. In Knödel and Schneider [4369], pages 141–148. CODEN COSPDM. ISBN 0-387-81606-2. ISSN 0344-8029. LCCN QA76.6 .P353. Essays in honor of W. Handler.

Rohrer:1981:MMI

[4134] H. Rohrer, H. Schreiber, and V. Sigmund. Micro-modules implementation of a learning pattern-recognition system. In Knödel and Schneider [4369], pages 149–157. CODEN COSPDM. ISBN 0-387-81606-2. ISSN 0344-8029. LCCN QA76.6 .P353. Essays in honor of W. Handler.

Schmitt:1981:MDS

[4135] A. Schmitt. Ein Modell für diskrete Systeme mit reellem Zeitparameter. (German) [A model of discrete systems with real time parameters]. In Knödel and Schneider [4369], pages 159–171. CODEN COSPDM. ISBN 0-387-81606-2. ISSN 0344-8029. LCCN QA76.6 .P353. Essays in honor of W. Handler.

Schneider:1981:ASS

[4136] H. J. Schneider. Algebraic specification of the software controlling a parcels distribution machine. In Knödel and Schneider [4369], pages 173–180. CO-DEN COSPDM. ISBN 0-387-81606-2. ISSN 0344-8029. LCCN QA76.6 .P353. Essays in honor of W. Handler.

Vollmar:1981:CAF

[4137] R. Vollmar. On cellular automata with a finite number of state changes. In Knödel and Schneider [4369], pages 181–191. CODEN COSPDM. ISBN 0-387-81606-2. ISSN 0344-8029. LCCN QA76.6 .P353. Essays in honor of W. Handler.

Wendler:1981:MDF

[4138] K. Wendler. Models to describe those features of cellular computer-nets which are relevant to the operating system. In Knödel and Schneider [4369], pages 193–203. CODEN COSPDM. ISBN 0-387-81606-2. ISSN 0344-8029. LCCN QA76.6 .P353. Essays in honor of W. Handler.

Loos:1982:I

[4139] R. Loos. Introduction. In Buchberger et al. [4370], pages 1–10. CODEN COSPDM. ISBN 0-387-81684-4. ISSN 0344-8029. LCCN QA155.7.E4 C65 1982. In cooperation with R. Albrecht.

Buchberger:1982:AS

[4140] B. Buchberger and R. Loos. Algebraic simplification. In Buchberger et al. [4370], pages 11–43. CODEN COSPDM. ISBN 0-387-81684-4. ISSN 0344-8029. LCCN QA155.7.E4 C65 1982. In cooperation with R. Albrecht.

Neubuser:1982:CGT

[4141] J. Neubüser. Computing with groups and their character tables. In Buchberger et al. [4370], pages 45–56. CO-DEN COSPDM. ISBN 0-387-81684-4. ISSN 0344-8029. LCCN QA155.7.E4 C65 1982. In cooperation with R. Albrecht.

Norman:1982:IFT

[4142] A. C. Norman. Integration in finite terms. In Buchberger et al. [4370], pages 57–69. CODEN COSPDM. ISBN 0-387-81684-4. ISSN 0344-8029. LCCN QA155.7.E4 C65 1982. In cooperation with R. Albrecht.

Lafon:1982:SFT

[4143] J. C. Lafon. Summation in finite terms. In Buchberger et al. [4370], pages 71– 77. CODEN COSPDM. ISBN 0-387-81684-4. ISSN 0344-8029. LCCN QA155.7.E4 C65 1982. In cooperation with R. Albrecht.

Collins:1982:QER

[4144] G. E. Collins. Quantifier elimination for real closed fields: a guide to the literature in computing. In Buchberger et al. [4370], pages 79–81. CODEN COSPDM. ISBN 0-387-81684-4. ISSN 0344-8029. LCCN QA155.7.E4 C65 1982. In cooperation with R. Albrecht.

Collins:1982:RZP

[4145] G. E. Collins and R. Loos. Real zeros of polynomials. In Buchberger et al. [4370], pages 83–94. CODEN COSPDM. ISBN 0-387-81684-4. ISSN 0344-8029. LCCN QA155.7.E4 C65 1982. In cooperation with R. Albrecht.

Kaltofen:1982:FP

[4146] E. Kaltofen. Factorization of polynomials. In Buchberger et al. [4370], pages 95–113. CODEN COSPDM. ISBN 0-387-81684-4. ISSN 0344-8029. LCCN QA155.7.E4 C65 1982. In cooperation with R. Albrecht.

Loos:1982:GPR

[4147] R. Loos. Generalized polynomial remainder sequences. In Buchberger et al. [4370], pages 115–137. CODEN COSPDM. ISBN 0-387-81684-4. ISSN 0344-8029. LCCN QA155.7.E4 C65 1982. In cooperation with R. Albrecht.

Lauer:1982:CHI

[4148] M. Lauer. Computing by homomorphic images. In Buchberger et al. [4370], pages 139–168. CODEN COSPDM. ISBN 0-387-81684-4. ISSN 0344-8029. LCCN QA155.7.E4 C65 1982. In cooperation with R. Albrecht.

Norman:1982:CTE

[4149] A. C. Norman. Computing in transcendental extensions. In Buchberger et al. [4370], pages 169–172. CODEN COSPDM. ISBN 0-387-81684-4. ISSN 0344-8029. LCCN QA155.7.E4 C65 1982. In cooperation with R. Albrecht.

Loos:1982:CAE

[4150] R. Loos. Computing in algebraic extensions. In Buchberger et al. [4370], pages 173–188. CODEN COSPDM. ISBN 0-387-81684-4. ISSN 0344-8029. LCCN QA155.7.E4 C65 1982. In cooperation with R. Albrecht.

Collins:1982:ABA

[4151] G. E. Collins, M. Mignotte, and F. Winkler. Arithmetic in basic algebraic domains. In Buchberger et al. [4370], pages 189–220. CODEN COSPDM. ISBN 0-387-81684-4. ISSN 0344-8029. LCCN QA155.7.E4 C65 1982. In cooperation with R. Albrecht.

vanHulzen:1982:CAS

[4152] J. A. van Hulzen and J. Calmet. Computer algebra systems. In Buchberger et al. [4370], pages 221–243. CODEN COSPDM. ISBN 0-387-81684-4. ISSN 0344-8029. LCCN QA155.7.E4 C65 1982. In cooperation with R. Albrecht.

Calmet:1982:CAA

[4153] J. Calmet and J. A. van Hulzen. Computer algebra applications. In Buchberger et al. [4370], pages 245–258. CO-DEN COSPDM. ISBN 0-387-81684-4. ISSN 0344-8029. LCCN QA155.7.E4 C65 1982. In cooperation with R. Albrecht.

Mignotte:1982:SUB

[4154] M. Mignotte. Some useful bounds. In Buchberger et al. [4370], pages 259– 263. CODEN COSPDM. ISBN 0-387-81684-4. ISSN 0344-8029. LCCN QA155.7.E4 C65 1982. In cooperation with R. Albrecht.

Anonymous:1982:ASI

[4155] Anonymous. Author and subject index. In Buchberger et al. [4370], pages 265–283. CODEN COSPDM. ISBN 0-387-81684-4. ISSN 0344-8029. LCCN QA155.7.E4 C65 1982. In cooperation with R. Albrecht.

Boehmer:1984:DCA

[4156] K. Boehmer, P. Hemker, and H. J. Stetter. The defect correction approach. In Böhmer and Stetter [4372], pages 1–32. CODEN COSPDM. ISBN 0-387-81832-4. ISSN 0344-8029. LCCN QA297.5 .D44 1984.

Frank:1984:DCA

[4157] R. Frank, J. Hertling, and H. Lehner. Defect correction algorithms for stiff ordinary differential equations. In Böhmer and Stetter [4372], pages 33– 41. CODEN COSPDM. ISBN 0-387-81832-4. ISSN 0344-8029. LCCN QA297.5 .D44 1984.

Reinhardt:1984:PDD

[4158] H. J. Reinhardt. On a principal of direct defect correction based on a posteriori error estimates. In Böhmer and Stetter [4372], pages 43–66. CODEN COSPDM. ISBN 0-387-81832-4. ISSN 0344-8029. LCCN QA297.5 .D44 1984.

Chatelin:1984:SNI

[4159] F. Chatelin. Simultaneous Newton’s iteration for the eigenproblem. In Böhmer and Stetter [4372], pages 67– 74. CODEN COSPDM. ISBN 0-387-81832-4. ISSN 0344-8029. LCCN QA297.5 .D44 1984.

Mandel:1984:STL

[4160] J. Mandel. On some two-level iterative methods. In Böhmer and Stetter [4372], pages 75–88. CODEN COSPDM. ISBN 0-387-81832-4. ISSN 0344-8029. LCCN QA297.5 .D44 1984.

Hackbusch:1984:LDC

[4161] W. Hackbusch. Local defect correction method and domain decomposition techniques. In Böhmer and Stetter [4372], pages 89–113. CODEN COSPDM. ISBN 0-387-81832-4. ISSN 0344-8029. LCCN QA297.5 .D44 1984.

McCormick:1984:FAC

[4162] Steve McCormick. Fast adaptive composite grid (FAC) methods. In Böhmer and Stetter [4372], pages 115–121. CO-DEN COSPDM. ISBN 0-387-81832-4. ISSN 0344-8029. LCCN QA297.5 .D44 1984.

Hemker:1984:MDCa

[4163] P. W. Hemker. Mixed defect correction iteration for the solution of a singular perturbation problem. In Böhmer and Stetter [4372], pages 123–145. CODEN COSPDM. ISBN 0-387-81832-4. ISSN 0344-8029. LCCN QA297.5 .D44 1984.

Rump:1984:SLN

[4164] S. M. Rump. Solution of linear and nonlinear algebraic problems with sharp, guaranteed bounds. In Böhmer and Stetter [4372], pages 147–168. CO-DEN COSPDM. ISBN 0-387-81832-4. ISSN 0344-8029. LCCN QA297.5 .D44 1984.

Kaucher:1984:RCV

[4165] E. Kaucher and W. L. Miranker. Residual correction and validation in functoids. In Böhmer and Stetter [4372], pages 169–192. CODEN COSPDM. ISBN 0-387-81832-4. ISSN 0344-8029. LCCN QA297.5 .D44 1984.

Bohmer:1984:DCH

[4166] K. Böhmer, W. Gross, B. Schmitt, and R. Schwarz. Defect corrections and Hartree–Fock method. In Böhmer and Stetter [4372], pages 193–209. CODEN COSPDM. ISBN 0-387-81832-4. ISSN 0344-8029. LCCN QA297.5 .D44 1984.

Pereyra:1984:DCS

[4167] V. Pereyra. Deferred corrections software and its application to seismic ray tracing. In Böhmer and Stetter [4372], pages 211–226. CODEN COSPDM. ISBN 0-387-81832-4. ISSN 0344-8029. LCCN QA297.5 .D44 1984.

Schonauer:1984:NEE

[4168] W. Schönauer, W. Schnepf, and K. Raith. Numerical engineering: experiences in designing PDE software with selfadaptive variable step size/variable order difference methods. In Böhmer and Stetter [4372], pages 227–242. CO-DEN COSPDM. ISBN 0-387-81832-4. ISSN 0344-8029. LCCN QA297.5 .D44 1984.

Kulisch:1988:ARV

[4169] U. Kulisch and H. J. Stetter. Automatic result verification. In Kulisch and Stetter [4373], pages 1–6. CODEN COSPDM. ISBN 0-387-82063-9. ISSN 0344-8029. LCCN QA297 .S392 1988. Based on papers presented at the conference on computer arithmetic and scientific computation held Sep. 30–Oct. 2, 1987 in Karlsruhe, FRG.

Schroder:1988:MPV

[4170] J. Schröder. A method for producing verified results for two-point boundary value problems. In Kulisch and Stetter [4373], pages 9–22. CODEN COSPDM. ISBN 0-387-82063-9. ISSN 0344-8029. LCCN QA297 .S392 1988. Based on papers presented at the conference on computer arithmetic and scientific computation held Sep. 30–Oct. 2, 1987 in Karlsruhe, FRG.

Weissinger:1988:KDM

[4171] J. Weissinger. A kind of difference method for enclosing solutions of ordinary linear boundary value problems. In Kulisch and Stetter [4373], pages 23–31. CODEN COSPDM. ISBN 0-387-82063-9. ISSN 0344-8029. LCCN QA297 .S392 1988. Based on papers presented at the conference on computer arithmetic and scientific computation held Sep. 30–Oct. 2, 1987 in Karlsruhe, FRG.

Jansson:1988:SVM

[4172] Ch. Jansson. A self-validating method for solving linear programming problems with interval input data. In Kulisch and Stetter [4373], pages 33– 45. CODEN COSPDM. ISBN 0-387-82063-9. ISSN 0344-8029. LCCN QA297 .S392 1988. Based on papers presented at the conference on computer arithmetic and scientific computation held Sep. 30–Oct. 2, 1987 in Karlsruhe, FRG.

Mayer:1988:ESL

[4173] G. Mayer. Enclosing the solutions of linear equations by interval iterative processes. In Kulisch and Stetter [4373], pages 47–58. CODEN COSPDM. ISBN 0-387-82063-9. ISSN 0344-8029. LCCN QA297 .S392 1988. Based on papers presented at the conference on computer arithmetic and scientific computation held Sep. 30–Oct. 2, 1987 in Karlsruhe, FRG.

Alefeld:1988:EQS

[4174] G. Alefeld. Errorbounds for quadratic systems of nonlinear equations using the precise scalar product. In Kulisch and Stetter [4373], pages 59–67. CO-DEN COSPDM. ISBN 0-387-82063-9. ISSN 0344-8029. LCCN QA297 .S392 1988. Based on papers presented at the conference on computer arithmetic and scientific computation held Sep. 30–Oct. 2, 1987 in Karlsruhe, FRG.

Behnke:1988:IEG

[4175] H. Behnke. Inclusion of eigenvalues of general eigenvalue problems for matrices. In Kulisch and Stetter [4373], pages 69–78. CODEN COSPDM. ISBN 0-387-82063-9. ISSN 0344-8029. LCCN QA297 .S392 1988. Computer Arithmetic and Scientific Computation.

Ohsmann:1988:VIE

[4176] M. Ohsmann. Verified inclusion for eigenvalues of certain difference and differential equations. In Kulisch and Stetter [4373], pages 79–88. CODEN COSPDM. ISBN 0-387-82063-9. ISSN 0344-8029. LCCN QA297 .S392 1988. Computer Arithmetic and Scientific Computation.

Ams:1988:VVI

[4177] A. Ams and W. Klein. VIB — verified inclusions of critical bending vibrations. In Kulisch and Stetter [4373], pages 91–98. CODEN COSPDM. ISBN 0-387-82063-9. ISSN 0344-8029. LCCN QA297 .S392 1988. Based on papers presented at the conference on computer arithmetic and scientific computation held Sep. 30–Oct. 2, 1987 in Karlsruhe, FRG.

Cordes:1988:STP

[4178] D. Cordes. Stability test for periodic differential equations on digital computers with applications. In Kulisch and Stetter [4373], pages 99–110. CO-DEN COSPDM. ISBN 0-387-82063-9. ISSN 0344-8029. LCCN QA297 .S392 1988. Based on papers presented at the conference on computer arithmetic and scientific computation held Sep. 30–Oct. 2, 1987 in Karlsruhe, FRG.

Adams:1988:PSO

[4179] E. Adams, A. Holzmüller, and D. Straub. The periodic solutions of the Oregonator and verification of results. In Kulisch and Stetter [4373], pages 111–122. CODEN COSPDM. ISBN 0-387-82063-9. ISSN 0344-8029. LCCN QA297 .S392 1988. Based on papers presented at the conference on computer arithmetic and scientific computation held Sep. 30–Oct. 2, 1987 in Karlsruhe, FRG.

Ottmann:1988:APG

[4180] Th. Ottmann, G. Thiemt, and Ch. Ullrich. On arithmetical problems of geometric algorithms in the plane. In Kulisch and Stetter [4373], pages 123– 136. CODEN COSPDM. ISBN 0-387-82063-9. ISSN 0344-8029. LCCN QA297 .S392 1988. Based on papers presented at the conference on computer arithmetic and scientific computation held Sep. 30–Oct. 2, 1987 in Karlsruhe, FRG.

Lohner:1988:PEP

[4181] R. Lohner. Precise evaluation of polynomials in several variables. In Kulisch and Stetter [4373], pages 139–148. CO-DEN COSPDM. ISBN 0-387-82063-9. ISSN 0344-8029. LCCN QA297 .S392 1988. Based on papers presented at the conference on computer arithmetic and scientific computation held Sep. 30–Oct. 2, 1987 in Karlsruhe, FRG.

Fischer:1988:EAE

[4182] H. C. Fischer, G. Schumacher, and R. Haggenmuller. Evaluation of arithmetic expressions with guaranteed high accuracy. In Kulisch and Stetter [4373], pages 149–158. CODEN COSPDM. ISBN 0-387-82063-9. ISSN 0344-8029. LCCN QA297 .S392 1988. Computer Arithmetic and Scientific Computation.

Braune:1988:SFR

[4183] Klaus Braune. Standard functions for real and complex point and interval arguments with dynamic accuracy. In Kulisch and Stetter [4373], pages 159– 184. CODEN COSPDM. ISBN 0-387-82063-9. ISSN 0344-8029. LCCN QA297 .S392 1988. Based on papers presented at the conference on computer arithmetic and scientific computation held Sep. 30–Oct. 2, 1987 in Karlsruhe, FRG.

Kramer:1988:ISF

[4184] W. Krämer. Inverse standard functions for real and complex point and interval arguments with dynamic accuracy. In Kulisch and Stetter [4373], pages 185–212. CODEN COSPDM. ISBN 0-387-82063-9. ISSN 0344-8029. LCCN QA297 .S392 1988. Computer Arithmetic and Scientific Computation.

Stetter:1988:IAF

[4185] H. J. Stetter. Inclusions algorithms with functions as data. In Kulisch and Stetter [4373], pages 213–224. CODEN COSPDM. ISBN 0-387-82063-9. ISSN 0344-8029. LCCN QA297 .S392 1988. Based on papers presented at the conference on computer arithmetic and scientific computation held Sep. 30–Oct. 2, 1987 in Karlsruhe, FRG.

Bleher:1988:FSS

[4186] J. H. Bleher, S. M. Rump, U. Kulisch, M. Metzger, Ch. Ullrich, and W. Walter. FORTRAN-SC: a study of a FORTRAN extension for engineering/ scientific computation with access to ACRITH. In Kulisch and Stetter [4373], pages 227–244. CODEN COSPDM. ISBN 0-387-82063-9. ISSN 0344-8029. LCCN QA297 .S392 1988. Based on papers presented at the conference on computer arithmetic and scientific computation held Sep. 30–Oct. 2, 1987 in Karlsruhe, FRG.

Proskurowski:1990:ECT

[4187] A. Proskurowski and M. M. Systo. Efficient computations in tree-like graphs. In Tinhofer et al. [4374], pages 1– 15. CODEN COSPDM. ISBN 0-387-82177-5 (New York), 3-211-82177-5 (Wien). ISSN 0344-8029. LCCN QA166 .C618 1990.

Mohring:1990:GPR

[4188] R. H. Möhring. Graph problems related to gate matrix layout and PLA folding. In Tinhofer et al. [4374], pages 17–51. CODEN COSPDM. ISBN 0-387-82177-5 (New York), 3-211-82177-5 (Wien). ISSN 0344-8029. LCCN QA166 .C618 1990.

Nishizeki:1990:PGP

[4189] T. Nishizeki. Planar graph problems. In Tinhofer et al. [4374], pages 53– 68. CODEN COSPDM. ISBN 0-387-82177-5 (New York), 3-211-82177-5 (Wien). ISSN 0344-8029. LCCN QA166 .C618 1990.

Mayr:1990:BPA

[4190] E. W. Mayr. Basic parallel algorithms in graph theory. In Tinhofer et al. [4374], pages 69–91. CODEN COSPDM. ISBN 0-387-82177-5 (New York), 3-211-82177-5 (Wien). ISSN 0344-8029. LCCN QA166 .C618 1990.

Helmbold:1990:APS

[4191] D. Helmbold and E. W. Mayr. Applications of parallel scheduling algorithms to families of perfect graphs. In Tinhofer et al. [4374], pages 93–107. CO-DEN COSPDM. ISBN 0-387-82177-5 (New York), 3-211-82177-5 (Wien). ISSN 0344-8029. LCCN QA166 .C618 1990.

Faigle:1990:OG

[4192] U. Faigle and R. Schrader. Orders and graphs. In Tinhofer et al. [4374], pages 109–124. CODEN COSPDM. ISBN 0-387-82177-5 (New York), 3-211-82177-5 (Wien). ISSN 0344-8029. LCCN QA166 .C618 1990.

Noltemeier:1990:DPO

[4193] H. Noltemeier. Dynamic partial orders and generalized heaps. In Tinhofer et al. [4374], pages 125–139. CO-DEN COSPDM. ISBN 0-387-82177-5 (New York), 3-211-82177-5 (Wien). ISSN 0344-8029. LCCN QA166 .C618 1990.

Faigle:1990:CC

[4194] U. Faigle and G. Turan. Communication complexity. In Tinhofer et al. [4374], pages 141–153. CODEN COSPDM. ISBN 0-387-82177-5 (New York), 3-211-82177-5 (Wien). ISSN 0344-8029. LCCN QA166 .C618 1990.

Rote:1990:PPG

[4195] G. Rote. Path problems in graphs.

In Tinhofer et al. [4374], pages 155– 189. CODEN COSPDM. ISBN 0-387-82177-5 (New York), 3-211-82177-5 (Wien). ISSN 0344-8029. LCCN QA166 .C618 1990.

deWerra:1990:HGC

[4196] D. de Werra. Heuristics for graph coloring. In Tinhofer et al. [4374], pages 191–208. CODEN COSPDM. ISBN 0-387-82177-5 (New York), 3-211-82177-5 (Wien). ISSN 0344-8029. LCCN QA166 .C618 1990.

Frieze:1990:PAG

[4197] Alan M. Frieze. Probabilistic analysis of graph algorithms. In Tinhofer et al. [4374], pages 209–233. CODEN COSPDM. ISBN 0-387-82177-5 (New York), 3-211-82177-5 (Wien). ISSN 0344-8029. LCCN QA166 .C618 1990.

Tinhofer:1990:GGU

[4198] G. Tinhofer. Generating graphs uniformly at random. In Tinhofer et al. [4374], pages 235–255. CODEN COSPDM. ISBN 0-387-82177-5 (New York), 3-211-82177-5 (Wien). ISSN 0344-8029. LCCN QA166 .C618 1990.

Monien:1990:EOI

[4199] B. Monien and H. Sudborough. Embedding one interconnection network in another. In Tinhofer et al. [4374], pages 257–282. CODEN COSPDM. ISBN 0-387-82177-5 (New York), 3-211-82177-5 (Wien). ISSN 0344-8029. LCCN QA166 .C618 1990.

Barnhill:1993:CRB

[4200] R. E. Barnhill, G. Farin, and Q. Chen. Constant-radius blending of parametric surfaces. In Farin et al. [4375], pages 1–20. CODEN COSPDM. ISBN 0-387-82399-9, 3-211-82399-9. ISSN 0344-8029. LCCN QA448.D38 G458 1993. Based on lectures given at the 1st Dagstuhl-seminar, Dagstuhl, Germany.

Bloor:1993:FSO

[4201] M. I. G. Bloor and M. J. Wilson. Functionality in solids obtained from partial differential equations. In Farin et al. [4375], pages 21–42. CODEN COSPDM. ISBN 0-387-82399-9, 3-211-82399-9. ISSN 0344-8029. LCCN QA448.D38 G458 1993. Based on lectures given at the 1st Dagstuhlseminar, Dagstuhl, Germany.

Brandenburg:1993:FOO

[4202] W. Brandenburg and B. Wördenweber.

Featuremodelling with an objectoriented approach. In Farin et al. [4375], pages 43–57. CODEN COSPDM. ISBN 0-387-82399-9, 3-211-82399-9. ISSN 0344-8029. LCCN QA448.D38 G458 1993. Based on lectures given at the 1st Dagstuhlseminar, Dagstuhl, Germany.

Degen:1993:BAP

[4203] W. L. F. Degen. Best approximations of parametric curves by splines. In Farin et al. [4375], pages 59–73. CO-DEN COSPDM. ISBN 0-387-82399-9, 3-211-82399-9. ISSN 0344-8029. LCCN QA448.D38 G458 1993. Based on lectures given at the 1st Dagstuhlseminar, Dagstuhl, Germany.

Brunet:1993:MSA

[4204] P. Brunet, I. Navazo, and A. Vinacua.

A modelling scheme for the approximate representation of closed surfaces. In Farin et al. [4375], pages 75–90. CO-DEN COSPDM. ISBN 0-387-82399-9, 3-211-82399-9. ISSN 0344-8029. LCCN QA448.D38 G458 1993. Based on lectures given at the 1st Dagstuhlseminar, Dagstuhl, Germany.

Foley:1993:CBD

[4205] T. A. Foley, S. Dayanand, and R. Santhanam. Cross boundary derivatives for transfinite triangular patches. In Farin et al. [4375], pages 91–100. CO-DEN COSPDM. ISBN 0-387-82399-9, 3-211-82399-9. ISSN 0344-8029. LCCN QA448.D38 G458 1993. Based on lectures given at the 1st Dagstuhlseminar, Dagstuhl, Germany.

Goodman:1993:RCS

[4206] T. N. T. Goodman, B. H. Ong, and K. Unsworth. Reconstruction of C^{1} closed surfaces with branching. In Farin et al. [4375], pages 101–115. CO-DEN COSPDM. ISBN 0-387-82399-9, 3-211-82399-9. ISSN 0344-8029. LCCN QA448.D38 G458 1993. Based on lectures given at the 1st Dagstuhlseminar, Dagstuhl, Germany.

Gregory:1993:HOC

[4207] J. A. Gregory and K. H. Lau. High order continuous polygonal patches. In Farin et al. [4375], pages 117–132. CO-DEN COSPDM. ISBN 0-387-82399-9, 3-211-82399-9. ISSN 0344-8029. LCCN QA448.D38 G458 1993. Based on lectures given at the 1st Dagstuhlseminar, Dagstuhl, Germany.

Hagen:1993:VDS

[4208] H. Hagen and G.-P. Bonneau. Variational design of smooth rational Bézier surfaces. In Farin et al. [4375], pages 133–138. CODEN COSPDM. ISBN 0-387-82399-9, 3-211-82399-9. ISSN 0344-8029. LCCN QA448.D38 G458 1993. Based on lectures given at the 1st Dagstuhl-seminar, Dagstuhl, Germany.

Hamann:1993:CAT

[4209] B. Hamann. Curvature approximation for triangulated surfaces. In Farin et al. [4375], pages 139–153. CO-DEN COSPDM. ISBN 0-387-82399-9, 3-211-82399-9. ISSN 0344-8029. LCCN QA448.D38 G458 1993. Based on lectures given at the 1st Dagstuhlseminar, Dagstuhl, Germany.

Lasser:1993:CTP

[4210] D. Lasser. Composition of tensor product Bézier representations. In Farin et al. [4375], pages 155–172. CO-DEN COSPDM. ISBN 0-387-82399-9, 3-211-82399-9. ISSN 0344-8029. LCCN QA448.D38 G458 1993. Based on lectures given at the 1st Dagstuhlseminar, Dagstuhl, Germany.

Koch:1993:IEB

[4211] P. E. Koch and T. Lyche. Interpolation with exponential B-splines in tension. In Farin et al. [4375], pages 173–190. CODEN COSPDM. ISBN 0-387-82399-9, 3-211-82399-9. ISSN 0344-8029. LCCN QA448.D38 G458 1993. Based on lectures given at the 1st Dagstuhl-seminar, Dagstuhl, Germany.

Nielson:1993:CAI

[4212] G. Nielson. A characterization of an affine invariant triangulation. In Farin et al. [4375], pages 191–210. CO-DEN COSPDM. ISBN 0-387-82399-9, 3-211-82399-9. ISSN 0344-8029. LCCN QA448.D38 G458 1993. Based on lectures given at the 1st Dagstuhlseminar, Dagstuhl, Germany.

Noltemeier:1993:DSR

[4213] H. Noltemeier, K. Berbarg, and C. Zirkelbach. A data structure for representing and efficient querying large scenes of geometric objects: MB* trees. In Farin et al. [4375], pages 211–226. CODEN COSPDM. ISBN 0-387-82399-9, 3-211-82399-9. ISSN 0344-8029. LCCN QA448.D38 G458 1993. Based on lectures given at the 1st Dagstuhl-seminar, Dagstuhl, Germany.

Piper:1993:PLC

[4214] B. Piper. Properties of local coordinates based on Dirichlet tesselations. In Farin et al. [4375], pages 227–240. CO-DEN COSPDM. ISBN 0-387-82399-9, 3-211-82399-9. ISSN 0344-8029. LCCN QA448.D38 G458 1993. Based on lectures given at the 1st Dagstuhlseminar, Dagstuhl, Germany.

Pratt:1993:AFR

[4215] M. J. Pratt. Automated feature recognition and its role in product modelling. In Farin et al. [4375], pages 241–250. CODEN COSPDM. ISBN 0-387-82399-9, 3-211-82399-9. ISSN 0344-8029. LCCN QA448.D38 G458 1993. Based on lectures given at the 1st Dagstuhl-seminar, Dagstuhl, Germany.

Prautzsch:1993:ABT

[4216] H. Prautzsch. Approximate C^{r} blending with tensor product polynomials. In Farin et al. [4375], pages 251– 258. CODEN COSPDM. ISBN 0-387-82399-9, 3-211-82399-9. ISSN 0344-8029. LCCN QA448.D38 G458 1993. Based on lectures given at the 1st Dagstuhl-seminar, Dagstuhl, Germany.

Roller:1993:SII

[4217] D. Roller. Shape information in industry specific product data model. In Farin et al. [4375], pages 259–266. CO-DEN COSPDM. ISBN 0-387-82399-9, 3-211-82399-9. ISSN 0344-8029. LCCN QA448.D38 G458 1993. Based on lectures given at the 1st Dagstuhlseminar, Dagstuhl, Germany.

Ruprecht:1993:FFD

[4218] D. Ruprecht and H. Müller. Free form deformation with scattered data interpolation methods. In Farin et al. [4375], pages 267–281. CODEN COSPDM. ISBN 0-387-82399-9, 3-211-82399-9. ISSN 0344-8029. LCCN QA448.D38 G458 1993. Based on lectures given at the 1st Dagstuhl-seminar, Dagstuhl, Germany.

Schwarz:1993:SMB

[4219] W. Schwarz. C^{1} smoothing of multipatch Bézier surfaces. In Farin et al. [4375], pages 283–289. CO-DEN COSPDM. ISBN 0-387-82399-9, 3-211-82399-9. ISSN 0344-8029. LCCN QA448.D38 G458 1993. Based on lectures given at the 1st Dagstuhlseminar, Dagstuhl, Germany.

Wassum:1993:GCB

[4220] P. Wassum. Geometric continuity between adjacent rational Bézier surface patches. In Farin et al. [4375], pages 291–316. CODEN COSPDM. ISBN 0-387-82399-9, 3-211-82399-9. ISSN 0344-8029. LCCN QA448.D38 G458 1993. Based on lectures given at the 1st Dagstuhl-seminar, Dagstuhl, Germany.

Apostolatos:1993:UCM

[4221] N. Apostolatos. On a unified concept of mathematics. In Albrecht et al. [4376], pages 1–10. CODEN COSPDM.

ISBN 0-387-82451-0 (New York), 3-211-82451-0 (Vienna). ISSN 0344-8029. LCCN QA297 .V27 1993. Dedicated to Ulrich Kulisch on the occasion of his 60th birthday.

Atanassova:1993:GAC

[4222] L. Atanassova and J. Herzberger. A general approach to a class of singlestep methods for the simultaneous inclusion of polynomial zeros. In Albrecht et al. [4376], pages 11–19. CODEN COSPDM. ISBN 0-387-82451-0 (New York), 3-211-82451-0 (Vienna). ISSN 0344-8029. LCCN QA297 .V27 1993. Dedicated to Ulrich Kulisch on the occasion of his 60th birthday.

Dimitrova:1993:SPI

[4223] N. S. Dimitrova. On some properties of an interval Newton type method and its modification. In Albrecht et al. [4376], pages 21–32. CODEN COSPDM. ISBN 0-387-82451-0 (New York), 3-211-82451-0 (Vienna). ISSN 0344-8029. LCCN QA297 .V27 1993. Dedicated to Ulrich Kulisch on the occasion of his 60th birthday.

Dobner:1993:VSI

[4224] H.-J. Dobner. Verified solution of the integral equations for the twodimensional Dirichlet and Neumann problem. In Albrecht et al. [4376], pages 33–43. CODEN COSPDM. ISBN 0-387-82451-0 (New York), 3-211-82451-0 (Vienna). ISSN 0344-8029. LCCN QA297 .V27 1993. Dedicated to Ulrich Kulisch on the occasion of his 60th birthday.

Frommer:1993:TSI

[4225] A. Frommer and G. Mayer. Two-stage interval iterative methods. In Albrecht et al. [4376], pages 45–65. CODEN COSPDM. ISBN 0-387-82451-0 (New York), 3-211-82451-0 (Vienna). ISSN 0344-8029. LCCN QA297 .V27 1993. Dedicated to Ulrich Kulisch on the occasion of his 60th birthday.

Han:1993:CAS

[4226] Weimin Han and Florian A. Potra. Convergence acceleration for some rootfinding methods. In Albrecht et al. [4376], pages 67–78. CODEN COSPDM. ISBN 0-387-82451-0 (New York), 3-211-82451-0 (Vienna). ISSN 0344-8029. LCCN QA297 .V27 1993. Dedicated to Ulrich Kulisch on the occasion of his 60th birthday.

Heindl:1993:PEC

[4227] G. Heindl. A program for enclosing cumulative binomial probabilities. In Albrecht et al. [4376], pages 79–92. CO-DEN COSPDM. ISBN 0-387-82451-0 (New York), 3-211-82451-0 (Vienna). ISSN 0344-8029. LCCN QA297 .V27 1993. Dedicated to Ulrich Kulisch on the occasion of his 60th birthday.

Jahn:1993:EEH

[4228] K.-U. Jahn. Effective evaluation of Hausdorff distances for finite unions of hyperrectangles. In Albrecht et al. [4376], pages 93–100. CODEN COSPDM. ISBN 0-387-82451-0 (New York), 3-211-82451-0 (Vienna). ISSN 0344-8029. LCCN QA297 .V27 1993. Dedicated to Ulrich Kulisch on the occasion of his 60th birthday.

Kaucher:1993:VCF

[4229] E. Kaucher and C. Baumhof. A verified computation of Fourier-representations of solutions for functional equations. In Albrecht et al. [4376], pages 101– 115. CODEN COSPDM. ISBN 0-387-82451-0 (New York), 3-211-82451-0 (Vienna). ISSN 0344-8029. LCCN QA297 .V27 1993. Dedicated to Ulrich Kulisch on the occasion of his 60th birthday.

Kearfott:1993:CPG

[4230] B. Kearfott and K. Du. The cluster problem in global optimization: The univariate case. In Albrecht et al. [4376], pages 117–127. CODEN COSPDM. ISBN 0-387-82451-0 (New York), 3-211-82451-0 (Vienna). ISSN 0344-8029. LCCN QA297 .V27 1993. Dedicated to Ulrich Kulisch on the occasion of his 60th birthday.

Konig:1993:DES

[4231] S. König and C. P. Ullrich. Developing expert systems for validating numerics. In Albrecht et al. [4376], pages 129–146. CODEN COSPDM. ISBN 0-387-82451-0 (New York), 3-211-82451-0 (Vienna). ISSN 0344-8029. LCCN QA297 .V27 1993. Dedicated to Ulrich Kulisch on the occasion of his 60th birthday.

Kramer:1993:CIB

[4232] W. Krämer and B. Barth. Computation of interval bounds for Weierstrass’ elliptic function ℘(z). In Albrecht et al. [4376], pages 147–159. CO-DEN COSPDM. ISBN 0-387-82451-0 (New York), 3-211-82451-0 (Vienna). ISSN 0344-8029. LCCN QA297 .V27 1993. Dedicated to Ulrich Kulisch on the occasion of his 60th birthday.

Nakao:1993:SNE

[4233] Mitsuhiro T. Nakao. Solving nonlinear elliptic problems with result verification using an H^{−1} type residual iteration. In Albrecht et al. [4376], pages 161–173. CODEN COSPDM. ISBN 0-387-82451-0 (New York), 3-211-82451-0 (Vienna). ISSN 0344-8029. LCCN QA297 .V27 1993. Dedicated to Ulrich Kulisch on the occasion of his 60th birthday.

Neumaier:1993:WEE

[4234] A. Neumaier. The wrapping effect, ellipsoid arithmetic, stability and confidence regions. In Albrecht et al. [4376], pages 175–190. CODEN COSPDM. ISBN 0-387-82451-0 (New York), 3-211-82451-0 (Vienna). ISSN 0344-8029. LCCN QA297 .V27 1993. Dedicated to Ulrich Kulisch on the occasion of his 60th birthday.

Rump:1993:VSL

[4235] S. M. Rump. Validated solution of large linear systems. In Albrecht et al. [4376], pages 191–212. CODEN COSPDM. ISBN 0-387-82451-0 (New York), 3-211-82451-0 (Vienna). ISSN 0344-8029. LCCN QA297 .V27 1993. Dedicated to Ulrich Kulisch on the occasion of his 60th birthday.

Schwandt:1993:IBA

[4236] H. Schwandt. The interval Buneman algorithm for arbitrary block dimension. In Albrecht et al. [4376], pages 213–231. CODEN COSPDM. ISBN 0-387-82451-0 (New York), 3-211-82451-0 (Vienna). ISSN 0344-8029. LCCN QA297 .V27 1993. Dedicated to Ulrich Kulisch on the occasion of his 60th birthday.

Spreuer:1993:EVD

[4237] H. Spreuer and E. Adams. On the existence and the verified determination of homoclinic and heteroclinic orbits of the origin for the Lorenz equations. In Albrecht et al. [4376], pages 233–246. CODEN COSPDM. ISBN 0-387-82451-0 (New York), 3-211-82451-0 (Vienna). ISSN 0344-8029. LCCN QA297 .V27 1993. Dedicated to Ulrich Kulisch on the occasion of his 60th birthday.

Stetter:1993:VCA

[4238] H. J. Stetter. Verification in computer algebra systems. In Albrecht et al. [4376], pages 247–263. CODEN COSPDM. ISBN 0-387-82451-0 (New York), 3-211-82451-0 (Vienna). ISSN 0344-8029. LCCN QA297 .V27 1993. Dedicated to Ulrich Kulisch on the occasion of his 60th birthday.

Walter:1993:FXP

[4239] W. V. Walter. FORTRAN-XSC. A portable Fortran 90 module library for accurate and reliable scientific computing. In Albrecht et al. [4376], pages 265–285. CODEN COSPDM. ISBN 0-387-82451-0 (New York), 3-211-82451-0 (Vienna). ISSN 0344-8029. LCCN QA297 .V27 1993. Dedicated to Ulrich Kulisch on the occasion of his 60th birthday.

WolffvonGudenberg:1993:IAM

[4240] Jürgen Wolff von Gudenberg. Implementation of accurate matrix multiplication on the CM-2. In Albrecht et al. [4376], pages 287–291. CODEN COSPDM. ISBN 0-387-82451-0 (New York), 3-211-82451-0 (Vienna). ISSN 0344-8029. LCCN QA297 .V27 1993. Dedicated to Ulrich Kulisch on the occasion of his 60th birthday.

Barnhill:1995:POS

[4241] R. E. Barnhill and T. M. Frost. Parametric offset surface approximation. In Hagen et al. [4377], pages 1–34. CO-DEN COSPDM. ISBN 3-211-82666-1. ISSN 0344-8029. LCCN QA565 .G45 1995. Based on lectures given at the 2nd Dagstuhl seminar on geometric modelling.

Barry:1995:UPG

[4242] P. J. Barry and R. N. Goldman. Unimodal properties of generalized ball bases. In Hagen et al. [4377], pages 35–41. CODEN COSPDM. ISBN 3-211-82666-1. ISSN 0344-8029. LCCN QA565 .G45 1995. Based on lectures given at the 2nd Dagstuhl seminar on geometric modelling.

Bieri:1995:NPB

[4243] H.-P. Bieri. Nef polyhedra: a brief introduction. In Hagen et al. [4377], pages 43–60. CODEN COSPDM. ISBN 3-211-82666-1. ISSN 0344-8029. LCCN QA565 .G45 1995. Based on lectures given at the 2nd Dagstuhl seminar on geometric modelling.

Bloor:1995:CPS

[4244] M. I. G. Bloor and M. J. Wilson.

Complex PDE surface generation for analysis and manufacture. In Hagen et al. [4377], pages 61–77. CODEN COSPDM. ISBN 3-211-82666-1. ISSN 0344-8029. LCCN QA565 .G45 1995. Based on lectures given at the 2nd Dagstuhl seminar on geometric modelling.

Bonneau:1995:WER

[4245] G.-P. Bonneau. Weight estimation of rational Bézier curves and surfaces. In Hagen et al. [4377], pages 79–86. CO-DEN COSPDM. ISBN 3-211-82666-1. ISSN 0344-8029. LCCN QA565 .G45 1995. Based on lectures given at the 2nd Dagstuhl seminar on geometric modelling.

Brown:1995:UMK

[4246] J. M. Brown, M. I. G. Bloor, M. S. Bloor, and M. J. Wilson. The use of multiple knots for B-spline finite element approximations to PDE surfaces. In Hagen et al. [4377], pages 87–99. CODEN COSPDM. ISBN 3-211-82666-1. ISSN 0344-8029. LCCN QA565 .G45 1995. Based on lectures given at the 2nd Dagstuhl seminar on geometric modelling.

Brunnett:1995:GDT

[4247] G. Brunnett. Geometric design with trimmed surfaces. In Hagen et al. [4377], pages 101–115. CODEN COSPDM. ISBN 3-211-82666-1. ISSN 0344-8029. LCCN QA565 .G45 1995. Based on lectures given at the 2nd Dagstuhl seminar on geometric modelling.

Degen:1995:SOS

[4248] W. L. F. Degen. The shape of the Overhauser spline. In Hagen et al. [4377], pages 117–128. CODEN COSPDM. ISBN 3-211-82666-1. ISSN 0344-8029. LCCN QA565 .G45 1995. Based on lectures given at the 2nd Dagstuhl seminar on geometric modelling.

Eck:1995:LEF

[4249] M. Eck and J. Hadenfeld. Local energy fairing of B-spline curves. In Hagen et al. [4377], pages 129–147. CO-DEN COSPDM. ISBN 3-211-82666-1. ISSN 0344-8029. LCCN QA565 .G45 1995. Based on lectures given at the 2nd Dagstuhl seminar on geometric modelling.

Fischer:1995:IAT

[4250] R. Fischer and A. S. Vieira. Integrating analysis tools into the design process through constrained parametric structures. In Hagen et al. [4377], pages 149–161. CODEN COSPDM. ISBN 3-211-82666-1. ISSN 0344-8029. LCCN QA565 .G45 1995. Based on lectures given at the 2nd Dagstuhl seminar on geometric modelling.

Foley:1995:LRB

[4251] T. A. Foley, S. Dayanand, and D. Zeckzer. Localized radial basis methods using rational triangle patches. In Hagen et al. [4377], pages 163–176. CODEN COSPDM. ISBN 3-211-82666-1. ISSN 0344-8029. LCCN QA565 .G45 1995. Based on lectures given at the 2nd Dagstuhl seminar on geometric modelling.

Franke:1995:RKL

[4252] R. Franke, H. Hagen, and G. M. Nielson. Repeated knots in least squares multiquadric functions. In Hagen et al. [4377], pages 177–187. CODEN COSPDM. ISBN 3-211-82666-1. ISSN 0344-8029. LCCN QA565 .G45 1995. Based on lectures given at the 2nd Dagstuhl seminar on geometric modelling.

Hagen:1995:SCS

[4253] H. Hagen and S. Hahmann. Stability concept for surfaces. In Hagen et al. [4377], pages 189–198. CODEN COSPDM. ISBN 3-211-82666-1. ISSN 0344-8029. LCCN QA565 .G45 1995. Based on lectures given at the 2nd Dagstuhl seminar on geometric modelling.

Hamann:1995:QSB

[4254] B. Hamann and T. A. Foley. A quartic spline based on a variational approach. In Hagen et al. [4377], pages 199–210. CODEN COSPDM. ISBN 3-211-82666-1. ISSN 0344-8029. LCCN QA565 .G45 1995. Based on lectures given at the 2nd Dagstuhl seminar on geometric modelling.

Langbecker:1995:KBS

[4255] U. Langbecker and H. Nowacki. A knowledge-based system for geometric design. In Hagen et al. [4377], pages 211–226. CODEN COSPDM. ISBN 3-211-82666-1. ISSN 0344-8029. LCCN QA565 .G45 1995. Based on lectures given at the 2nd Dagstuhl seminar on geometric modelling.

Lasser:1995:BRT

[4256] D. Lasser and G.-P. Bonneau. Bézier representation of trim curves. In Hagen et al. [4377], pages 227–242. CO-DEN COSPDM. ISBN 3-211-82666-1. ISSN 0344-8029. LCCN QA565 .G45 1995. Based on lectures given at the 2nd Dagstuhl seminar on geometric modelling.

Muller:1995:CPR

[4257] H. Müller and B. Gnadl. Control point representations of trigonometrically specified curves and surfaces. In Hagen et al. [4377], pages 243–251. CO-DEN COSPDM. ISBN 3-211-82666-1. ISSN 0344-8029. LCCN QA565 .G45 1995. Based on lectures given at the 2nd Dagstuhl seminar on geometric modelling.

Pratt:1995:TOA

[4258] M. J. Pratt. Towards optimality in automated feature recognition. In Hagen et al. [4377], pages 253–274. CO-DEN COSPDM. ISBN 3-211-82666-1. ISSN 0344-8029. LCCN QA565 .G45 1995. Based on lectures given at the 2nd Dagstuhl seminar on geometric modelling.

Roller:1995:SMC

[4259] D. Roller. Solid modeling with constrained form features. In Hagen et al. [4377], pages 275–284. CODEN COSPDM. ISBN 3-211-82666-1. ISSN 0344-8029. LCCN QA565 .G45 1995. Based on lectures given at the 2nd Dagstuhl seminar on geometric modelling.

Sapidis:1995:HMS

[4260] N. S. Sapidis and P. D. Kaklis. A hybrid method for shape-preserving interpolation with curvature-continuous quintic splines. In Hagen et al. [4377], pages 285–301. CODEN COSPDM. ISBN 3-211-82666-1. ISSN 0344-8029. LCCN QA565 .G45 1995. Based on lectures given at the 2nd Dagstuhl seminar on geometric modelling.

Sequin:1995:SIF

[4261] C. H. Sequin, P.-Y. Chang, and H. P.

Moreton. Scale-invariant functionals for smooth curves and surfaces. In Hagen et al. [4377], pages 303–321. CO-DEN COSPDM. ISBN 3-211-82666-1. ISSN 0344-8029. LCCN QA565 .G45 1995. Based on lectures given at the 2nd Dagstuhl seminar on geometric modelling.

Verbarg:1995:CTD

[4262] K. Verbarg. The C-tree: a dynamically balanced spatial index. In Hagen et al. [4377], pages 323–340. CO-DEN COSPDM. ISBN 3-211-82666-1. ISSN 0344-8029. LCCN QA565 .G45 1995. Based on lectures given at the 2nd Dagstuhl seminar on geometric modelling.

Vigo:1995:PLA

[4263] M. Vigo and P. Brunet. Piecewise linear approximation of trimmed surfaces. In Hagen et al. [4377], pages 341–356. CODEN COSPDM. ISBN 3-211-82666-1. ISSN 0344-8029. LCCN QA565 .G45 1995. Based on lectures given at the 2nd Dagstuhl seminar on geometric modelling.

Warren:1995:EAE

[4264] J. Warren. An efficient algorithm for evaluating polynomials in the Pólya basis. In Hagen et al. [4377], pages 357– 361. CODEN COSPDM. ISBN 3-211-82666-1. ISSN 0344-8029. LCCN QA565 .G45 1995. Based on lectures given at the 2nd Dagstuhl seminar on geometric modelling.

Daniilidis:1996:AVM

[4265] K. Daniilidis. Attentive visual motion processing: Computations in the logpolar plane. In Kropatsch et al. [4378], pages 1–20. CODEN COSPDM. ISBN 3-211-82730-7. ISSN 0344-8029. LCCN TA1634 .T43 1996. URL http://www.

springer.at/springer.py?Page=30&
Key=338&id=17&cat=3.

Eckhardt:1996:ITD

[4266] U. Eckhardt and L. Latecki. Invariant thinning and distance transform. In Kropatsch et al. [4378], pages 21– 36. CODEN COSPDM. ISBN 3-211-82730-7. ISSN 0344-8029. LCCN TA1634 .T43 1996. URL http://www.

springer.at/springer.py?Page=30&
Key=338&id=17&cat=3.

Flusser:1996:RID

[4267] Jan Flusser, Tomáš Suk, and Stanislav Saic. Recognition of images degraded by linear motion blur without restoration. In Kropatsch et al. [4378], pages 37–51. CODEN COSPDM. ISBN 3-211-82730-7. ISSN 0344-8029. LCCN TA1634 .T43 1996. URL http://www.

springer.at/springer.py?Page=30&
Key=338&id=17&cat=3.

Gimelfarb:1996:SBT

[4268] G. L. Gimel’farb. Symmetric bi- and trinocular stereo: Tradeoffs between theoretical foundations and heuristics. In Kropatsch et al. [4378], pages 53– 72. CODEN COSPDM. ISBN 3-211-82730-7. ISSN 0344-8029. LCCN TA1634 .T43 1996. URL http://www.

springer.at/springer.py?Page=30&
Key=338&id=17&cat=3.

Klette:1996:SMC

[4269] R. Klette. Surface from motion — without and with calibration. In Kropatsch et al. [4378], pages 73– 98. CODEN COSPDM. ISBN 3-211-82730-7. ISSN 0344-8029. LCCN TA1634 .T43 1996. URL http://www.

springer.at/springer.py?Page=30&
Key=338&id=17&cat=3.

Kropatsch:1996:PPR

[4270] W. G. Kropatsch. Properties of pyramidal representations. In Kropatsch et al. [4378], pages 99–112. CO-DEN COSPDM. ISBN 3-211-82730-7. ISSN 0344-8029. LCCN TA1634 .T43

1996. URL http://www.springer.
at/springer.py?Page=30&Key=338&
id=17&cat=3.

Leonardis:1996:RAE

[4271] A. Leonardis. A robust approach to estimation of parametric models. In Kropatsch et al. [4378], pages 113– 130. CODEN COSPDM. ISBN 3-211-82730-7. ISSN 0344-8029. LCCN TA1634 .T43 1996. URL http://www.

springer.at/springer.py?Page=30&
Key=338&id=17&cat=3.

Roerdink:1996:CVM

[4272] J. B. T. M. Roerdink. Computer vision and mathematical morphology. In Kropatsch et al. [4378], pages 131– 148. CODEN COSPDM. ISBN 3-211-82730-7. ISSN 0344-8029. LCCN TA1634 .T43 1996. URL http://www.

springer.at/springer.py?Page=30&
Key=338&id=17&cat=3.

Schnoerr:1996:VAD

[4273] C. Schnoerr, R. Sprengel, and B. Neumann. A variational approach to the design of early vision algorithms. In Kropatsch et al. [4378], pages 149– 166. CODEN COSPDM. ISBN 3-211-82730-7. ISSN 0344-8029. LCCN TA1634 .T43 1996. URL http://www.

springer.at/springer.py?Page=30&
Key=338&id=17&cat=3.

Skarbek:1996:BCI

[4274] W. Skarbek. Banach constructor and image compression. In Kropatsch et al. [4378], pages 167–182. CO-DEN COSPDM. ISBN 3-211-82730-7. ISSN 0344-8029. LCCN TA1634 .T43

1996. URL http://www.springer.
at/springer.py?Page=30&Key=338&
id=17&cat=3.

Sloboda:1996:PLA

[4275] F. Sloboda and B. Zat’ko. Piecewise linear approximation of planar Jordan curves and arcs: Theory and applications. In Kropatsch et al. [4378], pages 183–200. CODEN COSPDM. ISBN 3-211-82730-7. ISSN 0344-8029. LCCN TA1634 .T43 1996. URL http://www.

springer.at/springer.py?Page=30&
Key=338&id=17&cat=3.

Solina:1996:SVP

[4276] F. Solina. Segmentation with volumetric part models. In Kropatsch et al. [4378], pages 201–220. CO-DEN COSPDM. ISBN 3-211-82730-7. ISSN 0344-8029. LCCN TA1634 .T43

1996. URL http://www.springer.
at/springer.py?Page=30&Key=338&
id=17&cat=3.

Weickert:1996:TFA

[4277] J. Weickert. Theoretical foundations of anisotropic diffusion in image processing. In Kropatsch et al. [4378], pages 221–236. CODEN COSPDM. ISBN 3-211-82730-7. ISSN 0344-8029. LCCN TA1634 .T43 1996. URL http://www.

springer.at/springer.py?Page=30&
Key=338&id=17&cat=3.

Weinshall:1996:SLV

[4278] D. Weinshall, M. Werman, and N. Tishby. Stability and likelihood of views of three dimensional objects. In Kropatsch et al. [4378], pages 237– 256. CODEN COSPDM. ISBN 3-211-82730-7. ISSN 0344-8029. LCCN TA1634 .T43 1996. URL http://www.

springer.at/springer.py?Page=30&
Key=338&id=17&cat=3.

Demko:1998:GTH

[4279] D. Demko. Generalization of two hypergraphs. algorithm of calculation of the greatest sub-hypergraph common to two hypergraphs annotated by semantic information. In Jolion and Kropatsch [4379], pages 1–10. CO-DEN COSPDM. ISBN 3-211-83121-5. ISSN 0344-8029. LCCN TA1650 .G73

1998. URL http://www.springer.
at/springer.py?Page=30&Key=339&
id=518&cat=3.

Englert:1998:RPP

[4280] R. Englert, A. B. Cremers, and J. Seelmann-Eggebert. Recognition of polymorphic patterns in parameterized graphs for 3D building reconstruction. In Jolion and Kropatsch [4379], pages 11–20. CODEN COSPDM. ISBN 3-211-83121-5. ISSN 0344-8029. LCCN TA1650 .G73 1998. URL http://www.

springer.at/springer.py?Page=30&
Key=339&id=518&cat=3.

Tupin:1998:GBR

[4281] F. Tupin, J.-F. Mangin, E. Pechersky, and J.-M. Nicolas. A graph-based representation to detect linear features. In Jolion and Kropatsch [4379], pages 21–32. CODEN COSPDM. ISBN 3-211-83121-5. ISSN 0344-8029. LCCN TA1650 .G73 1998. URL http://www.

springer.at/springer.py?Page=30&
Key=339&id=518&cat=3.

Salotti:1998:EDF

[4282] M. Salotti. Edge detection as finding the minimum cost path in a graph. In Jolion and Kropatsch [4379], pages 33–42. CODEN COSPDM. ISBN 3-211-83121-5. ISSN 0344-8029. LCCN TA1650 .G73 1998. URL http://www.

springer.at/springer.py?Page=30&
Key=339&id=518&cat=3.

Cordella:1998:STI

[4283] L. P. Cordella, P. Foggia, C. Sansone, and M. Vento. Subgraph transformations for the inexact matching of attributed relational graphs. In Jolion and Kropatsch [4379], pages 43– 52. CODEN COSPDM. ISBN 3-211-83121-5. ISSN 0344-8029. LCCN TA1650 .G73 1998. URL http://www.

springer.at/springer.py?Page=30&
Key=339&id=518&cat=3.

Shearer:1998:EGM

[4284] K. Shearer, H. Bunke, S. Venkatesh, and D. Kieronska. Efficient graph matching for video indexing. In Jolion and Kropatsch [4379], pages 53– 62. CODEN COSPDM. ISBN 3-211-83121-5. ISSN 0344-8029. LCCN TA1650 .G73 1998. URL http://www.

springer.at/springer.py?Page=30&
Key=339&id=518&cat=3.

Wendling:1998:IBS

[4285] L. Wendling and J. Desachy. Isomorphism between strong fuzzy relational graphs based on k-formulae. In Jolion and Kropatsch [4379], pages 63– 72. CODEN COSPDM. ISBN 3-211-83121-5. ISSN 0344-8029. LCCN TA1650 .G73 1998. URL http://www.

                                        springer.at/springer.py?Page=30&
                                        Key=339&id=518&cat=3.

Jahn:1998:GSG

[4286] H. Jahn. A graph structure for grey value and texture segmentation. In Jolion and Kropatsch [4379], pages 73– 82. CODEN COSPDM. ISBN 3-211-83121-5. ISSN 0344-8029. LCCN TA1650 .G73 1998. URL http://www.

                                        springer.at/springer.py?Page=30&
                                        Key=339&id=518&cat=3.

Brun:1998:DMF

[4287] L. Brun, J.-P. Domenger, and J.-P. Braquelaire. Discrete maps: a framework for region segmentation algorithms. In Jolion and Kropatsch [4379], pages 83–92. CODEN COSPDM. ISBN 3-211-83121-5. ISSN 0344-8029. LCCN TA1650 .G73 1998. URL http://www.

                                        springer.at/springer.py?Page=30&
                                        Key=339&id=518&cat=3.

Bertolino:1998:ISS

[4288] P. Bertolino and S. Ribas. Image sequence segmentation by a single evolutionary graph pyramid. In Jolion and Kropatsch [4379], pages 93– 100. CODEN COSPDM. ISBN 3-211-83121-5. ISSN 0344-8029. LCCN TA1650 .G73 1998. URL http://www.

                                        springer.at/springer.py?Page=30&
                                        Key=339&id=518&cat=3.

Kropatsch:1998:DGC

[4289] W. G. Kropatsch, M. Burge, S. Ben Yacoub, and N. Selmaoui. Dual graph contraction with LEDA. In Jolion and Kropatsch [4379], pages 101–110. CO-DEN COSPDM. ISBN 3-211-83121-5.

ISSN 0344-8029. LCCN TA1650 .G73

1998. URL http://www.springer.
at/springer.py?Page=30&Key=339&
id=518&cat=3.

Ducourthial:1998:IIA

[4290] B. Ducourthial, G. Constantinescu, and A. Merigot. Implementing image analysis with a graph-based parallel computing model. In Jolion and Kropatsch [4379], pages 111–122. CO-DEN COSPDM. ISBN 3-211-83121-5. ISSN 0344-8029. LCCN TA1650 .G73

1998. URL http://www.springer.
at/springer.py?Page=30&Key=339&
id=518&cat=3.

Pailloncy:1998:FRG

[4291] J.-G. Pailloncy and J.-M. Jolion. The frontier-region graph. In Jolion and Kropatsch [4379], pages 123–134. CO-DEN COSPDM. ISBN 3-211-83121-5. ISSN 0344-8029. LCCN TA1650 .G73

1998. URL http://www.springer.
at/springer.py?Page=30&Key=339&
id=518&cat=3.

Mottl:1998:OTP

[4292] V. V. Mottl, A. B. Blinov, A. V. Kopylov, and A. Kostin. A. Optimization techniques on pixel neighborhood graphs for image processing. In Jolion and Kropatsch [4379], pages 135– 145. CODEN COSPDM. ISBN 3-211-83121-5. ISSN 0344-8029. LCCN TA1650 .G73 1998. URL http://www.

springer.at/springer.py?Page=30&
Key=339&id=518&cat=3.

Albrecht:1998:AFR

[4293] G. Albrecht. Applications of a fruitful relation between a Dupin cyclide and a right circular cone. In Farin et al. [4380], pages 1–16. CODEN COSPDM.

ISBN 3-211-83207-6. ISSN 0344-8029. LCCN QA565 .G452 1998. URL

                                        http://www.springer.at/springer.
                                        py?Page=30&Key=340&id=589&cat=3.

Bajaj:1998:IV

[4294] C. L. Bajaj. Interrogative visualization. In Farin et al. [4380], pages 17–26. CODEN COSPDM. ISBN 3-211-83207-6. ISSN 0344-8029. LCCN QA565 .G452 1998. URL http://www.

                                        springer.at/springer.py?Page=30&
                                        Key=340&id=589&cat=3.

Bieri:1998:RCN

[4295] H. Bieri. Representation conversions for Nef polyhedra. In Farin et al. [4380], pages 27–38. CODEN COSPDM. ISBN 3-211-83207-6. ISSN 0344-8029. LCCN QA565 .G452 1998. URL http://www.

                                        springer.at/springer.py?Page=30&
                                        Key=340&id=589&cat=3.

Bloor:1998:SPA

[4296] M. I. G. Bloor and M. J. Wilson. The shape parametrisation of an aircraft engine nacelle and pylon. In Farin et al. [4380], pages 39–50. CO-DEN COSPDM. ISBN 3-211-83207-6. ISSN 0344-8029. LCCN QA565 .G452

                                        1998. URL http://www.springer.
                                        at/springer.py?Page=30&Key=340&
                                        id=589&cat=3.

Bonneau:1998:MAN

[4297] G. P. Bonneau. Multiresolution analysis with non-nested spaces. In Farin et al. [4380], pages 51–66. CO-DEN COSPDM. ISBN 3-211-83207-6. ISSN 0344-8029. LCCN QA565 .G452

                                        1998. URL http://www.springer.
                                        at/springer.py?Page=30&Key=340&
                                        id=589&cat=3.

Brunnett:1998:ODR

[4298] G. Brunnett and T. Schreiber. Optimal degree reduction of free form curves. In Farin et al. [4380], pages 67–76. CODEN COSPDM. ISBN 3-211-83207-6. ISSN 0344-8029. LCCN QA565 .G452 1998. URL http://www.

springer.at/springer.py?Page=30&
Key=340&id=589&cat=3.

DeCasteljau:1998:IMC

[4299] P. De Casteljau. Intersection methods of convergence. In Farin et al. [4380], pages 77–80. CODEN COSPDM. ISBN 3-211-83207-6. ISSN 0344-8029. LCCN QA565 .G452 1998. URL http://www.

springer.at/springer.py?Page=30&
Key=340&id=589&cat=3.

Degen:1998:ACM

[4300] W. L. F. Degen. Approximation of curves by a measure of shape. In Farin et al. [4380], pages 81–100. CO-DEN COSPDM. ISBN 3-211-83207-6. ISSN 0344-8029. LCCN QA565 .G452

1998. URL http://www.springer.
at/springer.py?Page=30&Key=340&
id=589&cat=3.

Ge:1998:GDR

[4301] Q. J. Ge and B. Ravani. Geometric design of rational Bézier line congruences and ruled surfaces using line geometry. In Farin et al. [4380], pages 101–120. CODEN COSPDM. ISBN 3-211-83207-6. ISSN 0344-8029. LCCN QA565 .G452 1998. URL http://www.

springer.at/springer.py?Page=30&
Key=340&id=589&cat=3.

Hagen:1998:VDP

[4302] H. Hagen and A. Nawotki. Variational design and parameter optimized surface fitting. In Farin et al. [4380], pages 121–134. CODEN COSPDM. ISBN 3-211-83207-6. ISSN 0344-8029. LCCN QA565 .G452 1998. URL http://www.

springer.at/springer.py?Page=30&
Key=340&id=589&cat=3.

Hahmann:1998:SIS

[4303] S. Hahmann. Shape improvement of surfaces. In Farin et al. [4380], pages 135–152. CODEN COSPDM. ISBN 3-211-83207-6. ISSN 0344-8029. LCCN QA565 .G452 1998. URL http://www.

springer.at/springer.py?Page=30&
Key=340&id=589&cat=3.

Johnstone:1998:PSC

[4304] J. K. Johnstone and K. R. Sloan. A philosophy for smooth contour reconstruction. In Farin et al. [4380], pages 153–164. CODEN COSPDM. ISBN 3-211-83207-6. ISSN 0344-8029. LCCN QA565 .G452 1998. URL http://www.

springer.at/springer.py?Page=30&
Key=340&id=589&cat=3.

Kadi:1998:DPA

[4305] Z. Kadi and A. Rockwood. The design of physically accurate fluid flow. In Farin et al. [4380], pages 165– 176. CODEN COSPDM. ISBN 3-211-83207-6. ISSN 0344-8029. LCCN QA565 .G452 1998. URL http://www.

springer.at/springer.py?Page=30&
Key=340&id=589&cat=3.

Kaklis:1998:QPM

[4306] P. D. Kaklis and G. D. Koras. A quadratic-programming method for removing shape-failures from tensorproduct B-spline surfaces. In Farin et al. [4380], pages 177–188. CO-DEN COSPDM. ISBN 3-211-83207-6.

ISSN 0344-8029. LCCN QA565 .G452

1998. URL http://www.springer.
at/springer.py?Page=30&Key=340&
id=589&cat=3.

Labenski:1998:FR

[4307] C. Labenski and B. Piper. Fill-in regions. In Farin et al. [4380], pages 189–210. CODEN COSPDM. ISBN 3-211-83207-6. ISSN 0344-8029. LCCN QA565 .G452 1998. URL http://www.

springer.at/springer.py?Page=30&
Key=340&id=589&cat=3.

Peters:1998:SPU

[4308] J. Peters. Smoothing polyhedra using trimmed bicubic patches. In Farin et al. [4380], pages 211–216. CO-DEN COSPDM. ISBN 3-211-83207-6. ISSN 0344-8029. LCCN QA565 .G452

1998. URL http://www.springer.
at/springer.py?Page=30&Key=340&
id=589&cat=3.

Prautzsch:1998:SA

[4309] H. Prautzsch and G. Umlauf. A G^{2}-subdivision algorithm. In Farin et al. [4380], pages 217–224. CO-DEN COSPDM. ISBN 3-211-83207-6. ISSN 0344-8029. LCCN QA565 .G452

1998. URL http://www.springer.
at/springer.py?Page=30&Key=340&
id=589&cat=3.

Vinacua:1998:OMS

[4310] A. Vinacua, I. Navazo, and P. Brunet.

Octtrees meet splines. In Farin et al. [4380], pages 225–234. CO-DEN COSPDM. ISBN 3-211-83207-6. ISSN 0344-8029. LCCN QA565 .G452

1998. URL http://www.springer.
at/springer.py?Page=30&Key=340&
id=589&cat=3.

Wolters:1998:DRB

[4311] H. J. Wolters, G. Wu, and G. Farin. Degree reduction of B-spline curves. In Farin et al. [4380], pages 235– 241. CODEN COSPDM. ISBN 3-211-83207-6. ISSN 0344-8029. LCCN QA565 .G452 1998. URL http://www.

                                        springer.at/springer.py?Page=30&
                                        Key=340&id=589&cat=3.

Aguilera:2001:COP

[4312] A. Aguilera and D. Ayala. Converting orthogonal polyhedra from extreme vertices model to B-rep and to alternating sum of volumes. In Brunnett et al. [4381], pages 1–18. ISBN 3-211-83603-9. ISSN 0344-8029. LCCN QA565 .G4523 2001. Papers from the 4th Dagstuhl Seminar on Geometric Modelling held May 1999.

Bajaj:2001:SSC

[4313] C. L. Bajaj and G. Xu. Smooth shell construction with mixed prism fat surfaces. In Brunnett et al. [4381], pages 19–36. ISBN 3-211-83603-9. ISSN 0344-8029. LCCN QA565 .G4523 2001. Papers from the 4th Dagstuhl Seminar on Geometric Modelling held May 1999.

Brunnett:2001:GMP

[4314] G. Brunnett. Geometric modeling of parallel curves on surfaces. In Brunnett et al. [4381], pages 37–54. ISBN 3-211-83603-9. ISSN 0344-8029. LCCN QA565 .G4523 2001. Papers from the 4th Dagstuhl Seminar on Geometric Modelling held May 1999.

Davies:2001:CVP

[4315] T. J. G. Davies, R. R. Martin, and A. Bowyer. Computing volume properties using low-discrepancy sequences. In Brunnett et al. [4381], pages 55–72. ISBN 3-211-83603-9. ISSN 0344-8029. LCCN QA565 .G4523 2001. Papers from the 4th Dagstuhl Seminar on Geometric Modelling held May 1999.

Elber:2001:BAS

[4316] G. Elber, G. Barequet, and M. S.

Kim. Bisectors and alpha-sectors of rational varieties. In Brunnett et al. [4381], pages 73–88. ISBN 3-211-83603-9. ISSN 0344-8029. LCCN QA565 .G4523 2001. Papers from the 4th Dagstuhl Seminar on Geometric Modelling held May 1999.

Floater:2001:PLW

[4317] M. S. Floater and E. G. Quak. Piecewise linear wavelets over type-2 triangulations. In Brunnett et al. [4381], pages 89–104. ISBN 3-211-83603-9. ISSN 0344-8029. LCCN QA565 .G4523 2001. Papers from the 4th Dagstuhl Seminar on Geometric Modelling held May 1999.

Frohlich:2001:FBM

[4318] M. Frohlich, H. Muller, C. Pillokat, and F. Weller. Feature-based matching of triangular meshes. In Brunnett et al. [4381], pages 105–118. ISBN 3-211-83603-9. ISSN 0344-8029. LCCN QA565 .G4523 2001. Papers from the 4th Dagstuhl Seminar on Geometric Modelling held May 1999.

Gabrielides:2001:ISP

[4319] N. C. Gabrielides and P. D. Kaklis. C_{4} interpolatory shape-preserving polynomial splines of variable degree. In Brunnett et al. [4381], pages 119–154. ISBN 3-211-83603-9. ISSN 0344-8029. LCCN QA565 .G4523 2001. Papers from the 4th Dagstuhl Seminar on Geometric Modelling held May 1999.

Goldman:2001:BDD

[4320] R. Goldman. Blossoming and divided difference. In Brunnett et al. [4381], pages 155–184. ISBN 3-211-83603-9. ISSN 0344-8029. LCCN QA565 .G4523 2001. Papers from the 4th Dagstuhl Seminar on Geometric Modelling held May 1999.

Hahmann:2001:LSM

[4321] S. Hahmann, G.-P. Bonneau, and R. Taleb. Localizing the 4-split method for G_{1} free-form surface fitting. In Brunnett et al. [4381], pages 185–198. ISBN 3-211-83603-9. ISSN 0344-8029. LCCN QA565 .G4523 2001. Papers from the 4th Dagstuhl Seminar on Geometric Modelling held May 1999.

Heckel:2001:SRU

[4322] B. Heckel, A. E. Uva, B. Hamann, and K. I. Joy. Surface reconstruction using adaptive clustering methods. In Brunnett et al. [4381], pages 199–218. ISBN 3-211-83603-9. ISSN 0344-8029. LCCN QA565 .G4523 2001. Papers from the 4th Dagstuhl Seminar on Geometric Modelling held May 1999.

Kos:2001:ATS

[4323] G. Kos. An algorithm to triangulate surfaces in 3D using unorganised point clouds. In Brunnett et al. [4381], pages 219–232. ISBN 3-211-83603-9. ISSN 0344-8029. LCCN QA565 .G4523 2001. Papers from the 4th Dagstuhl Seminar on Geometric Modelling held May 1999.

Mann:2001:CSP

[4324] S. Mann and T. Yeung. Cylindrical surface pasting. In Brunnett et al. [4381], pages 233–248. ISBN 3-211-83603-9. ISSN 0344-8029. LCCN QA565 .G4523 2001. Papers from the 4th Dagstuhl Seminar on Geometric Modelling held May 1999.

Michalik:2001:CBM

[4325] P. Michalik and B. Bruderlin. A constraint-based method for sculpting free-form surfaces. In Brunnett et al. [4381], pages 249–266. ISBN 3-211-83603-9. ISSN 0344-8029. LCCN QA565 .G4523 2001. Papers from the 4th Dagstuhl Seminar on Geometric Modelling held May 1999.

Milbrandt:2001:GMA

[4326] V. Milbrandt. A geometrically motivated affine invariant norm. In Brunnett et al. [4381], pages 267–280. ISBN 3-211-83603-9. ISSN 0344-8029. LCCN QA565 .G4523 2001. Papers from the 4th Dagstuhl Seminar on Geometric Modelling held May 1999.

Nawotki:2001:EWC

[4327] A. Nawotki. Exploiting wavelet coefficients for modifying functions. In Brunnett et al. [4381], pages 281–292. ISBN 3-211-83603-9. ISSN 0344-8029. LCCN QA565 .G4523 2001. Papers from the 4th Dagstuhl Seminar on Geometric Modelling held May 1999.

Robinson:2001:PRC

[4328] M. Robinson, M. I. G. Bloor, and M. J. Wilson. Parametric representation of complex mechanical parts using PDE surface generation. In Brunnett et al. [4381], pages 293–308. ISBN 3-211-83603-9. ISSN 0344-8029. LCCN QA565 .G4523 2001. Papers from the 4th Dagstuhl Seminar on Geometric Modelling held May 1999.

Schatzl:2001:DDT

[4329] R. Schatzl, H. Hagen, J. C. Barnes, B. Hamann, and K. I. Joy. Datadependent triangulation in the plane with adaptive knot placement. In Brunnett et al. [4381], pages 309–322. ISBN 3-211-83603-9. ISSN 0344-8029. LCCN QA565 .G4523 2001. Papers from the 4th Dagstuhl Seminar on Geometric Modelling held May 1999.

Varady:2001:ISR

[4330] T. Varady, P. Benko, G. Kos, and A. Rockwood. Implicit surfaces revisited. I. Patches. In Brunnett et al. [4381], pages 323–336. ISBN 3-211-83603-9. ISSN 0344-8029. LCCN QA565 .G4523 2001. Papers from the 4th Dagstuhl Seminar on Geometric Modelling held May 1999.

Warren:2001:RBF

[4331] J. Warren and H. Weimer. Radial basis functions, discrete differences, and bellshaped bases. In Brunnett et al. [4381], pages 337–351. ISBN 3-211-83603-9. ISSN 0344-8029. LCCN QA565 .G4523 2001. Papers from the 4th Dagstuhl Seminar on Geometric Modelling held May 1999.

Abels:2001:UAB

[4332] A. Abels and J. Herzberger. A unified approach for bounding the positive root of certain classes of polynomials with applications. In Alefeld and Chen [4382], page ?? ISBN 3-211-83673-X.

ISSN 0344-8029. LCCN QA297 .T66 2001.

Agarwal:2001:NVS

[4333] R. P. Agarwal and C. S. Ryoo. Numerical verifications of solutions for obstacle problems. In Alefeld and Chen [4382], page ?? ISBN 3-211-83673-X. ISSN 0344-8029. LCCN QA297 .T66 2001.

Alefeld:2001:ETK

[4334] G. E. Alefeld, F. A. Potra, and Z. Shen.

On the existence theorems of Kantorovich, Moore and Miranda. In Alefeld and Chen [4382], page ?? ISBN 3-211-83673-X. ISSN 0344-8029. LCCN QA297 .T66 2001.

Axelsson:2001:SRP

[4335] O. Axelsson. A survey of robust preconditioning methods. In Alefeld and Chen [4382], page ?? ISBN 3-211-83673-X. ISSN 0344-8029. LCCN QA297 .T66 2001.

Birgin:2001:BCO

[4336] E. G. Birgin and J. M. Martínez.

A box-constrained optimization algorithm with negative curvature directions and spectral projected gradients. In Alefeld and Chen [4382], page ?? ISBN 3-211-83673-X. ISSN 0344-8029. LCCN QA297 .T66 2001.

Breuer:2001:IEP

[4337] B. Breuer, M. Plum, and P. J.

McKenna. Inclusions and existence proofs for solutions of a nonlinear boundary value problem by spectral numerical methods. In Alefeld and Chen [4382], page ?? ISBN 3-211-83673-X. ISSN 0344-8029. LCCN QA297 .T66 2001.

Chen:2001:SGC

[4338] X. Chen. A superlinearly and globally convergent method for reaction and diffusion problems with a non-Lipschitzian operator. In Alefeld and Chen [4382], page ?? ISBN 3-211-83673-X. ISSN 0344-8029. LCCN QA297 .T66 2001.

Frommer:2001:LAI

[4339] A. Frommer and P. Spiteri. On linear asynchronous iterations when the spectral radius of the modulus matrix is one. In Alefeld and Chen [4382], page ?? ISBN 3-211-83673-X. ISSN 0344-8029. LCCN QA297 .T66 2001.

Ishihara:2001:IME

[4340] K. Ishihara. Iterative methods for eigenvalue problems with nondifferentiable normalized condition of a general complex matrix. In Alefeld and Chen [4382], page ?? ISBN 3-211-83673-X. ISSN 0344-8029. LCCN QA297 .T66 2001.

Liu:2001:GOQ

[4341] Y. Liu, K. L. Teo, and S. Ito. Global optimization in quadratic semi-infinite programming. In Alefeld and Chen [4382], page ?? ISBN 3-211-83673-X. ISSN 0344-8029. LCCN QA297 .T66 2001.

Marek:2001:ADM

[4342] I. Marek and P. Mayer. Aggregation/disaggregation methods for pcyclic Markov chains. In Alefeld and Chen [4382], page ?? ISBN 3-211-83673-X. ISSN 0344-8029. LCCN QA297 .T66 2001.

Mayer:2001:NWD

[4343] G. Mayer. A new way to describe the symmetric solution set S of linear interval systems. In Alefeld and Chen [4382], page ?? ISBN 3-211-83673-X. ISSN 0344-8029. LCCN QA297 .T66 2001.

Nakao:2001:GBO

[4344] M. T. Nakao and N. Yamamoto. A guaranteed bound of the optimal constant in the error estimates for linear triangular element. In Alefeld and Chen [4382], page ?? ISBN 3-211-83673-X. ISSN 0344-8029. LCCN QA297 .T66 2001.

Ogita:2001:FVS

[4345] T. Ogita, S. Oishi, and Y. Ushiro.

Fast verification of solutions for sparse monotone matrix equations. In Alefeld and Chen [4382], page ?? ISBN 3-211-83673-X. ISSN 0344-8029. LCCN QA297 .T66 2001.

Petkovic:2001:LLM

[4346] M. S. Petkovič, L. Petkovič, and D. Živkovič. Laguerre-like methods for the simultaneous approximation of polynomial zeros. In Alefeld and Chen [4382], page ?? ISBN 3-211-83673-X. ISSN 0344-8029. LCCN QA297 .T66 2001.

Qi:2001:SNM

[4347] L. Qi and G. Zhou. A smoothing Newton method for ball constrained variational inequalities with applications. In Alefeld and Chen [4382], page ?? ISBN 3-211-83673-X. ISSN 0344-8029. LCCN QA297 .T66 2001.

Yamashita:2001:RCL

[4348] N. Yamashita and M. Fukushima. On the rate of convergence of the Levenberg–Marquardt method. In Alefeld and Chen [4382], page ?? ISBN 3-211-83673-X. ISSN 0344-8029. LCCN QA297 .T66 2001.

Alefeld:2001:SSS

[4349] G. Alefeld, V. Kreinovich, and G. Mayer. On symmetric solution sets. In Herzberger [4383], pages 1–22. ISBN 3-211-83852-X. ISSN 0344-8029. LCCN QA427 .I53 2003. EUR$65. Based on papers presented at the International GAMM Workshop held in Munich, Germany, Dec. 15-18, 2000. This workshop was organized with the support of GAMM, the International Association of Applied Mathematics and Mechanics, on the occasion of J. Herzberger’s 60th birthday.

Atanassova:2001:MCA

[4350] L. Atanassova, N. Kyurkchiev, and T. Yamamoto. Methods for computing all roots of a polynomial simultaneously — known results and open problems. In Herzberger [4383], pages 23–42. ISBN 3-211-83852-X. ISSN 0344-8029. LCCN QA427 .I53 2003. EUR$65. Based on papers presented at the International GAMM Workshop held in Munich, Germany, Dec. 15-18, 2000. This workshop was organized with the support of GAMM, the International Association of Applied Mathematics and Mechanics, on the occasion of J. Herzberger’s 60th birthday.

Detmers:2001:NBE

[4351] F. Detmers and J. Herzberger. Narrow bounds for the effective rate of return concerning a special problem for annuities. In Herzberger [4383], pages 43–44. ISBN 3-211-83852-X. ISSN 0344-8029. LCCN QA427 .I53 2003. EUR$65. Based on papers presented at the International GAMM Workshop held in Munich, Germany, Dec. 15-18, 2000. This workshop was organized with the support of GAMM, the International Association of Applied Mathematics and Mechanics, on the occasion of J. Herzberger’s 60th birthday.

Fischer:2001:ADI

[4352] H. Fischer. Algorithmic differentiation with intervals. In Herzberger [4383], pages 45–66. ISBN 3-211-83852-X. ISSN 0344-8029. LCCN QA427 .I53 2003. EUR$65. Based on papers presented at the International GAMM Workshop held in Munich, Germany, Dec. 15-18, 2000. This workshop was organized with the support of GAMM, the International Association of Applied Mathematics and Mechanics, on the occasion of J. Herzberger’s 60th birthday.

Frischmuth:2001:CFN

[4353] K. Frischmuth and V. G. Tsybulin. Computation of a family of noncosymmetrical equilibria in a system of nonlinear parabolic equations. In Herzberger [4383], pages 67–82. ISBN 3-211-83852-X. ISSN 0344-8029. LCCN QA427 .I53 2003. EUR$65. Based on papers presented at the International GAMM Workshop held in Munich, Germany, Dec. 15-18, 2000. This workshop was organized with the support of GAMM, the International Association of Applied Mathematics and Mechanics, on the occasion of J. Herzberger’s 60th birthday.

Hari:2001:QCS

[4354] V. Hari and J. Matejas. Quadratic convergence of scaled iterates by Kogbetliantz method. In Herzberger [4383], pages 83–106. ISBN 3-211-83852-X. ISSN 0344-8029. LCCN QA427 .I53 2003. EUR$65. Based on papers presented at the International GAMM Workshop held in Munich, Germany, Dec. 15-18, 2000. This workshop was organized with the support of GAMM, the International Association of Applied Mathematics and Mechanics, on the occasion of J. Herzberger’s 60th birthday.

Heindl:2001:MCI

[4355] G. Heindl. On a method for computing inclusions of solutions of the basic GPS equations. In Herzberger [4383], pages 107–120. ISBN 3-211-83852-X. ISSN 0344-8029. LCCN QA427 .I53 2003. EUR$65. Based on papers presented at the International GAMM Workshop held in Munich, Germany, Dec. 15-18, 2000. This workshop was organized with the support of GAMM, the International Association of Applied Mathematics and Mechanics, on the occasion of J. Herzberger’s 60th birthday.

Herzberger:2001:CBP

[4356] J. Herzberger and A. Langer. Construction of bounds for the positive root of a general class of polynomials with applications. In Herzberger [4383], pages 121–134. ISBN 3-211-83852-X. ISSN 0344-8029. LCCN QA427 .I53 2003. EUR$65. Based on papers presented at the International GAMM Workshop held in Munich, Germany, Dec. 15-18, 2000. This workshop was organized with the support of GAMM, the International Association of Applied Mathematics and Mechanics, on the occasion of J. Herzberger’s 60th birthday.

Kulisch:2001:RNZ

[4357] U. Kulisch. Rounding near zero.

In Herzberger [4383], pages 135–142. ISBN 3-211-83852-X. ISSN 0344-8029. LCCN QA427 .I53 2003. EUR$65. Based on papers presented at the International GAMM Workshop held in Munich, Germany, Dec. 15-18, 2000. This workshop was organized with the support of GAMM, the International Association of Applied Mathematics and Mechanics, on the occasion of J. Herzberger’s 60th birthday.

Kyurkchiev:2001:NCS

[4358] N. Kyurkchiev. A note on the convergence of the SOR-like Weierstrass method. In Herzberger [4383], pages 143–150. ISBN 3-211-83852-X. ISSN 0344-8029. LCCN QA427 .I53 2003. EUR$65. Based on papers presented at the International GAMM Workshop held in Munich, Germany, Dec. 15-18, 2000. This workshop was organized with the support of GAMM, the International Association of Applied Mathematics and Mechanics, on the occasion of J. Herzberger’s 60th birthday.

Langemann:2001:BRA

[4359] D. Langemann. Boundary regularity aspects in solving contact problems. In Herzberger [4383], pages 151–164. ISBN 3-211-83852-X. ISSN 0344-8029. LCCN QA427 .I53 2003. EUR$65.

Based on papers presented at the International GAMM Workshop held in Munich, Germany, Dec. 15-18, 2000. This workshop was organized with the support of GAMM, the International Association of Applied Mathematics and Mechanics, on the occasion of J. Herzberger’s 60th birthday.

Ling:2001:CDO

[4360] Y. x. Ling, H. l. Cao, and H. f. Sheng. The convex-decomposable operator and inclusive algorithm. In Herzberger [4383], pages 165–170. ISBN 3-211-83852-X. ISSN 0344-8029. LCCN QA427 .I53 2003. EUR$65. Based on papers presented at the International GAMM Workshop held in Munich, Germany, Dec. 15-18, 2000. This workshop was organized with the support of GAMM, the International Association of Applied Mathematics and Mechanics, on the occasion of J. Herzberger’s 60th birthday.

Ogita:2001:FIR

[4361] T. Ogita, S. Oishi, and Y. Ushiro. Fast inclusion and residual iteration for solutions of matrix equations. In Herzberger [4383], pages 171–184. ISBN 3-211-83852-X. ISSN 0344-8029. LCCN QA427 .I53 2003. EUR$65. Based on papers presented at the International GAMM Workshop held in Munich, Germany, Dec. 15-18, 2000. This workshop was organized with the support of GAMM, the International Association of Applied Mathematics and Mechanics, on the occasion of J. Herzberger’s 60th birthday.

Petkovic:2001:SLM

[4362] L. D. Petkovic and M. S. Petkovic.

Schroder-like methods for the simultaneous inclusion of polynomial zeros. In Herzberger [4383], pages 185–198. ISBN 3-211-83852-X. ISSN 0344-8029. LCCN QA427 .I53 2003. EUR$65. Based on papers presented at the International GAMM Workshop held in Munich, Germany, Dec. 15-18, 2000. This workshop was organized with the support of GAMM, the International Association of Applied Mathematics and Mechanics, on the occasion of J. Herzberger’s 60th birthday.

Petkovic:2001:IRF

[4363] L. D. Petkovic, M. S. Petkovic, and D. Zivkovic. Interval rootfinding methods of Laguerre’s type. In Herzberger [4383], pages 199–212. ISBN 3-211-83852-X. ISSN 0344-8029. LCCN QA427 .I53 2003. EUR$65. Based on papers presented at the International GAMM Workshop held in Munich, Germany, Dec. 15-18, 2000. This workshop was organized with the support of GAMM, the International Association of Applied Mathematics and Mechanics, on the occasion of J. Herzberger’s 60th birthday.

Popivanov:2001:EBS

[4364] N. Popivanov and T. Popov. Exact behaviour of singularities of Protter’s problem for the 3-D wave equation. In Herzberger [4383], pages 213–236. ISBN 3-211-83852-X. ISSN 0344-8029. LCCN QA427 .I53 2003. EUR$65. Based on papers presented at the International GAMM Workshop held in Munich, Germany, Dec. 15-18, 2000. This workshop was organized with the support of GAMM, the International Association of Applied Mathematics and Mechanics, on the occasion of J. Herzberger’s 60th birthday.

Smelyakov:2001:CSL

[4365] S. V. Smelyakov. Construction of shortest line of restricted curvature in a non-singly-connected polygonal area. In Herzberger [4383], pages 237–244. ISBN 3-211-83852-X. ISSN 0344-8029. LCCN QA427 .I53 2003. EUR$65. Based on papers presented at the International GAMM Workshop held in Munich, Germany, Dec. 15-18, 2000. This workshop was organized with the support of GAMM, the International Association of Applied Mathematics and Mechanics, on the occasion of J. Herzberger’s 60th birthday.

Kulisch:1977:GCA

[4366] Rudolf Albrecht und Ulrich Kulisch, editor. Grundlagen der Computer-Arithmetik. (German) [Foundations of Computer Arithmetic], volume 1 of Computing. Supplementum. Springer, Wien / New York, 1977. ISBN 0-387-81410-8. LCCN QA 162 G924 1975. Diese Artikel stellen eine Auswahl von Vortragen dar, die auf einer vom 4. bis 8. August 1975 im ‘Mathematischen Forschungsinstitut Oberwolfach’ stattgefundenen Tagung gehalten wurden.

Schmidt:1978:KVZ

[4367] J. W. Schmidt and W. Nauber. Korrectur: Über Über Verfahren zur zweiseitigen Approximation inverser Elemente. (German) [correction: On methods for the two-sided approximation of inverse elements]. Numerische Mathematik, 31(3):320, September 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [956].

Alefeld:1980:FNC

[4368] G. Alefeld and R. D. Grigorieff, editors. Fundamentals of numerical computation (computer-oriented numerical analysis), volume 2 of Computing. Supplementum. Springer, Wien / New York, 1980. CODEN COSPDM. ISBN 0-387-81566-X. ISSN 0344-8029. LCCN QA297 .F84. In cooperation with R. Albrecht, U. Kulisch, and F. Stummel. “Mainly a collection of the invited lectures which were given during a conference … held in June 5–8, 1979, on the occasion of the centennial of the Technical University of Berlin.”.

Knodel:1981:PPR

[4369] Walter Knödel and Hans Jürgen Schneider, editors. Parallel processes and related automata = Parallele Prozesse und damit zusammenhängende Automaten, volume 3 of Computing. Supplementum. Springer, Wien / New York, 1981. CODEN COSPDM. ISBN 0-387-81606-2. ISSN 0344-8029. LCCN QA76.6 .P353. Essays in honor of W. Handler.

Buchberger:1982:CAS

[4370] B. (Bruno) Buchberger, G. E.

(George Edwin) Collins, and R. (Rudiger) Loos, editors. Computer algebra: Symbolic and algebraic computation, volume 4 of Computing. Supplementum. Springer, Wien / New York, 1982. CODEN COSPDM. ISBN 0-387-81684-4. ISSN 0344-8029. LCCN QA155.7.E4 C65 1982. In cooperation with R. Albrecht.

Buchberger:1983:CAS

[4371] B. (Bruno) Buchberger, G. E.

(George Edwin) Collins, and R. (Rudiger) Loos, editors. Computer algebra: Symbolic and algebraic computation, volume 4 of Computing. Supplementum. Springer, Wien / New York, 1983. CODEN COSPDM. ISBN 0-387-81776-X. ISSN 0344-8029. LCCN QA155.7.E4 C65 1983. In cooperation with R. Albrecht.

Bohmer:1984:DCM

[4372] K. (Klaus) Böhmer and H. J. (Hans J.) Stetter, editors. Defect correction methods: Theory and applications, volume 5 of Computing. Supplementum. Springer, Wien / New York, 1984. CO-DEN COSPDM. ISBN 0-387-81832-4. ISSN 0344-8029. LCCN QA297.5 .D44 1984.

Kulisch:1988:SCA

[4373] Ulrich Kulisch and Hans J. Stetter, editors. Scientific computation with automatic result verification, volume 6 of Computing. Supplementum. Springer, Wien / New York, 1988. CODEN COSPDM. ISBN 0-387-82063-9. ISSN 0344-8029. LCCN QA297 .S392 1988. Based on papers presented at the conference on computer arithmetic and scientific computation held Sep. 30–Oct. 2, 1987 in Karlsruhe, FRG.

Tinhofer:1990:CGT

[4374] Gottfried Tinhofer, R. (Rudolf) Albrecht, et al., editors. Computational graph theory, volume 7 of Computing. Supplementum. Springer, Wien / New York, 1990. ISBN 0-387-82177-5 (New York), 3-211-82177-5 (Wien). LCCN QA166 .C618 1990.

Farin:1993:GM

[4375] G. Farin, H. Hagen, H. Noltemeier, and W. Knödel, editors. Geometric modelling, volume 8 of Computing. Supplementum. Springer, Wien / New York, 1993. CODEN COSPDM. ISBN 0-387-82399-9, 3-211-82399-9. ISSN 0344-8029. LCCN QA448.D38 G458 1993. Based on lectures given at the 1st Dagstuhl-seminar, Dagstuhl, Germany.

Albrecht:1993:VNT

[4376] R. Albrecht, G. Alefeld, and H. J. Stetter, editors. Validation numerics: theory and applications, volume 9 of Computing. Supplementum. Springer, Wien / New York, 1993. CODEN COSPDM. ISBN 0-387-82451-0 (New York), 3-211-82451-0 (Vienna). ISSN 0344-8029. LCCN QA297 .V27 1993. Dedicated to Ulrich Kulisch on the occasion of his 60th birthday.

Hagen:1995:GMD

[4377] H. (Hans) Hagen, Gerald E. Farin, Hartmut Noltemeier, and R. Albrecht, editors. Geometric modelling, Dagstuhl, Germany 1993, volume 10 of Computing. Supplementum. Springer, Wien / New York, 1995. CODEN COSPDM. ISBN 3-211-82666-1. ISSN 0344-8029. LCCN QA565 .G45 1995. Based on lectures given at the 2nd Dagstuhl seminar on geometric modelling.

Kropatsch:1996:TFC

[4378] W. (Walter) Kropatsch, R. (Reinhard)

Klette, F. (Franc) Solina, and R. Albrecht, editors. Theoretical foundations of computer vision. Proceedings of the 7th workshop, Dagstuhl, Germany, March 1994, volume 11 of Computing. Supplementum. Springer, Wien / New York, 1996. CODEN COSPDM. ISBN 3-211-82730-7. ISSN 0344-8029. LCCN TA1634 .T43 1996. URL http://www.

                                        springer.at/springer.py?Page=30&
                                        Key=338&id=17&cat=3.

Jolion:1998:GBR

[4379] Jean-Michel Jolion and Walter G. Kropatsch, editors. Graph based representations in pattern recognition. Workshop, GbR ’97, Lyon, France, April 17–18, 1997, volume 12 of Computing. Supplementum. Springer, Wien / New York, 1998. CODEN COSPDM. ISBN 3-211-83121-5. ISSN 0344-8029. LCCN TA1650 .G73 1998. URL

                                        http://www.springer.at/springer.
                                        py?Page=30&Key=339&id=518&cat=3.

Farin:1998:GMD

[4380] Gerald Farin, Hanspeter Bieri, Guido Brunnett, and Tony DeRose, editors. Geometric modelling. 3rd Dagstuhl workshop, Dagstuhl, Germany, 1996, volume 13 of Computing. Supplementum. Springer, Wien / New York, 1998. CODEN COSPDM. ISBN 3-211-83207-6. ISSN 0344-8029. LCCN QA565 .G452 1998. URL http://www.

                                        springer.at/springer.py?Page=30&
                                        Key=340&id=589&cat=3.

Brunnett:2001:GMD

[4381] G. (Guido) Brunnett, H. (Hanspeter) Bieri, and G. (Gerald) Farin, editors. Geometric modelling: Dagstuhl 1999, volume 14. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-211-83603-9. ISSN 0344-8029. LCCN QA565 .G4523 2001. Papers from the 4th Dagstuhl Seminar on Geometric Modelling held May 1999.

Alefeld:2001:TNA

[4382] G. Alefeld and X. (Xiaojun) Chen, editors. Topics in numerical analysis with special emphasis on nonlinear problems, volume 15 of Computing. Supplementum. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-211-83673-X. ISSN 0344-8029. LCCN QA297 .T66 2001.

Herzberger:2003:IMN

[4383] J. Herzberger, editor. Inclusion Methods for Nonlinear Problems: With Applications in Engineering, Economics and Physics, volume 16 of Computing. Supplementum. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 3-211-83852-X. ISSN 0344-8029. LCCN QA427 .I53 2003. EUR$65. Based on papers presented at the International GAMM Workshop held in Munich, Germany, Dec. 15-18, 2000. This workshop was organized with the support of GAMM, the International Association of Applied Mathematics and Mechanics, on the occasion of J. Herzberger’s 60th birthday.

Beliakov:2013:EIBb

[4384] G. Beliakov, D. Creighton, M. Johnstone, and T. Wilkin. Efficient implementation of Bailey and Borwein pseudo-random number generator based on normal numbers. Computer Physics Communications, 184 (8):1999–2004, August 2013. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0010465513001276. See

also [2903].